A Complete Bibliography of Publications in Mathematics and Computers in Simulation: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 1.00

Title word cross-reference

(0, m, s) [392]. (2 + 1) [2174, 1240, 279, 1170, 710, 1309]. (3 + 1) [2116, 2081, 845, 556, 876, 729]. (S − 1, S) [1195]. 19 [2223]. 1 < ρ ≤ 2 [2289]. _{2} [643, 260, 779, 2175, 951]. _{3} [779]. _{6} [643]. A [1860, 497]. α [1225]. β [1715]. C^{0} [1581, 2137]. C^{1} [2137]. C^{2} [769, 2178, 289, 478]. D [103, 435]. F [1342]. H > 1/2 [673]. H_{2} [2185]. H_{∞} [1260, 1301, 692, 35, 853, 999, 1882, 1817, 1219, 2052]. i_{g} [424]. k [698, 1930]. κ [1802]. L1 [483, 719]. L^{1} [1573]. L^{2} [71, 2118]. L_{2} [495]. L_{∞} [495]. M [845, 1170, 750]. S^{2} [153]. H(div) [71]. μ [1000]. N [1813, 72, 929, 750, 1532]. n + 1 [1111]. ω [2059]. p [627, 1376, 1327, 466]. P 1 [1097]. ψ [583, 1589, 1507, 818, 1653]. q [1691, 2093, 435]. S [1117, 1625, 1614]. SI_{R}I_{C}V [1451]. T [1117]. θ [2112, 298, 895]. v_{gs} [424]. y′′ = f(x, y) [1574]. z^{p} + log c^{t} [1625, 1282]. |K|(p, p) [1846].

-breather [750]. -Caputo [583, 1507, 818, 1653]. -circulant [2059]. -convexity [1625, 1614]. -covering [1715]. -dimensional [2081, 845, 824, 2174, 879, 1240, 279, 1170, 1532, 556, 2116, 710, 876, 1309, 729, 72]. -Galerkin [483]. -induced [2223]. -Laplacian [627, 466]. -logistic [1802]. -lump [845, 1170]. -method [298, 895]. -nets [392]. -operators [540]. -Padovan-like [698]. -PID [1260]. -policy [1813, 929, 1342]. -Riemann [1589]. -robust [1225]. -rung [1691, 2093]. -soliton [750]. -stabilization [1000]. -stable [1860].

-th [1376]. -weighted [1930].

/H [779]. /passivity [692].

1 [929]. 15MW [1960]. 19 [653, 834, 1013, 912, 1941, 856, 727, 2061, 744, 1159, 2230, 1229, 350, 928, 858, 1256, 1510, 857, 1523, 499, 1471, 2064]. 191 [1751]. 197 [1750]. 1D [625, 1285, 1076, 274, 1305, 1894]. 1st [592].

2 [1094, 2083, 1058, 1087, 923, 1368, 1917, 1939]. 2-D [923, 1917, 1939]. 2-LTR [1094]. 2020 [68, 154, 257, 40, 121, 143, 54, 215]. 2021 [378, 639, 316, 296, 459, 341, 571, 528]. 2022 [745, 854, 988, 684, 661, 808, 830, 728, 792, 946, 913, 885]. 2023 [1164, 1298, 1421, 1072, 1022, 1237, 1273, 1133, 1204, 1393, 1360, 1332, 1683]. 2024 [1531, 1851, 1489, 1446, 1607, 1647, 1502, 1572, 1781, 1709]. 2025 [2012, 2149, 2237, 1937, 1891, 2075, 2104, 1959, 2050, 2214, 2186, 2169]. 2026 [2304]. 208 [1752]. 2D [625, 1567, 424, 1558, 1401, 1147, 114, 1981, 58, 349, 2284, 274, 1346, 1407, 1637, 1903, 368, 1187, 2034, 2179, 1155, 488, 1316, 938, 1179]. 2D-CLCM [2284]. 2D-integro-differential [1401]. 2D-wave [1147]. 2D-Zettle [2179]. 2nd [1749, 114, 945]. 2nd-order [114].

3 [79]. 3-D [79]. 3-dimensional [491]. 3-species [577]. 3d [8, 2068, 185, 1774, 1812, 155, 622, 1496, 1641, 2200, 126, 291, 274, 1331, 1407, 206, 742, 1924, 1816, 1954, 488, 938]. 3D-chaotic [291]. 3D-cube [206].

4-leg [5]. 4D [1247]. 4H [426]. 4H-SiC [426].

5D [2257].

70th [610].

80 [348, 389]. 8T [206]. 8T-LE [206].

978-0-3673-7768-7 [592]. 978-0-4291-5545-1 [592]. 978-1-4987-3106-5 [592].

ABC [1906, 858, 2179]. ABC-fractional [858]. Ablowitz [1406]. absence [638]. absolute [1846]. absorption [1685]. abstract [767]. abundant [957]. AC [450]. AC/DC [450]. accelerate [1235]. accelerated [1795, 2014]. Acceleration [76, 1466, 2015, 1432]. accelerator [1768]. account [2121]. accumulation [1284]. Accuracy [583, 1035, 844, 383, 1860, 1324, 1313, 2178, 1191, 264, 896]. accurate [1318, 220, 1289, 1429, 2013, 1561, 2290, 1150]. acid [241]. acoustic [1830, 1805]. acquired [401]. acquisition [28]. acquisitions [1688]. activation [1780, 2039, 461, 819, 843]. activator [550]. Active [2028, 1464, 1772, 1816, 675, 1732]. activities [1230]. activity [2067, 1906, 2151, 2135, 2092]. Actor [1966]. actuator [94, 1031, 1793, 872, 1882, 1308]. actuators [223]. Adams [1361, 336]. Adams-type [1361]. adaptation [124]. Adapted [2252, 1842]. Adaptive [2303, 1932, 1966, 2055, 1203, 1300, 1450, 310, 361, 1308, 1362, 1635, 1494, 2051, 1209, 1073, 1913, 127, 663, 2030, 94, 76, 1593, 2247, 1995, 1915, 578, 549, 732, 1095, 2082, 159, 1415, 1963, 1160, 1988, 1911, 853, 758, 1793, 1651, 1320, 675, 955, 386, 122, 1248, 1580, 2231, 1560, 1461, 1659, 2097, 820, 1027]. adaptivity [1648]. AddACO [1549]. additional [2176, 1667, 1949, 225, 1976]. additive [109, 673]. ADE [515]. ADE-FDTD [515]. adherence [1420]. adhesion [97]. ADI [114, 1252, 765, 1142]. adiabatic [1846]. adjoint [886]. adjustment [2003, 1464]. ADM [505]. admission [1292, 1342]. Adomian [532, 100]. Advance [269]. Advanced [1512, 2302, 1491, 26]. advection [1238, 129, 1554, 343, 1307, 1355, 1149, 993, 1881, 2013, 1410, 1081, 209, 1399, 2120]. advection-dominated [1881]. advertisement [1231, 2198]. advertisements [1575]. Aedes [967]. Aegypti [967]. aerodynamic [2267]. aerosol [1074]. affected [355, 2005, 2206, 1621]. affine [2052, 1870]. AFLC [25]. after [1816]. after-balancing [1816]. Ag [643, 951]. against [1067, 1530, 1651]. age [1993, 1066, 1077, 1184, 1831, 382, 866, 1796, 2029, 476, 1519, 649, 1980, 1420, 1425, 1638]. age-dependent [1519]. age-difference [476]. age-since [1420]. age-space [1638]. age-structured [1993, 382, 866, 1796, 2029, 649, 1980, 1425]. agent [902, 107, 781, 1788, 1913, 667, 1021, 2256, 1793, 1104, 122, 2187, 1635]. agent-based [902]. agents [1800]. agglomeration [686]. aggregated [454]. aggregation [644, 131]. agile [2153]. Aging [2155]. agricultural [1552]. agriculture [458]. ahead [1767]. aided [2128]. AIDS [1883, 2151, 1664, 1302, 603, 308, 649, 2154, 1420]. air [2158, 1850, 1621]. aircraft [183]. airway [250]. airway/lung [250]. Airy [210]. Aitken [1974]. Al [779]. Algae [1034]. Algebraic [1558, 1458, 1135, 1874, 404, 2195, 996]. Algorithm [2028, 1437, 1242, 1345, 85, 1335, 216, 1544, 775, 749, 1080, 768, 627, 2285, 1657, 1206, 1813, 931, 1073, 1966, 1118, 1942, 2156, 1518, 1998, 963, 695, 256, 764, 2128, 155, 2024, 689, 1590, 130, 2252, 840, 935, 1197, 679, 1038, 1218, 635, 1023, 9, 381, 732, 1148, 1095, 1631, 2219, 1005, 897, 159, 947, 102, 1734, 184, 1103, 1674, 1459, 1519, 712, 1369, 867, 1312, 1006, 1152, 333, 21, 361, 803, 183, 1549, 307, 1106, 1075, 738, 1225, 1011, 994, 1320, 1181, 1064, 2179, 675, 574, 887, 1088, 930, 264, 1275, 1416, 1652, 218, 460, 353, 1276, 1636, 499, 1364, 2228, 1266, 2221, 2120]. algorithm [249, 1212, 1063, 1599, 1974, 2094, 2119, 2060, 2245, 503, 782, 689, 1610, 579, 1412, 991]. algorithm-II [991]. Algorithmic [2022, 1484]. algorithms [902, 1771, 553, 1930, 18, 2006, 600, 197, 1470, 1171, 1815, 724, 471, 72, 616, 1496, 1168, 860, 238, 1144, 2208, 1387, 1190, 2198, 1293, 564, 1762, 282, 834]. Allee [570, 721, 965, 1951, 1919, 2196, 2001, 1892, 914, 1178, 1115, 1615, 2293]. Allen [1226, 1463, 1208, 1853, 990]. Allmaras [1675]. allocation [559, 1518, 1173, 300]. almost [1341, 502]. alone [18]. along [79, 877]. alternating [1403, 1977]. alternative [705, 1172, 654, 2253]. alternatives [893]. Altruistic [1312]. alumina [366, 733]. Alzheimer [2260]. ambit [1472]. ameliorated [1181]. American [615, 1893, 248, 1103, 2043]. American-style [2043]. ammonia [787]. among [1314, 1240]. Amoroso [678]. Amplitude [530, 611]. anaerobic [468]. analogue [1431]. analyses [836, 1289, 690]. analysing [1058]. Analysis [429, 1958, 1987, 1538, 1426, 2151, 1748, 47, 638, 903, 1047, 401, 937, 1379, 356, 308, 794, 763, 862, 1976, 2080, 1738, 1114, 1993, 116, 709, 1045, 775, 485, 1365, 1771, 182, 1584, 560, 1972, 1878, 123, 968, 498, 839, 922, 220, 1646, 912, 1932, 1117, 18, 2217, 1409, 96, 971, 1336, 284, 1650, 1663, 223, 1367, 1745, 1975, 1581, 2163, 197, 931, 1731, 1831, 1054, 647, 2033, 1436, 1721, 145, 2270, 2113, 2076, 1554, 823, 1899, 541, 2240, 783, 1395, 1264, 1752, 674, 323, 748, 56, 2205, 290, 2144, 1451, 1226, 1537, 1883, 1723, 1876, 1146, 704, 108, 1912, 1211]. analysis [1247, 966, 496, 332, 2294, 1551, 489, 334, 1497, 1159, 734, 1261, 1397, 1970, 655, 779, 304, 1472, 644, 590, 1096, 1877, 1796, 2011, 2152, 2029, 2165, 2181, 279, 162, 819, 1277, 2018, 2158, 2291, 629, 1158, 562, 393, 126, 1951, 1648, 1149, 929, 599, 1854, 1728, 1622, 1526, 2202, 1049, 78, 1563, 692, 137, 87, 1136, 434, 1119, 565, 1142, 2201, 477, 1703, 1968, 1892, 604, 105, 1829, 770, 2054, 601, 1637, 287, 710, 1903, 858, 425, 1491, 1548, 1511, 539, 738, 921, 1317, 1422, 540, 1430, 2239, 939, 1255, 1059, 2160, 165, 420, 1420, 1835, 435, 493, 1236, 1685, 24, 1364, 1499, 370]. analysis [2231, 1571, 795, 1362, 1244, 1471, 1844, 1638, 2025, 1704, 2283, 2220, 1885, 864, 1686, 1628, 821, 466, 2105, 987, 361]. Analytic [323, 2041, 348, 1, 389, 1565]. Analytical [318, 1977, 86, 441, 1967, 1208, 1878, 437, 1624, 426, 823, 1845, 1595, 1854, 619, 2301]. Analyze [7, 2243]. Analyzing [643, 2296, 1875, 2005]. ANFIS [709, 10, 1334, 1551]. angiogenesis [1010]. angle [116, 444]. angles [1287]. anhysteretic [1657]. Anisotropic [362, 1968, 77, 1490, 1051, 952, 1307, 1696, 1641, 219, 1093, 1353]. anisotropy [597]. ANN [17]. annuities [2193]. annuity [700]. annular [1168]. annulus [753]. ANOVA [784]. Ant [963, 1549]. anthropogenic [1722]. Anti [1275, 922, 580, 1219, 1898, 1817]. Anti-diagonalization [1275]. anti-disturbance [1898]. anti-predator [580]. anti-retroviral [922]. anti-synchronization [1219, 1817]. antibody [704, 855]. antigen [855]. antiretroviral [1514]. antithetic [340]. any [202]. aperiodically [1593]. Aphid [963]. appearing [1041]. appendages [2007]. Application [2262, 908, 967, 880, 898, 1574, 45, 355, 1293, 758, 2244, 228, 1544, 1469, 769, 972, 575, 1733, 2091, 973, 1941, 600, 1354, 1198, 1189, 1224, 2172, 1284, 1211, 1038, 2294, 652, 238, 9, 2243, 2200, 2288, 124, 163, 1415, 1217, 2158, 13, 1826, 310, 1609, 5, 1772, 803, 1869, 2256, 1186, 287, 1156, 494, 1901, 178, 1825, 2045, 676, 545, 1523, 2016, 1183, 864, 1583, 1894, 77, 79, 1875, 718, 1384, 701]. Applications [2, 979, 1725, 772, 514, 1606, 786, 1717, 1878, 1715, 998, 419, 1897, 510, 1780, 1587, 1991, 1764, 1228, 1060, 1556, 1076, 654, 1765, 2202, 934, 1653, 2057, 517, 1967, 1804, 1741, 723, 996, 1931, 1011, 430, 1732, 1375, 990, 2295, 976, 1042, 1289, 1980]. Applied [609, 189, 31, 1770, 184, 83, 206, 1610, 438, 1509, 209, 1088, 446]. Applying [392]. Approach [1684, 1562, 977, 1033, 247, 1014, 861, 998, 825, 769, 302, 1026, 1401, 195, 1385, 1138, 1536, 55, 2177, 1815, 1908, 199, 731, 1945, 967, 1899, 778, 1998, 1594, 471, 300, 256, 201, 2126, 1280, 557, 1537, 660, 252, 614, 1723, 63, 697, 1565, 2002, 16, 47, 1691, 1989, 1028, 162, 1217, 1782, 2158, 1665, 1488, 869, 812, 1474, 1475, 322, 762, 1672, 604, 1546, 1609, 2093, 288, 1384, 88, 564, 1125, 700, 1986, 1297, 287, 1792, 2236, 1509, 1116, 1101, 180, 1834, 669, 460, 2188, 2293, 272, 1938]. approaches [1725, 961, 832, 894, 1826, 1854, 2135]. approximant [478]. Approximate [551, 1222, 1737, 323, 1246, 1098, 181]. approximating [958, 987]. Approximation [510, 1268, 2057, 1993, 133, 267, 1180, 1033, 558, 1344, 1484, 1452, 220, 468, 1926, 71, 624, 2118, 1224, 1441, 382, 1496, 1716, 527, 1918, 301, 2166, 1411, 1424, 1366, 479, 1665, 1791, 519, 1024, 1093, 1724, 131, 368, 1156, 1909, 811, 1215, 1964, 386, 406, 1245, 756, 798, 2043]. approximations [1712, 730, 599, 587]. April [68, 378, 745, 1164, 1531, 2012]. aquatic [644, 1865]. Aquila [1181]. Arabia [182]. Arbitrage [1773]. Arbitrarily [1495]. arbitrary [776, 1093, 478, 1275, 533]. arbitrary-order [533]. arch [1548]. Arched [164]. Archimedes [1038]. architecture [642]. area [1823]. areas [1980, 2005]. argon [1398]. argument [1306]. arid [12]. ARIMA [2044]. arisen [504]. arising [1616, 1503, 543, 1633, 2173]. arithmetic [418]. arm [2128]. Arnold [2039]. Arnoldi [656]. array [10]. Arrhenius [1129, 932]. arrhythmia [736]. ART [1664]. arterial [1503]. arteries [1067]. artery [2076, 1466, 60, 61]. Artificial [1175, 694, 453, 1123, 1004, 1970, 2168, 1186, 497, 33]. Asian [340, 1859]. Aspects [800, 379, 1858]. Assam [225]. assembly [1340]. assess [1138, 1010]. Assessing [1365]. Assessment [1182, 1385, 1746, 2054, 2206, 178]. asset [114, 1713]. assets [1223]. assignment [2157]. assisted [1905, 361]. associate [403]. associated [698, 1431]. associative [785]. assumption [177]. asymmetric [874, 1396]. asymptomatic [2049, 939, 1428]. Asymptotic [1066, 2111, 1521, 1376, 1596, 1835, 1646, 1871, 2166, 1350, 735, 1087, 101]. Asymptotically [103, 1942]. Asymptotics [818, 2173]. asynchronous [1351]. atmospheric [1398, 2296]. atoms [957]. attacks [1365, 699, 2282, 1530, 2052, 1901, 1651, 1601, 2298]. attempt [1334]. attraction [2141, 2219]. attractor [1331]. attractors [919]. attribute [85]. Audit [1686]. augmentation [1834]. Augmented [1284]. August [154, 854, 1298, 2149]. auto [903]. auto-convolution [903]. Automatic [278, 194, 2258, 1823]. automaton [45]. automotive [1772, 676]. Autonomous [1518, 1040, 83, 353, 1276, 1012]. Autoregressive [395, 297, 2225, 1043, 2221]. autotoxicity [1929]. AUV [1694]. auxiliary [1163, 697, 1649, 371, 1636]. average [1778, 1885]. averaged [2175]. averse [62, 397, 1008]. aversion [1679]. aware [2291, 184]. awareness [1105]. AWENO [1660]. axial [1774]. axially [889]. axisymmetric [243, 79, 2227]. Ayala [604]. azimuthal [1287].

B [1724, 838, 2262, 278, 1787, 1620, 730, 1146, 156, 527, 1023, 1612, 1874, 948, 1411, 1576, 519, 289, 797, 905, 593, 1371, 266].

B-spline [1724, 278, 1620, 156, 527, 1874, 1411, 1576, 519, 797, 593]. B-spline-like [1787]. B-splines [838, 2262, 1023, 1612, 289]. BA [2160]. back [438]. back-to-back [438]. background [1404, 831, 41]. backgrounds [1956]. backlogged [2198]. Bäcklund [2081]. Backpropagation [1610]. backstepping [1560, 820]. Backward [922, 1839, 125, 1662, 2103, 1803, 1902, 1245]. baculovirus [1552]. Badger [689]. badly [1483]. badly-shaped [1483]. Bakhvalov [1939]. Bakhvalov-type [1939]. balance [131]. Balanced [1462, 71, 1660, 1517, 1879, 387, 1629, 2037]. Balancing [1908, 1654, 415, 1816]. balking [2080]. Ball [840, 1197]. BAM [162, 670, 1255, 403, 1000]. Banach [1470]. banded [751]. bandgap [424]. Bangladesh [734, 2049]. barnacle [1544]. Barndorff [1542]. barrier [1265, 1398, 426, 1144]. barriers [1144]. bars [19, 628]. Bartlett [181]. Barycentric [1148, 888, 1253]. base [643, 1409]. based [85, 838, 977, 1547, 1654, 1860, 775, 902, 241, 782, 1771, 2051, 1775, 429, 1717, 1972, 968, 2226, 484, 998, 23, 1821, 1086, 7, 890, 1567, 132, 1932, 1761, 1503, 470, 125, 1789, 15, 1413, 194, 1620, 243, 84, 2280, 1905, 1458, 521, 2177, 1073, 781, 453, 1067, 550, 231, 306, 1998, 2180, 42, 1594, 320, 888, 847, 471, 1272, 1395, 695, 1447, 97, 56, 94, 2157, 451, 76, 2126, 28, 469, 2024, 1590, 1778, 1031, 1912, 1716, 164, 966, 364, 666, 1841, 595, 1038, 520, 2039, 860, 915, 1995, 836, 1023, 421, 2130, 80, 304, 784]. based [2232, 732, 960, 1095, 1036, 1690, 2029, 1973, 1782, 2238, 1963, 2284, 1827, 629, 1408, 1516, 291, 2168, 1149, 1525, 2278, 1642, 1301, 22, 25, 204, 2246, 1462, 173, 1684, 176, 2198, 1768, 621, 1327, 379, 654, 1449, 2027, 78, 2139, 1911, 352, 1369, 692, 418, 762, 1380, 759, 458, 1113, 1619, 1858, 718, 649, 2093, 1384, 2040, 443, 1816, 454, 35, 1741, 1186, 2052, 2014, 287, 283, 797, 1061, 884, 1187, 1165, 1139, 351, 112, 1317, 747, 1320, 180, 760, 999, 1104, 1464, 1979, 2160, 587, 930, 317, 717, 1592, 1652, 1904, 648, 1842, 353, 1276, 1694, 1399, 1738, 1685, 2228, 1882]. based [1814, 2197, 2231, 980, 249, 275, 1063, 1599, 1974, 2094, 2119, 1950, 282, 1688, 843, 1220, 1686, 2026, 821, 2171, 732]. bases [194, 1746]. basin [86]. Basis [166, 514, 518, 243, 244, 1354, 135, 527, 860, 149, 329, 654, 906, 1996, 1609, 368]. batch [2247]. bathtub [1405]. bathtub-shaped [1405]. batteries [419, 413, 415]. battery [1383, 10, 414, 416, 417]. Bautin [817]. Bayesian [181, 997, 184, 1257, 1986, 529]. Bazykin [1988]. BBM [2190, 1152]. BBM-KdV [2190]. BCC [515]. BCC-Grid [515]. BCM [642]. BDF [2220]. BDF-spectral [2220]. BDF2 [1154, 2276, 2292, 1372]. BDF2-FEM [2276]. bead [1641]. beam [560, 612, 832, 931, 480, 1366, 482, 1548, 613, 165, 1114]. beam-arch [1548]. beams [889, 224, 75, 2251]. Becker [592]. Beddington [833]. beds [1997, 2293]. bee [1186]. before [1013]. Behavior [1624, 1855, 580, 2020, 2144, 708, 2215, 1808, 655, 222, 1359, 1348, 1143, 1526, 87, 603, 21, 1822, 2293, 1961]. Behavioral [427]. behaviors [1239, 957, 134, 818, 960, 813]. behaviour [1582, 1492, 555]. behaviours [309]. Bell [299]. BEM [2201]. Benchmarking [658]. Bending [477]. benefits [894, 2193]. Benjamin [1474, 729]. bent [1126]. Bernoulli [832, 1813, 931, 320, 1366, 762, 1078, 2251, 1056, 2066, 2080]. Bernstein [1811, 1717, 518, 1458, 931, 892, 552, 1384]. Bertrand [2131]. Bessel [1550, 2203, 1268, 2265, 814, 1944]. Best [540, 191, 1609]. between [182, 424, 193, 1720, 2035, 1007, 476, 1291, 1773, 235, 1713, 586]. Beverton [1515]. beyond [2111]. Bezier [552, 786, 1811, 1717, 892, 663, 479]. BFGS [2185]. Bi [628, 450, 1080, 2021, 1173, 764, 1587, 1794, 2000, 2204, 1870]. Bi-affine [1870]. bi-directional [450, 764]. bi-fuzzy [1587]. bi-Laplacian [2021]. bi-level [2000]. bi-objective [1080, 1173, 1794]. Bi-penalty [628]. bi-virus [2204]. bias [95, 475]. bias-corrected [475]. bicycle [245]. bicycle-sharing [245]. bidimensional [2037]. bidirectional [785, 1772, 403]. bifurcated [2076]. Bifurcation [498, 1745, 1723, 262, 644, 1158, 1951, 1141, 1563, 1571, 1244, 1365, 922, 962, 1708, 2031, 2061, 2257, 2270, 1151, 138, 2144, 1451, 1803, 1839, 1023, 1397, 2087, 2165, 1157, 2199, 1622, 1049, 1892, 1349, 2218, 766, 315, 545, 386, 403, 953, 817, 370, 795, 1661, 1916]. Bifurcations [868, 657, 783, 829, 2250, 2106, 690, 802, 1115, 547, 1798]. Bifuzzy [1138]. biharmonic [1029, 1968]. bilateral [1841, 1966]. Bilinear [2081, 879]. bin [1942]. binary [1344, 299, 2228]. binding [1291]. Binomial [869, 1782]. bio [948, 1970, 1034, 1549]. bio-fertilizer [1034]. bio-heat [948]. bio-inspired [1549]. bio-mathematical [1970]. biocell [221]. biochemical [179]. Bioconvection [761, 932, 266]. Bioconvective [1129]. biodegradation [468, 178]. biodigester [1279]. biofilms [2266, 1927]. biological [2134, 1591, 137, 1427]. biology [1989]. Biomagnetic [908]. biomass [1929]. biomechanical [185]. biomedical [61]. bioreactor [2070]. biosorption [2266]. Biot [238]. bipartite [1017, 589]. bird’s [1455]. bird’s-eye [1455]. Birefringence [1186]. birth [1945]. birthday [610]. Bismut [1484]. Bistability [1949]. Biswas [356]. bit [2139, 503]. bit-level [503]. bitemporal [2005]. bitplane [1075]. bits [1583]. bivariate [1897, 654, 478, 1382, 1984]. BKK [824]. BKP [1240]. black [935, 1960, 1623, 114, 2274, 409]. black-start [1960]. blade [278]. Blaheta [610]. Blasius [532, 1465]. blast [1415]. blending [191]. Bloch [1189]. block [2258, 635, 1405, 762, 1619, 1727, 251, 472]. block-pulse [762]. blood [1333, 594, 1503, 1067, 741, 590, 2168, 1466, 758, 61, 346]. blow [2277, 1161]. blow-up [2277, 1161]. blowflies [108, 1704]. blowup [153]. BNN [1610]. Board [36, 37, 38, 52, 66, 91, 119, 141, 152, 171, 188, 213, 255, 294, 314, 339, 376, 411, 433, 457, 509, 526, 568, 569, 608, 634, 636, 659, 682, 715, 743, 790, 806, 828, 852, 883, 910, 911, 944, 986, 1020, 1070, 1110, 1162, 1202, 1234, 1271, 1294, 1296, 1330, 1358, 1391, 1419, 1444, 1480, 1487, 1500, 1529, 1570, 1605, 1645, 1682, 1707, 1755, 1756, 1779, 1849, 1889, 1936, 1957, 2010, 2048, 2073, 2102, 2147, 2167, 2184, 2212, 2235]. Boca [592]. body [1304, 874, 1466]. Bogdanov [829, 2250]. Bogoyavlenshii [1170]. Boltzmann [859, 1009, 1398, 311, 1439, 1067, 58, 1875, 142, 2140, 1608, 352, 217, 355, 2227]. Bona [1474, 729]. bond [1516]. bond-based [1516]. bonded [618]. bone [641]. bonobo [1518]. Book [592]. Boolean [214]. boom [2142]. Boomerang [2267]. boost [11, 5, 642]. boosted [935]. boosting [1563, 1840]. border [1256]. borne [2236]. Bose [1981, 2173]. bosonic [957]. both [922, 1662, 1831, 2121, 1492, 914, 235]. bound [1080, 511, 990, 1870, 1519]. boundaries [55]. Boundary [777, 2007, 1079, 1045, 2088, 390, 626, 625, 627, 796, 318, 160, 2272, 1313, 1401, 1147, 1122, 1650, 84, 1520, 1581, 1123, 236, 480, 1589, 2099, 660, 622, 1655, 1496, 650, 1168, 1465, 917, 371, 47, 992, 467, 1263, 2053, 224, 163, 561, 482, 735, 1326, 2085, 562, 335, 305, 515, 813, 400, 1407, 535, 563, 767, 1297, 1166, 283, 800, 1491, 1676, 891, 1317, 2239, 1420, 1245, 899, 1938]. boundary-domain [335]. boundary-layer [1465, 47]. Bounded [1301, 1260, 1050]. boundedness [904, 1327, 2295]. bounds [135, 1373, 2115].

Boussinesq [879, 1826, 264, 1192]. box [2129, 993, 1220, 2026]. Box-Jenkins [2026]. boxes [1280, 269, 2179]. brain [238]. branch [1080, 1519]. Bratu [998, 472, 915]. Bratu-type [998]. break [2198]. breakage [686, 2014]. breakdown [1834]. breakwaters [1973]. breast [1004, 81, 1102, 2188]. Breather [1309, 1670, 750, 2173]. Breather-to-soliton [1309]. Breathers [831, 299, 556, 729]. breathing [250]. breeding [1185]. Bridge [2028]. bridges [438]. brief [366]. Bringing [185]. Brinkman [380, 1558, 71]. Broer [2286, 824]. broken [19]. Brownian [1452, 708, 917, 671, 1859, 901]. brucellosis [2163, 2206]. Brusselator [2215, 74, 1049]. BSDE [1920]. BSLM [1658]. bubble [76, 829]. bubble-type [76]. buckling [1537]. budworm [706]. building [470, 2126]. buildup [180]. bulk [1931]. bulk-surface [1931]. Bulkley [2140]. bunch [2019]. bundle [717]. Buongiorno [643]. BURA [616]. Burger [1620]. Burgers’ [1658, 1267, 488, 838, 1285, 1899, 892, 860, 1251, 1289, 548, 813, 1474, 1792, 2290, 815]. buried [63]. bus [448, 1765]. bust [2142]. butterfly [1095, 1088, 2094]. buyer [1008]. BVP [1589, 1633]. BVPs [1895, 521, 945, 1168, 934, 1727].

C [1751, 643, 1900, 672, 2119]. C-means [2119]. cable [1553, 1339, 1885]. cable-driven [1553]. cables [1553]. Caged [2032]. Cahn [1452, 1226, 1795, 1463, 1808, 319, 884, 1208, 1853, 1954, 990, 2287, 1718, 2220]. calculation [496, 1459, 428, 649, 1177, 1561, 2259]. calculations [613]. calculus [793, 1620, 1825]. Calendar [51, 65, 90, 118, 140, 151, 187, 212, 254, 293, 313, 338, 375, 432, 508, 525, 607, 633, 637, 681, 714, 726, 789, 805, 827, 851, 882, 943, 985, 1019, 1069, 1109, 1130, 1131, 1201, 1233, 1270, 1295, 1329, 1357, 1389, 1418, 1443, 1479, 1486, 1528, 1569, 1604, 1644, 1681, 1706, 1754, 1777, 1848, 1888, 1922, 1935, 2009, 2047, 2072, 2101, 2133, 2146, 2183, 2211, 2234, 410, 2268, 170, 456]. calibration [242, 1103, 1617]. call [2207]. Calogero [1170]. Can [1988, 1185]. Canard [1916]. cancer [1925, 773, 1004, 2004, 1955, 1190, 1102, 2188]. cannibalism [2086]. cannibalistic [2255]. canopy [176]. cantilever [480, 812, 165]. capabilities [1900]. capacitance [1177]. capacity [559, 841, 802]. capital [791, 2152, 2124, 1435]. capital-induced [2152]. capsid [1030]. capsule [582]. Caputo [462, 1016, 583, 2105, 967, 888, 1807, 1507, 542, 381, 638, 57, 818, 1867, 2018, 1408, 1653, 265, 767, 1311, 1221, 596, 1630, 1156, 1255, 587, 2213, 822, 1362, 1062]. Caputo-Fabrizio [638]. Caputo-space [1807]. Caputo-type [57, 1408, 596]. Carbon [1782, 2122, 1231, 1722, 2135]. cardiac [2275]. cardinal [1370]. cardinalities [1942]. Carlo [1542, 1394, 145, 1910, 898, 1453, 340, 73, 2044, 392]. Carreau [79, 1263, 800]. carry [1892]. carry-over [1892]. carrying [802, 803]. cart [1308]. cascades [278]. case [1567, 550, 744, 1197, 877, 741, 638, 1414, 1457, 361, 2049, 2263, 403, 920]. cases [1176, 857]. Casson [761, 227]. casting [59]. cat [1911]. catalysed [395]. Cauchy [541, 1089]. caused [1906, 245]. causes [763]. cavitating [146]. cavities [1866]. cavity [677, 2140, 686, 539]. CD4 [1283, 1803, 584]. Cell [736, 7, 922, 15, 6, 1520, 325, 2240, 97, 2155, 418, 603, 5, 436, 34, 112]. Cell-to-cell [736, 922]. cell/supercapacitor [436]. cells [465, 1283, 1925, 1054, 1803, 638, 866, 2029, 1190, 584]. cellular [485, 748, 103, 740, 642, 1698, 45, 1065, 1499]. cement [675]. censored [763]. censoring [1405, 688]. Center [1477]. centered [2059]. centers [1477]. centres [371]. certain [1477, 1060]. cervical [1955]. CEV [1265]. CFD [2176, 1067]. CGSS [600]. chain [559, 1091, 62, 791, 1679, 270, 657, 577, 1504, 693, 1677, 1105, 1951, 1143, 690, 1865, 1819, 1243, 691, 99, 907, 397, 1810]. chains [545]. Challenges [4]. chameleon [1197]. Change [1368, 1542, 55, 3, 2005]. changed [1435]. changing [749]. channel [1303, 559, 1091, 426, 1105, 1290, 473, 1634, 905, 99, 907, 346, 2259]. Chaos [766, 683, 1013, 1584, 2131, 223, 1745, 580, 138, 1906, 1723, 1828, 2199, 1331, 128, 350, 994, 386, 2094, 217]. Chaotic [712, 241, 919, 1862, 2091, 285, 1092, 781, 231, 138, 469, 2281, 2247, 732, 1076, 2284, 291, 1190, 878, 2034, 351, 1837, 2179, 701, 1842, 1025, 864, 1220, 1583, 1894]. Chaplygin [2015]. character [1445]. characteristic [1477, 1975, 1307, 863, 2015, 774]. Characteristics [671, 1333, 1441, 2199, 1002, 2079]. Characterization [964, 539]. characterizations [1375]. characterized [1949]. charge [1127, 1476]. charge-conservative [1476]. charged [1304]. charging [1763, 449]. Chebyshev [998, 1307, 835, 2189, 2277, 711, 158, 1179, 1501]. chef [1369]. chef-based [1369]. chemical [79, 1049, 537]. chemically [1129]. Chemo [1710]. chemostat [1555, 1141, 1810]. chemotactic [1701]. chemotaxis [2135, 259, 2077, 1953, 2170]. chemotherapeutic [638, 1427]. chemotherapy [773]. Chickenpox [1884]. childhood [405]. Chile’ [727]. chili [1221]. chimp [840]. Chimpanzee [949]. China [3, 877, 1003, 2206, 920]. chip [656]. chlorophyll [497]. choice [927]. cholera [1597, 862]. Chun [345]. CIR [1995, 387, 226]. circle [1799]. circuit [424, 1774, 821]. circuits [442]. circulant [2059]. circular [1469, 327]. circulation [326, 993]. city [2107]. civil [2126]. clamped [477]. class [1440, 1821, 1016, 995, 1585, 1092, 1586, 1913, 2225, 1447, 469, 2281, 1999, 130, 915, 551, 927, 1915, 1060, 903, 2278, 1348, 1352, 1491, 878, 495, 417, 1633, 488, 822, 1543, 1938, 1284]. classical [14, 233, 1829]. Classification [1684, 2126, 2188, 1540]. classified [1743]. CLCM [2284]. Clenshaw [1235]. Clifford [969]. Clifford-valued [969]. climate [1806, 3]. climate-related [1806]. climatic [12]. clogging [245]. close [1324]. Closed [475, 677, 86, 1595, 930, 2290]. Closed-form [475, 1595, 2290]. closed-loop [930]. clothoids [193]. cloud [1813]. clouds [1924]. Clough [2137]. cluster [1076, 1677, 122]. clustering [1690, 1911, 1741, 2119]. clusterization [192]. CNN [2019, 1816]. CNN-based [1816]. CNTs [959]. CO [260, 1720, 912, 2083, 1229, 1214, 1274, 803, 430, 2175]. co-dynamics [1229]. Co-infection [1720, 912, 2083, 1274]. co-operation [803]. co-opetition [1214]. co-simulation [430]. Coalitional [1789]. Coarse [1290]. Coarse-grained [1290]. coastal [2122]. CoCoSo [1962]. CODAS [1691]. code [393]. codimension [2250]. coefficient [2180, 535, 2259]. coefficients [536, 1559, 2081, 381, 1990, 310, 2013, 1804, 1052, 1187, 372, 1713, 2245]. coexistence [875, 1988]. coexisting [585]. Cohen [1362]. Coiflet [677, 2290]. coil [446]. coils [427]. coinfection [1745]. cointegrating [310]. cold [957]. collaborating [555]. collapse [251]. collective [1789, 1836, 1908, 309]. collective-motion [1836]. collector [14, 12, 707]. college [317]. collisional [2014]. Collocation [832, 694, 320, 836, 838, 1079, 620, 1454, 160, 284, 2138, 43, 521, 931, 1135, 343, 1252, 1845, 1729, 1505, 1591, 2277, 1023, 2288, 903, 1148, 1872, 2029, 224, 1134, 377, 1728, 207, 1619, 1672, 1384, 771, 1634, 2297, 1278, 1179, 814, 798]. colony [1549, 1186]. Color [2284, 291, 1842]. combating [773]. Combination [773, 2091, 716, 1514, 1253]. Combinatorial [2300, 1840]. combined [764, 917, 1474, 641, 2117, 994, 1088, 843]. combining [173]. commensurate [2199]. Comment [348, 389]. Comments [897]. commercial [2151]. commodity [1457]. common [1112, 1556, 1139]. communication [94, 950]. communications [973, 2099]. community [1789, 1256]. commuter [326]. Comoving [1676]. Compact [264, 1403, 1035, 1386, 1899, 1306, 1041, 1610, 1267, 2190, 1152, 765, 1142, 372, 488, 1316, 815, 2245]. Comparative [17, 22, 968, 2242, 929, 1854, 366, 564, 1927, 821]. Comparison [450, 1771, 424, 14, 1975, 1354, 1007, 452, 653, 278, 448, 1987, 1336, 1598, 1845, 1191, 2197, 1524, 1774]. compartment [317]. compartmental [1941, 2031, 1426]. Compatibility [1011, 32]. Competing [555, 763, 688]. Competition [2024, 875, 668, 2020, 280, 230, 87, 604, 755]. Competition-based [2024]. competitive [875, 1702, 1521]. complementarity [1158, 204]. complete [523, 396]. Complex [1997, 2200, 1667, 1736, 1099, 1533, 501, 109, 991, 1067, 2141, 164, 966, 1852, 549, 1089, 1095, 1017, 142, 137, 1219, 2022, 88, 884, 101, 1338, 1181, 271, 572, 493, 669, 2295, 589, 1212, 1866]. complex-valued [109, 966, 1219, 101, 1212, 1866]. complexities [1919]. Complexity [1677, 365, 791, 833, 363, 225, 1099, 315, 1798, 2086]. complications [941]. component [705, 721, 1049, 1257]. component-by-component [705]. components [2226, 1597, 1760]. composed [408]. Composite [998, 2221, 367, 224, 1466, 21, 864]. Compositions [972]. composting [221]. comprehensive [402]. compressibility [333]. compressible [1539, 1056, 112]. compression [269, 1290]. compressive [2228]. Compromising [579]. Comput [348, 1752, 389, 1751, 1750]. Computation [605, 553, 2134, 2105, 1239, 1481, 1595, 335, 2250, 83, 1466, 1967, 932, 392, 181]. Computational [626, 510, 594, 609, 704, 2002, 590, 2057, 61, 775, 1567, 776, 1945, 823, 660, 232, 542, 1007, 2215, 2171, 1497, 987, 621, 812, 434, 1341, 1166, 383, 1509, 1900, 1427, 412, 1749, 379, 1016]. computationally [55, 2077]. computations [1235]. compute [768, 1554, 1187]. computer [1738]. computers [393]. Computing [1940, 189, 1180, 709, 1987, 1650, 1813, 2203, 364, 1551, 1886, 73, 1152, 954, 286, 506, 2302, 2216, 584, 1371]. concave [2238]. concentration [497]. concept [1821, 1352]. concrete [321, 665, 801, 287]. condensates [1981]. condensation [2173]. condition [625, 129, 515, 1420]. Conditional [2172, 1771, 1107, 1967, 1043]. conditions [1121, 1079, 1045, 1477, 390, 786, 796, 1401, 12, 776, 777, 1123, 1589, 2099, 622, 1655, 1496, 1168, 779, 163, 1532, 535, 767, 876, 980]. conducting [796]. conduction [536, 1761, 1490, 1381]. conductive [28]. conductivity [1567, 237, 1263]. conductors [196, 1726]. conduit [537]. cone [951, 235, 1685]. Conference [510, 609, 1683, 2057, 1749]. configuration [1769, 1765, 445]. confined [236, 800]. conformable [462]. conformable-Caputo [462]. conformal [164]. conforming [1668]. conic [1901]. conic-type [1901]. conjectural [960]. conjugate [977, 979, 557, 1459, 1869, 2299, 1266]. conjugation [1927]. conjunct [1898]. connected [10, 16, 982, 1835, 435]. connection [1958, 1804]. Connections [1114]. connectivity [1618]. consciousness [1185]. consensus [781, 1788, 1913, 667, 1021, 1546, 2256, 1104, 122, 2298]. consequences [1687]. Conservation [342, 398, 1140, 1906, 57, 599, 259, 1399]. Conservative [1192, 1318, 987, 1476, 1853, 1744, 46]. conserving [1495]. considerations [58]. considering [1771, 559, 1073, 1831, 2157, 334, 2082, 1677, 1105, 801, 907, 1244, 1671]. Consistent [1713, 1146]. consisting [2278]. constant [2217, 785, 727, 1994, 960, 1373, 1877, 2278, 688, 1961, 2259]. constant-stress [688]. constituent [2175]. Constrained [2004, 1203, 557, 2024, 1512, 2252, 2032, 1793, 1320, 1636, 1744, 1592]. Constraint [598, 791, 83]. constraints [991, 104, 110, 1450, 1387, 800, 1116, 1870, 1618]. construct [1562, 1239]. constructed [552, 1844, 1704]. Constructing [1107]. Construction [705, 289, 2226, 1354, 478]. consumer [1678, 1396]. consumer-resource [1396]. Consumers [422]. consumption [182, 1775, 1789, 1555, 1767, 1003, 665, 2058, 645]. contact [628, 1622, 2251]. contact-impact [628]. contagion [1319, 647, 157, 1671]. Contagious [2142, 1176]. containing [1973, 2015]. contaminant [178]. content [1415]. context [515, 1343, 873, 2058]. context-customized [873]. continuity [786]. Continuous [1185, 1546, 782, 1761, 62, 1928, 1714, 1306, 1808, 1242, 947, 2032, 955, 2070, 1364, 1916, 261]. Continuous-time [1185, 1546, 62, 2032, 955]. continuously [671, 144]. continuum [653, 205]. contract [1345]. contraction [1303, 532]. contractions [1693]. contractive [1470]. contracts [1806]. contrast [545]. contribution [302, 1545]. Control [1462, 1960, 975, 1402, 977, 2110, 1553, 1632, 2088, 2051, 626, 23, 1016, 912, 1626, 1958, 26, 1117, 1413, 696, 1836, 1147, 973, 1319, 1184, 1662, 2103, 1720, 1581, 2163, 1763, 1966, 773, 647, 781, 2113, 699, 1151, 231, 306, 1884, 1788, 1586, 1790, 138, 1913, 1906, 1260, 1272, 72, 1203, 94, 1153, 2157, 1593, 229, 2281, 1883, 614, 2254, 130, 1876, 936, 1778, 1031, 1839, 1920, 449, 1323, 1841, 104, 16, 1038, 281, 2123, 2242, 1397, 1575, 655, 578, 1286, 667, 2004, 222, 1096, 1036, 1230, 1053, 2082, 1359, 1124, 1450, 1963, 950, 102, 554, 1302, 2168, 221, 2278, 1823, 1301, 423]. control [22, 504, 1552, 1190, 1327, 989, 1331, 1449, 1656, 2027, 1698, 2095, 4, 369, 1379, 762, 1861, 1292, 718, 20, 1773, 2282, 5, 794, 1274, 2256, 1342, 35, 1611, 1530, 766, 144, 853, 2206, 1898, 878, 115, 1338, 758, 1901, 351, 34, 1923, 703, 1651, 675, 1732, 1351, 872, 93, 999, 1464, 1820, 2070, 2160, 2056, 930, 98, 572, 1416, 1601, 1904, 1917, 1882, 980, 412, 533, 1308, 1560, 846, 1635, 1471, 1659, 474, 848, 1000, 1938, 2260, 1621, 820, 1943, 1642, 592]. controllability [2289, 1538, 1611, 98, 1098]. controlled [17, 1693]. controller [96, 2281, 10, 642, 716, 974, 675, 955, 122, 1027]. controllers [1654, 1092]. Controlling [2131, 473, 385, 1514]. controls [448, 1740]. convection [761, 380, 677, 722, 311, 2180, 1126, 1505, 2129, 1263, 1337, 1261, 246, 167, 787, 563, 1378, 753, 1155, 2162, 1266, 933, 1730, 1939]. Convective [537, 779, 954, 951, 800]. convectively [534, 905]. conventional [17, 22]. Convergence [839, 197, 1721, 1554, 1912, 1920, 1728, 1223, 1214, 1829, 2239, 1364, 2204, 226, 1993, 1344, 2217, 1739, 1666, 823, 927, 1915, 1902, 673, 1149, 204, 217, 1491, 387, 1629, 1317, 1278, 933, 1939]. convergent [769, 900, 167, 2098, 1052, 797, 1377, 843, 976]. conversion [441]. conversions [1309]. converter [1654, 11, 2202, 1760, 5, 1732]. converter-based [1654]. converters [1761, 440, 421, 642, 34, 435]. convex [1508, 1482, 557, 1100, 239, 447, 2244, 1870]. convexity [1625, 1614]. conveying [1129, 741]. convolution [903]. convolutional [228, 1866]. cooperation [1690, 1667, 1613, 1949, 691]. cooperative [2003, 94, 1809, 1492, 1533, 1635]. coordinating [1230]. coordination [559, 1966, 1773]. COPRAS [1684]. COPRAS-based [1684]. corner [197]. corona [1211]. correct [29]. corrected [475, 1885]. correction [95, 1867, 1658, 1947]. corrector [550, 358, 628, 579, 331, 1035, 1630]. correlation [2177, 1218]. Corrigendum [1752, 1751, 1750]. cortical [662, 2092]. cosine [588, 869, 1599]. Cost [1085, 2131, 1813, 614, 1007, 929, 1865, 1342, 420, 1310, 817]. costs [1468, 228]. cottontail [2110]. Coulomb [1496, 1127]. count [1176]. country [1256]. Couple [1303]. coupled [247, 71, 1585, 859, 1035, 1554, 880, 892, 155, 2099, 2081, 2254, 1262, 156, 527, 2174, 1089, 2181, 1956, 2292, 1152, 1120, 530, 2282, 75, 2044, 1548, 1931, 1215, 1853, 815, 1844, 1461, 2127, 2222]. Coupling [1523, 295, 1668, 549, 1373, 110, 1464, 2187, 1461]. couplings [1017, 2125]. Cournot [971, 1880, 960]. Cournot-type [1880]. CoV [2083, 1058, 1087]. covariance [675, 392]. cover [2005]. coverage [1159, 1071, 1422, 1523, 809]. covering [1715]. coverings [484]. COVID [653, 834, 1013, 912, 1941, 856, 727, 2061, 744, 2223, 1159, 2230, 1229, 350, 928, 858, 1256, 1510, 857, 1523, 499, 1471, 2064, 1314]. COVID- [2223]. COVID-19 [653, 834, 1013, 912, 1941, 856, 727, 2061, 744, 1159, 2230, 1229, 350, 928, 858, 1256, 1510, 857, 1523, 499, 1471, 2064, 1314]. Cox [1967]. cpKP [1240]. cpKP-BKP [1240]. crack [28]. cracks [321, 1505]. craft [281]. Crank [1150]. CRC [592]. Credit [1319, 365, 647, 157, 397, 1671]. criminal [1365]. crisis [1343]. criteria [285, 231, 1449, 1656, 978, 2209, 106, 807, 1744]. criterion [1173, 1065, 275]. Critic [1966, 1963]. critical [1940, 309, 1199]. crop [554]. crop-pest [554]. Cross [1001, 1613, 1129, 1184, 776, 577, 1218, 2196, 1099, 1256, 2293, 1423]. cross-border [1256]. cross-correlation [1218]. Cross-diffusion [1613, 577, 2196, 1099, 1423]. cross-diffusion-ODE [2293]. cross-infection [1184]. cross-nano-material [1129]. cross-section [776]. crossover [1248]. crowd [724]. Crowley [1667, 1949]. crude [1582, 1457]. cruise [94]. cryptography [2039]. crystal [1154]. crystals [543]. CSP [454]. CT [147]. CTL [603, 601]. CTLs [1367, 1953]. Cu [779]. Cu-Al [779]. Cubature [654]. cube [206, 503]. Cubic [2036, 156, 838, 2262, 462, 1811, 769, 192, 1598, 1624, 323, 1463, 1874, 2249, 2140, 289, 398, 46]. Cucker [1700, 2125]. CUDA [352]. cultural [85]. cumulative [2280]. curative [1632]. curing [665]. currency [1207]. current [1761, 1051, 1778, 951, 24]. currents [1694]. curvature [958, 1676, 346]. curve [786, 1996, 1724, 1825, 261]. curved [461, 741, 2297, 1954]. curves [1657, 1458, 2118, 663, 2164, 840, 1197, 280]. curvilinear [1009]. cusps [829]. customer [1073, 464]. customers [2080]. customized [873]. cut [112, 1564]. cut-cell-based [112]. cutting [197]. CVNs [1541]. CWOA [1212]. cyber [699, 1651, 93, 2298]. cycle [731, 1088, 1916]. cycles [1799, 2229, 1999, 2142]. cylinder [908, 1129, 473, 800]. cylindrical [48, 408, 1056]. cytotoxic [334, 1720].

D [1437, 79, 923, 1900, 1917, 1939]. D-MOEDA [1437]. D-study [1900]. D3Q27 [2140]. dam [801]. damage [2126, 801]. damaged [1548]. Damping [1205, 778, 2157, 573, 496, 1366, 1801]. Daniel [2099]. Darboux [1550]. Darcian [753]. Darcy [380, 71, 954]. Darwinian [1515, 2001]. Data [2195, 1881, 653, 514, 1743, 192, 1926, 1969, 973, 1941, 600, 1838, 1538, 663, 1441, 1414, 667, 2232, 2011, 1030, 2135, 1449, 1656, 974, 1257, 265, 1861, 1819, 2256, 2054, 1986, 1964, 2154, 2092, 872, 763, 1025, 1882, 2119, 2026]. Data-driven [2195, 1881, 2011, 2135, 1986]. Day [1767]. Day-ahead [1767]. DC [1654, 2003, 450, 448, 421, 642, 1642, 423, 1760, 2231]. DC-DC [1654]. DC-link [423]. DC/DC [1760]. DCP [95]. deal [769, 542]. dealiasing [383]. Dealing [588]. DeAngelis [833]. death [1945, 530, 2193]. decay [2161]. deceased [834, 2076]. deceleration [874]. December [257, 639, 988, 1421, 1851, 2237]. decentralised [422, 440]. decentralized [2051, 110]. deception [2282, 1901, 1651]. Decision [2093, 1037, 1217, 173]. decision-making [1037]. decisions [1310]. decoder [1767]. Decomposition [2067, 532, 861, 1097, 618, 1251, 246, 100, 2213, 1241]. Decoupled [1333, 1778, 1476, 1566, 2276, 1196, 1372]. decreasing [511]. dedicated [424, 34]. Deduction [275]. Deep [1503, 454, 1547, 1763, 464, 1920, 898, 1005, 645, 1592, 848]. defense [875, 783, 1864]. deferred [1867, 1947]. defined [1475]. definite [996]. deformation [480, 1805, 165]. deformed [1064]. degeneracy [1976]. degenerate [1477, 797, 1583]. degenerated [862]. degradation [1764, 179]. degraded [26]. degree [1787, 1712, 1876, 840, 1197, 1217, 654, 478]. degree-graded [654]. Delay [128, 1338, 485, 1508, 2279, 1117, 2217, 962, 96, 62, 107, 785, 668, 1602, 2141, 833, 1788, 320, 1729, 232, 1723, 262, 2294, 1952, 1218, 740, 489, 844, 1786, 1711, 1577, 1902, 162, 673, 2278, 391, 1327, 1656, 2027, 1492, 106, 1322, 69, 1311, 530, 1349, 2282, 2218, 1611, 670, 1678, 766, 691, 298, 703, 939, 999, 1245, 2063, 980, 275, 795, 1661, 277, 336, 1704, 1938, 1501, 820, 1894]. delay-based [2027]. Delay-dependent [1338, 1508, 1711, 1656, 980]. delay-induced [668, 833, 1786, 1492]. Delayed [802, 2091, 971, 270, 1054, 1151, 231, 580, 1593, 2144, 1451, 136, 2087, 1286, 1596, 816, 1141, 974, 1379, 1219, 2000, 1873, 916, 315, 1255, 2079, 98, 309, 2263, 493, 2295, 1025, 1362, 1635, 2089, 1000, 901]. delays [109, 1347, 737, 1367, 1745, 1708, 1880, 727, 868, 1966, 1135, 531, 699, 94, 2083, 103, 108, 966, 1841, 1700, 354, 1157, 1579, 1376, 280, 1301, 391, 941, 692, 1136, 969, 1388, 101, 1511, 111, 706, 981, 2263, 403, 1065, 2187, 370, 1817, 980, 1571, 846, 1062, 1000]. delivery [1303]. delta [2189]. demand [1978, 1231, 2108, 452, 2250, 99, 920]. demands [1195]. demographic [2291]. Dengue [774, 2144, 1036, 2049, 794, 1274, 585]. denoising [662, 2213, 2044]. denoising-ARIMA [2044]. denominator [927]. dense [261]. densities [2287]. density [1050, 2020, 674, 997, 2248, 364, 1760, 795, 1890, 809]. density-dependent [2020, 1890]. dental [147]. Deoxyribonucleic [241]. Dependence [926]. dependency [924]. Dependent [2088, 889, 380, 1508, 123, 1404, 694, 1945, 1386, 2156, 707, 1559, 1264, 1752, 2020, 1447, 1456, 1856, 900, 1261, 924, 222, 965, 1711, 819, 2208, 1990, 391, 1519, 1656, 1024, 1858, 932, 1176, 1338, 865, 763, 2070, 1984, 750, 2295, 1882, 1353, 346, 980, 1062, 1890]. depending [1997]. depends [1231]. deployment [174]. deposition [465]. derivation [424]. Derivative [472, 80, 304, 593, 889, 1550, 1086, 995, 2113, 1258, 1640, 1507, 638, 784, 1574, 2288, 818, 1867, 2041, 1370, 1653, 596, 1156, 587, 2213, 1062]. Derivative-based [80, 304, 784]. Derivative-orthogonal [593]. derivatives [1016, 583, 2122, 2121, 166]. DES-PSO [326]. desalination [33]. descent [2203]. described [1457, 398]. describing [538, 1595, 993, 1621]. description [1928]. descriptor [1031]. Design [364, 25, 1027, 60, 2179, 955, 31, 437, 1409, 1585, 70, 426, 699, 1840, 449, 1100, 326, 1631, 354, 642, 423, 83, 974, 1726, 27, 120, 1530, 144, 34, 1320, 983, 2231, 846, 425]. designs [445]. destination [1256]. detecting [55, 24]. Detection [1769, 446, 2019, 28, 1037, 288, 2263, 821]. detectivity [1839]. detects [1445]. deteriorating [1231, 2198, 1820]. Determination [1986, 2030, 414, 529]. determine [415]. determined [849, 547]. Determining [564]. deterministic [1137, 1451, 1302, 519, 1832, 1210, 916, 529, 261]. developed [776, 1459]. developing [2280, 1256]. Development [1494, 959, 1462, 2044, 174, 642, 1878]. developments [510, 2057]. deviation [991]. deviation-Yager [991]. device [127]. devices [412]. DFIG [23]. DG [1353]. diabetes [961, 1538, 2168, 873]. diabetic [1314, 1086, 2223, 2243, 941, 758]. Diagnosis [1004, 424, 972, 704, 1037, 412]. Diagonalization [1447, 1275]. diagonally [2140]. Diamond [1093, 385]. DIC [630]. dictionaries [1791]. dielectric [1398]. Difference [998, 538, 267, 825, 330, 1336, 1122, 114, 1285, 2059, 1534, 1403, 694, 1035, 1602, 1123, 1132, 1386, 343, 1899, 888, 1193, 904, 1306, 1795, 364, 1465, 860, 602, 1588, 1989, 1277, 409, 476, 1325, 2207, 167, 1305, 1041, 1267, 2190, 1152, 535, 765, 1142, 1466, 319, 1448, 1923, 209, 739, 372, 1192, 263, 373, 1245, 488, 2162, 810, 1321, 1843, 1730, 1950, 1150, 2245, 1885]. difference/compact [765, 1142]. differences [1897, 166]. differencing [2013]. Different [1725, 1163, 302, 1702, 1239, 236, 582, 108, 1397, 590, 2108, 1191, 690]. differentiability [1982, 1352]. differentiable [106, 2187]. Differential [1335, 2171, 1562, 1547, 1440, 1079, 1473, 2189, 2036, 1584, 1469, 870, 1733, 1503, 2038, 125, 1401, 284, 1650, 1799, 325, 694, 849, 2289, 2257, 550, 42, 1559, 320, 780, 1669, 2229, 1182, 2042, 1999, 1507, 1809, 2277, 652, 602, 1612, 844, 381, 1145, 1497, 711, 1989, 2243, 1350, 1148, 1872, 1902, 149, 1782, 1158, 1408, 1392, 1674, 1765, 2202, 934, 1639, 1380, 759, 1429, 1573, 1113, 2112, 1311, 1829, 1384, 1804, 1169, 596, 1630, 331, 981, 739, 1299, 1361, 648, 218, 1266, 822, 533, 1467, 1843, 2222, 277, 336, 1434, 1098, 798, 1150, 1501, 2264]. differentiation [1484, 125, 2258, 1028]. differentiation-enhanced [2258]. Diffused [1812]. diffusion [1402, 1344, 1045, 775, 390, 560, 553, 1482, 2279, 1733, 1238, 583, 2285, 1209, 1336, 799, 995, 84, 1708, 722, 615, 362, 1403, 1241, 1436, 1855, 952, 1554, 343, 2067, 888, 1189, 349, 605, 616, 1306, 1748, 577, 1797, 104, 1505, 262, 1952, 1355, 486, 1588, 1701, 1261, 2130, 1001, 246, 483, 1036, 1796, 1577, 1277, 591, 1827, 401, 377, 562, 1513, 1734, 219, 167, 2196, 1141, 379, 925, 1369, 1613, 523, 787, 1379, 2255, 535, 1093, 1800, 563, 1322, 1250, 2013, 719, 1052, 1378, 1099, 1225, 2253, 1410, 298, 1600, 1629, 1081, 209, 1422, 372, 406, 862, 1396, 1255, 1984, 2056, 587, 1055, 1420, 1953]. diffusion [2090, 1245, 1155, 1399, 2293, 1423, 1278, 2120, 933, 1179, 1844, 1673, 1578, 1730, 1890, 2025, 1939, 1938, 1524, 1628]. diffusion-type [1250]. diffusions [1048]. diffusive [2279, 1650, 617, 2191, 752, 1001, 1174, 1892, 1678, 706, 1059, 547, 755, 1890, 2260]. diffusivity [800]. digital [2126, 440]. Dimension [1910, 1160, 78, 1205]. Dimensional [137, 1493, 558, 390, 918, 796, 870, 583, 2136, 330, 1398, 2059, 1035, 1386, 343, 892, 1559, 2266, 1189, 471, 1845, 2107, 72, 2081, 845, 1090, 2215, 841, 915, 1251, 1875, 824, 2174, 246, 879, 1240, 1095, 1973, 279, 1170, 1217, 1827, 1532, 556, 2116, 1325, 1674, 1833, 2140, 1223, 982, 1658, 491, 1093, 1858, 563, 2265, 710, 1061, 1222, 1676, 1253, 887, 1381, 1382, 876, 1309, 729, 347, 810, 2127, 2245, 503]. dimensionality [1536]. dimensions [1440, 1983, 1132, 650, 893, 765, 1142, 2154]. diode [426, 1825]. dioxide [1722, 2109]. Dirac [1254, 2265, 2127]. Direct [1082, 23, 1586, 414, 400, 1475, 1733, 1051, 208, 1979]. Directed [1437]. direction [1403, 2253, 1364, 1635]. directional [450, 764]. directions [1477]. Dirichlet [1581, 535, 1420]. disaster [300]. disasters [2080]. disc [919, 1933, 1703]. discharge [1398]. discontinuities [1324, 1316]. discontinuity [1999, 1548]. Discontinuous [301, 190, 1761, 2272, 1799, 722, 1668, 1172, 49, 1999, 1952, 248, 57, 1539, 871, 396, 2034, 1979, 587, 1055, 2287, 2245]. discouragement [1813, 1430]. Discrete [653, 2269, 2128, 1433, 1875, 1573, 1844, 782, 839, 1336, 737, 1662, 1880, 311, 2270, 469, 1146, 2252, 697, 1431, 1808, 2294, 1242, 1167, 360, 1801, 1595, 987, 1833, 2106, 1327, 1219, 1873, 1909, 111, 1901, 298, 703, 2045, 386, 1375, 1382, 2080, 1842, 1065, 1744, 1524, 869]. discrete-delay [703]. Discrete-Ordinate-Lattice-Boltzmann [1875]. discrete-time [1662, 1880, 469, 1327, 1219, 1873, 111, 1901, 703, 386, 2080, 1744]. Discretisation [1924, 184]. Discretised [1258]. discretization [620, 72, 2035, 841, 2082, 504, 523, 1653, 865, 2120, 1844]. discretizations [84]. Disease [1274, 592, 1632, 1720, 1884, 2223, 1037, 399, 1414, 1397, 222, 794, 939, 2160, 1425, 1571, 2260, 585]. diseases [1151, 1896, 405, 1575, 1850, 2236, 1621]. disk [44, 1197, 105, 235, 837, 1750]. dispersal [1066]. Dispersion [629, 1082, 2191, 1307, 1149, 303, 2296]. dispersion-type [1082]. dispersive [216, 2159, 1056, 1128]. displacement [1731, 629]. displacement-based [629]. Dissipation [573, 1454, 1687, 1263, 239]. Dissipation-preserving [573, 239]. dissipative [461, 779, 416, 1291]. Dissipativity [1461, 1208, 980]. Distance [1587, 70, 1699, 1825]. distancing [1852, 928]. Distributed [1913, 2003, 2176, 1117, 1336, 737, 2006, 1436, 1193, 349, 966, 489, 578, 1286, 1700, 1460, 1525, 391, 1305, 1041, 759, 1773, 530, 1678, 1225, 760, 1104, 1834, 795, 1062, 1661]. distributed-order [1193, 349, 1460, 1041, 1225]. Distribution [1437, 553, 1770, 1789, 612, 1404, 1073, 344, 674, 1090, 1431, 666, 520, 414, 1948, 1405, 214, 1087, 1773, 522, 2054, 723, 1375, 1427, 475]. distributions [2023, 359, 55, 678, 863, 1060, 1257, 1127, 2221]. Disturbance [781, 1904, 1898, 1351]. disturbances [1702, 447, 878, 1904, 2089]. Dive [185]. diverge [385]. divergence [71]. divergence-free [71]. diving [185]. divorced [2302]. DMLPG [543]. DNS [383]. dockless [245]. documents [29]. Does [2259]. Domain [2205, 861, 1097, 618, 1123, 1050, 2137, 1818, 246, 335, 523, 2201, 1187, 1634, 500, 2297]. domains [1493, 1550, 646, 945, 1400, 1505, 1168, 239, 1290, 2297]. dome [1320]. dominated [991, 1881]. dominating [1930, 1206]. double [2111, 1745, 1160, 1919, 742, 105, 1178, 1115, 2216, 1762, 1615, 2070, 1636]. double-bioreactor [2070]. double-stage [1762]. double-strains [1745]. doublets [1007]. DQM [1874]. drafting [236]. drag [1002]. dragonfly [272]. dragonfly-inspired [272]. drift [871]. drive [26, 1058, 447]. driven [1553, 1926, 363, 2223, 281, 2011, 2195, 2140, 2135, 1881, 686, 1986, 1246, 2188, 2060]. drivetrain [687]. droop [2003]. droplet [1805]. drug [1303, 601]. drug-resistance [601]. drugs [1427]. DTC [17, 22, 25]. DTC SVM [22]. Dual [462, 559, 622, 1091, 594, 1319, 49, 992, 2200, 2192, 1579, 1105, 1087, 523, 99, 907, 2161]. Dual-channel [559, 1091, 1105, 99, 907]. dual-memristor [2200]. dual-porosity-Stokes [2192]. dual-time [49]. Dual-wave [462]. Duan [532]. duct [1377]. due [44, 58, 905, 837, 1750]. Duffing [778, 323, 2288, 683]. Dufort [347]. Dufour [237]. duopoly [2131, 960, 1214]. duration [1043]. during [1013, 594, 295, 1591, 992, 1834, 2204]. dust [364]. DY [1266]. DY-HS [1266]. dyadic [791, 1679]. Dynamic [1367, 1883, 1556, 1970, 960, 1877, 2011, 2152, 1919, 2201, 601, 165, 1822, 1428, 2298, 1471, 848, 2086, 864, 2088, 450, 1972, 2023, 297, 1941, 6, 856, 1977, 157, 1340, 1723, 2224, 2039, 222, 1053, 2219, 159, 416, 1034, 369, 1119, 603, 2236, 1857, 1964, 760, 930, 1601, 2197, 1499, 843, 449, 491]. Dynamical [1409, 1663, 791, 2163, 833, 363, 2144, 1146, 1247, 957, 2291, 813, 2209, 939, 961, 1147, 962, 365, 1094, 1585, 2121, 1239, 2020, 549, 1886, 1005, 1376, 1387, 1348, 1990, 989, 269, 2241, 207, 1498, 396, 1293, 1611, 1338, 417, 493, 1842, 2185, 2293, 1673, 669]. dynamically [1146, 530]. Dynamics [570, 345, 62, 1880, 1896, 2083, 1701, 2174, 965, 1096, 1040, 357, 1863, 816, 812, 87, 1178, 905, 1810, 1961, 1015, 1621, 902, 761, 749, 1333, 2153, 1802, 664, 1862, 2122, 1051, 107, 1536, 1720, 776, 1928, 1831, 531, 1855, 2270, 1812, 1788, 1669, 687, 1153, 1946, 2096, 2301, 2002, 332, 1997, 1952, 334, 741, 2242, 734, 866, 824, 1575, 2243, 2200, 2029, 1801, 74, 1670, 1597, 1833, 2001, 1667, 2135, 1141, 941, 1229, 233, 1492, 2255, 2286, 565, 1174, 265, 914, 1214, 2049, 794, 1274, 1288, 1221, 227, 2302, 1980, 1971, 1736, 1074, 2206, 2034, 1071, 1533, 1523, 1976, 1953, 572, 755, 1425, 2197, 2170]. dynamics [2127, 809, 1628, 210, 908, 1482]. dynamo [919].

eagle [742, 1063]. Earthquake [801]. eastern [2110]. easy [925, 34]. EBHSV [2110]. EBHSV-infected [2110]. eBook [592]. echinococcosis [1708, 1153]. echo [452]. eco [1482, 833, 1178]. eco-epidemic [833, 1178]. eco-epidemiological [1482]. ecoepidemic [783]. ecological [1363, 2270, 644, 2199]. ecology [1828]. economic [1663, 1230, 934, 1686]. economic-epidemiological [1663]. ed [592]. edema [238]. edge [168]. Edition [2057]. Editorial [36, 37, 38, 52, 66, 91, 119, 141, 152, 171, 188, 213, 255, 294, 314, 339, 376, 411, 433, 457, 509, 526, 568, 569, 608, 634, 636, 659, 682, 715, 743, 790, 806, 828, 852, 883, 910, 911, 944, 986, 1020, 1070, 1110, 1162, 1202, 1234, 1271, 1294, 1296, 1330, 1358, 1391, 1419, 1444, 1480, 1487, 1500, 1529, 1570, 1605, 1645, 1682, 1707, 1755, 1756, 1779, 1849, 1889, 1936, 1957, 2010, 2048, 2073, 2102, 2147, 2167, 2184, 2212, 2235]. Effect [70, 1914, 1865, 225, 425, 1256, 2120, 1890, 116, 1632, 429, 570, 594, 270, 580, 752, 1010, 2053, 721, 1039, 1337, 644, 960, 965, 2291, 1951, 2196, 1515, 2001, 690, 1613, 2078, 128, 2255, 1850, 1892, 914, 308, 1288, 1178, 1115, 1736, 287, 8, 1099, 1510, 1615, 2293, 1423, 1499]. Effective [2065, 1616, 998, 302, 232, 542, 1989, 1677, 458]. Effects [1555, 1074, 1396, 2110, 1710, 1862, 1663, 1720, 997, 237, 917, 1517, 1687, 757, 1124, 1596, 1919, 1552, 1832, 1879, 2117, 1892, 753, 1533, 266, 1428, 1012, 346, 846, 837, 1750, 704, 1949, 2263]. efficacy [1427]. efficiency [1809, 1191, 1288, 564, 383]. Efficient [720, 902, 193, 1806, 155, 924, 2192, 2074, 1513, 2013, 1964, 2193, 1842, 497, 1345, 838, 1402, 558, 2068, 241, 1318, 2226, 768, 77, 302, 2038, 55, 325, 1436, 823, 1281, 1463, 1649, 2247, 1918, 1481, 1289, 126, 1149, 1380, 1619, 1727, 490, 2077, 1869, 1297, 596, 1061, 884, 1165, 886, 1317, 500, 1195, 249]. effort [1105, 2198, 99]. EG [366]. Ehrenfest [1114]. Eigendecomposition [440]. eigenfunctions [490]. eigenproblems [1561, 810]. eigenvalue [2021, 2074, 646]. eigenvalues [1180, 1522, 490, 886, 1561]. eigenvectors [1438]. eighth [1666]. Einstein [1981, 2173]. EKF [1731, 2064]. elastic [889, 619, 1553, 625, 628, 2201, 477, 1056, 112, 2161, 1290]. elasticity [541, 240, 2108, 1191, 384, 893, 700]. elastodynamic [1637]. elastodynamics [629, 621]. elastoplastic [1481]. elastostatic [1637]. election [949]. electric [7, 1383, 1767, 13, 20, 5]. Electrical [1764, 450, 442, 1775, 14, 453, 449, 1595, 452, 434, 1177]. electricity [182, 1978, 920]. ELECTRIMACS [434, 1757, 412, 1758]. electro [413, 2157, 2276]. electro-hydraulic [2157]. electro-hydrodynamic [2276]. electro-thermal [413]. electrode [415]. electrohydrodynamics [1196]. Electromagnetic [127, 32, 428, 21, 1866]. electromagnetics [1354]. electromechanical [223, 1622]. electron [1137]. electronics [416]. electroosmotic [355]. electrostatic [1127]. Element [658, 1565, 836, 133, 619, 1045, 761, 1452, 1097, 1394, 956, 1490, 623, 2217, 1147, 1029, 2285, 892, 1307, 1742, 1189, 127, 1203, 63, 1748, 2002, 595, 486, 360, 1261, 2175, 57, 239, 1047, 1366, 126, 1648, 219, 384, 504, 621, 59, 1346, 1474, 2292, 259, 1968, 1881, 1629, 1081, 613, 53, 598, 574, 1476, 1564, 2056, 1954, 1618, 1939, 1434, 282, 466, 1126]. element-free [1748, 1047]. element/interpolation [1474]. elementary [399]. elements [77, 1483, 1598, 2033, 49, 248, 2175, 629, 48, 1111, 282]. elevator [13]. eliminate [1185]. Elimination [19, 437, 439, 168]. elite [185, 1095]. ellipse [2267]. ellipsoidal [1804]. elliptic [1897, 2180, 1168, 467, 831, 2074, 335, 305, 1407, 1166, 408, 53, 263]. elliptical [753]. elongated [643]. embedded [451, 534, 890, 441]. ember [177]. Emden [1297, 998, 1918, 1078, 286, 120, 576, 506, 771, 161, 1165]. emergency [724]. emerging [1918]. EMHD [2076]. emission [1074]. emissions [1722]. emissivity [1986]. emphasising [1058]. empirical [1175, 2270]. emulator [416]. enclosed [2040]. enclosure [380, 311, 1126, 1875, 1039, 1337]. Encoded [1014]. encoder [1767]. encoding [2284, 1212]. encryption [241, 2091, 1280, 2247, 2200, 1076, 2284, 291, 269, 1075, 2179, 1183, 2094, 503, 864, 1220, 1583, 1894]. Endemic [829, 746, 2236]. enemy [554, 1976]. energetic [224]. energies [4]. ENERGY [412, 2, 645, 2127, 1771, 1140, 7, 1958, 1789, 1759, 453, 422, 1281, 1587, 697, 461, 1767, 2276, 482, 1134, 1495, 2250, 1642, 1143, 665, 2058, 35, 1127, 1468, 1561, 430, 598, 760, 1476, 2062, 2161, 1954, 373, 1814, 1543, 1718, 1757]. energy-conserving [1495]. Energy-preserving [2127, 1281, 697, 1134, 2062, 373, 1814]. energy-related [453]. energy-stable [2276]. engine [185]. engineering [2126, 1840, 2219, 712, 1369, 867, 1006, 1610, 434, 434, 609, 1749, 1758]. Enhanced [85, 231, 840, 2258, 2219, 1326, 1449, 1181, 460, 353, 867]. enhancement [2129, 1349, 545]. Enhancing [1819]. ENO [1811]. Enriched [2224, 1742, 329, 1564]. enrichment [2033]. Ensemble [1962, 902, 1699, 1649, 2093, 1181, 2188, 2060, 1688]. entanglement [138]. enterprise [1671]. entropic [88]. Entropy [105, 837, 1750, 1129, 2280, 991, 1383, 1912, 461, 1039, 1794, 954]. environment [510, 1066, 2191, 1994, 1962, 2261, 1083, 2057, 2093, 1388, 1890, 474]. environmental [1862, 1315, 1105, 1914, 565, 2206, 672, 1751]. environments [749, 1832, 916]. EP [540]. Epidemic [2000, 1584, 2279, 1066, 1077, 1855, 2191, 833, 674, 1146, 1211, 405, 1997, 2011, 2291, 401, 1664, 1302, 1665, 2041, 1597, 937, 1955, 265, 308, 1178, 1971, 865, 1923, 1422, 386, 1374, 1523, 67, 572, 2204, 809, 586]. epidemiological [834, 1482, 1663, 1946, 1563, 491, 1071, 774, 2283]. epidemiology [1740]. equal [282]. equal-order [282]. equalities [304]. equation [838, 133, 1033, 388, 720, 1163, 1616, 2088, 1318, 462, 1550, 1452, 768, 918, 132, 302, 318, 1238, 583, 125, 2217, 842, 330, 1147, 2285, 129, 964, 1336, 1620, 799, 114, 1285, 1624, 566, 722, 2159, 776, 1403, 1602, 1436, 2017, 1721, 1123, 1132, 343, 1899, 888, 1189, 1252, 471, 349, 1846, 323, 328, 1281, 2128, 1807, 1226, 2081, 1640, 1433, 542, 573, 1748, 108, 697, 1716, 1463, 156, 2301, 650, 472, 299, 652, 1355, 1023, 1612, 1251, 1588, 1167, 1289, 711, 1261, 2174, 548, 879, 1047, 1089, 1373, 1240, 1254, 1577, 279, 482, 1277, 1170, 1782, 2274, 102, 831, 377, 556, 2116, 1595]. equation [1525, 1734, 1990, 599, 813, 1933, 1474, 523, 923, 940, 893, 1161, 1267, 685, 217, 2190, 2286, 765, 1142, 1573, 1113, 303, 2265, 286, 120, 506, 100, 1125, 1804, 771, 131, 710, 1792, 884, 1225, 1634, 1410, 1600, 1629, 209, 398, 1215, 581, 1406, 2069, 500, 1208, 2056, 1055, 717, 264, 750, 876, 1309, 2173, 1954, 271, 990, 729, 1192, 648, 669, 1339, 347, 2162, 2290, 756, 810, 795, 1321, 2127, 814, 1434, 798, 2220, 2245, 153, 1524, 46, 210, 1156]. equation- [1156]. equations [1562, 979, 1547, 1082, 1440, 1079, 1473, 1719, 1606, 2189, 2036, 390, 620, 861, 1097, 1469, 1821, 1897, 220, 870, 956, 1490, 1733, 1503, 2038, 284, 243, 84, 694, 1135, 849, 1241, 2114, 1035, 1554, 550, 880, 823, 892, 1239, 1998, 2180, 1559, 320, 780, 1193, 1660, 1566, 1182, 557, 2099, 232, 505, 1507, 1689, 2300, 1306, 1262, 1795, 527, 1259, 1505, 2037, 2277, 1918, 602, 836, 551, 844, 381, 1145, 1497, 824, 1989, 2243, 239, 134, 246, 483, 903, 818, 1424, 1350, 1148, 1872, 1902, 2276, 1886, 163, 142, 544, 1495, 1801, 1956, 591, 1827, 409, 1408, 1826, 673, 1392, 1325, 1648]. equations [1674, 305, 205, 1728, 1305, 925, 1639, 322, 1658, 2292, 1380, 759, 1429, 535, 319, 2112, 1311, 1829, 1078, 1384, 2013, 1169, 1052, 596, 996, 161, 1061, 1378, 1630, 1222, 2253, 1165, 811, 921, 298, 2244, 331, 1253, 739, 1299, 598, 372, 1361, 1564, 373, 1245, 1694, 488, 2063, 1814, 1266, 1278, 1353, 822, 41, 1316, 533, 2120, 1737, 1467, 815, 1372, 1179, 1843, 1730, 1950, 277, 336, 1098, 1944, 282, 1150, 1501, 1885, 2264, 2273]. equidistribution [900]. equilibria [1169]. Equilibrium [1582, 746, 260, 2171, 110, 978, 1800, 753, 976]. equipped [1103, 1614]. equity [700]. Equivalence [1291]. equivalent [1774, 435]. ergodic [1211, 1087]. ergodicity [359]. Ericksen [1372]. Error [1336, 1277, 1142, 814, 1524, 2272, 1581, 1436, 72, 2205, 135, 1284, 1912, 1218, 1472, 1373, 1254, 2115, 1648, 511, 1658, 1267, 2292, 1968, 921, 1600, 975, 1245, 2025, 2220, 1885]. errors [2082]. escalating [1365]. escape [807]. ESG [1671]. esterification [395]. estimate [2292, 2040, 1524]. estimates [2272, 1926, 1581, 95, 1254, 1267, 1600, 1245, 814]. Estimating [2141, 1824, 687, 490, 886]. Estimation [553, 1437, 367, 1948, 447, 1519, 31, 1761, 1806, 470, 501, 1620, 678, 512, 1050, 2172, 666, 740, 1405, 1541, 27, 1429, 649, 516, 522, 801, 443, 2052, 1368, 2154, 2092, 857, 180, 975, 497]. estimations [1246]. estimator [354]. estimators [553, 1107, 475]. ETD [615]. ethylene [1685]. Euclidean [1825]. Euler [558, 1550, 832, 931, 1660, 1566, 2164, 1366, 1902, 2251, 1236, 226, 1844]. European [114, 1582, 248, 2207]. EV [1763]. evacuation [724]. evaluate [7]. Evaluating [512, 358, 2236]. Evaluation [321, 1383, 14, 12, 2126, 1127]. Event [369, 1530, 1793, 1651, 2187, 1659, 2123, 1053, 1541, 1359, 1596, 853, 1901, 999, 1104, 1601, 1904, 1635, 1938]. Event-triggered [369, 1530, 1793, 1651, 2187, 1659, 2123, 1053, 1541, 1359, 1596, 853, 1901, 999, 1104, 1601, 1904, 1635, 1938]. Events [51, 65, 90, 118, 140, 151, 187, 212, 254, 293, 313, 338, 375, 432, 508, 525, 607, 633, 637, 681, 714, 726, 789, 805, 827, 851, 882, 943, 985, 1019, 1069, 1109, 1130, 1131, 1201, 1233, 1270, 1295, 1329, 1357, 1389, 1418, 1443, 1479, 1486, 1528, 1569, 1604, 1644, 1681, 1706, 1754, 1777, 1848, 1888, 1922, 1935, 2009, 2047, 2072, 2101, 2133, 2146, 2183, 2211, 2234, 410, 2268, 170, 2022, 456]. evolution [834, 1057, 42, 295, 818, 245, 2116, 940, 765, 576, 506, 218, 899, 1974, 1335]. evolutionary [1771, 1594, 764, 2024, 1590, 364, 1636]. evolving [2032, 2098]. EWSO [1840]. Exact [388, 348, 1048, 1, 389, 613, 669, 1173, 651, 308]. example [857]. excellent [394]. exchange [1143, 2155, 1207, 148, 1617]. Excitability [2109]. excitation [442, 63, 1886, 683]. excited [889, 1205]. exercise [317]. exhaust [9]. exhibiting [501]. Existence [625, 1198, 969, 1633, 153, 877, 2251]. exogenous [297, 1839, 395]. exoskeleton [718, 1464]. expanded [861, 1974]. expansion [514, 1241, 2114, 1213, 784, 1187]. expansions [2111]. expectation [2221]. expectations [2131]. expected [1008]. experimental [1654, 1538, 22]. experimentation [185]. experiments [1496, 630, 425]. Explicit [382, 277, 1226, 1991, 203, 399, 1495, 2207, 1513, 801, 1250, 1178, 811, 1361, 373, 1814, 976]. Exploiting [416]. exploration [2028, 947, 1951]. explored [757]. Exploring [1625, 961, 2031, 655, 2077]. explosion [1633]. Exponential [560, 1219, 111, 768, 1347, 737, 1586, 1281, 966, 405, 1874, 1948, 1286, 2115, 1301, 274, 1257, 2095, 2282, 2013, 1169, 981, 2079, 1947, 1025, 2127, 1939, 277, 901]. exponential-Rayleigh [1948]. exponentially [1574, 2098]. exponentially-fitted [1574]. exposed [321]. expressions [1977]. Extended [623, 1714, 1625, 1029, 566, 880, 1239, 328, 1962, 1691, 279, 2155, 1119, 1967, 858, 387, 807, 264, 876, 980, 1335]. Extending [129, 1818]. Extension [1106]. external [1304, 297, 677, 1702, 323, 1904]. Extinction [402, 1087, 67, 1327]. extragradient [702, 976]. extrahepatic [1499]. extrapolation [1306, 1111]. extremal [2164]. extreme [1945, 2055, 1907, 1113]. eye [1455].

fabrication [465]. Fabrizio [967, 638, 1221, 587]. facilitate [1988]. fact [268]. Factor [1445, 1877, 1103, 178, 1688, 843]. factor-variable [843]. factorizations [2092]. factors [2158, 1794]. fading [675, 1015]. failure [287, 763]. failures [1264, 1752, 1882]. fake [268]. false [19]. Falsi [1112, 1445]. families [2286]. family [956, 678, 2137, 48]. farms [1759]. Fast [390, 1463, 1041, 535, 765, 424, 575, 448, 1899, 695, 1893, 1748, 578, 1877, 1144, 233, 1142, 1225, 1339, 1308, 537]. Fault [1768, 31, 1926, 2261, 1292, 675, 24, 1560, 821]. Fault-tolerant [1768, 2261, 1292, 675]. faults [1769, 1793, 675, 872, 1308, 1659]. FBNES [2171]. FDM [358]. FDTD [515]. FE [349, 146]. fear [270, 668, 580, 965, 690, 1832, 230, 1613, 128, 1914, 1865, 1949, 1736, 916, 691, 1533, 817]. Feature [1684, 42, 1699, 1962, 1691, 2093, 1580, 1599]. featuring [1906]. February [40, 316, 684, 1072, 1489, 1937]. feed [452, 545]. feed-forward [452, 545]. feedback [1117, 1813, 231, 1586, 1551, 2171, 2261, 1690, 2168, 1327, 369, 1342, 853, 1245, 980]. feedforward [1547]. Feldstein [310]. FEM [643, 839, 2114, 483, 2276, 951, 60]. female [1955]. femtosecond [398]. fermentation [2070]. Fermi [1821]. Ferrite [487]. ferroelectrics [201]. ferrofluid [273, 2227]. ferromagnetic [487]. fertilizer [1034]. fever [2144]. Feynman [1990]. fiber [2099, 717, 2173]. fibers [1439, 2081]. Fibonacci [343, 1665, 2278, 1380]. Fick [79]. fictitious [618]. Field [425, 129, 677, 1774, 1977, 1154, 368, 1476, 1853, 1685, 2287, 2197, 2227, 466]. Field-Effect [425]. field-lattice [2227]. fields [1140, 55, 1660, 1472, 1552]. fifth [544, 1727]. fifth-order [544, 1727]. Filippov [1952]. filled [582, 1126]. film [465, 959, 59]. filter [2155, 675]. filterbanks [198]. filtered [1778]. filtering [362, 481, 495, 2026]. finance [1741]. financial [1806, 962, 1723, 652, 166, 1343, 1967, 766, 2034]. financing [62]. find [1477]. finder [1445]. Finding [1930, 979, 1940, 1657, 1206, 2171]. fingering [1689]. Finite [1045, 485, 1452, 998, 1123, 327, 63, 1506, 2018, 621, 1698, 106, 481, 131, 739, 1255, 1601, 701, 1183, 589, 466, 1402, 619, 538, 267, 761, 1550, 1097, 1394, 2148, 1897, 825, 77, 1490, 623, 2217, 1147, 1029, 1336, 285, 114, 1285, 1534, 1403, 2033, 1602, 952, 1386, 611, 343, 892, 1307, 1742, 888, 1189, 127, 1203, 1107, 904, 2258, 2002, 364, 1465, 595, 841, 860, 486, 1588, 248, 1261, 2175, 57, 239, 1366, 1277, 1450, 629, 409, 126, 1325, 2207, 167, 384, 391, 504, 59, 1305, 1474, 1223, 1879, 1291, 1142, 1968, 1466, 1881, 1448, 1637, 1629, 1923, 1081, 703, 209]. finite [613, 53, 598, 574, 406, 1476, 1564, 1820, 2056, 1954, 271, 1192, 263, 488, 1065, 980, 1618, 810, 1843, 1950, 1939, 1434, 282, 1150, 2245, 1885, 836, 166]. finite-difference [343, 209]. Finite-time [485, 327, 1506, 2018, 1698, 106, 481, 1601, 701, 1183, 589, 285, 1450, 703, 271, 1065, 980]. Finite-time/fixed-time [1698]. finite-volume [2148]. Fins [267]. fire [321, 123, 1975, 1906]. firefly [159]. fires [176]. fireworks [930]. First [510, 1876, 1048, 1424, 535, 1573, 857, 2221]. first-order [2221]. fish [1865]. Fisher [220, 1602, 1716, 1792]. fishery [1533]. fit [1026, 499, 2259]. fitted [727, 1574, 1223, 1609, 372]. Fitting [181, 663, 1746, 1784, 2288, 1924]. fitting-techniques [2288]. FitzHugh [363, 1045]. Fixed [1788, 2281, 878, 2089, 1708, 647, 1774, 541, 1556, 2085, 1963, 2246, 2298, 589, 1974, 474]. fixed-point [1774]. Fixed-time [1788, 2281, 1708, 647, 1963, 2246, 1698, 2298, 589, 474]. Fixed/predefined [2089]. fixing [615]. FL [592]. flame [1209, 588, 42, 1580]. flat [671]. flexibility [1217]. flexible [1129, 324, 1286, 1176, 1762]. flight [1090]. floating [5]. Flocking [107, 309, 1690]. floor [176]. Flow [417, 116, 908, 643, 1303, 2303, 968, 1469, 1454, 1558, 796, 1932, 468, 594, 273, 1503, 2136, 623, 6, 2017, 513, 1067, 44, 2076, 2156, 1668, 2030, 1845, 1395, 2107, 324, 660, 2248, 252, 367, 1565, 237, 1126, 1465, 917, 2129, 1517, 47, 1992, 358, 79, 461, 1687, 1263, 2053, 1039, 779, 2175, 590, 757, 1460, 671, 534, 492, 1854, 2202, 978, 787, 2190, 733, 954, 1466, 951, 1881, 366, 770, 932, 800, 1377, 1676, 1002, 235, 113, 266, 1364, 2287, 346, 153, 837, 1750, 2240]. flows [1009, 859, 49, 1649, 1539, 1005, 1608, 1672, 686, 1535, 383, 1543, 2227]. fluctuation [96]. fluctuations [565, 1290, 672, 1751]. Fluid [908, 968, 625, 1932, 1627, 776, 611, 1845, 480, 660, 252, 708, 146, 47, 358, 741, 1263, 1039, 671, 534, 492, 355, 954, 1672, 801, 227, 770, 905, 112, 235, 346, 2222]. fluid-saturated [1039]. fluids [597, 237, 2053, 1460, 1670, 2140]. fluorophore [1673]. flux [1774, 614, 175, 342, 492, 2244]. fluxes [536]. FMVol [1868]. Fock [462]. focusing [1774]. Fokker [330, 1836, 591, 1990, 1634, 41]. follower [1928]. following [667, 1021, 716, 2256]. food [270, 1172, 657, 577, 1504, 258, 1951, 690, 1667, 1865, 225, 1736, 691, 1976, 1810]. food-limited [258]. foods [1949]. foot [164]. FOQVNNs [846]. foraging [1063]. force [1129, 429, 741, 1687, 1039, 1608, 1519]. Forced [224, 1263, 75]. Forces [1412, 1304, 83, 893, 473, 2251, 444]. forces/moment [83]. Forchheimer [954]. forcing [323]. forecast [1775]. Forecasting [727, 1978, 1279, 453, 847, 1767, 1003, 124, 2230, 452, 454, 2044]. foreign [1207, 1617]. forensic [1320]. forensic-based [1320]. forest [176]. forestry [1906]. fork [1441]. form [1811, 863, 879, 1595, 1498, 2290, 475]. formal [653]. formation [825, 1272, 327, 1994, 577, 262, 1818, 1786, 1700, 2165, 2125, 1835]. formations [1099, 1931]. forming [1958]. forms [302, 1539]. formula [838, 1484, 125, 2067, 1326, 1990, 511, 1658, 1697]. formulated [1401, 1920, 1152]. formulation [1895, 1721, 541, 1784, 982, 598]. formulations [1313, 1496]. forward [1662, 2103, 452, 545]. foundation [63]. Four [1715, 2110, 1625, 2128, 1736, 1633]. four-population [2110]. four-species [1736]. four-step [1625, 2128]. Fourier [359, 1559, 1790, 573, 79, 1411, 1373, 1254, 1154, 310, 1868, 539, 46]. Fourier-informed [1411]. fourth [1180, 1318, 244, 1285, 1581, 1666, 1132, 1447, 1203, 371, 358, 2074, 535, 286, 921, 1361, 373]. fourth-order [1318, 244, 1285, 1132, 2074, 535, 921, 1361, 373]. Fowler [286, 506, 771, 1297, 161]. Fox [1060]. FPGA [458, 1894]. FPK [1886]. fractal [1547, 1695, 1227, 1086, 877, 1828, 1034, 596, 2154, 1540, 2025]. fractal-fractional [877]. fractals [1007, 807, 1693]. Fractional [1838, 1151, 1906, 1302, 2095, 403, 1894, 1493, 1547, 1121, 1180, 2262, 1440, 720, 485, 1227, 1632, 390, 1318, 462, 1194, 1550, 1452, 968, 1086, 1016, 918, 1238, 583, 2122, 125, 1409, 2285, 1620, 2105, 284, 995, 84, 114, 2059, 931, 1403, 1439, 270, 2121, 1436, 2289, 1123, 1132, 2113, 550, 2141, 880, 823, 343, 1899, 1884, 1559, 1790, 138, 780, 1258, 1193, 1538, 1669, 1189, 1260, 471, 349, 1521, 1224, 1182, 290, 1281, 1807, 1589, 660, 616, 232, 1284, 1723, 1507, 542, 573, 1748, 697, 1795, 577, 1247, 1323, 1505, 1465, 405, 2277, 652, 1355, 1588, 877, 381, 1145, 1497, 2242]. fractional [638, 711, 1828, 2087, 2243, 578, 239, 134, 548, 483, 818, 1350, 1148, 1460, 1577, 2011, 2165, 279, 163, 1277, 1495, 1782, 1867, 2018, 2199, 404, 2195, 1827, 377, 1596, 1408, 1034, 673, 1392, 1665, 1149, 1525, 987, 2278, 2041, 1734, 1674, 1190, 599, 1331, 1526, 1041, 1370, 925, 692, 1024, 940, 207, 1267, 1136, 762, 1380, 759, 356, 535, 1653, 1142, 1120, 265, 319, 1873, 718, 767, 1311, 1829, 490, 2077, 1448, 1221, 1125, 2013, 1611, 670, 719, 1052, 601, 766, 596, 1859, 1630, 1156, 738, 1225, 1410, 1600, 1985, 1215, 739, 2154, 1299, 372, 574, 500, 887, 1208, 1361, 1255, 1246, 1104, 1381]. fractional [2062, 2056, 587, 1055, 2016, 2090, 373, 493, 1339, 2295, 122, 1065, 2063, 2162, 1557, 370, 1278, 2213, 822, 463, 899, 275, 1321, 1635, 1461, 1718, 1661, 2127, 2089, 2025, 336, 1434, 1938, 798, 920, 1524, 1885, 820, 46, 858]. Fractional-order [1151, 1302, 403, 485, 1227, 1238, 2122, 1439, 2141, 343, 138, 1260, 1723, 1145, 2242, 2087, 578, 2011, 2199, 1596, 1190, 692, 1136, 762, 1120, 1873, 718, 670, 601, 766, 493, 122, 1065, 275, 1635, 1461, 1938, 820]. fractonomials [1024]. fracture [1565, 1119]. fractured [1668]. fractures [251]. fragile [699, 740, 495, 1901, 999, 980]. frame [1778, 266]. frames [2164]. framework [2088, 2226, 2153, 2177, 1112, 326, 1481, 1257, 2032, 873, 2005, 1139]. Franchi [196]. Frankel [347]. Fredholm [1719, 1606, 1733, 1998, 551, 711, 1639, 322, 1737]. Fredholm-type [1733]. Free [496, 1126, 1770, 194, 1940, 71, 1650, 677, 1520, 1534, 168, 1537, 1748, 1481, 1261, 1047, 1539, 224, 126, 1672, 1971, 1535, 753, 2160, 2056]. free-boundary [1650, 1520]. freely [2032]. freely-evolving [2032]. frequency [7, 195, 1354, 496, 443, 1320, 1732]. frequency-constrained [1320]. friction [2259]. Friedkin [498]. front [615]. front-fixing [615]. frying [992]. fuel [123, 7, 6, 2155, 418, 5, 436]. fuel-type-dependent [123]. Full [1900, 438]. Fully [697, 360, 1372, 1440, 839, 776, 2276, 1909, 298]. fully-decoupled [2276]. Fully-discrete [697, 839]. Function [214, 1303, 1026, 793, 1235, 777, 138, 674, 748, 1912, 2039, 860, 1100, 1786, 149, 561, 819, 1326, 831, 1667, 906, 925, 1609, 522, 2093, 771, 408, 368, 706, 1837, 1693, 1317, 60, 2298, 1882, 822, 795, 809, 843, 901, 166]. function-dependent [1882]. function-finite [860]. Functional [1114, 1725, 514, 570, 1044, 1172, 546, 1096, 310, 1449, 1949, 517, 1078, 916, 1099, 1785, 2092, 547, 1810, 1199]. functionally [1537]. functionals [234, 614, 667]. functioning [709]. functions [1695, 190, 2131, 1555, 243, 244, 197, 1780, 1453, 527, 927, 863, 1060, 784, 224, 482, 329, 2278, 1791, 511, 762, 2265, 1187, 1059, 392]. fundamental [1895, 1973, 938]. furnace [1415]. fusion [354]. future [2128]. Fuzzy [1027, 1342, 287, 1320, 872, 1121, 1553, 485, 484, 23, 972, 1117, 1385, 1587, 1031, 332, 1699, 1841, 1962, 2039, 385, 1691, 11, 1037, 2261, 711, 693, 342, 1083, 1982, 716, 974, 2209, 458, 1861, 718, 2093, 481, 1125, 1530, 2052, 1352, 1898, 981, 1065, 1032, 1308, 1062, 1635, 1659, 1730, 2119]. fuzzy-valued [1121].

G [929, 901]. G-Brownian [901].

G-Lyapunov [901]. G. [592]. Ga [352]. Gabor [1187]. gain [1789, 96, 421, 1245]. gains [2003, 1530]. gait [1464]. Galerkin [620, 2272, 722, 1307, 1668, 1189, 49, 1748, 156, 527, 1126, 301, 248, 1261, 57, 483, 1047, 1539, 1392, 1498, 1968, 1600, 1629, 1215, 1979, 2056, 587, 1055, 2287, 41, 933, 1434, 46]. Galerkin-characteristic [1307]. Galois [1819]. game [2131, 1789, 1594, 1963, 1677, 397, 2267]. game-based [1789]. game-inspired [2267]. Games [2171, 936, 1809, 693, 2108, 934, 712]. gamma [344, 366, 723]. GaN [1760]. gang [1365]. Gannet [1006]. Gao [612]. Gardner [1616]. Gas [297, 796, 582, 9, 1003, 430, 2222]. gas-fluid [2222]. Gasoil [1175]. Gauss [1134, 1139]. Gaussian [1235, 363, 1886, 819, 1369]. Gaussian-wavelet-type [819]. GDP [182]. gear [1774, 1488]. Gegenbauer [1586, 1559, 1790, 320, 1525, 771]. Gegenbauer-wavelet [771]. gene [692]. General [1609, 807, 390, 1367, 1258, 748, 1784, 203, 47, 1497, 546, 1541, 357, 401, 1734, 941, 73, 811, 122, 275, 1844]. generalised [1534, 1125]. generalist [570, 1832, 1349, 755]. generalization [75, 1099]. Generalized [698, 1508, 722, 1241, 135, 989, 1637, 2197, 810, 336, 267, 388, 786, 970, 318, 1231, 785, 727, 1239, 1224, 1589, 2172, 1146, 2301, 650, 1259, 299, 1588, 1145, 39, 1289, 1263, 824, 80, 548, 549, 879, 1872, 110, 1170, 1495, 1522, 2018, 831, 1408, 2116, 813, 1474, 340, 1370, 923, 217, 2112, 1614, 2218, 307, 596, 1792, 1630, 1099, 209, 398, 598, 688, 67, 1309, 729, 263, 1321, 1843, 1950, 2264, 976]. generalized-Caputo [1630]. generally [408, 1834]. generated [829, 166, 954]. generating [1942, 45]. Generation [1614, 1693, 1454, 1978, 2065, 1789, 344, 3, 461, 1039, 1823, 905, 807, 1064, 420, 1685, 837, 1750]. generation/absorption [1685]. generator [18, 14, 2232]. generators [1562, 278, 9]. Generic [2232, 1788]. genetic [749, 1692, 1073, 991, 1998, 1190, 21, 675, 499, 834]. geological [2175]. Geometric [2181, 359, 597, 70, 520, 1576, 86, 384, 1924, 1116]. geometrical [658, 1746]. Geometrically [1469]. geometries [1469, 662, 884]. geometry [56, 2220]. GEPSO [307]. GFVM [801]. ghost [149]. Gibbs [1815]. Gilbert [153]. Gilpin [604]. Ginzburg [1689, 669]. GIS [582]. given [1048]. glacier [2136]. glazed [14]. Global [123, 737, 1855, 748, 1946, 2096, 95, 489, 866, 1058, 222, 74, 1087, 1174, 1873, 670, 916, 2206, 101, 1059, 1953, 2170, 1000, 153, 1077, 1347, 285, 1975, 199, 836, 334, 2219, 1160, 712, 1369, 867, 969, 1186, 1964, 2079, 572, 1652, 460, 1412, 1974, 261]. Globally [2098, 159]. glucose [1086, 2168, 758]. glycol [1685]. glyphosate [179]. GNAT [890]. GNAT-embedded [890]. gold [741, 1466]. Golden [742, 1512, 1652]. golden-sine [1652]. Gompertz [723]. Good [172, 705, 173]. goodness [1026]. Goodwin [2218]. Gooseneck [1544]. Gordon [1281, 1254, 2115, 2127, 133, 462, 2159, 1123, 573, 697, 1373, 599, 906, 100, 2069, 373]. Gordon-Schrödinger [1373]. governing [114]. government [1091, 1677, 1243]. governmental [693]. governor [1943]. Gower [570, 1172, 1952, 1786, 965, 1877, 2196, 1892, 914, 953, 1961]. GPGPU [1979]. GPGPU-based [1979]. GPU [77, 1481]. GPUs [242]. graded [799, 1537, 486, 1277, 654, 1410]. gradient [977, 979, 695, 1107, 2128, 557, 1991, 487, 1459, 1869, 1139, 2299, 1266, 1543]. gradient-zeroing [2128]. gradients [2037, 383, 1139]. Graffi [196]. grained [1290]. Grammian [876]. granular [1982, 1352]. grape [2019]. graph [45, 1686]. graphene [515, 425]. Graphical [445, 2301]. graphs [1206, 1438, 589]. gravitational [1660]. gravity [879, 2139, 801]. Greedy [1931, 1206, 2156, 1791]. Greeks [2043]. green [1091, 791, 693, 1105, 1293, 1243, 99, 771, 1139, 1317]. Grey [1813, 1715, 1284, 2294, 1003, 124, 329, 2016, 920, 1160]. Grid [1462, 1958, 1285, 453, 1340, 10, 595, 16, 483, 1674, 423, 4, 1773, 131, 1064, 1732, 53, 887, 420, 435, 515]. grid-connected [435]. grid-forming [1958]. grid-tied [423, 1732]. grids [1042, 1465, 1539, 409, 2194, 1093, 1377]. Grossberg [1362]. ground [63]. groundwater [623, 2017, 1668, 1649].

Group [875, 1184, 783, 279, 1217, 401, 1814, 1864]. groups [1743, 268, 2098]. growing [1925]. growth [1802, 1558, 1057, 2134, 1745, 748, 904, 1786, 1574, 379, 685, 1374, 1102]. guarantee [1671]. guaranteed [2193]. guarantees [1626]. Gudermannian [2302]. guide [142]. guided [2019]. guideline [173]. gyrostat [285].

H [779, 951, 643, 260]. Haar [1079, 1454, 842, 161, 1165, 2264]. habitat [225, 1099, 315]. habitats [1185]. Hadamard [2105, 1867, 1311, 1434]. HAGTM [2243]. Haldane [1555, 1810]. Hall [951]. Halton [926]. HAM [44, 671, 1703]. Hamiltonian [1799, 1134, 1498, 2062, 1467]. Hamiltonians [2249]. Hammerstein [2026, 1719, 2225, 1218]. handling [1905]. harbor [86]. hardback [592]. hardware [416, 228]. harmful [604]. Harmonic [1661, 153, 1097, 1026, 437, 439, 504, 1410, 895, 683, 821]. Harris [679, 1100, 1415, 1063]. Hartmann [1377]. harvest [1877]. harvested [1363]. harvesting [1172, 577, 1952, 690, 1914, 1865, 1949, 1115, 1736, 1533, 2293, 1961]. Hashimoto [1970]. having [897]. Hawk [679]. hawks [1100, 1415]. HBV [674, 1796]. HCNNs [531]. HCV [1174, 1499]. headlamps [70]. healthcare [1663, 1256]. heart [736]. Heat [1685, 536, 116, 1454, 918, 796, 1137, 582, 660, 175, 992, 358, 9, 1875, 948, 1263, 1039, 534, 492, 1854, 523, 733, 1703, 283, 8, 905, 235, 1381, 347, 2222, 153]. heat-conducting [796]. heated [889, 1337, 534, 905]. heating [311]. heavy [2023, 359, 55, 2266, 1856]. heavy-tailed [2023, 359, 55, 1856]. Hedging [148]. Hegselmann [498]. Heisenberg [1791]. Helmholtz [1624, 778, 650, 1089, 810, 2245]. helminth [276]. help [1185]. hemodynamics [590, 61]. Hepatitis [1751, 730, 1146, 1519, 1371, 672]. herbivore [2270, 547]. herd [87, 2293, 1961]. Hermite [838, 1787, 1235, 1969, 780, 1537, 202, 552, 551, 1672, 814]. Hermite-type [1537, 1672]. Hermitian [768]. Herschel [2140]. hesitant [1691, 2093, 1138, 1587]. Heston [1806, 242, 39, 248, 1103]. heterogeneity [1850, 862, 1396, 1621]. heterogeneous [2303, 1066, 2191, 44, 1788, 1913, 2030, 614, 1696, 136, 414, 1407, 1093, 1800, 1290, 1368, 2034, 1979, 1864, 1890]. heterosexual [2151]. Heuristic [286, 1371, 1930, 18, 1242, 519, 1293, 803, 2302, 1412]. hidden [1547, 1861]. hiding [2228]. Hierarchical [2225, 1690, 2202]. High [160, 888, 471, 1262, 1387, 1152, 443, 865, 2162, 815, 1718, 1494, 889, 1473, 558, 390, 1483, 1987, 737, 2059, 1534, 1386, 1899, 624, 892, 1788, 780, 1566, 49, 1090, 2254, 704, 1306, 103, 1912, 900, 1095, 1134, 1495, 2085, 166, 1768, 1474, 1041, 1760, 535, 1903, 1377, 372, 264, 1814, 1321, 1950, 2245]. high-dimensional [2059, 1386, 1090, 1095]. high-intensity [889]. High-order [160, 888, 471, 1262, 815, 1718, 1494, 1473, 390, 1483, 737, 1534, 1899, 624, 1788, 1566, 49, 1306, 103, 1134, 1495, 1474, 1041, 535, 372, 1814, 1321, 1950, 2245]. High-performance [1152, 1987]. Higher [620, 77, 125, 842, 229, 1670, 998, 1983, 486, 2085, 831, 2250, 1448, 1165, 1561, 1410, 1155]. Higher-order [1670, 2085, 831, 1448, 1410]. highest [1028]. Highly [302, 1561, 1402, 220, 2030, 1424, 814, 1944]. highway [70]. Hilbert [1046, 446]. Hilfer [2243]. Hilliard [1452, 1795, 1808, 319, 884, 1954, 2287, 1718, 2220]. Hindmarsh [974]. Hirota [2249, 2301, 1309]. historical [29]. hitting [1945]. HIV [922, 1283, 1367, 1745, 1094, 1720, 1838, 382, 2083, 1883, 2151, 1803, 332, 866, 2029, 1664, 1302, 1514, 1379, 565, 603, 308, 649, 601, 584, 2154, 1822, 1420, 1953, 1638]. HIV-HTLV-I [1720]. HIV/AIDS [1883, 2151, 1664, 1302, 603, 308, 649, 2154, 1420]. HIV/SARS [2083]. HIV/SARS-CoV-2 [2083]. Hodgkin [2181]. hodograph [2118, 2164]. Holder [1080, 1519]. holes [2297]. Holling [570, 825, 875, 1044, 2061, 1115, 916, 1099, 706, 1798, 2086]. Holling-type [916]. holonomic [1387, 83]. Holt [1515]. Homann [79]. home [1173]. homoclinic [1198]. Homogeneity [514]. Homogeneous [44, 2131, 796, 2229]. homogenization [631]. homology [501, 1723]. Homotopy [407, 1084, 775, 677, 1609, 2014, 2290, 2264, 105]. Honey [689]. Hopf [962, 1708, 2031, 2061, 2257, 1151, 138, 829, 1451, 1157, 802, 1892, 2218, 766, 315, 403, 817]. Hopf-bifurcation [2031, 1892]. Hopf-Hopf [2061]. Hopfield [737, 819, 1596]. hopscotch [1250]. Horioka [310]. horizon [1586, 1809, 934, 1820]. horizontal [2037, 204, 227, 753, 537]. horizontally [266]. hospital [1997, 1244]. hospitalization [1997]. host [922, 2083, 794, 1059]. Hot [1686]. Hough [29, 1924]. household [1773]. hp} [1729]. hp}-version [1729]. HPM [915]. HPV [1955]. HS [1266]. HTF [707]. HTLV [1720, 334]. HTLV-I [334]. Huang [446]. Human [256, 1906, 734, 476, 250, 1185, 1850, 2267]. human-caused [1906]. humans [862]. hunger [712]. hunting [1054, 1613, 1949, 691, 1533]. Huxley [1289, 2181]. HVDC [1958, 306]. Hybrid [1197, 472, 326, 276, 1216, 846, 977, 979, 1606, 643, 1303, 442, 890, 14, 623, 1265, 1620, 33, 1118, 557, 505, 1716, 1551, 487, 1145, 1687, 1085, 779, 1359, 751, 1670, 13, 1149, 1854, 1327, 762, 733, 1619, 873, 951, 35, 436, 688, 2299, 750, 502, 2298, 2287, 1266, 61, 1967]. hybridizable [1668]. hybridization [442, 803]. hydra [2255]. hydrate [685, 2222]. hydraulic [58, 2157, 251]. hydrodesulfurization [1175]. hydrodynamic [2276, 204, 473, 2150]. hydrodynamics [1853, 1812, 200]. hydrogen [787]. hydrophobic [355]. hydropower [3]. hydrostatic [1660]. Hyers [1311]. hyperbola [1799]. hyperbolic [1402, 1134, 342, 1182]. hyperbolic-NILT [1182]. hyperchaotic [2257, 1247, 2200, 1183]. hyperinfectivity [1059]. hyperinterpolation [640]. hyperspectral [2005]. hypertension [590]. hyperthermia [594]. hypothesis [1056]. hysteresis [2061, 631]. hysteretic [201].

IAP [95]. IAP-DCP [95]. ICMMCMSE [1749]. ICMMCMSE-2020 [1749]. IDA [2231]. IDA-PBC [2231]. identifiability [894]. Identification [687, 626, 1458, 2180, 2225, 614, 1218, 421, 1766, 1506, 163, 416, 1488, 418, 630, 144, 857, 1032, 2283, 2026]. ideological [1883]. IEFG [1191]. II [570, 1044, 991, 916, 706]. III [2061]. ill [695, 2035]. ill-posed [695, 2035]. IMACS [51, 65, 90, 118, 140, 151, 187, 212, 254, 293, 50, 64, 89, 117, 139, 150, 169, 186, 211, 253, 292, 313, 338, 375, 432, 456, 508, 525, 607, 633, 637, 312, 337, 374, 431, 455, 507, 524, 567, 606, 632, 681, 714, 726, 789, 805, 827, 851, 882, 943, 985, 1019, 680, 713, 725, 788, 804, 826, 850, 881, 909, 942, 984, 1018, 1069, 1109, 1130, 1131, 1201, 1233, 1270, 1295, 1329, 1357, 1389, 1418, 1443, 1068, 1108, 1200, 1232, 1269, 1328, 1356, 1390, 1417]. IMACS [1442, 1479, 1486, 1528, 1569, 1604, 1644, 1681, 1706, 1754, 1777, 1848, 1888, 1478, 1485, 1527, 1568, 1603, 1643, 1680, 1705, 1753, 1776, 1847, 1887, 1922, 1935, 2009, 2047, 2072, 2101, 2133, 2146, 2183, 2211, 2234, 1921, 1934, 2008, 2046, 2071, 2100, 2132, 2145, 2182, 2210, 2233, 410, 2268, 170]. Image [2228, 2094, 241, 2091, 973, 198, 1280, 2247, 2039, 1076, 2238, 2284, 291, 269, 2139, 835, 30, 1819, 1816, 1075, 2179, 545, 1842, 1183, 2213, 2267, 503, 864, 1220, 1583, 1894]. images [1455, 2247, 81, 2005, 1540]. imaging [208, 1673]. IMEX [2192, 1399]. imitation [2156]. immeasurable [27]. Immersed [480, 236, 53, 1316]. immersion [992]. immigration [1515, 2106, 2001]. immigrations [1227]. immiscible [1395, 2143]. immobile [1149]. immune [1367, 1054, 2113, 262, 603, 2117, 601, 402]. immunity [748, 401, 1563, 1499]. immunotherapy [1710, 1514]. Impact [664, 668, 1172, 1722, 2230, 2196, 69, 928, 619, 1365, 1584, 1925, 1363, 2121, 628, 1159, 2242, 1575, 214, 1492, 1322, 268, 939, 420, 2293, 2204, 1769, 1533]. Impacts [3, 1597, 1249, 691]. impedance [421, 418, 435]. implement [416, 925, 34]. Implementation [1462, 1394, 77, 1545, 242, 1783, 469, 11, 25, 1027, 1829, 228, 1979, 1894, 1080, 198]. implementing [2061, 358]. implications [2270, 2165, 1597]. implicit [1878, 1403, 1566, 49, 2248, 203, 1513, 319, 2253, 387, 811, 2069, 1361, 1236, 1316, 815]. implied [1144]. Importance [600, 2280]. imprecise [690]. imprecisely [1475]. Improved [2033, 1660, 1100, 1373, 1449, 1656, 1535, 158, 994, 1088, 271, 1599, 2119, 1324, 1118, 1239, 42, 578, 1359, 342, 1160, 1369, 418, 1429, 1293, 670, 112, 264, 1416, 218, 1065, 1399]. Improvement [95, 534, 23, 1035, 229, 35, 1364]. improvements [1557]. Improving [1483, 1735, 1652]. Impulse [391]. Impulse-dependent [391]. impulses [1596, 98]. Impulsive [974, 604, 115, 2289, 382, 2123, 1286, 222, 1096, 1040, 1124, 1579, 391, 1311, 2296, 1338, 98, 1601, 1904, 701, 1012, 846, 1000]. imputation [600]. in-time [1646]. in-wheel [20]. inaccurate [675]. INAR [1431, 1176]. inception [645]. incidence [1482, 1367, 1745, 1720, 833, 748, 1796, 401, 937, 1059, 67, 1499]. incidences [1174]. incipient [1769]. inclined [677, 1687, 1466]. including [889, 450, 603]. inclusion [772]. inclusions [1046, 1714]. incommensurate [2199]. incomplete [847, 1214]. incompressible [1932, 49, 1346, 1672, 408, 2287, 1372]. incorporating [1839, 1288, 1099]. Increasing [1739, 168, 896]. Incremental [288]. incubation [2144, 69]. independent [2280, 2274, 2092]. index [1171]. indexed [2092]. India [1302, 1510, 2154]. Indian [1955]. indications [19]. indicator [1187]. indices [80]. indirect [32]. individual [1908, 2058, 2204]. Indonesia [1838, 1826]. induced [668, 2061, 833, 480, 577, 1786, 1001, 1021, 2152, 1951, 1595, 128, 1492, 802, 1349, 1099, 706, 1423, 272, 2223]. induction [31, 18, 19, 447, 821]. inductionless [1476]. industrial [1175]. industrial-scale [1175]. inequalities [702, 784, 1081]. inequality [1508, 1991, 104, 304, 106, 670]. inertia [560]. Inertial [772, 702, 1347, 1593, 966, 162, 1136, 2079, 1817, 1000]. inertial-type [1136]. infected [2110, 856, 1094, 866, 2029]. infection [922, 912, 1283, 1077, 1184, 1720, 748, 382, 2083, 334, 1230, 1955, 1519, 1514, 603, 2049, 1274, 2206, 584, 939, 1822, 1420]. infection-age [1077]. infections [1884, 1563, 603, 276, 1428]. Infectious [746, 592, 1151, 783, 1896, 1397, 1575, 857, 2160, 1425, 1621]. infective [746, 939, 572]. infectives [308]. Inference [688, 23, 2011, 1964]. Inferential [1188]. infinite [1993, 1550, 2067, 1586, 1696, 1809, 1286, 357, 934, 511, 1311, 1634]. infinite-horizon [1809, 934]. infinitely [408]. inflated [1176]. inflection [857]. Influence [2035, 2108, 993, 787, 444, 672, 1751, 1688, 1884, 237, 651, 565, 1927, 2255]. influenced [1919]. influences [1685]. Influential [179]. influenza [1184, 399]. Inform [592]. Information [396, 194, 147, 1677, 1214, 586]. information-based [194]. informed [1821, 2224, 1411, 2241, 2045]. infrastructure [2209]. infusion [2168].

InGaN [15]. Ingersoll [1967]. inhibitor [550]. inhibitory [1698]. inhomogeneous [1161, 1299, 938]. Initial [1083, 1860, 129, 1867, 1982, 813, 1432]. initialization [1815]. injection [443]. injection-based [443]. injections [2124, 1435]. inlet [776]. Inner [1943, 1475]. innovation [553, 1688]. Innovative [189, 2278, 2143, 2016, 412, 1280]. inpainting [1808]. Input [1692, 93, 31, 123, 297, 2225, 1260, 1031, 966, 2294, 820, 395, 1943]. input-output [2294]. Input-to-state [1692, 93, 966]. inputs [27, 1793]. inrush [251]. insect [1552]. insertion [1290]. inserts [116]. Insight [227]. Insights [741, 1470, 2301, 2002, 1256]. inspection [742]. inspiration [645]. inspired [1544, 963, 935, 656, 1549, 2109, 460, 272, 2097, 2267]. instability [1786, 2174, 1595, 2143]. instantaneous [1124, 2198]. insulation [582]. insulin [1086, 2168]. insurance [2177]. integer [2121, 1512, 1247, 982, 1216, 2221]. integer-valued [2221]. integrability [556]. integrable [1640, 1956]. Integral [1060, 2273, 1695, 1079, 1719, 1606, 132, 1313, 2203, 2059, 1135, 1998, 1586, 234, 605, 505, 1841, 551, 1268, 957, 578, 304, 903, 1424, 335, 1728, 322, 106, 767, 670, 771, 1061, 1222, 1253, 1737, 814, 1944, 27]. integral-algebraic [1135]. integrals [511, 1177]. Integrated [413, 576, 842, 1663]. Integrating [1763, 1536, 2213, 2135, 2190]. Integration [895, 675, 1324, 768, 424, 132, 2114, 422, 1191, 400, 1619, 1637, 392, 2290, 1467, 1578, 1111]. integrator [2115, 2112, 2127]. integrators [77, 1134, 274]. integro [2036, 870, 1733, 1401, 325, 2289, 2042, 1507, 711, 1872, 1902, 1392, 1674, 1639, 1573, 2112, 1829, 648, 1843, 1098, 1150]. integro-differential [2036, 870, 1733, 325, 2289, 2042, 1507, 711, 1872, 1902, 1674, 1639, 1573, 2112, 1829, 648, 1843, 1098, 1150]. integro-partial [1392]. intelligence [1004, 576]. Intelligent [716, 954, 1584, 1409, 812, 685, 584]. intense [1818]. intensity [889, 1925, 1984]. inter [34, 24, 821]. inter-turn [821]. interacting [1994]. Interaction [2015, 643, 498, 625, 1627, 2113, 146, 1670, 69, 801, 112]. interactions [664, 1965, 2078, 41]. interactive [2027]. interchange [385, 326]. interconnected [2051, 1823, 1898, 703]. interconnection [2157]. interconnects [656]. interest [39, 1984]. interface [1029, 360, 1907, 112, 53, 1564, 263, 1316]. interfaces [160, 1545, 2076]. interfacial [1608]. interference [875, 428]. interferon [1499]. interior [1581, 1668, 897, 1392, 563, 1341, 1166, 1985, 502]. interleaved [438, 5]. interleaving [440]. intermediate [1002]. intermittent [1593, 2281, 1533]. internal [414, 1366, 905]. International [510, 609, 2057, 1143, 148, 1683, 1749]. Interplay [586, 2000]. interpolant [191, 518]. interpolants [289, 2040]. interpolating [1719, 1811, 1717]. interpolation [1695, 2036, 190, 132, 769, 1787, 1712, 550, 1287, 1998, 888, 234, 1050, 2164, 1300, 935, 16, 551, 1148, 1576, 1474, 641, 1619, 1253, 1599, 835]. interpolatory [552]. interpretable [113]. interruptible [2080]. intersection [193, 2007]. Interval [1553, 332, 666, 1962, 418, 1120, 1898, 930, 980]. interval-valued [1962]. intervention [1243]. interventions [693]. intra [1075]. intracellular [1367, 1745, 2083]. intraguild [2084]. intrahepatic [1499]. intraspecific [668, 230, 755]. Introducing [333]. introduction [1138]. Intuitionistic [1385, 1962]. Invariant [258, 1896]. invasion [2110, 2002, 1864]. invasive [1363, 2170, 446]. inventory [300, 136, 2198, 1195, 1820]. Inverse [2099, 102, 1567, 1733, 1287, 2128, 666, 1326, 180, 1381, 2090, 975, 2138]. inversion [494, 843]. inverted [1308]. inverter [447, 1772]. inverters [439, 1778]. investigate [1185]. Investigating [2017, 1105, 1591, 1879, 60]. Investigation [1823, 20, 1611, 968, 1279, 2121, 1668, 2223, 1818, 1504, 59, 1805, 733, 1320, 1427]. investment [731, 99]. investments [1293]. involved [604]. involving [1545, 785, 992, 1497, 1040, 1065, 370]. ion [419, 413, 416, 415]. iron [1415]. irradiation [1591]. irregular [1493, 646, 945, 1875, 239, 539, 739]. irrigation [2088]. IS-LM-AS [1526]. ISBN [592]. Ising [199]. iso [445]. iso-performance [445]. Isogeometric [1830, 1932, 43, 2130]. isolated [611, 1732]. isolation [1884]. isothermal [2129]. Issue [510, 2, 172, 434, 412, 609]. issues [198]. Itô [505, 1061, 1222, 1697]. Itô-Volterra [505, 1222]. Italian [1941]. Italy [2110, 744]. items [1231]. iterated [1206, 1693]. Iteration [2171, 1625, 216, 1470, 1556, 2201, 1614, 1052, 1432, 891, 1101, 1974, 1501]. iterations [1282]. Iterative [1272, 772, 1016, 195, 1940, 1983, 1739, 777, 2156, 2225, 200, 260, 440, 915, 9, 561, 479, 544, 1576, 504, 934, 288, 1416, 1870, 1618, 1673, 2245]. Iteratively [1642]. Ito [2301, 1309]. IV [875, 1798].

jackal [1512]. Jacobi [1728, 1024, 408]. Jafarimoghaddam [532]. January [296, 661, 1022, 1446, 1891, 2304]. JBS [426]. Jeffrey [461]. Jenkins [2026]. jets [1209]. Jimbo [876]. Johnsen [498]. join [1441]. joining [1334]. joint [1116, 2193]. Julia [1625, 1614]. July [143, 459, 830, 1273, 1647, 2104]. Jump [1800, 1861, 1926, 1692, 615, 699, 2067, 605, 1593, 1048, 1582, 1513, 369, 481, 853, 1901, 387, 1984, 1578]. jump-diffusion [2067, 1984, 1578]. jump-extended [387]. jumped [355]. jumps [1809, 1915, 357, 1859]. June [121, 808, 1237, 1607, 2075]. Jungck [1470, 807].

Kac [1990]. Kadomtsev

[328, 556, 750, 729]. Kahan [2249]. Kalecki [731]. Kalman [2155, 675]. Kampar [1826]. Kannan [1693]. Kansa [2138, 1168, 371, 923]. Kármán [1548, 1114]. Karush [1121]. Kaup [824, 2286, 1125]. Kawahara [1616, 302, 756]. Kawarada [2105]. Kaziranga [225]. KdV [964, 1262, 2190]. KdV-like [964]. Keller [2240, 2129, 2077]. kelp [2293]. Kepler [2219]. kernel [1719, 1534, 1905, 1716, 1874, 1370, 771, 887, 648, 795, 1843, 814, 1944, 756]. kernel-based [1716]. kernels [1606, 2036, 1669, 405, 765, 1150]. kidney [1037]. kiln [675]. Kimura [2249]. kind [1606, 1112, 1729, 1918, 551, 903, 1424, 1737]. kinds [603]. kinematic [1287, 474]. kinematics [2128]. Kinetic [1928, 1871, 1213, 932, 1844]. kinetics [1129, 513, 1323, 685]. kissing [236]. KKT [1121]. Klein [462, 1123, 1281, 697, 1373, 1254, 2115, 599, 100, 2069, 2127]. knapsack [1590]. knot [1411]. knots [519]. knowledge [960, 1686]. Korf [379]. Korteweg [1033, 247, 1167, 1801, 1215, 581, 2290, 1950]. Kotera [1023]. Koyna [801]. KP [566]. Krause [498]. Kronecker [84]. Kuala [774]. Kuhn [1121]. Kukles [2229]. Kupershmidt [824, 1125]. Kuramoto [1612]. Kutta [1307, 1574, 158, 1494, 1262, 277]. Kutta-Nyström [1494]. Kuznetsov [720, 279]. kW [417]. kW-class [417].

L1

[2217, 2285, 799, 548, 1142, 1630, 2063, 1524]. L1-Predictor-Corrector [1630]. L1/finite [2217]. L1/LDG [548]. L1/spectral [2285]. L3 [1156]. Label [1684, 1691]. labor [2152]. LabVIEW [28]. LabVIEW-based [28].

Ladik [1406]. lag [594, 752, 992, 1429, 1338, 2161]. lagoon [1879]. Lagrange [1897, 1627, 697, 377, 835, 1900, 1101]. Lagrange-quadratic [377]. Lagrangian [892, 1081]. lags [94]. Laguerre [1283, 771]. Laguerre-wavelet [771]. Lakshmanan [2099]. laminar [590]. land [2005]. Landau [1689, 530, 669, 153]. landfill [468]. landing [183]. lane [757, 998, 1918, 1078, 120, 576, 1297, 1165]. languages [45]. Laplace [844, 1424, 1350, 1933, 2014]. Laplacian [1194, 627, 2021, 2059, 1193, 2062, 46, 466]. Lapwood [380]. Large [2014, 1743, 724, 687, 260, 818, 1429, 703, 165]. large-scale [260, 703]. Larmor [1189]. laser [889, 1591]. LASSO [310]. latency [1831, 1638]. latent [1803, 1741, 283, 1822]. latently [866, 2029]. lateral [2096]. Lattice [311, 1875, 1009, 972, 1398, 2177, 1067, 705, 957, 142, 1595, 2140, 1608, 352, 355, 2263, 2227, 859, 1439, 58]. lattice-based [2177]. lattices [1844]. Laudation [610]. law [1126, 57, 1399]. laws [79, 599, 398, 342]. Lax [575]. layer [619, 1547, 1091, 318, 660, 1696, 704, 1465, 917, 47, 1263, 2053, 1506, 1670, 1341, 1548, 1468, 502]. layered [2201]. layers [897, 563, 21, 1166, 1939]. layout [982]. LBIE [1400]. LBM [582]. LC [1778]. LC-filtered [1778]. LDG [548]. leachate [468]. Leader [667, 1928, 1021, 1415, 2256, 949]. Leader-following [667, 1021, 2256]. Leadership [2097, 555]. leading [1413]. leak [1655]. leakage [485, 391, 111, 1065, 1817]. leakages [1404]. Learning [1761, 1516, 2241, 1493, 1014, 1821, 1503, 501, 1536, 1763, 199, 2156, 1824, 2055, 1272, 898, 1852, 1907, 1005, 1160, 742, 1113, 873, 454, 1186, 2244, 1592, 1652, 1276, 1248, 2188, 1599, 1688, 1966, 512, 732, 848]. Learning-based [1761, 454, 1592]. Least [1246, 2269, 1784, 1167, 360, 2139, 322, 1139]. least-squares [360, 1139]. Leffler [793, 405, 846]. left [2128, 763]. left-truncated [763]. leg [5]. Legendre [1180, 956, 931, 1729, 1591, 283, 738, 891, 1339, 759]. Leishmaniasis [538]. lemmas [1301]. length [1218]. Leslie [1952, 570, 1172, 1786, 965, 1877, 2196, 1892, 914, 1115, 1798, 953, 1961, 1372]. Leslie-type [1798]. Level [1726, 330, 1813, 791, 185, 1945, 1048, 2130, 1119, 2000, 1773, 287, 1106, 368, 1464, 249, 503]. level-dependent [1945]. levels [422]. Levenberg [1103]. Levèque [318]. levy [361, 363, 1090, 357, 1990, 1909, 1435, 1467]. levy-assisted [361]. Lewis [181]. LFP [419]. LG [1044]. LG-Holling [1044]. Li [419, 416, 415]. Li-ion [419, 416, 415]. library [1868]. lid [2140, 686]. lid-driven [686]. Lidstone [521]. Lidstone-based [521]. Lie [918, 2098, 1814]. Lie-group [1814]. life [1026, 2177, 2193]. lifecycle [1857]. lifespan [1993]. lifetime [2155]. Lifshitz [153]. light [889]. lightning [32]. like [698, 1787, 1331, 1185, 1869, 272, 964]. likelihood [1806, 2158, 2221, 2026]. limb [718, 1464]. Limit [1799, 2229, 1999, 1161, 1928, 1701, 757, 2115]. limitations [894]. limited [258, 1425, 1738]. limiter [342, 2168]. limiter-based [2168]. limiting [2244]. Lindley [1375]. line [328, 1999, 1007, 415, 288]. line-doublets [1007]. Linear [2053, 1700, 1917, 1440, 1606, 870, 575, 2038, 129, 2006, 2103, 1585, 849, 731, 1241, 1354, 2033, 541, 888, 1871, 1260, 1172, 2042, 1306, 203, 520, 635, 1588, 844, 2166, 1687, 1350, 110, 404, 751, 1158, 1516, 204, 1190, 690, 982, 2095, 1914, 207, 855, 217, 106, 1498, 1703, 1609, 1804, 1216, 1222, 811, 298, 1253, 676, 1299, 53, 1476, 1835, 1954, 2185, 533, 463, 899, 1372, 517]. Linear-Phase-Type [517]. Linearising [1657]. linearity [952]. linearity-preserving [952]. linearization [775]. linearization-based [775]. linearized [1452, 925, 2292]. Linearly [2069, 1566, 1236]. lines [958, 428, 75, 32]. link [1593, 423]. linked [700]. Liouville [1180, 2262, 1224, 1589, 1326, 1522, 490, 886, 899]. lipase [395]. liquid [643, 1415, 787, 543]. liquids [2143]. lithium [413]. LM [1526, 2026]. LMAPS [244]. LMFS [477]. LMI [967, 604]. load [182, 2232, 2230, 452, 1658, 444]. loading [619, 1010]. Local [1926, 2217, 242, 860, 1532, 958, 753, 1181, 587, 1055, 282, 190, 2272, 325, 823, 836, 335, 1964, 309, 2228]. Localised [754]. Localized [243, 1438, 1179, 2215, 1240, 1973, 938]. locally [12, 906]. locating [975]. location [1173, 300, 1195]. location-inventory [1195]. lockdown [2230]. logarithmic [1719, 1701]. logic [385, 1032]. logistic [1802, 1650, 748, 1574, 1327, 115, 1785, 1374, 795, 1745]. logistics [1073]. Long [2020, 134, 956, 842, 1808, 2192, 408, 177, 1215, 760, 1245, 2127, 1950, 1767]. long-range [177]. long-term [760]. long-time [1808, 2127]. Longitudinal [1488, 408]. lookback [1265]. loop [421, 930, 1943]. looped [667, 1449]. looped-functional [1449]. Lorentz [1129, 741, 1687]. Lorenz [1409, 1092]. loss [2017, 1977, 2022, 397, 1008, 446]. loss-averse [397, 1008]. losses [1856]. Lotka [1331, 2020, 280, 1399]. love [1157, 132, 1056]. Low [1744, 1925, 1787, 1354, 686, 444]. low-degree [1787]. low-intensity [1925]. lower [1257, 1464]. lowest [282]. LPV [447, 1488]. LS [1516]. LS-SVR [1516]. LSTM [645]. LTBI [1839]. LTNE [753]. LTR [1094]. LTV [996]. lubrication [204]. Lucas [970, 1061, 1011]. lump [566, 845, 1240, 1170, 556]. lumped [1954]. Lumpur [774]. Lundqvist [379]. lung [250]. Lur’e [231, 1025, 1744]. Lyapunov [1876, 718, 1384, 717, 901]. Lyapunov-based [718]. lymphocytes [334, 1720].

M [1368, 349, 929, 1368]. M/G/1 [929]. M/M/2 [1368]. M1 [1871]. machine [31, 1014, 1821, 501, 1536, 1824, 1907, 447, 1291, 458, 1113, 873, 1342, 2244, 444, 445, 1834, 821, 512]. machines [442, 1769, 2055, 19, 443]. machining [929, 2216]. MACMAS [510]. macro [1598]. macro-elements [1598]. macroeconomic [1526]. macrophages [1803]. macroscopic [1928, 2107]. magnet [1770, 1768, 443, 444, 24]. magnet-free [1770]. magnetic [908, 1140, 594, 1657, 677, 1774, 1977, 444, 1685, 837, 1750]. magnetisation [1657]. magnetization [446]. magnetized [932]. magneto [279]. magneto-plasmas [279]. magnetoelastic [1696]. magnetohydrodynamic [1476]. magnetohydrodynamics [917]. Magnus [1213]. Mahony [1474, 729]. maintenance [2022]. making [1037, 1217, 2093]. malaria [1646, 734, 1511, 1428, 1980]. Malaysia [774]. malfunctioning [2066]. Malliavin [1868]. malware [655, 1540]. mammography [1004]. management [1771, 1775, 7, 1763, 453, 1767, 13, 1642, 2058, 1293, 35, 1732]. manager [33]. Mandelbrot [1625, 1614, 1282]. MANFIS [948]. Manganese [487]. manipulator [1287, 1560]. Mann [1282]. manufacturing [1091]. map [2275, 2039, 2200, 1076, 2179, 1844, 153, 1220]. Mapping [1505, 532, 1532]. maps [291, 2249, 1614, 396]. Marangoni [800]. March [54, 341, 728, 1133, 1502, 1959]. marching [1797, 1089]. Marcus [1467]. Marine [1610]. market [652, 1457, 2142, 1863, 2034]. marketing [1946]. markets [1013, 1457]. Markov [1861, 699, 1593, 1809, 1457, 340, 2032, 2282]. Markov-switching [1457]. Markovian [1692, 382, 136, 1551, 1506, 1541, 357, 369, 481, 853, 1901, 2066, 226]. Marquardt [1103]. Martin [1667, 1949]. Maruyama [558, 226]. mass [513, 992, 948, 1039, 246, 1954]. mass-preserving [246]. massively [1394]. massless [2115]. Master [1562, 1086, 231, 872]. matching [1171]. MATCOM [510, 609]. Material [630, 643, 1129, 1770, 2130, 1726]. materials [28, 614, 1119, 408, 1056, 412]. Math [348, 1752, 389, 1751, 1750]. Mathematical [510, 468, 609, 439, 1545, 1683, 1287, 541, 1395, 744, 1211, 334, 1159, 1034, 221, 1343, 565, 2057, 858, 1102, 1420, 1638, 538, 388, 1227, 1314, 273, 832, 1279, 736, 399, 1414, 1828, 1970, 554, 1597, 1622, 1229, 418, 2000, 2302, 1980, 1256, 1015, 1749, 1511]. Mathieu [364]. MATLAB [2134]. MATLAB(R) [1868]. matrices [768, 2138, 847, 135, 591, 1011, 1275]. Matrix [2091, 1827, 1834, 972, 1401, 1780, 1565, 1481, 126, 1665, 1525, 204, 759, 1078, 1384, 2098, 596, 494, 717, 843]. Matrix-augmentation [1834]. matrix-free [1481]. Maturation [1349, 1322, 1678]. MAV [2027]. max [1546]. Maximal [857]. maximization [2221]. maximizing [1008]. Maximum [2026, 1602, 1226, 1920, 1766, 2158, 990]. maximum-principle-satisfying [1602]. Maxwell [1097, 155, 660, 142, 1325, 534, 1353]. May [792, 1204, 1572, 2050]. MCDM [1691]. MCMC [359, 1626]. MDA [1252]. mean [1026, 991, 1893, 1794, 2124, 2197, 1578, 809, 1676]. Mean-curvature [1676]. mean-entropy [1794]. mean-field [2197]. mean-reverting [809]. mean-semiabsolute [991]. means [1303, 663, 1177, 425, 2119]. measles [1669, 1665]. measure [1542, 2280, 1896, 1699, 304, 258, 162, 30, 1012]. measurement [1014, 107, 354, 2026]. measurement-data [2026]. measurements [354, 1985]. measures [1632, 1884, 1587, 2172, 1785, 1510]. Mechanic [185]. mechanical [1010, 1591, 812, 1548, 1205]. mechanically [611]. mechanics [250, 1119, 1290]. Mechanism [251, 2111, 1690, 947, 1359, 534, 1104, 1671, 2097]. mechanisms [588, 1309]. mechanotherapy [1925]. media [2303, 1490, 722, 2031, 1307, 58, 2266, 2030, 252, 1696, 1159, 1575, 86, 329, 1041, 1407, 710, 1071, 1422, 1523, 2271, 1353, 809]. mediated [1613]. medical [972, 1816, 1176, 1425]. Mediterranean [1742]. medium [1558, 236, 534, 2201, 1732, 572]. medium-frequency [1732]. Mellin [700]. Mellitus [1086]. melting [352, 800]. membership [1882]. membrane [2155]. memories [514, 517]. Memory [973, 889, 785, 252, 1359, 357, 205, 1861, 2256, 1678, 675, 403, 1423, 1015, 1767]. memory-dependent [889]. memory-induced [1423]. memristive [485, 740, 1698, 1894]. memristor [966, 2200, 1301, 2109]. memristor-based [1301]. memristor-inspired [2109]. Meningitis [1074]. Menten [1323]. mergers [1688]. MERS [1211]. MERS-corona [1211]. mesh [133, 2303, 2065, 2217, 799, 2269, 1534, 76, 1537, 836, 486, 900, 2082, 1277, 1648, 1672, 1676, 1939]. mesh-based [836]. mesh-free [1534, 1537, 1672]. meshes [2148, 278, 952, 663, 1903, 1410, 739, 372, 1150]. Meshfree [1355, 2258, 1407, 1119, 1931]. meshless [1493, 1606, 243, 764, 1010, 1807, 1400, 1716, 2215, 836, 1167, 467, 1608, 641, 1222]. meta [1242, 1293, 803, 1412]. meta-heuristic [1242, 1293, 803, 1412]. metabolism [1086]. metaheuristic [1553, 782, 1171, 1815, 1908, 306, 963, 256, 689, 1334, 2198, 1312, 1006, 2188, 2066]. metaheuristic-based [306]. metaheuristic-driven [2188]. metal [1439, 2266]. metallic [1395]. metamodel [747]. metastatic [1057, 2134]. metasurface [848]. metering [385]. methane [685]. Method [472, 838, 1494, 1402, 979, 133, 538, 772, 1180, 1860, 1440, 1079, 1473, 2068, 388, 1484, 1163, 775, 390, 620, 1140, 861, 1097, 1333, 1454, 1016, 890, 220, 132, 956, 575, 1490, 437, 2272, 1238, 2136, 623, 194, 842, 1029, 1209, 1336, 1347, 1983, 84, 114, 1285, 1895, 2203, 2269, 618, 2275, 722, 1534, 2159, 615, 1171, 776, 859, 694, 1439, 777, 1436, 1721, 145, 1123, 1046, 1554, 550, 880, 1812, 343, 892, 1307, 1742, 407, 1239, 2180, 1559, 1790, 58, 888, 1189, 127, 200, 349, 695, 1447, 1173, 56, 236, 764, 1182, 1010, 1426, 76, 1729, 2126, 1807]. method [49, 2248, 232, 505, 1084, 1300, 1507, 1991, 542, 1400, 1306, 573, 1748, 2258, 1920, 1716, 2215, 1463, 1453, 156, 1126, 1962, 1505, 1591, 1465, 841, 1038, 1649, 1218, 702, 602, 2171, 146, 1995, 2129, 371, 81, 1612, 486, 1251, 844, 11, 1167, 1289, 467, 360, 711, 900, 1261, 2130, 1003, 1574, 2288, 57, 239, 548, 1047, 1089, 1154, 1577, 2082, 1973, 1902, 149, 224, 163, 142, 561, 482, 1326, 1867, 377, 751, 1408, 1516, 1392, 126, 1325, 1648, 1149, 1525, 335, 219, 415, 1823, 166, 504, 1728, 621, 1459, 59, 869, 1608, 1346, 934, 340, 1407, 352, 1223, 1726, 73, 2095, 523]. method [923, 207, 2292, 2007, 418, 641, 1380, 1119, 1498, 259, 355, 1968, 563, 1727, 288, 490, 1384, 100, 1869, 1177, 75, 1816, 719, 1052, 1637, 1432, 1166, 596, 2014, 161, 1061, 1378, 1903, 1630, 878, 1187, 2253, 1676, 1165, 891, 34, 387, 921, 1634, 298, 1600, 112, 1629, 1081, 895, 1101, 1253, 739, 53, 598, 372, 406, 1361, 1381, 1476, 1564, 2079, 1947, 2056, 2297, 2299, 1055, 717, 990, 263, 975, 1339, 1694, 1236, 1399, 2185, 2162, 1814, 1557, 2290, 1870, 1316, 756, 1618, 226, 810, 938, 933, 1412, 1578, 2127, 1974, 1939, 336, 814, 1434, 1944, 798, 1150, 2220, 2264, 503, 261, 46, 998, 1565, 532]. Method [836, 1875, 105, 275]. methodology [423, 2143, 951]. Methods [609, 1493, 1553, 2189, 1616, 1140, 1469, 626, 918, 424, 302, 160, 2038, 583, 125, 1712, 345, 1940, 284, 243, 1975, 43, 1739, 1666, 1135, 1354, 2033, 1602, 1774, 1067, 1714, 1383, 1586, 1668, 320, 2225, 1112, 512, 1203, 616, 1262, 1910, 203, 1797, 2277, 1355, 1915, 301, 2273, 2175, 1472, 483, 903, 1872, 1539, 2029, 1495, 1576, 2085, 2074, 2116, 1513, 204, 173, 1858, 604, 1290, 686, 1609, 1829, 2040, 1250, 771, 543, 539, 865, 811, 158, 2069, 53, 500, 1979, 2090, 1192, 648, 1278, 41, 1179, 277, 1501, 976, 1749]. metric [1438, 1693]. Mexico [1302]. MgO [951]. MgO/H [951]. mhd [761, 968, 1454, 47, 358, 1687, 948, 779, 1460, 492, 2143, 1346, 1377, 266]. Michaelis [1323]. micro [889, 1454, 708, 487, 420, 1622]. micro-beams [889]. Micro-Electromechanical [1622]. micro-polar [1454, 708]. micro-rotation [487]. microbial [1315, 2135]. microchannel [1395]. Microgrid [760, 1654, 448, 421, 1642]. microgrids [2003, 2231]. micromagnetorotational [2076]. microorganisms [1129, 1689]. micropolar [796, 358, 671]. microsimulation [760]. microstretch [617]. microstructure [295]. migration [325, 2240, 2152, 1256]. mill [2254]. Milovic [356]. mimetic [1325]. MIMO [1027, 2026]. min [1546]. min-max [1546]. minimal [478]. Minimax [2238]. minimisation [1987]. minimization [1992, 1106, 598, 1310]. Minimizing [665]. minimum [1930, 1206, 2193]. mining [464, 1686]. Minkowski [1699]. miscible [513]. mismatch [1310]. mismatched [451]. missing [663, 354]. Mitigating [2110, 1159]. Mittag [793, 405, 846]. Mittag-Leffler [793, 405, 846]. Miwa [876]. Mixed [692, 1027, 574, 761, 861, 2131, 1490, 623, 1029, 1184, 1708, 1403, 1880, 1991, 1168, 702, 2129, 1579, 629, 305, 982, 73, 523, 787, 1267, 954, 1859, 287, 706, 1081, 981, 598, 1817, 980, 846, 1199]. mixed-hybrid [623]. mixed-integer [982]. mixed-type [1267]. mixture [2023, 2300, 1257, 1741, 2221]. mixtures [582, 863]. MLPG [543]. MLS [1179]. MMAP [2022]. MMC [1768]. MMC-based [1768]. mobile [655, 950, 1149, 716, 1560, 474]. modality [588]. Mode [1642, 1553, 2131, 26, 781, 306, 229, 1778, 1841, 16, 578, 1154, 1963, 1698, 287, 351, 955, 122, 1308, 1335]. Model [285, 642, 1128, 2283, 1345, 1993, 889, 538, 746, 709, 348, 961, 754, 643, 1227, 1632, 380, 1314, 1584, 1542, 570, 181, 123, 1333, 1086, 1482, 922, 1802, 825, 912, 413, 2279, 796, 1283, 1066, 1862, 1806, 470, 2136, 623, 1969, 971, 1555, 2084, 1650, 1663, 832, 1623, 1941, 62, 6, 856, 1319, 1367, 1184, 1745, 1708, 1679, 1975, 2006, 1094, 785, 1137, 1838, 1905, 325, 931, 1880, 727, 1763, 868, 270, 2121, 2031, 991, 2061, 1855, 426, 453, 1774, 2191, 580, 833, 1884, 295, 157, 1538, 1906, 2266, 2240, 1871, 2030, 1896, 2107, 1172, 674, 748, 324, 752, 1153]. model [2157, 1441, 290, 260, 1426, 363, 382, 829, 1994, 2144, 1451, 2083, 1883, 2151, 845, 1284, 1803, 904, 730, 1146, 1778, 1946, 657, 367, 2096, 1839, 1912, 1211, 577, 364, 332, 104, 595, 841, 405, 262, 1997, 1808, 2294, 1952, 860, 385, 238, 414, 1551, 1995, 1517, 489, 992, 1701, 334, 399, 877, 39, 248, 948, 957, 866, 721, 1058, 1397, 1828, 1970, 1001, 1504, 1767, 1852, 1574, 222, 258, 82, 124, 644, 546, 960, 965, 757, 1096, 1036, 1154, 1877, 1700, 1796, 1711, 2011, 2152, 2029, 2165, 2219, 245, 1157, 2238, 2291, 357, 401, 1158, 1664, 1794, 1516, 1034, 329, 554, 1951, 1919, 1302, 1665, 1488, 1149, 2207]. model [987, 2041, 1597, 1103, 1457, 937, 2196, 2246, 1955, 871, 173, 176, 1515, 1833, 2106, 816, 690, 1988, 665, 941, 1526, 1049, 974, 1041, 340, 1210, 1229, 1257, 1087, 128, 1492, 441, 1379, 2255, 87, 641, 855, 217, 1291, 428, 356, 603, 1174, 2117, 259, 1675, 2000, 1892, 2058, 914, 604, 1865, 308, 1322, 649, 1214, 630, 1349, 1949, 801, 794, 1274, 225, 645, 2218, 1178, 1043, 1980, 1971, 1736, 1186, 2015, 1678, 601, 1176, 2014, 2044, 1903, 858, 691, 8, 2206, 1099, 1747, 1511, 738, 1857, 2034, 758, 387, 1071, 706, 1435, 2066, 1533, 1825, 1, 389, 417, 1422, 584, 1371, 1762]. model [1430, 2045, 676, 675, 2154, 955, 113, 574, 386, 862, 688, 1059, 1246, 1476, 1615, 1820, 1976, 2160, 1984, 2161, 587, 1822, 2150, 317, 309, 397, 1420, 1835, 1853, 1953, 2016, 547, 572, 1810, 953, 402, 1236, 1425, 1399, 1738, 2185, 1196, 499, 2287, 1015, 817, 1499, 2221, 370, 2213, 1571, 226, 1244, 1471, 1844, 1673, 1718, 1638, 1617, 1864, 1890, 2025, 2260, 920, 809, 1621, 843, 1628, 1916, 2227, 466, 1199, 1943, 1354, 242, 95]. model-assisted [1905]. model-based [1778, 173]. Modeling [749, 968, 465, 1925, 1057, 1683, 189, 1404, 513, 1742, 874, 707, 1669, 1590, 487, 693, 492, 1552, 59, 83, 2027, 2117, 21, 1510, 1732, 2070, 420, 2293, 24, 908, 786, 2153, 890, 594, 1733, 273, 2122, 1439, 1287, 512, 127, 1395, 736, 2248, 2223, 2215, 841, 1722, 1175, 2242, 1414, 1415, 416, 221, 515, 1343, 1996, 2155, 733, 2032, 565, 1292, 2077, 1342, 2302, 1741, 2236, 427, 2244, 581, 1406, 445, 180, 1523, 1102, 1557, 2222, 429, 1749, 1280, 2300, 592]. Modelling [510, 2176, 609, 1054, 172, 847, 97, 252, 734, 1965, 176, 978, 2078, 434, 2057, 1927, 435, 1365, 442, 468, 297, 174, 185, 422, 744, 175, 449, 1764, 173, 2209, 517, 2022, 287, 1256, 430, 1713, 395, 609, 412]. models [1929, 653, 388, 1710, 902, 450, 498, 1715, 1394, 1646, 1740, 1545, 1077, 1983, 995, 365, 776, 199, 2067, 2225, 1824, 1783, 605, 201, 997, 2300, 894, 924, 1124, 1513, 1348, 1863, 1622, 993, 78, 1563, 1800, 576, 276, 1535, 2296, 865, 1964, 1382, 1374, 67, 1917, 701, 218, 1618, 1578, 2283]. modern [1547, 1004]. modes [1761]. modification [216]. Modified [1620, 1581, 735, 2158, 1033, 247, 1163, 643, 1227, 1333, 302, 12, 1983, 1044, 2061, 880, 1239, 1906, 1112, 1090, 1946, 650, 1259, 136, 1218, 1874, 486, 2273, 1877, 1277, 2196, 511, 477, 1675, 100, 1549, 1115, 2015, 758, 2299, 930, 572, 756]. modifiers [342]. modular [982, 1760]. modulated [195]. modulation [2174, 1595, 303]. module [12, 1766, 2094]. modulus [204]. modulus-based [204]. MOEDA [1437]. Mohand [940]. moisture [2122]. moment [2280, 1458, 82, 1376, 1327, 83, 1842]. moment-based [1458]. moment-independent [2280]. moments [1304, 1657, 1967, 1177]. monitoring [297]. monkeypox [1414]. Monod [1555, 1810]. monopoly [2108]. monotone [979, 1046, 1714, 557, 1578, 976]. Monotonicity [652, 702]. Monte [1542, 1394, 145, 1910, 898, 1453, 340, 73, 2044, 392]. Monte-Carlo [145, 340]. Morlet [120]. Morocco [1879]. morphological [58]. morphology [2246]. mortality [1249]. MoS [643]. mosquito [734, 1185, 2236]. mosquito-borne [2236]. mosquitoes [967]. MOST [898]. moth [588, 42, 1580]. moth-flame [588, 42]. motility [1520, 2020]. Motion [2164, 977, 1304, 1140, 1452, 2176, 793, 1836, 874, 1153, 708, 917, 671, 1859, 901]. motor [26, 17, 22, 20]. motors [1770, 24]. mould [867]. mountain [627]. movement [1423]. moving [1122, 327, 1300, 992, 1687, 2053, 1144, 671, 562, 83, 283, 1676, 2239]. moving-barrier [1144]. MPPT [16, 11]. MPS [1535]. MRI [662]. MS [331]. MS-stability [331]. Multi [1335, 1437, 1691, 1039, 2130, 1160, 1684, 2106, 1911, 2007, 2058, 803, 2124, 2173, 1248, 1617, 2060, 85, 1494, 1440, 1079, 1473, 390, 1469, 465, 278, 972, 1941, 1385, 107, 588, 2103, 362, 1403, 1118, 781, 1788, 1913, 2225, 1193, 471, 300, 1593, 2024, 1197, 1748, 1453, 2247, 1218, 1251, 2175, 667, 1852, 1021, 1506, 1711, 1415, 401, 656, 2116, 1488, 1525, 987, 2041, 1656, 978, 2209, 1735, 1312, 893, 2000, 767, 2022, 506, 2256, 1125, 719, 1207, 1061, 2044, 368, 1468, 1676, 34, 1793, 1634, 1253, 1430, 430, 872, 1396, 1104, 1464, 2070, 2297, 1055, 990, 1652, 1008, 122, 1183, 1636, 2187]. multi [249, 1635, 2097, 1524, 2028]. Multi-Active-Bridge [2028]. multi-agent [107, 781, 1788, 1913, 667, 1021, 2256, 1793, 1104, 2187, 1635]. Multi-breather [2173]. multi-cell [34]. multi-constituent [2175]. multi-criteria [978, 2209]. multi-currency [1207]. multi-delay [1711]. multi-delayed [1593]. multi-dimensional [390, 471, 1251, 1061, 1676, 1253]. multi-dimensions [1440]. multi-domain [1634, 2297]. multi-energy [430]. multi-error [1218]. Multi-force [1039]. multi-forecasting [2044]. multi-fractional [987]. multi-fuzzy [972]. multi-gear [1488]. multi-group [401]. multi-image [2247]. Multi-Label [1684, 1691]. multi-layer [1506, 1468]. multi-leader [1415]. multi-level [368, 1464]. multi-link [1593]. Multi-material [2130]. multi-modality [588]. Multi-Objective [1437, 1911, 2007, 2058, 85, 362, 2024, 1453, 1735, 1312, 2070, 1636, 249, 2097]. Multi-operator [1335]. Multi-parameter [2106, 2041]. multi-patch [278, 1852, 1396]. Multi-period [2124, 300]. multi-phase [1469, 1941, 990]. multi-piecewise [2225]. multi-population [1652]. multi-rate [1656]. Multi-regime [1617]. multi-resolution [1079, 1473]. Multi-robot [803]. Multi-sample [1248]. multi-scroll [1183]. multi-shift [656]. multi-singular [506]. multi-slave [872]. multi-source [465]. multi-sourcing [1008]. multi-stage [2103]. multi-state [1385, 2022]. multi-step [1494]. Multi-strategy [2060, 1118, 1197]. multi-structure [2116]. Multi-swarm [1160]. multi-term [1403, 1193, 1748, 1525, 767, 719, 1055, 1524]. multiagent [2298]. multicellular [440]. multicomponent [195, 666]. multiderivative [1797]. multidimensional [388, 1080, 2205, 1716, 1990, 1519, 1474, 131, 1792, 1947]. multidirectional [919]. multifactor [1623]. multigrid [696]. multilayer [2055]. multilevel [1394, 439, 43, 2267]. multimodal [326, 1911, 1312]. multinode [1737]. multiphase [1608]. multiphysics [238]. Multiple [198, 328, 27, 1115, 729, 1547, 709, 1347, 699, 580, 1445, 1521, 327, 829, 1818, 1058, 1631, 556, 291, 173, 2140, 352, 1257, 128, 2022, 288, 183, 1388, 394, 111, 180, 315, 403, 1560, 846, 1461, 547]. multiple-delayed [315]. Multiple-order [328, 729]. multiple-relaxation [352]. multiple-relaxation-time [2140]. multiple-runway [183]. multiple-sized [1631]. multiple-solitons [556]. multiple-susceptible [1058]. multiplex [1579, 586]. multiplicative [651, 1467]. multiplicity [1641]. multiplier [1627, 697, 1101]. multipoint [400]. Multiquadric [234, 190]. MultiQuadrics [2138]. multiresolution [72]. Multiscale [1646, 936, 700, 861, 1333, 722, 2266, 78, 598]. multiscroll [919]. multisensor [354]. Multistability [819, 868, 657]. Multistep [1135, 2288, 1872]. multivariable [2294]. multivariate [1050, 863, 2221]. Müntz [759]. Müntz-Legendre [759]. Murnaghan [1056]. mutation [935, 1095, 1415, 1422, 1181, 1374, 272]. mutation-induced [272]. Mutualism [963].

N [2216, 1900]. N-policy [2216]. NaCl [260]. NaCl-system [260]. Nador [1879]. Nagumo [1045, 363]. Nakagami [475]. nano [643, 1129, 708, 704, 346]. nano-fluid [708, 346]. nano-material [643]. nano-sensors [704]. nanofluid [761, 1303, 1454, 44, 2076, 1395, 1126, 917, 2129, 79, 1687, 948, 779, 1854, 787, 733, 951, 105, 932, 800, 753, 537, 1685]. nanofluid-saturated [753]. nanofluids [487, 366, 266]. nanoindentation [1030]. nanolayer [2076]. nanoliquid [959]. nanomaterial [461]. nanomaterials [837, 1750]. nanoparticle [2076]. nanoparticles [1395, 657, 741, 1466, 1865, 61]. Nash [2171, 110]. National [225]. Natural [1126, 380, 510, 1455, 311, 764, 496, 1337, 1003, 554, 2057, 1976]. Nature [460, 1544, 963]. Nature-inspired [460, 963]. naval [1764]. Navier [620, 696, 2114, 1252, 1056, 1694, 282]. near [191, 79, 366, 729]. near-best [191]. nearly [2166]. needle [1687]. negative [2110, 344, 996, 1102, 392, 2080]. negativity [1602]. neighborhood [2219]. neighbour [764]. neighbourhood [309]. Neisseria [1074]. Nernst [2224]. nesting [1119]. nets [197, 392]. network [1547, 450, 297, 1238, 1620, 453, 1780, 168, 290, 2128, 464, 164, 1228, 2039, 1175, 1970, 2087, 1450, 13, 393, 452, 978, 2241, 1610, 718, 120, 645, 1116, 228, 2160, 474, 1866, 1628, 33, 494]. network-based [453, 718]. Networked [999, 354, 369, 2187]. networks [485, 1584, 1508, 1409, 109, 973, 1692, 1404, 1347, 737, 1702, 785, 2017, 2141, 1518, 1521, 1593, 1946, 103, 1912, 2224, 966, 740, 655, 1286, 214, 549, 1053, 1506, 1017, 1005, 162, 819, 1359, 1579, 2018, 1596, 1408, 452, 184, 1301, 391, 1698, 692, 1136, 1219, 1120, 969, 1873, 396, 2282, 88, 1724, 268, 2302, 2054, 1971, 670, 853, 101, 1468, 111, 1338, 495, 430, 386, 1255, 2079, 271, 572, 403, 493, 2295, 1065, 1817, 980, 1560, 589, 1362, 1062, 1461, 2089, 1000, 586, 2045]. networks-based [1408]. Neumann [616, 2265, 372]. Neural [33, 13, 2045, 1547, 485, 1508, 1626, 297, 1238, 109, 973, 1347, 737, 1702, 785, 2275, 2017, 453, 1521, 1780, 1593, 2128, 464, 103, 2224, 164, 966, 1228, 2039, 740, 1175, 1970, 2087, 1286, 1005, 162, 819, 1359, 1450, 2018, 1596, 1408, 452, 1301, 391, 2241, 1698, 1610, 1136, 1219, 1120, 969, 1873, 718, 2282, 120, 1724, 2302, 670, 853, 101, 111, 495, 1116, 228, 1255, 2079, 271, 403, 2295, 1065, 1817, 980, 1560, 1362, 1062, 1461, 2089, 474, 1000, 1866, 494]. Neuro [506, 584, 954, 576]. neuro-computing [954]. Neuro-evolution [506, 576]. Neuro-swarm [584]. neurodynamic [1592]. neuron [363, 974]. neurons [2109]. neutral [531, 320, 103, 844, 2087, 969, 1311, 811, 981]. neutral-type [2087, 969]. News [50, 64, 89, 117, 139, 150, 169, 186, 211, 253, 292, 312, 337, 374, 431, 455, 507, 524, 567, 606, 632, 680, 713, 725, 788, 804, 826, 850, 881, 909, 942, 984, 1018, 1068, 1108, 1200, 1232, 1269, 1328, 1356, 1390, 1417, 1442, 1478, 1485, 1527, 1568, 1603, 1643, 1680, 1705, 1753, 1776, 1847, 1887, 1921, 1934, 2008, 2046, 2071, 2100, 2132, 2145, 2182, 2210, 2233, 268]. Newsvendor [1679, 62, 365]. Newton [1774, 550, 2225, 2085, 1869, 1599]. Newton-like [1869]. Newtonian [597, 776, 237, 741, 2053, 227, 346]. Nicholson [108, 1704]. Nickle [487]. Nicolson [1150]. Niels [592]. Nielsen [1542]. nighttime [70]. NILT [1182]. nine [1769]. nine-phase [1769]. Nitsche [620, 1029]. Nizhnik [651]. NLS [1814]. NLS-type [1814]. NMC [419]. no [30, 555]. nodal [49, 1191]. Node [1093, 384]. Node-Diamond [1093]. Noise [1021, 1862, 2225, 363, 1228, 1951, 651, 2255, 20, 1909, 675, 402, 1467, 474]. Noise-induced [1021, 1951]. noise-polluted [474]. noises [1990]. Noisy [1413, 749]. Non [1717, 991, 1602, 699, 260, 740, 495, 1901, 1064, 1314, 2148, 1958, 1624, 597, 1585, 1905, 325, 776, 1135, 731, 1241, 1354, 611, 778, 1669, 1395, 1994, 2099, 1999, 237, 841, 1100, 2129, 1588, 1687, 2053, 1040, 1124, 416, 1158, 409, 329, 1387, 2246, 1190, 690, 2198, 83, 1914, 641, 1291, 106, 1703, 1819, 996, 2236, 886, 753, 593, 1510, 406, 999, 1339, 1012, 2187, 1557, 346, 980, 446, 683, 1583, 976, 741, 1886, 227]. non-autonomous [1040, 1012]. non-capacity [841]. non-chain [1819]. non-constant [1994]. non-convex [1100]. non-Darcian [753]. non-degenerate [1583]. non-diabetic [1314]. non-differentiable [106, 2187]. non-dissipative [416, 1291]. Non-dominated [991]. non-endemic [2236]. non-equilibrium [753]. Non-fragile [699, 740, 495, 1901, 999, 980]. Non-Gaussian [1886]. non-grey [329]. non-holonomic [1387, 83]. non-instantaneous [1124, 2198]. non-invasive [446]. non-isothermal [2129]. Non-iterative [260]. non-linear [731, 1241, 1354, 1588, 1687, 1190, 690, 1914, 1703]. non-local [325]. non-monotone [976]. non-negative [996]. Non-negativity-preserving [1602]. non-Newtonian [597, 776, 237, 2053, 346, 741, 227]. non-paraxial [1624]. non-particle [2246]. non-perturbative [778]. non-pharmaceutical [1510]. non-polynomial [1557]. non-probability [1905]. non-regular [1999]. non-self-adjoint [886]. non-singular [1669]. non-smooth [1588, 1158, 1339, 683]. non-spherical [1395]. Non-stationary [1064]. non-symmetric [406]. Non-uniform [1717, 2148, 1958, 611, 409, 1703, 593]. non-vanishing [1135]. non-zero [2099]. nonautonomous [691]. nonconforming [2292]. nonconservative [2107]. nonconstant [591, 1249]. nonconvex [1352, 2299]. Nonfragile [1541]. nonholonomic [716]. nonidentical [370]. Nonlinear [2270, 1441, 2254, 2301, 461, 1240, 2001, 979, 1440, 1473, 1719, 1163, 1045, 775, 1042, 2051, 1314, 2226, 498, 1016, 870, 1490, 297, 1503, 2272, 1409, 612, 96, 1077, 1620, 1983, 1624, 618, 1739, 1666, 2031, 777, 1035, 781, 1774, 1132, 407, 1239, 2180, 1913, 157, 471, 2030, 349, 695, 1426, 1729, 557, 469, 2081, 505, 845, 1084, 1640, 542, 573, 1262, 2258, 156, 527, 1259, 364, 2277, 299, 551, 927, 2123, 2174, 57, 239, 82, 483, 879, 1047, 1053, 1796, 1577, 162, 561, 544, 735, 1450, 1495, 1801, 1956, 1963, 831, 401, 959, 1392, 1387, 2116, 1595, 987, 384, 1674, 937, 2143, 812, 1474, 906, 1041]. nonlinear [925, 1639, 923, 1161, 2292, 2286, 1027, 1174, 2125, 303, 1609, 1949, 1078, 286, 506, 2256, 1115, 1736, 543, 144, 710, 1222, 800, 1898, 891, 1901, 387, 1651, 1253, 398, 584, 1371, 1633, 2069, 574, 500, 887, 1361, 1059, 2160, 2271, 67, 2173, 1904, 701, 122, 1399, 2063, 1114, 2290, 275, 815, 1673, 1843, 1938, 798, 1150, 820, 46]. nonlinear-flow [2030]. nonlinearities [1163, 853, 1548, 1744]. nonlinearity [1285, 398, 1114]. Nonlocal [2222, 1066, 1401, 799, 2159, 1855, 2191, 1589, 1808, 2085, 87, 765, 2253, 755, 1179]. nonmatching [2269]. nonmonotonic [1096]. nonparallel [428]. nonprobability [1824]. nonrelativistic [2115]. nonresponse [522]. Nonsingularity [2138]. nonsmooth [1885, 1916]. Nonstandard [267, 1923, 538, 825, 904, 730, 399]. Nonsymmetric [1392]. Nontransitivity [997]. Nonuniform [2285, 1193, 1042, 2194]. nonzero [934, 2171]. nonzero-sum [934, 2171]. Noor [807]. norm [1260, 1629]. normal [863, 250, 2007]. Normalized [1787, 200, 289]. note [822, 2043]. Novel [1584, 995, 1895, 2243, 1595, 120, 1814, 1562, 1860, 1544, 786, 1806, 1409, 1657, 1536, 1118, 1899, 1594, 963, 256, 1807, 1280, 1991, 1912, 1247, 2294, 47, 1037, 1145, 2242, 2087, 2288, 124, 1005, 544, 2284, 1525, 1656, 2139, 1369, 867, 873, 1125, 1630, 921, 2179, 1276, 2188, 2213, 275, 1212, 1412, 2025, 2220, 920, 1513]. November [215, 946, 1393, 1781, 2214]. Novikov [651]. NSFD [927, 1833]. nucleation [1981]. number [1743, 228, 857]. numbers [892, 1217, 686, 1377, 1002]. numeric [1974]. Numerical [1719, 1324, 1140, 1283, 2136, 1077, 1122, 2203, 1666, 2114, 1812, 343, 2240, 2030, 1846, 2042, 650, 841, 917, 1213, 2129, 1874, 381, 1167, 1337, 1796, 1460, 2029, 1801, 1956, 562, 2208, 1826, 329, 1665, 1149, 1734, 1854, 1622, 1024, 1475, 1805, 923, 2190, 733, 1619, 686, 932, 771, 640, 891, 811, 1002, 266, 717, 648, 1236, 756, 533, 2025, 1098, 798, 838, 216, 1033, 1606, 560, 1482, 220, 302, 796, 870, 468, 318, 2038, 1057, 2134, 1836, 1209, 2021, 1520, 1137, 325, 615, 931, 1436, 1721, 550, 1981, 707, 2180, 295, 1668, 1746, 1224, 260, 382, 155, 708, 1090]. numerical [1640, 730, 1946, 1716, 1463, 156, 364, 532, 602, 1995, 1517, 47, 992, 1915, 877, 1289, 1263, 900, 2087, 1989, 1641, 134, 1577, 1973, 1028, 897, 404, 1408, 86, 673, 1513, 219, 599, 621, 1049, 934, 1210, 925, 1639, 322, 428, 762, 1380, 356, 1429, 565, 563, 649, 1829, 286, 490, 1384, 2077, 75, 2251, 1297, 394, 161, 283, 1061, 1378, 1903, 1630, 1491, 1909, 1511, 738, 886, 1410, 1422, 1208, 502, 1798, 1399, 1738, 1467, 1578, 2267, 896, 2043, 1650, 1783, 1845, 1612, 1370, 1858]. numerically [2036, 1792]. nursing [1173]. nutrient [1001]. Nyström [1494, 472, 1574]. NZSDG [2171].

O [260, 779, 643, 779, 951]. OA [1242]. obesity [2117]. object [227]. Objective [1437, 205, 85, 1080, 362, 1173, 2024, 1453, 950, 1794, 1911, 1735, 1312, 2007, 2058, 2070, 1636, 249, 2097]. objectives [1773]. objects [1455]. observability [98, 463, 899]. observables [2134]. observation [1327, 1435]. observations [605]. observed [763]. Observer [2051, 1841, 31, 781, 469, 1031, 447, 27, 2098, 1898, 1104, 1904, 1659]. Observer-based [2051, 1841, 781, 1904]. observers [701]. obstacle [1203]. obstacles [859]. obtain [1145, 1101]. obtained [217]. obtuse [206]. occupancy [2232]. occurrence [1703]. occurring [96]. ocean [1694]. oceanography [2173]. October [571, 913, 1360, 2186]. octree [1030]. octree-type [1030]. odd [1712, 521]. odd-degree [1712]. odd-order [521]. ODE [1239, 2293]. ODEs [1494, 1666]. ODIN [327]. odour [2121]. Off [415, 1608, 1862, 1960]. Off-lattice [1608]. Off-line [415]. off-shore [1960]. offshore [729]. oil [1582, 1457]. Oldroyd [266]. Oldroyd-B [266]. oligopoly [1880]. omicron [2242]. on-chip [656]. onchocerciasis [276]. oncolytic [2004]. One [796, 1398, 1862, 1983, 2121, 1559, 2266, 915, 2265, 347, 1015]. One-dimensional [796, 1398, 1559, 2266, 2265, 347]. one-predator [1862]. online [512, 290]. only [834]. onset [1689]. open [1470, 1783, 2259]. open-channel [2259]. operating [1761, 355, 1292]. operation [1960, 803, 596, 1248]. operational [135, 1665, 1525, 759, 1011, 907]. operations [1279]. operator [979, 1335, 2036, 1318, 1042, 2059, 967, 1714, 1538, 1669, 1906, 1370, 265, 1609, 1221, 858, 990, 1321, 2097, 2222, 466]. operators [1042, 518, 1470, 1714, 631, 103, 405, 1460, 1152, 406, 1737, 540]. opetition [1214]. opinion [464, 2209]. opinions [2204]. opportunities [555]. Opposition [42, 1095, 1599, 732]. Opposition-based [42, 1095, 1599, 732]. optical [2099, 2081, 957, 303]. optically [1490, 329]. optics [906, 2173]. Optimal [1091, 2272, 612, 1184, 1662, 43, 306, 1323, 104, 1852, 1036, 1254, 1230, 409, 476, 907, 1632, 2226, 626, 1016, 912, 1558, 1740, 1983, 1319, 1720, 2163, 1966, 773, 1942, 2113, 874, 1586, 1790, 1203, 1883, 614, 1400, 840, 1839, 1397, 1575, 2004, 950, 377, 2278, 504, 1190, 83, 1514, 1726, 2095, 762, 20, 794, 928, 1274, 2206, 1923, 445, 93, 1820, 2070, 2160, 2056, 1416, 1917, 2185, 533, 1471, 2260, 1621, 747]. optimality [1121]. optimisation [1534]. Optimization [782, 1770, 2261, 2198, 1459, 951, 1592, 977, 1121, 1553, 2028, 1860, 1544, 442, 1080, 1775, 465, 470, 15, 1759, 1813, 2021, 362, 1908, 1073, 1118, 1518, 42, 1594, 963, 695, 1264, 1752, 300, 324, 1107, 764, 2024, 689, 1655, 1090, 1512, 2252, 840, 935, 898, 1453, 679, 1334, 1038, 1100, 1242, 2130, 732, 1095, 2219, 447, 947, 1160, 1103, 1190, 1519, 1911, 712, 1369, 1735, 867, 1312, 1006, 1610, 1475, 2007, 434, 1292, 361, 1342, 1549, 307, 1186, 1352, 2124, 1216, 2066, 1205, 1181, 1195, 1088, 2299, 949, 1652, 975, 218, 460, 353, 1276, 1248, 1636, 272, 1266, 249, 1212, 1412, 1578, 1974]. optimization [2119, 2097, 2060, 2267, 261, 679, 1085, 1106]. Optimizations [2, 60]. optimize [2088, 1293]. Optimized [419, 1251, 1129, 364]. optimizer [588, 1415, 742, 1580, 249, 2267, 1840, 1160]. optimizing [1320]. Option [1623, 114, 615, 2067, 605, 2205, 1782, 1103, 1457, 869, 340, 1310]. options [1265, 1456, 1893, 1582, 248, 1144, 2207, 1513, 700, 1859, 1984, 2043]. orbit [1198, 807]. orbits [2216]. Orchard [1242]. Order [1354, 1494, 1180, 2262, 1860, 1440, 1473, 1344, 485, 1227, 1632, 390, 1318, 620, 998, 1086, 890, 77, 1483, 160, 2272, 1238, 2122, 125, 1409, 842, 1336, 244, 114, 737, 1285, 1581, 1534, 1739, 521, 931, 1439, 1666, 270, 2121, 777, 1436, 2289, 1132, 2113, 1151, 2141, 880, 343, 1899, 624, 1788, 138, 888, 1193, 1906, 1260, 471, 349, 1566, 1447, 1203, 2035, 945, 328, 49, 229, 1723, 1306, 103, 1262, 1912, 202, 527, 1247, 520, 1355, 371, 486, 358, 1145, 2178, 2242, 900, 2087, 578, 1460, 1577, 2011, 2165, 544, 735, 1134, 1495, 2199, 2085, 831, 1596, 1670, 2074, 1387, 1302, 1149, 1525]. order [2278, 166, 1190, 400, 1526, 1474, 1041, 1407, 692, 1136, 762, 759, 535, 1498, 1120, 1573, 1873, 718, 396, 1727, 286, 120, 576, 506, 1341, 1448, 1221, 1250, 2013, 670, 601, 766, 1903, 738, 1225, 1509, 1165, 886, 865, 921, 1410, 1128, 581, 1406, 1299, 372, 1361, 1104, 1382, 587, 1055, 990, 729, 373, 403, 493, 1245, 502, 1155, 122, 1065, 2162, 1814, 2221, 1278, 1316, 533, 275, 815, 1321, 1062, 1635, 1461, 1718, 1950, 1938, 282, 798, 1150, 2245, 1885, 820, 1894, 896]. order-dependent [1062]. ordered [972]. orders [1666]. ordinal [516]. ordinary [1562, 2038, 1361]. Ordinate [1875]. organic [2135]. orientation [159]. original [2249]. Ornstein [2151, 1822, 1810, 809]. orthogonal [558, 160, 1623, 1252, 2278, 654, 593, 1382, 1842]. orthogonalized [1326]. orthopair [1691, 2093]. orthotropic [1119]. oscillation [407, 2096, 86, 530]. oscillations [2131, 2031]. oscillator [778, 895, 1661, 683]. oscillators [1084, 530, 1609]. oscillatory [1315, 2203, 364, 1268, 2200, 1424, 1956, 814, 1944]. oscillatory-type [2203]. osculatory [1787]. Ostrowski [345]. other [1240]. outbreak [834, 727, 1159, 857]. Outer [1475, 549, 1943]. outliers [1257]. outlook [708]. Output [1351, 1586, 2294, 421, 369, 1214, 853, 2185, 1461]. output-coupled [1461]. over-determined [849]. overfilled [1866]. Overhead [32]. overland [1517]. oversaturated [385]. overset [1340]. oxygen [752, 221]. oxytactic [1129].

Padé [1918]. Padovan [698]. Pages [68, 154, 257, 40, 121, 143, 54, 215, 378, 639, 316, 296, 459, 341, 571, 528, 745, 854, 988, 684, 661, 808, 830, 728, 792, 946, 913, 885, 1164, 1298, 1421, 1072, 1022, 1237, 1273, 1133, 1204, 1393, 1360, 1332, 1531, 1851, 1489, 1446, 1607, 1647, 1502, 1572, 1781, 1709, 2012, 2149, 2237, 1937, 1891, 2075, 2104, 1959, 2050, 2214, 2186, 2169]. pair [71, 1470, 386]. pairwise [2197]. Pakistan [265]. Pancreas [2168]. pandemic [1013, 350, 928]. pandemics [1941]. panel [8]. pantograph [232, 134, 576]. paperback [592]. papillomavirus [476]. parabola [1799]. parabolic [861, 839, 707, 1193, 602, 360, 167, 504, 1429, 1341, 797, 1378, 598, 1947, 502, 1155, 277, 1098]. paradigm [584, 1192, 1540]. parallel [1553, 1394, 1627, 1398, 724, 242, 1441, 136, 1038, 635, 642, 1765, 1610, 1075, 1979, 282, 2245, 1447]. Parallel-in-Time [1447]. Parameter [501, 614, 238, 563, 1378, 2154, 218, 1402, 839, 132, 470, 793, 194, 1940, 678, 1228, 1504, 1711, 416, 2041, 2106, 1996, 1924, 1166, 1985, 180, 1155, 2283, 494]. parameter-free [194]. Parameter-robust [238]. Parameter-uniform [563]. parameter-variable [1228]. parameters [834, 1657, 856, 1058, 2011, 163, 2241, 418, 649, 499, 1866]. parametric [2040, 1125, 747]. parametrised [894]. parametrization [1825]. parareal [2120]. Paraskevopoulos [1011]. paraxial [1624]. Pareto [1441, 1809]. parity [1235]. Park [225]. Parkinson [1571]. parsimonious [2023]. partial [1440, 1469, 1503, 1559, 780, 1640, 704, 381, 1392, 1674, 2198, 1113, 648, 1266, 533, 1434, 1150, 2264]. partial-integrable [1640]. partially [513, 53, 763]. participation [2142]. Particle [145, 200, 81, 1140, 2246, 1190, 1768, 686, 307, 1216, 1248, 2119, 2026, 1106]. particles [908, 236, 245, 1002]. partition [1300, 906, 206, 1858]. partitioned [1649]. partitions [2137]. partly [369]. partners [476]. parts [687, 1726, 1299]. Pascal [698]. pass [627]. passage [1048]. Passive [679, 429, 878]. Passivity [1120, 351, 2157, 692, 1255]. passivity-based [2157]. past [859, 358, 779, 105]. Pasternak [613]. patch [278, 1852, 937, 1396, 1704]. patch-constructed [1704]. patches [1300, 479]. Path [353, 2068, 1456, 716, 742, 1276, 1694, 2060]. path-dependent [1456]. pathogen [1059, 2263]. patient [1067, 1464]. patient-specific [1067]. patients [1094, 2223, 758]. Pattern [825, 577, 1786, 2125, 1628, 1314, 1994, 262, 1001, 2165, 142, 1099, 1931, 180, 1396, 953, 2228]. Patterns [2135, 919, 2084, 1315, 995, 752, 1049, 1613, 1678, 706, 1615, 547, 2260]. payment [1231].

PBC [2231]. PCA [1725, 517]. PCB [441]. PCB-embedded [441]. PD [1151]. PDE [1929, 839, 1585, 1967, 1898, 2045]. PDEs [1493, 244, 1386, 1447, 2205, 335, 274, 1858, 2297]. PDR [2176]. PDT [2052]. peak [182]. peakon [669]. pedestrian [2096]. PEM [6]. PEMFC [470]. PEMFCs [1764]. penalization [859, 884]. penalized [53]. penalty [1581, 1668, 628, 2238, 1392, 1372]. penalty-based [2238]. pendulum [1433, 1308]. pendulum-cart [1308]. Penetration [1427]. pentadiagonal [575]. perception [2019]. perception-guided [2019]. perfectly [112]. perforated [116, 722]. perforation [116]. Performance [18, 12, 1551, 836, 929, 1430, 14, 1987, 1182, 2261, 1735, 1152, 8, 445, 1688]. performed [352]. peridynamic [1721, 1516, 1179, 2222]. period [300, 2144, 2124]. Periodic [265, 706, 1012, 1559, 1790, 323, 1631, 1532, 274, 813, 530, 1074, 905, 1435, 669, 683]. periodic-peakon [669]. periodically [1337]. periodicity [115]. periods [1563, 1834]. Permanence [546]. permanent [443, 444, 24]. permeable [537]. permutation [2156]. perpetual [1893]. Persistence [1327, 115, 402]. persistent [501, 1723]. Personal [585]. perspective [1545]. perspectives [1284]. perturbation [1558, 1094, 407, 1846, 1084, 214, 2181, 2014, 683]. perturbations [1555, 1017, 1040, 1667, 1169]. perturbative [778]. perturbed [1079, 839, 694, 1966, 860, 602, 486, 900, 1989, 735, 167, 1955, 1639, 1429, 1968, 563, 1341, 1448, 1166, 797, 1378, 351, 502, 1155, 933, 1939]. Pest [1124, 554, 1976]. pest-natural [1976]. pest-taxis [1976]. pesticides [178]. Petrov [1600, 1215]. Petviashvili [328, 556, 750, 729]. Pfaffianization [876]. PFF [30]. PFF-RVM [30]. phagocytosis [97]. pharmaceutical [1510]. pharmacokinetics [2260]. Phase [517, 2227, 1013, 1469, 594, 18, 1941, 6, 513, 1769, 295, 260, 10, 992, 1992, 1154, 1956, 423, 438, 303, 1290, 1772, 2015, 208, 1732, 1476, 2161, 1853, 990, 2287, 466]. phase-equilibrium [260]. phase-field [1476, 1853, 2287]. phase-lag [2161]. phase-separated [1290]. Phelps [102]. phenomena [216, 736, 1918]. phenomenon [2105, 800]. phosphorylation [2111]. Photovoltaic [1766, 1825, 1775, 14, 451, 1765, 2202, 8, 1762, 1732, 420]. physical [123, 465, 707, 1918, 770, 1747, 1651, 983, 93, 317]. physics [388, 1821, 2224, 176, 205, 2241, 543, 2173, 2045]. physics-based [176]. physics-informed [1821, 2224, 2241, 2045]. physiology [1989]. phytoplankton [1040, 1322]. PI [989, 1943]. Picard [891, 1282, 1501]. picture [2209]. PID [1260, 703]. PIDE [1874, 887]. piecewise [1799, 785, 1287, 2225, 2137, 1999, 1306, 1053, 1907, 2278, 1863, 1278, 814, 1916]. piecewise-smooth [1916]. piezoelectric [812]. PINN [1731]. PINN-EKF [1731]. Pinning [2282]. pipe [116]. pipeline [63]. pipelines [428]. pitch [1038]. PLA [60]. placement [1411, 519]. planar [2118, 945, 1999]. Planck [330, 1836, 2224, 591, 1990, 1634, 41]. plane [1558]. planes [1940]. Planing [281]. plankton [752, 1001, 816, 1865]. planktonic [1597]. planning [2068, 300, 742, 760, 353, 1276, 1694, 2060]. plant [1632, 2270, 222, 69, 287, 1088, 547]. plant-herbivore [2270]. plants [2165, 221, 1221]. plaque [60]. plasma [1398, 155, 2173]. plasmas [279]. plasmid [1927]. Plasmodium [1511]. plastic [112]. plate [1398, 237, 671, 1290, 473, 738]. plates [1537, 496, 477]. platform [174]. PM [26]. PMSG [1462, 1882]. PMSG-based [1882]. PMSM [17, 22, 25, 446]. PMSMs [1943]. pneumatically [281]. poaching [225]. Poincaré [1566]. Point [192, 893, 761, 1454, 160, 1774, 541, 1272, 695, 72, 1556, 400, 934, 1924, 1341, 1297, 1368, 1491, 891, 1317, 1633, 1762, 857, 406, 502, 1155, 1974]. point-to-point [1272]. points [1940, 149, 540]. Pointwise [1267]. Poiseuille [1558]. Poisson [1550, 658, 1585, 776, 1252, 1809, 2224, 836, 73, 1316]. Poissonian [2238]. Pol [323, 364]. polar [1194, 1454, 1287, 708, 2220]. polarized [268]. polarizer [595]. policies [99, 907]. policy [1091, 1663, 1813, 1231, 1334, 1551, 1085, 929, 2198, 1292, 2022, 1342, 2216, 1195]. pollutant [180]. polluted [474]. pollution [1733, 1742, 102, 1850, 2296, 1621]. poly [301, 1024]. poly-fractonomials [1024]. poly-sinc [301]. polyclonal [855]. polygonal [133, 2065, 952, 1537, 523, 1903, 1187, 1111]. polyharmonic [1895]. polymer [59]. Polynomial [202, 2226, 243, 244, 931, 2257, 705, 2229, 997, 135, 1876, 16, 1996, 1530, 1557, 1278, 864, 217]. polynomials [1897, 1586, 299, 551, 467, 762, 1380, 759, 1061, 738, 1011, 1900, 1382]. polytopic [1260]. popular [2142]. population [1993, 2110, 1314, 2153, 1315, 856, 1994, 904, 1722, 1058, 258, 1955, 1552, 941, 1312, 2255, 225, 131, 1964, 1652]. populations [1896, 914, 1374]. poroelasticity [624]. porosity [2176, 2192]. porous [380, 2303, 1558, 1307, 58, 2266, 2030, 2107, 252, 237, 1039, 1631, 86, 534, 1041, 1703, 473, 710, 1002, 753, 2161]. Porsezian [2099]. port [1498]. port-Hamiltonian [1498]. portfolio [1014, 991, 1794, 2124, 1578]. portfolios [148]. posed [695, 2035]. posedness [2161, 1953]. position [443]. positive [1995, 399, 1726]. Positivity [904, 387, 727, 1517]. positivity-preserving [1517]. possibilistic [991]. post [1622]. post-contact [1622]. posteriori [1581, 1648, 1600]. potential [498, 426, 2300, 1326, 178]. pouch [413]. Powell [1746]. Power [1462, 32, 35, 427, 450, 424, 1761, 26, 33, 1977, 1518, 229, 1126, 1497, 421, 1767, 1766, 897, 416, 13, 1823, 1768, 1765, 2202, 742, 441, 434, 1760, 428, 436, 1762, 444, 1732, 1088, 420, 843, 2202]. power-law [1126]. power-of-type-2 [897]. powered [1932]. powerful [247]. powers [2298]. powertrain [676]. pp [592]. PPH [2194]. Practical [894, 2250, 901, 1631]. Practices [172, 173]. PRBS [421]. PRBS-based [421]. Pre [442, 1760, 2253, 300, 1696, 178]. pre-disaster [300]. pre-factor [178]. Pre-optimization [442]. Pre-sizing [1760]. pre-stressed [1696]. Pre-tensorial [2253]. Preassigned [1017]. Preassigned-time [1017]. precast [665]. precession [1189]. Precise [1522, 1177]. precision [704, 1028, 2085, 1557]. Preconditioned [479, 1576]. preconditioner [2059]. preconditioners [84, 43]. preconditioning [1827]. predation [1965, 1288, 706]. Predator [1864, 902, 570, 825, 664, 1862, 2084, 1044, 668, 1363, 1151, 580, 1896, 97, 1952, 721, 1828, 1786, 546, 965, 1096, 1877, 1965, 357, 1158, 1919, 1833, 2106, 1832, 1988, 230, 1185, 233, 1613, 2078, 128, 1492, 802, 1914, 2255, 87, 2292, 1892, 914, 1349, 1949, 1288, 1388, 1115, 1736, 916, 1099, 315, 1615, 1249, 1798, 953, 1012, 755, 1961, 1423, 1015, 817, 370, 1890, 2086, 1916, 1199, 1227]. predator-prey [1864, 1916]. predator-prey-like [1185]. predator-taxis [825]. Predators [1610, 2121, 1172, 1015]. Predefined [1702, 2089]. Predefined-time [1702]. predetermination [445]. predict [318, 295, 1658]. predict-correction [1658]. Predicting [1003, 920]. Prediction [948, 2188, 1978, 1413, 1057, 2294, 2261, 1610, 2155, 458, 873, 645, 113, 1088, 2064]. prediction-based [1413]. predictions [184]. Predictive [1462, 1584, 1409, 1731, 1763, 675, 1943]. Predictor [1630, 550, 358, 628, 758, 331, 1035]. Preface [92, 172, 1757, 1758]. preferences [422, 693, 1773]. prehistorical [925]. premise [27]. PrEP [1664]. Prescribed [1053]. Prescribed-time [1053]. prescribing [776]. presence [638, 4, 787, 1914, 1865, 1099]. preservation [1293]. preserving [1602, 1132, 952, 624, 1871, 471, 1566, 1281, 1226, 904, 573, 697, 1262, 203, 1797, 1995, 1517, 2178, 239, 246, 1154, 1134, 591, 2249, 1305, 1152, 581, 2062, 264, 990, 373, 2063, 1814, 41, 1321, 2127]. Press [592]. pressure [2136, 696, 1404, 1252, 355, 251]. pressure-affected [355]. pressure-dependent [1404]. pressureless [2150]. prevention [1884]. preventive [1632, 2022]. prey [902, 570, 825, 664, 1862, 2084, 1044, 270, 668, 1363, 2121, 1151, 580, 783, 1896, 1172, 97, 1952, 721, 1828, 1786, 546, 965, 1096, 1877, 1965, 357, 1158, 1919, 1833, 2106, 1832, 1988, 230, 1185, 233, 1613, 2078, 128, 1492, 802, 1914, 2255, 87, 2292, 1892, 914, 1349, 1949, 1288, 1388, 1115, 916, 1099, 315, 1615, 1249, 1798, 953, 1012, 755, 1961, 1015, 817, 370, 1864, 1890, 2086, 1916, 1199, 1227, 817]. prey-taxis [1249, 953, 1864]. price [1978, 1231, 605, 2108, 1457, 1143, 2198]. prices [2205]. Pricing [1893, 39, 248, 1859, 99, 1345, 1091, 1623, 114, 791, 615, 2177, 2067, 1582, 240, 1782, 1144, 1513, 1103, 166, 869, 340, 1223, 1857, 1984, 1617]. primitives [1924]. principle [1336, 1602, 1226, 1920, 1791, 1186, 990, 975, 1524]. principles [1567]. priori [1600]. PRISMA [2005]. prismatic [384]. privacy [2247]. PRNs [1280]. Probabilistic [145, 470, 214, 101, 1116, 999]. probabilities [1945, 73]. probability [1905, 1856, 1060, 517]. problem [267, 1180, 514, 2303, 625, 627, 1836, 1029, 2105, 1520, 1581, 2113, 2156, 541, 1203, 1173, 56, 1224, 2042, 155, 622, 1496, 1590, 1839, 650, 532, 104, 915, 992, 1992, 1037, 2004, 2288, 1867, 102, 562, 167, 2246, 305, 2198, 813, 1343, 978, 563, 288, 1341, 1448, 183, 1342, 1297, 719, 1166, 283, 1106, 1216, 208, 1985, 2239, 1381, 2056, 1834, 2090, 1008, 502, 1155, 1266, 1939]. Problems [631, 536, 1402, 977, 772, 1121, 2262, 1860, 782, 626, 998, 1016, 1567, 1733, 160, 2272, 1627, 1401, 1787, 696, 1987, 71, 1122, 2021, 2269, 615, 1470, 777, 1118, 952, 1307, 1586, 1790, 963, 1668, 695, 324, 2035, 646, 616, 689, 1090, 1991, 1512, 2252, 1840, 2258, 1920, 1100, 371, 486, 1242, 1481, 628, 467, 360, 900, 1556, 1089, 1095, 2219, 1907, 897, 561, 735, 947, 1326, 1522, 2085, 1191, 1160, 2074, 219, 384, 204, 504, 1083, 1982, 1459, 32, 400, 1911, 1407, 1726, 712, 1369, 867, 2095, 1312, 1006, 1024, 762, 1619, 1800, 1968, 1881, 767, 490, 1250, 1549, 1352]. problems [1432, 797, 1491, 1377, 1676, 1509, 891, 886, 1923, 1317, 1181, 53, 372, 1195, 1947, 2299, 263, 1416, 1114, 938, 933, 2119, 976, 1592]. procedure [439, 1195]. procedures [873]. process [465, 1503, 1439, 236, 2151, 1431, 1085, 1556, 1105, 665, 59, 379, 1369, 2032, 1614, 1967, 1822, 1810, 1236, 809, 395]. processes [1725, 1344, 553, 1945, 582, 1472, 1299]. processing [1987, 198, 1441, 1765, 2202]. producing [1040, 816, 1322]. product [133, 132, 84, 479, 1326, 1105, 1243, 1857]. production [1775, 1003, 1766, 454]. productivity [1838, 458, 1800]. Professor [610]. profile [1175, 2232]. profiles [1066, 824, 1986]. profitability [419]. programming [982, 576, 1371, 1870]. programs [452, 1116]. progress [2304]. progressive [1405, 479, 688]. Projectile [793]. projection [1649, 656, 1829, 1139]. projection-based [1139]. projective [2091, 846]. projectors [1719]. PROMETHEE [1699]. promising [1574]. promotes [875]. proof [439]. propagating [408, 2271]. propagation [1616, 196, 1624, 617, 2275, 655, 1325, 303, 398, 1215, 1738, 1950, 1628, 2064]. propensities [1824]. Proper [1241]. properties [961, 908, 926, 707, 492, 205, 770, 786, 2172]. proportional [531, 103, 966, 740, 1062, 27]. proportions [1743]. Proposal [1453]. proposed [1470, 1239]. prospects [4]. protection [2247]. protective [585]. protocol [1546]. proton [2155]. provinces [2206]. proximal [1991]. proximity [540]. proxy [304]. proxy-measure [304]. prudent [1965]. pseudo [2189, 573, 1373, 1254, 1154, 884]. pseudo-spectral [2189, 573, 1373, 1254, 1154, 884]. Pseudoconvex [1592]. pseudoinverse [847]. pseudoinversion [1780]. pseudomonotone [976]. Pseudoprimality [970]. pseudospectral [220, 1586, 1559, 1790, 2082, 1792]. PSO [1092, 326, 25]. PSO-tuned [1092]. ptychographic [208]. public [2209]. pull [1679]. pulse [1624, 762, 398]. pulsed [402]. pump [355]. purchasing [1310]. pure [1216]. push [1679]. push-pull [1679]. puzzle [310]. PV [419, 14, 12, 33, 10, 16, 423, 2202, 8]. PV-system [16]. PV-wind [33]. PV/T [12]. PWM [438]. Pythagorean [2118, 2164]. Pythagorean-hodograph [2118, 2164]. Python [352].

q [2243]. q-HAGTM [2243]. QSC [163]. Quadratic [935, 462, 2131, 191, 2103, 2257, 2137, 1999, 1910, 552, 2175, 732, 1076, 377, 951, 576, 797, 1371, 1917, 2185, 1870]. Quadrature [1539, 1235, 1612, 149, 511, 1944]. Quadrature-free [1539]. quadrotor [578]. quadruple [1177]. Qualitative [1831, 1526, 74, 540, 1802, 1289]. quality [1746, 2158, 184, 30]. Quantifying [1013]. quantile [600, 1043]. quantiles [2172]. quantitative [1802, 1521]. Quantization [1783]. Quantized [1104, 1017].

Quantized-observer [1104]. Quantum [272, 279]. Quantum-like [272]. quarantine [937]. Quasi [518, 1050, 493, 1719, 2036, 1811, 1717, 132, 191, 769, 575, 1712, 1470, 1945, 1998, 234, 130, 1465, 2166, 2085, 289, 2040, 1338, 1381, 392, 1362, 1062, 1910]. quasi-birth [1945]. quasi-contractive [1470]. Quasi-interpolant [518, 191]. quasi-interpolants [289, 2040]. quasi-interpolating [1719, 1811, 1717]. Quasi-interpolation [1050, 2036, 132, 769, 1712, 1998, 234]. quasi-Monte [392, 1910]. quasi-pentadiagonal [575]. quasi-reversibility [1381]. Quasi-synchronization [493, 1338, 1362]. quasi-Toeplitz [2166]. quasi-uniform [1465, 1062]. quasi-velocities [130]. quasilinear [627]. quasilinearization [161, 1253]. quasistatic [624]. Quaternion [1697, 737, 162, 1873, 2089, 1000]. quaternion-valued [162, 1873, 1000]. Quenching [2105, 223, 295]. question [1185]. queue [1085, 1368]. Queueing [1292, 709, 1441, 1551, 529, 2066, 1430, 2080]. Queuing [983]. quintic [1624, 323, 527, 1612].

R [1319]. R&D [971, 1677]. Rach [532]. Radial [368, 1409, 860, 467, 149, 906, 1561, 444, 166]. Radiant [175]. radiation [1490, 358, 741, 1687, 329, 1830, 492, 1703, 837, 1750]. radiative [968, 1875, 948, 779, 82, 1460, 671, 951]. radiative-dissipative [779]. Radim [610]. radius [1400]. Railway [1731, 450]. rainfall [181, 95]. ramp [385]. Random [344, 558, 1555, 1122, 1650, 55, 997, 2300, 63, 924, 354, 1143, 1369, 853, 1299, 1820, 1713, 2138, 2191]. randomly [96]. range [749, 177]. rank [407, 1823]. ranking [1383, 1580]. Raphson [1774]. rapid [1974]. rapidly [511]. ratcheting [597]. rate [968, 1482, 196, 1720, 727, 833, 748, 1997, 937, 1143, 816, 205, 941, 1656, 857, 148, 1617]. rates [1264, 1752, 39, 1541, 369, 1207]. rating [2202]. ratio [442, 1284, 222, 965]. ratio-dependent [965]. rational [964, 1586, 888, 845, 552, 1614, 1187, 1253]. rational-expansion-based [1187]. Raton [592]. Rayleigh [778, 1948, 1661, 683]. RBF [190, 1168, 371, 836, 1858]. RBF-based [1858]. re [1334]. re-attempt [1334]. reaction [1045, 775, 390, 1238, 2217, 995, 1708, 1554, 343, 1306, 1797, 104, 1952, 486, 79, 1261, 2130, 1036, 401, 925, 1379, 2255, 563, 2013, 1410, 298, 1629, 1422, 372, 862, 1255, 1844, 2025, 1938, 1628]. reaction-subdiffusion [2217]. reactions [44, 1049, 855, 537]. reactive [1129, 671, 1732]. reactivity [2076]. reactor [1175, 144]. ready [287]. ready-mixed [287]. real [653, 796, 1555, 1838, 991, 451, 28, 1414, 1301, 83, 867, 265, 2015, 1747, 2154]. real-time [451, 28, 83, 1747]. real-world [991]. realistic [1794]. realization [724]. Realizing [549]. reassignment [195]. reboot [2261, 1292]. reciprocal [1823]. reciprocally [1508]. recognising [1924]. reconfiguration [88]. Reconstructing [2274]. Reconstruction [536, 1866, 1013, 1042, 190, 192, 662, 1660, 56, 2208, 621, 1139]. recovery [2061, 2144, 557, 9, 941, 1292]. rectangular [142]. Recurrent [494]. recursive [847, 1918]. recycling [1001, 1105]. red [1221]. Redox [417]. reduce [287, 228]. Reduced [1406, 1382, 890, 1354, 2181, 217, 581]. Reduced-order [1406, 1382, 581]. reduction [85, 918, 1536, 2280, 840, 1197, 1910, 78, 20, 1205, 1128, 2185, 1354]. reductions [824]. redundant [1287, 1264, 1752, 2022, 2216]. reference [1302, 30, 276, 758, 955, 2231, 1943]. referred [643]. refinement [2303, 1932, 2065, 1926, 76, 2082]. Refining [1791, 197]. Refocusing [621]. reformulation [1252]. refuge [1862, 270, 1172, 1504, 1919, 1832, 1988, 1492, 691]. refurbishing [1105]. regime [1345, 258, 2115, 1103, 340, 1967, 1984, 1617]. regime-switching [1345, 1103, 340, 1984]. regimes [530, 1713]. region [12, 426, 2141, 1689, 79, 1852, 921]. region-specific [1852]. Regional [463, 899]. regions [1710, 2030, 2236]. regression [1969, 2006, 600, 1910, 520, 779, 2158, 310, 322, 1785]. Regula [1445, 1112]. Regular [2034, 1999]. Regularization [2090, 2035, 873, 1985]. regularized [956, 842, 2006, 1845]. regulation [2168]. regulations [1231]. regulator [1620]. regulatory [1692, 692]. rehabilitation [718, 1464]. reinfection [1839]. reinforced [947]. Reinforcement [1966, 848, 1852, 1276]. reinsurance [1345, 1856]. rejection [1351]. relapse [1831]. related [970, 1806, 1945, 453, 1722, 1359]. Relations [782]. relationship [182, 97]. relaxation [859, 1554, 2140, 352, 1543]. Reliability [1264, 1752, 1405, 1385, 1138, 666, 1948, 4, 218]. Reliable [1740, 730, 2256, 216, 1426, 481]. reluctance [20]. remainder [1573]. remodeling [641]. removing [2277]. Renewable [2, 1978, 1958, 1587, 4, 412]. renewal [1879]. repair [1264, 1752, 1342, 1834]. repairable [1813, 2261, 929, 1188, 2080]. reparametrization [958]. repeat [1563]. Repeated [1785]. replenishment [1293]. Replicating [1675]. replicator [2197]. Reply [389]. report [2263]. reporting [173]. Representation [2265, 1717, 1787, 194, 1598]. representations [1077]. represented [701]. reproduction [202, 721, 857]. Research [724, 464, 317]. reservoir [1886]. reservoirs [1803]. residential [419, 1767, 2232, 2058]. residual [1026, 2280, 295, 1497, 645]. residue [228]. resilience [2054]. resistance [2176, 1766, 601]. resistances [414]. resistive [1725, 514, 517]. resizing [835]. Resolution [2037, 1079, 1473, 1057, 2021, 780, 1712]. resolved [383]. resonance [1661]. Resonant [545, 1163]. resonator [441]. resource [1771, 1906, 1678, 1396]. resources [2020, 2057, 760, 1425, 1738, 1244, 510]. respect [822]. respiratory [1850, 1621]. response [570, 875, 1367, 1044, 2061, 1824, 1172, 546, 1096, 452, 1667, 137, 1949, 916, 601, 1099, 1953, 1810, 1798, 951]. responses [1591, 603, 547, 1199]. Responsibility [1686]. restart [1869]. restoration [2238]. restore [1284]. restricted [1992]. restrictions [1521]. results [785, 531, 2289, 1845, 469, 966, 532, 1286, 22, 981, 1633, 1470]. retailer [397]. retaining [1189]. retrial [2261, 1342, 2216, 1430, 1195]. retrieval [208]. retroviral [922]. return [907]. returns [2108]. reversal [621]. reversibility [1381]. Reversible [2247]. reversion [1893]. reverting [809]. Review [4, 592, 252, 894]. revolution [1064]. reweighted [1784, 288]. rework [324]. Reynolds [892, 686, 1002]. rhinoceros [225]. Riccati [2088, 996]. Rich [1952, 1833, 1492]. Richards’ [1721, 2088]. Richardson [1306, 1111]. Ricker [2106, 2001]. Ricker-type [2106]. Riemann [1589, 899]. Riemannian [2299, 2185]. Riesz [1318, 995, 1807, 1052, 373]. right [2128, 763]. right-censored [763]. rigid [196, 874, 1509]. ring [2226, 367, 1819]. ripple [444]. Ripplet200b [321]. rise [1722]. risk [1771, 1014, 62, 1319, 1679, 647, 157, 2172, 1723, 1876, 693, 287, 2236, 178, 1435, 148, 1671, 1014]. risk-averse [62]. risk-based [1771]. risk-sensitive [1876]. risks [1582, 763, 688]. River [1826]. RKL [348, 389, 1]. RLW [813]. RNA [2284]. road [367, 93]. Robin [1045, 167, 400]. robot [1553, 2128, 950, 716, 83, 803, 2060, 474]. robotic [1287]. robots [1735, 718]. Robust [470, 109, 1260, 602, 1588, 1286, 354, 2168, 1898, 703, 122, 1560, 474, 820, 896, 71, 300, 1031, 238, 1497, 1359, 173, 989, 2027, 1879, 718, 5, 1341, 1448, 1125, 1792, 1225, 34, 1410, 981, 2179, 2092, 1220]. Robustness [88, 658]. rock [2136]. rock-glacier [2136]. rod [408, 1056]. Rodrigues [2164]. Roe [2037]. Roesser [1917]. Rogue [1259, 328, 1670]. Role [721, 1664, 230, 1703, 1288, 1388, 2031, 334, 1127, 1667]. roll [2254]. rolling [2254]. root [1940, 1445]. root-finding [1940]. roots [394]. Rose [974]. Rosenau [813]. Rosenau-RLW-Burgers [813]. Ross [1967]. Rosseland [329]. Rössler [1198]. rotated [1171]. rotating [889, 761, 1304, 273, 44, 1977, 1981, 1778, 787, 1703, 951, 235, 266, 837, 1750]. rotation [487]. rotational [643, 560]. rotatory [874]. rotor [19, 2027]. rotors [1770]. rough [85, 484]. Route [1073]. routes [223, 783]. routing [978]. RRE [896]. RSCSCS [1052]. rubella [265]. Rubik [503]. Ruin [1856]. rule [2128, 287, 309]. rules [1324, 705, 654]. rumor [1711, 1210]. rung [1691, 2093]. Runge [1494, 1307, 1262, 1574, 158, 277]. runway [183]. RVM [30].

S [1280, 11, 269, 974, 2179, 1220]. S-box [1220]. S-Boxes [2179, 1280, 269]. Sabin [1746]. safety [1731]. Said [840]. sailing [2007]. sale [1759]. Salerno [1595]. sales [99]. salesman [1549]. salinity [2248]. salp [361, 994, 1063, 1599]. salt [2126, 1689]. sample [529, 1248]. Sampled [1025, 973, 667, 1449, 1656, 974, 1861, 2256, 872, 1882]. Sampled-data [1025, 667, 1449, 1656, 974, 1861, 2256, 872]. sampled-data-based [1882]. samples [1905, 1540]. Sampling [94, 600, 1815, 73].

Sampling-based [94]. sand [1911]. sandwich [1696]. SARS [1058, 1087]. SARS-CoV-2 [2083, 1058, 1087]. satisfaction [1073]. satisfying [1602]. Satsuma [2301, 1309]. saturable [2271]. saturated [1745, 1720, 829, 1039, 1796, 1865, 1793, 753]. saturation [1260, 1031, 2206, 1499, 1943]. Saudi [182]. SAV [471, 1795]. Sawada [1023]. SBA* [2068]. SBVPs [927]. SC-Grid [515]. scaffolds [1631]. scalable [6]. scalar [697, 1649]. scale [724, 168, 687, 260, 863, 1175, 2108, 1971, 703, 2160]. scale-free [168, 1971, 2160]. scales [1702]. scaling [1870, 1445]. scam [1210]. scarce [2026]. scattered [1926]. scattering [2099, 1866]. scenarios [585]. schedule [2080]. scheduling [2156, 324, 420]. scheme [1625, 1547, 1082, 558, 1606, 241, 1318, 1009, 870, 575, 2217, 330, 1836, 1147, 799, 2059, 1137, 1403, 1035, 1132, 1386, 823, 1899, 1871, 1845, 1566, 1226, 1640, 697, 1795, 2037, 552, 2247, 1995, 1588, 399, 2178, 628, 246, 1373, 1254, 1541, 2276, 1277, 2284, 673, 291, 2207, 2194, 167, 1833, 1608, 1305, 1041, 925, 1267, 319, 797, 884, 1222, 1491, 1931, 1377, 1901, 209, 1111, 373, 502, 1155, 347, 2063, 2287, 1353, 815, 1372, 1321, 1673, 1730, 1950, 2245, 1524, 1885, 1220]. schemes [2148, 1624, 952, 1193, 1660, 1281, 904, 730, 202, 927, 957, 1411, 2192, 544, 591, 1191, 2190, 535, 765, 1142, 1573, 383, 1923, 331, 2062, 1954, 488, 1196, 1814, 1543, 1718, 1843]. Schiff [1170]. Scholes [1623, 114, 2274, 409]. Schrödinger [1163, 1318, 768, 1132, 823, 1239, 471, 1281, 1807, 2081, 1640, 1262, 156, 527, 1259, 299, 2174, 1047, 1373, 2115, 1495, 1956, 831, 923, 1161, 1291, 303, 1561, 398, 2239, 2173, 2063, 1321, 46]. Schwarz [1554]. Science [1749]. Sciences [609, 189]. Scientific [189]. score [2093]. Scott [2207]. scrambled [926, 392]. scrambling [1075]. screening [308]. scroll [1183]. scrutinizing [2143]. SCW [1016].

SCW-iterative-computational [1016]. SDEs [558, 1915]. Sea [1742]. SEAIQR [2160]. search [1173, 2219, 816, 712, 1312, 1181, 930, 1364, 817, 446, 1475]. seasonal [1862, 2163, 1667, 920]. Second [1885, 1494, 2262, 1606, 560, 566, 727, 777, 1899, 527, 551, 1577, 735, 1407, 1498, 1573, 120, 576, 1341, 1250, 2013, 1509, 886, 1299, 990, 1245, 502, 1737, 1150, 2057]. Second-order [1885, 1494, 777, 1577, 735, 576, 1341, 2013, 1509, 886, 1299, 990, 1150]. section [776]. sector [1744]. Secure [1220, 973, 1651, 2179]. security [83, 1819]. sediment [841]. seepage [2222]. Segel [2240, 2077]. Segel-Flow [2240]. segmentation [81, 1816, 2267]. Segmented [919, 6, 1975]. SEICR [1174]. SEIHDR [2011]. SEIR [1740, 856, 1855, 1665, 937, 858, 1374, 67, 499]. Selangor [774]. selecting [310]. Selection [1684, 42, 1699, 1962, 1691, 1794, 2093, 1931, 1580, 1599, 2097, 2283]. selective [437]. self [2051, 1789, 2053, 2058, 303, 886, 2298, 153]. self-consumption [1789, 2058]. self-similar [2053, 153]. self-triggered [2051, 2298]. selfish [978]. semi [1878, 12, 892, 2248, 1696, 1306, 1175, 1541, 86, 1854, 481, 853, 811, 298, 1081, 53]. semi-analytical [1878, 1854]. semi-arid [12]. semi-closed [86]. semi-empirical [1175]. semi-implicit [2248]. semi-infinite [1696]. semi-Lagrangian [892, 1081]. semi-linear [1306, 811, 298, 53]. semi-Markovian [1541, 481, 853]. semiabsolute [991]. semilinear [133, 602, 305, 274, 1429, 1903, 277]. semimartingale [1246]. Senegal [1980]. sense [542, 1467]. sensing [2228]. sensitive [1876, 355, 516]. sensitivities [422]. Sensitivity [78, 425, 961, 123, 912, 365, 1975, 56, 1701, 1058, 80, 304, 621]. sensor [31, 429, 655, 853, 675, 1659]. sensorless [443]. sensors [1585, 704]. separated [1799, 1290]. September [528, 885, 1332, 1709, 2169]. septic [1023]. sequence [2156]. sequence-dependent [2156]. sequences [2028, 698, 970, 926]. sequencing [241]. sequential [695, 767, 576, 1371]. sequestration [2122, 2175]. Series [863, 436, 1725, 834, 1235, 2114, 453, 1497, 1766, 2202, 1610, 2265, 1741, 1176, 2070]. serotonin [871]. serotypes [1274, 585]. server [709, 1085, 1292, 529, 1430]. servers [136, 1368, 2066]. service [709, 1441, 1085, 907]. services [1363]. servo [2157]. set [85, 1206, 1587, 2130, 1726, 630, 1282, 1138]. sets [1625, 484, 1930, 332, 1614]. setting [1207]. settling [391]. settling-time [391]. setup [2156, 1590]. several [1940, 1024]. severely [2206]. severity [2223]. sex [308]. sex-structured [308]. SH [1696]. SH-wave [1696]. shallow [2037, 2244, 1215, 2290]. shallow-water [2244]. Shannon [1383]. Shape [1655, 2178, 539, 2021, 777, 2076, 561, 1996, 1290, 346]. Shape-preserving [2178]. shaped [2189, 1483, 55, 1405]. shapes [86]. shaping [421]. shareholding [1679]. sharing [245]. Shark [1840]. sharp [112]. sharp-interface [112]. shear [480]. shear-induced [480]. sheet [358, 779, 959, 534, 492, 515]. Shepard [1737]. Shephard [1542]. shift [656, 1349]. Shifted [931, 738]. shifts [1429]. Shilnikov [1198]. shirking [358]. shock [2015]. shop [2156, 324]. shore [1960]. Short [1978, 1767, 424, 2230, 113, 821]. short-circuit [424, 821]. Short-Term [1767, 1978, 2230]. shunting [1698]. SI [746]. SiC [426]. sidewalk [245]. sideways [1985]. sight [70]. Signal [557, 1987, 2275]. signals [195]. signature [24]. signatures [821]. signed [589]. Significance [1466, 761]. significant [741, 2139, 770]. signum [1837]. SIHR [1997]. Silica [733]. silico [250, 1818]. silicon [1415]. similar [2053, 153]. similarity [1171, 1587]. simple [1445, 178, 1837, 333, 579]. simplicial [2033, 2291]. simplified [123, 1762]. Simplifying [1101]. simply [477]. Simul [1750]. simulate [155, 355]. simulated [972, 1126, 60]. Simulating [1455, 1209, 1649, 9, 1672, 1535, 112, 924]. Simulation [348, 510, 15, 189, 597, 1752, 146, 1472, 22, 494, 389, 1751, 908, 643, 761, 1542, 968, 1482, 890, 1932, 318, 1657, 1650, 1941, 1398, 2275, 174, 1279, 311, 513, 1287, 1812, 1742, 707, 541, 300, 324, 582, 451, 1640, 1048, 1946, 2300, 1565, 385, 238, 1874, 487, 1641, 1149, 1830, 2140, 59, 1608, 379, 340, 1370, 1379, 434, 2190, 356, 733, 565, 771, 2044, 1903, 1422, 540, 430, 180, 1976, 266, 1738, 412, 2025, 2057]. simulation-optimization [300]. Simulations [142, 352, 619, 2176, 77, 71, 776, 1067, 1906, 480, 260, 877, 2175, 1030, 987, 1210, 1747, 1511, 177, 61, 585]. simultaneous [496]. simultaneously [1386]. sinc [301, 648]. since [1831, 1420]. sine [133, 588, 573, 1827, 906, 1370, 1652, 2159, 373]. single [709, 2281, 414, 423, 1188, 183, 529, 1825, 1732, 872, 145]. single- [183]. single-diode [1825]. single-master [872]. single-phase [1732]. single-server [529]. Singular [849, 2036, 623, 1669, 1874, 2181, 561, 482, 1728, 765, 1078, 506, 1491, 1225, 1633, 887, 1361, 2150, 648, 1843, 1150]. singularities [1712]. Singularity [2277, 1127]. Singularity-removing [2277]. singularly [1079, 839, 694, 860, 602, 486, 900, 1989, 735, 167, 1639, 1429, 1968, 563, 1341, 1448, 1166, 797, 1378, 502, 1155, 933, 1939]. sink [1703]. sintering [1439]. sinusoidal [864]. sinusoidal-polynomial [864]. SIR [2279, 1066, 1077, 727, 829, 104, 405, 2041, 1971, 1236, 1244]. SIS [1584, 1319, 2191, 1071, 572]. SIS/R [1319]. SISO [435]. sites [1185]. situation [309]. Sivashinsky [1612]. six [18, 384]. six-node [384]. six-phase [18]. sixth [2272, 1666, 1316]. sixth-order [2272, 1316]. size [529]. sized [1631]. sizes [236]. sizing [1760]. skeleton [29]. skew [768, 29, 863, 1275]. skew-Hermitian [768]. skew-normal [863]. skew-symmetric [1275]. skill [2252]. slave [1086, 231, 872]. Sliding [1153, 16, 1963, 1642, 1553, 781, 306, 229, 1778, 1841, 578, 1698, 1379, 351, 955, 122, 1308, 491]. Sliding-mode [1963]. slime [867]. slip [625, 622, 1496, 461, 1687, 779]. slope [1744]. slow [1877, 233]. Smale [1700, 2125]. small [132, 1373, 1429, 564]. small-world [564]. Smart [10, 4]. Smith [1786, 758]. smoking [1665]. Smooth [1462, 1588, 1053, 1158, 1863, 310, 1339, 683, 1916, 998]. smoothed [1119, 200]. smoothing [192, 756]. smoothness [2178]. snapshot [890]. Sobol’ [80]. Sobolev [1285, 1041, 1903]. Sobolev-type [1285]. SoC [451]. SoC-based [451]. Social [782, 2031, 290, 1946, 1575, 1852, 928, 268, 2302]. socio [1230]. socio-economic [1230]. socioecological [1972]. soft [1365, 484, 972, 1037, 2216]. Software [1585, 2134, 1067, 218]. soil [2135, 179]. solar [465, 15, 12]. solid [1304, 384, 366, 1216, 112]. solitary [348, 1, 389, 669]. Soliton [1163, 1239, 2081, 845, 957, 1240, 2286, 750, 1309]. Solitons [299, 1056, 1259, 2174, 1240, 1956, 556, 303, 2271]. Solow [2152]. solute [1333, 710]. Solution [940, 1125, 1299, 838, 1993, 619, 216, 1719, 1616, 380, 1454, 625, 1821, 796, 318, 1503, 2136, 1409, 1122, 677, 1520, 325, 343, 1845, 2042, 616, 845, 156, 650, 332, 1213, 1995, 551, 1612, 1588, 1145, 2166, 1167, 1289, 2087, 1516, 671, 329, 2041, 813, 1407, 322, 207, 1161, 259, 951, 308, 1881, 100, 75, 2014, 996, 161, 283, 1630, 1222, 1225, 891, 613, 648, 502, 533, 1737]. solutions [653, 979, 388, 720, 1163, 2189, 348, 462, 302, 870, 1283, 501, 964, 1895, 566, 2159, 1239, 1559, 323, 646, 328, 2081, 2301, 532, 1808, 2171, 957, 824, 1641, 1350, 1973, 735, 1170, 1522, 2116, 1595, 1525, 651, 762, 2286, 2265, 1967, 286, 2251, 408, 706, 1, 389, 1361, 750, 669, 1339, 2290, 938, 1885, 210, 896]. Solvability [1589]. solve [85, 1562, 538, 1606, 782, 2036, 1469, 1836, 1171, 1182, 2252, 1497, 1148, 1886, 1475, 1380, 1549, 1061, 1116, 1253, 1316]. solved [224]. solver [658, 1627, 696, 1795, 1565, 2195, 1879, 1619, 208]. solvers [1987]. Solving [1821, 1733, 1401, 2156, 945, 232, 1005, 482, 1326, 2085, 274, 1933, 1113, 1592, 1494, 977, 772, 1547, 1860, 1771, 1333, 1016, 132, 575, 842, 243, 1983, 244, 615, 1739, 1666, 880, 1998, 1586, 1559, 1790, 963, 320, 780, 324, 56, 2035, 2128, 49, 689, 622, 1090, 1590, 1991, 904, 1512, 1840, 1463, 527, 836, 1251, 1242, 467, 711, 1556, 1989, 1574, 246, 1095, 1539, 561, 544, 1134, 1827, 1191, 1648, 504, 1728, 934, 1911, 1312, 1006, 1639, 1658, 759, 259, 1829, 1727, 120, 1297, 719, 596, 1106, 1509, 298, 1317, 1371, 739, 1381, 717, 2090, 271, 263, 1416, 1339, 488, 1266, 1179]. solving [2119, 336, 2245, 1501, 2264, 843, 976]. Some [1567, 1984, 2062, 2038, 1363, 784, 1980, 1509, 1923]. Soret [237]. Soret-Dufour [237]. sorting [991]. sound [560]. source [465, 1783, 841, 358, 925, 1703, 1772, 2015, 436, 2090]. source/sink [1703]. sources [2208, 179]. sourcing [1008]. SP [1614]. SP-iteration [1614]. Space [1439, 599, 390, 1318, 1013, 1550, 77, 583, 84, 780, 1189, 1252, 349, 1173, 1807, 573, 697, 326, 381, 1504, 1460, 1827, 2208, 2255, 1429, 259, 1924, 1052, 1792, 1225, 1600, 1693, 1208, 1382, 373, 1364, 1278, 1353, 2120, 1638, 2127, 1950, 1434]. space-dependent [2208]. space-fractional [390, 697, 1827, 1208, 373, 1278]. space-time [2255, 1950]. spaces [71, 1470, 1046, 705, 2137, 207]. Spalart [1675]. Spanish [834]. sparse [194, 199]. Spatial [1423, 754, 902, 1620, 1132, 752, 1994, 1795, 1796, 1141, 865, 706, 862, 1396, 1953, 1621]. spatial-dependent [865]. spatially [501, 611, 535, 1178, 2092, 1864, 1890]. Spatio [1051, 1798, 2018]. Spatio-temporal [1798, 2018]. Spatiotemporal [914, 1678, 547, 2260, 2275, 1001, 1249, 755, 1613]. SPDEs [1213]. Special [510, 2, 172, 434, 412, 1494, 609, 1302, 276, 1412]. species [668, 1363, 1151, 577, 1701, 690, 1667, 1988, 1492, 259, 604, 1736, 691, 2170]. specific [1067, 1852]. Spectral [1454, 1867, 217, 979, 1440, 2189, 1194, 1469, 918, 956, 2285, 2159, 1455, 1721, 1586, 1845, 573, 360, 711, 2273, 1373, 1254, 1154, 2074, 1728, 1459, 884, 921, 1634, 1600, 500, 1947, 2297, 1339, 46, 2220]. spectral-conjugate [979]. spectrogram [195]. spectrum [1595]. speed [23, 583, 124, 732, 757, 817]. SPH [1313, 146, 896]. sphere [761, 787, 366]. spherical [1287, 1395, 82, 640, 921, 1633]. spheroidal [1804]. spider [935]. spider-inspired [935]. spillover [960]. spillovers [971, 1677]. spin [487]. spinning [105]. Spiral [1049]. spline [1719, 2036, 278, 160, 1787, 1926, 1969, 1620, 284, 1623, 1998, 1252, 2137, 156, 527, 552, 1874, 1411, 1576, 377, 519, 1933, 207, 2040, 1724, 1816, 478, 797, 593, 1557]. splines [838, 2262, 1811, 1717, 769, 1932, 193, 192, 521, 1023, 1612, 289]. Split [2112, 1556, 1902, 1772]. Split-step [2112, 1902]. splitting [7, 583, 842, 1046, 1714, 2180, 1995, 1373, 1254, 1827, 204, 990]. SpMV [1481]. spot [1686]. spread [123, 856, 1975, 1302, 1210, 1510, 1927]. spreader [8]. spreading [1711, 268, 586]. spruce [706]. square [2269, 311, 473]. squares [1784, 1167, 360, 322, 1139, 1246]. squares-support [322]. Stability [220, 1117, 96, 284, 223, 647, 2113, 580, 783, 752, 290, 1226, 1803, 108, 1397, 2165, 162, 1157, 2199, 393, 280, 1892, 1169, 386, 2160, 67, 755, 1499, 370, 1704, 653, 1710, 485, 560, 1508, 922, 2148, 1483, 1413, 109, 971, 1692, 1347, 1708, 785, 1436, 611, 1260, 1521, 748, 229, 1876, 966, 203, 1797, 489, 877, 2053, 1058, 214, 1376, 1532, 1596, 1149, 2194, 1526, 1449, 1656, 1087, 692, 128, 106, 1573, 969, 1873, 604, 1322, 1311, 1349, 1611, 916, 101, 555, 2296, 111, 209, 981, 537, 676, 93, 2079, 572, 2295, 1065, 1399, 1744, 2063, 1015, 2231, 275, 901]. Stability [323, 2087, 655, 331]. Stabilization [2123, 1861, 1938, 1117, 96, 391, 989, 1449, 1656, 604, 996, 1000]. stabilized [1097, 2114, 1795, 628, 906, 1881, 862, 282, 2175]. stabilizer [1261, 2056]. stabilizer-free [2056]. Stable [303, 1196, 538, 1860, 1606, 448, 103, 635, 399, 2192, 2276, 335, 2077, 1250, 1792, 1476, 1954, 2287, 1543, 1718]. Stackelberg [936, 397]. stacking [1218]. stage [709, 971, 2103, 2024, 1085, 1677, 1492, 1773, 1388, 1762]. stage-structure [1492]. stage-structured [1388]. stages [746]. stagnation [761, 1454, 79]. stand [18]. stand-alone [18]. standard [265]. star [55, 1769]. star-shaped [55]. start [1960]. STATCOM [229]. State [27, 1117, 1692, 1385, 1203, 966, 740, 1541, 354, 819, 1450, 452, 1291, 2022, 2052, 93, 1382, 2070, 2295, 1798, 1179, 1943, 2088]. state-dependent [819, 2070, 2295]. state-feedback [1117]. state-space [1382]. states [1994, 1622, 1656, 499, 1199, 1414]. Static [2251, 391, 1119]. station [450, 1763, 449]. Stationary [1315, 674, 953, 696, 1087, 235, 1064]. Statistical [1028, 1964]. Statistically [656]. statistics [520]. stator [24, 821]. steady [625, 1994, 354, 1377, 1798, 1179, 282, 2264]. steady-state [354, 1798]. steel [295]. steepest [2203]. steered [1056]. Stefan [1122, 1676]. steganography [2139, 1842]. stem [1925]. Stencil [1534]. Stenger [511]. stenosed [346]. stenosis [1466]. stenotic [61]. stent [60]. Step [2045, 1494, 1625, 1983, 2128, 1154, 1902, 354, 1791, 2292, 579, 2112, 1816, 1718]. Step-by-step [2045]. stepped [612]. stepping [49, 1931, 1339]. stepsize [1991, 1556]. stick [622, 1496, 803]. stiff [1797, 274, 2241, 1619, 811]. stirred [144]. Stochastic [365, 1094, 242, 240, 871, 1143, 1498, 1971, 1909, 1071, 676, 1374, 795, 902, 1584, 1452, 129, 109, 1265, 1692, 1044, 2103, 2177, 699, 2067, 1783, 1896, 674, 1107, 382, 1893, 1451, 2151, 1512, 2252, 1582, 1809, 1920, 1211, 898, 1213, 39, 2123, 258, 546, 1877, 1506, 1902, 1017, 1886, 245, 1040, 1450, 357, 1664, 673, 2207, 1348, 1990, 651, 1832, 1327, 379, 1223, 1087, 1291, 565, 603, 1873, 2112, 308, 649, 1311, 2282, 700, 1169, 916, 1207, 1061, 1222, 101, 1338, 331, 895, 872, 1984, 1822, 1416, 1601, 1810, 1917, 1012, 1236, 1025, 2298, 41, 1467, 1471, 809, 901, 1199]. stochastically [549, 1955]. stock [1013, 1231, 2142]. stockless [1863]. Stokes [620, 696, 71, 2114, 1252, 622, 1655, 1496, 836, 2192, 1648, 1564, 1694, 282]. Storage [1775, 1759, 1773, 35, 412]. stories [2142]. strain [408]. strains [1745]. strang [842]. strangle [1893]. strategic [1334]. strategies [2280, 1908, 76, 622, 2242, 1856, 1397, 1575, 1852, 1230, 937, 1768, 742, 438, 1293, 1274, 2206, 555, 1128, 1181, 572, 585]. strategy [1679, 2163, 1118, 1340, 1197, 449, 222, 732, 1053, 110, 1331, 2027, 718, 1773, 794, 1869, 1464, 1063, 2060]. strategy-based [732]. stratified [1039]. Straughan [196]. Streeter [102]. strength [666, 1948, 2187]. Strengthening [173]. stress [1303, 1315, 295, 666, 1948, 688]. stressed [1696]. stretchable [105]. stretching [908, 358, 779, 959, 534, 492, 800, 266]. stretching/shirking [358]. strings [451]. strip [619]. stripe [1240]. stripes [1818]. stroke [32]. Strong [203, 1797, 1915, 1322, 2001, 383, 1220]. strongly [1084, 1609, 976]. Structural [1032, 1171, 137, 1320]. Structure [1132, 591, 1305, 581, 41, 625, 1627, 618, 624, 471, 1566, 1441, 1262, 2294, 1154, 2116, 184, 1492, 1152, 801, 60, 264, 729, 1321, 146]. Structure-preserving [1305, 581, 471, 1566, 1262, 1154, 1152, 264, 1321]. structured [1993, 382, 866, 1796, 2029, 308, 649, 1388, 1980, 1377, 1425, 1638]. structures [754, 321, 1239, 1631, 1030, 137, 1819]. Stuart [530]. student [317]. Studies [690, 86, 361, 1688]. Study [216, 273, 1238, 236, 1696, 438, 710, 834, 247, 448, 129, 1347, 2059, 1720, 2121, 550, 823, 1981, 17, 744, 1994, 1197, 1247, 1323, 1722, 917, 2129, 992, 877, 1263, 1414, 1337, 1828, 1460, 1956, 404, 1191, 1034, 554, 250, 219, 1457, 22, 1955, 1854, 1188, 923, 2049, 366, 770, 543, 1002, 209, 2154, 2079, 920, 821, 1900]. Studying [1304, 1980]. Sturm [1180, 2262, 1224, 1326, 1522, 490, 886]. style [2043]. sub [388, 1336, 343, 888, 1400, 535, 1245, 1239]. sub-diffusion [1336, 343, 888, 535, 1245]. sub-domains [1400]. sub-equation [388]. sub-ODE [1239]. subcritical [802]. subdiffusion [2217, 799, 163]. subdivision [202, 552, 2178, 2194]. subfamilies [345]. subgrid [2244]. subject [129, 365, 2103, 1260, 94, 1876, 1031, 2022]. subjected [1304, 677]. submerged [1455, 1973]. subsidies [1677]. subsidy [1091]. subsoil [613]. subspace [695, 1931, 2092]. substrate [595]. substrate-based [595]. succession [2097]. suction [741]. Sugeno [27, 1861, 481, 981, 701]. sum [1963, 934, 2171]. summarization [464]. summation [1235]. super [769, 955]. super-convergent [769]. super-twisting [955]. supercapacitor [436]. Supercloseness [1261]. Superconvergent [2273, 289]. supercritical [802]. superior [407]. superlinear [1445]. superpopulation [512]. superposition [1240]. Supervised [1493]. supervisor [422]. supplier [397]. supplier-Stackelberg [397]. Supply [1310, 559, 1091, 62, 791, 1679, 693, 1677, 1105, 2250, 1143, 1768, 1243, 99, 907, 397]. support [173, 322, 458]. supported [477, 1187]. supporting [174]. suppress [967]. Surface [951, 643, 1746, 2096, 1463, 1649, 461, 741, 879, 1963, 515, 1996, 1816, 1535, 1931]. surface-based [1963]. surface-groundwater [1649]. surfaces [2040, 1954]. surfactant [1853]. surge [1857]. Surrogate [747]. surveys [512, 1824]. susceptibility [1523]. susceptible [1058, 1374, 746]. suspended [959]. suspension [679]. sustainability [1896, 2045]. Sustained [1642]. Sutterby [1750, 837]. SVR [1516]. Swan [2152]. swaps [39, 240]. Swarm [1106, 782, 1197, 1160, 1190, 1911, 1735, 361, 307, 1216, 994, 584, 1276, 1248, 249, 1063, 1599, 2119]. swarm-based [782]. swarming [1928]. switch [1413, 128]. Switched [1882, 973, 819, 1963, 20, 2282, 495, 1351, 1659]. switched-type [973]. switches [1713]. switching [1345, 1725, 437, 580, 382, 258, 1506, 1541, 1124, 357, 1596, 280, 1103, 1457, 340, 1967, 1288, 2052, 676, 1351, 1984, 1817, 226]. switching-based [2052]. swiveling [70]. Sylvester [2128, 1384, 271]. Symbolic [1239, 1595, 392]. symmetric [919, 1326, 1134, 1387, 406, 1275]. symmetrical [555]. symmetrically [1304]. symmetries [918]. symmetrized [1344]. Symmetry [824, 2116, 82, 599]. Symplectic [1467]. Synchronisation [1462]. synchronizability [168]. Synchronization [1593, 1136, 1086, 2091, 285, 737, 1702, 231, 138, 1521, 469, 1286, 549, 1053, 1506, 1017, 1579, 2018, 974, 1698, 1219, 1873, 396, 2282, 670, 2109, 853, 878, 1468, 1338, 351, 1601, 493, 1183, 1025, 2187, 1817, 1032, 589, 846, 1362, 1062, 1461, 2089]. synchronous [442, 1770, 26, 443, 444, 24]. synergetic [469]. synergy [1731]. synthesis [588]. System [2, 144, 2110, 247, 1045, 1775, 1972, 23, 7, 273, 2091, 1409, 962, 1813, 33, 107, 1044, 1585, 1092, 668, 1363, 1054, 1035, 2270, 1198, 699, 1151, 138, 783, 1264, 1752, 2020, 1441, 2205, 260, 229, 28, 1655, 1723, 2254, 2215, 1247, 10, 679, 136, 1218, 385, 1037, 2261, 824, 1786, 2273, 1989, 2192, 2115, 2181, 245, 1040, 2199, 2284, 74, 2195, 1158, 1670, 987, 2250, 280, 929, 1823, 1190, 651, 2001, 1667, 1832, 1728, 1141, 230, 1346, 269, 812, 906, 2209, 233, 1613, 1188, 802, 1914, 1152, 458, 106, 491, 1292, 2125, 2022, 1967, 1078, 506, 2077, 1388, 35, 1169, 1115]. system [916, 766, 287, 529, 1378, 2216, 1548, 351, 1837, 1256, 1693, 703, 1205, 1430, 228, 983, 93, 315, 1464, 2062, 2161, 1249, 1853, 2016, 669, 701, 1798, 1842, 1183, 1012, 755, 1961, 2170, 1032, 1308, 1704, 2086, 864, 1583, 1894, 427, 16]. systematic [423]. systemic [2172]. systems [1473, 775, 2051, 575, 1117, 2038, 501, 96, 1799, 1385, 1138, 6, 285, 1662, 2103, 1585, 1739, 849, 781, 2289, 2257, 1046, 1714, 231, 306, 1788, 1913, 1260, 1340, 2229, 2157, 1587, 469, 2281, 1999, 1876, 1031, 2224, 364, 1797, 1841, 635, 2166, 2123, 667, 1021, 1053, 1886, 1005, 354, 1134, 1450, 1963, 404, 1376, 751, 1387, 2278, 1990, 423, 504, 1327, 989, 1449, 1656, 1185, 982, 2241, 27, 369, 1475, 207, 1498, 1027, 1093, 481, 2256, 1804, 1611, 1530, 2098, 2052, 543, 996, 1898, 878, 115, 1901, 1793, 1651, 981, 1762, 430, 676, 1351, 872, 999, 1396, 1104, 1382, 98, 2080, 2263]. systems [990, 1416, 1601, 1904, 2295, 701, 122, 1025, 1744, 2185, 2298, 2187, 412, 463, 899, 275, 1635, 1659, 1938, 901, 820, 2026, 1622, 1861].

T [12, 1720, 1054, 1803, 11, 334, 866, 2029, 974, 584]. T-cells [1054, 584]. T-lymphocytes [1720, 334]. tab [281]. Taguchi [1038]. tailed [2023, 359, 55, 1856]. Tailored [406]. Takagi [27, 1861, 481, 981, 701]. Takens [829, 2250]. taking [2121]. tandem [15]. tanh [880]. tank [1973, 144]. Tanner [825, 2086]. target [327, 83, 603]. targeted [559, 641]. targets [1365]. tau [1440, 1728]. tau-collocation [1728]. tax [1231]. taxis [825, 1976, 1249, 953, 1864]. Taylor [1241, 1289, 771]. Taylor-wavelet [771]. TB [912, 1831]. TD [2201]. TD-BEM [2201]. TDNNS [629]. teaching [1652]. teaching-learning-based [1652]. technique [133, 782, 1878, 1905, 952, 407, 780, 915, 1918, 1145, 628, 269, 906, 1639, 1475, 2201, 1078, 286, 1341, 1448, 1177, 394, 1165, 886, 1410, 2066, 1317, 1032, 533, 2267, 896]. techniques [720, 190, 1958, 1280, 1505, 2130, 2288, 1619, 516, 2077, 2216]. technology [1293, 1186, 1688]. Teissier [1431]. telegraph [1525, 1142]. teleoperation [1841, 872]. Teleoperators [1966]. temperature [594, 196, 611, 707, 2248, 2037, 1722, 1175, 1305, 355, 932, 1985]. temperature-dependent [932]. temperature-rate [196]. temperature-sensitive [355]. tempered [1899, 1189, 377, 1843]. template [1171]. temporal [1051, 818, 2018, 2154, 1798]. temporary [401, 1185]. tension [1874]. tensionless [63]. Tensor [646, 1620, 479, 1325, 406]. tensor-train [1325]. tensorial [2253]. terahertz [595]. Term [1767, 979, 485, 2279, 2176, 1978, 799, 1403, 1193, 1306, 1748, 841, 1350, 2230, 1525, 767, 2015, 719, 113, 760, 2160, 1055, 1266, 1524]. terminal [578, 177, 1308]. terminal-velocity [177]. terms [1313, 323, 130, 831, 2265]. ternary [1303, 2076, 2129, 1854]. Test [1026, 520]. Testing [1743, 164]. tests [1567, 1783]. tetrahedral [1539]. Text [1684]. TF [505]. TGV [2238]. th [1376, 1327]. THB [1932]. THB-splines-powered [1932]. their [1715, 1646, 278, 600, 469, 654, 893, 1375, 1309]. theorem [532, 652]. Theoretical [1626, 2144, 511, 1858, 1422]. theories [877]. theory [85, 1544, 1536, 1594, 1846, 1211, 350, 396, 1275, 723, 1114]. therapy [922, 2117, 402]. Thermal [2076, 665, 321, 413, 1009, 318, 14, 237, 1764, 2129, 358, 741, 1687, 1263, 590, 329, 492, 1726, 1703, 800, 753, 1633]. thermal-free [753]. thermally [355]. thermally-jumped [355]. thermistor [1641]. thermo [889, 1591, 770]. thermo-elastic [889]. thermo-mechanical [1591]. thermo-physical [770]. thermodynamic [201]. thermoelastic [560, 562]. thermoelasticity [617, 2208]. thermohaline [993]. thermophoresis [708, 917, 671]. thermoviscoelastoplastic [1305]. thick [1490, 329, 613]. thickness [612, 44, 227]. thin [465, 1537, 1687, 629, 959, 477, 21]. thin-film [465]. thin-walled [629]. Third [472, 1666, 1729, 2178, 903, 400, 1727, 506]. third-kind [1729, 903]. third-order [2178, 400, 1727]. Thomas [754, 1821]. Three [1973, 544, 2140, 1772, 2271, 1843, 979, 1860, 1163, 484, 1091, 583, 793, 125, 1598, 1739, 1880, 1714, 2215, 10, 246, 82, 757, 2181, 1325, 1049, 982, 1257, 1658, 438, 765, 1142, 1297, 1169, 691, 810, 503]. three-component [1049, 1257]. Three-dimensional [1973, 2140, 583, 2215, 246, 1325, 982, 1658, 810, 503]. three-layer [1091]. three-moment [82]. three-operator [1714]. Three-phase [1772, 438]. three-point [1297]. three-term [979]. three-timescale [2181]. Threshold [1425, 1399, 809, 1077, 1855, 1655, 1059]. thresholding [2092, 2267]. Thresholds [1199, 1349]. throughout [1857]. Thyroiditis [1970]. tides [1879]. tied [423, 1732]. tiger [225]. tight [1291]. tight-binding [1291]. Tikhonov [1985]. tilt [2027]. tilt-rotor [2027]. Time [380, 1550, 2275, 1218, 950, 1741, 1468, 346, 1062, 1725, 1440, 720, 485, 462, 620, 1508, 1097, 768, 1646, 77, 2091, 1117, 583, 2217, 2285, 109, 96, 2105, 1347, 62, 285, 737, 1708, 1662, 1702, 859, 1403, 1880, 727, 694, 868, 1966, 2114, 647, 1436, 453, 1123, 1386, 2141, 2156, 231, 1788, 1559, 780, 1189, 349, 1780, 2020, 1447, 327, 752, 451, 1807, 49, 28, 469, 2281, 1048, 542, 367, 1748, 1797, 1841, 1505, 262, 1808, 2294, 1952, 1992, 1588, 381, 628, 638, 900, 1261, 1286, 134, 548, 483, 818, 1373, 1254, 1053, 1506, 1700, 1577, 2192, 2165, 1017, 279, 162, 1277, 1450, 2018, 1963, 2274]. time [1376, 377, 1392, 1149, 1525, 1674, 1301, 1990, 391, 2246, 504, 599, 2140, 621, 1327, 83, 310, 1185, 352, 2241, 1698, 1610, 2255, 1267, 1136, 2292, 2032, 1429, 1219, 106, 2201, 1619, 259, 1858, 969, 319, 1873, 1546, 1448, 481, 1388, 2013, 670, 719, 1176, 2014, 996, 1106, 1792, 101, 1156, 1747, 878, 1931, 1225, 111, 1901, 1410, 298, 1600, 1435, 703, 981, 1215, 2045, 955, 372, 574, 386, 999, 887, 1255, 1381, 2056, 587, 1055, 750, 2080, 2090, 271, 1601, 1339, 1245, 701, 1183, 1065, 1744, 2298, 2063, 2187, 2290, 1353, 1817, 2213, 980, 1571, 463, 275, 589, 1718, 1661, 2127, 2089, 1950]. time [474, 1000, 1938, 1524, 1885, 820, 1894, 1447]. time-changed [1435]. Time-delay [1218, 1117, 96, 262, 2294, 275, 1894]. time-delayed [2091, 231, 1000]. time-delays [1376]. Time-dependent [380, 1062, 1386, 1559, 1990, 1858, 750, 1353]. time-efficient [1619]. time-fractional [1440, 720, 462, 2105, 1403, 1436, 1123, 780, 542, 1748, 1505, 1588, 548, 1277, 1392, 1525, 1674, 1267, 319, 1448, 2013, 719, 1156, 1410, 1600, 1215, 372, 1381, 587, 2090, 2063, 2213, 1718, 1524, 1885]. time-harmonic [504]. time-independent [2274]. time-integration [2290]. time-order [2165]. time-splitting [1373, 1254]. time-stepping [1931]. time-variable [298]. time-varying [1508, 109, 1347, 737, 2141, 1788, 1780, 1841, 1301, 310, 2241, 106, 996, 101, 111, 981, 999, 271, 2187, 1817, 980, 1000, 1938]. time/fixed [1698]. timer [1456]. times [1834]. timescale [2181, 1879]. Timoshenko [1548, 1114]. tire [1747]. tissue [594, 1591, 1805]. tissues [1427, 1499]. Tocher [2137]. Toeplitz [575, 635, 2166]. tolerant [1228, 2261, 1768, 1292, 675, 1560]. Tool [1759, 662]. top [2140, 1736]. topic [464]. Topological [56, 2054, 430, 396]. topologies [1788]. Topology [764, 2130, 1506, 88, 1835]. TOPSIS [1383]. torque [23, 444, 1464, 1943]. torque-based [1464]. torques [83]. Torrey [1189]. tortuosity [58]. Total [2289, 862]. tourism [1972]. tourism-based [1972]. township [1073]. toxic [1388]. toxicity [2170]. toxin [1906, 1040, 816, 1322, 547, 1423]. toxin-determined [547]. toxin-producing [1040]. tQSSA [2111]. Tracking [2246, 696, 781, 1913, 1272, 2157, 130, 1963, 83, 1793, 1651, 1762, 1351, 1416]. tracking-type [696]. traction [1772]. Trade [1862, 365, 397]. Trade-off [1862]. traders [2034]. traditional [1067]. traffic [2107, 367, 757, 113]. train [1325]. training [1464]. Trajectory [2157, 130, 1272, 975]. trans [1151]. trans-species [1151]. Transfer [427, 536, 116, 513, 582, 660, 1883, 992, 1875, 1263, 1039, 82, 534, 1854, 2241, 235, 1685, 2222]. transform [1695, 321, 359, 29, 2099, 844, 1497, 1424, 1350, 1827, 869, 2139, 940, 1125, 700, 446]. transformation [295, 2081, 2014, 1927]. transformed [1469, 203, 719, 1524]. transformer [1724]. transformers [34]. transforms [2203, 1268, 1060, 1842]. Transient [709, 329]. transistors [424, 425]. transition [1541, 369, 1675]. transitioning [176]. transitions [1928, 1951]. translation [1042]. Transmission [2049, 1221, 572, 1314, 922, 1347, 1838, 1831, 2061, 306, 1153, 2144, 734, 354, 1850, 1980, 1074, 2154, 672, 1751, 857, 939, 1428]. transpiration [2053]. transport [2303, 1137, 1742, 1871, 841, 326, 733, 710, 1909, 177, 93, 747]. transportation [324, 1992, 948, 2209, 1106, 1216]. transverse [619]. travel [367, 1549]. Traveling [1482, 2279, 2263, 720, 196, 1023, 1701, 2286]. trawl [1472]. treat [1066]. treat-age [1066]. treated [1051]. Treating [522]. treatment [594, 2061, 829, 638, 2004, 1514, 2160, 1102, 1420]. treatments [1575, 1664]. tree [1344, 1782, 176, 869]. Trees [782]. Trefftz [1353]. trend [317, 1686]. trends [510, 2057]. Tresca [625]. tri [1504]. tri-trophic [1504]. trial [1931]. triangle [49]. triangles [698, 640]. Triangular [1897, 1458, 1539, 1791, 131]. triangulation [751, 478]. triangulations [1746, 206]. tridiagonal [635, 2166]. tridiagonalization [126]. triggered [2051, 2123, 1053, 1541, 1359, 1596, 369, 1530, 853, 1901, 1793, 1651, 999, 1104, 1601, 1904, 2298, 2187, 1635, 1659, 1938]. trigonometric [786, 1874]. trim [281]. triple [1533, 1102]. triply [1631]. trivariate [191]. trophic [1504]. trough [707]. truncated [1235, 344, 2095, 763, 1236]. truncation [2205]. TT-M [349]. tuberculosis [2031, 1839, 877, 1229]. Tucker [1121]. tumbling [236]. tumor [1710, 594, 1057, 2134, 2240, 2002, 262, 638, 2117, 1102, 402]. tumor-immune [2117, 402]. tumors [1051, 81]. tumour [1054, 2113]. tumour-immune [2113]. tuna [1276]. tuned [675, 1092]. Tunnel [757, 539]. tuples [997]. turbine [297, 687, 1038, 1960]. Turing [2084, 1615, 1798]. turn [24, 821]. turning [1341, 502, 1155]. twenty [2254]. twenty-high [2254]. twin [2126]. twisting [955]. Two [1616, 870, 330, 1285, 1281, 136, 1992, 483, 1134, 1103, 2040, 500, 1954, 976, 1493, 619, 1654, 746, 839, 918, 302, 448, 160, 1862, 2136, 971, 1813, 1983, 6, 114, 859, 868, 2121, 1035, 513, 1132, 422, 343, 892, 1239, 1189, 1252, 1260, 783, 1845, 695, 2020, 2107, 323, 236, 2083, 2024, 108, 650, 841, 1612, 1701, 1875, 957, 1085, 2087, 1574, 2288, 1350, 1154, 354, 1157, 1217, 1677, 1827, 1670, 280, 1823, 1674, 2143, 1833, 690, 1667, 400, 941, 934, 1223, 1726, 438, 1119, 603, 1093, 259, 1858, 563, 767, 1773, 1274, 2015, 1166, 1106, 1222, 1491, 891, 1317, 1985, 1510, 53]. two [887, 1381, 1382, 2287, 1015, 370, 1571, 2127, 2245, 2227, 585]. two-area [1823]. two-delayed [2087]. two-derivative [1574, 2288]. Two-dimensional [870, 1493, 918, 2136, 330, 1035, 343, 892, 1189, 2107, 1875, 1217, 1827, 1674, 1093, 1858, 563, 1381, 1382, 2127, 2245]. two-external [323]. Two-factor [1103]. Two-grid [1285, 483, 1674, 53, 887]. two-layer [619, 1670]. Two-level [330, 1813, 1119, 1106]. two-material [1726]. two-mode [1154]. two-norm [1260]. two-parameter [839, 1166, 1985]. Two-phase [1992, 6, 513, 2015, 2287, 2227]. two-point [160, 695, 934, 1491, 891, 1317]. two-prey [1862]. two-relaxation-time [859]. two-species [1701, 259]. two-stage [971, 2024, 1677, 1773]. two-step [1983, 354]. two-term [1350]. Type [1982, 2110, 772, 1553, 1082, 558, 1045, 570, 123, 968, 998, 1086, 1733, 696, 973, 196, 875, 1044, 1285, 2203, 1470, 1880, 531, 2061, 1198, 888, 76, 1537, 697, 332, 472, 1808, 702, 915, 1918, 381, 1060, 1786, 2087, 1641, 57, 897, 819, 1030, 1408, 1826, 554, 167, 2106, 400, 1267, 1136, 969, 1672, 1250, 771, 1169, 1115, 916, 596, 161, 1630, 1898, 1099, 758, 1901, 1361, 1798, 1961, 1814, 1950, 1939, 961, 2168, 1083, 517]. type-1 [1086, 2168]. Type-2 [1982, 1553, 332, 1898, 1083]. type-III [2061]. type-IV [1798]. types [484, 1715, 590, 690, 2271].

UAT [1933]. UAV [742, 1560]. UAVs [1272]. Uhlenbeck [2151, 1822, 1810, 809].

Ulam [1311]. ultra [2272]. Ultrafast [449]. ultrasound [1925, 81]. Unbalanced [1462]. unbiased [1107]. unbounded [737, 1123, 1301, 500]. Uncertain [2103, 1243, 1820, 2051, 109, 1788, 631, 652, 1286, 1541, 1782, 2158, 1376, 1794, 391, 217, 670, 878, 111, 351, 703, 99, 347, 820]. uncertainties [1260, 693, 1579, 1219, 1120, 420, 1560]. Uncertainty [180, 559, 2280, 2157, 179]. Uncertainty-based [180]. Unconditional [611, 2292]. unconditionally [538, 1052, 1476, 2287, 1543]. unconstrained [977, 1459]. Uncritical [268]. underactuated [327]. underdamped [895]. underground [428]. underwater [353, 1276]. undirected [1835]. unemployment [489, 1343, 2000]. uneven [1466]. uneven/composite [1466]. unidimensional [1098]. unification [1139]. Unified [2006, 47, 1546, 351]. Uniform [2115, 2295, 933, 1939, 1717, 2148, 1646, 1958, 2217, 611, 1465, 1874, 1373, 409, 1703, 563, 1378, 1629, 593, 1155, 1685, 1062, 1150]. Uniformly [1377, 167, 1305, 797]. unilateral [511]. Unique [996]. uniqueness [969, 2251]. unit [1188]. United [1414]. units [1739, 2022]. unity [1300, 906, 1858]. universal [532]. Unknown [31, 1031, 605, 687, 27, 369, 1651, 1904, 2298, 1635]. unmanned [2068]. unoccupied [1997]. unreliable [709, 1551, 1430, 1085]. unreported [856]. Unsteady [779, 643, 968, 917, 358, 2192, 492, 1266, 61, 2264]. unstructured [1594, 841, 1377]. unsymmetric [2138]. update [1272]. updating [1618]. upgrading [407]. upper [1257, 718]. upper-limb [718]. upstream [473]. upwelling [497]. Upwind [1730]. urban [2107, 2158, 180, 420]. urban-porous [2107]. urbanization [256]. US$105.00 [592]. US$74.95 [592]. USA [2242]. Use [184, 71]. used [1185, 20, 738]. users [1773]. Using [29, 1235, 1010, 1886, 1190, 499, 834, 2110, 133, 1625, 1553, 216, 1606, 1163, 961, 643, 1365, 442, 380, 321, 190, 861, 181, 1897, 220, 1930, 2176, 297, 192, 1117, 1238, 2122, 793, 18, 1336, 1623, 588, 597, 1838, 185, 1171, 1815, 1731, 694, 1966, 1241, 1123, 2113, 44, 1287, 1742, 1239, 707, 1586, 58, 1538, 1824, 127, 236, 764, 1426, 1280, 252, 1723, 464, 95, 1007, 2215, 1453, 10, 679, 332, 16, 1218, 281, 414, 146, 2129, 551, 1004, 1612, 1691, 11, 301, 1818, 1167, 248, 948, 467, 957, 900, 1414, 1970, 2130, 1767, 1003, 447, 342]. using [1030, 1191, 1516, 1034, 329, 1665, 2168, 1488, 166, 1833, 1459, 59, 1608, 1933, 2241, 1735, 1024, 1475, 322, 923, 940, 2155, 2007, 2190, 458, 2201, 477, 1619, 289, 951, 1614, 21, 105, 2282, 801, 100, 803, 1125, 1986, 543, 2044, 2216, 858, 8, 1075, 539, 1901, 34, 1837, 1205, 1964, 1111, 2154, 180, 218, 2188, 24, 1114, 756, 533, 2222, 848, 1866, 2245, 1524, 261]. utility [1008]. utilization [2278]. utilizing [269].

V [1825]. vacation

[709, 1334, 1551, 1085, 929, 2022, 2066, 1430]. vacationing [1292]. vacations [136, 2080]. vaccination [1584, 1941, 2061, 1058, 937, 928, 1510, 1244, 2242]. vaccinations [2064]. vaccine [476]. validated [417]. Validation [1067, 2270, 2030, 1453, 630]. Valuation [1265, 2207, 1457, 1806, 1513, 2193]. value [1860, 1771, 627, 160, 2272, 84, 777, 72, 650, 371, 467, 561, 735, 1867, 2085, 305, 1083, 1982, 813, 400, 1407, 767, 1297, 1432, 1491, 891, 1317, 1014]. value-at-risk [1771, 1014]. valued [1121, 109, 2033, 966, 1962, 162, 1219, 969, 1873, 101, 2221, 1212, 1000, 1866]. values [1945]. Valuing [1456]. valve [223]. vanadium [2109, 417]. vanishing [1135, 1729]. vapor [465]. vaporization [1805]. VAR [182]. variability [1914]. Variable [237, 1440, 908, 23, 931, 1966, 44, 1559, 1264, 1752, 2248, 2081, 697, 1997, 1038, 1649, 1228, 1218, 1355, 381, 732, 1154, 1460, 1577, 2011, 1149, 492, 816, 692, 2292, 535, 1873, 227, 770, 2013, 2015, 1052, 283, 298, 372, 587, 1055, 2193, 2287, 817, 1278, 1718, 798, 843]. variable-density [2248]. variable-order [1440, 1355, 1460, 1577, 1873, 587, 1055, 1278, 798]. variable-speed [732]. variable-step [1154, 1718]. variable-time-step [2292]. variables [558, 123, 1252, 997, 924, 27, 1024, 516]. variance [1107, 39, 240, 700, 2124, 1578]. variant [29]. variants [1087, 1510, 2064]. variate [344]. variates [1626, 340]. variation [960]. variational [216, 1333, 1567, 77, 1991, 702, 1024, 604, 1432, 1081, 1101, 2171]. variations [58]. various [1239, 2247, 1918, 894, 86]. varistructure [124]. varying [1508, 109, 1347, 737, 2141, 1788, 1780, 1841, 1286, 1301, 310, 2241, 1563, 106, 1530, 996, 101, 111, 981, 999, 271, 2187, 1817, 2231, 980, 1000, 1938, 494]. varying-based [2231]. Vasicek [1246]. vector [1725, 1567, 2033, 1036, 1956, 322, 1658, 458, 69, 794, 368, 901]. vector-valued [2033]. vegetation [1929, 1517, 2165, 1348]. vehicle [7, 70, 679, 20, 5, 353, 1276]. vehicles [2068, 1383, 327, 130, 449, 1488]. velocities [130]. velocity [696, 129, 2125, 177]. VEM [1483, 2065]. Verhulst [1574]. verification [1654]. version [1729, 1323, 867, 477, 1549, 1697]. versus [515]. Vertical [2122, 1731, 237, 1126, 1677]. Vertically [2175]. Very [1144, 558, 132, 1534]. Veselov [651]. vesicles [1290]. vessel [611, 1633]. via [388, 761, 302, 470, 1401, 2136, 501, 1439, 967, 2156, 1669, 168, 1593, 2281, 697, 1910, 1962, 299, 1213, 1874, 1289, 2123, 1424, 1350, 1053, 162, 1595, 2168, 2198, 1656, 1726, 1698, 106, 604, 473, 670, 1222, 853, 1898, 878, 891, 811, 1282, 1693, 2179, 93, 1382, 1601, 1310, 122, 1032, 1659, 1000, 798, 901]. vibrating [2096]. vibration [2254, 496, 126, 20, 75]. Vibrational [1548, 1304]. Vibrations [889, 832, 224]. VIEs [1729]. Vieta [2278, 1011]. view [1455]. viewpoint [1365]. Villari [429]. VIM [2171]. Viral [1030, 748, 1946, 866, 2029, 1519, 1850, 2064]. virotherapy [868, 773]. Virtual [133, 185, 1648, 219, 1346, 1765, 2292, 658]. virulent [393]. Virus [603, 653, 922, 730, 1221, 1422, 1371, 1374, 1738, 2204, 585]. virus-to-cell [922]. viruses [704, 2004]. Visceral [538]. viscoelastic [931, 611, 367, 496, 1366, 1460, 2143, 2140, 770, 738, 2161]. viscoplastic [1845]. viscosity [772, 694, 237, 487, 932]. viscosity-type [772]. viscous [1454, 796, 496, 47, 1687, 1263, 1672, 1954, 815]. Visualization [147]. vivax [1511]. Vlasov [1585]. VLSI [983]. vol [240]. vol-of-vol [240]. volatilities [1223]. Volatility [242, 1868, 1265, 2067, 1783, 605, 1893, 2274, 2207, 1741, 1984]. Voltage [1654, 2231, 448, 447, 423, 446]. Voltage-balancing [1654]. Volterra [870, 2020, 2042, 505, 904, 1874, 711, 2273, 903, 1424, 1872, 280, 1728, 1331, 1573, 2112, 1829, 1061, 1222, 1253, 1399, 814, 1944]. volume [1402, 619, 2148, 952, 841, 1223, 1879, 131, 884, 1127, 859]. volumes [147, 2040, 1064]. Vortex [1812, 1981]. vorticity [1812, 142]. Vries [2290, 1033, 247, 1167, 1801, 1215, 581, 1950]. vs [2191]. VSC [306]. VSC-HVDC [306].

w [346]. w-shape [346]. walk [1369]. walk-based [1369]. walking [2096]. wall [611, 1337]. walled [629]. walls [355]. Walsh [705, 392]. Walter [948]. Walter-B [948]. Walters [905]. Wang [1197]. waning [1563]. waning-boosting [1563]. washoff [180]. waste [468]. water [216, 14, 1404, 6, 542, 2037, 879, 2165, 102, 184, 1879, 366, 2054, 2244, 1215, 251, 353, 1685, 2290]. water-plants [2165]. waters [1455]. Wave [617, 216, 720, 348, 1616, 1318, 462, 1482, 956, 842, 1147, 566, 1403, 1035, 880, 1239, 328, 1696, 542, 824, 239, 1973, 2115, 142, 1495, 1670, 1325, 2190, 1804, 2015, 1056, 1156, 905, 1, 389, 1215, 574, 500, 2150, 669, 353, 2290, 1321, 2127, 1950]. wave-guide [142]. waveform [437, 1554, 1791]. waveguides [1089]. Wavelet [2092, 1079, 1473, 1454, 194, 842, 677, 780, 127, 1507, 935, 1591, 819, 1525, 120, 771, 161, 283, 2044, 1165, 2290, 2264, 1181, 771]. wavelet-homotopy [677, 2264]. wavelets [870, 1283, 1623, 320, 47, 1289, 1665, 1078, 593]. waves [2279, 196, 1259, 1023, 1701, 879, 1240, 556, 2286, 408, 1128, 857, 2263]. wavy [1126]. Weak [784, 1902, 558, 1484, 625, 2272, 1261, 1968, 1629, 2056, 795, 933]. weakly [2036, 1728, 765, 1225, 1128, 887, 1361, 648, 1150]. weather [2232]. web [1736]. WECS [1462]. wedge [2129, 2053]. Weibull [666, 688]. weight [527]. weighted [1930, 705, 1824, 1152, 815, 2025]. weighting [1905]. weights [1160, 1461]. Well [2161, 1660, 2128, 2037, 1517, 1879, 1953]. well-balanced [1660, 1517, 1879]. Well-posedness [2161, 1953]. Wendroff [575]. WENO [1082, 190, 1811, 1717]. WENO-based [1717]. wetlands [2122]. whale [1118, 732, 947, 460, 1212]. wheel [20]. wheeled [716, 474]. wheels [83]. White [1840, 363, 2255, 402]. Whitham [2286]. whole [2201]. wide [1016, 424, 2092]. widow [935]. Wiener [1236]. Wigner [145]. wildfire [2005]. wildfire-affected [2005]. wildfires [175]. wildlife [1722]. WILL [275]. Williamson [2129, 741]. wind [664, 1759, 33, 687, 1038, 1960, 2078, 2060]. winding [1769, 24, 821]. winning [555]. wire [595]. wire-grid [595]. wireless [1518, 655, 427]. within [922, 668, 2076, 147, 2083, 1677, 2032]. within-host [922, 2083]. without [702, 530]. Wolbachia [967]. Wolf [1160, 1813]. workflow [174, 185]. working [136, 1834, 2080]. workplace [1838]. Workspace [174, 185]. world [991, 867, 564]. Wright [1060]. Wronskian [566, 876]. WTI [1457]. WTS [1882].

X [1335]. X-MODE [1335]. XFEM [1545]. xi [592]. XOR [1217, 1583]. XV [189]. XVA [1207].

yacht [2007]. Yager [991]. yellow [1221].

Z [2110, 554]. Z-type [2110, 554]. zag [957]. Zakharov [720, 279, 987]. zebra [1818]. Zero [2257, 2099, 1963, 1176]. Zero-Hopf [2257]. zero-inflated [1176]. zerosum [1963]. zeroing [1780, 2128, 1228, 2039, 271, 474]. zeros [210, 1235]. Zettle [2179]. Zhangians [1780]. Zhejiang [920]. zig [957]. zig-zag [957]. Zinc [487]. ZNN [2246, 843]. zooplankton [1040, 816, 1322].

References

Triki:2009:EAS

[1] Houria Triki and Thiab R. Taha. Exact analytic solitary wave solutions for the RKL model. Mathematics and Computers in Simulation, 80(4): 849–854, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                              //www.sciencedirect.com/science/
                                              article/pii/S0378475409002833. See

comment [348] and reply [389].

Becherif:2020:SIO

[2] Mohamed Becherif, Haitham S. Ramadan, Haoping Wang, Ala’a H. Al-Muhtaseb, and Zhifu Mi. Special issue on optimizations for renewable energy system applications. Mathematics and Computers in Simulation, 167:1–3, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541930206X.

Fan:2020:ICC

[3] Jing-Li Fan, Jia-Wei Hu, Xian Zhang, Ling-Si Kong, Fengyu Li, and Zhifu Mi. Impacts of climate change on hydropower generation in China. Mathematics and Computers in Simulation, 167:4–18, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300065.

Ourahou:2020:RSG

[4] M. Ourahou, W. Ayrir, B. El Hassouni, and A. Haddi. Review on smart grid control and reliability in presence of renewable energies: Challenges and prospects. Mathematics and Computers in Simulation, 167:19–31, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303045.

Saadi:2020:RCL

[5] R. Saadi, M. Y. Hammoudi, O. Kraa, M. Y. Ayad, and M. Bahri. A robust control of a 4-leg floating interleaved boost converter for fuel cell electric vehicle application. Mathematics and Computers in Simulation, 167:32–47, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541930268X.

Chen:2020:DSS

[6] Jian Chen, Lianghui Huang, Chizhou Yan, and Zhiyang Liu. A dynamic scalable segmented model of PEM fuel cell systems with twophase water flow. Mathematics and Computers in Simulation, 167:48– 64, January 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301186.

Badji:2020:AEE

[7] Abderrezak Badji, Djaffar Ould Abdeslam, Mohamed Becherif, Fouad Eltoumi, and Nacereddine Benamrouche. Analyze and evaluate of energy management system for fuel cell electric vehicle based on frequency splitting. Mathematics and Computers in Simulation, 167:65– 77, January 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300722.

Soliman:2020:MEU

[8] Aly M. A. Soliman, Hamdy Hassan, Mahmoud Ahmed, and Shinichi Ookawara. A 3d model of the effect of using heat spreader on the performance of photovoltaic panel (PV). Mathematics and Computers in Simulation, 167:78–91, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830123X.

Khaled:2020:IAS

[9] Mahmoud Khaled, Rabih Murr, Hicham El Hage, Mohamad Ramadan, Haitham Ramadan, and Mohamad Becherif. An iterative algorithm for simulating heat recovery from exhaust gas — application on generators. Mathematics and Computers in Simulation, 167:92–103, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300971.

Ismail:2020:SBC

[10] Mohamed M. Ismail and Ahmed F. Bendary. Smart battery controller using ANFIS for three phase grid connected PV array system. Mathematics and Computers in Simulation, 167:104–118, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300946.

Khabou:2020:MIB

[11] H. Khabou, M. Souissi, and A. Aitouche. MPPT implementation on boost converter by using T–S fuzzy method. Mathematics and Computers in Simulation, 167:119–134, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475418301228.

Boumaaraf:2020:PEL

[12] Billel Boumaaraf, Houria Boumaaraf, Mohamed El-Amine Slimani, Selma Tchoketch-Kebir, Mohamed Salah Aitcheikh, and Khaled Touafek. Performance evaluation of a locally modified PV module to a PV/T solar collector under climatic conditions of semi-arid region. Mathematics and Computers in Simulation, 167:135–154, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302678.

Maamir:2020:NNP

[13] M. Maamir, O. Charrouf, A. Betka, M. Sellali, and M. Becherif. Neural network power management for hybrid electric elevator application. Mathematics and Computers in Simulation, 167:155–175, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302629.

Boumaaraf:2020:CET

[14] Billel Boumaaraf, Khaled Touafek, Mohamed Salah Ait-cheikh, and Mohamed El Amine Slimani. Comparison of electrical and thermal performance evaluation of a classical PV generator and a water glazed hybrid photovoltaic–thermal collector. Mathematics and Computers in Simulation, 167:176–193, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302131.

Boudaoud:2020:SOT

[15] Chahrazad Boudaoud, Abdelkader Hamdoune, and Zehor Allam. Simulation and optimization of a tandem solar cell based on InGaN. Mathematics and Computers in Simulation, 167:194–201, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418302179.

Jendoubi:2020:SMC

[16] Abdessattar Jendoubi, Faouzi Tlili, and Faouzi Bacha. Sliding mode control for a grid connected PVsystem using interpolation polynomial MPPT approach. Mathematics and Computers in Simulation, 167: 202–218, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302617.

Ghamri:2020:CSA

[17] A. Ghamri, R. Boumaaraf, M. T. Benchouia, H. Mesloub, A. Goléa, and N. Goléa. Comparative study of ANN DTC and conventional DTC controlled PMSM motor. Mathematics and Computers in Simulation, 167: 219–230, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302605.

Bouhadjra:2020:PAS

[18] Dyhia Bouhadjra, Aissa Kheldoun, and Ali Zemouche. Performance analysis of stand-alone six-phase induction generator using heuristic algorithms. Mathematics and Computers in Simulation, 167:231–249, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302095.

Guellout:2020:EBR

[19] O. Guellout, A. Rezig, S. Touati, and A. Djerdir. Elimination of broken rotor bars false indications in induction machines. Mathematics and Computers in Simulation, 167: 250–266, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302083.

Rezig:2020:IOC

[20] Ali Rezig, Wafa Boudendouna, Abdesslem Djerdir, and Abdoul N’Diaye. Investigation of optimal control for vibration and noise reduction in-wheel switched reluctance motor used in electric vehicle. Mathematics and Computers in Simulation, 167:267– 280, January 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419301922.

Reddaf:2020:MEB

[21] Abdelmalek Reddaf, Fatima Djerfaf, Karim Ferroudji, Mounir Boudjerda, Khaled Hamdi-Chérif, and Islem Bouchachi. Modeling of electromagnetic behavior of composite thin layers using genetic algorithm. Mathematics and Computers in Simulation, 167:281–295, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418303306.

Mesloub:2020:CSC

[22] H. Mesloub, R. Boumaaraf, M. T. Benchouia, A. Goléa, N. Goléa, and K. Srairi. Comparative study of conventional DTC and DTC SVM based control of PMSM motor — simulation and experimental results. Mathematics and Computers in Simulation, 167:296–307, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301484.

Ayrir:2020:DTC

[23] W. Ayrir, M. Ourahou, B. El Hassouni, and A. Haddi. Direct torque control improvement of a variable speed DFIG based on a fuzzy inference system. Mathematics and Computers in Simulation, 167:308– 324, January 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418301265.

Yassa:2020:MDS

[24] N. Yassa and M. Rachek. Modeling and detecting the stator winding inter turn fault of permanent magnet synchronous motors using stator current signature analysis. Mathematics and Computers in Simulation, 167:325–339, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418300983.

Mesloub:2020:DID

[25] H. Mesloub, M. T. Benchouia, R. Boumaaraf, A. Goléa, N. Goléa, and M. Becherif. Design and implementation of DTC based on AFLC and PSO of a PMSM. Mathematics and Computers in Simulation, 167: 340–355, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541830096X.

Bethoux:2020:ACP

[26] O. Béthoux, E. Labouré, E. Berthelot, A. Kolli, and A. De Bernardinis. An advanced control for a PM synchronous motor drive in power degraded mode. Mathematics and Computers in Simulation, 167:356– 371, January 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300497.

Ouhib:2020:SUI

[27] Lamia Ouhib. State and unknown inputs estimation for Takagi–Sugeno systems with immeasurable premise variables: Proportional Multiple Integral observer design. Mathematics and Computers in Simulation, 167: 372–380, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418301253.

Hamel:2020:LBR

[28] M. Hamel and H. Mohellebi. A LabVIEW-based real-time acquisition system for crack detection in conductive materials. Mathematics and Computers in Simulation, 167:381– 388, January 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418300442.

Boudraa:2020:USH

[29] Omar Boudraa, Walid Khaled Hidouci, and Dominique Michelucci. Using skeleton and Hough transform variant to correct skew in historical documents. Mathematics and Computers in Simulation, 167:389– 403, January 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301739.

Sadou:2020:PRN

[30] Besma Sadou, Atidel Lahoulou, and Toufik Bouden. PFF-RVM: a new no reference image quality measure. Mathematics and Computers in Simulation, 167:404–414, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475418303008.

Amrane:2020:UIO

[31] Ahmed Amrane, Abdelkader Larabi, and Abdel Aitouche. Unknown input observer design for fault sensor estimation applied to induction machine. Mathematics and Computers in Simulation, 167:415– 428, January 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475418302507.

Mziou:2020:ECP

[32] Nassima M’ziou. Electromagnetic compatibility problems of indirect lightning stroke on Overhead Power Lines. Mathematics and Computers in Simulation, 167:429–442, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475418300934.

Charrouf:2020:ANN

[33] O. Charrouf, A. Betka, S. Abdeddaim, and A. Ghamri. Artificial Neural Network power manager for hybrid PV-wind desalination system. Mathematics and Computers in Simulation, 167:443–460, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302599.

Tamizi:2020:EIR

[34] K. Tamizi, O. Béthoux, and E. Labouré. An easy to implement and robust design control method dedicated to multi-cell converters using inter cell transformers. Mathematics and Computers in Simulation, 167: 461–477, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419300527.

Sellali:2020:PMI

[35] M. Sellali, A. Betka, and A. Djerdir. Power management improvement of hybrid energy storage system based on H_{∞} control. Mathematics and Computers in Simulation, 167:478– 494, January 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541930151X.

Anonymous:2020:EBa

[36] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 167:495, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302770.

Anonymous:2020:EBb

[37] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 167:496, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302782.

Anonymous:2020:EBc

[38] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 167:ifc, January 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302708.

Kim:2020:PGV

[39] See-Woo Kim and Jeong-Hoon Kim. Pricing generalized variance swaps under the Heston model with stochastic interest rates. Mathematics and Computers in Simulation, 168: 1–27, February 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302289.

Anonymous:2020:PF

[40] Anonymous. Pages 1–216 (February 2020). Mathematics and Computers in Simulation, 168:1–216, February 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Zanella:2020:SPS

[41] Mattia Zanella. Structure preserving stochastic Galerkin methods for Fokker–Planck equations with background interactions. Mathematics and Computers in Simulation, 168: 28–47, February 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302277.

Elaziz:2020:OBM

[42] Mohamed Abd Elaziz, Ahmed A. Ewees, Rehab Ali Ibrahim, and Songfeng Lu. Opposition-based moth-flame optimization improved by differential evolution for feature selection. Mathematics and Computers in Simulation, 168:48– 75, February 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302150.

Cho:2020:OMP

[43] Durkbin Cho. Optimal multilevel preconditioners for isogeometric collocation methods. Mathematics and Computers in Simulation, 168:76– 89, February 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302320.

Doh:2020:HHR

[44] Deog-Hee Doh, M. Muthtamilselvan, B. Swathene, and E. Ramya. Homogeneous and heterogeneous reactions in a nanofluid flow due to a rotating disk of variable thickness using HAM. Mathematics and Computers in Simulation, 168:90–110, February 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302344.

Praba:2020:AGC

[45] B. Praba and R. Saranya. Application of the graph cellular automaton in generating languages. Mathematics and Computers in Simulation, 168:111–121, February 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302253.

Zou:2020:CFS

[46] Guang an Zou, Bo Wang, and Tony W. H. Sheu. On a conservative Fourier spectral Galerkin method for cubic nonlinear Schrödinger equation with fractional Laplacian. Mathematics and Computers in Simulation, 168: 122–134, February 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302356.

Karkera:2020:AGU

[47] Harinakshi Karkera, Nagaraj N. Katagi, and Ramesh B. Kudenatti. Analysis of general unified MHD boundary-layer flow of a viscous fluid — a novel numerical approach through wavelets. Mathematics and Computers in Simulation, 168:135– 154, February 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419302332.

Rezaiee-Pajand:2020:FCE

[48] M. Rezaiee-Pajand, A. Aftabi S., and M. S. Kazemiyan. A family of cylindrical elements. Mathematics and Computers in Simulation, 168: 155–172, February 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302368.

Hajihassanpour:2020:IDT

[49] M. Hajihassanpour and K. Hejranfar. An implicit dual-time stepping highorder nodal discontinuous Galerkin method for solving incompressible flows on triangle elements. Mathematics and Computers in Simulation, 168: 173–214, February 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302514.

Anonymous:2020:NIa

[50] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 168:215, February 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303003.

Anonymous:2020:ICEa

[51] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 168:216, February 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419303015.

Anonymous:2020:EBd

[52] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 168:ifc, February 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302952.

Wang:2020:TGM

[53] Yang Wang, Yanping Chen, and Yunqing Huang. A two-grid method for semi-linear elliptic interface problems by partially penalized immersed finite element methods. Mathematics and Computers in Simulation, 169: 1–15, March 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303180.

Anonymous:2020:PM

[54] Anonymous. Pages 1–168 (March 2020). Mathematics and Computers in Simulation, 169:1–168, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Cheng:2020:CEA

[55] Tsung-Lin Cheng and Jheng-Ting Wang. A computationally efficient approach on detecting star-shaped change boundaries in random fields with heavy-tailed distributions. Mathematics and Computers in Simulation, 169:16–25, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303118.

Ghezaiel:2020:TSA

[56] Emna Ghezaiel and Maatoug Hassine. Topological sensitivity analysis based method for solving a geometry reconstruction problem. Mathematics and Computers in Simulation, 169:26–50, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302915.

Li:2020:DGF

[57] Changpin Li and Zhen Wang. The discontinuous Galerkin finite element method for Caputo-type nonlinear conservation law. Mathematics and Computers in Simulation, 169:51– 73, March 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302903.

Espinoza-Andaluz:2020:HTV

[58] Mayken Espinoza-Andaluz, Víctor Velasco-Galarza, and Alex Romero-Vera. On hydraulic tortuosity variations due to morphological considerations in 2D porous media by using the Lattice Boltzmann method. Mathematics and Computers in Simulation, 169:74–87, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541930285X.

Mu:2020:MSI

[59] Yue Mu, Lianqiang Hang, Guoqun Zhao, Xiaona Wang, Youlei Zhou, and Zhihao Cheng. Modeling and simulation for the investigation of polymer film casting process using finite element method. Mathematics and Computers in Simulation, 169:88–102, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302666.

Torki:2020:DOP

[60] Mohammad Mehdi Torki, Shadi Hassanajili, and Marc Mehrzad Jalisi. Design optimizations of PLA stent structure by FEM and investigating its function in a simulated plaque artery. Mathematics and Computers in Simulation, 169:103–116, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302654.

Zaman:2020:CBS

[61] Akbar Zaman, Nasir Ali, and Ambreen Afsar Khan. Computational biomedical simulations of hybrid nanoparticles on unsteady blood hemodynamics in a stenotic artery. Mathematics and Computers in Simulation, 169:117–132, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302642.

Chen:2020:DRA

[62] Jianxin Chen, Tonghua Zhang, and Yongwu Zhou. Dynamics of a risk-averse newsvendor model with continuous-time delay in supply chain financing. Mathematics and Computers in Simulation, 169:133– 148, March 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419302630.

Hsu:2020:FEA

[63] Yang Shang Hsu. Finite element approach of the buried pipeline on tensionless foundation under random ground excitation. Mathematics and Computers in Simulation, 169:149– 165, March 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419302587.

Anonymous:2020:NIb

[64] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 169:166, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303271.

Anonymous:2020:ICEb

[65] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 169:167, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419303283.

Anonymous:2020:EBe

[66] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 169:ifc, March 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303222.

Wei:2020:SES

[67] Fengying Wei and Rui Xue. Stability and extinction of SEIR epidemic models with generalized nonlinear incidence. Mathematics and Computers in Simulation, 170:1– 15, April 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302538.

Anonymous:2020:PAa

[68] Anonymous. Pages 1–414 (April 2020). Mathematics and Computers in Simulation, 170:1–414, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ray:2020:IID

[69] Santanu Ray and Fahad Al Basir. Impact of incubation delay in plant– vector interaction. Mathematics and Computers in Simulation, 170: 16–31, April 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541930254X.

DeSantos-Berbel:2020:EVS

[70] César De Santos-Berbel and Maria Castro. Effect of vehicle swiveling headlamps and highway geometric design on nighttime sight distance. Mathematics and Computers in Simulation, 170:32–50, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302563.

Carvalho:2020:UDB

[71] Pablo G. S. Carvalho, Philippe R. B. Devloo, and Sônia M. Gomes. On the use of divergence balanced H(div)– L^{2} pair of approximation spaces for divergence-free and robust simulations of Stokes, coupled Stokes–Darcy and Brinkman problems. Mathematics and Computers in Simulation, 170:51–78, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302551.

Garcia:2020:DEC

[72] J. García, J. A. Padilla, J. Ruiz, and J. C. Trillo. N-dimensional error control multiresolution algorithms for the point value discretization. Mathematics and Computers in Simulation, 170:79–97, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302198.

Ong:2020:GMC

[73] Seng-Huat Ong, Wen-Jau Lee, and Yeh-Ching Low. A general method of computing mixed Poisson probabilities by Monte Carlo sampling. Mathematics and Computers in Simulation, 170:98–106, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302575.

Llibre:2020:GQD

[74] Jaume Llibre and Clàudia Valls. Global qualitative dynamics of the Brusselator system. Mathematics and Computers in Simulation, 170:107– 114, April 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302526.

Sarker:2020:GML

[75] Pratik Sarker and Uttam K. Chakravarty. A generalization of the method of lines for the numerical solution of coupled, forced vibration of beams. Mathematics and Computers in Simulation, 170:115–142, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303143.

Guo:2020:ASB

[76] Wei Guo, Yufeng Nie, and Weiwei Zhang. Acceleration strategies based on bubble-type adaptive mesh refinement method. Mathematics and Computers in Simulation, 170:143–163, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303179.

Bartelt:2020:HOF

[77] M. Bartelt, O. Klöckner, J. Dietzsch, and M. Groß. Higher order finite elements in space and time for anisotropic simulations with variational integrators. Application of an efficient GPU implementation. Mathematics and Computers in Simulation, 170:164–204, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302861.

Nikishova:2020:SAB

[78] Anna Nikishova, Giovanni E. Comi, and Alfons G. Hoekstra. Sensitivity analysis based dimension reduction of multiscale models. Mathematics and Computers in Simulation, 170:205–220, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303167.

Khan:2020:DAC

[79] Mair Khan, T. Salahuddin, M. Y. Malik, Anum Tanveer, Arif Hussain, and Ali S. Alqahtani. 3-D axisymmetric Carreau nanofluid flow near the Homann stagnation region along with chemical reaction: Application Fourier’s and Fick’s laws. Mathematics and Computers in Simulation, 170:221–235, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303325.

Lamboni:2020:DBG

[80] Matieyendou Lamboni. Derivativebased generalized sensitivity indices and Sobol’ indices. Mathematics and Computers in Simulation, 170: 236–256, April 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303209.

Karunanayake:2020:PMS

[81] N. Karunanayake, P. Aimmanee, W. Lohitvisate, and S. S. Makhanov. Particle method for segmentation of breast tumors in ultrasound images. Mathematics and Computers in Simulation, 170:257–284, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847541930312X.

Li:2020:NTM

[82] Ruo Li, Weiming Li, and Lingchao Zheng. A nonlinear three-moment model for radiative transfer in spherical symmetry. Mathematics and Computers in Simulation, 170:285– 299, April 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303374.

My:2020:MCR

[83] Chu Anh My, Stanislav S. Makhanov, Nguyen A. Van, and Vu M. Duc. Modeling and computation of real-time applied torques and non-holonomic constraint forces/moment, and optimal design of wheels for an autonomous security robot tracking a moving target. Mathematics and Computers in Simulation, 170:300– 315, April 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303350.

Chen:2020:KPB

[84] Hao Chen and Qiuyue Huang. Kronecker product based preconditioners for boundary value method discretizations of space fractional diffusion equations. Mathematics and Computers in Simulation, 170:316– 331, April 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303404.

Abdolrazzagh-Nezhad:2020:ECA

[85] Majid Abdolrazzagh-Nezhad, Homa Radgohar, and Seyede Najme Salimian. Enhanced cultural algorithm to solve multi-objective attribute reduction based on rough set theory. Mathematics and Computers in Simulation, 170:332–350, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303386.

Magdalena:2020:ANS

[86] I. Magdalena and H. Q. Rif’atin. Analytical and numerical studies for harbor oscillation in a semi-closed basin of various geometric shapes with porous media. Mathematics and Computers in Simulation, 170: 351–365, April 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303337.

Peng:2020:DAP

[87] Yahong Peng and Guoying Zhang. Dynamics analysis of a predator–prey model with herd behavior and nonlocal prey competition. Mathematics and Computers in Simulation, 170:366–378, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303453.

Safaei:2020:RTR

[88] F. Safaei, H. Yeganloo, and R. Akbar. Robustness on topology reconfiguration of complex networks: an entropic approach. Mathematics and Computers in Simulation, 170:379–409, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303477.

Anonymous:2020:NIc

[89] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 170:410, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303635.

Anonymous:2020:ICEc

[90] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 170:411–414, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475419303647.

Anonymous:2020:EBf

[91] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 170:ifc, April 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303581.

Cao:2020:P

[92] Jinde Cao, Ramachandran Raja, and Xiaodi Li. Preface. Mathematics and Computers in Simulation, 171:1–2, May 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419303349.

Wang:2020:ISS

[93] Yaqi Wang, Jianquan Lu, Jinde Cao, Wei Huang, Jianhua Guo, and Yun Wei. Input-to-state stability of the road transport system via cyber– physical optimal control. Mathematics and Computers in Simulation, 171:3–12, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301545.

Gong:2020:SBC

[94] Jian Gong, Jinde Cao, Yuan Zhao, Yun Wei, Jianhua Guo, and Wei Huang. Sampling-based cooperative adaptive cruise control subject to communication delays and actuator lags. Mathematics and Computers in Simulation, 171:13–25, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303155.

Humphries:2020:IRE

[95] Usa Humphries, Pramet Kaewmesri, Pariwate Varnakovida, and Prungchan Wongwises. Improvement in rainfall estimates using bias correction with the IAP-DCP Global Model. Mathematics and Computers in Simulation, 171:26–35, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302939.

Cao:2020:SSA

[96] Yang Cao, R. Sriraman, and R. Samidurai. Stability and stabilization analysis of nonlinear time-delay systems with randomly occurring controller gain fluctuation. Mathematics and Computers in Simulation, 171: 36–51, May 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419300825.

Georgiou:2020:MPB

[97] F. Georgiou and N. Thamwattana. Modelling phagocytosis based on cell–cell adhesion and prey– predator relationship. Mathematics and Computers in Simulation, 171: 52–64, May 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302873.

Wu:2020:COI

[98] Shuchen Wu, Xiaohui Sun, Xiaodi Li, and Haipeng Wang. On controllability and observability of impulsive control systems with delayed impulses. Mathematics and Computers in Simulation, 171:65–78, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301090.

Wang:2020:PPD

[99] Limin Wang and Qiankun Song. Pricing policies for dual-channel supply chain with green investment and sales effort under uncertain demand. Mathematics and Computers in Simulation, 171:79–93, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302393.

Saelao:2020:SKG

[100] Jeerawan Saelao and Natsuda Yokchoo. The solution of Klein–Gordon equation by using modified Adomian decomposition method. Mathematics and Computers in Simulation, 171: 94–102, May 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419303131.

Sriraman:2020:GAS

[101] R. Sriraman, Yang Cao, and R. Samidurai. Global asymptotic stability of stochastic complex-valued neural networks with probabilistic timevarying delays. Mathematics and Computers in Simulation, 171:103– 118, May 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301120.

Long:2020:IAS

[102] Bui Ta Long. Inverse algorithm for Streeter–Phelps equation in water pollution control problem. Mathematics and Computers in Simulation, 171:119–126, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303519.

Huang:2020:ASH

[103] Chuangxia Huang, Renli Su, Jinde Cao, and Songlin Xiao. Asymptotically stable high-order neutral cellular neural networks with proportional delays and D operators. Mathematics and Computers in Simulation, 171:127–135, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301983.

Jang:2020:OCP

[104] Junyoung Jang, Hee-Dae Kwon, and Jeehyun Lee. Optimal control problem of an SIR reaction–diffusion model with inequality constraints. Mathematics and Computers in Simulation, 171:136–151, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302307.

Renuka:2020:EAN

[105] A. Renuka, M. Muthtamilselvan, Deog-Hee Doh, and Gyeong-Rae Cho. Entropy analysis and nanofluid past a double stretchable spinning disk using Homotopy Analysis Method. Mathematics and Computers in Simulation, 171:152–169, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419301570.

Puangmalai:2020:FTS

[106] Jirapong Puangmalai, Jakkrapong Tongkum, and Thaned Rojsiraphisal. Finite-time stability criteria of linear system with non-differentiable time-varying delay via new integral inequality. Mathematics and Computers in Simulation, 171:170– 186, May 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419302113.

Chen:2020:FDM

[107] Maoli Chen and Xiao Wang. Flocking dynamics for multi-agent system with measurement delay. Mathematics and Computers in Simulation, 171:187–200, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302836.

Huang:2020:SAN

[108] Chuangxia Huang, Xiaoguang Yang, and Jinde Cao. Stability analysis of Nicholson’s blowflies equation with two different delays. Mathematics and Computers in Simulation, 171:201–206, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302927.

Cao:2020:RSU

[109] Yang Cao, R. Sriraman, N. Shyamsundarraj, and R. Samidurai. Robust stability of uncertain stochastic complex-valued neural networks with additive time-varying delays. Mathematics and Computers in Simulation, 171:207–220, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301752.

Liu:2020:DSG

[110] Ping Liu, Zao Fu, Jinde Cao, Yun Wei, Jianhu Guo, and Wei Huang. A decentralized strategy for generalized Nash equilibrium with linear coupling constraints. Mathematics and Computers in Simulation, 171: 221–232, May 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419302022.

Suntonsinsoungvon:2020:ESD

[111] E. Suntonsinsoungvon and S. Udpin. Exponential stability of discrete-time uncertain neural networks with multiple time-varying leakage delays. Mathematics and Computers in Simulation, 171:233–245, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303106.

Tao:2020:NIC

[112] Liang Tao and Xiao-Long Deng. A new and improved cut-cell-based sharpinterface method for simulating compressible fluid elastic–perfectly plastic solid interaction. Mathematics and Computers in Simulation, 171: 246–263, May 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419301740.

Wang:2020:IMS

[113] Wei Wang, Hanyu Zhang, Tong Li, Jianhua Guo, Jianhua Guo, Wei Huang, Yun Wei, and Jinde Cao. An interpretable model for short term traffic flow prediction. Mathematics and Computers in Simulation, 171:264–278, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303726.

Chen:2020:OAF

[114] Wen Chen and Song Wang. A 2ndorder ADI finite difference method for a 2D fractional Black–Scholes equation governing European two asset option pricing. Mathematics and Computers in Simulation, 171: 279–293, May 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419303192.

Sun:2020:ICP

[115] Li Sun, Haitao Zhu, and Yanhui Ding. Impulsive control for persistence and periodicity of logistic systems. Mathematics and Computers in Simulation, 171:294–305, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302046.

Acherjee:2020:EAP

[116] Simul Acherjee, Ujjwal Kumar Deb, and Md. Moniruzzaman Bhuyan. The effect of the angle of perforation on perforated inserts in a pipe flow for heat transfer analysis. Mathematics and Computers in Simulation, 171:306–314, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303040.

Anonymous:2020:NId

[117] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 171:315, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303842.

Anonymous:2020:ICEd

[118] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 171:316–319, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419303854.

Anonymous:2020:EBg

[119] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 171:ifc, May 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303775.

Sabir:2020:NDM

[120] Zulqurnain Sabir, Hafiz Abdul Wahab, Muhammad Umar, Mehmet Giyas Sakar, and Muhammad Asif Zahoor Raja. Novel design of Morlet wavelet neural network for solving second order Lane–Emden equation. Mathematics and Computers in Simulation, 172:1–14, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420300069.

Anonymous:2020:PJa

[121] Anonymous. Pages 1–346 (June 2020). Mathematics and Computers in Simulation, 172:1–346, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Yaghoubi:2020:RCC

[122] Zahra Yaghoubi. Robust cluster consensus of general fractional-order nonlinear multi agent systems via adaptive sliding mode controller. Mathematics and Computers in Simulation, 172:15–32, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420300033.

Asensio-Sevilla:2020:GSA

[123] M. I. Asensio-Sevilla, M. T. Santos-Martín, D. Álvarez-León, and L. Ferragut-Canals. Global sensitivity analysis of fuel-type-dependent input variables of a simplified physical fire spread model. Mathematics and Computers in Simulation, 172:33–44, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300021.

Li:2020:NVG

[124] Shoujun Li, Yanzi Miao, Guangyu Li, and Muhammad Ikram. A novel varistructure grey forecasting model with speed adaptation and its application. Mathematics and Computers in Simulation, 172: 45–70, June 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542030001X.

Bonab:2020:HOM

[125] Zahra Farzaneh Bonab and Mohammad Javidi. Higher order methods for fractional differential equation based on fractional backward differentiation formula of order three. Mathematics and Computers in Simulation, 172:71–89, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303933.

Malatip:2020:EMT

[126] A. Malatip, N. Prasomsuk, C. Siriparu, N. Paoprasert, and S. Otarawanna.

An efficient matrix tridiagonalization method for 3D finite element analysis of free vibration. Mathematics and Computers in Simulation, 172: 90–110, June 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541930391X.

Filippi:2020:EDM

[127] Miguel G. Filippi, Patrick Kuo-Peng, and Marcelo G. Vanti. Electromagnetic device modeling using a new adaptive wavelet finite element method. Mathematics and Computers in Simulation, 172:111–133, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303908.

Panday:2020:DIM

[128] Pijush Panday, Sudip Samanta, Nikhil Pal, and Joydev Chattopadhyay. Delay induced multiple stability switch and chaos in a predator–prey model with fear effect. Mathematics and Computers in Simulation, 172:134– 158, June 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303751.

Calatayud:2020:ESL

[129] J. Calatayud, J.-C. Cortés, F. A. Dorini, and M. Jornet. Extending the study on the linear advection equation subject to stochastic velocity field and initial condition. Mathematics and Computers in Simulation, 172:159–174, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541930374X.

Herman:2020:TTC

[130] Przemysław Herman and Wojciech Adamski. Trajectory tracking control algorithm in terms of quasi-velocities for a class of vehicles. Mathematics and Computers in Simulation, 172:175–190, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303714.

Singh:2020:FVA

[131] Mehakpreet Singh, Randhir Singh, Sukhjit Singh, Gagandeep Singh, and Gavin Walker. Finite volume approximation of multidimensional aggregation population balance equation on triangular grid. Mathematics and Computers in Simulation, 172:191–212, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303556.

Barrera:2020:QIP

[132] D. Barrera, F. El Mokhtari, M. J. Ibáñez, and D. Sbibih. A quasiinterpolation product integration based method for solving Love’s integral equation with a very small parameter. Mathematics and Computers in Simulation, 172:213–223, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419303544.

Adak:2020:VEM

[133] D. Adak and S. Natarajan. Virtual element method for semilinear sine–Gordon equation over polygonal mesh using product approximation technique. Mathematics and Computers in Simulation, 172:224– 243, June 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419303532.

Li:2020:LTN

[134] Dongfang Li and Chengjian Zhang. Long time numerical behaviors of fractional pantograph equations. Mathematics and Computers in Simulation, 172:244–257, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303507.

Guimaraes:2020:GOM

[135] O. Guimarães, W. Labecca, and José Roberto C. Piqueira. Generalized operational matrices and error bounds for polynomial basis. Mathematics and Computers in Simulation, 172: 258–272, June 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419303490.

Jeganathan:2020:TPH

[136] K. Jeganathan and M. Abdul Reiyas. Two parallel heterogeneous servers Markovian inventory system with modified and delayed working vacations. Mathematics and Computers in Simulation, 172:273–304, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475419303489.

Patel:2020:DAS

[137] Reena R. Patel, David S. Thompson, Guillermo A. Riveros, Wayne D. Hodo, John F. Peters, and Felipe J. Acosta. Dimensional analysis of structural response in complex biological structures. Mathematics and Computers in Simulation, 172:305–320, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303465.

Eshaghi:2020:HBC

[138] Shiva Eshaghi, Reza Khoshsiar Ghaziani, and Alireza Ansari. Hopf bifurcation, chaos control and synchronization of a chaotic fractional-order system with chaos entanglement function. Mathematics and Computers in Simulation, 172:321–340, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303428.

Anonymous:2020:NIe

[139] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 172:341, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300343.

Anonymous:2020:ICEe

[140] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 172:342–345, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475420300355.

Anonymous:2020:EBh

[141] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 172:ifc, June 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030029X.

Liu:2020:SRW

[142] Yanhong Liu and Huimin Wang. Simulations of the rectangular waveguide pattern in the complex Maxwell vorticity equations by lattice Boltzmann method. Mathematics and Computers in Simulation, 173:1– 15, July 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420300239.

Anonymous:2020:PJb

[143] Anonymous. Pages 1–120 (July 2020). Mathematics and Computers in Simulation, 173:1–120, July 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Simorgh:2020:SIC

[144] Abolfazl Simorgh, Abolhassan Razminia, and Vladimir I. Shiryaev. System identification and control design of a nonlinear continuously stirred tank reactor. Mathematics and Computers in Simulation, 173: 16–31, July 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420300173.

Dimov:2020:PAS

[145] I. Dimov and M. Savov. Probabilistic analysis of the Single Particle Wigner Monte-Carlo method. Mathematics and Computers in Simulation, 173:32–50, July 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030015X.

Kalateh:2020:SCF

[146] Farhoud Kalateh and Ali Koosheh. Simulation of cavitating fluid–Structure interaction using SPH–FE method. Mathematics and Computers in Simulation, 173:51–70, July 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300264.

Gonzalez:2020:VDI

[147] C. González, G. Albrecht, and M. Paluszny. Visualization of dental information within CT volumes. Mathematics and Computers in Simulation, 173:71–84, July 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300161.

Yu:2020:HER

[148] Xing Yu, Wei Guo Zhang, Yong Jun Liu, Xinxin Wang, and Chao Wang. Hedging the exchange rate risk for international portfolios. Mathematics and Computers in Simulation, 173: 85–104, July 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420300537.

Lin:2020:RBF

[149] Ji Lin, Yuxiang Zhao, Daniel Watson, and C. S. Chen. The radial basis function differential quadrature method with ghost points. Mathematics and Computers in Simulation, 173:105–114, July 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420300100.

Anonymous:2020:NIf

[150] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 173:115, July 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420300719.

Anonymous:2020:ICEf

[151] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 173:116–119, July 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475420300720.

Anonymous:2020:EBi

[152] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 173:ifc, July 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300665.

Zhong:2020:GEL

[153] Penghong Zhong, Ganshan Yang, and Xuan Ma. Global existence of Landau– Lifshitz–Gilbert equation and selfsimilar blowup of harmonic map heat flow on S^{2}. Mathematics and Computers in Simulation, 174:1–18, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2020:PAb

[154] Anonymous. Pages 1–248 (August 2020). Mathematics and Computers in Simulation, 174:1–248, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Hamiaz:2020:ENA

[155] Adnane Hamiaz, Xavier Ferrieres, and Olivier Pascal. Efficient numerical algorithm to simulate a 3D coupled Maxwell–plasma problem. Mathematics and Computers in Simulation, 174: 19–31, August 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Iqbal:2020:CBS

[156] Azhar Iqbal, Nur Nadiah Abd Hamid, and Ahmad Izani Md. Ismail. Cubic B-spline Galerkin method for numerical solution of the coupled nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 174:32–44, August 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Fanelli:2020:NDM

[157] Viviana Fanelli and Lucia Maddalena. A nonlinear dynamic model for credit risk contagion. Mathematics and Computers in Simulation, 174:45–58, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tang:2020:IRK

[158] Xiao Tang and Aiguo Xiao. Improved Runge–Kutta–Chebyshev methods. Mathematics and Computers in Simulation, 174:59–75, August 2020. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liu:2020:DAF

[159] Jingsen Liu, Yinan Mao, Xiaozhen Liu, and Yu Li. A dynamic adaptive firefly algorithm with globally orientation. Mathematics and Computers in Simulation, 174:76–101, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Bhal:2020:HOO

[160] Santosh Kumar Bhal, P. Danumjaya, and G. Fairweather. High-order orthogonal spline collocation methods for twopoint boundary value problems with interfaces. Mathematics and Computers in Simulation, 174:102–122, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Singh:2020:HWQ

[161] Randhir Singh, Vandana Guleria, and Mehakpreet Singh. Haar wavelet quasilinearization method for numerical solution of Emden–Fowler type equations. Mathematics and Computers in Simulation, 174:123–133, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liu:2020:SAB

[162] Jin Liu, Jigui Jian, and Baoxian Wang. Stability analysis for BAM quaternionvalued inertial neural networks with time delay via nonlinear measure approach. Mathematics and Computers in Simulation, 174:134–152, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liu:2020:QMF

[163] Jun Liu, Hongfei Fu, and Jiansong Zhang. A QSC method for fractional subdiffusion equations with fractional boundary conditions and its application in parameters identification. Mathematics and Computers in Simulation, 174:153–174, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ibrahim:2020:AFB

[164] Rabha W. Ibrahim. Arched foot based on conformal complex neural network testing. Mathematics and Computers in Simulation, 174:175–182, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Wei:2020:DLD

[165] H. Wei, Q. X. Pan, O. B. Adetoro, E. Avital, Y. Yuan, and P. H. Wen. Dynamic large deformation analysis of a cantilever beam. Mathematics and Computers in Simulation, 174:183–204, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420300616.

Milovanovi:2020:HOM

[166] Slobodan Milovanovi and Lina von Sydow. A high order method for pricing of financial derivatives using Radial Basis Function generated Finite Differences. Mathematics and Computers in Simulation, 174:205–217, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mbroh:2020:UCF

[167] Nana Adjoah Mbroh, Suares Clovis Oukouomi Noutchie, and Rodrigue Yves M. pika Massoukou. A uniformly convergent finite difference scheme for Robin type singularly perturbed parabolic convection diffusion problem. Mathematics and Computers in Simulation, 174:218–232, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Garza-Gonzalez:2020:ISS

[168] E. Garza-González, C. Posadas-Castillo, D. López-Mancilla, and A. G. Soriano-Sánchez. Increasing synchronizability in a scale-free network via edge elimination. Mathematics and Computers in Simulation, 174:233–243, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Anonymous:2020:NIg

[169] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 174:244, August 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Zhong:2020:ICE

[170] Penghong Zhong, Ganshan Yang, and Xuan Ma. IMACS calendar of events. Mathematics and Computers in Simulation, 174:245– 248, August 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420300926.

Anonymous:2020:EBj

[171] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 174:ifc, August 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300860.

Elsawah:2020:PSI

[172] Sondoss Elsawah and Anthony J. Jakeman. Preface for the special issue on good modelling practices. Mathematics and Computers in Simulation, 175:1–2, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300112.

Moallemi:2020:SGM

[173] Enayat A. Moallemi, Sondoss Elsawah, and Michael J. Ryan. Strengthening ‘good’ modelling practices in robust decision support: a reporting guideline for combining multiple model-based methods. Mathematics and Computers in Simulation, 175:3–24, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419301508.

Cleary:2020:WWP

[174] Paul W. Cleary, David Thomas, Lachlan Hetherton, Matt Bolger, James E. Hilton, and Damien Watkins. Workspace: a workflow platform for supporting development and deployment of modelling and simulation. Mathematics and Computers in Simulation, 175:25–61, September 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303441.

Hilton:2020:RHF

[175] James E. Hilton, Justin E. Leonard, Raphaele Blanchi, Glenn J. Newnham, Kimberley Opie, Anthony Power, Chris Rucinski, and William Swedosh. Radiant heat flux modelling for wildfires. Mathematics and Computers in Simulation, 175: 62–80, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541930223X.

Moinuddin:2020:MTF

[176] K. A. M. Moinuddin and D. Sutherland. Modelling of tree fires and fires transitioning from the forest floor to the canopy with a physicsbased model. Mathematics and Computers in Simulation, 175:81– 95, September 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419301995.

Thomas:2020:TVA

[177] C. M. Thomas, J. J. Sharples, and J. P. Evans. The terminal-velocity assumption in simulations of longrange ember transport. Mathematics and Computers in Simulation, 175: 96–107, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541930237X.

Tang:2020:SPF

[178] Fiona H. M. Tang, Daniele la Cecilia, R. Willem Vervoort, Nicholas V. Coleman, Chris Conoley, and Federico Maggi. A simple pre-factor for contaminant biodegradation potential and its application to pesticides risk assessment. Mathematics and Computers in Simulation, 175: 108–120, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302381.

laCecilia:2020:ISU

[179] Daniele la Cecilia and Federico Maggi. Influential sources of uncertainty in glyphosate biochemical degradation in soil. Mathematics and Computers in Simulation, 175:121–139, September 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420300045.

Wang:2020:UBP

[180] Qian Wang, Rui Zou, Alvi Khalid, and Tao Yang. Uncertainty-based parameter estimation for urban pollutant buildup and washoff simulation using a multiple pattern inverse modeling approach. Mathematics and Computers in Simulation, 175: 140–152, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302241.

Aryal:2020:FBL

[181] Nanda R. Aryal and Owen D. Jones. Fitting the Bartlett–Lewis rainfall model using Approximate Bayesian Computation. Mathematics and Computers in Simulation, 175:153–163, September 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475419303313.

Alsaedi:2020:RBE

[182] Yasir Hamad Alsaedi and Gurudeo Anand Tularam. The relationship between electricity consumption, peak load and GDP in Saudi Arabia: a VAR analysis. Mathematics and Computers in Simulation, 175: 164–178, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302101.

Salehipour:2020:ASM

[183] Amir Salehipour. An algorithm for single- and multiple-runway aircraft landing problem. Mathematics and Computers in Simulation, 175:179– 191, September 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541930309X.

Mayfield:2020:USA

[184] Helen J. Mayfield, Edoardo Bertone, Carl Smith, and Oz Sahin. Use of a structure aware discretisation algorithm for Bayesian networks applied to water quality predictions. Mathematics and Computers in Simulation, 175:192–201, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302204.

Cohen:2020:DMB

[185] Raymond C. Z. Cohen, Simon M. Harrison, and Paul W. Cleary. Dive Mechanic: Bringing 3D virtual experimentation using biomechanical modelling to elite level diving with the Workspace workflow engine. Mathematics and Computers in Simulation, 175:202–217, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300811.

Anonymous:2020:NIh

[186] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 175:218, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301038.

Anonymous:2020:ICEg

[187] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 175:219–222, September 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542030104X.

Anonymous:2020:EBk

[188] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 175:ifc, September 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420300963.

Carillo:2020:ASC

[189] Sandra Carillo, Costanza Conti, Daniela Mansutti, Francesca Pitolli, and Rosa Maria Spitaleri. Applied scientific computing XV: Innovative modeling and simulation in sciences. Mathematics and Computers in Simulation, 176:1–3, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301269.

Arandiga:2020:RDF

[190] Francesc Aràndiga, Rosa Donat, Lucia Romani, and Milvia Rossini. On the reconstruction of discontinuous functions using multiquadric RBF–WENO local interpolation techniques. Mathematics and Computers in Simulation, 176:4–24, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300252.

Barrera:2020:TNB

[191] D. Barrera, C. Dagnino, M. J. Ibáñez, and S. Remogna. A trivariate near-best blending quadratic quasi-interpolant. Mathematics and Computers in Simulation, 176:25– 35, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303064.

Bertolazzi:2020:PDR

[192] Enrico Bertolazzi, Marco Frego, and Francesco Biral. Point data reconstruction and smoothing using cubic splines and clusterization. Mathematics and Computers in Simulation, 176:36–56, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301130.

Bertolazzi:2020:EIB

[193] Enrico Bertolazzi, Paolo Bevilacqua, and Marco Frego. Efficient intersection between splines of clothoids. Mathematics and Computers in Simulation, 176:57–72, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302897.

Bruni:2020:APF

[194] V. Bruni, L. Della Cioppa, and D. Vitulano. An automatic and parameter-free information-based method for sparse representation in wavelet bases. Mathematics and Computers in Simulation, 176:73–95, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302848.

Bruni:2020:IAS

[195] Vittoria Bruni, Michela Tartaglione, and Domenico Vitulano. An iterative approach for spectrogram reassignment of frequency modulated multicomponent signals. Mathematics and Computers in Simulation, 176:96–119, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303398.

Carillo:2020:PTR

[196] Sandra Carillo and Pedro M. Jordan. On the propagation of temperaturerate waves and traveling waves in rigid conductors of the Graffi–Franchi– Straughan type. Mathematics and Computers in Simulation, 176:120– 133, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420300240.

Conti:2020:CAC

[197] Costanza Conti, Nira Dyn, and Lucia Romani. Convergence analysis of corner cutting algorithms refining nets of functions. Mathematics and Computers in Simulation, 176: 134–146, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300197.

Cotronei:2020:MFI

[198] Mariantonia Cotronei, Dörte Rüweler, and Tomas Sauer. Multiple filterbanks for image processing: Implementation issues. Mathematics and Computers in Simulation, 176:147– 159, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847541930343X.

DeCanditiis:2020:GAL

[199] Daniela De Canditiis. A global approach for learning sparse Ising models. Mathematics and Computers in Simulation, 176:160–170, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420300513.

Francomano:2020:NIS

[200] Elisa Francomano and Marta Paliaga. A normalized iterative Smoothed Particle Hydrodynamics method. Mathematics and Computers in Simulation, 176:171–180, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419303052.

Giorgi:2020:TAH

[201] C. Giorgi and A. Morro. A thermodynamic approach to hysteretic models in ferroelectrics. Mathematics and Computers in Simulation, 176: 181–194, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302290.

Huning:2020:PRH

[202] Svenja Hüning. Polynomial reproduction of Hermite subdivision schemes of any order. Mathematics and Computers in Simulation, 176:195– 205, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303568.

Izzo:2020:SSP

[203] Giuseppe Izzo and Zdzislaw Jackiewicz. Strong stability preserving implicit–explicit transformed general linear methods. Mathematics and Computers in Simulation, 176:206– 225, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303416.

Mezzadri:2020:CMB

[204] Francesco Mezzadri and Emanuele Galligani. On the convergence of modulus-based matrix splitting methods for horizontal linear complementarity problems in hydrodynamic lubrication. Mathematics and Computers in Simulation, 176:226– 242, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420300215.

Morro:2020:ORE

[205] A. Morro and C. Giorgi. Objective rate equations and memory properties in continuum physics. Mathematics and Computers in Simulation, 176:243–253, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475419302319.

Padron:2020:PAO

[206] Miguel A. Padrón and Ángel Plaza. The 8T-LE partition applied to the obtuse triangulations of the 3D-cube. Mathematics and Computers in Simulation, 176:254–265, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420300185.

Pellegrino:2020:CMS

[207] E. Pellegrino, L. Pezza, and F. Pitolli. A collocation method in spline spaces for the solution of linear fractional dynamical systems. Mathematics and Computers in Simulation, 176:266– 278, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475419303520.

Sissouno:2020:DSP

[208] Nada Sissouno, Florian Boßmann, Frank Filbir, Mark Iwen, Maik Kahnt, Rayan Saab, Christian Schroer, and Wolfgang zu Castell. A direct solver for the phase retrieval problem in ptychographic imaging. Mathematics and Computers in Simulation, 176:

292–300, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303362.

Tinoco-Guerrero:2020:SSG

[209] G. Tinoco-Guerrero, F. J. Domínguez-Mota, and J. G. Tinoco-Ruiz. A study of the stability for a generalized finite-difference scheme applied to the advection–diffusion equation. Mathematics and Computers in Simulation, 176:301–311, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420300380.

Zullo:2020:DZS

[210] F. Zullo. On the dynamics of the zeros of solutions of the Airy equation. Mathematics and Computers in Simulation, 176:312–318, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419302174.

Anonymous:2020:NIi

[211] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 176:319, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030135X.

Anonymous:2020:ICEh

[212] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 176:320–323, October 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301361.

Anonymous:2020:EBl

[213] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 176:ifc, October 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301282.

Li:2020:FPI

[214] Xiaodong Li, Haitao Li, Yalu Li, and Xinrong Yang. Function perturbation impact on stability in distribution of probabilistic Boolean networks. Mathematics and Computers in Simulation, 177:1–12, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301208.

Anonymous:2020:PN

[215] Anonymous. Pages 1–678 (November 2020). Mathematics and Computers in Simulation, 177:1–678, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ahmad:2020:SNS

[216] Hijaz Ahmad, Aly R. Seadawy, and Tufail A. Khan. Study on numerical solution of dispersive water wave phenomena by using a reliable modification of variational iteration algorithm. Mathematics and Computers in Simulation, 177:13–23, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301178.

Poette:2020:SCG

[217] Gaël Poëtte. Spectral convergence of the generalized Polynomial Chaos reduced model obtained from the uncertain linear Boltzmann equation. Mathematics and Computers in Simulation, 177:24–45, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030121X.

Yaghoobi:2020:POS

[218] Tahere Yaghoobi. Parameter optimization of software reliability models using improved differential evolution algorithm. Mathematics and Computers in Simulation, 177:46– 62, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420301142.

Mazzia:2020:NSV

[219] Annamaria Mazzia. A numerical study of the virtual element method in anisotropic diffusion problems. Mathematics and Computers in Simulation, 177:63–85, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030118X.

Balyan:2020:SAH

[220] L. K. Balyan, A. K. Mittal, M. Kumar, and M. Choube. Stability analysis and highly accurate numerical approximation of Fisher’s equations using pseudospectral method. Mathematics and Computers in Simulation, 177:86–104, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301233.

Martalo:2020:MMO

[221] Giorgio Martalò, Cesidio Bianchi, Bruno Buonomo, Massimo Chiappini, and Vincenzo Vespri. Mathematical modeling of oxygen control in biocell composting plants. Mathematics and Computers in Simulation, 177: 105–119, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301245.

Li:2020:GDB

[222] Wenjie Li, Lihong Huang, Zhenyuan Guo, and Jinchen Ji. Global dynamic behavior of a plant disease model with ratio dependent impulsive control strategy. Mathematics and Computers in Simulation, 177: 120–139, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300835.

Chang:2020:SAR

[223] Shun-Chang Chang. Stability analysis, routes to chaos, and quenching chaos in electromechanical valve actuators. Mathematics and Computers in Simulation, 177:140–151, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420301488.

Liu:2020:FFV

[224] Chein-Shan Liu and Botong Li. Forced and free vibrations of composite beams solved by an energetic boundary functions collocation method. Mathematics and Computers in Simulation, 177: 152–168, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301476.

Saikia:2020:EHC

[225] Munmi Saikia, Atasi Patra Maiti, and Anuradha Devi. Effect of habitat complexity on rhinoceros and tiger population model with additional food and poaching in Kaziranga National Park, Assam. Mathematics and Computers in Simulation, 177: 169–191, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301191.

Zhang:2020:CEM

[226] Zhenzhong Zhang, Tiandao Zhou, Xinghu Jin, and Jinying Tong. Convergence of the Euler–Maruyama method for CIR model with Markovian switching. Mathematics and Computers in Simulation, 177:192– 210, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420301257.

Salahuddin:2020:IDN

[227] T. Salahuddin, Nazim Siddique, and Maryam Arshad. Insight into the dynamics of the Non-Newtonian Casson fluid on a horizontal object with variable thickness. Mathematics and Computers in Simulation, 177:211– 231, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301579.

Valueva:2020:ARN

[228] M. V. Valueva, N. N. Nagornov, P. A. Lyakhov, G. V. Valuev, and N. I. Chervyakov. Application of the residue number system to reduce hardware costs of the convolutional neural network implementation. Mathematics and Computers in Simulation, 177: 232–243, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301580.

Halder:2020:HOS

[229] Asim Halder, Nitai Pal, and Debasish Mondal. Higher order sliding mode STATCOM control for power system stability improvement. Mathematics and Computers in Simulation, 177: 244–262, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301592.

Mukherjee:2020:RFP

[230] Debasis Mukherjee. Role of fear in predator–prey system with intraspecific competition. Mathematics and Computers in Simulation, 177:263– 275, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542030152X.

Duan:2020:EMS

[231] Wenyong Duan, Yan Li, Yi Sun, Jian Chen, and Xiaodong Yang. Enhanced master–slave synchronization criteria for chaotic Lur’e systems based on time-delayed feedback control. Mathematics and Computers in Simulation, 177:276–294, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301221.

Hashemi:2020:SFP

[232] M. S. Hashemi, A. Atangana, and S. Hajikhah. Solving fractional pantograph delay equations by an effective computational method. Mathematics and Computers in Simulation, 177:295–305, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301518.

Owen:2020:SFD

[233] Livia Owen and Johan Matheus Tuwankotta. On slow–fast dynamics in a classical predator–prey system. Mathematics and Computers in Simulation, 177:306–315, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301622.

Gao:2020:MQI

[234] Wenwu Gao, Xia Zhang, and Xuan Zhou. Multiquadric quasi-interpolation for integral functionals. Mathematics and Computers in Simulation, 177: 316–328, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301415.

Turkyilmazoglu:2020:FFH

[235] Mustafa Turkyilmazoglu. On the fluid flow and heat transfer between a cone and a disk both stationary or rotating. Mathematics and Computers in Simulation, 177:329– 340, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301166.

Ghosh:2020:SDK

[236] Sudeshna Ghosh and Manish Kumar. Study of drafting, kissing and tumbling process of two particles with different sizes using immersed boundary method in a confined medium. Mathematics and Computers in Simulation, 177:341–357, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301555.

Idowu:2020:VTC

[237] A. S. Idowu and B. O. Falodun. Variable thermal conductivity and viscosity effects on non-Newtonian fluids flow through a vertical porous plate under Soret-Dufour influence. Mathematics and Computers in Simulation, 177:358–384, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301609.

Ju:2020:PRM

[238] Guoliang Ju, Mingchao Cai, Jingzhi Li, and Jing Tian. Parameterrobust multiphysics algorithms for Biot model with application in brain edema simulation. Mathematics and Computers in Simulation, 177:385– 403, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301531.

Li:2020:DPF

[239] Meng Li, Mingfa Fei, Nan Wang, and Chengming Huang. A dissipationpreserving finite element method for nonlinear fractional wave equations on irregular convex domains. Mathematics and Computers in Simulation, 177: 404–419, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301646.

Kim:2020:SEV

[240] Seong-Tae Kim and Jeong-Hoon Kim. Stochastic elasticity of vol-of-vol and pricing of variance swaps. Mathematics and Computers in Simulation, 177: 420–440, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301154.

Alghafis:2020:EIE

[241] Abdullah Alghafis, Faiza Firdousi, Majid Khan, Syeda Iram Batool, and Muhammad Amin. An efficient image encryption scheme based on chaotic and Deoxyribonucleic acid sequencing. Mathematics and Computers in Simulation, 177:441–466, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301749.

Ferreiro-Ferreiro:2020:NCH

[242] Ana María Ferreiro-Ferreiro, José A. García-Rodríguez, Luis Souto, and Carlos Vázquez. A new calibration of the Heston Stochastic Local Volatility Model and its parallel implementation on GPUs. Mathematics and Computers in Simulation, 177: 467–486, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301129.

Chang:2020:LMM

[243] Wanru Chang, C. S. Chen, Xiao-Yan Liu, and J. Li. Localized meshless methods based on polynomial basis functions for solving axisymmetric equations. Mathematics and Computers in Simulation, 177: 487–499, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301658.

Chen:2020:LSF

[244] C. S. Chen, Shu-Hui Shen, Fangfang Dou, and J. Li. The LMAPS for solving fourth-order PDEs with polynomial basis functions. Mathematics and Computers in Simulation, 177: 500–515, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301725.

Liu:2020:ECS

[245] Junrong Liu, Qihong Duan, and Wen-Xiu Ma. The evolution of a clogging sidewalk caused by a dockless bicycle-sharing system: a stochastic particles model. Mathematics and Computers in Simulation, 177: 516–526, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301567.

Li:2020:MPD

[246] Ran Li, Zhongguo Zhou, Lin Li, Yan Wang, Hao Pan, Ruiqi Dong, and Jing Zhou. The mass-preserving domain decomposition scheme for solving three-dimensional convection– diffusion equations. Mathematics and Computers in Simulation, 177: 527–555, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301634.

Akinyemi:2020:PAS

[247] Lanre Akinyemi and Shaheed N. Huseen. A powerful approach to study the new modified coupled Korteweg– de Vries system. Mathematics and Computers in Simulation, 177:556– 567, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301804.

Kozpinar:2020:PEA

[248] Sinem Kozpınar, Murat Uzunca, and Bülent Karasözen. Pricing European and American options under Heston model using discontinuous Galerkin finite elements. Mathematics and Computers in Simulation, 177:568– 587, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301816.

Zhang:2020:EMO

[249] XuWei Zhang, Hao Liu, LiangPing Tu, and Jian Zhao. An efficient multiobjective optimization algorithm based on level swarm optimizer. Mathematics and Computers in Simulation, 177: 588–602, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301841.

Marconi:2020:SSA

[250] Silvia Marconi and Claudio De Lazzari. In silico study of airway/lung mechanics in normal human breathing. Mathematics and Computers in Simulation, 177:603–624, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301737.

Wang:2020:MWI

[251] Jing Wang, Shucai Li, Liping Li, Shaoshuai Shi, Zongqing Zhou, and Shuguang Song. Mechanism of water inrush in fractures and block collapse under hydraulic pressure. Mathematics and Computers in Simulation, 177:625–642, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301877.

Hashan:2020:MFF

[252] Mahamudul Hashan, Labiba Nusrat Jahan, Tareq-Uz-Zaman, Syed Imtiaz, and M. Enamul Hossain. Modelling of fluid flow through porous media using memory approach: a review. Mathematics and Computers in Simulation, 177:643–673, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301853.

Anonymous:2020:NIj

[253] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 177:674, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302020.

Anonymous:2020:ICEi

[254] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 177:675–678, November 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302032.

Anonymous:2020:EBm

[255] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 177:ifc, November 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030197X.

Ghasemian:2020:HUA

[256] Hadi Ghasemian, Fahimeh Ghasemian, and Hamed Vahdat-Nejad. Human urbanization algorithm: a novel metaheuristic approach. Mathematics and Computers in Simulation, 178:1–15, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301828.

Anonymous:2020:PD

[257] Anonymous. Pages 1–672 (December 2020). Mathematics and Computers in Simulation, 178:1–672, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Li:2020:IMS

[258] Dagen Li and Meng Liu. Invariant measure of a stochastic foodlimited population model with regime switching. Mathematics and Computers in Simulation, 178:16–26, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420301932.

Rabbani:2020:STC

[259] Attia Rabbani and Waqas Ashraf. A space time conservation element and solution element method for solving two-species chemotaxis model. Mathematics and Computers in Simulation, 178:27–45, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301920.

Grunwald:2020:NIP

[260] Norbert Grunwald, Jobst Maßmann, Olaf Kolditz, and Thomas Nagel. Noniterative phase-equilibrium model of the H_{2}O–CO_{2}–NaCl-system for largescale numerical simulations. Mathematics and Computers in Simulation, 178:46–61, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030183X.

Ziadi:2020:DMC

[261] Raouf Ziadi, Abdelatif Bencherif-Madani, and Rachid Ellaia. A deterministic method for continuous global optimization using a dense curve. Mathematics and Computers in Simulation, 178:62– 91, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420301907.

Jia:2020:BPF

[262] Yunfeng Jia. Bifurcation and pattern formation of a tumor–immune model with time-delay and diffusion. Mathematics and Computers in Simulation, 178:92–108, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302123.

Xing:2020:GFD

[263] Yanan Xing, Lina Song, Xiaoming He, and Changxin Qiu. A generalized finite difference method for solving elliptic interface problems. Mathematics and Computers in Simulation, 178: 109–124, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030207X.

Wongsaijai:2020:CSP

[264] B. Wongsaijai, C. Oonariya, and K. Poochinapan. Compact structurepreserving algorithm with high accuracy extended to the improved Boussinesq equation. Mathematics and Computers in Simulation, 178: 125–150, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301610.

Qureshi:2020:PDR

[265] Sania Qureshi. Periodic dynamics of rubella epidemic under standard and fractional Caputo operator with real data from Pakistan. Mathematics and Computers in Simulation, 178:

151–165, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301919.

Waqas:2020:NSB

[266] Hassan Waqas, Muhammad Imran, Sajjad Hussain, Farooq Ahmad, Ilyas Khan, Kottakkaran Sooppy Nisar, and A. Othman Almatroud. Numerical simulation for bioconvection effects on MHD flow of Oldroyd-B nanofluids in a rotating frame stretching horizontally. Mathematics and Computers in Simulation, 178: 166–182, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301944.

Aderogba:2020:NFD

[267] A. A. Aderogba, O. O. Fabelurin, S. O. Akindeinde, A. O. Adewumi, and B. S. Ogundare. Nonstandard finite difference approximation for a generalized Fins problem. Mathematics and Computers in Simulation, 178: 183–191, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302135.

SanMartin:2020:UPG

[268] Jesús San Martín, Fátima Drubi, and Daniel Rodríguez Pérez. Uncritical polarized groups: the impact of spreading fake news as fact in social networks. Mathematics and Computers in Simulation, 178: 192–206, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302159.

Naseer:2020:AIE

[269] Yasir Naseer, Tariq Shah, Attaullah, and Adnan Javeed. Advance image encryption technique utilizing compression, dynamical system and S-boxes. Mathematics and Computers in Simulation, 178:207–217, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302081.

Das:2020:DFO

[270] Meghadri Das and G. P. Samanta. A delayed fractional order food chain model with fear effect and prey refuge. Mathematics and Computers in Simulation, 178:218–245, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302172.

Xiao:2020:IFT

[271] Lin Xiao, Qian Yi, Qiuyue Zuo, and Yongjun He. Improved finitetime zeroing neural networks for timevarying complex Sylvester equation solving. Mathematics and Computers in Simulation, 178:246–258, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302160.

Yu:2020:QLM

[272] Caiyang Yu, Zhennao Cai, Xiaojia Ye, Mingjing Wang, Xuehua Zhao, Guoxi Liang, Huiling Chen, and Chengye Li. Quantum-like mutationinduced dragonfly-inspired optimization approach. Mathematics and Computers in Simulation, 178:259– 289, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420302147.

Bhandari:2020:SFF

[273] Anupam Bhandari. Study of ferrofluid flow in a rotating system through mathematical modeling. Mathematics and Computers in Simulation, 178: 290–306, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302202.

Montanelli:2020:SPS

[274] Hadrien Montanelli and Niall Bootland. Solving periodic semilinear stiff PDEs in 1D, 2D and 3D with exponential integrators. Mathematics and Computers in Simulation, 178: 307–327, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030210X.

Zhang:2020:NGS

[275] Zhe Zhang, Jing Zhang, Zhaoyang Ai, FanYong Cheng, and Feng Liu. A novel general stability criterion of time-delay fractional-order nonlinear systems based on WILL Deduction Method. Mathematics and Computers in Simulation, 178:328– 344, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302214.

Safan:2020:HMH

[276] Muntaser Safan. Hybrid models for helminth infections with special reference to onchocerciasis. Mathematics and Computers in Simulation, 178: 345–365, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302251.

Zhao:2020:EER

[277] Jingjun Zhao, Rui Zhan, and Yang Xu. Explicit exponential Runge–Kutta methods for semilinear parabolic delay differential equations. Mathematics and Computers in Simulation, 178: 366–381, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302275.

Bastl:2020:AGM

[278] Bohumír Bastl and Kristýna Michálková. Automatic generators of multi-patch B-spline meshes of blade cascades and their comparison. Mathematics and Computers in Simulation, 178: 382–406, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302305.

Liu:2020:GAT

[279] Jian-Gen Liu, Xiao-Jun Yang, Yi-Ying Feng, and Ping Cui. On group analysis of the time fractional extended (2 + 1)-dimensional Zakharov–Kuznetsov equation in quantum magneto-plasmas. Mathematics and Computers in Simulation, 178: 407–421, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302329.

Matsumoto:2020:SSC

[280] Akio Matsumoto and Ferenc Szidarovszky. Stability switching curves in a Lotka–Volterra competition system with two delays. Mathematics and Computers in Simulation, 178: 422–438, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302184.

Jokar:2020:PCC

[281] H. Jokar, H. Zeinali, and M. H. Tamaddondar. Planing craft control using pneumatically driven trim tab. Mathematics and Computers in Simulation, 178:439–463, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301889.

Zheng:2020:LPS

[282] Bo Zheng and Yueqiang Shang. Local and parallel stabilized finite element algorithms based on the lowest equal-order elements for the steady Navier–Stokes equations. Mathematics and Computers in Simulation, 178: 464–484, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030238X.

Singh:2020:LWB

[283] Jitendra Singh, Jitendra, and Kabindra Nath Rai. Legendre wavelet based numerical solution of variable latent heat moving boundary problem. Mathematics and Computers in Simulation, 178:485–500, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302226.

Cardone:2020:SAS

[284] Angelamaria Cardone and Dajana Conte. Stability analysis of spline collocation methods for fractional differential equations. Mathematics and Computers in Simulation, 178:501– 514, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302330.

Chen:2020:MCG

[285] Yun Chen, Yanyi Xu, Qian Lin, and Xiyong Zhang. Model and criteria on the global finite-time synchronization of the chaotic gyrostat systems. Mathematics and Computers in Simulation, 178:515–533, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302238.

Sabir:2020:HCT

[286] Zulqurnain Sabir, Sahar Saoud, Muhammad Asif Zahoor Raja, Hafiz Abdul Wahab, and Adnène Arbi. Heuristic computing technique for numerical solutions of nonlinear fourth order Emden–Fowler equation. Mathematics and Computers in Simulation, 178:534–548, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030224X.

Simsek:2020:FFM

[287] Barış Şimşek and Yusuf Tansel İç. Fuzzy failure mode and effect analysis application to reduce risk level in a ready-mixed concrete plant: a fuzzy rule based system modelling approach. Mathematics and Computers in Simulation, 178:549– 587, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420302287.

Sabo:2020:IMM

[288] Kristian Sabo, Danijel Grahovac, and Rudolf Scitovski. Incremental method for multiple line detection problem — iterative reweighted approach. Mathematics and Computers in Simulation, 178:588–602, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302421.

Rahouti:2020:CSQ

[289] A. Rahouti, A. Serghini, and A. Tijini. Construction of superconvergent quasi-interpolants using new normalized C^{2} cubic B-splines. Mathematics and Computers in Simulation, 178: 603–624, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302378.

Graef:2020:SAF

[290] John R. Graef, Lingju Kong, Andrew Ledoan, and Min Wang. Stability analysis of a fractional online social network model. Mathematics and Computers in Simulation, 178: 625–645, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302408.

Malik:2020:CMI

[291] Dania Saleem Malik and Tariq Shah. Color multiple image encryption scheme based on 3D-chaotic maps. Mathematics and Computers in Simulation, 178:646–666, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302342.

Anonymous:2020:NIk

[292] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 178:667, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302585.

Anonymous:2020:ICEj

[293] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 178:668–671, December 2020. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420302597.

Anonymous:2020:EBn

[294] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 178:ifc, December 2020. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302536.

Esfahani:2021:NMC

[295] Ali Kouhi Esfahani, Mahdi Babaei, and Saeid Sarrami-Foroushani. A numerical model coupling phase transformation to predict microstructure evolution and residual stress during quenching of 1045 steel. Mathematics and Computers in Simulation, 179:1–22, January 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302445.

Anonymous:2021:PJa

[296] Anonymous. Pages 1–284 (January 2021). Mathematics and Computers in Simulation, 179:1–284, January 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

BenRahmoune:2021:GTM

[297] Mohamed Ben Rahmoune, Ahmed Hafaifa, Abdellah Kouzou, XiaoQi Chen, and Ahmed Chaibet. Gas turbine monitoring using neural network dynamic nonlinear autoregressive with external exogenous input modelling. Mathematics and Computers in Simulation, 179:23– 47, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302457.

Tang:2021:FDM

[298] Changyang Tang and Chengjian Zhang. A fully discrete θ-method for solving semi-linear reaction– diffusion equations with time-variable delay. Mathematics and Computers in Simulation, 179:48–56, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302470.

Jiang:2021:SBG

[299] Yan Jiang and Qi-Xing Qu. Solitons and breathers for a generalized nonlinear Schrödinger equation via the binary Bell polynomials. Mathematics and Computers in Simulation, 179:57–68, January 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302494.

Ghasemi:2021:RSO

[300] Peiman Ghasemi and Kaveh Khalili-Damghani. A robust simulationoptimization approach for pre-disaster multi-period location–allocation–inventory planning. Mathematics and Computers in Simulation, 179:69–95, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420302500.

Khalil:2021:DGM

[301] Omar A. Khalil and Gerd Baumann. Discontinuous Galerkin methods using poly-sinc approximation. Mathematics and Computers in Simulation, 179:96–110, January 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302512.

Bashan:2021:HEA

[302] Ali Başhan. Highly efficient approach to numerical solutions of two different forms of the modified Kawahara equation via contribution of two effective methods. Mathematics and Computers in Simulation, 179:111– 125, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420302718.

Rizvi:2021:SPO

[303] Syed Tahir Raza Rizvi, Salah Ud-Din Khan, Mohsan Hassan, Ishrat Fatima, and Shahab Ud-Din Khan. Stable propagation of optical solitons for nonlinear Schrödinger equation with dispersion and self phase modulation. Mathematics and Computers in Simulation, 179:126–136, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302792.

Lamboni:2021:DBI

[304] Matieyendou Lamboni. Derivativebased integral equalities and inequality: a proxy-measure for sensitivity analysis. Mathematics and Computers in Simulation, 179:137– 161, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302706.

Misur:2021:MBV

[305] Marin Mišur. On the mixed boundary value problem for semilinear elliptic equations. Mathematics and Computers in Simulation, 179:162– 177, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542030269X.

Ebrahim:2021:OMB

[306] M. A. Ebrahim, M. N. Ahmed, H. S. Ramadan, M. Becherif, and J. Zhao. Optimal metaheuristic-based sliding mode control of VSC-HVDC transmission systems. Mathematics and Computers in Simulation, 179:178– 193, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302779.

Sedighizadeh:2021:GNG

[307] Davoud Sedighizadeh, Ellips Masehian, Mostafa Sedighizadeh, and Hossein Akbaripour. GEPSO: a new generalized particle swarm optimization algorithm. Mathematics and Computers in Simulation, 179:194– 212, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302743.

Rathinasamy:2021:AES

[308] A. Rathinasamy, M. Chinnadurai, and S. Athithan. Analysis of exact solution of stochastic sex-structured HIV/ AIDS epidemic model with effect of screening of infectives. Mathematics and Computers in Simulation, 179: 213–237, January 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302846.

Wu:2021:FBD

[309] Jun Wu and Yicheng Liu. Flocking behaviours of a delayed collective model with local rule and critical neighbourhood situation. Mathematics and Computers in Simulation, 179:238–252, January 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302822.

Neto:2021:ALS

[310] David Neto. Adaptive LASSO for selecting Fourier coefficients in a functional smooth time-varying cointegrating regression: an application to the Feldstein–Horioka puzzle. Mathematics and Computers in Simulation, 179: 253–264, January 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302780.

Dadvand:2021:LBS

[311] Abdolrahman Dadvand, Sina Hassanzadeh Saraei, Soheila Ghoreishi, and Ali J. Chamkha. Lattice Boltzmann simulation of natural convection in a square enclosure with discrete heating. Mathematics and Computers in Simulation, 179:265– 278, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542030272X.

Anonymous:2021:NIa

[312] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 179:279, January 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303013.

Anonymous:2021:ICEa

[313] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 179:280–283, January 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303025.

Anonymous:2021:EBa

[314] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 179:ifc, January 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302962.

Wang:2021:HBM

[315] Shufan Wang, Haopeng Tang, and Zhihui Ma. Hopf bifurcation of a multiple-delayed predator–prey system with habitat complexity. Mathematics and Computers in Simulation, 180:1–23, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302767.

Anonymous:2021:PF

[316] Anonymous. Pages 1–424 (February 2021). Mathematics and Computers in Simulation, 180:1–424, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Weng:2021:RCS

[317] Xiaoyu Weng, Longxing Qi, and Pan Tang. Research on college students’ physical exercise trend based on compartment model. Mathematics and Computers in Simulation, 180:24–42, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302834.

Belhocine:2021:ASN

[318] Ali Belhocine and Wan Zaidi Wan Omar. Analytical solution and numerical simulation of the generalized Levèque equation to predict the thermal boundary layer. Mathematics and Computers in Simulation, 180:43–60, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302755.

Ran:2021:IDS

[319] Maohua Ran and Xiaoyi Zhou. An implicit difference scheme for the time-fractional Cahn–Hilliard equations. Mathematics and Computers in Simulation, 180:61–71, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302883.

Faheem:2021:CMB

[320] Mo Faheem, Akmal Raza, and Arshad Khan. Collocation methods based on Gegenbauer and Bernoulli wavelets for solving neutral delay differential equations. Mathematics and Computers in Simulation, 180: 72–92, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302858.

Andrushia:2021:ETC

[321] A. Diana Andrushia, N. Anand, and G. Prince Arulraj. Evaluation of thermal cracks on fire exposed concrete structures using Ripplet˘200b transform. Mathematics and Computers in Simulation, 180: 93–113, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302731.

Parand:2021:NAN

[322] K. Parand, A. A. Aghaei, M. Jani, and A. Ghodsi. A new approach to the numerical solution of Fredholm integral equations using least squaressupport vector regression. Mathematics and Computers in Simulation, 180: 114–128, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302809.

Ghaleb:2021:AAS

[323] A. F. Ghaleb, M. S. Abou-Dina, G. M. Moatimid, and M. H. Zekry. Analytic approximate solutions of the cubic– quintic Duffing–van der Pol equation with two-external periodic forcing terms: Stability analysis. Mathematics and Computers in Simulation, 180:129–151, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302664.

Gheisariha:2021:SOM

[324] Elmira Gheisariha, Madjid Tavana, Fariborz Jolai, and Meysam Rabiee. A simulation–optimization model for solving flexible flow shop scheduling problems with rework and transportation. Mathematics and Computers in Simulation, 180:152– 178, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542030286X.

Colombi:2021:IDN

[325] Annachiara Colombi, Silvia Falletta, Marco Scianna, and Letizia Scuderi. An integro-differential non-local model for cell migration and its efficient numerical solution. Mathematics and Computers in Simulation, 180: 179–204, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302871.

Khattak:2021:HPF

[326] Afaq Khattak and Arshad Hussain. Hybrid DES-PSO framework for the design of commuters’ circulation space at multimodal transport interchange. Mathematics and Computers in Simulation, 180:205– 229, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302925.

Ghommam:2021:FTC

[327] Jawhar Ghommam, Maarouf Saad, and Faisal Mnif. Finite-time circular formation around a moving target with multiple underactuated ODIN vehicles. Mathematics and Computers in Simulation, 180:230– 250, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302937.

Guo:2021:MOL

[328] Jutong Guo, Jingsong He, Maohua Li, and Dumitru Mihalache.

Multiple-order line rogue wave solutions of extended Kadomtsev– Petviashvili equation. Mathematics and Computers in Simulation, 180: 251–257, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303141.

Malek:2021:NSR

[329] Mustapha Malek, Nouh Izem, M. Shadi Mohamed, Mohammed Seaid, and Mohamed Wakrim. Numerical solution of Rosseland model for transient thermal radiation in non-grey optically thick media using enriched basis functions. Mathematics and Computers in Simulation, 180:258– 275, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302913.

Butt:2021:TLD

[330] Muhammad Munir Butt. Two-level difference scheme for the two-dimensional Fokker–Planck equation. Mathematics and Computers in Simulation, 180: 276–288, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303098.

Tocino:2021:MSP

[331] A. Tocino, R. Zeghdane, and M. J. Senosiaín. On the MS-stability of predictor–corrector schemes for stochastic differential equations. Mathematics and Computers in Simulation, 180:289–305, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303128.

Jafelice:2021:IAH

[332] R. M. Jafelice and W. A. Lodwick. Interval analysis of the HIV dynamics model solution using type-2 fuzzy sets. Mathematics and Computers in Simulation, 180:306– 327, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302895.

Qin:2021:ICS

[333] Jian Qin, Huachen Pan, M. M. Rahman, Xiaoqing Tian, and Zefei Zhu. Introducing compressibility with SIM-PLE algorithm. Mathematics and Computers in Simulation, 180:328– 353, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303189.

Khajanchi:2021:MAG

[334] Subhas Khajanchi, Sovan Bera, and Tapan Kumar Roy. Mathematical analysis of the global dynamics of a HTLV-I infection model, considering the role of cytotoxic T-lymphocytes. Mathematics and Computers in Simulation, 180:354–378, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303177.

Marinelli:2021:SCL

[335] L. Ponzellini Marinelli, N. Caruso, and M. Portapila. A stable computation on local boundary-domain integral method for elliptic PDEs. Mathematics and Computers in Simulation, 180:379–400, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302949.

Zhao:2021:GAM

[336] Jingjun Zhao, Xingzhou Jiang, and Yang Xu. Generalized Adams method for solving fractional delay differential equations. Mathematics and Computers in Simulation, 180:401– 419, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303153.

Anonymous:2021:NIb

[337] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 180:420, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303633.

Anonymous:2021:ICEb

[338] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 180:421–424, February 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303645.

Anonymous:2021:EBb

[339] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 180:ifc, February 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030358X.

Noorani:2021:GAV

[340] Idin Noorani, Farshid Mehrdoust, and Abdelaziz Nasroallah. A generalized antithetic variates Monte-Carlo simulation method for pricing of Asian option in a Markov regime-switching model. Mathematics and Computers in Simulation, 181:1–15, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303190.

Anonymous:2021:PM

[341] Anonymous. Pages 1–646 (March 2021). Mathematics and Computers in Simulation, 181:1–646, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Lochab:2021:IFL

[342] Ruchika Lochab and Vivek Kumar. An improved flux limiter using fuzzy modifiers for Hyperbolic Conservation Laws. Mathematics and Computers in Simulation, 181: 16–37, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303207.

Dwivedi:2021:NST

[343] Kushal Dhar Dwivedi and Jagdev Singh. Numerical solution of twodimensional fractional-order reaction advection sub-diffusion equation with finite-difference Fibonacci collocation method. Mathematics and Computers in Simulation, 181:38– 50, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303165.

Devroye:2021:RVG

[344] Luc Devroye. Random variate generation for the truncated negative gamma distribution. Mathematics and Computers in Simulation, 181: 51–56, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542030313X.

Campos:2021:DSO

[345] B. Campos and P. Vindel. Dynamics of subfamilies of Ostrowski– Chun methods. Mathematics and Computers in Simulation, 181:57– 81, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303281.

Zaman:2021:TDN

[346] Akbar Zaman and Ambreen Afsar Khan. Time dependent non-Newtonian nano-fluid (blood) flow in w-shape stenosed channel; with curvature effects. Mathematics and Computers in Simulation, 181:82– 97, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542030327X.

Yang:2021:DFS

[347] Xiangfeng Yang and Dan A. Ralescu. A Dufort–Frankel scheme for onedimensional uncertain heat equation. Mathematics and Computers in Simulation, 181:98–112, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475420303359.

Al-Ghafri:2021:CEA

[348] K. S. Al-Ghafri and E. V. Krishnan. Comment on: “Exact analytic solitary wave solutions for the RKL model” [Math. Comput. Simulation 80 (4) (2009) 849–854]. Mathematics and Computers in Simulation, 181:113–116, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301750. See

[1] and reply [389].

Gao:2021:TMF

[349] Xinghua Gao, Baoli Yin, Hong Li, and Yang Liu. TT-M FE method for a 2D nonlinear time distributedorder and space fractional diffusion equation. Mathematics and Computers in Simulation, 181:117– 137, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542030330X.

Postavaru:2021:CPC

[350] O. Postavaru, S. R. Anton, and A. Toma. COVID-19 pandemic and chaos theory. Mathematics and Computers in Simulation, 181:138– 149, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303396.

Takhi:2021:PBS

[351] Hocine Takhi, Karim Kemih, Lazaros Moysis, and Christos Volos. Passivity based sliding mode control and synchronization of a perturbed uncertain unified chaotic system. Mathematics and Computers in Simulation, 181:150–169, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303311.

Noyola-Garcia:2021:SGM

[352] Benjamín Salomón Noyola-García and Suemi Rodriguez-Romo. Simulations of Ga melting based on multiple-relaxation time lattice Boltzmann method performed with CUDA in Python. Mathematics and Computers in Simulation, 181:170– 191, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303347.

Yan:2021:PPA

[353] Zheping Yan, Jinzhong Zhang, and Jialing Tang. Path planning for autonomous underwater vehicle based on an enhanced water wave optimization algorithm. Mathematics and Computers in Simulation, 181:192– 241, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303293.

Liu:2021:RMF

[354] Wenqiang Liu, Guili Tao, and Chen Shen. Robust measurement fusion steady-state estimator design for multisensor networked systems with random two-step transmission delays and missing measurements. Mathematics and Computers in Simulation, 181:242–283, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303219.

Rahmati:2021:ALB

[355] A. R. Rahmati, H. Khorasanizadeh, and M. R. Arabyarmohammadi. Application of lattice Boltzmann method to simulate a pressure-affected electroosmotic pump with hydrophobic thermally-jumped walls and temperature-sensitive operating fluid. Mathematics and Computers in Simulation, 181:284–297, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475420303384.

Prakash:2021:ANS

[356] Amit Prakash and Hardish Kaur. Analysis and numerical simulation of fractional Biswas–Milovic model. Mathematics and Computers in Simulation, 181:298–315, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303268.

Lu:2021:DSM

[357] Chun Lu. Dynamics of a stochastic Markovian switching predator–prey model with infinite memory and general Lévy jumps. Mathematics and Computers in Simulation, 181:316– 332, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303414.

Khader:2021:EUM

[358] M. M. Khader and Ram Prakash Sharma. Evaluating the unsteady MHD micropolar fluid flow past stretching/shirking sheet with heat source and thermal radiation: Implementing fourth order predictor– corrector FDM. Mathematics and Computers in Simulation, 181:333– 350, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303220.

Belomestny:2021:FTM

[359] Denis Belomestny and Leonid Iosipoi. Fourier transform MCMC, heavytailed distributions, and geometric ergodicity. Mathematics and Computers in Simulation, 181:351– 363, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303451.

Kumar:2021:FDL

[360] N. Kishore Kumar and Pankaj Biswas. Fully discrete least-squares spectral element method for parabolic interface problems. Mathematics and Computers in Simulation, 181:364– 379, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303402.

Ren:2021:ALA

[361] Hao Ren, Jun Li, Huiling Chen, and ChenYang Li. Adaptive levy-assisted salp swarm algorithm: Analysis and optimization case studies. Mathematics and Computers in Simulation, 181:380–409, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303323.

Cuevas:2021:ADF

[362] Erik Cuevas, Héctor Becerra, and Alberto Luque. Anisotropic diffusion filtering through multi-objective optimization. Mathematics and Computers in Simulation, 181:410– 429, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303438.

Guo:2021:DCF

[363] Yongfeng Guo, Linjie Wang, Qiang Dong, and Xiaojuan Lou. Dynamical complexity of FitzHugh–Nagumo neuron model driven by Lévy noise and Gaussian white noise. Mathematics and Computers in Simulation, 181:430–443, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303372.

Jadoon:2021:DEO

[364] Ihtesham Jadoon, Muhammad Asif Zahoor Raja, Muhammad Junaid, Ashfaq Ahmed, Ata ur Rehman, and Muhammad Shoaib. Design of evolutionary optimized finite difference based numerical computing for dust density model of nonlinear van der Pol Mathieu’s oscillatory systems. Mathematics and Computers in Simulation, 181:444–470, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030344X.

Chen:2021:SSD

[365] Jianxin Chen, Tonghua Zhang, and Yong wu Zhou. Stochastic sensitivity and dynamical complexity of newsvendor models subject to trade credit. Mathematics and Computers in Simulation, 181:471– 486, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303463.

Salahuddin:2021:BCS

[366] T. Salahuddin, S. Sakinder, Sayer Obaid Alharbi, and Zahra Abdelmalek. A brief comparative study of gamma alumina–water and gamma alumina– EG nanofluids flow near a solid sphere. Mathematics and Computers in Simulation, 181:487–500, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475420303517.

Hu:2021:ETT

[367] Zejing Hu, M. N. Smirnova, Yongliang Zhang, N. N. Smirnov, and Zuojin Zhu. Estimation of travel time through a composite ring road by a viscoelastic traffic flow model. Mathematics and Computers in Simulation, 181:501–521, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303335.

Smolik:2021:RBF

[368] Michal Smolik and Vaclav Skala. Radial basis function and multi-level 2D vector field approximation. Mathematics and Computers in Simulation, 181:522–538, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303499.

Pan:2021:ETD

[369] Suying Pan, Jin Zhou, and Zhiyong Ye. Event-triggered dynamic output feedback control for networked Markovian jump systems with partly unknown transition rates. Mathematics and Computers in Simulation, 181:539–561, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303426.

Yuan:2021:SBA

[370] Jun Yuan, Lingzhi Zhao, Chengdai Huang, and Min Xiao. Stability and bifurcation analysis of a fractional predator–prey model involving two nonidentical delays. Mathematics and Computers in Simulation, 181:562–580, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303530.

Karageorghis:2021:KRM

[371] Andreas Karageorghis, Demetriana Tappoura, and C. S. Chen. The Kansa RBF method with auxiliary boundary centres for fourth order boundary value problems. Mathematics and Computers in Simulation, 181:581–597, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303505.

Wang:2021:HOC

[372] Yuan-Ming Wang. A high-order compact difference method on fitted meshes for Neumann problems of timefractional reaction–diffusion equations with variable coefficients. Mathematics and Computers in Simulation, 181:598–623, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303542.

Xing:2021:EFO

[373] Zhiyong Xing, Liping Wen, and Wansheng Wang. An explicit fourth-order energy-preserving difference scheme for the Riesz space-fractional Sine– Gordon equations. Mathematics and Computers in Simulation, 181:624– 641, March 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303475.

Anonymous:2021:NIc

[374] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 181:642, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030389X.

Anonymous:2021:ICEc

[375] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 181:643–646, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475420303906.

Anonymous:2021:EBc

[376] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 181:ifc, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303840.

Luo:2021:LQS

[377] Wei-Hua Luo, Xian-Ming Gu, Liu Yang, and Jing Meng. A Lagrangequadratic spline optimal collocation method for the time tempered fractional diffusion equation. Mathematics and Computers in Simulation, 182:1–24, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303566.

Anonymous:2021:PA

[378] Anonymous. Pages 1–918 (April 2021). Mathematics and Computers in Simulation, 182:1–918, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Nafidi:2021:SDP

[379] Ahmed Nafidi and Abdenbi El Azri. A stochastic diffusion process based on the Lundqvist–Korf growth: Computational aspects and simulation. Mathematics and Computers in Simulation, 182:25–38, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303700.

Amin:2021:TDS

[380] Fahs Amin, Ali Zakeri, and Adrien Wanko. Time-dependent solution for natural convection in a porous enclosure using the Darcy–Lapwood– Brinkman model. Mathematics and Computers in Simulation, 182:39– 65, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303529.

Kheybari:2021:NAC

[381] Samad Kheybari. Numerical algorithm to Caputo type time–space fractional partial differential equations with variable coefficients. Mathematics and Computers in Simulation, 182:66–85, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303682.

Guo:2021:ENA

[382] Wenjuan Guo and Qimin Zhang. Explicit numerical approximation for an impulsive stochastic age-structured HIV infection model with Markovian switching. Mathematics and Computers in Simulation, 182:86– 115, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303554.

Sinhababu:2021:ACE

[383] Arijit Sinhababu and Sathyanarayana Ayyalasomayajula. Accuracy and computational efficiency of dealiasing schemes for the DNS of under resolved flows with strong gradients. Mathematics and Computers in Simulation, 182:116–142, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303712.

Meftah:2021:SNP

[384] Kamel Meftah, Lakhdar Sedira, and Ayoub Ayadi. A six-node prismatic solid finite element for geometric nonlinear problems in elasticity. Mathematics and Computers in Simulation, 182:143–164, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303724.

Jovanovic:2021:FLS

[385] Aleksandar Jovanović, Katarina Kukić, and Aleksandar Stevanović. A fuzzy logic simulation model for controlling an oversaturated diverge diamond interchange and ramp metering system. Mathematics and Computers in Simulation, 182:165– 181, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303785.

Wang:2021:SBC

[386] Xinhe Wang, Zhen Wang, Junwei Lu, and Bo Meng. Stability, bifurcation and chaos of a discrete-time pair approximation epidemic model on adaptive networks. Mathematics and Computers in Simulation, 182:182–194, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303694.

Tan:2021:PCB

[387] Jianguo Tan, Yang Chen, Weiwei Men, and Yongfeng Guo. Positivity and convergence of the balanced implicit method for the nonlinear jump-extended CIR model. Mathematics and Computers in Simulation, 182:195–210, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303797.

Akinyemi:2021:ESG

[388] Lanre Akinyemi, Mehmet Şenol, and Olaniyi S. Iyiola. Exact solutions of the generalized multidimensional mathematical physics models via subequation method. Mathematics and Computers in Simulation, 182:211– 233, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303670.

Triki:2021:RCE

[389] Houria Triki and Thiab R. Taha. Reply to ‘Comment on: “Exact analytic solitary wave solutions for the RKL model” [Math. Comput. Simulation 80 (4) (2009) 849–854]’. Mathematics and Computers in Simulation, 182:234, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303256. See

[1, 348].

Almushaira:2021:FHO

[390] M. Almushaira, H. Bhatt, and A. M. Al-rassas. Fast high-order method for multi-dimensional space-fractional reaction–diffusion equations with general boundary conditions. Mathematics and Computers in Simulation, 182:235–258, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303943.

Miaadi:2021:IDS

[391] Foued Miaadi and Xiaodi Li. Impulsedependent settling-time for finite time stabilization of uncertain impulsive static neural networks with leakage delay and distributed delays. Mathematics and Computers in Simulation, 182:259–276, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303967.

Wiart:2021:WFS

[392] Jaspar Wiart and Elaine Wong. Walsh functions, scrambled (0, m, s)-nets, and negative covariance: Applying symbolic computation to quasi-Monte Carlo integration. Mathematics and Computers in Simulation, 182:277– 295, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303815.

Magagula:2021:SAV

[393] Vusi Mpendulo Magagula and S’yanda Nkanyiso Mungwe. Stability analysis of a virulent code in a network of computers. Mathematics and Computers in Simulation, 182:296– 315, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303980.

Sharma:2021:ENT

[394] Janak Raj Sharma and Sunil Kumar. An excellent numerical technique for multiple roots. Mathematics and Computers in Simulation, 182:316–324, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030402X.

Zulkeflee:2021:AEI

[395] Siti Asyura Zulkeflee, Fakhrony Sholahudin Rohman, Suhairi Abd Sata, and Norashid Aziz. Autoregressive exogenous input modelling for lipase catalysed esterification process. Mathematics and Computers in Simulation, 182:325–339, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420304006.

Rocha:2021:ITS

[396] J. Leonel Rocha and S. Carvalho. Information theory, synchronization and topological order in complete dynamical networks of discontinuous maps. Mathematics and Computers in Simulation, 182:340– 352, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304018.

Wu:2021:LAR

[397] Chengfeng Wu, Xin Liu, and Annan Li. A loss-averse retailer–supplier supply chain model under trade credit in a supplier-Stackelberg game. Mathematics and Computers in Simulation, 182:353–365, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303803.

Trofimov:2021:CLF

[398] Vyacheslav A. Trofimov, Svetlana Stepanenko, and Alexander Razgulin. Conservation laws of femtosecond pulse propagation described by generalized nonlinear Schrödinger equation with cubic nonlinearity. Mathematics and Computers in Simulation, 182:366–396, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420304055.

Khalsaraei:2021:PES

[399] Mohammad Mehdizadeh Khalsaraei, Ali Shokri, Higinio Ramos, and Shahin Heydari. A positive and elementary stable nonstandard explicit scheme for a mathematical model of the influenza disease. Mathematics and Computers in Simulation, 182:397– 410, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304079.

Nasir:2021:DIT

[400] Nadirah Mohd Nasir, Zanariah Abdul Majid, Fudziah Ismail, and Norfifah Bachok. Direct integration of the thirdorder two point and multipoint Robin type boundary value problems. Mathematics and Computers in Simulation, 182:411–427, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303992.

Luo:2021:AGM

[401] Yantao Luo, Long Zhang, Zhidong Teng, and Tingting Zheng. Analysis of a general multi-group reaction– diffusion epidemic model with nonlinear incidence and temporary acquired immunity. Mathematics and Computers in Simulation, 182:428– 455, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420303955.

Yang:2021:EPT

[402] Huan Yang, Yuanshun Tan, Jin Yang, and Zijian Liu. Extinction and persistence of a tumor-immune model with white noise and pulsed comprehensive therapy. Mathematics and Computers in Simulation, 182: 456–470, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304080.

Xu:2021:FOB

[403] Changjin Xu, Zixin Liu, Maoxin Liao, Peiluan Li, Qimei Xiao, and Shuai Yuan. Fractional-order bidirectional associate memory (BAM) neural networks with multiple delays: the case of Hopf bifurcation. Mathematics and Computers in Simulation, 182: 471–494, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420304547.

Lorin:2021:NSF

[404] Emmanuel Lorin and Simon Tian. A numerical study of fractional linear algebraic systems. Mathematics and Computers in Simulation, 182:495–513, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420304043.

Jena:2021:SEM

[405] Rajarama Mohan Jena, Snehashish Chakraverty, and Dumitru Baleanu. SIR epidemic model of childhood diseases through fractional operators with Mittag-Leffler and exponential kernels. Mathematics and Computers in Simulation, 182:514– 534, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420304110.

Wang:2021:TFP

[406] Yihong Wang. Tailored finite point method for the approximation of diffusion operators with non-symmetric diffusion tensor. Mathematics and Computers in Simulation, 182:535– 554, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420304146.

El-Dib:2021:HPM

[407] Yusry O. El-Dib. Homotopy perturbation method with rank upgrading technique for the superior nonlinear oscillation. Mathematics and Computers in Simulation, 182:555– 565, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304122.

Silambarasan:2021:LSW

[408] Rathinavel Silambarasan, Haci Mehmet Baskonus, R. Vijay Anand, M. Dinakaran, Balamurugan Balusamy, and Wei Gao. Longitudinal strain waves propagating in an infinitely long cylindrical rod composed of generally incompressible materials and its Jacobi elliptic function solutions. Mathematics and Computers in Simulation, 182:566–602, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420304031.

Lyu:2021:ONU

[409] Jisang Lyu, Eunchae Park, Sangkwon Kim, Wonjin Lee, Chaeyoung Lee, Sungha Yoon, Jintae Park, and Junseok Kim. Optimal non-uniform finite difference grids for the Black– Scholes equations. Mathematics and Computers in Simulation, 182:690– 704, April 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304596.

Chen:2021:ICE

[410] Zhijie Chen, Runze Xu, and Zhanwen Yang. IMACS calendar of events. Mathematics and Computers in Simulation, 182:914–917, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475420304882.

Anonymous:2021:EBd

[411] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 182:ifc, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420304821.

Zamboni:2021:SIE

[412] Walter Zamboni, Giovanni Petrone, Maria Carmela Di Piazza, Éric Monmasson, and Benoît Robyns. Special issue ELECTRIMACS 2019 ENERGY Modelling and computational simulation for control and diagnosis in renewable energy systems, energy storage, innovative devices and materials. Mathematics and Computers in Simulation, 183:1–4, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302676.

Barcellona:2021:IET

[413] Simone Barcellona and Luigi Piegari. Integrated electro-thermal model for pouch lithium ion batteries. Mathematics and Computers in Simulation, 183:5–19, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420300847.

Juston:2021:DDS

[414] Maxime Juston, Nicolas Damay, Christophe Forgez, Guy Friedrich, Stephane Vivier, Karrick Mergo Mbeya, and Bogdan Vulturescu. Direct determination of a single battery internal resistances distribution using a heterogeneous model. Mathematics and Computers in Simulation, 183:20–33, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300227.

Mbeya:2021:LMD

[415] Karrick Mergo Mbeya, Nicolas Damay, Guy Friedrich, Christophe Forgez, and Maxime Juston. Off-line method to determine the electrode balancing of Li-ion batteries. Mathematics and Computers in Simulation, 183: 34–47, May 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420300525.

Luna:2021:EDM

[416] M. Luna, M. C. Di Piazza, G. La Tona, A. Accetta, and M. Pucci. Exploiting dynamic modeling, parameter identification, and power electronics to implement a non-dissipative Liion battery hardware emulator. Mathematics and Computers in Simulation, 183:48–65, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301543.

Trovo:2021:VDM

[417] Andrea Trovò, Piergiorgio Alotto, Monica Giomo, Federico Moro, and Massimo Guarnieri. A validated dynamical model of a kW-class Vanadium Redox Flow Battery. Mathematics and Computers in Simulation, 183:66–77, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303702.

Petrone:2021:IMM

[418] Giovanni Petrone, Giovanni Spagnuolo, Walter Zamboni, and Raffaele Siano. An improved mathematical method for the identification of fuel cell impedance parameters based on the interval arithmetic. Mathematics and Computers in Simulation, 183:78–96, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301427.

Ayuso:2021:OPL

[419] Pablo Ayuso, Hector Beltran, Jorge Segarra-Tamarit, and Emilio Pérez. Optimized profitability of LFP and NMC Li-ion batteries in residential PV applications. Mathematics and Computers in Simulation, 183: 97–115, May 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420300501.

Wen:2021:MPP

[420] Xin Wen, Dhaker Abbes, and Bruno François. Modeling of photovoltaic power uncertainties for impact analysis on generation scheduling and cost of an urban micro grid. Mathematics and Computers in Simulation, 183:116–128, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300628.

Khodamoradi:2021:PBL

[421] Aram Khodamoradi, Guangyuan Liu, Paolo Mattavelli, Tuomas Messo, and Hossein Abedini. PRBS-based loop gain identification and output impedance shaping in DC microgrid power converters. Mathematics and Computers in Simulation, 183:129–141, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301439.

Durillon:2021:CSP

[422] Benoit Durillon, Florentin Salomez, Arnaud Davigny, Sabine Kazmierczak, Hervé Barry, Christophe Saudemont, and Benoît Robyns. Consumers’ sensitivities and preferences modelling and integration in a decentralised two levels energy supervisor. Mathematics and Computers in Simulation, 183:142–157, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300422.

Merai:2021:SDM

[423] M. Merai, M. W. Naouar, I. Slama-Belkhodja, and E. Monmasson. A systematic design methodology for DClink voltage control of single phase grid-tied PV systems. Mathematics and Computers in Simulation, 183:158–170, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030166X.

Barazi:2021:CBI

[424] Yazan Barazi, Nicolas Rouger, and Frédéric Richardeau. Comparison between i_{g} integration and v_{gs} derivation methods dedicated to fast shortcircuit 2D diagnosis for wide bandgap power transistors. Mathematics and Computers in Simulation, 183:171– 186, May 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420301713.

Spinelli:2021:SAG

[425] Giovanni Spinelli, Patrizia Lamberti, Vincenzo Tucci, Francisco Pasadas, and David Jiménez. Sensitivity analysis of a Graphene Field-Effect Transistors by means of Design of Experiments. Mathematics and Computers in Simulation, 183: 187–197, May 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420302068.

DiBenedetto:2021:DCR

[426] Luigi Di Benedetto, Gian Domenico Licciardo, and Alfredo Rubino. On the design of the channel region in 4H-SiC JBS diode through an analytical model of the potential barrier. Mathematics and Computers in Simulation, 183:198–207, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302366.

Stoyka:2021:BMW

[427] Kateryna Stoyka, Gennaro Di Mambro, Nicola Femia, Antonio Maffucci, Salvatore Ventre, and Fabio Villone. Behavioral modeling of Wireless Power Transfer System coils. Mathematics and Computers in Simulation, 183:208–220, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300057.

Popoli:2021:NMC

[428] Arturo Popoli, Andrea Cristofolini, and Leonardo Sandrolini. A numerical model for the calculation of electromagnetic interference from power lines on nonparallel underground pipelines. Mathematics and Computers in Simulation, 183:221– 233, May 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420300549.

Apicella:2021:AMP

[429] Valerio Apicella, Carmine Stefano Clemente, Daniele Davino, Damiano Leone, and Ciro Visone. Analysis and Modeling of a passive force sensor based on Villari effect. Mathematics and Computers in Simulation, 183:234–243, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420300203.

Vaccariello:2021:TMG

[430] Enrico Vaccariello, Pierluigi Leone, Flavio G. Canavero, and Igor S. Stievano. Topological modelling of gas networks for co-simulation applications in multi-energy systems. Mathematics and Computers in Simulation, 183:244–253, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475419303921.

Anonymous:2021:NId

[431] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 183:254, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420304237.

Anonymous:2021:ICEd

[432] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 183:255–258, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475420304249.

Anonymous:2021:EBe

[433] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 183:ifc, May 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542030416X.

Petrone:2021:SIE

[434] Giovanni Petrone, Walter Zamboni, Efstratios Batzelis, Patrizio Manganiello, and Ramon Blasco-Gimenez. Special issue ELECTRI-MACS 2019 ENGINEERING modelling and computational simulation for analysis and optimization in electrical power engineering. Mathematics and Computers in Simulation, 184:1–4, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302688.

Xiao:2021:MAE

[435] Qi Xiao, Paolo Mattavelli, Aram Khodamoradi, and Poh Chiang Loh. Modelling and analysis of equivalent SISO d–q impedance of gridconnected converters. Mathematics and Computers in Simulation, 184: 5–20, June 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301701.

Siangsanoh:2021:SHF

[436] A. Siangsanoh, M. Bahrami, W. Kaewmanee, R. Gavagsaz-ghoachani, M. Phattanasak, J. P. Martin, B. Nahid-Mobarakeh, M. Weber, S. Pierfederici, G. Maranzana, and S. Didierjean. Series hybrid fuel cell/ supercapacitor power source. Mathematics and Computers in Simulation, 184:21–40, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300276.

Bhadra:2021:AMS

[437] Sraddhanjoli Bhadra and Hirak Patangia. An analytical method of switching waveform design for selective harmonic elimination. Mathematics and Computers in Simulation, 184: 41–54, June 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420301440.

Patin:2021:SIP

[438] Nicolas Patin, Zakaria Chmeit, Georges Salloum, and Rita Mbayed. Study of interleaved PWM strategies applied to two back-to-back threephase full bridges. Mathematics and Computers in Simulation, 184: 55–68, June 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420300458.

Buccella:2021:MPH

[439] Concettina Buccella, Maria Gabriella Cimoroni, and Carlo Cecati. Mathematical proof of a harmonic elimination procedure for multilevel inverters. Mathematics and Computers in Simulation, 184:69–81, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302317.

Hillesheim:2021:EDI

[440] Miguel Mannes Hillesheim, Marc Cousineau, Miguel Vivert, Guillaume Aulagnier, and Guillaume Gateau. Eigendecomposition of a digital iterative decentralised interleaving for multicellular converters. Mathematics and Computers in Simulation, 184:82–105, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030241X.

Pascal:2021:AMR

[441] Yoann Pascal, Mickaël Petit, Denis Labrousse, and François Costa. Analytical model of a resonator for PCBembedded power conversion. Mathematics and Computers in Simulation, 184:106–117, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302263.

Amara:2021:POH

[442] Yacine Amara, Sami Hlioui, Hamid Ben Ahmed, and Mohamed Gabsi. Pre-optimization of hybridization ratio in hybrid excitation synchronous machines using electrical circuits modelling. Mathematics and Computers in Simulation, 184:118– 136, June 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301464.

Scicluna:2021:HFI

[443] K. Scicluna, C. Spiteri Staines, and R. Raute. High frequency injection-based sensorless position estimation in permanent magnet synchronous machines. Mathematics and Computers in Simulation, 184: 137–152, June 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542030063X.

Uygun:2021:ILA

[444] Emre Uygun, Michel Hecquet, Abdelmounaïm Tounzi, Daniel Depernet, Vincent Lanfranchi, Serge Bruno, and Thierry Tollance. Influence of the load angle on magnetic radial forces and torque ripple of a low power permanent magnet synchronous machine. Mathematics and Computers in Simulation, 184:153–164, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420301798.

Vivier:2021:GPO

[445] Stephane Vivier. Graphical predetermination of optimal machine designs by iso-performance configuration modeling. Mathematics and Computers in Simulation, 184:165– 183, June 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420300641.

Zhang:2021:DML

[446] Jian Zhang, Abdelmounaim Tounzi, Abdelkader Benabou, and Yvonnick Le Menach. Detection of magnetization loss in a PMSM with Hilbert Huang transform applied to noninvasive search coil voltage. Mathematics and Computers in Simulation, 184:184–195, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300483.

Lin-Shi:2021:EIV

[447] Xuefang Lin-Shi, Paolo Massioni, and Jean-Yves Gauthier. Estimation of inverter voltage disturbances for induction machine drive using LPV observer with convex optimization. Mathematics and Computers in Simulation, 184:196–209, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301956.

BenGhorbal:2021:CST

[448] Manel Jebali Ben Ghorbal, Sonia Moussa, Jihen Arbi Ziani, and Ilhem Slama-Belkhodja. A comparison study of two DC microgrid controls for a fast and stable DC bus voltage. Mathematics and Computers in Simulation, 184:210–224, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300471.

Iannuzzi:2021:UCS

[449] Diego Iannuzzi and Pasquale Franzese. Ultrafast charging station for electrical vehicles: Dynamic modelling, design and control strategy. Mathematics and Computers in Simulation, 184:225–243, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030149X.

Almaksour:2021:CDM

[450] Khaled Almaksour, Youssef Krim, N’guessan Kouassi, Nicolas Navarro, Bruno François, Tony Letrouvé, Christophe Saudemont, Lionel Taunay, and Benoit Robyns. Comparison of dynamic models for a DC railway electrical network including an AC/DC bi-directional power station. Mathematics and Computers in Simulation, 184:244–266, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301865.

Guarino:2021:SBE

[451] Antonio Guarino, Éric Monmasson, and Giovanni Spagnuolo. SoC-based embedded real-time simulation of mismatched photovoltaic strings. Mathematics and Computers in Simulation, 184:267–281, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301452.

Mansoor:2021:CES

[452] Muhammad Mansoor, Francesco Grimaccia, Sonia Leva, and Marco Mussetta. Comparison of echo state network and feed-forward neural networks in electrical load forecasting for demand response programs. Mathematics and Computers in Simulation, 184:282–293, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302391.

DiPiazza:2021:ANN

[453] A. Di Piazza, M. C. Di Piazza, G. La Tona, and M. Luna. An artificial neural network-based forecasting model of energy-related time series for electrical grid management. Mathematics and Computers in Simulation, 184:294–305, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301695.

Segarra-Tamarit:2021:DLB

[454] Jorge Segarra-Tamarit, Emilio Pérez, Eric Moya, Pablo Ayuso, and Hector Beltran. Deep learning-based forecasting of aggregated CSP production. Mathematics and Computers in Simulation, 184:306–318, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847542030046X.

Anonymous:2021:NIe

[455] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 184:319, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420304420.

Anonymous:2021:ICEe

[456] Anonymous. IMACS Calendar of Events. Mathematics and Computers in Simulation, 184:320–323, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475420304432.

Anonymous:2021:EBf

[457] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 184:ifc, June 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420304353.

Prabakaran:2021:FBE

[458] G. Prabakaran, D. Vaithiyanathan, and Madhavi Ganesan. FPGA based effective agriculture productivity prediction system using fuzzy support vector machine. Mathematics and Computers in Simulation, 185:1–16, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420304663.

Anonymous:2021:PJb

[459] Anonymous. Pages 1–818 (July 2021). Mathematics and Computers in Simulation, 185:1–818, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Yan:2021:NIA

[460] Zheping Yan, Jinzhong Zhang, Jia Zeng, and Jialing Tang. Natureinspired approach: an enhanced whale optimization algorithm for global optimization. Mathematics and Computers in Simulation, 185: 17–46, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420304638.

Khan:2021:NDS

[461] M. Ijaz Khan and Faris Alzahrani. Nonlinear dissipative slip flow of Jeffrey nanomaterial towards a curved surface with entropy generation and activation energy. Mathematics and Computers in Simulation, 185:47– 61, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304614.

Alquran:2021:DWS

[462] Marwan Alquran, Feras Yousef, Farah Alquran, Tukur A. Sulaiman, and Abdullahi Yusuf. Dual-wave solutions for the quadratic–cubic conformable-Caputo time-fractional Klein–Fock– Gordon equation. Mathematics and Computers in Simulation, 185:62– 76, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304699.

Zguaid:2021:ROL

[463] Khalid Zguaid, Fatima Zahrae El Alaoui, and Ali Boutoulout. Regional observability for linear time fractional systems. Mathematics and Computers in Simulation, 185: 77–87, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304687.

Hong:2021:RCO

[464] Ming Hong and Heyong Wang. Research on customer opinion summarization using topic mining and deep neural network. Mathematics and Computers in Simulation, 185: 88–114, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542030464X.

Bakhta:2021:MOF

[465] Athmane Bakhta and Julien Vidal. Modeling and optimization of the fabrication process of thin-film solar cells by multi-source physical vapor deposition. Mathematics and Computers in Simulation, 185:115– 133, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304717.

Zou:2021:FEA

[466] Guang an Zou, Xue Wang, and Tony W. H. Sheu. Finite element analysis of a new phase field model with p-Laplacian operator. Mathematics and Computers in Simulation, 185: 134–152, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304924.

Ku:2021:SEB

[467] Cheng-Yu Ku, Jing-En Xiao, Chih-Yu Liu, and Der-Guey Lin. On solving elliptic boundary value problems using a meshless method with radial polynomials. Mathematics and Computers in Simulation, 185: 153–173, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304675.

Belhachmi:2021:MMN

[468] Zakaria Belhachmi, Zoubida Mghazli, and Salih Ouchtout. Mathematical modelling and numerical approximation of a leachate flow in the anaerobic biodegradation of waste in a landfill. Mathematics and Computers in Simulation, 185: 174–193, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304754.

Hamiche:2021:NSR

[469] Hamid Hamiche, Hocine Takhi, Manal Messadi, Karim Kemih, Ouerdia Megherbi, and Maamar Bettayeb. New synchronization results for a class of nonlinear discrete-time chaotic systems based on synergetic observer and their implementation. Mathematics and Computers in Simulation, 185:194–217, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420304778.

Blanco-Cocom:2021:RPE

[470] Luis Blanco-Cocom, Salvador Botello-Rionda, L. C. Ordoñez, and S. Ivvan Valdez. Robust parameter estimation of a PEMFC via optimization based on probabilistic model building. Mathematics and Computers in Simulation, 185:218–237, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420304766.

Fu:2021:HOS

[471] Yayun Fu, Dongdong Hu, and Yushun Wang. High-order structure-preserving algorithms for the multi-dimensional fractional nonlinear Schrödinger equation based on the SAV approach. Mathematics and Computers in Simulation, 185:238–255, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420304808.

Jator:2021:BHT

[472] S. N. Jator, D. C. Mayo, and M. T. Omojola. Block Hybrid Third Derivative Nyström type method for Bratu’s equation. Mathematics and Computers in Simulation, 185: 256–271, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420304936.

Saraei:2021:CHF

[473] Sina Hassanzadeh Saraei, Ali Chamkha, and Abdolrahman Dadvand. Controlling the hydrodynamic forces on a square cylinder in a channel via an upstream porous plate. Mathematics and Computers in Simulation, 185:272–288, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420304729.

Zhao:2021:RZN

[474] Lv Zhao, Jie Jin, and Jianqiang Gong. Robust zeroing neural network for fixed-time kinematic control of wheeled mobile robot in noisepolluted environment. Mathematics and Computers in Simulation, 185: 289–307, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304961.

Zhao:2021:CFE

[475] Jun Zhao, SungBum Kim, and Hyoung-Moon Kim. Closed-form estimators and bias-corrected estimators for the Nakagami distribution. Mathematics and Computers in Simulation, 185:308–324, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420304912.

Madhu:2021:OVH

[476] Kalyanasundaram Madhu and Mo’tassem Al-arydah. Optimal vaccine for human papillomavirus and age-difference between partners. Mathematics and Computers in Simulation, 185:325– 346, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000033.

Qu:2021:BAS

[477] Wenzhen Qu, Linlin Sun, and Po-Wei Li. Bending analysis of simply supported and clamped thin elastic plates by using a modified version of the LMFS. Mathematics and Computers in Simulation, 185: 347–357, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420304997.

Serghini:2021:CBS

[478] A. Serghini. A construction of a bivariate C^{2} spline approximant with minimal degree on arbitrary triangulation. Mathematics and Computers in Simulation, 185:358– 371, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421000045.

Liu:2021:PPI

[479] Chengzhi Liu, Zhongyun Liu, and Xuli Han. Preconditioned progressive iterative approximation for tensor product Bézier patches. Mathematics and Computers in Simulation, 185:372–383, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000021.

Ghosh:2021:IBS

[480] Sudeshna Ghosh. Immersed boundary simulations of fluid shear-induced deformation of a cantilever beam. Mathematics and Computers in Simulation, 185:384–402, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542100001X.

Sakthivel:2021:FTR

[481] R. Sakthivel, V. T. Suveetha, V. Nithya, and R. Sakthivel. Finite-time reliable filtering for Takagi–Sugeno fuzzy semi-Markovian jump systems. Mathematics and Computers in Simulation, 185:403–418, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420305048.

Liu:2021:SSB

[482] Chein-Shan Liu and Botong Li. Solving a singular beam equation by the method of energy boundary functions. Mathematics and Computers in Simulation, 185:419– 435, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000057.

Li:2021:TGM

[483] Qingfeng Li, Yanping Chen, Yunqing Huang, and Yang Wang. Twogrid methods for nonlinear time fractional diffusion equations by L1-Galerkin FEM. Mathematics and Computers in Simulation, 185:436– 451, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420305012.

Atef:2021:TTS

[484] Mohammed Atef and Shokry I. Nada. On three types of soft fuzzy coverings based rough sets. Mathematics and Computers in Simulation, 185:452–467, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030478X.

Ali:2021:FTS

[485] M. Syed Ali, G. Narayanan, Sumit Saroha, Bandana Priya, and Ganesh Kumar Thakur. Finite-time stability analysis of fractional-order memristive fuzzy cellular neural networks with time delay and leakage term. Mathematics and Computers in Simulation, 185:468–485, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000082.

Kaushik:2021:MGM

[486] Aditya Kaushik, Vijayant Kumar, Manju Sharma, and Nitika Sharma. A modified graded mesh and higher order finite element method for singularly perturbed reaction–diffusion problems. Mathematics and Computers in Simulation, 185:486–496, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475421000069.

Khan:2021:MSM

[487] M. Ijaz Khan, Sumaira Qayyum, Shahid Farooq, Yu-Ming Chu, and Seifedine Kadry. Modeling and simulation of micro-rotation and spin gradient viscosity for ferromagnetic hybrid (manganese zinc ferrite, nickle zinc ferrite) nanofluids. Mathematics and Computers in Simulation, 185:497–509, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000070.

Yang:2021:CCF

[488] Xiaojia Yang, Yongbin Ge, and Bin Lan. A class of compact finite difference schemes for solving the 2D and 3D Burgers’ equations. Mathematics and Computers in Simulation, 185:510–534, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000100.

Kaslik:2021:GSA

[489] Eva Kaslik, Mihaela Neamţu, and Loredana Flavia Vesa. Global stability analysis of an unemployment model with distributed delay. Mathematics and Computers in Simulation, 185:535–546, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000112.

Sadabad:2021:ENM

[490] Mahnaz Kashfi Sadabad and Aliasghar Jodayree Akbarfam. An efficient numerical method for estimating eigenvalues and eigenfunctions of fractional Sturm–Liouville problems. Mathematics and Computers in Simulation, 185:547–569, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000094.

Pulecio-Montoya:2021:SDD

[491] Ana María Pulecio-Montoya, Luis Eduardo López-Montenegro, and Jhon Cerón-Caicedo. Sliding Dynamic in a 3-dimensional epidemiological system. Mathematics and Computers in Simulation, 185:570– 582, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542100032X.

Megahed:2021:MMF

[492] Ahmed M. Megahed, M. Gnaneswara Reddy, and W. Abbas. Modeling of MHD fluid flow over an unsteady stretching sheet with thermal radiation, variable fluid properties and heat flux. Mathematics and Computers in Simulation, 185: 583–593, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421000124.

Xu:2021:QSA

[493] Liguang Xu, Xiaoyan Chu, and Hongxiao Hu. Quasi-synchronization analysis for fractional-order delayed complex dynamical networks. Mathematics and Computers in Simulation, 185:594–613, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000331.

Stanimirovic:2021:SVP

[494] Predrag Stanimirović, Dimitris Gerontitis, Panagiotis Tzekis, Ratikanta Behera, and Jajati Keshari Sahoo. Simulation of Varying Parameter Recurrent Neural Network with application to matrix inversion. Mathematics and Computers in Simulation, 185:614–628, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000355.

Tai:2021:NFF

[495] Weipeng Tai, Dandan Zuo, Zuxing Xuan, Jianping Zhou, and Zhen Wang. Non-fragile L_{2}–L_{∞} filtering for a class of switched neural networks. Mathematics and Computers in Simulation, 185:629–645, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000318.

Jafari:2021:FVA

[496] Nasrin Jafari and Mojtaba Azhari. Free vibration analysis of viscoelastic plates with simultaneous calculation of natural frequency and viscous damping. Mathematics and Computers in Simulation, 185:646– 659, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000367.

Zhu:2021:EEC

[497] Jianxin Zhu, Haocai Huang, Honghong Lin, Xiaojie Lei, and Ruizhong Zheng. Efficient estimation of chlorophyll a concentration in artificial upwelling. Mathematics and Computers in Simulation, 185:660– 675, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420304791.

Atas:2021:BAF

[498] Fatma Ataş, Ali Demirci, and Cihangir Özemir. Bifurcation analysis of Friedkin–Johnsen and Hegselmann– Krause models with a nonlinear interaction potential. Mathematics and Computers in Simulation, 185: 676–686, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421000276.

Yarsky:2021:UGA

[499] P. Yarsky. Using a genetic algorithm to fit parameters of a COVID-19 SEIR model for US states. Mathematics and Computers in Simulation, 185:687–695, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000392.

Wang:2021:TES

[500] Nan Wang and Dongyang Shi. Two efficient spectral methods for the nonlinear fractional wave equation in unbounded domain. Mathematics and Computers in Simulation, 185: 696–718, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421000380.

Calcina:2021:PES

[501] Sabrina S. Calcina and Marcio Gameiro. Parameter estimation in systems exhibiting spatially complex solutions via persistent homology and machine learning. Mathematics and Computers in Simulation, 185:719–732, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000306.

Yadav:2021:ASO

[502] Swati Yadav and Pratima Rai. An almost second order hybrid scheme for the numerical solution of singularly perturbed parabolic turning point problem with interior layer. Mathematics and Computers in Simulation, 185:733–753, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000343.

Zhu:2021:TDB

[503] Hegui Zhu, Lewen Dai, Yating Liu, and Lijun Wu. A three-dimensional bit-level image encryption algorithm with Rubik’s cube method. Mathematics and Computers in Simulation, 185:754–770, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000483.

Mirchi:2021:NIM

[504] Hamid Mirchi and Davod Khojasteh Salkuyeh. A new iterative method for solving the systems arisen from finite element discretization of a time-harmonic parabolic optimal control problems. Mathematics and Computers in Simulation, 185:771– 782, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000525.

Hashemi:2021:ATH

[505] Seyyed Amjad Samareh Hashemi and Habibollah Saeedi. ADM–TF hybrid method for nonlinear Itô-Volterra integral equations. Mathematics and Computers in Simulation, 185:783– 798, July 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000422.

Sabir:2021:NEC

[506] Zulqurnain Sabir, Muhammad Asif Zahoor Raja, Chaudry Masood Khalique, and Canan Unlu. Neuro-evolution computing for nonlinear multi-singular system of third order Emden–Fowler equation. Mathematics and Computers in Simulation, 185:799–812, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475421000434.

Anonymous:2021:NIf

[507] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 185:813, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000744.

Anonymous:2021:ICEf

[508] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 185:814–817, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475421000756.

Anonymous:2021:EBg

[509] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 185:ifc, July 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000690.

Barrera:2021:MSI

[510] Domingo Barrera, Sara Remogna, and Driss Sbibih. MATCOM special issue MACMAS 2019: First International Conference on Mathematical and Computational Modelling, Approximation and Simulation: New trends, recent developments and applications in environment and natural resources. Mathematics and Computers in Simulation, 186:1–2, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420304742.

Okayama:2021:MSQ

[511] Tomoaki Okayama and Shu Hanada. A modified Stenger’s quadrature formula for infinite integrals of unilateral rapidly decreasing functions and its theoretical error bound. Mathematics and Computers in Simulation, 186:3–18, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030080X.

Ferri-Garcia:2021:EML

[512] Ramón Ferri-García, Luis Castro-Martín, and María del Mar Rueda. Evaluating Machine Learning methods for estimation in online surveys with superpopulation modeling. Mathematics and Computers in Simulation, 186:19–28, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420300793.

Dia:2021:MSP

[513] Ben Mansour Dia, Bilal Saad, and Mazen Saad. Modeling and simulation of partially miscible two-phase flow with kinetics mass transfer. Mathematics and Computers in Simulation, 186:29–40, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301506.

Aguilera:2021:HPB

[514] Ana M. Aguilera, Christian Acal, M. Carmen Aguilera-Morillo, Francisco Jiménez-Molinos, and Juan B. Roldán. Homogeneity problem for basis expansion of functional data with applications to resistive memories. Mathematics and Computers in Simulation, 186:41–51, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301774.

Moreno:2021:BGV

[515] E. Moreno, P. González, R. Emadi, J. B. Roldán, and E. A. Michael. BCC-Grid versus SC-Grid in the modeling of a sheet of graphene as a surface boundary condition in the context of ADE-FDTD. Mathematics and Computers in Simulation, 186: 52–61, August 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301786.

Rueda:2021:NET

[516] M. Rueda, B. Cobo, and P. F. Perri. New estimation techniques for ordinal sensitive variables. Mathematics and Computers in Simulation, 186:62–70, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302196.

Ruiz-Castro:2021:LPT

[517] Juan E. Ruiz-Castro, Christian Acal, Ana M. Aguilera, M. Carmen Aguilera-Morillo, and Juan B. Roldán. Linear-Phase-Type probability modelling of functional PCA with applications to resistive memories. Mathematics and Computers in Simulation, 186:71– 79, August 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302354.

Bouhiri:2021:QIO

[518] S. Bouhiri, A. Lamnii, M. Lamnii, and A. Zidna. Quasi-interpolant operators in Bernstein basis. Mathematics and Computers in Simulation, 186:80–90, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302299.

Michel:2021:NDH

[519] D. Michel and A. Zidna. A new deterministic heuristic knots placement for B-spline approximation. Mathematics and Computers in Simulation, 186:91–102, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302482.

Jimenez-Gamero:2021:TGD

[520] M. D. Jiménez-Gamero and M. V. Alba-Fernández. A test for the geometric distribution based on linear regression of order statistics. Mathematics and Computers in Simulation, 186:103–123, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302901.

Costabile:2021:LBC

[521] F. A. Costabile, M. I. Gualtieri, and A. Napoli. Lidstone-based collocation splines for odd-order BVPs. Mathematics and Computers in Simulation, 186:124–135, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303104.

Rueda:2021:TNE

[522] M. Rueda, S. Martínez, and M. Illescas. Treating nonresponse in the estimation of the distribution function. Mathematics and Computers in Simulation, 186:136–144, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303244.

Paquet:2021:CDD

[523] Luc Paquet and Réda Korikache. The complete discretization of the dual mixed method for the heat diffusion equation in a polygonal domain. Mathematics and Computers in Simulation, 186:145–160, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303360.

Anonymous:2021:NIg

[524] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 186:161, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000215.

Anonymous:2021:ICEg

[525] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 186:162–165, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475421000227.

Anonymous:2021:EBh

[526] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 186:ifc, August 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000148.

Iqbal:2021:GAQ

[527] Azhar Iqbal, Nur Nadiah Abd Hamid, Ahmad Izani Md. Ismail, and Muhammad Abbas. Galerkin approximation with quintic B-spline as basis and weight functions for solving second order coupled nonlinear Schrödinger equations. Mathematics and Computers in Simulation, 187:1–16, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000513.

Anonymous:2021:PS

[528] Anonymous. Pages 1–736 (September 2021). Mathematics and Computers in Simulation, 187:1–736, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Singh:2021:BSS

[529] Saroja Kumar Singh, Sarat Kumar Acharya, Frederico R. B. Cruz, and Roberto C. Quinino. Bayesian sample size determination in a single-server deterministic queueing system. Mathematics and Computers in Simulation, 187:17–29, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000495.

Roopnarain:2021:ADO

[530] Ryan Roopnarain and S. Roy Choudhury. Amplitude death, oscillation death, and periodic regimes in dynamically coupled Landau–Stuart oscillators with and without distributed delay. Mathematics and Computers in Simulation, 187:30– 50, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000458.

Deng:2021:NRD

[531] Yunke Deng, Chuangxia Huang, and Jinde Cao. New results on dynamics of neutral type HCNNs with proportional delays. Mathematics and Computers in Simulation, 187: 51–59, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000409.

Jafarimoghaddam:2021:UBP

[532] A. Jafarimoghaddam, N. C. Roşca, A. V. Roşca, and I. Pop. The universal Blasius problem: New results by Duan–Rach Adomian Decomposition Method with Jafarimoghaddam contraction mapping theorem and numerical solutions. Mathematics and Computers in Simulation, 187: 60–76, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000537.

Zarepour:2021:NSA

[533] Mazyar Zarepour and Ghasem Barid Loghmani. Numerical solution of arbitrary-order linear partial differential equations using an optimal control technique. Mathematics and Computers in Simulation, 187:77– 96, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421000471.

Megahed:2021:IHT

[534] Ahmed M. Megahed. Improvement of heat transfer mechanism through a Maxwell fluid flow over a stretching sheet embedded in a porous medium and convectively heated. Mathematics and Computers in Simulation, 187: 97–109, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000574.

Pu:2021:FHO

[535] Zhe Pu, Maohua Ran, and Hong Luo. Fast and high-order difference schemes for the fourth-order fractional subdiffusion equations with spatially variable coefficient under the first Dirichlet boundary conditions. Mathematics and Computers in Simulation, 187: 110–133, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000562.

Abdelhamid:2021:RHT

[536] Talaat Abdelhamid, Ammar H. Elsheikh, Olatunji Mumini Omisore, N. A. Saeed, T. Muthuramalingam, Ronglinag Chen, and Md. Mahbub Alam. Reconstruction of the heat transfer coefficients and heat fluxes in heat conduction problems. Mathematics and Computers in Simulation, 187: 134–154, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000501.

Umavathi:2021:CSP

[537] Jawali C. Umavathi and Ali J. Chamkha. Convective stability of a permeable nanofluid inside a horizontal conduit: Fast chemical reactions. Mathematics and Computers in Simulation, 187:155–170, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000550.

Adamu:2021:USN

[538] Elias M. Adamu, Kailash C. Patidar, and Andriamihaja Ramanantoanina. An unconditionally stable nonstandard finite difference method to solve a mathematical model describing Visceral Leishmaniasis. Mathematics and Computers in Simulation, 187: 171–190, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100046X.

Sun:2021:SCM

[539] Shangqu Sun, Peng He, Gang Wang, Weiteng Li, Hongbo Wang, Diyang Chen, and Fei Xu. Shape characterization methods of irregular cavity using Fourier analysis in tunnel. Mathematics and Computers in Simulation, 187:191–214, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000549.

Usurelu:2021:BPP

[540] Gabriela Ioana Usurelu and Teodor Turcanu. Best proximity points of (EP)-operators with qualitative analysis and simulation. Mathematics and Computers in Simulation, 187: 215–230, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000616.

Ellabib:2021:MAS

[541] Abdellatif Ellabib, Abdeljalil Nachaoui, and Abdessamad Ousaadane. Mathematical analysis and simulation of fixed point formulation of Cauchy problem in linear elasticity. Mathematics and Computers in Simulation, 187: 231–247, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000598.

Hosseini:2021:ECM

[542] Kamyar Hosseini, Mousa Ilie, Mohammad Mirzazadeh, Abdullahi Yusuf, Tukur Abdulkadir Sulaiman, Dumitru Baleanu, and Soheil Salahshour. An effective computational method to deal with a time-fractional nonlinear water wave equation in the Caputo sense. Mathematics and Computers in Simulation, 187:248–260, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000604.

Shokri:2021:SNS

[543] Ali Shokri and Erfan Bahmani. A study of nonlinear systems arising in the physics of liquid crystals, using MLPG and DMLPG methods. Mathematics and Computers in Simulation, 187:261–281, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100063X.

Liu:2021:TNF

[544] Chein-Shan Liu, Essam R. El-Zahar, and Chih-Wen Chang. Three novel fifth-order iterative schemes for solving nonlinear equations. Mathematics and Computers in Simulation, 187: 282–293, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000665.

Wang:2021:RBF

[545] Wenlong Wang, Xiao Lin, and Chunrui Zhang. Resonant bifurcation of feedforward chains and application in image contrast enhancement. Mathematics and Computers in Simulation, 187: 294–307, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000781.

Li:2021:PSP

[546] Shangzhi Li and Shangjiang Guo. Permanence of a stochastic prey– predator model with a general functional response. Mathematics and Computers in Simulation, 187:308– 336, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421000641.

Xiang:2021:SPD

[547] Nan Xiang, Qidong Wu, and Aying Wan. Spatiotemporal patterns of a diffusive plant–herbivore model with toxin-determined functional responses: Multiple bifurcations. Mathematics and Computers in Simulation, 187: 337–356, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000860.

Li:2021:LLM

[548] Changpin Li, Dongxia Li, and Zhen Wang. L1/LDG method for the generalized time-fractional Burgers equation. Mathematics and Computers in Simulation, 187:357–378, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542100080X.

Li:2021:RGO

[549] Wang Li, Lingzhi Zhao, Hongjun Shi, Donghua Zhao, and Yongzheng Sun. Realizing generalized outer synchronization of complex dynamical networks with stochastically adaptive coupling. Mathematics and Computers in Simulation, 187:379– 390, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000653.

Douaifia:2021:NIB

[550] Redouane Douaifia, Samir Bendoukha, and Salem Abdelmalek. A Newton interpolation based predictor–corrector numerical method for fractional differential equations with an activator– inhibitor case study. Mathematics and Computers in Simulation, 187: 391–413, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000847.

Karamollahi:2021:ASN

[551] Nasibeh Karamollahi, Mohammad Heydari, and Ghasem Barid Loghmani. Approximate solution of nonlinear Fredholm integral equations of the second kind using a class of Hermite interpolation polynomials. Mathematics and Computers in Simulation, 187:414–432, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000902.

Jena:2021:HIS

[552] Mahendra Kumar Jena. A Hermite interpolatory subdivision scheme constructed from quadratic rational Bernstein–Bezier spline. Mathematics and Computers in Simulation, 187:

433–448, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000938.

Arenas:2021:EDA

[553] Zochil González Arenas, Juan Carlos Jimenez, Li-Vang Lozada-Chang, and Roberto Santana. Estimation of distribution algorithms for the computation of innovation estimators of diffusion processes. Mathematics and Computers in Simulation, 187: 449–467, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000926.

Mandal:2021:MSC

[554] Dibyendu Sekhar Mandal, Abdennasser Chekroun, Sudip Samanta, and Joydev Chattopadhyay. A mathematical study of a crop-pest–natural enemy model with Z-type control. Mathematics and Computers in Simulation, 187:468–488, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000896.

Stefani:2021:CCN

[555] Silvana Stefani, Marcel Ausloos, Concepción González-Concepción, Adeyemi Sonubi, Ma Candelaria Gil-Fariña, Celina Pestano-Gabino, and Enrico Moretto. Competing or collaborating, with no symmetrical behaviour: Leadership opportunities and winning strategies under stability. Mathematics and Computers in Simulation, 187:

489–504, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000884.

Ma:2021:NDK

[556] Yu-Lan Ma, Abdul-Majid Wazwaz, and Bang-Qing Li. A new (3 + 1)- dimensional Kadomtsev–Petviashvili equation and its integrability, multiplesolitons, breathers and lump waves. Mathematics and Computers in Simulation, 187:505–519, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475421000872.

Halilu:2021:SRC

[557] Abubakar Sani Halilu, Arunava Majumder, Mohammed Yusuf Waziri, and Kabiru Ahmed. Signal recovery with convex constrained nonlinear monotone equations through conjugate gradient hybrid approach. Mathematics and Computers in Simulation, 187: 520–539, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000951.

Akahori:2021:EWE

[558] Jirô Akahori, Masahiro Kinuya, Takashi Sawai, and Tomooki Yuasa. An efficient weak Euler–Maruyama type approximation scheme of very high dimensional SDEs by orthogonal random variables. Mathematics and Computers in Simulation, 187: 540–565, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000859.

Asl-Najafi:2021:DCS

[559] Javad Asl-Najafi, Saeed Yaghoubi, and Fatemeh Zand. Dual-channel supply chain coordination considering targeted capacity allocation under uncertainty. Mathematics and Computers in Simulation, 187:566– 585, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542100094X.

Aouadi:2021:ESN

[560] Moncef Aouadi and Maria Inês M. Copetti. Exponential stability and numerical analysis of a thermoelastic diffusion beam with rotational inertia and second sound. Mathematics and Computers in Simulation, 187: 586–613, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001014.

Liu:2021:BSF

[561] Chein-Shan Liu, Essam R. El-Zahar, and Chih-Wen Chang. A boundary shape function iterative method for solving nonlinear singular boundary value problems. Mathematics and Computers in Simulation, 187: 614–629, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001051.

Madureira:2021:NAT

[562] Rodrigo L. R. Madureira, Mauro A. Rincon, and Moncef Aouadi. Numerical analysis for a thermoelastic diffusion problem in moving boundary. Mathematics and Computers in Simulation, 187:630–655, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001075.

Rao:2021:PUN

[563] S. Chandra Sekhara Rao and Abhay Kumar Chaturvedi. Parameteruniform numerical method for a two-dimensional singularly perturbed convection–reaction–diffusion problem with interior and boundary layers. Mathematics and Computers in Simulation, 187:656–686, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000914.

Sanchez:2021:DES

[564] Allan G. S. Sánchez, C. Posadas-Castillo, and E. Garza-González. Determining efficiency of small-world algorithms: a comparative approach. Mathematics and Computers in Simulation, 187:687–699, September 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001038.

Qi:2021:MMA

[565] Haokun Qi and Xinzhu Meng. Mathematical modeling, analysis and numerical simulation of HIV: the influence of stochastic environmental fluctuations on dynamics. Mathematics and Computers in Simulation, 187: 700–719, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001026.

Cheng:2021:WLW

[566] Li Cheng, Yi Zhang, Wen-Xiu Ma, and Jian-Ya Ge. Wronskian and lump wave solutions to an extended second KP equation. Mathematics and Computers in Simulation, 187: 720–731, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000999.

Anonymous:2021:NIh

[567] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 187:732, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001427.

Anonymous:2021:EBi

[568] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 187:733–736, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001439.

Anonymous:2021:EBj

[569] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 187:ifc, September 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001373.

Arancibia–Ibarra:2021:DLG

[570] Claudio Arancibia-Ibarra and José Flores. Dynamics of a Leslie–Gower predator–prey model with Holling type II functional response, Allee effect and a generalist predator. Mathematics and Computers in Simulation, 188:1–22, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001105.

Anonymous:2021:PO

[571] Anonymous. Pages 1–614 (October 2021). Mathematics and Computers in Simulation, 188:1–614, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Xie:2021:TDG

[572] Yingkang Xie and Zhen Wang. Transmission dynamics, global stability and control strategies of a modified SIS epidemic model on complex networks with an infective medium. Mathematics and Computers in Simulation, 188:23–34, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100104X.

Hu:2021:DPF

[573] Dongdong Hu, Wenjun Cai, Zhuangzhi Xu, Yonghui Bo, and Yushun Wang. Dissipation-preserving Fourier pseudospectral method for the space fractional nonlinear sine–Gordon equation with damping. Mathematics and Computers in Simulation, 188: 35–59, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001099.

Wang:2021:MFE

[574] Jinfeng Wang, Baoli Yin, Yang Liu, Hong Li, and Zhichao Fang. Mixed finite element algorithm for a nonlinear time fractional wave model. Mathematics and Computers in Simulation, 188:60–76, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001130.

Belhaj:2021:FMS

[575] Skander Belhaj, Fahd Hcini, and Maher Moakher. A fast method for solving quasi-pentadiagonal Toeplitz linear systems and its application to the Lax–Wendroff scheme. Mathematics and Computers in Simulation, 188:77–86, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001063.

Sabir:2021:IIN

[576] Zulqurnain Sabir, Muhammad Asif Zahoor Raja, Hafiz Abdul Wahab, Gilder Cieza Altamirano, Yu-Dong Zhang, and Dac-Nhuong Le. Integrated intelligence of neuro-evolution with sequential quadratic programming for second-order Lane–Emden pantograph models. Mathematics and Computers in Simulation, 188: 87–101, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001117.

Iqbal:2021:PFI

[577] Naveed Iqbal, Ranchao Wu, and Wael W. Mohammed. Pattern formation induced by fractional crossdiffusion in a 3-species food chain model with harvesting. Mathematics and Computers in Simulation, 188: 102–119, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001166.

Labbadi:2021:IAF

[578] Moussa Labbadi and Hassan El Moussaoui. An improved adaptive fractional-order fast integral terminal sliding mode control for distributed quadrotor. Mathematics and Computers in Simulation, 188: 120–134, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001142.

Rahman:2021:CCS

[579] M. M. Rahman. Compromising with corrector step of SIM-PLE Algorithm. Mathematics and Computers in Simulation, 188:135– 163, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542100118X.

Dubey:2021:SSC

[580] Balram Dubey, Sajan, and Ankit Kumar. Stability switching and chaos in a multiple delayed prey– predator model with fear effect and anti-predator behavior. Mathematics and Computers in Simulation, 188: 164–192, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001129.

Uzunca:2021:SPR

[581] Murat Uzunca, Bülent Karasözen, and Süleyman Yıldız. Structurepreserving reduced-order modeling of Korteweg–de Vries equation. Mathematics and Computers in Simulation, 188:193–211, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001178.

Guan:2021:LSH

[582] Xiangyu Guan, Naiqiu Shu, and Hui Peng. LBM simulation of heat transfer processes inside GIS capsule filled with different insulation gas mixtures. Mathematics and Computers in Simulation, 188:212–225, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001087.

Bohaienko:2021:ASS

[583] V. O. Bohaienko. Accuracy and speed of splitting methods for threedimensional space–time fractional diffusion equation with ψ-Caputo derivatives. Mathematics and Computers in Simulation, 188:226–240, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001221.

Umar:2021:NSI

[584] Muhammad Umar, Zulqurnain Sabir, Muhammad Asif Zahoor Raja, J. F. Gómez Aguilar, Fazli Amin, and Muhammad Shoaib. Neuro-swarm intelligent computing paradigm for nonlinear HIV infection model with CD4+ T-cells. Mathematics and Computers in Simulation, 188:241– 253, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001269.

daCruz:2021:PPS

[585] Artur M. C. Brito da Cruz and Helena Sofia Rodrigues. Personal protective strategies for dengue disease: Simulations in two coexisting virus serotypes scenarios. Mathematics and Computers in Simulation, 188: 254–267, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001208.

Zhu:2021:IBE

[586] Linhe Zhu, Wenshan Liu, and Zhengdi Zhang. Interplay between epidemic and information spreading on multiplex networks. Mathematics and Computers in Simulation, 188:268– 279, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542100135X.

Wei:2021:LDG

[587] Leilei Wei and Wenbo Li. Local discontinuous Galerkin approximations to variable-order time-fractional diffusion model based on the Caputo– Fabrizio fractional derivative. Mathematics and Computers in Simulation, 188:280–290, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001191.

Chen:2021:DMM

[588] Chengcheng Chen, Xianchang Wang, Helong Yu, Mingjing Wang, and Huiling Chen. Dealing with multi-modality using synthesis of moth-flame optimizer with sine cosine mechanisms. Mathematics and Computers in Simulation, 188:291–318, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001245.

Zhang:2021:FTF

[589] Wanli Zhang, Xinsong Yang, Shiju Yang, and Ahmed Alsaedi. Finitetime and fixed-time bipartite synchronization of complex networks with signed graphs. Mathematics and Computers in Simulation, 188:319– 329, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001312.

Li:2021:CHT

[590] Yong-Min Li, Shahab Naghdi Sedeh, Davood Toghraie, and As’ad Alizadeh. Computational hemodynamics and thermal analysis of laminar blood flow for different types of hypertension. Mathematics and Computers in Simulation, 188:330– 341, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001294.

Loy:2021:SPS

[591] Nadia Loy and Mattia Zanella. Structure preserving schemes for Fokker– Planck equations with nonconstant diffusion matrices. Mathematics and Computers in Simulation, 188:342– 362, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001464.

Stadtlander:2021:BRM

[592] Christian T. K.-H. Stadtländer. Book review: Modeling to Inform Infectious Disease Control, Becker Niels G.. CRC Press, Boca Raton, FL (2019), 1st ed., xi + 208 pp., US$74.95, ISBN 978-0-3673-7768-7 (paperback); US$105.00, ISBN 978-1-4987-3106-5 (hardback); US$74.95, ISBN 978-0-4291-5545-1 (eBook). Mathematics and Computers in Simulation, 188: 363–367, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001282.

Theodosiou:2021:DON

[593] T. C. Theodosiou. Derivativeorthogonal non-uniform B-spline wavelets. Mathematics and Computers in Simulation, 188:368–388, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475421001300.

Bera:2021:CME

[594] A. Bera, S. Dutta, J. C. Misra, and G. C. Shit. Computational modeling of the effect of blood flow and dual phase lag on tissue temperature during tumor treatment by magnetic hyperthermia. Mathematics and Computers in Simulation, 188: 389–403, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542100152X.

Jaworski:2021:FEM

[595] Nazariy Jaworski, Nazariy Andrushchak, Mykhailo Lobur, and Marek Iwaniec. A finite element model of terahertz substrate-based wire-grid polarizer. Mathematics and Computers in Simulation, 188:404–414, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001233.

Shloof:2021:EOM

[596] A. M. Shloof, N. Senu, A. Ahmadian, and Soheil Salahshour. An efficient operation matrix method for solving fractal–fractional differential equations with generalized Caputo-type fractional–fractal derivative. Mathematics and Computers in Simulation, 188:415–435, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001506.

Chrispell:2021:SRN

[597] J. C. Chrispell, E. W. Jenkins, and P. Westerbaan. Simulation of ratcheting non-Newtonian fluids using a geometric anisotropy. Mathematics and Computers in Simulation, 188: 436–454, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001518.

Wang:2021:CEM

[598] Yiran Wang, Eric Chung, and Lina Zhao. Constraint energy minimization generalized multiscale finite element method in mixed formulation for parabolic equations. Mathematics and Computers in Simulation, 188: 455–475, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001348.

Mohammadizadeh:2021:STF

[599] Fatemeh Mohammadizadeh, Saeede Rashidi, and S. Reza Hejazi. Space– time fractional Klein–Gordon equation: Symmetry analysis, conservation laws and numerical approximations. Mathematics and Computers in Simulation, 188:476–497, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001336.

Cheng:2021:ISI

[600] Hao Cheng. Importance sampling imputation algorithms in quantile regression with their application in CGSS data. Mathematics and Computers in Simulation, 188: 498–508, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001324.

Shi:2021:DAF

[601] Ruiqing Shi, Ting Lu, and Cuihong Wang. Dynamic analysis of a fractional-order model for HIV with drug-resistance and CTL immune response. Mathematics and Computers in Simulation, 188:509– 536, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421001531.

Kabeto:2021:RNM

[602] Masho Jima Kabeto and Gemechis File Duressa. Robust numerical method for singularly perturbed semilinear parabolic differential difference equations. Mathematics and Computers in Simulation, 188:537–547, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001762.

Qi:2021:VDB

[603] Kai Qi, Daqing Jiang, Tasawar Hayat, and Ahmed Alsaedi. Virus dynamic behavior of a stochastic HIV/AIDS infection model including two kinds of target cell infections and CTL immune responses. Mathematics and Computers in Simulation, 188:548– 570, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001804.

Rao:2021:ISS

[604] Ruofeng Rao, Xinsong Yang, Rongqiang Tang, Yulin Zhang, Xinggui Li, and Lei Shi. Impulsive stabilization and stability analysis for Gilpin–Ayala competition model involved in harmful species via LMI approach and variational methods. Mathematics and Computers in Simulation, 188:571– 590, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001567.

Georgiev:2021:CUV

[605] Slavi G. Georgiev and Lubin G. Vulkov. Computation of the unknown volatility from integral option price observations in jump–diffusion models. Mathematics and Computers in Simulation, 188:591–608, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001798.

Anonymous:2021:NIi

[606] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 188:609, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421002111.

Anonymous:2021:ICEh

[607] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 188:610–613, October 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421002123.

Anonymous:2021:EBk

[608] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 188:ifc, October 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421002056.

Blaheta:2021:MSI

[609] Radim Blaheta, Jaroslav Haslinger, Stanislav Sysala, Peter Arbenz, and Johannes Kraus. MATCOM special issue Modelling 2019: International Conference on Mathematical Modelling and Computational Methods in Applied Sciences and Engineering. Mathematics and Computers in Simulation, 189:1–2, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001154.

Sysala:2021:LBP

[610] Stanislav Sysala. Laudation for the 70th birthday of Professor Radim Blaheta. Mathematics and Computers in Simulation, 189:3– 4, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001579.

Dostalik:2021:UFA

[611] Mark Dostalík, Vít Průša, and Judith Stein. Unconditional finite amplitude stability of a viscoelastic fluid in a mechanically isolated vessel with spatially non-uniform wall temperature. Mathematics and Computers in Simulation, 189:5–20, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420301683.

Burkotova:2021:OTD

[612] Jana Burkotová, Jitka Machalová, and Jana Radová. Optimal thickness distribution of stepped nonlinear Gao beam. Mathematics and Computers in Simulation, 189:21– 35, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420301671.

Vala:2021:EST

[613] J. Vala, I. Němec, and A. Vaněčková. Exact solution of a thick beam on Pasternak subsoil in finite element calculations. Mathematics and Computers in Simulation, 189:36– 54, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420301762.

Haslinger:2021:PIH

[614] Jaroslav Haslinger, Radim Blaheta, and Raino A. E. Mäkinen. Parameter identification for heterogeneous materials by optimal control approach with flux cost functionals. Mathematics and Computers in Simulation, 189:55–68, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420302111.

Company:2021:FFE

[615] Rafael Company, Vera N. Egorova, and Lucas Jódar. A front-fixing ETD numerical method for solving jump–diffusion American option pricing problems. Mathematics and Computers in Simulation, 189:69– 84, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420302433.

Harizanov:2021:NFD

[616] Stanislav Harizanov, Nikola Kosturski, Svetozar Margenov, and Yavor Vutov. Neumann fractional diffusion problems: BURA solution methods and algorithms. Mathematics and Computers in Simulation, 189: 85–98, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420302469.

Chirila:2021:WPD

[617] Adina Chirilă and Marin Marin. Wave propagation in diffusive microstretch thermoelasticity. Mathematics and Computers in Simulation, 189:99– 113, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420302810.

Chorfi:2021:FDD

[618] Amina Chorfi, Jonas Koko, and Olivier Bodart. A fictitious domain decomposition method for a nonlinear bonded structure. Mathematics and Computers in Simulation, 189: 114–125, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420303116.

Adamek:2021:TLE

[619] V. Adámek, A. Berezovski, M. Mračko, and R. Kolman. A two-layer elastic strip under transverse impact loading: Analytical solution, finite element, and finite volume simulations. Mathematics and Computers in Simulation, 189: 126–140, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303487.

Anselmann:2021:HOG

[620] Mathias Anselmann and Markus Bause. Higher order Galerkin– collocation time discretization with Nitsche’s method for the Navier– Stokes equations. Mathematics and Computers in Simulation, 189:141– 162, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420303827.

Mracko:2021:FEM

[621] M. Mračko, J. Kober, R. Kolman, Z. Převorovský, A. Tkachuk, and J. Plešek. Finite element method based computational time reversal in elastodynamics: Refocusing, reconstruction and its numerical sensitivity. Mathematics and Computers in Simulation, 189:163–190, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303979.

Haslinger:2021:DSS

[622] Jaroslav Haslinger, Radek Kučera, Taoufik Sassi, and Václav Šátek. Dual strategies for solving the Stokes problem with stick–slip boundary conditions in 3D. Mathematics and Computers in Simulation, 189:191– 206, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475420304705.

Brezina:2021:EFE

[623] Jan Březina and Pavel Exner. Extended finite element method in mixedhybrid model of singular groundwater flow. Mathematics and Computers in Simulation, 189:207–236, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475420304730.

Egger:2021:SPH

[624] H. Egger and M. Sabouri. On the structure preserving high-order approximation of quasistatic poroelasticity. Mathematics and Computers in Simulation, 189:237–252, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542030495X.

Ayed:2021:EWS

[625] Hela Ayed, Leonardo Baffico, and Taoufik Sassi. Existence of a weak solution to a steady 2D fluid–1D elastic structure interaction problem with Tresca slip boundary condition. Mathematics and Computers in Simulation, 189:253–275, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475420305000.

Axelsson:2021:CMB

[626] Owe Axelsson, Michal Béreš, and Radim Blaheta. Computational methods for boundary optimal control and identification problems. Mathematics and Computers in Simulation, 189:

276–290, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000586.

Bailova:2021:MPA

[627] Michaela Bailová and Jiří Bouchala. A mountain pass algorithm for quasilinear boundary value problem with p-Laplacian. Mathematics and Computers in Simulation, 189:291–304, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000811.

Kolman:2021:BPS

[628] Radek Kolman, Ján Kopačka, José A. González, S. S. Cho, and K. C. Park. Bi-penalty stabilized technique with predictor–corrector time scheme for contact-impact problems of elastic bars. Mathematics and Computers in Simulation, 189:305– 324, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421000987.

Lukas:2021:DAD

[629] Dalibor Lukáš and Joachim Schöberl. Dispersion analysis of displacementbased and TDNNS mixed finite elements for thin-walled elastodynamics. Mathematics and Computers in Simulation, 189:325–338, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542100121X.

Rojicek:2021:MMI

[630] Jaroslav Rojíček, Martin Čermák, Radim Halama, Zbyněk Paška, and Milan Vaško. Material model identification from set of experiments and validation by DIC. Mathematics and Computers in Simulation, 189: 339–367, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001257.

Francu:2021:PUH

[631] Jan Franců. Problems with uncertain hysteresis operators and homogenization. Mathematics and Computers in Simulation, 189:368–379, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001543.

Anonymous:2021:NIj

[632] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 189:380, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100166X.

Anonymous:2021:ICEi

[633] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 189:381–384, November 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421001671.

Anonymous:2021:EBl

[634] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 189:ifc, November 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001592.

Kamra:2021:SPA

[635] Rabia Kamra and S. Chandra Sekhara Rao. A stable parallel algorithm for block tridiagonal Toeplitz–block– Toeplitz linear systems. Mathematics and Computers in Simulation, 190(??): 1415–1440, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542100269X.

Anonymous:2021:EBm

[636] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 190(??):1441, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421002986.

Anonymous:2021:ICEj

[637] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 190(??):1442– 1445, December 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421002998.

Kumar:2021:ATC

[638] S. Kumar, Subir Das, and S. H. Ong. Analysis of tumor cells in the absence and presence of chemotherapeutic treatment: the case of Caputo-Fabrizio time fractional derivative. Mathematics and Computers in Simulation, 190:1–14, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001786.

Anonymous:2021:PD

[639] Anonymous. Pages 1–1446 (December 2021). Mathematics and Computers in Simulation, 190:1–1446, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Sommariva:2021:NHS

[640] A. Sommariva and M. Vianello. Numerical hyperinterpolation over spherical triangles. Mathematics and Computers in Simulation, 190:15– 22, December 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001749.

Peyroteo:2021:NNT

[641] M. M. A. Peyroteo, J. Belinha, and R. M. Natal Jorge. A new non targeted bone remodeling model combined with an interpolation meshless method. Mathematics and Computers in Simulation, 190:23– 37, December 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001816.

Marcelo:2021:MDC

[642] Alexandre Borges Marcelo, Flávio Alessandro Serrão Gonçalves, and Fernando Pinhabel Marafão. Model development and controller design for DC–DC Boost converters in BCM with parallel cellular architecture. Mathematics and Computers in Simulation, 190:38–56, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001750.

Ali:2021:AIH

[643] Bagh Ali, Sajjad Hussain, Mohammad Shafique, Danial Habib, and Ghulam Rasool. Analyzing the interaction of hybrid base liquid C _{2} H _{6} O _{2}–H _{2} O with hybrid nanomaterial Ag–MoS _{2} for unsteady rotational flow referred to an elongated surface using modified Buongiorno’s model: FEM simulation. Mathematics and Computers in Simulation, 190:57–74, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100183X.

Li:2021:BAN

[644] Xinxin Li, Hengguo Yu, Chuanjun Dai, Zengling Ma, Qi Wang, and Min Zhao. Bifurcation analysis of a new aquatic ecological model with aggregation effect. Mathematics and Computers in Simulation, 190:75– 96, December 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001877.

Salam:2021:ECP

[645] Abdulwahed Salam and Abdelaaziz El Hibaoui. Energy consumption prediction model with deep inception residual network inspiration and LSTM. Mathematics and Computers in Simulation, 190:97–109, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001774.

Guimaraes:2021:TSI

[646] Osvaldo Guimarães, William Labecca, and José R. C. Piqueira. Tensor solutions in irregular domains: Eigenvalue problems. Mathematics and Computers in Simulation, 190: 110–130, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001907.

Dehshalie:2021:SAF

[647] Maziar Ebrahimi Dehshalie, Meisam Kabiri, and Mahyar Ebrahimi Dehshali. Stability analysis and fixed-time control of credit risk contagion. Mathematics and Computers in Simulation, 190:131–139, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001956.

Xu:2021:NSP

[648] Da Xu. Numerical solution of partial integro-differential equation with a weakly singular kernel based on Sinc methods. Mathematics and Computers in Simulation, 190:140– 158, December 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001865.

Ren:2021:SAS

[649] Jie Ren, Qimin Zhang, Xining Li, Feilong Cao, and Ming Ye. A stochastic age-structured HIV/AIDS model based on parameters estimation and its numerical calculation. Mathematics and Computers in Simulation, 190: 159–180, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001555.

IvanyshynYaman:2021:NSG

[650] Olha Ivanyshyn Yaman and Gazi Özdemir. Numerical solution of a generalized boundary value problem for the modified Helmholtz equation in two dimensions. Mathematics and Computers in Simulation, 190: 181–191, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001853.

Mohammed:2021:IMN

[651] W. W. Mohammed and M. El-Morshedy. The influence of multiplicative noise on the stochastic exact solutions of the Nizhnik–Novikov– Veselov system. Mathematics and Computers in Simulation, 190:192– 202, December 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001932.

Jin:2021:MTU

[652] Ting Jin and Xiangfeng Yang. Monotonicity theorem for the uncertain fractional differential equation and application to uncertain financial market. Mathematics and Computers in Simulation, 190:203–221, December 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001890.

Abdel-Gawad:2021:DCM

[653] Hamdy I. Abdel-Gawad and Ahmed H. Abdel-Gawad. Discrete and continuum models of COVID-19 virus, formal solutions, stability and comparison with real data. Mathematics and Computers in Simulation, 190: 222–230, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001889.

Naserizadeh:2021:CRB

[654] L. Naserizadeh, M. Hadizadeh, and A. Amiraslani. Cubature rules based on a bivariate degree-graded alternative orthogonal basis and their applications. Mathematics and Computers in Simulation, 190:231–245, December 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001737.

Kumari:2021:EBM

[655] Sangeeta Kumari and Ranjit Kumar Upadhyay. Exploring the behavior of malware propagation on mobile wireless sensor networks: Stability and control analysis. Mathematics and Computers in Simulation, 190: 246–269, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001981.

Malik:2021:SIM

[656] Rahila Malik, Mehboob Alam, Shah Muhammad, Rashida Hussain, Asghar Ali, Namra Akram, Faisal Zaid Duraihem, and Anwar Ul Haq. Statistically inspired multi-shift Arnoldi projection for on-chip interconnects. Mathematics and Computers in Simulation, 190:418–428, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421001968.

Hossain:2021:BMF

[657] Mainul Hossain, N. C. Pati, Saheb Pal, Sourav Rana, Nikhil Pal, and G. C. Layek. Bifurcations and multistability in a food chain model with nanoparticles. Mathematics and Computers in Simulation, 190:808– 825, December 2021. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421002408.

Attene:2021:BGR

[658] Marco Attene, Silvia Biasotti, Silvia Bertoluzza, Daniela Cabiddu, Marco Livesu, Giuseppe Patanè, Micol Pennacchio, Daniele Prada, and Michela Spagnuolo. Benchmarking the geometrical robustness of a Virtual Element Poisson solver. Mathematics and Computers in Simulation, 190: 1392–1414, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002706.

Anonymous:2021:EBn

[659] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 190:ifc, December 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002937.

Hanif:2022:CAB

[660] Hanifa Hanif. A computational approach for boundary layer flow and heat transfer of fractional Maxwell fluid. Mathematics and Computers in Simulation, 191:1–13, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475421002743.

Anonymous:2022:PJa

[661] Anonymous. Pages 1–350 (January 2022). Mathematics and Computers in Simulation, 191:1–350, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Dassi:2022:DTR

[662] Franco Dassi, Julia M. Kroos, L. Gerardo-Giorda, and Simona Perotto. A denoising tool for the reconstruction of cortical geometries from MRI. Mathematics and Computers in Simulation, 191:14– 32, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421002688.

Fortes:2022:FMD

[663] M. A. Fortes and E. Medina. Fitting missing data by means of adaptive meshes of Bézier curves. Mathematics and Computers in Simulation, 191:33–48, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002755.

Barman:2022:IWD

[664] Dipesh Barman, Jyotirmoy Roy, and Shariful Alam. Impact of wind in the dynamics of prey– predator interactions. Mathematics and Computers in Simulation, 191: 49–81, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542100272X.

Mostafavi:2022:TMP

[665] Seyed Alireza Mostafavi and Zahra Joneidi. Thermal model of precast concrete curing process: Minimizing energy consumption. Mathematics and Computers in Simulation, 191:82–94, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002779.

Jana:2022:IEM

[666] Nabakumar Jana and Samadrita Bera. Interval estimation of multicomponent stress–strength reliability based on inverse Weibull distribution. Mathematics and Computers in Simulation, 191:95–119, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002767.

Lavanya:2022:LFC

[667] S. Lavanya and S. Nagarani. Leaderfollowing consensus of multi-agent systems with sampled-data control and looped functionals. Mathematics and Computers in Simulation, 191:120– 133, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421002792.

Das:2022:IFD

[668] Bijoy Kumar Das, Debgopal Sahoo, and G. P. Samanta. Impact of fear in a delay-induced predator–prey system with intraspecific competition within predator species. Mathematics and Computers in Simulation, 191: 134–156, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002822.

Xu:2022:ESW

[669] Guoan Xu, Yi Zhang, and Jibin Li. Exact solitary wave and periodic-peakon solutions of the complex Ginzburg– Landau equation: Dynamical system approach. Mathematics and Computers in Simulation, 191:157– 167, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421002846.

Shafiya:2022:GSU

[670] M. Shafiya, G. Nagamani, and D. Dafik. Global synchronization of uncertain fractional-order BAM neural networks with time delay via improved fractional-order integral inequality. Mathematics and Computers in Simulation, 191:168– 186, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421002780.

Mabood:2022:CTB

[671] F. Mabood, MD. Shamshuddin, and S. R. Mishra. Characteristics of thermophoresis and Brownian motion on radiative reactive micropolar fluid flow towards continuously moving flat plate: HAM solution. Mathematics and Computers in Simulation, 191: 187–202, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002810.

Vujovic:2022:IEF

[672] Vuk Vujović. Influence of environmental fluctuations on Hepatitis C transmission. Mathematics and Computers in Simulation, 191:203– 218, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421002858. See

corrigendum [1751].

Mahmoudi:2022:CNS

[673] Fatemeh Mahmoudi and Mahdieh Tahmasebi. The convergence of a numerical scheme for additive fractional stochastic delay equations with H > 1/2. Mathematics and Computers in Simulation, 191:219– 231, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421002871.

Ge:2022:SDD

[674] Junyan Ge, Wenjie Zuo, and Daqing Jiang. Stationary distribution and density function analysis of a stochastic epidemic HBV model. Mathematics and Computers in Simulation, 191:232–255, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002809.

Veerasamy:2022:IGA

[675] Gomathi Veerasamy, Ramkumar Kannan, RakeshKumar Siddharthan, Guruprasath Muralidharan, Venkatesh Sivanandam, and Rengarajan Amirtharajan. Integration of genetic algorithm tuned adaptive fading memory Kalman filter with model predictive controller for active fault-tolerant control of cement kiln under sensor faults with inaccurate noise covariance. Mathematics and Computers in Simulation, 191:256–277, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421002731.

Vargas:2022:SSS

[676] Alessandro N. Vargas, Constantin F. Caruntu, João Y. Ishihara, and Hassane Bouzahir. Stochastic stability of switching linear systems with application to an automotive powertrain model. Mathematics and Computers in Simulation, 191:278– 287, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421002834.

Chen:2022:CWH

[677] Qing-Bo Chen and Hang Xu. Coiflet wavelet-homotopy solution of free convection in a closed cavity subjected to an inclined external magnetic field. Mathematics and Computers in Simulation, 191:288–308, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542100286X.

Combes:2022:PEA

[678] Catherine Combes and Hon Keung Tony Ng. On parameter estimation for Amoroso family of distributions. Mathematics and Computers in Simulation, 191:309– 327, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421002548.

Issa:2022:PVS

[679] Mohamed Issa and Anas Samn. Passive vehicle suspension system optimization using Harris Hawk Optimization algorithm. Mathematics and Computers in Simulation, 191: 328–345, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003037.

Anonymous:2022:NIa

[680] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 191:346, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003323.

Anonymous:2022:ICEa

[681] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 191:347–350, January 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003335.

Anonymous:2022:EBa

[682] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 191:ifc, January 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003281.

Zhou:2022:CRD

[683] Liangqiang Zhou and Fangqi Chen. Chaos of the Rayleigh–Duffing oscillator with a non-smooth periodic perturbation and harmonic excitation. Mathematics and Computers in Simulation, 192:1–18, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002913.

Anonymous:2022:PF

[684] Anonymous. Pages 1–542 (February 2022). Mathematics and Computers in Simulation, 192:1–542, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Perez-Moroyoqui:2022:IEM

[685] René Pérez-Moroyoqui, Oscar Ibáñez-Orozco, and Suemi Rodríguez-Romo. An intelligent equation for methane hydrate growth kinetics. Mathematics and Computers in Simulation, 192:19–32, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002883.

Roy:2022:NMP

[686] N. Roy, R. Dürr, A. Bück, J. Kumar, and S. Sundar. Numerical methods for particle agglomeration and breakage in lid-driven cavity flows at low Reynolds numbers. Mathematics and Computers in Simulation, 192:33–49, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003025.

Golnary:2022:IDD

[687] Farshad Golnary and Hamed Moradi. Identification of the dynamics of the drivetrain and estimating its unknown parts in a large scale wind turbine. Mathematics and Computers in Simulation, 192:50– 69, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421002895.

Wang:2022:ICS

[688] Liang Wang, Yogesh Mani Tripathi, Chandrakant Lodhi, and Xuanjia Zuo. Inference for constantstress Weibull competing risks model under generalized progressive hybrid censoring. Mathematics and Computers in Simulation, 192:70– 83, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003049.

Hashim:2022:HBA

[689] Fatma A. Hashim, Essam H. Houssein, Kashif Hussain, Mai S. Mabrouk, and Walid Al-Atabany. Honey Badger Algorithm: New metaheuristic algorithm for solving optimization problems. Mathematics and Computers in Simulation, 192:84– 110, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421002901.

Mondal:2022:SDT

[690] Bapin Mondal, Uttam Ghosh, Md Sadikur Rahman, Pritam Saha, and Susmita Sarkar. Studies of different types of bifurcations analyses of an imprecise two species food chain model with fear effect and non-linear harvesting. Mathematics and Computers in Simulation, 192:111–135, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421003098.

Sk:2022:DNM

[691] Nazmul Sk, Pankaj Kumar Tiwari, and Samares Pal. A delay nonautonomous model for the impacts of fear and refuge in a three species food chain model with hunting cooperation. Mathematics and Computers in Simulation, 192:136–166, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421003086.

Padmaja:2022:MPB

[692] N. Padmaja and P. Balasubramaniam. Mixed H_{∞}/passivity based stability analysis of fractional-order gene regulatory networks with variable delays. Mathematics and Computers in Simulation, 192:167–181, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542100313X.

Li:2022:MGS

[693] Juntao Li and Pengfei Liu. Modeling green supply chain games with governmental interventions and risk preferences under fuzzy uncertainties. Mathematics and Computers in Simulation, 192:182–200, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003104.

Daba:2022:CMU

[694] Imiru Takele Daba and Gemechis File Duressa. Collocation method using artificial viscosity for time dependent singularly perturbed differential– difference equations. Mathematics and Computers in Simulation, 192: 201–220, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003207.

Gao:2022:FTP

[695] Guangyu Gao, Bo Han, and Shanshan Tong. A fast two-point gradient algorithm based on sequential subspace optimization method for nonlinear ill-posed problems. Mathematics and Computers in Simulation, 192: 221–245, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003190.

Butt:2022:MSS

[696] Muhammad Munir Butt. On a multigrid solver for stationary Navier– Stokes velocity–pressure tracking-type control problems. Mathematics and Computers in Simulation, 192:246– 264, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421003153.

Huang:2022:FDE

[697] Qiong-Ao Huang, Gengen Zhang, and Bing Wu. Fully-discrete energypreserving scheme for the spacefractional Klein–Gordon equation via Lagrange multiplier type scalar auxiliary variable approach. Mathematics and Computers in Simulation, 192: 265–277, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421003177.

Anatriello:2022:GPT

[698] Giuseppina Anatriello, László Németh, and Giovanni Vincenzi. Generalized Pascal’s triangles and associated k-Padovan-like sequences. Mathematics and Computers in Simulation, 192: 278–290, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421003219.

Divya:2022:NFC

[699] H. Divya, R. Sakthivel, S. A. Karthick, and C. Aouiti. Non-fragile control design for stochastic Markov jump system with multiple delays and cyber attacks. Mathematics and Computers in Simulation, 192: 291–302, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421003165.

Seo:2022:MSE

[700] Jun-Ho Seo and Jeong-Hoon Kim. Multiscale stochastic elasticity of variance for options and equity linked annuity; a Mellin transform approach. Mathematics and Computers in Simulation, 192:303–320, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003189.

Yacine:2022:FTI

[701] Zedjiga Yacine, Hamid Hamiche, Saïd Djennoune, and Saïd Mammar. Finitetime impulsive observers for nonlinear systems represented by Takagi– Sugeno models: Application to a chaotic system. Mathematics and Computers in Simulation, 192:321– 352, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003232.

Jolaoso:2022:IET

[702] Lateef O. Jolaoso, Yekini Shehu, and Jen-Chih Yao. Inertial extragradient type method for mixed variational inequalities without monotonicity. Mathematics and Computers in Simulation, 192:353–369, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003256.

Tharanidharan:2022:RFT

[703] V. Tharanidharan, R. Sakthivel, Yong Ren, and S. Marshal Anthoni. Robust finite-time PID control for discretetime large-scale interconnected uncertain system with discrete-delay. Mathematics and Computers in Simulation, 192:370–383, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003141.

Hosseini-Ara:2022:CAH

[704] Reza Hosseini-Ara, Ali Mokhtarian, Amir Hossein Karamrezaei, and Davood Toghraie. Computational analysis of high precision nano-sensors for diagnosis of viruses: Effects of partial antibody layer. Mathematics and Computers in Simulation, 192: 384–398, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003244.

Ebert:2022:CGP

[705] Adrian Ebert, Peter Kritzer, Onyekachi Osisiogu, and Tetiana Stepaniuk. Construction of good polynomial lattice rules in weighted Walsh spaces by an alternative component-by-component construction. Mathematics and Computers in Simulation, 192: 399–419, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003220.

Tang:2022:PSS

[706] Xiaosong Tang. Periodic solutions and spatial patterns induced by mixed delays in a diffusive spruce budworm model with Holling II predation function. Mathematics and Computers in Simulation, 192:420– 429, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003360.

ElKouche:2022:MNS

[707] Amal El Kouche and Francisco Ortegón Gallego. Modeling and numerical simulation of a parabolic trough collector using an HTF with temperature dependent physical properties. Mathematics and Computers in Simulation, 192:430–451, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100344X.

Hazarika:2022:BMT

[708] Silpi Hazarika and Sahin Ahmed. Brownian motion and thermophoresis behavior on micro-polar nano-fluid — a numerical outlook. Mathematics and Computers in Simulation, 192: 452–463, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003359.

Ahuja:2022:TAA

[709] Anjali Ahuja, Anamika Jain, and Madhu Jain. Transient analysis and ANFIS computing of unreliable single server queueing model with multiple stage service and functioning vacation. Mathematics and Computers in Simulation, 192:464– 490, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421003268.

Singh:2022:SAN

[710] Anup Singh, Subir Das, and S. H. Ong. Study and analysis of nonlinear (2 + 1)-dimensional solute transport equation in porous media. Mathematics and Computers in Simulation, 192:491–500, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421003128.

Kumar:2022:CSM

[711] Sachin Kumar, Juan J. Nieto, and Bashir Ahmad. Chebyshev spectral method for solving fuzzy fractional Fredholm–Volterra integrodifferential equation. Mathematics and Computers in Simulation, 192: 501–513, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421003463.

Onay:2022:CHG

[712] Funda Kutlu Onay and Salih Berkan Aydemir. Chaotic hunger games search optimization algorithm for global optimization and engineering problems. Mathematics and Computers in Simulation, 192:514– 536, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421003426.

Anonymous:2022:NIb

[713] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 192:537, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003712.

Anonymous:2022:ICEb

[714] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 192:538–541, February 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003724.

Anonymous:2022:EBb

[715] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 192:ifc, February 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003669.

Mondal:2022:ICN

[716] Suman Mondal, Ranjit Ray, Srinivasa Reddy N., and Sambhunath Nandy. Intelligent controller for nonholonomic wheeled mobile robot: a fuzzy path following combination. Mathematics and Computers in Simulation, 193(??):533– 555, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542100392X.

Win:2022:NMB

[717] Aung Naing Win and Mingming Li. Numerical method based on fiber bundle for solving Lyapunov matrix equation. Mathematics and Computers in Simulation, 193(??):556– 566, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003943.

Razzaghian:2022:FNN

[718] Amir Razzaghian. A fuzzy neural network-based fractional-order Lyapunov-based robust control strategy for exoskeleton robots: Application in upper-limb rehabilitation. Mathematics and Computers in Simulation, 193(??):567–583, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003876.

She:2022:TMS

[719] Mianfu She, Dongfang Li, and Hai wei Sun. A transformed L1 method for solving the multi-term time-fractional diffusion problem. Mathematics and Computers in Simulation, 193(??): 584–606, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004018.

Akram:2022:ETT

[720] Ghazala Akram, Maasoomah Sadaf, Muhammad Abbas, Iqra Zainab, and Syeda Rijaa Gillani. Efficient techniques for traveling wave solutions of time-fractional Zakharov–Kuznetsov equation. Mathematics and Computers in Simulation, 193(??):607– 622, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004006.

Kumar:2022:RAE

[721] Udai Kumar and Partha Sarathi Mandal. Role of Allee effect on prey– predator model with component Allee effect for predator reproduction. Mathematics and Computers in Simulation, 193(??):623–665, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003918.

Chung:2022:GMD

[722] Eric T. Chung, Uygulaana Kalachikova, Maria Vasilyeva, and Valentin Alekseev. Generalized multiscale discontinuous Galerkin method for convection–diffusion equation in perforated media. Mathematics and Computers in Simulation, 193(??):666– 688, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003979.

Shama:2022:GGD

[723] M. S. Shama, Sanku Dey, Emrah Altun, and Ahmed Z. Afify. The Gamma–Gompertz distribution: Theory and applications. Mathematics and Computers in Simulation, 193 (??):689–712, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542100389X.

Cui:2022:RRP

[724] Xiaoting Cui, Jingwei Ji, Xuehe Bai, Yin Cao, and Tong Wu. Research and realization of parallel algorithms for large scale crowd evacuation in emergency. Mathematics and Computers in Simulation, 193 (??):713–724, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421003906.

Anonymous:2022:NIc

[725] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 193(??):725, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421004262.

Anonymous:2022:ICEc

[726] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 193(??):726– 729, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421004274.

Cumsille:2022:FCC

[727] Patricio Cumsille, Óscar Rojas-Díaz, Pablo Moisset de Espanés, and Paula Verdugo-Hernández. Forecasting COVID-19 Chile’ second outbreak by a generalized SIR model with constant time delays and a fitted positivity rate.

Mathematics and Computers in Simulation, 193:1–18, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003451.

Anonymous:2022:PMa

[728] Anonymous. Pages 1–730 (March 2022). Mathematics and Computers in Simulation, 193:1–730, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Xie:2022:MOB

[729] Yingying Xie and Lingfei Li. Multipleorder breathers for a generalized (3+1)- dimensional Kadomtsev–Petviashvili Benjamin–Bona–Mahony equation near the offshore structure. Mathematics and Computers in Simulation, 193:19–31, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003116.

Hoang:2022:RAH

[730] Manh Tuan Hoang. Reliable approximations for a hepatitis B virus model by nonstandard numerical schemes. Mathematics and Computers in Simulation, 193:32– 56, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003499.

DeCesare:2022:NLA

[731] Luigi De Cesare and Mario Sportelli. A non-linear approach to Kalecki’s investment cycle. Mathematics and Computers in Simulation, 193:57– 70, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003529.

Li:2022:CSB

[732] Maodong Li, Guanghui Xu, Qiang Lai, and Jie Chen. A chaotic strategy-based quadratic Opposition-Based Learning adaptive variable-speed whale optimization algorithm. Mathematics and Computers in Simulation, 193: 71–99, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003542.

Preeti:2022:NIH

[733] Preeti and Odelu Ojjela. Numerical investigation of heat transport in Alumina–Silica hybrid nanofluid flow with modeling and simulation. Mathematics and Computers in Simulation, 193:100–122, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003517.

Kuddus:2022:MAH

[734] Md Abdul Kuddus and Azizur Rahman. Modelling and analysis of human–mosquito malaria transmission dynamics in Bangladesh. Mathematics and Computers in Simulation, 193:123–138, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003505.

Liu:2022:MASa

[735] Chein-Shan Liu and Chih-Wen Chang. Modified asymptotic solutions for second-order nonlinear singularly perturbed boundary value problems. Mathematics and Computers in Simulation, 193:139–152, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475421003566.

Garza:2022:CCM

[736] Gabriel López Garza, Aurelio Nicolás Mata, Graciela Román Alonso, José Rafael Godínez Fernández, and Miguel Alfonso Castro García. Cell-to-cell mathematical modeling of arrhythmia phenomena in the heart. Mathematics and Computers in Simulation, 193:153–172, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003530.

Chen:2022:GES

[737] Yonghui Chen, Xian Zhang, and Yu Xue. Global exponential synchronization of high-order quaternion Hopfield neural networks with unbounded distributed delays and timevarying discrete delays. Mathematics and Computers in Simulation, 193:173–189, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003645.

Sun:2022:SLP

[738] Lin Sun, Yiming Chen, Rongqi Dang, Gang Cheng, and Jiaquan Xie. Shifted Legendre polynomials algorithm used for the numerical analysis of viscoelastic plate with a fractional order model. Mathematics and Computers in Simulation, 193:190–203, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100358X.

Vargas:2022:FDM

[739] Antonio M. Vargas. Finite difference method for solving fractional differential equations at irregular meshes. Mathematics and Computers in Simulation, 193:204– 216, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542100361X.

Karnan:2022:NFS

[740] A. Karnan and G. Nagamani. Nonfragile state estimation for memristive cellular neural networks with proportional delay. Mathematics and Computers in Simulation, 193:217– 231, March 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003608.

Khan:2022:IDB

[741] Umair Khan, A. Zaib, A. Ishak, Sakhinah Abu Bakar, I. L. Animasaun, and Se-Jin Yook. Insights into the dynamics of blood conveying gold nanoparticles on a curved surface when suction, thermal radiation, and Lorentz force are significant: the case of Non-Newtonian Williamson fluid. Mathematics and Computers in Simulation, 193:250–268, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003761.

Pan:2022:GEO

[742] Jeng-Shyang Pan, Ji-Xiang Lv, Li-Jun Yan, Shao-Wei Weng, Shu-Chuan Chu, and Jian-Kai Xue. Golden eagle optimizer with double learning strategies for 3D path planning of UAV in power inspection. Mathematics and Computers in Simulation, 193:509–532, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003967.

Anonymous:2022:EBc

[743] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 193:ifc, March 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004213.

Ghosh:2022:MMC

[744] Jayanta Kumar Ghosh, Sudhanshu Kumar Biswas, Susmita Sarkar, and Uttam Ghosh. Mathematical modelling of COVID-19: a case study of Italy. Mathematics and Computers in Simulation, 194:1–18, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004043.

Anonymous:2022:PAa

[745] Anonymous. Pages 1–786 (April 2022). Mathematics and Computers in Simulation, 194:1–786, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ahmetolan:2022:SIS

[746] Semra Ahmetolan, Ayse Humeyra Bilge, Ali Demirci, and Ayse Peker Dobie. A Susceptible–Infectious (SI) model with two infective stages and an endemic equilibrium. Mathematics and Computers in Simulation, 194: 19–35, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421003992.

Torregrosa:2022:SPM

[747] Sergio Torregrosa, Victor Champaney, Amine Ammar, Vincent Herbert, and Francisco Chinesta. Surrogate parametric metamodel based on Optimal Transport. Mathematics and Computers in Simulation, 194: 36–63, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421004067.

Gharahasanlou:2022:GSA

[748] Tohid Kasbi Gharahasanlou, Vahid Roomi, and Zeynab Hemmatzadeh. Global stability analysis of viral infection model with logistic growth rate, general incidence function and cellular immunity. Mathematics and Computers in Simulation, 194:64– 79, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004110.

Amirjanov:2022:MDC

[749] Adil Amirjanov and Konstantin Sobolev. Modeling the dynamics of a changing range genetic algorithm in noisy environments. Mathematics and Computers in Simulation, 194:80–88, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003980.

Wu:2022:SBH

[750] Jianping Wu. N-soliton, M-breather and hybrid solutions of a timedependent Kadomtsev–Petviashvili equation. Mathematics and Computers in Simulation, 194:89–96, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475421003621.

Luo:2022:HTM

[751] Wei-Hua Luo, Xian-Ming Gu, and Bruno Carpentieri. A hybrid triangulation method for banded linear systems. Mathematics and Computers in Simulation, 194:97– 108, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004080.

Gokce:2022:SSP

[752] Aytül Gökçe, Samire Yazar, and Yadigar Sekerci. Stability of spatial patterns in a diffusive oxygen–plankton model with time lag effect. Mathematics and Computers in Simulation, 194:109–123, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100402X.

Tayebi:2022:LTN

[753] Tahar Tayebi, Ali J. Chamkha, Hakan F. Öztop, and Lynda Bouzeroura. Local thermal non-equilibrium (LTNE) effects on thermal-free convection in a nanofluid-saturated horizontal elliptical non-Darcian porous annulus. Mathematics and Computers in Simulation, 194:124–140, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475421004079.

AlSaadi:2022:LSS

[754] Fahad Al Saadi, Annette Worthy, Haifaa Alrihieli, and Mark Nelson. Localised spatial structures in the Thomas model. Mathematics and Computers in Simulation, 194:141– 158, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003955.

Yang:2022:SSD

[755] Feng Yang and Yongli Song. Stability and spatiotemporal dynamics of a diffusive predator–prey system with generalist predator and nonlocal intraspecific competition. Mathematics and Computers in Simulation, 194:159–168, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004092.

Zara:2022:NAM

[756] Aiman Zara, Shafiq Ur Rehman, Fayyaz Ahmad, Salima Kouser, and Anjum Pervaiz. Numerical approximation of modified Kawahara equation using Kernel smoothing method. Mathematics and Computers in Simulation, 194:169–184, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004109.

Li:2022:TSL

[757] Zhengming Li, M. N. Smirnova, Yongliang Zhang, N. N. Smirnov, and Zuojin Zhu. Tunnel speed limit effects on traffic flow explored with a three lane model. Mathematics and Computers in Simulation, 194: 185–197, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004122.

Taherinasab:2022:AMR

[758] Sina Taherinasab, Saedeh Soleimaniasl, and Saleh Taherinasab. Application of model reference adaptive control and modified Smith predictor to control blood glucose in type 1 diabetic patients. Mathematics and Computers in Simulation, 194:198–209, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002159.

Pourbabaee:2022:NOM

[759] Marzieh Pourbabaee and Abbas Saadatmandi. A new operational matrix based on Müntz-Legendre polynomials for solving distributed order fractional differential equations. Mathematics and Computers in Simulation, 194:210–235, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004195.

Wang:2022:MDE

[760] Beibei Wang, Li Chen, Jiale Wang, and Shengnan Zhao. Microgrid distributed energy resources planning based on a long-term dynamic microsimulation. Mathematics and Computers in Simulation, 194:236– 253, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004134.

Ali:2022:BSM

[761] Bagh Ali, Anum Shafiq, Abdul Manan, Abderrahim Wakif, and Sajjad Hussain. Bioconvection: Significance of mixed convection and mhd on dynamics of Casson nanofluid in the stagnation point of rotating sphere via finite element simulation. Mathematics and Computers in Simulation, 194:254–268, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004158.

Postavaru:2022:NAB

[762] Octavian Postavaru and Antonela Toma. A numerical approach based on fractional-order hybrid functions of block-pulse and Bernoulli polynomials for numerical solutions of fractional optimal control problems. Mathematics and Computers in Simulation, 194:269–284, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004353.

Wang:2022:ADL

[763] Liang Wang, Yogesh Mani Tripathi, Sanku Dey, Chunfang Zhang, and Ke Wu. Analysis of dependent lefttruncated and right-censored competing risks data with partially observed failure causes. Mathematics and Computers in Simulation, 194:285–307, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100433X.

Goncalves:2022:TOU

[764] D. C. Gonçalves, J. D. F. Lopes, R. D. S. G. Campilho, and J. Belinha. Topology optimization using a natural neighbour meshless method combined with a bi-directional evolutionary algorithm. Mathematics and Computers in Simulation, 194: 308–328, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004171.

Qiao:2022:FAD

[765] Leijie Qiao, Da Xu, Bo Tang, and Jun Zhou. Fast ADI difference/compact difference schemes for the nonlocal evolution equation with weakly singular kernels in three dimensions. Mathematics and Computers in Simulation, 194:329–347, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421004390.

Shi:2022:CHB

[766] Jianping Shi, Ke He, and Hui Fang. Chaos, Hopf bifurcation and control of a fractional-order delay financial system. Mathematics and Computers in Simulation, 194:348– 364, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421004432.

Rezapour:2022:TAC

[767] Sh. Rezapour, S. Kumar, M. Q. Iqbal, A. Hussain, and S. Etemad. On two abstract Caputo multi-term sequential fractional boundary value problems under the integral conditions. Mathematics and Computers in Simulation, 194:365–382, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421004146.

Bader:2022:EAC

[768] Philipp Bader, Sergio Blanes, Fernando Casas, and Muaz Seydaoğlu. An efficient algorithm to compute the exponential of skew-Hermitian matrices for the time integration of the Schrödinger equation. Mathematics and Computers in Simulation, 194:383–400, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004365.

Barrera:2022:NAD

[769] D. Barrera, S. Eddargani, M. J. Ibáñez, and A. Lamnii. A new approach to deal with C^{2} cubic splines and its application to superconvergent quasi-interpolation. Mathematics and Computers in Simulation, 194:401–415, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004377.

Salahuddin:2022:SSF

[770] T. Salahuddin, Nazim Siddique, Mair Khan, and Mohamed Altanji. A significant study on flow analysis of viscoelastic fluid with variable thermophysical properties. Mathematics and Computers in Simulation, 194: 416–429, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004304.

Shahni:2022:NSE

[771] Julee Shahni and Randhir Singh. Numerical simulation of Emden–Fowler integral equation with Green’s function type kernel by Gegenbauer-wavelet, Taylor-wavelet and Laguerre-wavelet collocation methods. Mathematics and Computers in Simulation, 194: 430–444, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421004420.

Adamu:2022:IVT

[772] A. Adamu, D. Kitkuan, A. Padcharoen, C. E. Chidume, and P. Kumam. Inertial viscosity-type iterative method for solving inclusion problems with applications. Mathematics and Computers in Simulation, 194:445–459, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421004419.

Das:2022:CVC

[773] Anusmita Das, Hemanta Kr. Sarmah, Debashish Bhattacharya, Kaushik Dehingia, and Kamyar Hosseini. Combination of virotherapy and chemotherapy with optimal control for combating cancer. Mathematics and Computers in Simulation, 194:460– 488, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421004389.

Tay:2022:DEC

[774] Chai Jian Tay, Muhammad Fakhruddin, Ilham Saiful Fauzi, Su Yean Teh, Muhammad Syamsuddin, Nuning Nuraini, and Edy Soewono. Dengue epidemiological characteristic in Kuala Lumpur and Selangor, Malaysia. Mathematics and Computers in Simulation, 194:489–504, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004407.

Al-Qudah:2022:LBC

[775] Alaa Al-Qudah, Zaid Odibat, and Nabil Shawagfeh. A linearizationbased computational algorithm of homotopy analysis method for nonlinear reaction–diffusion systems. Mathematics and Computers in Simulation, 194:505–522, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004341.

Craven:2022:PEM

[776] Brent A. Craven, Mohammad M. Faghih, Kenneth I. Aycock, and Ebrahim M. Kolahdouz. A Poisson equation method for prescribing fully developed non-Newtonian inlet conditions for computational fluid dynamics simulations in models of arbitrary cross-section. Mathematics and Computers in Simulation, 194: 523–538, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004055.

Deng:2022:BSF

[777] Aimin Deng, Ji Lin, and Chein-Shan Liu. Boundary shape function iterative method for nonlinear second-order boundary value problems with nonlinear boundary conditions. Mathematics and Computers in Simulation, 194:539–551, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542100447X.

El-Dib:2022:DHR

[778] Yusry O. El-Dib. The damping Helmholtz–Rayleigh–Duffing oscillator with the non-perturbative approach. Mathematics and Computers in Simulation, 194:552–562, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475421004481.

Kumbhakar:2022:UMR

[779] Bidyasagar Kumbhakar and Susmay Nandi. Unsteady MHD radiativedissipative flow of Cu-Al_{2}O_{3} /H_{2}O hybrid nanofluid past a stretching sheet with slip and convective conditions: a regression analysis. Mathematics and Computers in Simulation, 194:563–587, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421004523.

Faheem:2022:HRH

[780] Mo Faheem, Arshad Khan, and Akmal Raza. A high resolution Hermite wavelet technique for solving space–time-fractional partial differential equations. Mathematics and Computers in Simulation, 194:588– 609, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421004468.

Derakhshannia:2022:DOB

[781] Mehran Derakhshannia and Seyyed Sajjad Moosapour. Disturbance observerbased sliding mode control for consensus tracking of chaotic nonlinear multi-agent systems. Mathematics and Computers in Simulation, 194:610–628, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004511.

Alimoradi:2022:TSR

[782] Mahmoud Alimoradi, Hossein Azgomi, and Ali Asghari. Trees Social Relations Optimization Algorithm: a new swarm-based metaheuristic technique to solve continuous and discrete optimization problems. Mathematics and Computers in Simulation, 194:629–664, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004444.

Fu:2022:SBA

[783] Shuaiming Fu, Jianfeng Luo, and Yi Zhao. Stability and bifurcations analysis in an ecoepidemic system with prey group defense and two infectious routes. Mathematics and Computers in Simulation, 194:665–690, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004183.

Lamboni:2022:WDB

[784] Matieyendou Lamboni. Weak derivativebased expansion of functions: ANOVA and some inequalities. Mathematics and Computers in Simulation, 194:691–718, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004535.

Chiu:2022:NSR

[785] Kuo-Shou Chiu and Tongxing Li. New stability results for bidirectional associative memory neural networks model involving generalized piecewise constant delay. Mathematics and Computers in Simulation, 194: 719–743, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542100450X.

Ammad:2022:NGT

[786] Muhammad Ammad, Md Yushalify Misro, and Ahmad Ramli. A novel generalized trigonometric Bézier curve: Properties, continuity conditions and applications to the curve modeling. Mathematics and Computers in Simulation, 194:744– 763, April 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004456.

Patil:2022:IMC

[787] P. M. Patil, Sunil Benawadi, and Bandari Shanker. Influence of mixed convection nanofluid flow over a rotating sphere in the presence of diffusion of liquid hydrogen and ammonia. Mathematics and Computers in Simulation, 194:764–781, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004560.

Anonymous:2022:NId

[788] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 194:782, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000179.

Anonymous:2022:ICEd

[789] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 194:783–786, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475422000180.

Anonymous:2022:EBd

[790] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 194:ifc, April 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200012X.

Chen:2022:DCP

[791] Jianxin Chen, Junhao Zheng, Tonghua Zhang, Rui Hou, and Yong wu Zhou. Dynamical complexity of pricing and green level for a dyadic supply chain with capital constraint. Mathematics and Computers in Simulation, 195:1–21, May 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421004493.

Anonymous:2022:PMb

[792] Anonymous. Pages 1–232 (May 2022). Mathematics and Computers in Simulation, 195:1–232, May 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Bokhari:2022:PMU

[793] Ahmed Bokhari, Rachid Belgacem, Sunil Kumar, Dumitru Baleanu, and Salih Djilali. Projectile motion using three parameter Mittag-Leffler function calculus. Mathematics and Computers in Simulation, 195:22–30, May 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421004547.

Saha:2022:AHV

[794] Sangeeta Saha and Guruprasad Samanta. Analysis of a host–vector dynamics of a dengue disease model with optimal vector control strategy. Mathematics and Computers in Simulation, 195:31–55, May 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421004559.

Zhang:2022:SBD

[795] Xiaofeng Zhang and Rong Yuan. Stochastic bifurcation and density function analysis of a stochastic logistic equation with distributed delay and weak kernel. Mathematics and Computers in Simulation, 195:

56–70, May 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004572.

Basic-Sisko:2022:ODM

[796] Angela Bašić-Šiško, Ivan Dražić, and Loredana Simčić. One-dimensional model and numerical solution to the viscous and heat-conducting micropolar real gas flow with homogeneous boundary conditions. Mathematics and Computers in Simulation, 195: 71–87, May 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004584.

Singh:2022:UCQ

[797] Satpal Singh, Devendra Kumar, and Higinio Ramos. A uniformly convergent quadratic B-spline based scheme for singularly perturbed degenerate parabolic problems. Mathematics and Computers in Simulation, 195: 88–106, May 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421004602.

Zheng:2022:NAN

[798] Xiangcheng Zheng. Numerical approximation for a nonlinear variableorder fractional differential equation via a collocation method. Mathematics and Computers in Simulation, 195:107–118, May 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000052.

Chaudhary:2022:LSG

[799] Sudhakar Chaudhary and Pari J. Kundaliya. L1 scheme on graded mesh for subdiffusion equation with nonlocal diffusion term. Mathematics and Computers in Simulation, 195: 119–137, May 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000064.

Song:2022:ATD

[800] Ying-Qing Song, Hassan Waqas, Kamel Al-Khaled, Umar Farooq, Soumaya Gouadria, M. Imran, Sami Ullah Khan, M. Ijaz Khan, Sumaira Qayyum, and Qiu-Hong Shi. Aspects of thermal diffusivity and melting phenomenon in Carreau nanofluid flow confined by nonlinear stretching cylinder with convective Marangoni boundary constraints. Mathematics and Computers in Simulation, 195:138– 150, May 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000015.

Sabbagh-Yazdi:2022:EDE

[801] Saeed-Reza Sabbagh-Yazdi, Arwin Farhoud, Amir Mohammadi, and Elahe Yazdinasabkarani. Earthquake damage estimation of the Koyna concrete gravity dam using explicit GFVM model considering fluid– structure interaction. Mathematics and Computers in Simulation, 195: 151–170, May 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000027.

Pati:2022:DCC

[802] N. C. Pati and Bapan Ghosh. Delayed carrying capacity induced subcritical and supercritical Hopf bifurcations in a predator–prey system. Mathematics and Computers in Simulation, 195:171–196, May 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000088.

Sahu:2022:MRC

[803] Bandita Sahu, Pradipta Kumar Das, and Manas Ranjan Kabat. Multirobot co-operation for stick carrying application using hybridization of meta-heuristic algorithm. Mathematics and Computers in Simulation, 195:197–226, May 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000106.

Anonymous:2022:NIe

[804] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 195:227, May 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000490.

Anonymous:2022:ICEe

[805] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 195:228–231, May 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475422000507.

Anonymous:2022:EBe

[806] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 195:ifc, May 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200043X.

Tassaddiq:2022:GEC

[807] Asifa Tassaddiq. General escape criteria for the generation of fractals in extended Jungck–Noor orbit. Mathematics and Computers in Simulation, 196:1–14, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422000039.

Anonymous:2022:PJb

[808] Anonymous. Pages 1–358 (June 2022). Mathematics and Computers in Simulation, 196:1–358, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Zhou:2022:TDD

[809] Baoquan Zhou, Daqing Jiang, Bingtao Han, and Tasawar Hayat. Threshold dynamics and density function of a stochastic epidemic model with media coverage and mean-reverting Ornstein–Uhlenbeck process. Mathematics and Computers in Simulation, 196:15–44, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000246.

Zhang:2022:GFD

[810] Juan Zhang, Rong-Juin Shuy, Chiung-Lin Chu, and Chia-Ming Fan. Generalized finite difference method for three-dimensional eigenproblems of Helmholtz equation. Mathematics and Computers in Simulation, 196: 45–67, June 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000076.

Tan:2022:NAS

[811] Zengqiang Tan and Chengjian Zhang. Numerical approximation to semilinear stiff neutral equations via implicit–explicit general linear methods. Mathematics and Computers in Simulation, 196:68–87, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475422000234.

Naz:2022:DNC

[812] Sidra Naz, Muhammad Asif Zahoor Raja, Aneela Kausar, Aneela Zameer, Ammara Mehmood, and Muhammad Shoaib. Dynamics of nonlinear cantilever piezoelectric–mechanical system: an intelligent computational approach. Mathematics and Computers in Simulation, 196:88– 113, June 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000210.

Mouktonglang:2022:DBS

[813] Thanasak Mouktonglang, Suriyon Yimnet, Nattakorn Sukantamala, and Ben Wongsaijai. Dynamical behaviors of the solution to a periodic initial–boundary value problem of the generalized Rosenau-RLW-Burgers equation. Mathematics and Computers in Simulation, 196:114– 136, June 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422000040.

Zhao:2022:EEP

[814] Longbin Zhao and Pengde Wang. Error estimates of piecewise Hermite collocation method for highly oscillatory Volterra integral equation with Bessel kernel. Mathematics and Computers in Simulation, 196: 137–150, June 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422000271.

Zhang:2022:HOI

[815] Xu Zhang, Yanqun Jiang, Yinggang Hu, and Xun Chen. Highorder implicit weighted compact nonlinear scheme for nonlinear coupled viscous Burgers’ equations. Mathematics and Computers in Simulation, 196:151–165, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200009X.

Mondal:2022:DDT

[816] Sudeshna Mondal and Guruprasad Samanta. Dynamics of a delayed toxin producing plankton model with variable search rate of zooplankton. Mathematics and Computers in Simulation, 196:166–191, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000350.

Yu:2022:HBB

[817] Fei Yu and Yuanshi Wang. Hopf bifurcation and Bautin bifurcation in a prey–predator model with prey’s fear cost and variable predator search speed. Mathematics and Computers in Simulation, 196:192– 209, June 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000398.

Li:2022:ALT

[818] Zhiqiang Li. Asymptotics and large time behaviors of fractional evolution equations with temporal ψ-Caputo derivative. Mathematics and Computers in Simulation, 196:210– 231, June 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000362.

Liu:2022:MASb

[819] Yang Liu, Zhen Wang, and Xia Huang. Multistability analysis of statedependent switched Hopfield neural networks with the Gaussian-wavelet-type activation function. Mathematics and Computers in Simulation, 196:232–250, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000349.

Zirkohi:2022:RAB

[820] Majid Moradi Zirkohi. Robust adaptive backstepping control of uncertain fractional-order nonlinear systems with input time delay. Mathematics and Computers in Simulation, 196:251–272, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000337.

Zorig:2022:CSS

[821] Assam Zorig, Shahin Hedayati Kia, Aissa Chouder, and Abdelhamid Rabhi. A comparative study for stator winding inter-turn short-circuit fault detection based on harmonic analysis of induction machine signatures. Mathematics and Computers in Simulation, 196:273–288, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000325.

Zaky:2022:NCC

[822] M. A. Zaky, A. S. Hendy, and D. Suragan. A note on a class of Caputo fractional differential equations with respect to another function. Mathematics and Computers in Simulation, 196:289–295, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000283.

Dubey:2022:EAS

[823] Ved Prakash Dubey, Jagdev Singh, Ahmed M. Alshehri, Sarvesh Dubey, and Devendra Kumar. An efficient analytical scheme with convergence analysis for computational study of local fractional Schrödinger equations. Mathematics and Computers in Simulation, 196:296–318, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475422000222.

Kumar:2022:SRG

[824] Sachin Kumar, Shubham Kumar Dhiman, and Astha Chauhan. Symmetry reductions, generalized solutions and dynamics of wave profiles for the (2+1)- dimensional system of Broer–Kaup– Kupershmidt (BKK) equations. Mathematics and Computers in Simulation, 196:319–335, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000374.

Banda:2022:PFH

[825] Heather Banda, Michael Chapwanya, and Phindile Dumani. Pattern formation in the Holling–Tanner predator–prey model with predatortaxis. A nonstandard finite difference approach. Mathematics and Computers in Simulation, 196:336– 353, June 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422000416.

Anonymous:2022:NIf

[826] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 196:354, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422000982.

Anonymous:2022:ICEf

[827] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 196:355–358, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475422000994.

Anonymous:2022:EBf

[828] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 196:ifc, June 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422000933.

Gupta:2022:EBM

[829] R. P. Gupta and Arun Kumar. Endemic bubble and multiple cusps generated by saturated treatment of an SIR model through Hopf and Bogdanov–Takens bifurcations. Mathematics and Computers in Simulation, 197:1–21, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422000386.

Anonymous:2022:PJc

[830] Anonymous. Pages 1–418 (July 2022). Mathematics and Computers in Simulation, 197:1–418, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Lou:2022:BEF

[831] Yu Lou and Yi Zhang. Breathers on elliptic function background for a generalized nonlinear Schrödinger equation with higher-order terms. Mathematics and Computers in Simulation, 197:22–31, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000544.

Cayan:2022:CAM

[832] Seda Çayan, B. Burak Özhan, and Mehmet Sezer. Collocation approaches to the mathematical model of an Euler– Bernoulli beam vibrations. Mathematics and Computers in Simulation, 197:32–44, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000404.

Dutta:2022:DCD

[833] Protyusha Dutta, Debgopal Sahoo, Sudeshna Mondal, and Guruprasad Samanta. Dynamical complexity of a delay-induced eco-epidemic model with Beddington–DeAngelis incidence rate. Mathematics and Computers in Simulation, 197:45–90, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000556.

Acosta-Gonzalez:2022:ECE

[834] Eduardo Acosta-González, Julián Andrada-Félix, and Fernando Fernández-Rodríguez. On the evolution of the COVID-19 epidemiological parameters using only the series of deceased. A study of the Spanish outbreak using Genetic Algorithms. Mathematics and Computers in Simulation, 197:91–104, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200060X.

Occorsio:2022:LCI

[835] Donatella Occorsio, Giuliana Ramella, and Woula Themistoclakis. Lagrange– Chebyshev Interpolation for image resizing. Mathematics and Computers in Simulation, 197:105– 126, July 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422000295.

Karakan:2022:PAM

[836] İsmet Karakan, Ceren Gürkan, and Cem Avcı. Performance analyses of mesh-based local Finite Element Method and meshless global RBF Collocation Method for solving Poisson and Stokes equations. Mathematics and Computers in Simulation, 197:127–150, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422000684.

urRahman:2022:EGS

[837] Mujeeb ur Rahman, Tasawar Hayat, Sohail A. Khan, and A. Alsaedi. Entropy generation in Sutterby nanomaterials flow due to rotating disk with radiation and magnetic effects. Mathematics and Computers in Simulation, 197:151–165, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000647. See

corrigendum [1750].

Abdullah:2022:ECM

[838] Muhammad Abdullah, Muhammad Yaseen, and Manuel De la Sen. An efficient collocation method based on Hermite formula and cubic Bsplines for numerical solution of the Burgers’ equation. Mathematics and Computers in Simulation, 197:166– 184, July 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000660.

Avijit:2022:CAF

[839] D. Avijit and S. Natesan. Convergence analysis of a fully-discrete FEM for singularly perturbed twoparameter parabolic PDE. Mathematics and Computers in Simulation, 197:185–206, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000581.

Hu:2022:ECO

[840] Gang Hu, Wanting Dou, Xiaofeng Wang, and Muhammad Abbas. An enhanced chimp optimization algorithm for optimal degree reduction of Said–Ball curves. Mathematics and Computers in Simulation, 197: 207–252, July 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000313.

Jelti:2022:NMT

[841] S. Jelti and M. Boulerhcha. Numerical modeling of two dimensional noncapacity model for sediment transport by an unstructured finite volume method with a new discretization of the source term. Mathematics and Computers in Simulation, 197:253–276, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000659.

Bulut:2022:HOH

[842] Fatih Bulut, Ömer Oruç, and Alaattin Esen. Higher order Haar wavelet method integrated with strang splitting for solving regularized long wave equation. Mathematics and Computers in Simulation, 197:277– 290, July 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000593.

Zhu:2022:CPA

[843] Jingcan Zhu, Jie Jin, Weijie Chen, and Jianqiang Gong. A combined power activation function based convergent factor-variable ZNN model for solving dynamic matrix inversion. Mathematics and Computers in Simulation, 197:291–307, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000714.

Kerr:2022:ALT

[844] Gilbert Kerr and Gilberto González-Parra. Accuracy of the Laplace transform method for linear neutral delay differential equations. Mathematics and Computers in Simulation, 197:308–326, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000702.

He:2022:LSS

[845] Xue-Jiao He and Xing Lü. M-lump solution, soliton solution and rational solution to a (3 + 1)-dimensional nonlinear model. Mathematics and Computers in Simulation, 197:327– 340, July 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000672.

Zhang:2022:HCD

[846] Hai Zhang, Yuhong Cheng, Hongmei Zhang, Weiwei Zhang, and Jinde Cao. Hybrid control design for Mittag-Leffler projective synchronization on FOQVNNs with multiple mixed delays and impulsive effects. Mathematics and Computers in Simulation, 197:341–357, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422000751.

Filelis-Papadopoulos:2022:MFB

[847] Christos K. Filelis-Papadopoulos, Panagiotis E. Kyziropoulos, John P. Morrison, and Philip O’Reilly. Modelling and forecasting based on recursive incomplete pseudoinverse matrices. Mathematics and Computers in Simulation, 197:358–376, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475422000738.

Zhao:2022:DMC

[848] Ying Zhao, Liang Li, Stéphane Lanteri, and Jonathan Viquerat. Dynamic metasurface control using Deep Reinforcement Learning. Mathematics and Computers in Simulation, 197: 377–395, July 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422000696.

Dassios:2022:SDS

[849] Ioannis Dassios, Georgios Tzounas, Muyang Liu, and Federico Milano. Singular over-determined systems of linear differential equations. Mathematics and Computers in Simulation, 197:396–412, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422000568.

Anonymous:2022:NIg

[850] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 197:413, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001343.

Anonymous:2022:ICEg

[851] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 197:414–417, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475422001355.

Anonymous:2022:EBg

[852] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 197:ifc, July 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200129X.

Song:2022:SSM

[853] Xingxing Song, Hongqian Lu, Yao Xu, and Wuneng Zhou. H_{∞} synchronization of semi-Markovian jump neural networks with random sensor nonlinearities via adaptive event-triggered output feedback control. Mathematics and Computers in Simulation, 198:1–19, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000611.

Anonymous:2022:PAb

[854] Anonymous. Pages 1–526 (August 2022). Mathematics and Computers in Simulation, 198:1–526, August 2022.

CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Pfeil:2022:LMP

[855] Tamás Pfeil and Blanka Herbály. A linear model for polyclonal antibody– antigen reactions. Mathematics and Computers in Simulation, 198:20– 30, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542200057X.

Chen:2022:SMU

[856] Ziren Chen, Lin Feng, Harold A. Lay, Khaled Furati, and Abdul Khaliq. SEIR model with unreported infected population and dynamic parameters for the spread of COVID-19. Mathematics and Computers in Simulation, 198:31–46, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000787.

Waku:2022:MRN

[857] J. Waku, K. Oshinubi, and J. Demongeot. Maximal reproduction number estimation and identification of transmission rate from the first inflection point of new infectious cases waves: COVID-19 outbreak example. Mathematics and Computers in Simulation, 198:47–64, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000763.

Sintunavarat:2022:MAE

[858] Wutiphol Sintunavarat and Ali Turab. Mathematical analysis of an extended SEIR model of COVID-19 using the ABC-fractional operator. Mathematics and Computers in Simulation, 198:65–84, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000623.

Cui:2022:CTR

[859] Xiongwei Cui, Zhikai Wang, Xiongliang Yao, Minghao Liu, and Fulin Yu. A coupled two-relaxation-time Lattice Boltzmann–Volume penalization method for flows past obstacles. Mathematics and Computers in Simulation, 198:85–105, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421004596.

Jiwari:2022:LRB

[860] Ram Jiwari. Local radial basis function-finite difference based algorithms for singularly perturbed Burgers’ model. Mathematics and Computers in Simulation, 198:106– 126, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000775.

Arshad:2022:MDD

[861] Muhammad Arshad, Rukhsana Jabeen, and Suliman Khan. A multiscale domain decomposition approach for parabolic equations using expanded mixed method. Mathematics and Computers in Simulation, 198:127– 150, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000635.

Wang:2022:ADR

[862] Jinliang Wang, Wenjing Wu, and Toshikazu Kuniya. Analysis of a degenerated reaction–diffusion cholera model with spatial heterogeneity and stabilized total humans. Mathematics and Computers in Simulation, 198:151–171, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000799.

Kim:2022:SFC

[863] SungBum Kim and Hyoung-Moon Kim. Series form of the characteristic functions of scale mixtures of multivariate skew-normal distributions. Mathematics and Computers in Simulation, 198:172–187, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000866.

Zhu:2022:DAI

[864] Hegui Zhu, Jiangxia Ge, Wentao Qi, Xiangde Zhang, and Xiaoxiong Lu. Dynamic analysis and image encryption application of a sinusoidal-polynomial composite chaotic system. Mathematics and Computers in Simulation, 198:188–210, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000829.

Takacs:2022:HOD

[865] Bálint Takács and Yiannis Hadjimichael. High order discretization methods for spatial-dependent epidemic models. Mathematics and Computers in Simulation, 198:211– 236, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542200074X.

Kumar:2022:GDA

[866] Manoj Kumar and Syed Abbas. Global dynamics of an age-structured model for HIV viral dynamics with latently infected T cells. Mathematics and Computers in Simulation, 198:237– 252, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542200088X.

Ornek:2022:NVS

[867] Bülent Nafi Örnek, Salih Berkan Aydemir, Timur Düzenli, and Bilal Özak. A novel version of slime mould algorithm for global optimization and real world engineering problems: Enhanced slime mould algorithm. Mathematics and Computers in Simulation, 198:253–288, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000830.

Dai:2022:BMV

[868] Qinrui Dai, Mengjie Rong, and Ren Zhang. Bifurcations and multistability in a virotherapy model with two time delays. Mathematics and Computers in Simulation, 198:289– 311, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422000817.

Muroi:2022:BTM

[869] Yoshifumi Muroi and Shintaro Suda. Binomial tree method for option pricing: Discrete cosine transform approach. Mathematics and Computers in Simulation, 198:312–331, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475422000854.

Behera:2022:TDW

[870] S. Behera and S. Saha Ray. Twodimensional wavelets scheme for numerical solutions of linear and nonlinear Volterra integro-differential equations. Mathematics and Computers in Simulation, 198:332–358, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475422000726.

Milosevic:2022:SSM

[871] Marija Milošević. Stochastic serotonin model with discontinuous drift. Mathematics and Computers in Simulation, 198:359–374, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422000891.

Visakamoorthi:2022:FSD

[872] B. Visakamoorthi and P. Muthukumar. Fuzzy sampled-data control for single-master multi-slave teleoperation systems with stochastic actuator faults. Mathematics and Computers in Simulation, 198:375– 387, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001045.

Rajagopal:2022:NHM

[873] Aghila Rajagopal, Sudan Jha, Ramachandran Alagarsamy, Shio Gai Quek, and Ganeshsree Selvachandran. A novel hybrid machine learning framework for the prediction of diabetes with context-customized regularization and prediction procedures. Mathematics and Computers in Simulation, 198:388–406, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200091X.

El-Sabaa:2022:MOD

[874] F. M. El-Sabaa, T. S. Amer, A. A. Sallam, and I. M. Abady. Modeling of the optimal deceleration for the rotatory motion of asymmetric rigid body. Mathematics and Computers in Simulation, 198:407– 425, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000908.

Castillo-Alvino:2022:GDP

[875] Hamlet Humberto Castillo-Alvino and Marcos Marvá. Group defense promotes coexistence in interference competition: the Holling type IV competitive response. Mathematics and Computers in Simulation, 198:426– 445, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422000842.

Wu:2022:WGC

[876] Jianping Wu. Wronskian and Grammian conditions, and Pfaffianization of an extended (3 + 1)-dimensional Jimbo–Miwa equation. Mathematics and Computers in Simulation, 198:446–454, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422000805.

Khan:2022:CSF

[877] Hasib Khan, Khurshaid Alam, Haseena Gulzar, Sina Etemad, and Shahram Rezapour. A case study of fractalfractional tuberculosis model in China: Existence and stability theories along with numerical simulations. Mathematics and Computers in Simulation, 198:455–473, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001082.

Su:2022:FTC

[878] Haipeng Su, Runzi Luo, Jiaojiao Fu, and Meichun Huang. Fixed time control and synchronization of a class of uncertain chaotic systems with disturbances via passive control method. Mathematics and Computers in Simulation, 198:474– 493, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001094.

Li:2022:BFN

[879] Liu-Qing Li, Yi-Tian Gao, Xin Yu, Cui-Cui Ding, and Dong Wang. Bilinear form and nonlinear waves of a (1+1)-dimensional generalized Boussinesq equation for the gravity waves over water surface. Mathematics and Computers in Simulation, 198:494– 508, August 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422000878.

Dubey:2022:AME

[880] Shweta Dubey and S. Chakraverty. Application of modified extended tanh method in solving fractional order coupled wave equations. Mathematics and Computers in Simulation, 198:509–520, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001069.

Anonymous:2022:NIh

[881] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 198:521, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001872.

Anonymous:2022:ICEh

[882] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 198:522–525, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475422001884.

Anonymous:2022:EBh

[883] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 198:ifc, August 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001823.

Sinhababu:2022:PSB

[884] Arijit Sinhababu and Anirban Bhattacharya. A pseudo-spectral based efficient volume penalization scheme for Cahn–Hilliard equation in complex geometries. Mathematics and Computers in Simulation, 199:1–24, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422001148.

Anonymous:2022:PS

[885] Anonymous. Pages 1–468 (September 2022). Mathematics and Computers in Simulation, 199:1–468, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Taher:2022:ENT

[886] Anis Haytham Saleh Taher. An efficient numerical technique for estimating eigenvalues of second-order non-self-adjoint Sturm–Liouville problems. Mathematics and Computers in Simulation, 199:25–37, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001136.

Wang:2022:TTG

[887] Furong Wang, Xuehua Yang, Haixiang Zhang, and Lijiao Wu. A time two-grid algorithm for the two dimensional nonlinear fractional PIDE with a weakly singular kernel. Mathematics and Computers in Simulation, 199:38–59, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001033.

Fahimi-khalilabad:2022:HOF

[888] Iraj Fahimi-khalilabad, Safar Irandoustpakchin, and Somayeh Abdi-mazraeh. High-order finite difference method based on linear barycentric rational interpolation for Caputo type sub-diffusion equation. Mathematics and Computers in Simulation, 199: 60–80, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001070.

Abouelregal:2022:VAE

[889] Ahmed E. Abouelregal, Fawzy A. Mohammed, Moez Benhamed, Adam Zakria, and Ibrahim-Elkhalil Ahmed. Vibrations of axially excited rotating micro-beams heated by a highintensity laser in light of a thermoelastic model including the memorydependent derivative. Mathematics and Computers in Simulation, 199: 81–99, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001197.

Bai:2022:ROM

[890] Feng Bai and Yi Wang. A reduced order modeling method based on GNAT-embedded hybrid snapshot simulation. Mathematics and Computers in Simulation, 199:100– 132, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001057.

Tafakkori–Bafghi:2022:NST

[891] M. Tafakkori-Bafghi, G. B. Loghmani, and M. Heydari. Numerical solution of two-point nonlinear boundary value problems via Legendre– Picard iteration method. Mathematics and Computers in Simulation, 199: 133–159, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001240.

El-Amrani:2022:SLB

[892] Mofdi El-Amrani, Bassou Khouya, and Mohammed Seaid. A semi-Lagrangian Bernstein–Bézier finite element method for two-dimensional coupled Burgers’ equations at high Reynolds numbers. Mathematics and Computers in Simulation, 199:160– 181, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001100.

Peng:2022:PFE

[893] Q. Peng and F. J. Vermolen. Point forces in elasticity equation and their alternatives in multi dimensions. Mathematics and Computers in Simulation, 199:182–201, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001239.

Lam:2022:PIP

[894] Nicholas N. Lam, Paul D. Docherty, and Rua Murray. Practical identifiability of parametrised models: a review of benefits and limitations of various approaches. Mathematics and Computers in Simulation, 199: 202–216, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001227.

Tocino:2022:ISU

[895] A. Tocino, Y. Komori, and T. Mitsui. Integration of the stochastic underdamped harmonic oscillator by the θ-method. Mathematics and Computers in Simulation, 199: 217–230, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001112.

daSilva:2022:RRT

[896] L. P. da Silva, C. H. Marchi, M. Meneguette, and A. C. Foltran. Robust RRE technique for increasing the order of accuracy of SPH numerical solutions. Mathematics and Computers in Simulation, 199: 231–252, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001185.

Liseikin:2022:CNA

[897] V. D. Liseikin. Comments on a numerical algorithm for problems having interior power-of-type-2 layers. Mathematics and Computers in Simulation, 199:253–256, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001264.

Inage:2022:AMC

[898] Sin ichi Inage and Hana Hebishima. Application of Monte Carlo stochastic optimization (MOST) to deep learning. Mathematics and Computers in Simulation, 199:257–271, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422001124.

Zguaid:2022:RBO

[899] Khalid Zguaid and Fatima-Zahrae El Alaoui. Regional boundary observability for Riemann–Liouville linear fractional evolution systems. Mathematics and Computers in Simulation, 199: 272–286, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001252.

Kumar:2022:HOC

[900] Sunil Kumar, Sumit, and Jesus Vigo-Aguiar. A high order convergent numerical method for singularly perturbed time dependent problems using mesh equidistribution. Mathematics and Computers in Simulation, 199: 287–306, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001276.

Zhu:2022:PES

[901] Dejun Zhu. Practical exponential stability of stochastic delayed systems with G-Brownian motion via vector G-Lyapunov function. Mathematics and Computers in Simulation, 199: 307–316, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001409.

Albi:2022:EES

[902] Giacomo Albi, Roberto Chignola, and Federica Ferrarese. Efficient ensemble stochastic algorithms for agentbased models with spatial predator– prey dynamics. Mathematics and Computers in Simulation, 199:317– 340, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422001215.

Li:2022:ACM

[903] Yuping Li, Zhanwen Yang, and Hui Liang. Analysis of collocation methods for a class of thirdkind auto-convolution Volterra integral equations. Mathematics and Computers in Simulation, 199:341– 358, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422001380.

Hoang:2022:PBP

[904] Manh Tuan Hoang. Positivity and boundedness preserving nonstandard finite difference schemes for solving Volterra’s population growth model. Mathematics and Computers in Simulation, 199:359–373, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422001410.

Tanveer:2022:DWB

[905] Anum Tanveer, Mair Khan, T. Salahuddin, Basem Al Alwan, and Abdelfattah Amari. Dynamics of Walters’ B fluid due to periodic wave in a convectively heated channel with internal heat generation. Mathematics and Computers in Simulation, 199: 374–393, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001203.

Nikan:2022:LSR

[906] O. Nikan and Z. Avazzadeh. A locally stabilized radial basis function partition of unity technique for the sine–Gordon system in nonlinear optics. Mathematics and Computers in Simulation, 199:394–413, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001483.

Wang:2022:OOP

[907] Qiang Wang, Jie Wu, and Nenggui Zhao. Optimal operational policies of a dual-channel supply chain considering return service. Mathematics and Computers in Simulation, 199: 414–437, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001513.

Alam:2022:ABF

[908] Jahangir Alam, M. G. Murtaza, E. E. Tzirtzilakis, and M. Ferdows. Application of Biomagnetic Fluid Dynamics modeling for simulation of flow with magnetic particles and variable fluid properties over a stretching cylinder. Mathematics and Computers in Simulation, 199:438–462, September 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001501.

Anonymous:2022:NIi

[909] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 199:463, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002166.

Anonymous:2022:EBi

[910] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 199:464–467, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002178.

Anonymous:2022:EBj

[911] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 199:ifc, September 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002117.

Bandekar:2022:CIM

[912] Shraddha Ramdas Bandekar and Mini Ghosh. A co-infection model on TB — COVID-19 with optimal control and sensitivity analysis. Mathematics and Computers in Simulation, 200:1–31, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001392.

Anonymous:2022:PO

[913] Anonymous. Pages 1–608 (October 2022). Mathematics and Computers in Simulation, 200:1–608, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Rana:2022:SDL

[914] Sourav Rana, Sabyasachi Bhattacharya, and Sudip Samanta. Spatiotemporal dynamics of Leslie–Gower predator–prey model with Allee effect on both populations. Mathematics and Computers in Simulation, 200:32–49, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001537.

Jyoti:2022:ITB

[915] Jyoti and Mandeep Singh. An iterative technique based on HPM for a class of one dimensional Bratu’s type problem. Mathematics and Computers in Simulation, 200:50– 64, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001495.

Shao:2022:GSD

[916] Yuanfu Shao. Global stability of a delayed predator–prey system with fear and Holling-type II functional response in deterministic and stochastic environments. Mathematics and Computers in Simulation, 200:65– 77, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001550.

Kalpana:2022:NSC

[917] G. Kalpana, K. R. Madhura, and Ramesh B. Kudenatti. Numerical study on the combined effects of Brownian motion and thermophoresis on an unsteady magnetohydrodynamics nanofluid boundary layer flow. Mathematics and Computers in Simulation, 200:78–96, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001525.

Bakhshandeh-Chamazkoti:2022:LSR

[918] Rohollah Bakhshandeh-Chamazkoti and Mohsen Alipour. Lie symmetries reduction and spectral methods on the fractional two-dimensional heat equation. Mathematics and Computers in Simulation, 200:97– 107, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001574.

Azam:2022:SDD

[919] Anam Azam, Muhammad Aqeel, and Danish Ali Sunny. Segmented disc dynamo with symmetric multidirectional patterns of multiscroll chaotic attractors. Mathematics and Computers in Simulation, 200:108– 127, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001434.

Zhou:2022:NSF

[920] Wenhao Zhou, Hailin Li, and Zhiwei Zhang. A novel seasonal fractional grey model for predicting electricity demand: a case study of Zhejiang in China. Mathematics and Computers in Simulation, 200:128– 147, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001422.

Tan:2022:NSM

[921] Ting Tan, Lan Li, and Jing An. A novel spectral method and error analysis for fourth-order equations in a spherical region. Mathematics and Computers in Simulation, 200:148–161, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001598.

Bai:2022:BBS

[922] Ning Bai and Rui Xu. Backward bifurcation and stability analysis in a within-host HIV model with both virus-to-cell infection and cell-to-cell transmission, and antiretroviral therapy. Mathematics and Computers in Simulation, 200:162– 185, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001628.

Pathak:2022:NSG

[923] Maheshwar Pathak, Pratibha Joshi, and Kottakkaran Sooppy Nisar. Numerical study of generalized 2-D nonlinear Schrödinger equation using Kansa method. Mathematics and Computers in Simulation, 200: 186–198, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001720.

Lamboni:2022:EDM

[924] Matieyendou Lamboni. Efficient dependency models: Simulating dependent random variables. Mathematics and Computers in Simulation, 200:199–217, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.math.utah.edu/
                                        pub/tex/bib/prng.bib;   https:/
                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422001604.

Omran:2022:EIL

[925] A. K. Omran, M. A. Zaky, A. S. Hendy, and V. G. Pimenov. An easy to implement linearized numerical scheme for fractional reaction– diffusion equations with a prehistorical nonlinear source function. Mathematics and Computers in Simulation, 200:218–239, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001562.

Dong:2022:DPS

[926] Gracia Y. Dong and Christiane Lemieux. Dependence properties of scrambled Halton sequences. Mathematics and Computers in Simulation, 200:240–262, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001586.

Kayenat:2022:CDF

[927] Sheerin Kayenat and Amit Kumar Verma. On the choice of denominator functions and convergence of NSFD schemes for a class of nonlinear SBVPs. Mathematics and Computers in Simulation, 200:263– 284, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001616.

Saha:2022:IOV

[928] Sangeeta Saha, Guruprasad Samanta, and Juan J. Nieto. Impact of optimal vaccination and social distancing on COVID-19 pandemic. Mathematics and Computers in Simulation, 200:285–314, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001677.

Meena:2022:PCC

[929] Rakesh Kumar Meena, Madhu Jain, Assif Assad, Rachita Sethi, and Deepika Garg. Performance and cost comparative analysis for M/G/1 repairable machining system with Npolicy vacation. Mathematics and Computers in Simulation, 200:315– 328, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001549.

Wei:2022:MFA

[930] Wenqi Wei, Haibin Ouyang, Steven Li, Xuanbo Zhao, and Dexuan Zou. A modified fireworks algorithm with dynamic search interval based on closed-loop control. Mathematics and Computers in Simulation, 200:329– 360, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422001689.

Cui:2022:SBL

[931] Yuhuan Cui, Jingguo Qu, Cundi Han, Gang Cheng, Wei Zhang, and Yiming Chen. Shifted Bernstein– Legendre polynomial collocation algorithm for numerical analysis of viscoelastic Euler–Bernoulli beam with variable order fractional model. Mathematics and Computers in Simulation, 200:361–376, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200177X.

Shahid:2022:NCM

[932] A. Shahid, H. L. Huang, M. M. Bhatti, and M. Marin. Numerical computation of magnetized bioconvection nanofluid flow with temperature-dependent viscosity and Arrhenius kinetic. Mathematics and Computers in Simulation, 200:377–392, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001744.

Zhang:2022:UCW

[933] Jin Zhang and Xiaowei Liu. Uniform convergence of a weak Galerkin method for singularly perturbed convection– diffusion problems. Mathematics and Computers in Simulation, 200:393– 403, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001653.

Nikooeinejad:2022:NIM

[934] Z. Nikooeinejad, M. Heydari, and G. B. Loghmani. A numerical iterative method for solving two-point BVPs in infinite-horizon nonzerosum differential games: Economic applications. Mathematics and Computers in Simulation, 200:404– 427, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001641.

Hu:2022:QIB

[935] Gang Hu, Bo Du, Huinan Li, and Xupeng Wang. Quadratic interpolation boosted black widow spiderinspired optimization algorithm with wavelet mutation. Mathematics and Computers in Simulation, 200:428– 467, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001732.

Herty:2022:MCS

[936] Michael Herty, Sonja Steffensen, and Anna Thünen. Multiscale control of Stackelberg games. Mathematics and Computers in Simulation, 200: 468–488, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001707.

Meng:2022:ASE

[937] Lan Meng and Wei Zhu. Analysis of SEIR epidemic patch model with nonlinear incidence rate, vaccination and quarantine strategies. Mathematics and Computers in Simulation, 200:489–503, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001690.

Zhang:2022:LMF

[938] Junli Zhang, Chenchen Yang, Hui Zheng, Chia-Ming Fan, and Ming-Fu Fu. The localized method of fundamental solutions for 2D and 3D inhomogeneous problems. Mathematics and Computers in Simulation, 200: 504–524, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001665.

Wang:2022:DAI

[939] Ning Wang, Longxing Qi, and Guangyi Cheng. Dynamical analysis for the impact of asymptomatic infective and infection delay on disease transmission. Mathematics and Computers in Simulation, 200:525– 556, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001719.

Patra:2022:SFE

[940] A. Patra, P. Baliarsingh, and H. Dutta. Solution to fractional evolution equation using Mohand transform. Mathematics and Computers in Simulation, 200:557–570, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200163X.

Nasir:2022:DDP

[941] Hanis Nasir. On the dynamics of a diabetic population model with two delays and a general recovery rate of complications. Mathematics and Computers in Simulation, 200: 571–602, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001768.

Anonymous:2022:NIj

[942] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 200:603, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002373.

Anonymous:2022:ICEi

[943] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 200:604–607, October 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422002385.

Anonymous:2022:EBk

[944] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 200:ifc, October 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002324.

Guimaraes:2022:SOB

[945] O. Guimarães, W. Labecca, and José R. C. Piqueira. Solving 2nd order BVPs in planar irregular domains. Mathematics and Computers in Simulation, 201:1–22, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001938.

Anonymous:2022:PN

[946] Anonymous. Pages 1–774 (November 2022). Mathematics and Computers in Simulation, 201:1–774, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liu:2022:REM

[947] Jianxun Liu, Jinfei Shi, Fei Hao, and Min Dai. A reinforced exploration mechanism whale optimization algorithm for continuous optimization problems. Mathematics and Computers in Simulation, 201: 23–48, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001756.

Krishna:2022:PBH

[948] S. Gopi Krishna, M. Shanmugapriya, and P. Senthil Kumar. Prediction of bio-heat and mass transportation in radiative MHD Walter-B nanofluid using MANFIS model. Mathematics and Computers in Simulation, 201: 49–67, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001793.

Wibowo:2022:CLE

[949] Ferry Wahyu Wibowo, Eko Sediyono, and Hindriyanto Dwi Purnomo. Chimpanzee leader election optimization. Mathematics and Computers in Simulation, 201:68–95, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422001951.

Loheac:2022:TOC

[950] Jérôme Lohéac, Vineeth Satheeskumar Varma, and Irinel Constantin Morărescu. Time optimal control for a mobile robot with a communication objective. Mathematics and Computers in Simulation, 201: 96–120, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002014.

Rana:2022:FSQ

[951] Puneet Rana and Gaurav Gupta. FEM solution to quadratic convective and radiative flow of Ag–MgO/H_{2}O hybrid nanofluid over a rotating cone with Hall current: Optimization using Response Surface Methodology. Mathematics and Computers in Simulation, 201:121–140, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002002.

Dong:2022:LPT

[952] Cheng Dong and Tong Kang. A linearity-preserving technique for finite volume schemes of anisotropic diffusion problems on polygonal meshes. Mathematics and Computers in Simulation, 201:141–162, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001999.

Yan:2022:SPB

[953] Xiao Yan, Yimamu Maimaiti, and Wenbin Yang. Stationary pattern and bifurcation of a Leslie–Gower predator– prey model with prey-taxis. Mathematics and Computers in Simulation, 201:163–192, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422001987.

Raja:2022:INC

[954] M. Asif Zahoor Raja, M. Shoaib, Ghania Zubair, M. Ijaz Khan, R. J. Punith Gowda, B. C. Prasannakumara, and Kamel Guedri. Intelligent neurocomputing for entropy generated Darcy–Forchheimer mixed convective fluid flow. Mathematics and Computers in Simulation, 201:193– 214, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422001926.

VieiraHollweg:2022:DCT

[955] Guilherme Vieira Hollweg, Paulo Jefferson Dias de Oliveira Evald, Deise Maria Cirolini Milbradt, Rodrigo Varella Tambara, and Hilton Abílio Gründling. Design of continuous-time model reference adaptive and super-twisting sliding mode controller. Mathematics and Computers in Simulation, 201: 215–238, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002026.

Behnood:2022:LSE

[956] Maryam Behnood and Ali Shokri. A Legendre spectral element method for the family of regularized long wave equations. Mathematics and Computers in Simulation, 201:239– 253, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422002075.

Kumar:2022:DBA

[957] Sachin Kumar and Amit Kumar. Dynamical behaviors and abundant optical soliton solutions of the cold bosonic atoms in a zig-zag optical lattice model using two integral schemes. Mathematics and Computers in Simulation, 201:254–274, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001975.

Malinen:2022:LRA

[958] Mika Malinen. Local reparametrization by approximating lines of curvature. Mathematics and Computers in Simulation, 201:275–290, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542200204X.

Maity:2022:DCS

[959] S. Maity, R. Krishanan, S. K. Singh, B. S. Dandapat, and D. Sen. Development of CNTs suspended thin nanoliquid film over a nonlinear stretching sheet. Mathematics and Computers in Simulation, 201:291– 304, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422001781.

Li:2022:DBC

[960] Hui Li, Ruichun He, and Wei Zhou. Dynamic behaviors in a Cournot duopoly model with knowledge spillover effect based on constant conjectural variation. Mathematics and Computers in Simulation, 201:305–323, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542200194X.

AlAli:2022:EDP

[961] Hannah Al Ali, Alireza Daneshkhah, Abdesslam Boutayeb, Noble Jahalamajaha Malunguza, and Zindoga Mukandavire. Exploring dynamical properties of a Type 1 diabetes model using sensitivity approaches. Mathematics and Computers in Simulation, 201: 324–342, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422001963.

Calis:2022:HBF

[962] Yasemin Çalış, Ali Demirci, and Cihangir Özemir. Hopf bifurcation of a financial dynamical system with delay. Mathematics and Computers in Simulation, 201:343– 361, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542200180X.

Eslami:2022:AAM

[963] N. Eslami, S. Yazdani, M. Mirzaei, and E. Hadavandi. Aphid–ant Mutualism: a novel nature-inspired metaheuristic algorithm for solving optimization problems. Mathematics and Computers in Simulation, 201: 362–395, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002038.

Campos:2022:CRS

[964] Brian D. Vasquez Campos. Characterization of rational solutions of a KdV-like equation. Mathematics and Computers in Simulation, 201: 396–416, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002208.

Li:2022:DRD

[965] Yajing Li, Mengxin He, and Zhong Li. Dynamics of a ratio-dependent Leslie– Gower predator–prey model with Allee effect and fear effect. Mathematics and Computers in Simulation, 201: 417–439, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002051.

Iswarya:2022:NRE

[966] M. Iswarya, R. Raja, J. Cao, M. Niezabitowski, J. Alzabut, and C. Maharajan. New results on exponential input-to-state stability analysis of memristor based complex-valued inertial neural networks with proportional and distributed delays. Mathematics and Computers in Simulation, 201: 440–461, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000379.

Dianavinnarasi:2022:ACF

[967] J. Dianavinnarasi, R. Raja, J. Alzabut, J. Cao, M. Niezabitowski, and O. Bagdasar. Application of Caputo– Fabrizio operator to suppress the Aedes Aegypti mosquitoes via Wolbachia: an LMI approach. Mathematics and Computers in Simulation, 201:462–485, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000410.

Asifa:2022:MSB

[968] Asifa, Poom Kumam, Asifa Tassaddiq, Wiboonsak Watthayu, Zahir Shah, and Talha Anwar. Modeling and simulation based investigation of unsteady MHD radiative flow of rate type fluid; a comparative fractional analysis. Mathematics and Computers in Simulation, 201:486–507, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000446.

Rajchakit:2022:EUG

[969] G. Rajchakit, R. Sriraman, C. P. Lim, and B. Unyong. Existence, uniqueness and global stability of Clifford-valued neutral-type neural networks with time delays. Mathematics and Computers in Simulation, 201: 508–527, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000628.

Andrica:2022:PRG

[970] Dorin Andrica and Ovidiu Bagdasar. Pseudoprimality related to the generalized Lucas sequences. Mathematics and Computers in Simulation, 201: 528–542, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000677.

Cao:2022:SAD

[971] Yang Cao, Renato Colucci, and Luca Guerrini. On the stability analysis of a delayed two-stage Cournot model with R&D spillovers. Mathematics and Computers in Simulation, 201: 543–554, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421000823.

Begam:2022:CLO

[972] S. Sabeena Begam and J. Vimala. Compositions on lattice ordered multifuzzy soft matrix and its simulated application in medical diagnosis. Mathematics and Computers in Simulation, 201:555–563, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000835.

Cao:2022:MSD

[973] Yang Cao, K. Udhayakumar, K. Pradeepa Veerakumari, and R. Rakkiyappan. Memory sampled data control for switched-type neural networks and its application in image secure communications. Mathematics and Computers in Simulation, 201:564–587, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475421000963.

Nirvin:2022:ISD

[974] Prasath Nirvin, Fathalla A. Rihan, Rajan Rakkiyappan, and Chandrasekar Pradeep. Impulsive sampled-data controller design for synchronization of delayed T–S fuzzy Hindmarsh– Rose neuron model. Mathematics and Computers in Simulation, 201: 588–602, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421000975.

Xu:2022:CPE

[975] Shaosheng Xu, Jinde Cao, Zhuoxuan Li, and Wenhua Qian. Control principle and error estimation for inverse trajectory method under locating error with optimization. Mathematics and Computers in Simulation, 201: 603–615, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421001270.

urRehman:2022:TGN

[976] Habib ur Rehman, Poom Kumam, Murat Özdemir, and Ibrahim Karahan. Two generalized non-monotone explicit strongly convergent extragradient methods for solving pseudomonotone equilibrium problems and applications. Mathematics and Computers in Simulation, 201:616–639, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421001725.

Abubakar:2022:HCG

[977] Auwal Bala Abubakar, Poom Kumam, Maulana Malik, and Abdulkarim Hassan Ibrahim. A hybrid conjugate gradient based approach for solving unconstrained optimization and motion control problems. Mathematics and Computers in Simulation, 201: 640–657, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421002214.

ONeill:2022:MEM

[978] Sam O’Neill, Ovidiu Bagdasar, Stuart Berry, Nicolae Popovici, and Ramachandran Raja. Modelling equilibrium for a multi-criteria selfish routing network equilibrium flow problem. Mathematics and Computers in Simulation, 201:658–669, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542100224X.

Abubakar:2022:NHT

[979] Auwal Bala Abubakar, Poom Kumam, Abdulkarim Hassan Ibrahim, Parin Chaipunya, and Sadiya Ali Rano. New hybrid three-term spectral-conjugate gradient method for finding solutions of nonlinear monotone operator equations with applications. Mathematics and Computers in Simulation, 201: 670–683, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542100255X.

Zamart:2022:NDD

[980] Chantapish Zamart, Thongchai Botmart, Wajaree Weera, and Suphachai Charoensin. New delay-dependent conditions for finite-time extended dissipativity based non-fragile feedback control for neural networks with mixed interval time-varying delays. Mathematics and Computers in Simulation, 201:684–713, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475421002573.

Tranthi:2022:NRR

[981] Janejira Tranthi, Thongchai Botmart, Wajaree Weera, Teerapong Lainchua, and Sirada Pinjai. New results on robust exponential stability of Takagi–Sugeno fuzzy for neutral differential systems with mixed time-varying delays. Mathematics and Computers in Simulation, 201: 714–738, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475421003475.

ONeill:2022:MIL

[982] Sam O’Neill, Paul Wrigley, and Ovidiu Bagdasar. A mixed-integer linear programming formulation for the modular layout of three-dimensional connected systems. Mathematics and Computers in Simulation, 201:739– 754, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475421003487.

Vanalakshmi:2022:QSV

[983] R. Vanalakshmi, S. Maragathasundari, and K. S. Dhanalakshmi. Queuing system in VLSI physical design. Mathematics and Computers in Simulation, 201:755–768, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542100416X.

Anonymous:2022:NIk

[984] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 201:769, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002695.

Anonymous:2022:ICEj

[985] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 201:770–773, November 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002701.

Anonymous:2022:EBl

[986] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 201:ifc, November 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002646.

Martinez:2022:NDM

[987] Romeo Martínez, Jorge E. Macías-Díaz, and Qin Sheng. A nonlinear discrete model for approximating a conservative multi-fractional Zakharov system: Analysis and computational simulations. Mathematics and Computers in Simulation, 202:1–21, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002282.

Anonymous:2022:PD

[988] Anonymous. Pages 1–576 (December 2022). Mathematics and Computers in Simulation, 202:1–576, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Munoz-Vazquez:2022:GPC

[989] Aldo Jonathan Muñoz-Vázquez, Oscar Martínez-Fuentes, and Guillermo Fernández-Anaya. Generalized PI control for robust stabilization of dynamical systems. Mathematics and Computers in Simulation, 202:22– 35, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422002427.

Xiao:2022:SOM

[990] Xufeng Xiao and Xinlong Feng. A second-order maximum bound principle preserving operator splitting method for the Allen–Cahn equation with applications in multiphase systems. Mathematics and Computers in Simulation, 202:36– 58, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422002269.

Deng:2022:NDS

[991] Xue Deng, Jiaxing Chen, Xu Wang, and Fengting Geng. Non-dominated sorting genetic algorithm-II for possibilistic mean-semiabsolute deviation-Yager entropy portfolio model with complex real-world constraints. Mathematics and Computers in Simulation, 202:59–78, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002099.

Kaur:2022:NSM

[992] Harpreet Kaur, K. N. Rai, and Subrahamanyam Upadhyay. A numerical study of moving boundary problem involving dual phase lag model of heat mass transfer during immersion frying. Mathematics and Computers in Simulation, 202:79– 100, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002270.

Neitzel:2022:IAB

[993] Leonie Neitzel and Edeltraud Gehrig. Influence of advection in box models describing thermohaline circulation. Mathematics and Computers in Simulation, 202:101–112, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002063.

Tawhid:2022:ISS

[994] Mohamed A. Tawhid and Abdelmonem M. Ibrahim. Improved salp swarm algorithm combined with chaos. Mathematics and Computers in Simulation, 202:113–148, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002415.

Che:2022:NPC

[995] Han Che, Yu-Lan Wang, and Zhi-Yuan Li. Novel patterns in a class of fractional reaction–diffusion models with the Riesz fractional derivative. Mathematics and Computers in Simulation, 202:149–163, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200249X.

Simos:2022:UNN

[996] Theodore E. Simos, Vasilios N. Katsikis, Spyridon D. Mourtas, and Predrag S. Stanimirović. Unique non-negative definite solution of the time-varying algebraic Riccati equations with applications to stabilization of LTV systems. Mathematics and Computers in Simulation, 202: 164–180, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002452.

Gorbunova:2022:NTR

[997] A. V. Gorbunova and A. V. Lebedev. Nontransitivity of tuples of random variables with polynomial density and its effects in Bayesian models. Mathematics and Computers in Simulation, 202:181–192, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002488.

Aydinlik:2022:EAB

[998] Soner Aydinlik, Ahmet Kiris, and Pradip Roul. An effective approach based on Smooth Composite Chebyshev Finite Difference Method and its applications to Bratu-type and higher order Lane–Emden problems. Mathematics and Computers in Simulation, 202:193–205, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002440.

Wang:2022:NCS

[999] Yan Wang, Zhiwei Zhao, Chao Ge, Lei Wang, and Yajuan Liu. Networked control systems with probabilistic timevarying delay based on event-triggered non-fragile H_{∞} control. Mathematics and Computers in Simulation, 202: 206–222, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002300.

Zhao:2022:GSQ

[1000] Rui Zhao, Baoxian Wang, and Jigui Jian. Global μ-stabilization of quaternion-valued inertial BAM neural networks with time-varying delays via time-delayed impulsive control. Mathematics and Computers in Simulation, 202:223–245, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002476.

Kumari:2022:CDI

[1001] Sarita Kumari, Satish Kumar Tiwari, and Ranjit Kumar Upadhyay. Cross diffusion induced spatiotemporal pattern in diffusive nutrient–plankton model with nutrient recycling. Mathematics and Computers in Simulation, 202:246–272, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002294.

Tao:2022:NSD

[1002] Shi Tao, Qing He, Xiaoping Yang, Jiahong Luo, and Xingxi Zhao. Numerical study on the drag and flow characteristics of porous particles at intermediate Reynolds numbers. Mathematics and Computers in Simulation, 202:273–294, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002567.

Lao:2022:PPC

[1003] Tongfei Lao and Yanrui Sun. Predicting the production and consumption of natural gas in China by using a new grey forecasting method. Mathematics and Computers in Simulation, 202:295–315, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200221X.

Karthiga:2022:DBC

[1004] R. Karthiga, K. Narasimhan, and Rengarajan Amirtharajan. Diagnosis of breast cancer for modern mammography using artificial intelligence. Mathematics and Computers in Simulation, 202:316–330, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002506.

Liao:2022:SFD

[1005] Guangyuan Liao and Limin Zhang.

Solving flows of dynamical systems by deep neural networks and a novel deep learning algorithm. Mathematics and Computers in Simulation, 202: 331–342, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002592.

Pan:2022:GOA

[1006] Jeng-Shyang Pan, Li-Gang Zhang, Ruo-Bin Wang, Václav Snášel, and Shu-Chuan Chu. Gannet optimization algorithm: a new metaheuristic algorithm for solving engineering optimization problems. Mathematics and Computers in Simulation, 202: 343–373, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002774.

Hussain:2022:CBC

[1007] Sardar Muhammad Hussain, Hasrat Hussain Shah, and Jong-Suk Ro. Comparison between computational cost of fractals using linedoublets. Mathematics and Computers in Simulation, 202:374–387, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422002580.

Xu:2022:MLA

[1008] Xinsheng Xu, Ping Ji, and Felix T. S. Chan. On maximizing a lossaverse buyer’s expected utility in a multi-sourcing problem. Mathematics and Computers in Simulation, 202: 388–404, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002464.

Barraza:2022:CLB

[1009] J. A. Reyes Barraza and R. Deiterding. A curvilinear lattice Boltzmann scheme for thermal flows. Mathematics and Computers in Simulation, 202: 405–420, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002579.

Guerra:2022:UMM

[1010] Ana Guerra, Jorge Belinha, and Renato Natal Jorge. Using a meshless method to assess the effect of mechanical loading in angiogenesis. Mathematics and Computers in Simulation, 202:421–441, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002518.

Talib:2022:CPA

[1011] Imran Talib, Zulfiqar Ahmad Noor, Zakia Hammouch, and Hammad Khalil. Compatibility of the Paraskevopoulos’s algorithm with operational matrices of Vieta–Lucas polynomials and applications. Mathematics and Computers in Simulation, 202:442– 463, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422002610.

Yang:2022:PMS

[1012] Jiangtao Yang. Periodic measure of a stochastic non-autonomous predator–prey system with impulsive effects. Mathematics and Computers in Simulation, 202:464–479, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002804.

Alves:2022:QCS

[1013] P. R. L. Alves. Quantifying chaos in stock markets before and during COVID-19 pandemic from the phase space reconstruction. Mathematics and Computers in Simulation, 202: 480–499, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003287.

Arian:2022:EVR

[1014] Hamid Arian, Mehrdad Moghimi, Ehsan Tabatabaei, and Shiva Zamani. Encoded Value-at-Risk: a machine learning approach for portfolio risk measurement. Mathematics and Computers in Simulation, 202: 500–525, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003172.

Yilmaz:2022:DST

[1015] Zeynep Yılmaz, Selahattin Maden, and Aytül Gökçe. Dynamics and stability of two predators–one prey mathematical model with fading memory in one predator. Mathematics and Computers in Simulation, 202: 526–539, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003263.

Baghani:2022:SIC

[1016] Omid Baghani. SCW-iterative-computational method for solving a wide class of nonlinear fractional optimal control problems with Caputo derivatives. Mathematics and Computers in Simulation, 202:540– 558, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003251.

Liang:2022:PTB

[1017] Tao Liang, Degang Yang, Li Lei, Wanli Zhang, and Ju Pan. Preassigned-time bipartite synchronization of complex networks with quantized couplings and stochastic perturbations. Mathematics and Computers in Simulation, 202: 559–570, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200324X.

Anonymous:2022:NIl

[1018] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 202:571, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003469.

Anonymous:2022:ICEk

[1019] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 202:572–575, December 2022. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422003470.

Anonymous:2022:EBm

[1020] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 202:ifc, December 2022. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200341X.

Li:2023:NIC

[1021] Wang Li, Haifeng Dai, Lingzhi Zhao, Donghua Zhao, and Yongzheng Sun. Noise-induced consensus of leaderfollowing multi-agent systems. Mathematics and Computers in Simulation, 203:1–11, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002853.

Anonymous:2023:PJa

[1022] Anonymous. Pages 1–982 (January 2023). Mathematics and Computers in Simulation, 203:1–982, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Karakoc:2023:CAB

[1023] Seydi Battal Gazi Karakoc, Asit Saha, and Derya Yıldırım Sucu. A collocation algorithm based on septic B-splines and bifurcation of traveling waves for Sawada–Kotera equation. Mathematics and Computers in Simulation, 203:12–27, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002907.

Pandey:2023:NAF

[1024] Divyansh Pandey, Rajesh K. Pandey, and R. P. Agarwal. Numerical approximation of fractional variational problems with several dependent variables using Jacobi poly-fractonomials. Mathematics and Computers in Simulation, 203:28–43, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002889.

Yang:2023:SDE

[1025] Te Yang, Zhen Wang, Jianwei Xia, and Hao Shen. Sampled-data exponential synchronization of stochastic chaotic Lur’e delayed systems. Mathematics and Computers in Simulation, 203:44–57, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002816.

Bera:2023:THM

[1026] Smaranika Bera, Dhrubasish Bhattacharyya, Ruhul Ali Khan, and Murari Mitra. Test for harmonic mean residual life function: a goodness of fit approach. Mathematics and Computers in Simulation, 203:58–70, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002786.

Puviyarasi:2023:DIA

[1027] B. Puviyarasi and C. Murukesh. Design and implementation of Adaptive Mixed Fuzzy Controller for MIMO nonlinear systems. Mathematics and Computers in Simulation, 203:71– 91, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002890.

Liptaj:2023:SAH

[1028] Andrej Liptaj. Statistical approach for highest precision numerical differentiation. Mathematics and Computers in Simulation, 203:92– 111, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002944.

Cai:2023:NME

[1029] Ying Cai, Jinru Chen, and Nan Wang. A Nitsche mixed extended finite element method for the biharmonic interface problem. Mathematics and Computers in Simulation, 203: 112–130, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002920.

Lopez-Ruiz:2023:VCN

[1030] Jorge López-Ruiz, Salvador Botello-Rionda, Rafael Herrera-Guzmán, and Mauricio Carrillo-Tripp. Viral capsid nanoindentation simulations using octree-type data structures. Mathematics and Computers in Simulation, 203:131–149, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002877.

Houda:2023:UIO

[1031] K. Houda, D. Saifia, M. Chadli, and S. Labiod. Unknown input observer based robust control for fuzzy descriptor systems subject to actuator saturation. Mathematics and Computers in Simulation, 203:150– 173, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542200283X.

Zand:2023:SSI

[1032] Behnam Zand, Pedram Ghaderi, and Fereidoun Amini. Structural system identification via synchronization technique and fuzzy logic. Mathematics and Computers in Simulation, 203:174–188, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002798.

Ahmad:2023:NAN

[1033] Fayyaz Ahmad, Shafiq Ur Rehman, and Aiman Zara. A new approach for the numerical approximation of modified Korteweg–de Vries equation. Mathematics and Computers in Simulation, 203:189–206, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002919.

Mahmood:2023:MSA

[1034] Tariq Mahmood, Mati ur Rahman, Muhammad Arfan, Sadaf-Ilyas Kayani, and Mei Sun. Mathematical study of Algae as a bio-fertilizer using fractal– fractional dynamic model. Mathematics and Computers in Simulation, 203:207–222, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002981.

Deng:2023:AIP

[1035] Dingwen Deng and Qiang Wu. Accuracy improvement of a Predictor– Corrector compact difference scheme for the system of two-dimensional coupled nonlinear wave equations. Mathematics and Computers in Simulation, 203:223–249, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003007.

Li:2023:OCD

[1036] Yazhi Li, Yan Wang, and Lili Liu. Optimal control of dengue vector based on a reaction–diffusion model. Mathematics and Computers in Simulation, 203:250–270, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002968.

Khalil:2023:NDS

[1037] Ahmed Mostafa Khalil, Ahmed Mohamed Zahran, and Rehab Basheer. A novel diagnosis system for detection of kidney disease by a fuzzy soft decision-making problem. Mathematics and Computers in Simulation, 203:271–305, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002841.

Jiang:2023:PAO

[1038] Shi-Jie Jiang, Shu-Chuan Chu, Fu-Min Zou, Jie Shan, Shi-Guang Zheng, and Jeng-Shyang Pan. A parallel Archimedes optimization algorithm based on Taguchi method for application in the control of variable pitch wind turbine. Mathematics and Computers in Simulation, 203: 306–327, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200297X.

Kumar:2023:MFE

[1039] Vinay Kumar, S. V. S. S. N. V.

G. Krishna Murthy, and B. V. Rathish Kumar. Multi-force effect on fluid flow, heat and mass transfer, and entropy generation in a stratified fluidsaturated porous enclosure. Mathematics and Computers in Simulation, 203:328–367, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422002956.

Liu:2023:DSN

[1040] He Liu, Chuanjun Dai, Hengguo Yu, Qing Guo, Jianbing Li, Aimin Hao, Jun Kikuchi, and Min Zhao. Dynamics of a stochastic non-autonomous phytoplankton–zooplankton system involving toxin-producing phytoplankton and impulsive perturbations. Mathematics and Computers in Simulation, 203:368–386, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002828.

Niu:2023:FHO

[1041] Yuxuan Niu, Yang Liu, Hong Li, and Fawang Liu. Fast highorder compact difference scheme for the nonlinear distributed-order fractional Sobolev model appearing in porous media. Mathematics and Computers in Simulation, 203:387– 407, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422003020.

Amat:2023:TOA

[1042] S. Amat, P. Ortiz, J. Ruiz, J. C. Trillo, and D. F. Yáñez. The translation operator. Applications to nonlinear reconstruction operators on nonuniform grids. Mathematics and Computers in Simulation, 203:408– 424, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002439.

Saulo:2023:QAC

[1043] Helton Saulo, Narayanaswamy Balakrishnan, and Roberto Vila. On a quantile autoregressive conditional duration model. Mathematics and Computers in Simulation, 203:425– 448, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003044.

Chen:2023:SPP

[1044] Xingzhi Chen, Baodan Tian, Xin Xu, Hailan Zhang, and Dong Li. A stochastic predator–prey system with modified LG-Holling type II functional response. Mathematics and Computers in Simulation, 203:449– 485, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422002865.

Al-Juaifri:2023:FEA

[1045] Ghassan A. Al-Juaifri and Akil J.

Harfash. Finite element analysis of nonlinear reaction–diffusion system of Fitzhugh–Nagumo type with Robin boundary conditions. Mathematics and Computers in Simulation, 203: 486–517, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200307X.

Dong:2023:NSM

[1046] Yunda Dong. A new splitting method for systems of monotone inclusions in Hilbert spaces. Mathematics and Computers in Simulation, 203:518– 537, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422002932.

Li:2023:AEF

[1047] Xiaolin Li and Shuling Li. Analysis of an element-free Galerkin method for the nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 203:538–552, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003019.

Herrmann:2023:ESF

[1048] Samuel Herrmann and Nicolas Massin. Exact simulation of the first passage time through a given level of jump diffusions. Mathematics and Computers in Simulation, 203:553– 576, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003093.

Nazimuddin:2023:SPN

[1049] A. K. M. Nazimuddin, M. Humayun Kabir, and M. Osman Gani. Spiral patterns and numerical bifurcation analysis in a three-component Brusselator model for chemical reactions. Mathematics and Computers in Simulation, 203:577–591, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200310X.

Gao:2023:QIM

[1050] Wenwu Gao, Jiecheng Wang, and Ran Zhang. Quasi-interpolation for multivariate density estimation on bounded domain. Mathematics and Computers in Simulation, 203:592– 608, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003081.

Castaneda:2023:STD

[1051] Antonio Rafael Selva Castañeda, Josue Mariño del Pozo, Erick Eduardo Ramirez-Torres, Eduardo José Roca Oria, Sorangel Bolaños Vaillant, Juan I. Montijano, and Luis Enrique Bergues Cabrales. Spatio temporal dynamics of direct current in treated anisotropic tumors. Mathematics and Computers in Simulation, 203:609–632, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003068.

She:2023:UCR

[1052] Zi-Hang She, Li-Min Qiu, and Wei Qu. An unconditionally convergent RSCSCS iteration method for Riesz space fractional diffusion equations with variable coefficients. Mathematics and Computers in Simulation, 203:633–646, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003056.

Li:2023:PTS

[1053] Xinna Li, Huaiqin Wu, and Jinde Cao. Prescribed-time synchronization in networks of piecewise smooth systems via a nonlinear dynamic eventtriggered control strategy. Mathematics and Computers in Simulation, 203:647–668, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003123.

Dehingia:2023:MAD

[1054] Kaushik Dehingia, Parthasakha Das, Ranjit Kumar Upadhyay, Arvind Kumar Misra, Fathalla A. Rihan, and Kamyar Hosseini. Modelling and analysis of delayed tumour–immune system with hunting T-cells. Mathematics and Computers in Simulation, 203:669–684, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003111.

Wei:2023:LDG

[1055] Leilei Wei and Huanhuan Wang. Local discontinuous Galerkin method for multi-term variable-order time fractional diffusion equation. Mathematics and Computers in Simulation, 203: 685–698, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003196.

Silambarasan:2023:SDW

[1056] Rathinavel Silambarasan and Adem Kılıçman. Solitons of dispersive wave steered from Navier–Bernoulli and Love’s hypothesis in cylindrical elastic rod with compressible Murnaghan’s materials. Mathematics and Computers in Simulation, 203:699– 720, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003160.

Bulai:2023:MMT

[1057] I. M. Bulai, M. C. De Bonis, C. Laurita, and V. Sagaria. Modeling metastatic tumor evolution, numerical resolution and growth prediction. Mathematics and Computers in Simulation, 203:721–740, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003032.

Kumar:2023:GSA

[1058] R. Prem Kumar, P. K. Santra, and G. S. Mahapatra. Global stability and analysing the sensitivity of parameters of a multiple-susceptible population model of SARS-CoV-2 emphasising vaccination drive. Mathematics and Computers in Simulation, 203:741–766, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003147.

Wang:2023:GTA

[1059] Jinliang Wang, Wenjing Wu, and Toshikazu Kuniya. Global threshold analysis on a diffusive host–pathogen model with hyperinfectivity and nonlinear incidence functions. Mathematics and Computers in Simulation, 203:767–802, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003159.

Kumar:2023:ITP

[1060] Anish Kumar and Sourav Das. Integral transforms and probability distributions for a certain class of Fox–Wright type functions and its applications. Mathematics and Computers in Simulation, 203:803– 825, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422003135.

Singh:2023:ENM

[1061] P. K. Singh and S. Saha Ray. An efficient numerical method based on Lucas polynomials to solve multidimensional stochastic Itô–Volterra integral equations. Mathematics and Computers in Simulation, 203:826– 845, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002993.

Zhang:2023:TDC

[1062] Hai Zhang, Yuhong Cheng, Weiwei Zhang, and Hongmei Zhang. Time-dependent and Caputo derivative order-dependent quasi-uniform synchronization on fuzzy neural networks with proportional and distributed delays. Mathematics and Computers in Simulation, 203:846– 857, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003214.

Zhang:2023:SSA

[1063] Xuncai Zhang, Shida Wang, Kai Zhao, and Yanfeng Wang. A salp swarm algorithm based on Harris Eagle foraging strategy. Mathematics and Computers in Simulation, 203: 858–877, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003202.

Ushakova:2023:NSG

[1064] Olga V. Ushakova and Natalya A.

Artyomova. Non-stationary grid generation algorithm for deformed volumes of revolution. Mathematics and Computers in Simulation, 203: 878–909, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003184.

Yang:2023:ICF

[1065] Zhanying Yang, Jie Zhang, Zhihui Zhang, and Jun Mei. An improved criterion on finite-time stability for fractional-order fuzzy cellular neural networks involving leakage and discrete delays. Mathematics and Computers in Simulation, 203: 910–925, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003305.

Bentout:2023:APN

[1066] Soufiane Bentout and Salih Djilali. Asymptotic profiles of a nonlocal dispersal SIR epidemic model with treat-age in a heterogeneous environment. Mathematics and Computers in Simulation, 203:926– 956, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422003226.

Djukic:2023:VLB

[1067] Tijana Djukic, Marko Topalovic, and Nenad Filipovic. Validation of lattice Boltzmann based software for blood flow simulations in complex patient-specific arteries against traditional CFD methods. Mathematics and Computers in Simulation, 203: 957–976, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003299.

Anonymous:2023:NIa

[1068] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 203:977, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003846.

Anonymous:2023:ICEa

[1069] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 203:978–981, January 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003858.

Anonymous:2023:EBa

[1070] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 203:ifc, January 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003792.

Tan:2023:SDS

[1071] Yiping Tan, Yongli Cai, Xiaoqin Wang, Zhihang Peng, Kai Wang, Ruoxia Yao, and Weiming Wang. Stochastic dynamics of an SIS epidemiological model with media coverage. Mathematics and Computers in Simulation, 204:1–27, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003330.

Anonymous:2023:PF

[1072] Anonymous. Pages 1–768 (February 2023). Mathematics and Computers in Simulation, 204:1–768, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Cui:2023:ROT

[1073] Huixia Cui, Jianlong Qiu, Jinde Cao, Ming Guo, Xiangyong Chen, and Sergey Gorbachev. Route optimization in township logistics distribution considering customer satisfaction based on adaptive genetic algorithm. Mathematics and Computers in Simulation, 204:28–42, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422002087.

Signing:2023:EPA

[1074] Francis Signing, Berge Tsanou, Jean Lubuma, and Samuel Bowong. Effects of periodic aerosol emission on the transmission dynamics of Neisseria Meningitis A. Mathematics and Computers in Simulation, 204: 43–70, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422002609.

Song:2023:PIE

[1075] Wei Song, Chong Fu, Yu Zheng, Ming Tie, Jun Liu, and Junxin Chen. A parallel image encryption algorithm using intra bitplane scrambling. Mathematics and Computers in Simulation, 204:71–88, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003317.

Liu:2023:CQC

[1076] Lingfeng Liu and Jie Wang. A cluster of 1D quadratic chaotic map and its applications in image encryption. Mathematics and Computers in Simulation, 204:89–114, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003329.

Cao:2023:NRG

[1077] Shuxing Cao, Zhijie Chen, and Zhanwen Yang. Numerical representations of global epidemic threshold for nonlinear infection-age SIR models. Mathematics and Computers in Simulation, 204:115–132, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003238.

S:2023:BWF

[1078] Kumbinarasaiah S., Manohara G., and Hariharan G. Bernoulli wavelets functional matrix technique for a system of nonlinear singular Lane Emden equations. Mathematics and Computers in Simulation, 204:133– 165, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003275.

Ahsan:2023:HWM

[1079] Muhammad Ahsan, Martin Bohner, Aizaz Ullah, Amir Ali Khan, and Sheraz Ahmad. A Haar wavelet multi-resolution collocation method for singularly perturbed differential equations with integral boundary conditions. Mathematics and Computers in Simulation, 204:166– 180, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003366.

Ammar:2023:BBA

[1080] Hamadi Ammar and Bechir Naffeti. A branch and bound algorithm for Holder bi-objective optimization. Implementation to multidimensional optimization. Mathematics and Computers in Simulation, 204: 181–201, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003354.

Tber:2023:SLM

[1081] Moulay Hicham Tber. A semi-Lagrangian mixed finite element method for advection–diffusion variational inequalities. Mathematics and Computers in Simulation, 204: 202–215, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200338X.

Ahmat:2023:DWS

[1082] Muyassar Ahmat and Jianxian Qiu. Direct WENO scheme for dispersiontype equations. Mathematics and Computers in Simulation, 204:216– 229, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422003561.

Mohapatra:2023:IVP

[1083] Dhabaleswar Mohapatra and S. Chakraverty. Initial value problems in Type-2 fuzzy environment. Mathematics and Computers in Simulation, 204:230– 242, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422003342.

He:2023:HPM

[1084] Ji-Huan He, Man-Li Jiao, Khaled A.

Gepreel, and Yasir Khan. Homotopy perturbation method for strongly nonlinear oscillators. Mathematics and Computers in Simulation, 204: 243–258, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003378.

Kumar:2023:COU

[1085] Anshul Kumar and Madhu Jain. Cost Optimization of an Unreliable server queue with two stage service process under hybrid vacation policy. Mathematics and Computers in Simulation, 204:259–281, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003391.

Babu:2023:MSS

[1086] N. Ramesh Babu and P. Balasubramaniam. Master–slave synchronization for glucose–insulin metabolism of type-1 diabetic Mellitus model based on new fractal–fractional order derivative. Mathematics and Computers in Simulation, 204:282– 301, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003603.

Omame:2023:GAS

[1087] Andrew Omame, Mujahid Abbas, and Anwarud Din. Global asymptotic stability, extinction and ergodic stationary distribution in a stochastic model for dual variants of SARS-CoV-2. Mathematics and Computers in Simulation, 204:302–336, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003585.

Wang:2023:IBO

[1088] Xiao Wang, Xiao-Xue Sun, Shu-Chuan Chu, Junzo Watada, and Jeng-Shyang Pan. Improved butterfly optimization algorithm applied to prediction of combined cycle power plant. Mathematics and Computers in Simulation, 204:337–353, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200355X.

Li:2023:CMM

[1089] Peng Li and Keying Liu. A coupled marching method for Cauchy problems of the Helmholtz equation in complex waveguides. Mathematics and Computers in Simulation, 204: 354–375, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003615.

He:2023:MLF

[1090] Quanqin He, Hao Liu, Guiyan Ding, and Liangping Tu. A modified Lévy flight distribution for solving high-dimensional numerical optimization problems. Mathematics and Computers in Simulation, 204:376– 400, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003639.

Barman:2023:OPP

[1091] Abhijit Barman, Pijus Kanti De, Ashis Kumar Chakraborty, Chee Peng Lim, and Rubi Das. Optimal pricing policy in a three-layer dual-channel supply chain under government subsidy in green manufacturing. Mathematics and Computers in Simulation, 204:401–429, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003548.

Dali:2023:CPT

[1092] Ali Dali, Samir Abdelmalek, Azzeddine Bakdi, and Maamar Bettayeb. A class of PSO-tuned controllers in Lorenz chaotic system. Mathematics and Computers in Simulation, 204: 430–449, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003627.

Quenjel:2023:NDA

[1093] El Houssaine Quenjel and Abdelaziz Beljadid. Node-diamond approximation of heterogeneous and anisotropic diffusion systems on arbitrary twodimensional grids. Mathematics and Computers in Simulation, 204:450– 472, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003573.

Chinnadurai:2023:SPL

[1094] M. Chinnadurai, Mohamed El Fatini, and A. Rathinasamy. Stochastic perturbation to 2-LTR dynamical model in HIV infected patients. Mathematics and Computers in Simulation, 204: 473–497, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003652.

Li:2023:OBB

[1095] Yu Li, Xiaomei Yu, and Jingsen Liu. An opposition-based butterfly optimization algorithm with adaptive elite mutation in solving complex high-dimensional optimization problems. Mathematics and Computers in Simulation, 204:498– 528, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422003664.

Li:2023:DAP

[1096] Wenjie Li, Ying Zhang, and Lihong Huang. Dynamics analysis of a predator–prey model with nonmonotonic functional response and impulsive control. Mathematics and Computers in Simulation, 204:529– 555, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422003688.

Asadzadeh:2023:SDD

[1097] M. Asadzadeh and L. Beilina. A stabilized P 1 domain decomposition finite element method for time harmonic Maxwell’s equations. Mathematics and Computers in Simulation, 204:556–574, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003597.

Zhao:2023:NAC

[1098] Hengzhi Zhao, Jiwei Zhang, and Jing Lu. Numerical approximate controllability for unidimensional parabolic integro-differential equations. Mathematics and Computers in Simulation, 204:575–596, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003676.

Souna:2023:CPF

[1099] Fethi Souna, Mustapha Belabbas, and Youssaf Menacer. Complex pattern formations induced by the presence of cross-diffusion in a generalized predator–prey model incorporating the Holling type functional response and generalization of habitat complexity effect. Mathematics and Computers in Simulation, 204: 597–618, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003706.

Kang:2023:IHH

[1100] Helei Kang, Renyun Liu, Yifei Yao, and Fanhua Yu. Improved Harris hawks optimization for non-convex function optimization and design optimization problems. Mathematics and Computers in Simulation, 204: 619–639, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003767.

Tomar:2023:SVI

[1101] Saurabh Tomar, Mehakpreet Singh, Kuppalapalle Vajravelu, and Higinio Ramos. Simplifying the variational iteration method: a new approach to obtain the Lagrange multiplier. Mathematics and Computers in Simulation, 204:640–644, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200369X.

Wei:2023:MMT

[1102] Hsiu-Chuan Wei. Mathematical modeling of tumor growth and treatment: Triple negative breast cancer. Mathematics and Computers in Simulation, 204:645–659, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003718.

Mehrdoust:2023:TFH

[1103] Farshid Mehrdoust, Idin Noorani, and Abdelouahed Hamdi. Twofactor Heston model equipped with regime-switching: American option pricing and model calibration by Levenberg–Marquardt optimization algorithm. Mathematics and Computers in Simulation, 204:660–678, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542200372X.

Wang:2023:QOB

[1104] Fei Wang, Ning Li, and Yongqing Yang. Quantized-observer based consensus for fractional order multiagent systems under distributed eventtriggered mechanism. Mathematics and Computers in Simulation, 204: 679–694, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003779.

Mandal:2023:IDC

[1105] Anindya Mandal and Brojeswar Pal.

Investigating dual-channel green supply chain considering refurbishing process and product recycling with environmental awareness effort. Mathematics and Computers in Simulation, 204:695–726, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003755.

Singh:2023:EPS

[1106] Gurwinder Singh and Amarinder Singh. Extension of Particle Swarm Optimization algorithm for solving two-level time minimization transportation problem. Mathematics and Computers in Simulation, 204:727– 742, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003895.

Goda:2023:CUG

[1107] Takashi Goda and Wataru Kitade. Constructing unbiased gradient estimators with finite variance for conditional stochastic optimization. Mathematics and Computers in Simulation, 204:743–763, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422003883.

Anonymous:2023:NIb

[1108] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 204:764, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004232.

Anonymous:2023:ICEb

[1109] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 204:765–768, February 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422004244.

Anonymous:2023:EBb

[1110] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 204:ifc, February 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004189.

Vengatesan:2023:ISP

[1111] S. Vengatesan, Sundararajan Natarajan, and P. V. Jeyakarthikeyan. n + 1 Integration scheme for polygonal elements using Richardson extrapolation. Mathematics and Computers in Simulation, 205(??):659– 677, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004116.

Fernandez-Diaz:2023:CFM

[1112] Julio M. Fernández-Díaz and César O.

Menéndez-Pérez. A common framework for modified Regula Falsi methods and new methods of this kind. Mathematics and Computers in Simulation, 205(??):678–696, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004335.

Quan:2023:SPD

[1113] Ho Dac Quan and Hieu Trung Huynh. Solving partial differential equation based on extreme learning machine. Mathematics and Computers in Simulation, 205(??):697– 708, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004323.

Yassopoulos:2023:ANT

[1114] Christopher Yassopoulos, J. N. Reddy, and Daniele Mortari. Analysis of nonlinear Timoshenko–Ehrenfest beam problems with von Kármán nonlinearity using the Theory of Functional Connections. Mathematics and Computers in Simulation, 205 (??):709–744, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004165.

Shang:2023:MBP

[1115] Zuchong Shang and Yuanhua Qiao. Multiple bifurcations in a predator– prey system of modified Holling and Leslie type with double Allee effect and nonlinear harvesting. Mathematics and Computers in Simulation, 205 (??):745–764, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004426.

Tassouli:2023:NNA

[1116] Siham Tassouli and Abdel Lisser. A neural network approach to solve geometric programs with joint probabilistic constraints. Mathematics and Computers in Simulation, 205 (??):765–777, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004384.

Bhuvaneshwari:2023:SSA

[1117] G. Bhuvaneshwari, M. Prakash, R. Rakkiyappan, and A. Manivannan. Stability and stabilization analysis of T –S fuzzy systems with distributed time-delay using statefeedback control. Mathematics and Computers in Simulation, 205(??): 778–793, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422004050.

Deng:2023:NIW

[1118] Huaijun Deng, Linna Liu, Jianyin Fang, Boyang Qu, and Quanzhen Huang. A novel improved whale optimization algorithm for optimization problems with multi-strategy and hybrid algorithm. Mathematics and Computers in Simulation, 205 (??):794–817, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004372.

Pu:2023:TLN

[1119] Nana Pu, Yifei Zhang, and Wentao Ma. A two-level nesting smoothed extended meshfree method for static and dynamic fracture mechanics analysis of orthotropic materials. Mathematics and Computers in Simulation, 205 (??):818–844, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004359.

Qiu:2023:PFO

[1120] Hongling Qiu, Jinde Cao, and Heng Liu. Passivity of fractional-order coupled neural networks with interval uncertainties. Mathematics and Computers in Simulation, 205(??):845– 860, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004438.

Agarwal:2023:KKT

[1121] Deepika Agarwal, Pitam Singh, and M. A. El Sayed. The Karush– Kuhn–Tucker (KKT) optimality conditions for fuzzy-valued fractional optimization problems. Mathematics and Computers in Simulation, 205 (??):861–877, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004396.

Casaban:2023:NDS

[1122] M.-C. Casabán, R. Company, and L. Jódar. Numerical difference solution of moving boundary random Stefan problems. Mathematics and Computers in Simulation, 205(??): 878–901, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004402.

Ding:2023:FDM

[1123] Peng Ding, Yubin Yan, Zongqi Liang, and Yuyuan Yan. Finite difference method for time-fractional Klein– Gordon equation on an unbounded domain using artificial boundary conditions. Mathematics and Computers in Simulation, 205(??):902– 925, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542200444X.

Liu:2023:PCS

[1124] Jingna Liu, Qi Qi, Bing Liu, and Shujing Gao. Pest control switching models with instantaneous and noninstantaneous impulsive effects. Mathematics and Computers in Simulation, 205(??):926–938, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004414.

Sartanpara:2023:SGF

[1125] Parthkumar P. Sartanpara and Ramakanta Meher. Solution of generalised fuzzy fractional Kaup–Kupershmidt equation using a robust multi parametric approach and a novel transform. Mathematics and Computers in Simulation, 205(??):939– 969, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004347.

Jain:2023:NCF

[1126] Sapna Jain and Rama Bhargava. Natural convection flow on a bent wavy vertical enclosure filled with powerlaw nanofluid simulated by Element Free Galerkin method. Mathematics and Computers in Simulation, 205 (??):970–986, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004475.

Sten:2023:EEE

[1127] Johan C.-E. Stén and Hanno Essén. On the evaluation of electrostatic energy of volume charge distributions and the role of the Coulomb singularity. Mathematics and Computers in Simulation, 205(??):987–996, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004499.

Torlo:2023:MOR

[1128] Davide Torlo and Mario Ricchiuto. Model order reduction strategies for weakly dispersive waves. Mathematics and Computers in Simulation, 205 (??):997–1028, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004487.

Ali:2023:BCR

[1129] Asgar Ali, Soumitra Sarkar, and Sanatan Das. Bioconvective chemically reactive entropy optimized crossnano-material conveying oxytactic microorganisms over a flexible cylinder with Lorentz force and Arrhenius kinetics. Mathematics and Computers in Simulation, 205(??):1029– 1051, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422004517.

Anonymous:2023:ICEc

[1130] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 205(??):1052, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475422004797.

Anonymous:2023:ICEd

[1131] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 205(??):1053– 1056, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004803.

Ding:2023:SPF

[1132] Hengfei Ding and Junhong Tian. Structure preserving fourth-order difference scheme for the nonlinear spatial fractional Schrödinger equation in two dimensions. Mathematics and Computers in Simulation, 205: 1–18, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003986.

Anonymous:2023:PMa

[1133] Anonymous. Pages 1–1056 (March 2023). Mathematics and Computers in Simulation, 205:1–1056, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liu:2023:THO

[1134] Changying Liu, Jiayin Li, Zhenqi Yang, Yumeng Tang, and Kai Liu. Two high-order energy-preserving and symmetric Gauss collocation integrators for solving the hyperbolic Hamiltonian systems. Mathematics and Computers in Simulation, 205: 19–32, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003937.

Darania:2023:MCM

[1135] P. Darania and S. Pishbin. Multistep collocation methods for integralalgebraic equations with non-vanishing delays. Mathematics and Computers in Simulation, 205:33–61, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475422003640.

Peng:2023:SAF

[1136] Qiu Peng and Jigui Jian. Synchronization analysis of fractionalorder inertial-type neural networks with time delays. Mathematics and Computers in Simulation, 205:62– 77, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004001.

Chrisment:2023:DNS

[1137] A. Chrisment, A. Debayle, J.-L. Feugeas, P. Loiseau, P.-E. Masson-Laborde, J. Mathiaud, Ph. Nicolaï, and V. Tikhonchuk. A deterministic numerical scheme for an electron heat transport model. Mathematics and Computers in Simulation, 205:78–97, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003913.

Chaube:2023:HBS

[1138] Shshank Chaube, Dheeraj Kumar Joshi, and Chandan Singh Ujarari. Hesitant Bifuzzy Set (an introduction): a new approach to assess the reliability of the systems. Mathematics and Computers in Simulation, 205:98–107, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003962.

Syrakos:2023:ULS

[1139] Alexandros Syrakos, Oliver Oxtoby, Eugene de Villiers, Stylianos Varchanis, Yannis Dimakopoulos, and John Tsamopoulos. A unification of leastsquares and Green–Gauss gradients under a common projection-based gradient reconstruction framework. Mathematics and Computers in Simulation, 205:108–141, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003743.

Arendt:2023:NME

[1140] V. Arendt. Numerical methods, energy conservation, and a new method for particle motion in magnetic fields. Mathematics and Computers in Simulation, 205:142– 185, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003925.

Mu:2023:BDD

[1141] Yu Mu and Zuxiong Li. Bifurcation dynamics of a delayed chemostat system with spatial diffusion. Mathematics and Computers in Simulation, 205:186–204, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003998.

Qiao:2023:EAF

[1142] Leijie Qiao, Wenlin Qiu, and Da Xu.

Error analysis of fast L1 ADI finite difference/compact difference schemes for the fractional telegraph equation in three dimensions. Mathematics and Computers in Simulation, 205:205–231, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004025.

Mittal:2023:SBE

[1143] Mandeep Mittal and Biswajit Sarkar. Stochastic behavior of exchange rate on an international supply chain under random energy price. Mathematics and Computers in Simulation, 205:232–250, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003731.

Lu:2023:VFA

[1144] Yu-Ming Lu and Yuh-Dauh Lyuu. Very fast algorithms for implied barriers and moving-barrier options pricing. Mathematics and Computers in Simulation, 205:251–271, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422003950.

Khirsariya:2023:NHT

[1145] Sagar R. Khirsariya, Snehal B. Rao, and Jignesh P. Chauhan. A novel hybrid technique to obtain the solution of generalized fractional-order differential equations. Mathematics and Computers in Simulation, 205: 272–290, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004141.

Hoang:2023:DAG

[1146] Manh Tuan Hoang. Dynamical analysis of a generalized hepatitis B epidemic model and its dynamically consistent discrete model. Mathematics and Computers in Simulation, 205:291–314, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004074.

Bzeih:2023:FES

[1147] Moussa Bzeih, Toufic El Arwadi, Ali Wehbe, Rodrigo L. R. Madureira, and Mauro A. Rincon. A finite element scheme for a 2D-wave equation with dynamical boundary control. Mathematics and Computers in Simulation, 205:315–339, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004013.

Li:2023:BIC

[1148] Jin Li, Xiaoning Su, and Kaiyan Zhao. Barycentric interpolation collocation algorithm to solve fractional differential equations. Mathematics and Computers in Simulation, 205: 340–367, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004062.

Marasi:2023:NST

[1149] H. R. Marasi and M. H. Derakhshan. Numerical simulation of time variable fractional order mobile– immobile advection–dispersion model based on an efficient hybrid numerical method with stability and convergence analysis. Mathematics and Computers in Simulation, 205:368– 389, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003974.

Zheng:2023:SOA

[1150] Zi-Yun Zheng and Yuan-Ming Wang. A second-order accurate Crank–Nicolson finite difference method on uniform meshes for nonlinear partial integrodifferential equations with weakly singular kernels. Mathematics and Computers in Simulation, 205:390– 413, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004037.

Du:2023:FOP

[1151] Wentong Du, Min Xiao, Jie Ding, Yi Yao, Zhengxin Wang, and Xinsong Yang. Fractional-order PD control at Hopf bifurcation in a delayed predator–prey system with transspecies infectious diseases. Mathematics and Computers in Simulation, 205:414–438, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004153.

Poochinapan:2023:HPC

[1152] Kanyuta Poochinapan and Ben Wongsaijai. High-performance computing of structure-preserving algorithm for the coupled BBM system formulated by weighted compact difference operators. Mathematics and Computers in Simulation, 205:439– 467, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422003949.

Gong:2023:SMC

[1153] Wei Gong and Zhanping Wang. Sliding motion control of Echinococcosis transmission dynamics model. Mathematics and Computers in Simulation, 205:468–482, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004098.

Li:2023:SPV

[1154] Dongfang Li, Xiaoxi Li, Ming Mei, and Wanqiu Yuan. A structure-preserving and variable-step BDF2 Fourier pseudo-spectral method for the twomode phase field crystal model. Mathematics and Computers in Simulation, 205:483–506, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004104.

Yadav:2023:PUH

[1155] Swati Yadav and Pratima Rai. A parameter uniform higher order scheme for 2D singularly perturbed parabolic convection–diffusion problem with turning point. Mathematics and Computers in Simulation, 205:507– 531, March 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422004128.

Srivastava:2023:LAC

[1156] Nikhil Srivastava and Vineet Kumar Singh. L3 approximation of Caputo derivative and its application to timefractional wave equation-(I). Mathematics and Computers in Simulation, 205:532–557, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004049.

Liu:2023:SHB

[1157] Yifan Liu, Jiazhi Cai, Haowen Xu, Minghe Shan, and Qingbin Gao. Stability and Hopf bifurcation of a love model with two delays. Mathematics and Computers in Simulation, 205:558–580, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200413X.

Luo:2023:BAN

[1158] Jianfeng Luo and Yi Zhao. Bifurcation analysis of a non-smooth prey– predator model by a differential linear complementarity system. Mathematics and Computers in Simulation, 205:581–599, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200430X.

Koutou:2023:MAI

[1159] Ousmane Koutou, Abou Bakari Diabaté, and Boureima Sangaré. Mathematical analysis of the impact of the media coverage in mitigating the outbreak of COVID-19. Mathematics and Computers in Simulation, 205:600–618, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004311.

Ma:2023:MSI

[1160] Shuidong Ma, Yiming Fang, Xiaodong Zhao, and Zhendong Liu. Multiswarm improved Grey Wolf Optimizer with double adaptive weights and dimension learning for global optimization problems. Mathematics and Computers in Simulation, 205:619–641, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004086.

Peng:2023:LBS

[1161] Congming Peng, Dun Zhao, and Qihong Shi. Limit of the blow-up solution for the inhomogeneous nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 205:642–658, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004360.

Anonymous:2023:EBc

[1162] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 205:ifc, March 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004748.

Akram:2023:SSR

[1163] Ghazala Akram, Maasoomah Sadaf, and M. Atta Ullah Khan. Soliton solutions of the resonant nonlinear Schrödinger equation using modified auxiliary equation method with three different nonlinearities. Mathematics and Computers in Simulation, 206:1–20, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004463.

Anonymous:2023:PAa

[1164] Anonymous. Pages 1–798 (April 2023). Mathematics and Computers in Simulation, 206:1–798, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Swati:2023:ETB

[1165] Swati, Mandeep Singh, and Karanjeet Singh. An efficient technique based on higher order Haar wavelet method for Lane–Emden equations. Mathematics and Computers in Simulation, 206:21–39, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004451.

Shiromani:2023:CMT

[1166] Ram Shiromani, Vembu Shanthi, and Higinio Ramos. A computational method for a two-parameter singularly perturbed elliptic problem with boundary and interior layers. Mathematics and Computers in Simulation, 206:40–64, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004529.

Kohnesara:2023:NSK

[1167] Sima Molaei Kohnesara and Ali Rahmani Firoozjaee. Numerical solution of Korteweg–de Vries equation using discrete least squares meshless method. Mathematics and Computers in Simulation, 206:65–76, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004505.

Jankowska:2023:KRA

[1168] Małgorzata A. Jankowska, Andreas Karageorghis, and C. S. Chen. Kansa– RBF algorithms for elliptic BVPs in annular domains with mixed boundary conditions. Mathematics and Computers in Simulation, 206:77– 104, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004554.

Shaikhet:2023:SEE

[1169] Leonid Shaikhet. Stability of equilibria of exponential type system of three differential equations under stochastic perturbations. Mathematics and Computers in Simulation, 206:105–117, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004578.

Liu:2023:LSD

[1170] Meng-Meng Liu. M-lump solutions to the (2 + 1)-dimensional generalized Calogero–Bogoyavlenshii– Schiff equation. Mathematics and Computers in Simulation, 206:118– 129, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542200458X.

Corona:2023:NMS

[1171] Gemma Corona, Oscar Maciel-Castillo, Juan Morales-Castañeda, Adrian Gonzalez, and Erik Cuevas. A new method to solve rotated template matching using metaheuristic algorithms and the structural similarity index. Mathematics and Computers in Simulation, 206:130–146, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004542.

Garcia:2023:IPR

[1172] Christian Cortés García. Impact of prey refuge in a discontinuous Leslie– Gower model with harvesting and alternative food for predators and linear functional response. Mathematics and Computers in Simulation, 206:147–165, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422004621.

Gargari:2023:ECS

[1173] Masoud Atashpaz Gargari and Rashed Sahraeian. An exact criterion space search method for a bi-objective nursing home location and allocation problem. Mathematics and Computers in Simulation, 206:166– 180, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004608.

Qi:2023:GDD

[1174] Ke Qi, Zhijun Liu, Lianwen Wang, and Yuming Chen. Global dynamics of a diffusive SEICR HCV model with nonlinear incidences. Mathematics and Computers in Simulation, 206:181–197, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004669.

Kordkheili:2023:ANN

[1175] Masoud Sheikhi Kordkheili and Farshad Rahimpour. Artificial neural network and semi-empirical modeling of industrial-scale Gasoil hydrodesulfurization reactor temperature profile. Mathematics and Computers in Simulation, 206:198– 215, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542200461X.

Shirozhan:2023:FIT

[1176] M. Shirozhan, Hassan S. Bakouch, and M. Mohammadpour. A flexible INAR(1) time series model with dependent zero-inflated count series and medical contagious cases. Mathematics and Computers in Simulation, 206:216–230, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004633.

Sarkarati:2023:PCE

[1177] Saeed Sarkarati, Mohammad Mehdi Tehranchi, and Esfandiar Mehrshahi. Precise calculation of electrical capacitance by means of quadruple integrals in method of moments technique. Mathematics and Computers in Simulation, 206:231– 240, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004645.

Sarangi:2023:DSE

[1178] B. P. Sarangi and S. N. Raw. Dynamics of a spatially explicit eco-epidemic model with double Allee effect. Mathematics and Computers in Simulation, 206:241–263, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004530.

Zhang:2023:LCM

[1179] Shangyuan Zhang and Yufeng Nie. Localized Chebyshev and MLS collocation methods for solving 2D steady state nonlocal diffusion and peridynamic equations. Mathematics and Computers in Simulation, 206:264– 285, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004670.

Aghazadeh:2023:LAM

[1180] A. Aghazadeh, Y. Mahmoudi, and F. D. Saei. Legendre approximation method for computing eigenvalues of fourth order fractional Sturm– Liouville problem. Mathematics and Computers in Simulation, 206:286– 301, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004566.

Turgut:2023:LSE

[1181] Oguz Emrah Turgut and Mert Sinan Turgut. Local search enhanced Aquila optimization algorithm ameliorated with an ensemble of Wavelet mutation strategies for complex optimization problems. Mathematics and Computers in Simulation, 206:302– 374, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004694.

Gonzalez-Calderon:2023:APH

[1182] Alfredo González-Calderón, Luis X.

Vivas-Cruz, M. A. Taneco-Hernández, and J. F. Gómez-Aguilar. Assessment of the performance of the hyperbolic-NILT method to solve fractional differential equations. Mathematics and Computers in Simulation, 206:375–390, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004712.

Yan:2023:FTS

[1183] Shaohui Yan, Binxian Gu, Ertong Wang, and Yu Ren. Finite-time synchronization of multi-scroll hyperchaotic system and its application in image encryption. Mathematics and Computers in Simulation, 206:391–409, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004657.

Chen:2023:OCI

[1184] Ya Chen, Juping Zhang, and Zhen Jin. Optimal control of an influenza model with mixed crossinfection by age group. Mathematics and Computers in Simulation, 206:410–436, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004682.

Ngwa:2023:CTP

[1185] Gideon A. Ngwa, Miranda I. Teboh-Ewungkem, and James A. Njongwe. Continuous-time predator-prey-like systems used to investigate the question: Can human consciousness help eliminate temporary mosquito breeding sites from around human habitats? Mathematics and Computers in Simulation, 206:437– 469, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004591.

Shao:2023:BLN

[1186] Peng Shao, Ying Liang, Guangquan Li, Xing Li, and Le Yang. Birefringence learning: a new global optimization technology model based on birefringence principle in application on artificial bee colony. Mathematics and Computers in Simulation, 206:470–486, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004700.

Sun:2023:REB

[1187] Ligang Sun, Roeland J. Dilz, and Martijn C. van Beurden. A rationalexpansion-based method to compute Gabor coefficients of 2D indicator functions supported on polygonal domain. Mathematics and Computers in Simulation, 206:487– 502, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004918.

Patawa:2023:ISS

[1188] Rohit Patawa and Pramendra Singh Pundir. Inferential study of single unit repairable system. Mathematics and Computers in Simulation, 206:503–516, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004906.

Feng:2023:GFE

[1189] Libo Feng, Fawang Liu, and Vo V. Anh.

Galerkin finite element method for a two-dimensional tempered time–space fractional diffusion equation with application to a Bloch–Torrey equation retaining Larmor precession. Mathematics and Computers in Simulation, 206:517–537, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004864.

Mohammadi:2023:UPS

[1190] Shaban Mohammadi and S. Reza Hejazi. Using particle swarm optimization and genetic algorithms for optimal control of non-linear fractional-order chaotic system of cancer cells. Mathematics and Computers in Simulation, 206:538–560, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004724.

Luo:2023:CSE

[1191] Xu-Bao Luo, D. M. Li, and Bin Liu. A comparison study of the efficiency and accuracy of IEFG in solving elasticity problems using different nodal integration schemes. Mathematics and Computers in Simulation, 206:561–587, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200489X.

Xie:2023:CFD

[1192] Jianqiang Xie, Quanxiang Wang, and Zhiyue Zhang. Conservative finite difference methods for the Boussinesq paradigm equation. Mathematics and Computers in Simulation, 206:588–613, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200492X.

Fardi:2023:NDS

[1193] M. Fardi, M. A. Zaky, and A. S.

Hendy. Nonuniform difference schemes for multi-term and distributed-order fractional parabolic equations with fractional Laplacian. Mathematics and Computers in Simulation, 206: 614–635, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004967.

Ansari:2023:SPF

[1194] Alireza Ansari and Mohammad Hossein Derakhshan. On spectral polar fractional Laplacian. Mathematics and Computers in Simulation, 206:636–663, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004955.

Wang:2023:EOP

[1195] Fong-Fan Wang. An efficient optimization procedure for locationinventory problems with (S − 1, S) policy and retrial demands. Mathematics and Computers in Simulation, 206:664–688, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004979.

Yao:2023:SDS

[1196] Hui Yao, Chuanju Xu, and Mejdi Azaiez. Stable and decoupled schemes for an electrohydrodynamics model. Mathematics and Computers in Simulation, 206:689–708, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475422004943.

Hu:2023:HCS

[1197] Gang Hu, Rui Yang, and Guo Wei. Hybrid chameleon swarm algorithm with multi-strategy: a case study of degree reduction for disk Wang–Ball curves. Mathematics and Computers in Simulation, 206:709– 769, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422004888.

Ding:2023:EHO

[1198] Yuting Ding and Liyuan Zheng. Existence of homoclinic orbit of Shilnikov type and the application in Rössler system. Mathematics and Computers in Simulation, 206:770– 779, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422005006.

Zou:2023:TCS

[1199] Xiaoling Zou, Qingwei Li, Wenhao Cao, and Jingliang Lv. Thresholds and critical states for a stochastic predator–prey model with mixed functional responses. Mathematics and Computers in Simulation, 206: 780–795, April 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422005031.

Anonymous:2023:NIc

[1200] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 206:796, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000228.

Anonymous:2023:ICEe

[1201] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 206:797, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847542300023X.

Anonymous:2023:EBd

[1202] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 206:ifc, April 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000174.

Garg:2023:AFE

[1203] Divay Garg and Kamana Porwal.

Adaptive finite element methods for a fourth order obstacle problem and a state constrained optimal control problem. Mathematics and Computers in Simulation, 207: 1–23, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422004980.

Anonymous:2023:PMb

[1204] Anonymous. Pages 1–576 (May 2023). Mathematics and Computers in Simulation, 207:1–576, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tomljanovic:2023:DOE

[1205] Zoran Tomljanović. Damping optimization of the excited mechanical system using dimension reduction. Mathematics and Computers in Simulation, 207:24–40, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422005043.

Casado:2023:IGA

[1206] A. Casado, S. Bermudo, A. D. López-Sánchez, and J. Sánchez-Oro. An iterated greedy algorithm for finding the minimum dominating set in graphs. Mathematics and Computers in Simulation, 207:41–58, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475422005055.

Simonella:2023:XMC

[1207] Roberta Simonella and Carlos Vázquez. XVA in a multi-currency setting with stochastic foreign exchange rates. Mathematics and Computers in Simulation, 207:59–79, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422005018.

Wang:2023:AND

[1208] Wansheng Wang and Yi Huang.

Analytical and numerical dissipativity for the space-fractional Allen– Cahn equation. Mathematics and Computers in Simulation, 207:80– 96, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004992.

Calegari:2023:ANM

[1209] Priscila C. Calegari, Alexandre M.

Roma, Luis C. C. Santos, and Guenther C. Krieger Filho. An adaptive numerical method for simulating diffusion flame jets. Mathematics and Computers in Simulation, 207: 97–110, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422005080.

Nwaibeh:2023:DMS

[1210] E. A. Nwaibeh and C. R. Chikwendu. A deterministic model of the spread of scam rumor and its numerical simulations. Mathematics and Computers in Simulation, 207:111–129, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422005110.

Hussain:2023:MAS

[1211] Shah Hussain, Osman Tunç, Ghaus ur Rahman, Hasib Khan, and Elissa Nadia. Mathematical analysis of stochastic epidemic model of MERS-corona & application of ergodic theory. Mathematics and Computers in Simulation, 207:130– 150, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422005109.

Zhang:2023:CNC

[1212] Jinzhong Zhang, Gang Zhang, Min Kong, Tan Zhang, Duansong Wang, and Rui Chen. CWOA: a novel complex-valued encoding whale optimization algorithm. Mathematics and Computers in Simulation, 207: 151–188, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422005092.

Kamm:2023:NSK

[1213] Kevin Kamm, Stefano Pagliarani, and Andrea Pascucci. Numerical solution of kinetic SPDEs via stochastic Magnus expansion. Mathematics and Computers in Simulation, 207:189–208, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542200516X.

Ren:2023:COD

[1214] Jing Ren, Hao Sun, Genjiu Xu, and Dongshuang Hou. Convergence of output dynamics in duopoly co-opetition model with incomplete information. Mathematics and Computers in Simulation, 207:209– 225, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422005134.

UlArifeen:2023:PGA

[1215] Shams Ul Arifeen and Sirajul Haq.

Petrov–Galerkin approximation of time-fractional coupled Korteweg–de Vries equation for propagation of long wave in shallow water. Mathematics and Computers in Simulation, 207:226–242, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422005158.

Singh:2023:HPS

[1216] Bikramjit Singh and Amarinder Singh.

Hybrid particle swarm optimization for pure integer linear solid transportation problem. Mathematics and Computers in Simulation, 207: 243–266, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422005067.

Liu:2023:TDA

[1217] Fang Liu, Ya-Ru Chen, and Da-Hai Zhou. A two-dimensional approach to flexibility degree of XOR numbers with application to group decision making. Mathematics and Computers in Simulation, 207:267– 287, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422005171.

Jing:2023:TDH

[1218] Shaoxue Jing. Time-delay Hammerstein system identification using modified cross-correlation method and variable stacking length multierror algorithm. Mathematics and Computers in Simulation, 207:288– 300, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422005195.

Priyanka:2023:ESA

[1219] K. Sri Raja Priyanka, G. Soundararajan, Ardak Kashkynbayev, and G. Nagamani. Exponential H_{∞} synchronization and anti-synchronization of delayed discrete-time complexvalued neural networks with uncertainties. Mathematics and Computers in Simulation, 207:301–321, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475423000022.

Zhu:2023:SIE

[1220] Shenli Zhu, Xiaoheng Deng, Wendong Zhang, and Congxu Zhu. Secure image encryption scheme based on a new robust chaotic map and strong S-box. Mathematics and Computers in Simulation, 207:322– 346, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475422005122.

Sajjad:2023:TDF

[1221] Assad Sajjad, Muhammad Farman, Ali Hasan, and Kottakkaran Sooppy Nisar.

Transmission dynamics of fractional order yellow virus in red chili plants with the Caputo–Fabrizio operator. Mathematics and Computers in Simulation, 207:347–368, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000046.

Solhi:2023:AST

[1222] Erfan Solhi, Farshid Mirzaee, and Shiva Naserifar. Approximate solution of two dimensional linear and nonlinear stochastic Itô-Volterra integral equations via meshless scheme. Mathematics and Computers in Simulation, 207:369–387, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000095.

Nyoumbi:2023:CFF

[1223] Christelle Dleuna Nyoumbi and Antoine Tambue. Convergence of a fitted finite volume method for pricing two dimensional assets with stochastic volatilities. Mathematics and Computers in Simulation, 207:388– 416, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000010.

Goel:2023:NAG

[1224] Eti Goel, Rajesh K. Pandey, S. Yadav, and Om P. Agrawal. A numerical approximation for generalized fractional Sturm–Liouville problem with application. Mathematics and Computers in Simulation, 207:417–436, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000034.

Sun:2023:RFA

[1225] Lu-Yao Sun, Siu-Long Lei, Hai-Wei Sun, and Jia-Li Zhang. An αrobust fast algorithm for distributedorder time–space fractional diffusion equation with weakly singular solution. Mathematics and Computers in Simulation, 207:437– 452, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423000113.

Ham:2023:SAM

[1226] Seokjun Ham and Junseok Kim. Stability analysis for a maximum principle preserving explicit scheme of the Allen–Cahn equation. Mathematics and Computers in Simulation, 207:453–465, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000277.

Ali:2023:FFO

[1227] Zeeshan Ali, Faranak Rabiei, and Kamyar Hosseini. A fractal–fractional-order modified Predator–Prey mathematical model with immigrations. Mathematics and Computers in Simulation, 207:466–481, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542300006X.

Jin:2023:NTP

[1228] Jie Jin, Weijie Chen, Lixin Qiu, Jingcan Zhu, and Haiyan Liu. A noise tolerant parameter-variable zeroing neural network and its applications. Mathematics and Computers in Simulation, 207:482–498, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000125.

Ojo:2023:MMC

[1229] Mayowa M. Ojo, Olumuyiwa James Peter, Emile Franc Doungmo Goufo, and Kottakkaran Sooppy Nisar. A mathematical model for the co-dynamics of COVID-19 and tuberculosis. Mathematics and Computers in Simulation, 207:499–520, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000149.

Li:2023:OSC

[1230] Tangjuan Li and Yanni Xiao. Optimal strategies for coordinating infection control and socio-economic activities. Mathematics and Computers in Simulation, 207:533–555, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423000289.

Chaudhari:2023:GPP

[1231] Urmila Chaudhari, Amrita Bhadoriya, Mrudul Y. Jani, and Biswajit Sarkar. A generalized payment policy for deteriorating items when demand depends on price, stock, and advertisement under carbon tax regulations. Mathematics and Computers in Simulation, 207:556–574, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542200502X.

Anonymous:2023:NId

[1232] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 207:575, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542300071X.

Anonymous:2023:ICEf

[1233] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 207:576, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475423000721.

Anonymous:2023:EBe

[1234] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 207:ifc, May 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000666.

Boyd:2023:UPA

[1235] John P. Boyd. Using parity to accelerate Hermite function computations: Zeros of truncated Hermite series, Gaussian quadrature and Clenshaw summation. Mathematics and Computers in Simulation, 207:

521–532, May 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422004931.

Yang:2023:NAL

[1236] Xiaochen Yang, Zhanwen Yang, and Chiping Zhang. Numerical analysis of the Linearly implicit Euler method with truncated Wiener process for the stochastic SIR model. Mathematics and Computers in Simulation, 208:1–14, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000101.

Anonymous:2023:PJb

[1237] Anonymous. Pages 1–746 (June 2023).

Mathematics and Computers in Simulation, 208:1–746, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Biswas:2023:SFO

[1238] Chetna Biswas, Anup Singh, Manish Chopra, and Subir Das. Study of fractional-order reaction–advection– diffusion equation using neural network method. Mathematics and Computers in Simulation, 208:15– 27, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475422005183.

El-Ganaini:2023:SCC

[1239] Shoukry El-Ganaini and Sachin Kumar. Symbolic computation to construct new soliton solutions and dynamical behaviors of various wave structures for two different extended and generalized nonlinear Schrödinger equations using the new improved modified generalized sub-ODE proposed method. Mathematics and Computers in Simulation, 208:28– 56, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000137.

Li:2023:NSA

[1240] Yan Li, Xiazhi Hao, Ruoxia Yao, Yarong Xia, and Yali Shen. Nonlinear superposition among lump soliton, stripe solitons and other nonlinear localized waves of the (2 + 1)-dimensional cpKP-BKP equation. Mathematics and Computers in Simulation, 208:57–70, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000307.

Deeb:2023:PGD

[1241] Ahmad Deeb, Omar Kalaoun, and Rafik Belarbi. Proper Generalized Decomposition using Taylor expansion for non-linear diffusion equations. Mathematics and Computers in Simulation, 208:71–94, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000071.

Kaveh:2023:OAO

[1242] Mehrdad Kaveh, Mohammad Saadi Mesgari, and Bahram Saeidian. Orchard Algorithm (OA): a new metaheuristic algorithm for solving discrete and continuous optimization problems. Mathematics and Computers in Simulation, 208:95–135, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422005146.

Shen:2023:UGP

[1243] Jiayu Shen, Jianxin Shi, Lingceng Gao, Qiang Zhang, and Kai Zhu. Uncertain green product supply chain with government intervention. Mathematics and Computers in Simulation, 208:136–156, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000368.

Zhang:2023:BAS

[1244] Jiajia Zhang and Yuanhua Qiao.

Bifurcation analysis of an SIR model considering hospital resources and vaccination. Mathematics and Computers in Simulation, 208:157– 185, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542300037X.

Xu:2023:LTE

[1245] Da Xu. The long time error estimates for the second order backward difference approximation to subdiffusion equations with boundary time delay and feedback gain. Mathematics and Computers in Simulation, 208:186–206, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000411.

Wang:2023:LSE

[1246] Jixia Wang, Xiaofang Xiao, and Chao Li. Least squares estimations for approximate fractional Vasicek model driven by a semimartingale. Mathematics and Computers in Simulation, 208:207–218, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000150.

Iskakova:2023:DSN

[1247] Kulpash Iskakova, Mohammad Mahtab Alam, Shabir Ahmad, Sayed Saifullah, Ali Akgül, and Gülnur Yılmaz. Dynamical study of a novel 4D hyperchaotic system: an integer and fractional order analysis. Mathematics and Computers in Simulation, 208:219–245, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000381.

Yang:2023:MSL

[1248] Xu Yang and Hongru Li. Multisample learning particle swarm optimization with adaptive crossover operation. Mathematics and Computers in Simulation, 208:246–282, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475422005079.

Wu:2023:IPT

[1249] Daiyong Wu, Youwei Yang, and Peng Wu. Impacts of prey-taxis and nonconstant mortality on a spatiotemporal predator–prey system. Mathematics and Computers in Simulation, 208:283–300, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000484.

Saleh:2023:NSE

[1250] Mahmoud Saleh, Endre Kovács, and Ádám Nagy. New stable, explicit, second order hopscotch methods for diffusion-type problems. Mathematics and Computers in Simulation, 208:301–325, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000435.

Kaushik:2023:ODM

[1251] Sonali Kaushik and Rajesh Kumar. Optimized decomposition method for solving multi-dimensional Burgers’ equation. Mathematics and Computers in Simulation, 208:326– 350, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000575.

Fisher:2023:AOS

[1252] Nick Fisher. ADI+MDA orthogonal spline collocation for the pressure Poisson reformulation of the Navier–Stokes equation in two space variables. Mathematics and Computers in Simulation, 208:351–365, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000344.

Torkaman:2023:CQM

[1253] Soraya Torkaman, Mohammad Heydari, and Ghasem Barid Loghmani. A combination of the quasilinearization method and linear barycentric rational interpolation to solve nonlinear multi-dimensional Volterra integral equations. Mathematics and Computers in Simulation, 208:366– 397, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423000538.

Li:2023:OEE

[1254] Jiyong Li. Optimal error estimates of a time-splitting Fourier pseudospectral scheme for the Klein–Gordon– Dirac equation. Mathematics and Computers in Simulation, 208:398– 423, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423000356.

Wang:2023:FTP

[1255] Chen Wang, Hai Zhang, Renyu Ye, Weiwei Zhang, and Hongmei Zhang. Finite time passivity analysis for Caputo fractional BAM reaction–diffusion delayed neural networks. Mathematics and Computers in Simulation, 208:424–443, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000563.

Terefe:2023:ECB

[1256] Y. A. Terefe, J. B. H. Njagarah, and S. M. Kassa. Effect of cross-border migration on the healthcare system of a destination community: Insights from mathematical modelling of COVID-19 in a developing country. Mathematics and Computers in Simulation, 208:444–479, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300054X.

Okhli:2023:TCM

[1257] Kheirolah Okhli and Mehdi Jabbari Nooghabi. On the threecomponent mixture of exponential distributions: a Bayesian framework to model data with multiple lower and upper outliers. Mathematics and Computers in Simulation, 208: 480–500, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000514.

Fan:2023:DGF

[1258] Enyu Fan, Changpin Li, and Martin Stynes. Discretised general fractional derivative. Mathematics and Computers in Simulation, 208:501– 534, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000447.

Izgi:2023:RWS

[1259] Zehra Pinar Izgi. Rogue waves and solitons of the generalized modified nonlinear Schrödinger equations. Mathematics and Computers in Simulation, 208:535–549, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000551.

Fiuzy:2023:RPC

[1260] Mohammad Fiuzy and Saeed Shamaghdari. Robust H_{∞}-PID control stability of fractional-order linear systems with polytopic and two-norm bounded uncertainties subject to input saturation. Mathematics and Computers in Simulation, 208:550– 581, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000393.

Kumar:2023:SAS

[1261] Naresh Kumar. Supercloseness analysis of a stabilizer free weak Galerkin finite element method for time dependent convection diffusion reaction equation. Mathematics and Computers in Simulation, 208:582– 602, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000587.

Huang:2023:HOR

[1262] Yifei Huang, Gang Peng, Gengen Zhang, and Hong Zhang. High-order Runge–Kutta structure-preserving methods for the coupled nonlinear Schrödinger– KdV equations. Mathematics and Computers in Simulation, 208:603– 618, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000459.

Kudenatti:2023:NSB

[1263] Ramesh B. Kudenatti, Noor E. Misbah, and Bharathi M. C. A numerical study on boundary layer flow of Carreau fluid and forced convection heat transfer with viscous dissipation and generalized thermal conductivity. Mathematics and Computers in Simulation, 208:619–636, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300040X.

Gao:2023:RAO

[1264] Shan Gao. Reliability analysis and optimization for a redundant system with dependent failures and variable repair rates. Mathematics and Computers in Simulation, 208:637– 659, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000605. See

corrigendum [1752].

Cao:2023:VBL

[1265] Jiling Cao, Jeong-Hoon Kim, Xi Li, and Wenjun Zhang. Valuation of barrier and lookback options under hybrid CEV and stochastic volatility. Mathematics and Computers in Simulation, 208:660–676, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000496.

Yu:2023:NDH

[1266] Yang Yu, Yu Wang, Rui Deng, and Yu Yin. New DY-HS hybrid conjugate gradient algorithm for solving optimization problem of unsteady partial differential equations with convection term. Mathematics and Computers in Simulation, 208:677–701, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000472.

Peng:2023:PEE

[1267] Xiangyi Peng, Da Xu, and Wenlin Qiu. Pointwise error estimates of compact difference scheme for mixed-type timefractional Burgers’ equation. Mathematics and Computers in Simulation, 208:702–726, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000630.

Khan:2023:AOB

[1268] Suliman Khan, Sakhi Zaman, Muhammad Arshad, Sharifah E. Alhazmi, Feroz Khan, and Jongee Park. Approximation of oscillatory Bessel integral transforms. Mathematics and Computers in Simulation, 208:727– 744, June 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423000423.

Anonymous:2023:NIe

[1269] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 208:745, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001015.

Anonymous:2023:ICEg

[1270] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 208:746, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423001027.

Anonymous:2023:EBf

[1271] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 208:ifc, June 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000964.

Fu:2023:ILC

[1272] Xingjian Fu and Jianshuai Peng. Iterative learning control for UAVs formation based on point-to-point trajectory update tracking. Mathematics and Computers in Simulation, 209:1–15, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000526.

Anonymous:2023:PJc

[1273] Anonymous. Pages 1–452 (July 2023).

Mathematics and Computers in Simulation, 209:1–452, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Saha:2023:DDO

[1274] Pritam Saha, Gopal Chandra Sikdar, Jayanta Kumar Ghosh, and Uttam Ghosh. Disease dynamics and optimal control strategies of a two serotypes dengue model with co-infection. Mathematics and Computers in Simulation, 209:16–43, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000848.

Wu:2023:ADT

[1275] Yunyun Wu and Yayun Li. Antidiagonalization theory and algorithm of matrices — from skew-symmetric matrices to arbitrary matrices. Mathematics and Computers in Simulation, 209:44–54, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000599.

Yan:2023:NRL

[1276] Zheping Yan, Jinyu Yan, Yifan Wu, Sijia Cai, and Hongxing Wang. A novel reinforcement learning based tuna swarm optimization algorithm for autonomous underwater vehicle path planning. Mathematics and Computers in Simulation, 209: 55–86, July 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000629.

Liu:2023:EAF

[1277] Li-Bin Liu, Lei Xu, and Yong Zhang. Error analysis of a finite difference scheme on a modified graded mesh for a time-fractional diffusion equation. Mathematics and Computers in Simulation, 209:87–101, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000800.

Yuan:2023:CPP

[1278] Wenping Yuan, Hui Liang, and Yanping Chen. On the convergence of piecewise polynomial collocation methods for variable-order space-fractional diffusion equations. Mathematics and Computers in Simulation, 209: 102–117, July 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000861.

Curletto:2023:MIS

[1279] Chiara Curletto, Lorenza Bulla, Loris Canovi, Fiorenza Demicheli, and Ezio Venturino. A mathematical investigation for the simulation and forecasting of a biodigester operations. Mathematics and Computers in Simulation, 209:118–152, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000617.

Haider:2023:IAT

[1280] Muhammad Imran Haider, Tariq Shah, Asif Ali, Dawood Shah, and Ijaz Khalid. An Innovative approach towards image encryption by using novel PRNs and S-boxes Modeling techniques. Mathematics and Computers in Simulation, 209:153– 168, July 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423000502.

Guo:2023:TEE

[1281] Yantao Guo and Yayun Fu. Two efficient exponential energy-preserving schemes for the fractional Klein– Gordon Schrödinger equation. Mathematics and Computers in Simulation, 209:169–183, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000794.

Tanveer:2023:MSM

[1282] Muhammad Tanveer, Waqas Nazeer, and Krzysztof Gdawiec. On the Mandelbrot set of z^{p} + log c^{t} via the Mann and Picard–Mann iterations. Mathematics and Computers in Simulation, 209:184–204, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542300085X.

Beler:2023:NSH

[1283] Ayşe Beler, Gökçe Özaltun Şimşek, and Sevin Gümgüm. Numerical solutions of the HIV infection model of CD4(+) cells by Laguerre wavelets. Mathematics and Computers in Simulation, 209:205–219, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000903.

He:2023:AFA

[1284] Jing He, Shuhua Mao, and Yuxiao Kang. Augmented fractional accumulation grey model and its application: Class ratio and restore error perspectives. Mathematics and Computers in Simulation, 209:220– 247, July 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000812.

Chen:2023:TGF

[1285] Hao Chen, Omid Nikan, Wenlin Qiu, and Zakieh Avazzadeh. Two-grid finite difference method for 1D fourth-order Sobolev-type equation with Burgers’ type nonlinearity. Mathematics and Computers in Simulation, 209:248–266, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000873.

Lakshmi:2023:RES

[1286] K. Pooja Lakshmi and T. Senthilkumar. Robust exponential synchronization results for uncertain infinite time varying distributed delayed neural networks with flexible delayed impulsive control. Mathematics and Computers in Simulation, 209:267–281, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000824.

Dupac:2023:MMS

[1287] Mihai Dupac. Mathematical modeling and simulation of the inverse kinematic of a redundant robotic manipulator using azimuthal angles and spherical polar piecewise interpolation. Mathematics and Computers in Simulation, 209:282–298, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000836.

Saha:2023:RPE

[1288] Sangeeta Saha, Debgopal Sahoo, and Guruprasad Samanta. Role of predation efficiency in prey–predator dynamics incorporating switching effect. Mathematics and Computers in Simulation, 209:299–323, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423000897.

Korkut:2023:AEN

[1289] Sıla Övgü Korkut. An accurate and efficient numerical solution for the generalized Burgers–Huxley equation via Taylor wavelets method: Qualitative analyses and Applications. Mathematics and Computers in Simulation, 209:324–341, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000927.

Rower:2023:CGM

[1290] David A. Rower and Paul J. Atzberger. Coarse-grained methods for heterogeneous vesicles with phase-separated domains: Elastic mechanics of shape fluctuations, plate compression, and channel insertion. Mathematics and Computers in Simulation, 209:342– 361, July 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000939.

Pomorski:2023:EBF

[1291] Krzysztof Pomorski. Equivalence between finite state stochastic machine, non-dissipative and dissipative tightbinding and Schrödinger model. Mathematics and Computers in Simulation, 209:362–407, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000915.

Rani:2023:QMO

[1292] Shobha Rani, Madhu Jain, and Rakesh Kumar Meena. Queueing modeling and optimization of a faulttolerant system with reboot, recovery, and vacationing server operating under admission control policy. Mathematics and Computers in Simulation, 209:408–425, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000885.

Saha:2023:AIM

[1293] Subrata Saha, Biswajit Sarkar, and Mitali Sarkar. Application of improved meta-heuristic algorithms for green preservation technology management to optimize dynamical investments and replenishment strategies. Mathematics and Computers in Simulation, 209:426–450, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000642.

Anonymous:2023:EBg

[1294] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 209:451, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001520.

Anonymous:2023:ICEh

[1295] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 209:452, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475423001532.

Anonymous:2023:EBh

[1296] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 209:ifc, July 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001477.

Shahni:2023:ENA

[1297] Julee Shahni, Randhir Singh, and Carlo Cattani. An efficient numerical approach for solving threepoint Lane–Emden–Fowler boundary value problem. Mathematics and Computers in Simulation, 210: 1–16, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423001143.

Anonymous:2023:PAb

[1298] Anonymous. Pages 1–680 (August 2023). Mathematics and Computers in Simulation, 210:1–680, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Villafuerte:2023:SPS

[1299] L. Villafuerte. Solution processes for second-order linear fractional differential equations with random inhomogeneous parts. Mathematics and Computers in Simulation, 210: 17–48, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001064.

Heryudono:2023:APU

[1300] Alfa Heryudono and Mehdi Raessi.

Adaptive partition of unity interpolation method with moving patches. Mathematics and Computers in Simulation, 210:49–65, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423001118.

Meng:2023:BRL

[1301] Xianhe Meng, Xian Zhang, and Yantao Wang. Bounded real lemmas and exponential H_{∞} control for memristorbased neural networks with unbounded time-varying delays. Mathematics and Computers in Simulation, 210: 66–81, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001192.

Mangal:2023:FOD

[1302] Shiv Mangal, O. P. Misra, and Joydip Dhar. Fractional-order deterministic epidemic model for the spread and control of HIV/AIDS with special reference to Mexico and India. Mathematics and Computers in Simulation, 210:82–102, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001131.

Alnahdi:2023:CST

[1303] Abeer S. Alnahdi, Saleem Nasir, and Taza Gul. Couple stress ternary hybrid nanofluid flow in a contraction channel by means of drug delivery function. Mathematics and Computers in Simulation, 210:103– 119, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423000940.

Amer:2023:SVM

[1304] T. S. Amer, F. M. El-Sabaa, A. A. Sallam, and I. M. Abady. Studying the vibrational motion of a rotating symmetrically charged solid body subjected to external forces and moments. Mathematics and Computers in Simulation, 210:120–146, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001088.

Nagata:2023:SPF

[1305] Takuto Nagata and Shuji Yoshikawa. Structure-preserving finite difference scheme for 1D thermoviscoelastoplastic equations under uniformly distributed temperature. Mathematics and Computers in Simulation, 210: 147–168, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001076.

Hou:2023:HOC

[1306] Bo Hou and Chengjian Zhang. A high-order compact difference method and its Richardson extrapolation for semi-linear reaction–diffusion equations with piecewise continuous argument in diffusion term. Mathematics and Computers in Simulation, 210:169–183, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001180.

El-Amrani:2023:NRK

[1307] Mofdi El-Amrani, Loubna Salhi, and Mohammed Seaid. A new Runge–Kutta–Chebyshev Galerkincharacteristic finite element method for advection–dispersion problems in anisotropic porous media. Mathematics and Computers in Simulation, 210:184–206, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000460.

Zeghlache:2023:AFF

[1308] Samir Zeghlache, Mohammed Zinelaabidine Ghellab, Ali Djerioui, Brahim Bouderah, and Mohamed Fouad Benkhoris. Adaptive fuzzy fast terminal sliding mode control for inverted pendulumcart system with actuator faults. Mathematics and Computers in Simulation, 210:207–234, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001106.

Wu:2023:BSC

[1309] Zhi-Jia Wu and Shou-Fu Tian. Breather-to-soliton conversions and their mechanisms of the (2 + 1)- dimensional generalized Hirota–Satsuma– Ito equation. Mathematics and Computers in Simulation, 210:235– 259, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423001222.

Xu:2023:SOP

[1310] Xinsheng Xu, Ping Ji, and Shuming Sang. Supply option purchasing decisions via mismatch cost minimization. Mathematics and Computers in Simulation, 210:260– 280, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542300126X.

Rhaima:2023:UHS

[1311] Mohamed Rhaima. Ulam–Hyers stability for an impulsive Caputo–Hadamard fractional neutral stochastic differential equations with infinite delay. Mathematics and Computers in Simulation, 210:281–295, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001258.

Ouyang:2023:APA

[1312] Haibin Ouyang, Jianhong Chen, Steven Li, Jianhua Xiang, and Zhi-Hui Zhan. Altruistic population algorithm: a metaheuristic search algorithm for solving multimodal multi-objective optimization problems. Mathematics and Computers in Simulation, 210: 296–319, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300109X.

Boregowda:2023:ASF

[1313] Parikshit Boregowda and Gui-Rong Liu. On the accuracy of SPH formulations with boundary integral terms. Mathematics and Computers in Simulation, 210:320– 345, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001234.

Anand:2023:NMM

[1314] Monalisa Anand, P. Danumjaya, and P. Raja Sekhara Rao. A nonlinear mathematical model on the Covid-19 transmission pattern among diabetic and non-diabetic population. Mathematics and Computers in Simulation, 210:346–369, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001210.

Chapwanya:2023:SOP

[1315] Michael Chapwanya and Phindile Dumani. Stationary and oscillatory patterns in microbial population under environmental stress. Mathematics and Computers in Simulation, 210:370–383, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001271.

Zapata:2023:CSO

[1316] M. Uh Zapata, R. Itza Balam, and J. Montalvo-Urquizo. A compact sixth-order implicit immersed interface method to solve 2D Poisson equations with discontinuities. Mathematics and Computers in Simulation, 210:384–407, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001179.

Tomar:2023:ETB

[1317] Saurabh Tomar, Soniya Dhama, Higinio Ramos, and Mehakpreet Singh. An efficient technique based on Green’s function for solving two-point boundary value problems and its convergence analysis. Mathematics and Computers in Simulation, 210:408– 423, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423001209.

Almushaira:2023:EFO

[1318] Mustafa Almushaira. An efficient fourth-order accurate conservative scheme for Riesz space fractional Schrödinger equation with wave operator. Mathematics and Computers in Simulation, 210:424– 447, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001246.

Chen:2023:CRC

[1319] Naixi Chen and Hong Fan. Credit risk contagion and optimal dual control — an SIS/R model. Mathematics and Computers in Simulation, 210:448–472, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001362.

Truong:2023:FAF

[1320] Dinh-Nhat Truong and Jui-Sheng Chou. Fuzzy adaptive forensicbased investigation algorithm for optimizing frequency-constrained structural dome design. Mathematics and Computers in Simulation, 210:473– 531, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542300112X.

Zhang:2023:HOS

[1321] Xi Zhang, Maohua Ran, Yang Liu, and Li Zhang. A high-order structurepreserving difference scheme for generalized fractional Schrödinger equation with wave operator. Mathematics and Computers in Simulation, 210:532–546, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001325.

Raw:2023:SSI

[1322] Sharada Nandan Raw and Sevak Ram Sahu. Strong stability with impact of maturation delay and diffusion on a toxin producing phytoplankton– zooplankton model. Mathematics and Computers in Simulation, 210:547– 570, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423001283.

J:2023:OCS

[1323] Kokila J., Vellappandi M., Meghana D., and Govindaraj V. Optimal control study on Michaelis–Menten kinetics — a fractional version. Mathematics and Computers in Simulation, 210:571–592, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001386.

Amat:2023:NIR

[1324] Sergio Amat, Zhilin Li, Juan Ruiz-Álvarez, Concepción Solano, and Juan C. Trillo. Numerical integration rules with improved accuracy close to discontinuities. Mathematics and Computers in Simulation, 210:593–614, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001374.

Manzini:2023:TTM

[1325] G. Manzini, P. M. D. Truong, R. Vuchkov, and B. Alexandrov. The tensor-train mimetic finite difference method for three-dimensional Maxwell’s wave propagation equations. Mathematics and Computers in Simulation, 210:615–639, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423001313.

Liu:2023:SSL

[1326] Chein-Shan Liu and Botong Li. Solving Sturm–Liouville inverse problems by an orthogonalized enhanced boundary function method and a product formula for symmetric potential. Mathematics and Computers in Simulation, 210:640–660, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001301.

Mukama:2023:PEB

[1327] Denis Sospeter Mukama, Mohammad Ghani, and Isambi Sailon Mbalawata. Persistence, Extinction, and boundedness in p th moment of hybrid stochastic logistic systems by delay feedback control based on discretetime observation. Mathematics and Computers in Simulation, 210:661– 677, August 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001398.

Anonymous:2023:NIf

[1328] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 210:678, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542300188X.

Anonymous:2023:ICEi

[1329] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 210:679, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475423001891.

Anonymous:2023:EBi

[1330] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 210:ifc, August 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001830.

Naik:2023:CCS

[1331] Manisha Krishna Naik, Chandrali Baishya, and P. Veeresha. A chaos control strategy for the fractional 3D Lotka–Volterra like attractor. Mathematics and Computers in Simulation, 211:1–22, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001428.

Anonymous:2023:PS

[1332] Anonymous. Pages 1–516 (September 2023). Mathematics and Computers in Simulation, 211:1–516, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Atrout:2023:DMC

[1333] Sabah Atrout, Md. Abdullah Al Mahbub, and Haibiao Zheng. Decoupled modified characteristics variational multiscale method for solving the blood solute dynamics model. Mathematics and Computers in Simulation, 211:23–56, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001404.

Jain:2023:AMO

[1334] Madhu Jain and Sibasish Dhibar. AN-FIS and metaheuristic optimization for strategic joining policy with reattempt and vacation. Mathematics and Computers in Simulation, 211: 57–84, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001295.

Aggarwal:2023:XME

[1335] Sakshi Aggarwal and Krishn K.

Mishra. X-MODE: Extended Multioperator Differential Evolution algorithm. Mathematics and Computers in Simulation, 211:85–108, September 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423000290.

Cao:2023:EAF

[1336] Dewei Cao and Hu Chen. Error analysis of a finite difference method for the distributed order subdiffusion equation using discrete comparison principle. Mathematics and Computers in Simulation, 211:109– 117, September 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423001659.

Kumar:2023:NSP

[1337] Anil Kumar and Pentyala Srinivasa Rao. Numerical study of periodically heated wall effect on natural convection in an enclosure. Mathematics and Computers in Simulation, 211: 118–133, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423000083.

Suo:2023:DDI

[1338] JingJing Suo, Hongxiao Hu, and Liguang Xu. Delay-dependent impulsive control for lag quasisynchronization of stochastic complex dynamical networks. Mathematics and Computers in Simulation, 211: 134–153, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001453.

Xu:2023:FTS

[1339] Yibin Xu, Yanqin Liu, Xiuling Yin, Libo Feng, Zihua Wang, and Qiuping Li. A fast time stepping Legendre spectral method for solving fractional Cable equation with smooth and non-smooth solutions. Mathematics and Computers in Simulation, 211: 154–170, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001635.

Fu:2023:NOG

[1340] Hao Fu, Jian Xia, and Shuling Tian.

A new overset grid assembly strategy for dynamic grid systems. Mathematics and Computers in Simulation, 211: 171–191, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300143X.

Sahoo:2023:ASO

[1341] Sanjay Ku Sahoo and Vikas Gupta.

An almost second-order robust computational technique for singularly perturbed parabolic turning point problem with an interior layer. Mathematics and Computers in Simulation, 211: 192–213, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001660.

Sanga:2023:FMC

[1342] Sudeep Singh Sanga and Gannamaneni Sai Charan. Fuzzy modeling and cost optimization for machine repair problem with retrial under admission control F -policy and feedback. Mathematics and Computers in Simulation, 211:214–240, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001416.

Njike-Tchaptchet:2023:MMU

[1343] Eric Rostand Njike-Tchaptchet and Calvin Tadmon. Mathematical modeling of the unemployment problem in a context of financial crisis. Mathematics and Computers in Simulation, 211:241–262, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001696.

Akahori:2023:COB

[1344] Jirô Akahori, Jie Yen Fan, and Yuri Imamura. On the convergence order of a binary tree approximation of symmetrized diffusion processes. Mathematics and Computers in Simulation, 211:263–277, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001350.

Abbaspour:2023:EAP

[1345] Manijeh Abbaspour, Kianoush Fathi Vajargah, and Parvin Azhdari. An efficient algorithm for pricing reinsurance contract under the regimeswitching model. Mathematics and Computers in Simulation, 211:278– 300, September 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423001738.

Naranjo-Alvarez:2023:VEM

[1346] S. Naranjo-Alvarez, L. Beirão da Veiga, V. A. Bokil, F. Dassi, V. Gyrya, and G. Manzini. The virtual element method for a 2D incompressible MHD system. Mathematics and Computers in Simulation, 211: 301–328, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001349.

Chang:2023:NMS

[1347] Shuang Chang, Yantao Wang, Xian Zhang, and Xin Wang. A new method to study global exponential stability of inertial neural networks with multiple time-varying transmission delays. Mathematics and Computers in Simulation, 211:329–340, September 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001623.

Meng:2023:DBC

[1348] Qingyan Meng, Yejuan Wang, and Peter E. Kloeden. The dynamical behavior of a class of stochastic vegetation models. Mathematics and Computers in Simulation, 211: 341–367, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001672.

Roy:2023:MDI

[1349] Jyotirmoy Roy, Subrata Dey, and Malay Banerjee. Maturation delay induced stability enhancement and shift of bifurcation thresholds in a predator–prey model with generalist predator. Mathematics and Computers in Simulation, 211:368– 393, September 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542300174X.

Li:2023:AST

[1350] Yuyu Li, Tongke Wang, and Guang hua Gao. The asymptotic solutions of twoterm linear fractional differential equations via Laplace transform. Mathematics and Computers in Simulation, 211:394–412, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001647.

Vijayakumar:2023:OTD

[1351] M. Vijayakumar, R. Sakthivel, F. Kong, and S. Marshal Anthoni. Output tracking and disturbance rejection control for switched systems under asynchronous switching. Mathematics and Computers in Simulation, 211: 413–429, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300160X.

Shi:2023:CNF

[1352] Fangfang Shi, Guoju Ye, Wei Liu, and Dafang Zhao. A class of nonconvex fuzzy optimization problems under granular differentiability concept. Mathematics and Computers in Simulation, 211:430–444, September 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001763.

Yuan:2023:STT

[1353] Long Yuan and Wenxiu Gong. A space–time Trefftz DG scheme for the time-dependent Maxwell equations in anisotropic media. Mathematics and Computers in Simulation, 211: 445–469, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001702.

Delagnes:2023:CRB

[1354] T. Delagnes, T. Henneron, S. Clenet, M. Fratila, and J. P. Ducreux. Comparison of reduced basis construction methods for Model Order Reduction, with application to non-linear low frequency electromagnetics. Mathematics and Computers in Simulation, 211: 470–488, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001611.

Ju:2023:MMV

[1355] Yuejuan Ju, Jiye Yang, Zhiyong Liu, and Qiuyan Xu. Meshfree methods for the variable-order fractional advection–diffusion equation. Mathematics and Computers in Simulation, 211:489–514, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001441.

Anonymous:2023:NIg

[1356] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 211:515, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002318.

Anonymous:2023:ICEj

[1357] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 211:516, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S037847542300232X.

Anonymous:2023:EBj

[1358] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 211:ifc, September 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002264.

Liu:2023:MBR

[1359] Yang Liu, Zhenzhen Zhang, Hao Chen, and Shouming Zhong. A memory behavior related hybrid event-triggered mechanism for an improved robust control on neural networks. Mathematics and Computers in Simulation, 212:1–20, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001787.

Anonymous:2023:PO

[1360] Anonymous. Pages 1–586 (October 2023). Mathematics and Computers in Simulation, 212:1–586, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Wang:2023:FOF

[1361] Yuan-Ming Wang and Bo Xie. A fourth-order fractional Adams-type implicit–explicit method for nonlinear fractional ordinary differential equations with weakly singular solutions. Mathematics and Computers in Simulation, 212:21– 48, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001726.

Zhang:2023:AQS

[1362] Hai Zhang, Xinbin Chen, Renyu Ye, Ivanka Stamova, and Jinde Cao. Adaptive quasi-synchronization analysis for Caputo delayed Cohen–Grossberg neural networks. Mathematics and Computers in Simulation, 212:49– 65, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001805.

Das:2023:IIS

[1363] Debabrata Das, T. K. Kar, and Debprasad Pal. The impact of invasive species on some ecological services in a harvested predator–prey system. Mathematics and Computers in Simulation, 212:66–90, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001799.

Ye:2023:CAF

[1364] Wenwen Ye and Shengping Li. Convergence analysis of flow direction algorithm in continuous search space and its improvement. Mathematics and Computers in Simulation, 212:91–121, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423001751.

Aliyu:2023:AIE

[1365] Major Murtala Bello Aliyu, Ali Audu Baidu, Bala Ma’aji Abdulhamid, Mohammed Olanrewaju Ibrahim, and Fu’ad Muhammad Mukhtar. Assessing the impact of escalating attacks on soft targets by criminal gang: a modelling viewpoint using bifurcation analysis. Mathematics and Computers in Simulation, 212: 122–137, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002008.

Li:2023:FEA

[1366] Yiqun Li and Hong Wang. A finite element approximation to a viscoelastic Euler–Bernoulli beam with internal damping. Mathematics and Computers in Simulation, 212:138– 158, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542300201X.

Chen:2023:DAH

[1367] Chong Chen and Yinggao Zhou. Dynamic analysis of HIV model with a general incidence, CTLs immune response and intracellular delays. Mathematics and Computers in Simulation, 212:159–181, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001970.

Singh:2023:CPE

[1368] Saroja Kumar Singh, Sarat Kumar Acharya, F. R. B. Cruz, and André L. F. Cançado. Change point estimation in an M/M/2 queue with heterogeneous servers. Mathematics and Computers in Simulation, 212: 182–194, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001775.

Onay:2023:NIC

[1369] Funda Kutlu Onay. A novel improved chef-based optimization algorithm with Gaussian random walk-based diffusion process for global optimization and engineering problems. Mathematics and Computers in Simulation, 212:195–223, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001957.

Odibat:2023:NFD

[1370] Zaid Odibat and Dumitru Baleanu.

A new fractional derivative operator with generalized cardinal sine kernel: Numerical simulation. Mathematics and Computers in Simulation, 212:224–233, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002033.

Umar:2023:HCS

[1371] Muhammad Umar, Zulqurnain Sabir, Muhammad Asif Zahoor Raja, Haci Mehmet Baskonus, Mohamed R. Ali, and Nehad Ali Shah. Heuristic computing with sequential quadratic programming for solving a nonlinear hepatitis B virus model. Mathematics and Computers in Simulation, 212: 234–248, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002045.

Zhang:2023:FDL

[1372] Xin Zhang, Danxia Wang, Jianwen Zhang, and Hongen Jia. Fully decoupled linear BDF2 scheme for the penalty incompressible Ericksen– Leslie equations. Mathematics and Computers in Simulation, 212:249– 266, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423002069.

Li:2023:IUE

[1373] Jiyong Li and Hongyu Fang. Improved uniform error bounds of a time-splitting Fourier pseudospectral scheme for the Klein–Gordon-Schrödinger equation with the small coupling constant. Mathematics and Computers in Simulation, 212:267– 288, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002021.

Wang:2023:SSE

[1374] Qi Wang, Kainan Xiang, Chunhui Zhu, and Lang Zou. Stochastic SEIR epidemic models with virus mutation and logistic growth of susceptible populations. Mathematics and Computers in Simulation, 212: 289–309, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002057.

Wang:2023:NCD

[1375] Shaochen Wang and Christian H.

Weiß. New characterizations of the (discrete) Lindley distribution and their applications. Mathematics and Computers in Simulation, 212:310– 322, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002082.

Lu:2023:ASM

[1376] Ziqiang Lu and Yuanguo Zhu. Asymptotic stability in p-th moment of uncertain dynamical systems with time-delays. Mathematics and Computers in Simulation, 212:323–335, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002100.

Sun:2023:UCS

[1377] Tao Sun and Yihong Wang. Uniformly convergent scheme for steady MHD duct flow problems with high Hartmann numbers on structured and unstructured grids. Mathematics and Computers in Simulation, 212:336– 359, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002124.

Singh:2023:PUN

[1378] Satpal Singh and Devendra Kumar. Parameter uniform numerical method for a system of singularly perturbed parabolic convection– diffusion equations. Mathematics and Computers in Simulation, 212:360– 381, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002094.

Pei:2023:ASD

[1379] Yongzhen Pei, Na Shen, Jingjing Zhao, Yuping Yu, and Yasong Chen. Analysis and simulation of a delayed HIV model with reaction–diffusion and sliding control. Mathematics and Computers in Simulation, 212: 382–405, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300215X.

Postavaru:2023:ENM

[1380] Octavian Postavaru. An efficient numerical method based on Fibonacci polynomials to solve fractional differential equations. Mathematics and Computers in Simulation, 212:406– 422, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001969.

Wang:2023:QRM

[1381] Yan Wang and Zhi Qian. A quasireversibility method for solving a two-dimensional time-fractional inverse heat conduction problem. Mathematics and Computers in Simulation, 212:423–440, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002173.

Wang:2023:ROS

[1382] Zhao-Hong Wang, Yao-Lin Jiang, and Kang-Li Xu. Reduced-order statespace models for two-dimensional discrete systems via bivariate discrete orthogonal polynomials. Mathematics and Computers in Simulation, 212: 441–456, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002148.

Dwivedi:2023:ERB

[1383] Pankaj Prasad Dwivedi and Dilip Kumar Sharma. Evaluation and ranking of battery electric vehicles by Shannon’s entropy and TOPSIS methods. Mathematics and Computers in Simulation, 212:457–474, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002185.

Sadek:2023:NAB

[1384] Lakhlifa Sadek, Ahmad Sami Bataineh, Osman Rasit Isik, Hamad Talibi Alaoui, and Ishak Hashim. A numerical approach based on Bernstein collocation method: Application to differential Lyapunov and Sylvester matrix equations. Mathematics and Computers in Simulation, 212:475– 488, October 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423002161.

Chachra:2023:IFA

[1385] Aayushi Chachra, Akshay Kumar, and Mangey Ram. Intuitionistic fuzzy approach to reliability assessment of multi-state systems. Mathematics and Computers in Simulation, 212: 489–503, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002197.

Doostaki:2023:NSC

[1386] Reza Doostaki, Mohammad Mehdi Hosseini, and Abbas Salemi. A new simultaneously compact finite difference scheme for high-dimensional time-dependent PDEs. Mathematics and Computers in Simulation, 212: 504–523, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002136.

Man:2023:HOS

[1387] Shumin Man, Qiang Gao, and Wanxie Zhong. High order symmetric algorithms for nonlinear dynamical systems with non-holonomic constraints. Mathematics and Computers in Simulation, 212:524–547, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002215.

Santra:2023:RMT

[1388] Nirapada Santra, Sangeeta Saha, and Guruprasad Samanta. Role of multiple time delays on a stagestructured predator–prey system in a toxic environment. Mathematics and Computers in Simulation, 212: 548–583, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002203.

Anonymous:2023:ICEk

[1389] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 212:584, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423002604.

Anonymous:2023:NIh

[1390] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 212:585, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002616.

Anonymous:2023:EBk

[1391] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 212:ifc, October 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002550.

Maji:2023:NIP

[1392] Sandip Maji and Srinivasan Natesan. Nonsymmetric interior penalty Galerkin method for nonlinear timefractional integro-partial differential equations. Mathematics and Computers in Simulation, 213:1–17, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423002392.

Anonymous:2023:PN

[1393] Anonymous. Pages 1–528 (November 2023). Mathematics and Computers in Simulation, 213:1–528, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Badia:2023:MPI

[1394] Santiago Badia, Jerrad Hampton, and Javier Principe. A massively parallel implementation of multilevel Monte Carlo for finite element models. Mathematics and Computers in Simulation, 213:18–39, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002239.

Gajbhiye:2023:MMA

[1395] Sneha Gajbhiye, Arundhati Warke, and Katta Ramesh. Mathematical modeling and analysis of immiscible metallic based nanofluid flow in a microchannel with nonspherical nanoparticles. Mathematics and Computers in Simulation, 213: 40–54, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002410.

Wang:2023:EHS

[1396] Shikun Wang and Yuanshi Wang.

Effects of heterogeneity and spatial pattern in multi-patch consumerresource systems with asymmetric diffusion. Mathematics and Computers in Simulation, 213:55–77, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002458.

Kumar:2023:SBA

[1397] Arjun Kumar, Ashvini Gupta, Uma S.

Dubey, and Balram Dubey. Stability and bifurcation analysis of an infectious disease model with different optimal control strategies. Mathematics and Computers in Simulation, 213:78–114, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002434.

Choe:2023:ODL

[1398] Yong Son Choe, Yong Jun Kim, Thae Nam Ri, and Tong Kuk Kim. One-dimensional lattice Boltzmann simulation of parallel plate dielectric barrier discharge plasma in atmospheric argon. Mathematics and Computers in Simulation, 213:115– 126, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423002240.

Yang:2023:TSI

[1399] Shiyuan Yang, Xing Liu, and Meng Zhang. Threshold stability of an improved IMEX numerical method based on conservation law for a nonlinear advection–diffusion Lotka– Volterra model. Mathematics and Computers in Simulation, 213:127– 144, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423002525.

Hosseinzadeh:2023:ORS

[1400] Hossein Hosseinzadeh and Ahmad Shirzadi. On optimal radius of subdomains in meshless LBIE method. Mathematics and Computers in Simulation, 213:145–160, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475423002495.

Borhanifar:2023:SID

[1401] A. Borhanifar, S. Shahmorad, E. Feizi, and D. Baleanu. Solving 2D-integro-differential problems with nonlocal boundary conditions via a matrix formulated approach. Mathematics and Computers in Simulation, 213: 161–176, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002641.

Aboussi:2023:HEF

[1402] Wassim Aboussi, Moussa Ziggaf, Imad Kissami, and Mohamed Boubekeur. A highly efficient finite volume method with a diffusion control parameter for hyperbolic problems. Mathematics and Computers in Simulation, 213: 177–193, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002422.

Cui:2023:ADI

[1403] Mingrong Cui. An alternating direction implicit compact finite difference scheme for the multi-term time-fractional mixed diffusion and diffusion–wave equation. Mathematics and Computers in Simulation, 213: 194–210, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300246X.

Chambon:2023:MPD

[1404] Camille Chambon, Olivier Piller, and Iraj Mortazavi. Modeling of pressuredependent background leakages in water distribution networks. Mathematics and Computers in Simulation, 213: 211–236, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002513.

Kumari:2023:REB

[1405] Rani Kumari, Yogesh Mani Tripathi, Rajesh Kumar Sinha, and Liang Wang. Reliability estimation for bathtubshaped distribution under block progressive censoring. Mathematics and Computers in Simulation, 213:237– 260, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002501.

Uzunca:2023:ROM

[1406] Murat Uzunca and Bülent Karasözen. Reduced-order modeling for Ablowitz– Ladik equation. Mathematics and Computers in Simulation, 213:261– 273, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002677.

Noormohammadi:2023:MMS

[1407] Nima Noormohammadi, Danial Afifi, Bijan Boroomand, and Omid Bateniparvar. A meshfree method for the solution of 2D and 3D second order elliptic boundary value problems in heterogeneous media. Mathematics and Computers in Simulation, 213: 274–301, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002471.

M:2023:NNB

[1408] Sivalingam S. M., Pushpendra Kumar, and Venkatesan Govindaraj. A neural networks-based numerical method for the generalized Caputo-type fractional differential equations. Mathematics and Computers in Simulation, 213:302–323, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002665.

Bukhari:2023:DAN

[1409] Ayaz Hussain Bukhari, Muhammad Shoaib, Adiqa Kausar Kiani, Naveed Ishtiaq Chaudhary, Muhammad Asif Zahoor Raja, and Chi-Min Shu. Dynamical analysis of nonlinear fractional order Lorenz system with a novel design of intelligent solution predictive radial base networks. Mathematics and Computers in Simulation, 213:324–347, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002483.

Taneja:2023:RHO

[1410] Komal Taneja, Komal Deswal, and Devendra Kumar. A robust higherorder numerical technique with graded and harmonic meshes for the timefractional diffusion–advection–reaction equation. Mathematics and Computers in Simulation, 213:348–373, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002537.

Lenz:2023:FIK

[1411] David Lenz, Oana Marin, Vijay Mahadevan, Raine Yeh, and Tom Peterka. Fourier-informed knot placement schemes for B-spline approximation. Mathematics and Computers in Simulation, 213:374–393, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423002227.

Zhang:2023:SFA

[1412] Wei Zhang, Ke Pan, Shigang Li, and Yagang Wang. Special Forces Algorithm: a novel meta-heuristic method for global optimization. Mathematics and Computers in Simulation, 213: 394–417, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002690.

Braverman:2023:NPB

[1413] E. Braverman and A. Rodkina. Noisy prediction-based control leading to stability switch. Mathematics and Computers in Simulation, 213:418– 443, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423002689.

Kumar:2023:CSM

[1414] Pushpendra Kumar, M. Vellappandi, Zareen A. Khan, Sivalingam S. M, Anthony Kaziboni, and V. Govindaraj. A case study of monkeypox disease in the United States using mathematical modeling with real data. Mathematics and Computers in Simulation, 213:444–465, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002707.

Liu:2023:MLH

[1415] Zhendong Liu, Yiming Fang, Le Liu, and Shuidong Ma. A multi-leader Harris hawks optimizer with adaptive mutation and its application for modeling of silicon content in liquid iron of blast furnace. Mathematics and Computers in Simulation, 213: 466–514, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002756.

Xing:2023:IIA

[1416] Shuangyun Xing, Yifan Liu, and Da-Yan Liu. An improved iterative algorithm for solving optimal tracking control problems of stochastic systems. Mathematics and Computers in Simulation, 213:515–526, November 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002744.

Anonymous:2023:NIi

[1417] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 213:527, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003051.

Anonymous:2023:ICEl

[1418] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 213:528, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475423003063.

Anonymous:2023:EBl

[1419] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 213:ifc, November 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003002.

Wu:2023:MAA

[1420] Peng Wu, Ran Zhang, and Anwarud Din. Mathematical analysis of an agesince infection and diffusion HIV/AIDS model with treatment adherence and Dirichlet boundary condition. Mathematics and Computers in Simulation, 214:1–27, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002719.

Anonymous:2023:PD

[1421] Anonymous. Pages 1–430 (December 2023). Mathematics and Computers in Simulation, 214:1–430, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tu:2023:RDE

[1422] Yunbo Tu and Xinzhu Meng. A reaction–diffusion epidemic model with virus mutation and media coverage: Theoretical analysis and numerical simulation. Mathematics and Computers in Simulation, 214: 28–67, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002781.

Ye:2023:SMM

[1423] Luhong Ye, Hongyong Zhao, and Daiyong Wu. Spatial movement with memory-induced cross-diffusion effect and toxin effect in predator. Mathematics and Computers in Simulation, 214:68–91, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300277X.

Li:2023:AHO

[1424] Bin Li, Hongchao Kang, Songliang Chen, and Shanjing Ren. On the approximation of highly oscillatory Volterra integral equations of the first kind via Laplace transform. Mathematics and Computers in Simulation, 214:92–113, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002732.

Yang:2023:TDA

[1425] Jin Yang, Zhuo Chen, Yuanshun Tan, Zijian Liu, and Robert A. Cheke. Threshold dynamics of an age-structured infectious disease model with limited medical resources. Mathematics and Computers in Simulation, 214:114–132, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002823.

Guirao:2023:ANC

[1426] Juan Luis García Guirao, Mansoor Alsulami, Haci Mehmet Baskonus, Esin Ilhan, and P. Veeresha. Analysis of nonlinear compartmental model using a reliable method. Mathematics and Computers in Simulation, 214: 133–151, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300280X.

Yadav:2023:PDE

[1427] K. S. Yadav and D. C. Dalal. Penetration and distribution efficacy of chemotherapeutic drugs in biological tissues: a computational investigation. Mathematics and Computers in Simulation, 214:152–171, December 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002859.

Xue:2023:DEA

[1428] Leilei Xue, Liping Sun, and Songbai Guo. Dynamic effects of asymptomatic infections on malaria transmission. Mathematics and Computers in Simulation, 214:172–182, December 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002835.

Priyadarshana:2023:ITA

[1429] S. Priyadarshana, J. Mohapatra, and S. R. Pattanaik. An improved time accurate numerical estimation for singularly perturbed semilinear parabolic differential equations with small space shifts and a large time lag. Mathematics and Computers in Simulation, 214:183–203, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002896.

V:2023:PAM

[1430] Saravanan V., Poongothai V., and Godhandaraman P. Performance analysis of a multi server retrial queueing system with unreliable server, discouragement and vacation model. Mathematics and Computers in Simulation, 214:204–226, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002884.

Irshad:2023:DAT

[1431] M. R. Irshad, P. Jodrá, A. Krishna, and R. Maya. On the discrete analogue of the Teissier distribution and its associated INAR(1) process. Mathematics and Computers in Simulation, 214:227–245, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002872.

Shirazian:2023:NAV

[1432] Mohammad Shirazian. A new acceleration of variational iteration method for initial value problems. Mathematics and Computers in Simulation, 214: 246–259, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002811.

Hisano:2023:DPE

[1433] Kanato Hisano and Nozomu Matsuura. Discrete pendulum equation. Mathematics and Computers in Simulation, 214:260–271, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003105.

Zhao:2023:GFE

[1434] Zhengang Zhao and Yunying Zheng. A Galerkin finite element method for the space Hadamard fractional partial differential equation. Mathematics and Computers in Simulation, 214: 272–289, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002768.

Teng:2023:TCL

[1435] Ye Teng and Zhimin Zhang. On a time-changed Lévy risk model with capital injections and periodic observation. Mathematics and Computers in Simulation, 214:290–314, December 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423002938.

Derakhshan:2023:ENM

[1436] Mohammad Hossein Derakhshan, Hamid Rezaei, and Hamid Reza Marasi. An efficient numerical method for the distributed order timefractional diffusion equation with error analysis and stability. Mathematics and Computers in Simulation, 214: 315–333, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002975.

Botello-Aceves:2023:DMO

[1437] Salvador Botello-Aceves, Arturo Hernandez-Aguirre, and S. Ivvan Valdez. The Directed Multi-Objective Estimation Distribution Algorithm (D-MOEDA). Mathematics and Computers in Simulation, 214:334–351, December 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542300294X.

Kravitz:2023:LEM

[1438] H. Kravitz, M. Brio, and J.-G. Caputo. Localized eigenvectors on metric graphs. Mathematics and Computers in Simulation, 214:352– 372, December 2023. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002914.

Dai:2023:SFO

[1439] Houping Dai, Yingxin Feng, Xuedan Wei, Dongdong Chen, Zhoushun Zheng, and Jianzhong Wang. Space fractional-order modeling for the sintering process of metal fibers via Lattice Boltzmann method. Mathematics and Computers in Simulation, 214: 373–387, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003099.

Ahmed:2023:FST

[1440] Hoda F. Ahmed and W. A. Hashem. A fully spectral tau method for a class of linear and nonlinear variableorder time-fractional partial differential equations in multi-dimensions. Mathematics and Computers in Simulation, 214:388–408, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003178.

Gorbunova:2023:NAC

[1441] A. V. Gorbunova and A. V. Lebedev. Nonlinear approximation of characteristics of a fork–join queueing system with Pareto service as a model of parallel structure of data processing. Mathematics and Computers in Simulation, 214:409–428, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003233.

Anonymous:2023:NIj

[1442] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 214:429, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003622.

Anonymous:2023:ICEm

[1443] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 214:430, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423003634.

Anonymous:2023:EBm

[1444] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 214:ifc, December 2023. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003579.

Fernandez-Diaz:2024:SSF

[1445] Julio M. Fernández-Díaz and César O.

Menéndez-Pérez. A superlinear Scaling Factor Regula Falsi root finder that detects the simple or multiple character of the root. Mathematics and Computers in Simulation, 215:1–20, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003294.

Anonymous:2024:PJa

[1446] Anonymous. Pages 1–640 (January 2024). Mathematics and Computers in Simulation, 215:1–640, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Garai:2024:DBP

[1447] Gobinda Garai and Bankim C. Mandal. Diagonalization based Parallelin-Time method for a class of fourth order time dependent PDEs. Mathematics and Computers in Simulation, 215:21–42, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003221.

Sahoo:2024:RHO

[1448] Sanjay Ku Sahoo, Vikas Gupta, and Shruti Dubey. A robust higher-order finite difference technique for a time-fractional singularly perturbed problem. Mathematics and Computers in Simulation, 215:43– 68, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003397.

Nguyen:2024:ISSa

[1449] Khanh Hieu Nguyen and Sung Hyun Kim. Improved stability and stabilization criteria of sampled-data control systems based on an enhanced looped-functional. Mathematics and Computers in Simulation, 215:69– 81, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003403.

Liu:2024:AFT

[1450] Xiaohua Liu, Mengling Li, and Pengyu Zeng. Adaptive finite-time neural network control for nonlinear stochastic systems with state constraints. Mathematics and Computers in Simulation, 215:82–97, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003300.

Hajri:2024:DDS

[1451] Youssra Hajri, Amina Allali, and Saida Amine. A delayed deterministic and stochastic SI_{R}I_{C}V model: Hopf bifurcation and stochastic analysis. Mathematics and Computers in Simulation, 215:98–121, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300321X.

Arezoomandan:2024:FEA

[1452] Mahdieh Arezoomandan and Ali R.

Soheili. Finite element approximation of the linearized stochastic Cahn–Hilliard equation with fractional Brownian motion. Mathematics and Computers in Simulation, 215:122– 145, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003270.

Inage:2024:PVO

[1453] Sin ichi Inage, Shouki Ohgi, and Yoshinori Takahashi. Proposal and validation of an optimization method using Monte Carlo method for multiobjective functions. Mathematics and Computers in Simulation, 215: 146–157, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003269.

Awati:2024:SHW

[1454] Vishwanath B. Awati, Akash Goravar, and Mahesh Kumar N. Spectral and Haar wavelet collocation method for the solution of heat generation and viscous dissipation in micropolar nanofluid for MHD stagnation point flow. Mathematics and Computers in Simulation, 215:158– 183, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003257.

Corbett:2024:SSI

[1455] Andrew Corbett, Jacqueline Christmas, Christopher Lawrence, and Joseph Feenan. Simulating spectral images of natural waters and submerged objects from a bird’s-eye view. Mathematics and Computers in Simulation, 215:184–207, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300318X.

Ha:2024:VTP

[1456] Mijin Ha, Donghyun Kim, and Ji-Hun Yoon. Valuing of timer pathdependent options. Mathematics and Computers in Simulation, 215:208– 227, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003361.

Mehrdoust:2024:VOP

[1457] Farshid Mehrdoust, Idin Noorani, and Juho Kanniainen. Valuation of option price in commodity markets described by a Markov-switching model: a case study of WTI crude oil market. Mathematics and Computers in Simulation, 215:228–269, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300335X.

Conti:2024:TBM

[1458] Costanza Conti and Mariantonia Cotronei. Triangular Bernstein moment-based identification of algebraic curves. Mathematics and Computers in Simulation, 215:270– 281, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003415.

Mrad:2024:OUP

[1459] Hatem Mrad and Seyyed Mojtaba Fakhari. Optimization of unconstrained problems using a developed algorithm of spectral conjugate gradient method calculation. Mathematics and Computers in Simulation, 215:282–290, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003208.

Li:2024:NSR

[1460] Nan Li, Xiaoping Wang, Huanying Xu, and Haitao Qi. Numerical study on radiative MHD flow of viscoelastic fluids with distributed-order and variableorder space fractional operators. Mathematics and Computers in Simulation, 215:291–305, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003117.

Zhang:2024:DSF

[1461] Xiulan Zhang, YiYu Liu, Hongling Qiu, and Heng Liu. Dissipativity and synchronization of fractional-order outputcoupled neural networks with multiple adaptive coupling weights. Mathematics and Computers in Simulation, 215:306–322, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003427.

Mishra:2024:PPC

[1462] Rupa Mishra and Tapas Kumar Saha. Predictive power control of PMSG based WECS: Development and implementation for smooth grid synchronisation, balanced and unbalanced grid. Mathematics and Computers in Simulation, 215:323– 337, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423003373.

Hwang:2024:FEN

[1463] Youngjin Hwang, Junxiang Yang, Gyeongyu Lee, Seokjun Ham, Seungyoon Kang, Soobin Kwak, and Junseok Kim. Fast and efficient numerical method for solving the Allen–Cahn equation on the cubic surface. Mathematics and Computers in Simulation, 215:338–356, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003191.

Wang:2024:ATB

[1464] Yu Wang, Yang Tian, Yida Guo, and Haoping Wang. Active torquebased gait adjustment multi-level control strategy for lower limb patient– exoskeleton coupling system in rehabilitation training. Mathematics and Computers in Simulation, 215: 357–381, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003464.

Jannelli:2024:FDM

[1465] Alessandra Jannelli. A finite difference method on quasi-uniform grids for the fractional boundarylayer Blasius flow. Mathematics and Computers in Simulation, 215:382– 398, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542300349X.

Raju:2024:SBA

[1466] C. S. K. Raju, H. Thameem Basha, N. F. M. Noor, Nehad Ali Shah, and Se-Jin Yook. Significance of body acceleration and gold nanoparticles through blood flow in an uneven/ composite inclined stenosis artery: a finite difference computation. Mathematics and Computers in Simulation, 215:399–419, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003324.

Zhan:2024:SNI

[1467] Qingyi Zhan, Jinqiao Duan, Xiaofan Li, and Yuhong Li. Symplectic numerical integration for Hamiltonian stochastic differential equations with multiplicative Lévy noise in the sense of Marcus. Mathematics and Computers in Simulation, 215: 420–439, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003385.

Sun:2024:TEC

[1468] Ruyi Sun, Jiaqi Chang, Hongmei Wang, Miaomiao Li, and Yongzheng Sun. Time and energy costs for synchronization of multi-layer networks. Mathematics and Computers in Simulation, 215:440–455, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423003488.

Assar:2024:GTS

[1469] Moein Assar and Brian Arthur Grimes. Geometrically transformed spectral methods to solve partial differential equations in circular geometries, application for multi-phase flow. Mathematics and Computers in Simulation, 215:456–475, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003452.

Copur:2024:NIP

[1470] Ayşegül Keten Çopur, Emirhan Hacıoğlu, and Faik Gürsoy. New insights on a pair of quasi-contractive operators in Banach spaces: Results on Jungck type iteration algorithms and proposed open problems. Mathematics and Computers in Simulation, 215:476–497, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300352X.

Zhang:2024:DAO

[1471] Ge Zhang, Zhiming Li, Anwarud Din, and Tao Chen. Dynamic analysis and optimal control of a stochastic COVID-19 model. Mathematics and Computers in Simulation, 215: 498–517, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003312.

Leonte:2024:SME

[1472] Dan Leonte and Almut E. D. Veraart. Simulation methods and error analysis for trawl processes and ambit fields. Mathematics and Computers in Simulation, 215:518– 542, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002987.

Ahsan:2024:HOM

[1473] Muhammad Ahsan, Weidong Lei, Martin Bohner, and Amir Ali Khan. A high-order multi-resolution wavelet method for nonlinear systems of differential equations. Mathematics and Computers in Simulation, 215:543– 559, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423003683.

Ngondiep:2024:HOC

[1474] Eric Ngondiep. A high-order combined finite element/interpolation approach for multidimensional nonlinear generalized Benjamin–Bona–Mahony– Burgers equation. Mathematics and Computers in Simulation, 215:560– 577, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423003786.

Panigrahi:2024:NAS

[1475] Paresh Kumar Panigrahi and Sukanta Nayak. Numerical approach to solve imprecisely defined systems using Inner Outer Direct Search optimization technique. Mathematics and Computers in Simulation, 215:578– 606, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423003518.

Wang:2024:DLU

[1476] Xiaorong Wang and Xiaodi Zhang. Decoupled, linear, unconditionally energy stable and charge-conservative finite element method for an inductionless magnetohydrodynamic phasefield model. Mathematics and Computers in Simulation, 215:607– 627, January 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003750.

Algaba:2024:CCF

[1477] Antonio Algaba, Cristóbal García, and Jaume Giné. Center conditions to find certain degenerate centers with characteristic directions. Mathematics and Computers in Simulation, 215: 628–638, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002860.

Anonymous:2024:NIa

[1478] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 215:639, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004020.

Anonymous:2024:ICEa

[1479] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 215:640, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423004032.

Anonymous:2024:EBa

[1480] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 215:ifc, January 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300397X.

Kiran:2024:EFM

[1481] Utpal Kiran, Deepak Sharma, and Sachin Singh Gautam. An efficient framework for matrix-free SpMV computation on GPU for elastoplastic problems. Mathematics and Computers in Simulation, 216(??):318– 346, February 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004160.

Bagheri:2024:TWE

[1482] Safieh Bagheri, Mohammad Hossein Akrami, Ghasem Barid Loghmani, and Mohammad Heydari. Traveling wave in an eco-epidemiological model with diffusion and convex incidence rate: Dynamics and numerical simulation. Mathematics and Computers in Simulation, 216(??): 347–366, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004275.

Berrone:2024:IHO

[1483] Stefano Berrone, Gioana Teora, and Fabio Vicini. Improving high-order VEM stability on badly-shaped elements. Mathematics and Computers in Simulation, 216(??):367– 385, February 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004287.

Akiyama:2024:WAB

[1484] Naho Akiyama and Toshihiro Yamada. A weak approximation for Bismut’s formula: an algorithmic differentiation method. Mathematics and Computers in Simulation, 216(??): 386–396, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003853.

Anonymous:2024:NIb

[1485] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 216(??):397, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004421.

Anonymous:2024:ICEb

[1486] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 216(??):398, February 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004433.

Anonymous:2024:EBb

[1487] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 216:ii, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004366.

Marashian:2024:LMI

[1488] Arash Marashian, Abolhassan Razminia, Vladimir I. Shiryaev, and Hamid R. Ossareh. Longitudinal model identification of multi-gear vehicles using an LPV approach. Mathematics and Computers in Simulation, 216:1–14, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003798.

Anonymous:2024:PF

[1489] Anonymous. Pages 1–398 (February 2024). Mathematics and Computers in Simulation, 216:1–398, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Belhamadia:2024:MFE

[1490] Youssef Belhamadia and Mohammed Seaid. A mixed finite element method for nonlinear radiation– conduction equations in optically thick anisotropic media. Mathematics and Computers in Simulation, 216:15– 29, February 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423003671.

Sriwastav:2024:ANS

[1491] Nikhil Sriwastav, Amit K. Barnwal, Higinio Ramos, Ravi P. Agarwal, and Mehakpreet Singh. Advanced numerical scheme and its convergence analysis for a class of two-point singular boundary value problems. Mathematics and Computers in Simulation, 216:30–48, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003749.

Pandey:2024:RDD

[1492] Soumik Pandey, Uttam Ghosh, Debashis Das, Sarbani Chakraborty, and Abhijit Sarkar. Rich dynamics of a delay-induced stage-structure prey– predator model with cooperative behaviour in both species and the impact of prey refuge. Mathematics and Computers in Simulation, 216: 49–76, February 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003841.

Abbaszadeh:2024:SLM

[1493] Mostafa Abbaszadeh, Mahmoud A.

Zaky, Ahmed S. Hendy, and Mehdi Dehghan. Supervised learning and meshless methods for two-dimensional fractional PDEs on irregular domains. Mathematics and Computers in Simulation, 216:77–103, February 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003348.

Abdulsalam:2024:DHO

[1494] Athraa Abdulsalam, Norazak Senu, Zanariah Abdul Majid, and Nik Mohd Asri Nik Long. Development of high-order adaptive multi-step Runge– Kutta-Nyström method for solving special second-order ODEs. Mathematics and Computers in Simulation, 216: 104–125, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003889.

Liu:2024:AHO

[1495] Yang Liu and Maohua Ran. Arbitrarily high-order explicit energyconserving methods for the generalized nonlinear fractional Schrödinger wave equations. Mathematics and Computers in Simulation, 216:126– 144, February 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542300383X.

Haslinger:2024:SPC

[1496] Jaroslav Haslinger, Radek Kučera, Kristina Motyčková, and Václav Šátek. Stokes problem with the Coulomb stick–slip boundary conditions in 3D: Formulations, approximation, algorithms, and experiments. Mathematics and Computers in Simulation, 216: 145–167, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003737.

Khirsariya:2024:RCA

[1497] Sagar R. Khirsariya, Jignesh P. Chauhan, and Snehal B. Rao. A robust computational analysis of residual power series involving general transform to solve fractional differential equations. Mathematics and Computers in Simulation, 216:168– 186, February 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423003890.

Pulch:2024:SGM

[1498] Roland Pulch. Stochastic Galerkin method and port-Hamiltonian form for linear dynamical systems of second order. Mathematics and Computers in Simulation, 216:187–197, February 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003865.

Yu:2024:SAH

[1499] Bin Yu, Jin Shi, Zeyu Xue, Mingzhe Yang, Xue Yang, and Yongmei Su. Stability analysis of HCV dynamic model with saturation incidence, cellular immunity and interferon effect in intrahepatic and extrahepatic tissues. Mathematics and Computers in Simulation, 216:301–317, February 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004172.

Anonymous:2024:EBc

[1500] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 217:ii, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004998.

Zhou:2024:CPI

[1501] Quan Zhou, Yinkun Wang, and Yicheng Liu. Chebyshev–Picard iteration methods for solving delay differential equations. Mathematics and Computers in Simulation, 217: 1–20, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423004238.

Anonymous:2024:PMa

[1502] Anonymous. Pages 1–446 (March 2024). Mathematics and Computers in Simulation, 217:1–446, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Bhaumik:2024:DLB

[1503] Bivas Bhaumik, Soumen De, and Satyasaran Changdar. Deep learning based solution of nonlinear partial differential equations arising in the process of arterial blood flow. Mathematics and Computers in Simulation, 217:21–36, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004469.

Kumbhakar:2024:IPS

[1504] Ruma Kumbhakar, Mainul Hossain, Sarbari Karmakar, and Nikhil Pal. An investigation of the parameter space in a tri-trophic food chain model with refuge. Mathematics and Computers in Simulation, 217: 37–59, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423004457.

Jang:2024:MTC

[1505] Bongsoo Jang and Hyunju Kim. Mapping techniques for collocation method of time-fractional convection– diffusion equations in domains with cracks. Mathematics and Computers in Simulation, 217:60–79, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423004500.

Li:2024:FTS

[1506] Ran Li, Chunmei Zhang, Hui Yang, and Huiling Chen. Finite-time synchronization and topology identification of stochastic multi-layer networks with Markovian switching. Mathematics and Computers in Simulation, 217:80–96, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004573.

Heydari:2024:NWM

[1507] M. H. Heydari and M. Razzaghi. A new wavelet method for fractional integro-differential equations with ψ-Caputo fractional derivative. Mathematics and Computers in Simulation, 217:97–108, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004585.

Arunagirinathan:2024:GDD

[1508] S. Arunagirinathan and T. H. Lee.

Generalized delay-dependent reciprocally convex inequality on stability for neural networks with timevarying delay. Mathematics and Computers in Simulation, 217:109– 120, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423004482.

Sunday:2024:CAS

[1509] Joshua Sunday, Ali Shokri, Nathaniel Mahwash Kamoh, Bwebum Cleofas Dang, and Nazim Idrisoglu Mahmudov. A computational approach to solving some applied rigid secondorder problems. Mathematics and Computers in Simulation, 217:121– 138, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423004536.

Vashishth:2024:MEN

[1510] Anil K. Vashishth and Komal Basaiti. Modeling the effect of nonpharmaceutical measures and vaccination on the spread of two variants of COVID-19 in India. Mathematics and Computers in Simulation, 217:139–168, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004408.

Suh:2024:MPV

[1511] Jiyeon Suh, Hee-Dae Kwon, and Jeehyun Lee. A model of plasmodium vivax malaria with delays: Mathematical analysis and numerical simulations. Mathematics and Computers in Simulation, 217:169– 187, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423004445.

Horng:2024:AGJ

[1512] Shih-Cheng Horng and Shieh-Shing Lin. Advanced golden jackal optimization for solving the constrained integer stochastic optimization problems. Mathematics and Computers in Simulation, 217:188– 201, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004561.

Maurya:2024:EPO

[1513] Vikas Maurya, Ankit Singh, Vivek S.

Yadav, and Manoj K. Rajpoot. Efficient pricing of options in jump– diffusion models: Novel implicit– explicit methods for numerical valuation. Mathematics and Computers in Simulation, 217:202–225, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423004603.

Nath:2024:OCA

[1514] Bhagya Jyoti Nath, Khadijeh Sadri, Hemanta Kumar Sarmah, and Kamyar Hosseini. An optimal combination of antiretroviral treatment and immunotherapy for controlling HIV infection. Mathematics and Computers in Simulation, 217:226– 243, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004470.

Mokni:2024:DBH

[1515] Karima Mokni and Mohamed Ch-Chaoui. A Darwinian Beverton–Holt model with immigration effect. Mathematics and Computers in Simulation, 217:244–261, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004494.

Ma:2024:LSB

[1516] Jie Ma, Zhiwei Yang, and Ning Du. Learning solution of a bondbased linear peridynamic model using LS-SVR method. Mathematics and Computers in Simulation, 217:262–272, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004524.

Karjoun:2024:WBP

[1517] Hasan Karjoun and Abdelaziz Beljadid. A well-balanced and positivitypreserving numerical model for overland flow under vegetation effects. Mathematics and Computers in Simulation, 217:273–293, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423004548.

Eid:2024:ABO

[1518] Heba F. Eid, Erik Cuevas, and Romany F. Mansour. Autonomous bonobo optimization algorithm for power allocation in wireless networks. Mathematics and Computers in Simulation, 217:294–310, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847542300455X.

Naffeti:2024:BBA

[1519] Bechir Naffeti, Hamadi Ammar, and Walid Ben Aribi. A Branch and Bound algorithm for multidimensional Holder optimization: Estimation of the age-dependent viral hepatitis A infection force. Mathematics and Computers in Simulation, 217:311– 326, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004597.

Chernik:2024:NSF

[1520] Vitaly Chernik and Pavel Buklemishev.

The numerical solution of the freeboundary cell motility problem. Mathematics and Computers in Simulation, 217:327–337, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004512.

Gao:2024:ASQ

[1521] Panqing Gao, Renyu Ye, Hai Zhang, Ivanka Stamova, and Jinde Cao. Asymptotic stability and quantitative synchronization of fractional competitive neural networks with multiple restrictions. Mathematics and Computers in Simulation, 217:338– 353, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004676.

Liu:2024:PES

[1522] Chein-Shan Liu, Yung-Wei Chen, and Chih-Wen Chang. Precise eigenvalues in the solutions of generalized Sturm–Liouville problems. Mathematics and Computers in Simulation, 217:354–373, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300469X.

Wang:2024:CMC

[1523] Xueli Wang and Suxia Zhang. Coupling media coverage and susceptibility for modeling epidemic dynamics: an application to COVID-19. Mathematics and Computers in Simulation, 217:374–394, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004615.

Zhou:2024:EET

[1524] Yongtao Zhou and Mingzhu Li. Error estimate of a transformed L1 scheme for a multi-term time-fractional diffusion equation by using discrete comparison principle. Mathematics and Computers in Simulation, 217:395– 404, March 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004718.

Marasi:2024:NMB

[1525] H. R. Marasi, M. H. Derakhshan, Amer A. Ghuraibawi, and Pushpendra Kumar. A novel method based on fractional order Gegenbauer wavelet operational matrix for the solutions of the multi-term time-fractional telegraph equation of distributed order. Mathematics and Computers in Simulation, 217:405–424, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004664.

Navarro:2024:QBF

[1526] Ciro Eduardo Bazán Navarro and Renato Mario Benazic Tomé. Qualitative behavior in a fractional order IS-LM-AS macroeconomic model with stability analysis. Mathematics and Computers in Simulation, 217:425–443, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004652.

Anonymous:2024:NIc

[1527] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 217:444, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005037.

Anonymous:2024:ICEc

[1528] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 217:445, March 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423005049.

Anonymous:2024:EBd

[1529] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 218:ii, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000089.

Selvaraj:2024:ETC

[1530] P. Selvaraj, O. M. Kwon, S. H. Lee, R. Sakthivel, and S. M. Lee. Eventtriggered control design with varying gains for polynomial fuzzy systems against DoS attacks. Mathematics and Computers in Simulation, 218:1–14, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004834.

Anonymous:2024:PA

[1531] Anonymous. Pages 1–724 (April 2024). Mathematics and Computers in Simulation, 218:1–724, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Luis:2024:LSC

[1532] Rafael Luís and Sandra Mendonça. Local stability conditions for a ndimensional periodic mapping. Mathematics and Computers in Simulation, 218:15–30, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004846.

Tian:2024:CDF

[1533] Yuan Tian, Huanmeng Li, and Kaibiao Sun. Complex dynamics of a fishery model: Impact of the triple effects of fear, cooperative hunting and intermittent harvesting. Mathematics and Computers in Simulation, 218:31–48, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004858.

Clain:2024:SKO

[1534] S. Clain and J. Figueiredo. Stencil and kernel optimisation for meshfree very high-order generalised finite difference method. Mathematics and Computers in Simulation, 218: 49–78, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004706.

Shobeyri:2024:IMM

[1535] Gholamreza Shobeyri. Improved MPS models for simulating free surface flows. Mathematics and Computers in Simulation, 218:79–97, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004767.

Chen:2024:NDR

[1536] Xiyuan Chen and Qiubao Wang. A novel dimensionality reduction approach by integrating dynamics theory and machine learning. Mathematics and Computers in Simulation, 218:98–111, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004883.

Hammou:2024:MFH

[1537] Amina Hammou, Youssef Hilali, Said Mesmoudi, Radouane Boujmal, and Oussama Bourihane. A mesh-free Hermite-type approach for buckling analysis of functionally graded polygonal thin plates. Mathematics and Computers in Simulation, 218:112– 132, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004925.

Farman:2024:ACD

[1538] Muhammad Farman, Aqeel Ahmad, Anum Zehra, Kottakkaran Sooppy Nisar, Evren Hincal, and Ali Akgul. Analysis and controllability of diabetes model for experimental data by using fractional operator. Mathematics and Computers in Simulation, 218:133–148, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004779.

Li:2024:QFF

[1539] Wanai Li. Quadrature-free forms of discontinuous Galerkin methods in solving compressible flows on triangular and tetrahedral grids. Mathematics and Computers in Simulation, 218:149–173, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004895.

Zelinka:2024:MSF

[1540] Ivan Zelinka, Miloslav Szczypka, Jan Plucar, and Nikolay Kuznetsov. From malware samples to fractal images: a new paradigm for classification. Mathematics and Computers in Simulation, 218:174–203, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004937.

Li:2024:NSE

[1541] Qiang Li, Jinling Liang, Weiqiang Gong, Kai Wang, and Jinling Wang. Nonfragile state estimation for semi-Markovian switching CVNs with general uncertain transition rates: an event-triggered scheme. Mathematics and Computers in Simulation, 218:204–222, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004913.

Arai:2024:MCS

[1542] Takuji Arai and Yuto Imai. Monte Carlo simulation for Barndorff–Nielsen and Shephard model under change of measure. Mathematics and Computers in Simulation, 218:223– 234, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004901.

Zhang:2024:CUE

[1543] Gengen Zhang, Jingyu Li, and Qiong-Ao Huang. A class of unconditionally energy stable relaxation schemes for gradient flows. Mathematics and Computers in Simulation, 218:235– 247, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004743.

Ahmed:2024:GBO

[1544] Marzia Ahmed, Mohd Herwan Sulaiman, Ahmad Johari Mohamad, and Mostafijur Rahman. Gooseneck barnacle optimization algorithm: a novel nature inspired optimization theory and application. Mathematics and Computers in Simulation, 218:248–265, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004329.

Cao-Rial:2024:MPX

[1545] M. T. Cao-Rial, C. Moreno, and P. Quintela. Mathematical perspective on XFEM implementation for models involving contribution on interfaces. Mathematics and Computers in Simulation, 218:266– 291, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423004731.

Rezaei:2024:CTM

[1546] Vahid Rezaei and Esmaeel Khanmirza. Continuous-time min-max consensus protocol: a unified approach. Mathematics and Computers in Simulation, 218:292–310, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004810.

Admon:2024:NMS

[1547] Mohd Rashid Admon, Norazak Senu, Ali Ahmadian, Zanariah Abdul Majid, and Soheil Salahshour. A new modern scheme for solving fractal– fractional differential equations based on deep feedforward neural network with multiple hidden layer. Mathematics and Computers in Simulation, 218:311–333, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004639.

Stojanovic:2024:VAC

[1548] Vladimir Stojanović, Jian Deng, Dunja Milić, and Marko D. Petković. Vibrational analysis of a coupled damaged Timoshenko beam-arch mechanical system with von Kármán nonlinearities and layer discontinuity. Mathematics and Computers in Simulation, 218:334–356, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300472X.

Scianna:2024:ABI

[1549] Marco Scianna. The AddACO: a bioinspired modified version of the ant colony optimization algorithm to solve travel salesman problems. Mathematics and Computers in Simulation, 218:357–382, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005116.

Ansari:2024:TSF

[1550] Alireza Ansari and Mohammad Hossein Derakhshan. Time–space fractional Euler–Poisson–Darboux equation with Bessel fractional derivative in infinite and finite domains. Mathematics and Computers in Simulation, 218:383–402, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423005086.

K:2024:PAA

[1551] Divya K. and Indhira K. Performance analysis and ANFIS computing of an unreliable Markovian feedback queueing model under a hybrid vacation policy. Mathematics and Computers in Simulation, 218: 403–419, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423005104.

Misra:2024:MEB

[1552] A. K. Misra, Akash Yadav, and Mohammad Sajid. Modeling the effects of baculovirus to control insect population in agricultural fields. Mathematics and Computers in Simulation, 218:420–434, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423005098.

Aghaseyedabdollah:2024:ITF

[1553] Mohammadhossein Aghaseyedabdollah, Mostafa Abedi, and Mahdi Pourgholi. Interval type-2 fuzzy sliding mode control for a cable-driven parallel robot with elastic cables using metaheuristic optimization methods. Mathematics and Computers in Simulation, 218:435–461, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005062.

Dong:2024:CAS

[1554] W. B. Dong and H. S. Tang. Convergence analysis of Schwarz waveform relaxation method to compute coupled advection–diffusion– reaction equations. Mathematics and Computers in Simulation, 218:462– 481, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004871.

Caraballo:2024:ERR

[1555] Tomás Caraballo, Javier López de-la Cruz, and Verónica Caraballo-Romero. Effects of real random perturbations on Monod and Haldane consumption functions in the chemostat model. Mathematics and Computers in Simulation, 218:482–497, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005050.

Kumar:2024:DSI

[1556] Ajay Kumar, Balwant Singh Thakur, and Mihai Postolache. Dynamic stepsize iteration process for solving split common fixed point problems with applications. Mathematics and Computers in Simulation, 218:498–511, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005128.

Yousif:2024:FNP

[1557] Majeed A. Yousif and Faraidun K. Hamasalh. The fractional nonpolynomial spline method: Precision and modeling improvements. Mathematics and Computers in Simulation, 218:512–525, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004949.

Basavaraj:2024:AGO

[1558] M. S. Basavaraj and D. L. Shivaraj Kumar. Algebraic growth of 2D optimal perturbation of a plane Poiseuille flow in a Brinkman porous medium. Mathematics and Computers in Simulation, 218:526–543, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542300486X.

Elgindy:2024:FGPa

[1559] Kareem T. Elgindy. Fourier– Gegenbauer pseudospectral method for solving time-dependent onedimensional fractional partial differential equations with variable coefficients and periodic solutions. Mathematics and Computers in Simulation, 218:544–555, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004950.

Zeghlache:2024:RAB

[1560] Samir Zeghlache, Hilal Rahali, Ali Djerioui, Loutfi Benyettou, and Mohamed Fouad Benkhoris. Robust adaptive backstepping neural networks fault tolerant control for mobile manipulator UAV with multiple uncertainties. Mathematics and Computers in Simulation, 218:556– 585, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005074.

Taher:2024:HAC

[1561] Anis Haytham Saleh Taher. Highly accurate calculation of higher energy eigenvalues for the radial Schrödinger eigenproblems. Mathematics and Computers in Simulation, 218:586– 599, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005153.

Abu-Ghuwaleh:2024:MGN

[1562] Mohammad Abu-Ghuwaleh, Rania Saadeh, and Rasheed Saffaf. Master generators: a novel approach to construct and solve ordinary differential equations. Mathematics and Computers in Simulation, 218:600– 623, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004780.

Opoku-Sarkodie:2024:BAW

[1563] R. Opoku-Sarkodie, F. A. Bartha, M. Polner, and G. Röst. Bifurcation analysis of waning-boosting epidemiological models with repeat infections and varying immunity periods. Mathematics and Computers in Simulation, 218:624–643, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542300513X.

Wang:2024:ECF

[1564] Kun Wang and Lin Mu. An enriched cut finite element method for Stokes interface equations. Mathematics and Computers in Simulation, 218:644–665, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423005232.

Hussain:2024:MSA

[1565] Sardar Muhammad Hussain, José Alberto Cuminato, Ivan Silvestre Paganini Marin, and Edson Cezar Wendland. A matrix solver approach for fracture flow simulation by Analytic Element Method. Mathematics and Computers in Simulation, 218:666–678, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423005219.

Gao:2024:DLI

[1566] Ruimin Gao, Dongfang Li, Ming Mei, and Dan Zhao. A decoupled, linearly implicit and high-order structure-preserving scheme for Euler– Poincaré equations. Mathematics and Computers in Simulation, 218: 679–703, April 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423005165.

Bandeira:2024:SCT

[1567] L. Bandeira and P. Pedregal. Some computational tests for inverse conductivity problems based on vector, variational principles: the 2D case. Mathematics and Computers in Simulation, 218:704–721, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005244.

Anonymous:2024:NId

[1568] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 218:722, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000120.

Anonymous:2024:ICEd

[1569] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 218:723, April 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424000132.

Anonymous:2024:EBe

[1570] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 219:ii, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000466.

Zeng:2024:BAP

[1571] Qiaoyun Zeng, Yanhong Zheng, and Dan Yi. Bifurcation analysis of a Parkinson’s disease model with two time delays. Mathematics and Computers in Simulation, 219: 1–11, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005141.

Anonymous:2024:PMb

[1572] Anonymous. Pages 1–624 (May 2024). Mathematics and Computers in Simulation, 219:1–624, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Qiu:2024:DRS

[1573] Wenlin Qiu, Yanping Chen, Xu Xiao, and Xiangcheng Zheng. Discrete L^{1} remainder stability of first and second order schemes for a Volterra integro-differential equation. Mathematics and Computers in Simulation, 219:12–27, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005190.

Lee:2024:PEF

[1574] K. C. Lee, R. Nazar, N. Senu, and A. Ahmadian. A promising exponentially-fitted two-derivative Runge–Kutta–Nyström method for solving y′′ = f(x, y): Application to Verhulst logistic growth model. Mathematics and Computers in Simulation, 219:28–49, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005256.

Kumar:2024:ISM

[1575] Arjun Kumar, Uma S. Dubey, and Balram Dubey. The impact of social media advertisements and treatments on the dynamics of infectious diseases with optimal control strategies. Mathematics and Computers in Simulation, 219:50–86, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005220.

Liu:2024:PGI

[1576] Chengzhi Liu, Yue Qiu, and Li Zhang.

Preconditioned geometric iterative methods for B-spline interpolation. Mathematics and Computers in Simulation, 219:87–100, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005177.

Li:2024:SON

[1577] Jing Li, Xinyue Kang, Xingyun Shi, and Yufei Song. A secondorder numerical method for nonlinear variable-order fractional diffusion equation with time delay. Mathematics and Computers in Simulation, 219:101–111, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005268.

Zhang:2024:MNI

[1578] Hanwen Zhang and Duy-Minh Dang.

A monotone numerical integration method for mean–variance portfolio optimization under jump-diffusion models. Mathematics and Computers in Simulation, 219:112–140, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423005189.

Liu:2024:ISM

[1579] Shuang Liu, Tianwei Xu, Qingyun Wang, and Yan Yang. The impulsive synchronization of multiplex networks with mixed delays and dual uncertainties. Mathematics and Computers in Simulation, 219:141– 163, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005207.

Yu:2024:ARM

[1580] Xiaobing Yu, Haoyu Wang, and Yangchen Lu. An adaptive ranking moth flame optimizer for feature selection. Mathematics and Computers in Simulation, 219:164– 184, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542300527X.

Chowdhury:2024:MIP

[1581] Sudipto Chowdhury, Divay Garg, and Ravina Shokeen. Modified C^{0} interior penalty analysis for fourth order Dirichlet boundary control problem and a posteriori error estimates. Mathematics and Computers in Simulation, 219:185–211, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005323.

Hu:2024:EPE

[1582] Zhihao Hu, Ben-Zhang Yang, Xin-Jiang He, and Jia Yue. Equilibrium pricing of European crude oil options with stochastic behaviour and jump risks. Mathematics and Computers in Simulation, 219:212– 230, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005293.

Zhu:2024:NDC

[1583] Hegui Zhu, Jiangxia Ge, Jinwen He, and Libo Zhang. A non-degenerate chaotic bits XOR system with application in image encryption. Mathematics and Computers in Simulation, 219:231–250, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300530X.

Anwar:2024:NIP

[1584] Nabeela Anwar, Iftikhar Ahmad, Adiqa Kausar Kiani, Muhammad Shoaib, and Muhammad Asif Zahoor Raja. Novel intelligent predictive networks for analysis of chaos in stochastic differential SIS epidemic model with vaccination impact. Mathematics and Computers in Simulation, 219:251–283, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005311.

Cisse:2024:SSD

[1585] Amadou Cissé and Mohamed Boutayeb. Software sensors design for a class of non linear coupled PDE systems: the Vlasov–Poisson dynamical system. Mathematics and Computers in Simulation, 219:284– 296, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005359.

Elgindy:2024:DIP

[1586] Kareem T. Elgindy and Hareth M. Refat. Direct integral pseudospectral and integral spectral methods for solving a class of infinite horizon optimal output feedback control problems using rational and exponential Gegenbauer polynomials. Mathematics and Computers in Simulation, 219:297–320, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005335.

Gupta:2024:DSM

[1587] Soniya Gupta, Dheeraj Kumar Joshi, Natasha Awasthi, Manish Pant, Bhagawati prasad Joshi, and Shshank Chaube. Distance and similarity measures of Hesitant bi-fuzzy set and its applications in renewable energy systems. Mathematics and Computers in Simulation, 219: 321–336, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005281.

Kedia:2024:RFD

[1588] Nikki Kedia, Anatoly A. Alikhanov, and Vineet Kumar Singh. Robust finite difference scheme for the non-linear generalized timefractional diffusion equation with nonsmooth solution. Mathematics and Computers in Simulation, 219:337– 354, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005414.

Haddouchi:2024:SGR

[1589] Faouzi Haddouchi and Mohammad Esmael Samei. Solvability of a generalized ψ-Riemann–Liouville fractional BVP under nonlocal boundary conditions. Mathematics and Computers in Simulation, 219:355– 377, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005360.

He:2024:MSK

[1590] Yichao He, Jinghong Wang, Xuejing Liu, Xizhao Wang, and Haibin Ouyang. Modeling and solving of knapsack problem with setup based on evolutionary algorithm. Mathematics and Computers in Simulation, 219:378–403, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005402.

Jangid:2024:LWC

[1591] Komal Jangid, Bhagwan Singh, and Santwana Mukhopadhyay. Legendre wavelet collocation method for investigating thermo-mechanical responses on biological tissue during laser irradiation. Mathematics and Computers in Simulation, 219:404– 423, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005372.

Wu:2024:SCP

[1592] Dawen Wu and Abdel Lisser. Solving Constrained Pseudoconvex Optimization Problems with deep learning-based neurodynamic optimization. Mathematics and Computers in Simulation, 219:424–434, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005396.

Guo:2024:SML

[1593] Beibei Guo and Yu Xiao. Synchronization of multi-link and multidelayed inertial neural networks with Markov jump via aperiodically intermittent adaptive control. Mathematics and Computers in Simulation, 219:435–453, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005463.

Escobar-Cuevas:2024:NOA

[1594] Héctor Escobar-Cuevas, Erik Cuevas, Jorge Gálvez, and Miguel Toski. A novel optimization approach based on unstructured evolutionary game theory. Mathematics and Computers in Simulation, 219:454–472, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423005347.

Mann:2024:NCF

[1595] Nikita Mann, Setu Rani, Sachin Kumar, and Raj Kumar. Novel closedform analytical solutions and modulation instability spectrum induced by the Salerno equation describing nonlinear discrete electrical lattice via symbolic computation. Mathematics and Computers in Simulation, 219:473–490, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005384.

Luo:2024:ASF

[1596] Lingao Luo, Lulu Li, and Wei Huang.

Asymptotic stability of fractionalorder Hopfield neural networks with event-triggered delayed impulses and switching effects. Mathematics and Computers in Simulation, 219:491– 504, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423005426.

Medda:2024:IPC

[1597] Rakesh Medda, Pankaj Kumar Tiwari, and Samares Pal. Impacts of planktonic components on the dynamics of cholera epidemic: Implications from a mathematical model. Mathematics and Computers in Simulation, 219:505–526, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423005451.

Cesek:2024:RCT

[1598] Ema Češek, Jan Grošelj, Andrej Kolar-Požun, Maruša Lekše, Gašper Domen Romih, Ada Šadl Praprotnik, and Matija Šteblaj. A representation and comparison of three cubic macro-elements. Mathematics and Computers in Simulation, 219: 527–543, May 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423005505.

Zhang:2024:ISS

[1599] Hongbo Zhang, Xiwen Qin, Xueliang Gao, Siqi Zhang, Yunsheng Tian, and Wei Zhang. Improved salp swarm algorithm based on Newton interpolation and cosine opposition-based learning for feature selection. Mathematics and Computers in Simulation, 219:544–558, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542300544X.

Tang:2024:PPE

[1600] Bo Tang, Wenting Mao, and Zhankuan Zeng. A priori and a posteriori error estimates of a space–time Petrov– Galerkin spectral method for timefractional diffusion equation. Mathematics and Computers in Simulation, 219:559–572, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542400003X.

Xing:2024:FTS

[1601] Xiaofei Xing, Huaiqin Wu, and Jinde Cao. Finite-time synchronization of impulsive stochastic systems with DoS attacks via dynamic event-triggered control. Mathematics and Computers in Simulation, 219:573–593, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005487.

Deng:2024:NNP

[1602] Dingwen Deng and Mengting Hu. Nonnegativity-preserving and maximumprinciple-satisfying finite difference methods for Fisher’s equation with delay. Mathematics and Computers in Simulation, 219:594–622, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424000259.

Anonymous:2024:NIe

[1603] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 219:623, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000508.

Anonymous:2024:ICEe

[1604] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 219:624, May 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847542400051X.

Anonymous:2024:EBf

[1605] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 220:ii, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400082X.

Akbari:2024:SMN

[1606] Tahereh Akbari, Mohsen Esmaeilbeigi, and Davoud Moazami. A stable meshless numerical scheme using hybrid kernels to solve linear Fredholm integral equations of the second kind and its applications. Mathematics and Computers in Simulation, 220:1–28, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000168.

Anonymous:2024:PJb

[1607] Anonymous. Pages 1–698 (June 2024). Mathematics and Computers in Simulation, 220:1–698, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Musavi:2024:LIF

[1608] Seyed Hossein Musavi, Mahmud Ashrafizaadeh, and Seyyed Meysam Khatoonabadi. Off-lattice interfacial force scheme for simulation of multiphase flows using meshless lattice Boltzmann method. Mathematics and Computers in Simulation, 220: 29–43, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000181.

Roy:2024:GAB

[1609] Tapas Roy and Dilip K. Maiti. General approach on the best fitted linear operator and basis function for homotopy methods and application to strongly nonlinear oscillators. Mathematics and Computers in Simulation, 220:44–64, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000041.

Pan:2024:PCM

[1610] Jeng-Shyang Pan, Zhen Zhang, Shu-Chuan Chu, Si-Qi Zhang, and Jimmy Ming-Tai Wu. A parallel compact Marine Predators Algorithm applied in time series prediction of Backpropagation neural network (BNN) and engineering optimization. Mathematics and Computers in Simulation, 220:65–88, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000247.

Selvam:2024:ICS

[1611] Anjapuli Panneer Selvam and Venkatesan Govindaraj. Investigation of controllability and stability of fractional dynamical systems with delay in control. Mathematics and Computers in Simulation, 220:89– 104, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000193.

Kaur:2024:KSE

[1612] Navneet Kaur and Varun Joshi. Kuramoto–Sivashinsky equation: Numerical solution using two quintic B-splines and differential quadrature method. Mathematics and Computers in Simulation, 220:105– 127, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423005438.

Pal:2024:CDM

[1613] Debjit Pal, Dipak Kesh, and Debasis Mukherjee. Cross-diffusion mediated Spatiotemporal patterns in a predator–prey system with hunting cooperation and fear effect. Mathematics and Computers in Simulation, 220:128–147, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000028.

Rawat:2024:GMJ

[1614] Shivam Rawat, Darshana J. Prajapati, Anita Tomar, and Krzysztof Gdawiec. Generation of Mandelbrot and Julia sets for generalized rational maps using SP-iteration process equipped with s-convexity. Mathematics and Computers in Simulation, 220:148–169, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423005475.

Wang:2024:TPP

[1615] Fatao Wang, Ruizhi Yang, and Xin Zhang. Turing patterns in a predator–prey model with double Allee effect. Mathematics and Computers in Simulation, 220:170– 191, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000272.

Ali:2024:TEM

[1616] Khalid K. Ali, Derya Yıldırım Sucu, and Seydi Battal Gazi Karakoc. Two effective methods for solution of the Gardner–Kawahara equation arising in wave propagation. Mathematics and Computers in Simulation, 220: 192–203, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000144.

Zhang:2024:MRF

[1617] Ziqing Zhang. Multi-regime foreign exchange rate model: Calibration and pricing. Mathematics and Computers in Simulation, 220: 204–218, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542400017X.

Zeng:2024:IMU

[1618] Min Zeng and Yongxin Yuan. An iterative method for updating finite element models with connectivity constraints. Mathematics and Computers in Simulation, 220:219– 236, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000284.

Qureshi:2024:NIS

[1619] Sania Qureshi, Higinio Ramos, Amanullah Soomro, Olusheye Aremu Akinfenwa, and Moses Adebowale Akanbi. Numerical integration of stiff problems using a new time-efficient hybrid block solver based on collocation and interpolation techniques. Mathematics and Computers in Simulation, 220:237–252, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000016.

Cao:2024:MHB

[1620] Baiheng Cao, Xuedong Wu, Yaonan Wang, and Zhiyu Zhu. Modified hybrid B-spline estimation based on spatial regulator tensor network for Burger equation with nonlinear fractional calculus. Mathematics and Computers in Simulation, 220:253–275, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000156.

Zhou:2024:DOC

[1621] Qi Zhou, Xining Li, Jing Hu, and Qimin Zhang. Dynamics and optimal control for a spatial heterogeneity model describing respiratory infectious diseases affected by air pollution. Mathematics and Computers in Simulation, 220:276– 295, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000363.

Naudet:2024:NBA

[1622] Charles J. Naudet and Alan E. Lindsay. Numerical bifurcation analysis of post-contact states in mathematical models of Micro-Electromechanical Systems. Mathematics and Computers in Simulation, 220:296–308, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424000296.

Cerna:2024:OPU

[1623] Dana Černá and Kateřina Fiňková.

Option pricing under multifactor Black–Scholes model using orthogonal spline wavelets. Mathematics and Computers in Simulation, 220: 309–340, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000338.

Chen:2024:BAS

[1624] Haiying Chen, Adele Shahi, Gurpreet Singh, Jalil Manafian, Baharak Eslami, and Naief Alabed Alkader. Behavior of analytical schemes with non-paraxial pulse propagation to the cubic–quintic nonlinear Helmholtz equation. Mathematics and Computers in Simulation, 220:341–356, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000429.

Adhikari:2024:EJM

[1625] Nabaraj Adhikari and Wutiphol Sintunavarat. Exploring the Julia and Mandelbrot sets of z^{p} + log c^{t} using a four-step iteration scheme extended with s-convexity. Mathematics and Computers in Simulation, 220:357–381, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542400020X.

Belomestny:2024:TGN

[1626] Denis Belomestny, Artur Goldman, Alexey Naumov, and Sergey Samsonov. Theoretical guarantees for neural control variates in MCMC. Mathematics and Computers in Simulation, 220:382–405, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000314.

Boffi:2024:PSF

[1627] Daniele Boffi, Fabio Credali, Lucia Gastaldi, and Simone Scacchi. A parallel solver for fluid– structure interaction problems with Lagrange multiplier. Mathematics and Computers in Simulation, 220: 406–424, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424000399.

Zhu:2024:PDA

[1628] Linhe Zhu, Siyi Chen, and Shuling Shen. Pattern dynamics analysis of a reaction–diffusion network propagation model. Mathematics and Computers in Simulation, 220: 425–444, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000375.

Tao:2024:UCW

[1629] Xia Tao, Jiaxiong Hao, and Yu Zhang.

The uniform convergence of a weak Galerkin finite element method in the balanced norm for reaction– diffusion equation. Mathematics and Computers in Simulation, 220:445– 461, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000405.

Sivalingam:2024:NLP

[1630] S. M. Sivalingam, Pushpendra Kumar, Hieu Trinh, and V. Govindaraj. A novel L1-Predictor-Corrector method for the numerical solution of the generalized-Caputo type fractional differential equations. Mathematics and Computers in Simulation, 220:462–480, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000302.

Li:2024:PAD

[1631] Yibao Li, Qing Xia, Seungyoon Kang, Soobin Kwak, and Junseok Kim. A practical algorithm for the design of multiple-sized porous scaffolds with triply periodic structures. Mathematics and Computers in Simulation, 220:481–495, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000521.

Ali:2024:ECP

[1632] Hegagi Mohamed Ali, Ismail Gad Ameen, and Yasmeen Ahmed Gaber. The effect of curative and preventive optimal control measures on a fractional order plant disease model. Mathematics and Computers in Simulation, 220:496–515, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000582.

Urus:2024:ERC

[1633] Nazia Urus and Amit Kumar Verma. Existence results for a class of four point nonlinear singular BVP arising in thermal explosion in a spherical vessel. Mathematics and Computers in Simulation, 220:516– 532, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424000533.

Tan:2024:MDS

[1634] Jia Tan and Tian jun Wang. A multidomain spectral collocation method for the Fokker–Planck equation in an infinite channel. Mathematics and Computers in Simulation, 220: 533–551, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424000612.

Zhang:2024:AFE

[1635] Xiulan Zhang, Jiangteng Shi, Heng Liu, and Fangqi Chen. Adaptive fuzzy event-triggered cooperative control for fractional-order delayed multi-agent systems with unknown control direction. Mathematics and Computers in Simulation, 220:552–566, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000545.

Yang:2024:DAO

[1636] Yongkuan Yang, Bing Yan, Xiangsong Kong, and Jing Zhao. A double auxiliary optimization constrained multi-objective evolutionary algorithm. Mathematics and Computers in Simulation, 220:567– 579, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000624.

Shi:2024:GFI

[1637] C. Z. Shi, H. Zheng, Y. C. Hon, and P. H. Wen. Generalized finite integration method for 2D elastostatic and elastodynamic analysis. Mathematics and Computers in Simulation, 220:580–594, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000636.

Zhang:2024:MAA

[1638] Lidong Zhang, Jinliang Wang, and Ran Zhang. Mathematical analysis for an age-space structured HIV model with latency. Mathematics and Computers in Simulation, 220: 595–617, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000661.

Panda:2024:NTS

[1639] Abhilipsa Panda, Jugal Mohapatra, Ilhame Amirali, Muhammet Enes Durmaz, and Gabil M. Amiraliyev. A numerical technique for solving nonlinear singularly perturbed Fredholm integro-differential equations. Mathematics and Computers in Simulation, 220:618–629, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000600.

He:2024:PIN

[1640] Tingxiao He, Yun Wang, and Yingnan Zhang. A partial-integrable numerical simulation scheme of the derivative nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 220:630–639, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000697.

Lahrache:2024:NSA

[1641] Manar Lahrache, Francisco Ortegón Gallego, and Mohamed Rhoudaf. 3D numerical simulation of an anisotropic bead type thermistor and multiplicity of solutions. Mathematics and Computers in Simulation, 220:640– 672, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000673.

Meenakshi:2024:ISS

[1642] RM. Meenakshi and K. Selvi. Iteratively Sustained Sliding Mode Control based energy management in a DC Microgrid. Mathematics and Computers in Simulation, 220:673– 695, June 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004299.

Anonymous:2024:NIf

[1643] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 220:696, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000855.

Anonymous:2024:ICEf

[1644] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 220:697, June 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424000879.

Anonymous:2024:EBg

[1645] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 221:ii, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001198.

Banasiak:2024:MMM

[1646] J. Banasiak and S. Y. Tchoumi. Multiscale malaria models and their uniform in-time asymptotic analysis. Mathematics and Computers in Simulation, 221:1–18, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000648.

Anonymous:2024:PJc

[1647] Anonymous. Pages 1–664 (July 2024). Mathematics and Computers in Simulation, 221:1–664, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Manzini:2024:PEA

[1648] Gianmarco Manzini and Annamaria Mazzia. A posteriori error analysis and mesh adaptivity for a virtual element method solving the Stokes equations. Mathematics and Computers in Simulation, 221:19–38, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000326.

Jiang:2024:ESA

[1649] Nan Jiang and Ying Li. An efficient scalar auxiliary variable partitioned projection ensemble method for simulating surface-groundwater flows. Mathematics and Computers in Simulation, 221:39–54, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000417.

Casaban:2024:RFB

[1650] M.-C. Casabán, R. Company, V. N.

Egorova, and L. Jódar. A random free-boundary diffusive logistic differential model: Numerical analysis, computing and simulation. Mathematics and Computers in Simulation, 221:55–78, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400065X.

Tian:2024:ETA

[1651] Yongjie Tian, Huiyan Zhang, Yongchao Liu, Ning Zhao, and Kalidass Mathiyalagan. Event-triggered adaptive secure tracking control for nonlinear cyber–physical systems against unknown deception attacks. Mathematics and Computers in Simulation, 221:79–93, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000715.

Xing:2024:ITL

[1652] Aosheng Xing, Yong Chen, Jinyi Suo, and Jie Zhang. Improving teachinglearning-based optimization algorithm with golden-sine and multi-population for global optimization. Mathematics and Computers in Simulation, 221: 94–134, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000570.

Pulido:2024:NDC

[1653] M. Aurora P. Pulido, J. Vanterler C.

Sousa, and E. Capelas de Oliveira. New discretization of ψ-Caputo fractional derivative and applications. Mathematics and Computers in Simulation, 221:135–158, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000491.

Afkar:2024:VBT

[1654] Mohammad Afkar, Roghayeh Gavagsaz-Ghoachani, Matheepot Phattanasak, and Serge Pierfederici. Voltagebalancing of two controllers for a DC-DC converter-based DC microgrid with experimental verification. Mathematics and Computers in Simulation, 221:159–179, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000685.

Haslinger:2024:SOS

[1655] Jaroslav Haslinger and Raino A. E. Mäkinen. Shape optimization for the Stokes system with threshold leak boundary conditions. Mathematics and Computers in Simulation, 221:180–196, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000739.

Nguyen:2024:ISSb

[1656] Khanh Hieu Nguyen and Sung Hyun Kim. Improved stability and stabilization criteria for multi-rate sampleddata control systems via novel delaydependent states. Mathematics and Computers in Simulation, 221:197– 209, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000752.

Carosi:2024:LAM

[1657] Daniele Carosi, Fabiana Zama, Alessandro Morri, and Lorella Ceschini. Linearising anhysteretic magnetisation curves: a novel algorithm for finding simulation parameters and magnetic moments. Mathematics and Computers in Simulation, 221:210– 221, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000776.

Park:2024:EPC

[1658] Sangbeom Park, Yonghyeon Jeon, Philsu Kim, and Soyoon Bak. An error predict-correction formula of the load vector in the BSLM for solving threedimensional Burgers’ equations. Mathematics and Computers in Simulation, 221:222–243, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000727.

Zhang:2024:ETC

[1659] Ziyu Zhang, Xinsong Yang, Hak-Keung Lam, and Zhengrong Xiang. Eventtriggered control for switched systems with sensor faults via adaptive fuzzy observer. Mathematics and Computers in Simulation, 221: 244–259, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424000764.

Fu:2024:IWB

[1660] Qingcheng Fu, Zhen Gao, Yaguang Gu, Peng Li, and Bao-Shan Wang. Improved well-balanced AWENO schemes with hydrostatic reconstruction for the Euler equations under gravitational fields. Mathematics and Computers in Simulation, 221:260– 280, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424000788.

Zhang:2024:HRB

[1661] Yufeng Zhang, Jing Li, Shaotao Zhu, and Zerui Ma. Harmonic resonance and bifurcation of fractional Rayleigh oscillator with distributed time delay. Mathematics and Computers in Simulation, 221: 281–297, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424000880.

Chen:2024:OCB

[1662] Xin Chen, Yue Yuan, Dongmei Yuan, and Xiao Ge. Optimal control for both forward and backward discretetime systems. Mathematics and Computers in Simulation, 221:298– 314, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424000892.

Cavalli:2024:DAH

[1663] Fausto Cavalli, Ahmad Naimzada, and Daniela Visetti. Dynamical analysis of healthcare policy effects in an integrated economicepidemiological model. Mathematics and Computers in Simulation, 221:315–336, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000910.

Luo:2024:RAP

[1664] Yantao Luo, Jianhua Huang, Zhidong Teng, and Qun Liu. Role of ART and PrEP treatments in a stochastic HIV/AIDS epidemic model. Mathematics and Computers in Simulation, 221:337–357, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000909.

Manohara:2024:NAF

[1665] G. Manohara and S. Kumbinarasaiah. Numerical approximation of fractional SEIR epidemic model of measles and smoking model by using Fibonacci wavelets operational matrix approach. Mathematics and Computers in Simulation, 221:358– 396, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000703.

Dang:2024:NMF

[1666] Quang A. Dang, Quang Long Dang, and T. Kim Quy Ngo. Numerical methods of fourth, sixth and eighth orders convergence for solving third order nonlinear ODEs. Mathematics and Computers in Simulation, 221:397–414, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000995.

Mondal:2024:CDT

[1667] Bapin Mondal, Ashraf Adnan Thirthar, Nazmul Sk, Manar A. Alqudah, and Thabet Abdeljawad. Complex dynamics in a two species system with Crowley–Martin response function: Role of cooperation, additional food and seasonal perturbations. Mathematics and Computers in Simulation, 221:415–434, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000958.

Etangsale:2024:NIC

[1668] Grégory Etangsale, Marwan Fahs, Vincent Fontaine, and Hussein Hoteit. A numerical investigation on coupling of conforming and hybridizable interior penalty discontinuous Galerkin methods for fractured groundwater flow problems. Mathematics and Computers in Simulation, 221: 435–460, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000740.

Farman:2024:MDM

[1669] Muhammad Farman, Changjin Xu, Aamir Shehzad, and Ali Akgul. Modeling and dynamics of measles via fractional differential operator of singular and non-singular kernels. Mathematics and Computers in Simulation, 221:461–488, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000983.

Ma:2024:HOH

[1670] Yu-Lan Ma and Bang-Qing Li. Higherorder hybrid rogue wave and breather interaction dynamics for the AB system in two-layer fluids. Mathematics and Computers in Simulation, 221:489–502, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400096X.

Zhang:2024:MEC

[1671] Jingyi Zhang and Xuejuan Liu.

The mechanism of enterprise credit guarantee risk contagion considering ESG. Mathematics and Computers in Simulation, 221:503– 514, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000946.

Rammane:2024:HTC

[1672] Mohammed Rammane, Oussama Elmhaia, Said Mesmoudi, Omar Askour, Abdeljalil Tri, Bouazza Braikat, and Noureddine Damil. A Hermitetype collocation mesh-free approach for simulating incompressible viscous fluid flows. Mathematics and Computers in Simulation, 221:515– 532, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001022.

Zhang:2024:FIN

[1673] Qiang Zhang and Jijun Liu. On fluorophore imaging by nonlinear diffusion model with dynamical iterative scheme. Mathematics and Computers in Simulation, 221:533– 549, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001034.

Mei:2024:TTG

[1674] Yusha Mei, Mingrong Cui, and Fanhai Zeng. A time two-grid algorithm for two-dimensional nonlinear time-fractional partial integrodifferential equations. Mathematics and Computers in Simulation, 221: 550–569, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001058.

Rahman:2024:RTM

[1675] M. M. Rahman, Hongqian Zhu, K. Hasan, and Sheng Chen. Replicating transition with modified Spalart– Allmaras model. Mathematics and Computers in Simulation, 221:570– 588, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424000971.

Sunayama:2024:CMM

[1676] Yosuke Sunayama, Julius Fergy Tiongson Rabago, and Masato Kimura. Comoving mesh method for multidimensional moving boundary problems: Mean-curvature flow and Stefan problems. Mathematics and Computers in Simulation, 221:589– 605, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001009.

Long:2024:CTS

[1677] Jianjun Long and Fenglian Wang.

Complexity of a two-stage R&D game within a cluster supply chain considering vertical R&D spillovers, effective information, and government subsidies. Mathematics and Computers in Simulation, 221:606– 621, July 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001071.

Shen:2024:SPD

[1678] Hao Shen and Yongli Song. Spatiotemporal patterns in a diffusive resource– consumer model with distributed memory and maturation delay. Mathematics and Computers in Simulation, 221:622–644, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400106X.

Chen:2024:NMD

[1679] Jianxin Chen, Rui Hou, Tonghua Zhang, and Yongwu Zhou. Newsvendor model for a dyadic supply chain with push-pull strategy under shareholding and risk aversion. Mathematics and Computers in Simulation, 221:645–662, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000922.

Anonymous:2024:NIg

[1680] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 221:663, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001253.

Anonymous:2024:ICEg

[1681] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 221:664, July 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475424001241.

Anonymous:2024:EBh

[1682] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 222:ii, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001708.

Cao:2024:ICMa

[1683] Jinde Cao, S. Saravanan, Ardak Kashkynbayev, M. Muthtamilselvan, and R. Rakkiyappan. International Conference on Mathematical Modeling 2023. Mathematics and Computers in Simulation, 222: 1–2, October 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004688.

Mohanrasu:2024:CBA

[1684] S. S. Mohanrasu, K. Janani, and R. Rakkiyappan. A COPRAS-based approach to multi-label feature selection for text classification. Mathematics and Computers in Simulation, 222:3–23, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003129.

Yashodha:2024:HTA

[1685] S. Yashodha, A. K. Abdul Hakeem, B. Ganga, and P. Renuka. Heat transfer analysis of water–ethylene glycol (50:50) based nanofluid over a cone with the influences of magnetic field and uniform heat generation/absorption. Mathematics and Computers in Simulation, 222:24– 37, August 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002793.

Zhu:2024:HSM

[1686] Jingjuan Zhu, Wenjie Zhang, Lingyun Lu, Yi Lu, and Duo Wang. Hot spot mining and trend analysis of Economic Responsibility Audit based on knowledge graph. Mathematics and Computers in Simulation, 222: 38–49, August 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423003555.

Kirusakthika:2024:MSE

[1687] S. Kirusakthika, S. Priya, A. K. Abdul Hakeem, and B. Ganga. MHD slip effects on (50:50) hybrid nanofluid flow over a moving thin inclined needle with consequences of non-linear thermal radiation, viscous dissipation, and inclined Lorentz force. Mathematics and Computers in Simulation, 222:50–66, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002963.

Zhou:2024:IFS

[1688] Jinwei Zhou and Qi Luo. Influence factor studies based on ensemble learning on the innovation performance of technology mergers and acquisitions. Mathematics and Computers in Simulation, 222:67–89, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002926.

Hingis:2024:GLE

[1689] Y. M. Gifteena Hingis and M. Muthtamilselvan. Ginzburg–Landau equations for the salt fingering region with the onset of microorganisms. Mathematics and Computers in Simulation, 222:90–109, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003245.

Li:2024:HCC

[1690] Saisai Li, Wenke Wang, Le Li, Chuangxia Huang, and Zhaoye Yao. Hierarchical clustering cooperation flocking based on feedback mechanism. Mathematics and Computers in Simulation, 222:110–147, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423003543.

Kavitha:2024:MLF

[1691] S. Kavitha, J. Satheeshkumar, and T. Amudha. Multi-label feature selection using q-rung orthopair hesitant fuzzy MCDM approach extended to CODAS. Mathematics and Computers in Simulation, 222: 148–173, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003282.

Cao:2024:ISS

[1692] Yang Cao, A. Chandrasekar, T. Radhika, and V. Vijayakumar. Inputto-state stability of stochastic Markovian jump genetic regulatory networks. Mathematics and Computers in Simulation, 222:174–187, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475423003336.

Thangaraj:2024:GFI

[1693] C. Thangaraj, D. Easwaramoorthy, Bilel Selmi, and Bhagwati Prasad Chamola. Generation of fractals via iterated function system of Kannan contractions in controlled metric space. Mathematics and Computers in Simulation, 222:188– 198, August 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003440.

Yan:2024:NPP

[1694] Xinhui Yan, Wenke Wang, Chuangxia Huang, and Le Li. A new path planning method for AUV based on the Navier–Stokes equations for ocean currents. Mathematics and Computers in Simulation, 222:199– 208, August 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542300366X.

Agathiyan:2024:ITF

[1695] A. Agathiyan, A. Gowrisankar, and Nur Aisyah Abdul Fataf. On the integral transform of fractal interpolation functions. Mathematics and Computers in Simulation, 222:209– 224, August 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003439.

Hemalatha:2024:SSW

[1696] K. Hemalatha, S. Kumar, and Ikhyun Kim. Study of SH-wave in a prestressed anisotropic magnetoelastic layer sandwich by heterogeneous semiinfinite media. Mathematics and Computers in Simulation, 222:225– 241, August 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003476.

Zeng:2024:QVI

[1697] Runtian Zeng, Qiankun Song, and Shuning Sun. Quaternion version of the Itô’s formula. Mathematics and Computers in Simulation, 222: 242–251, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003774.

Otkel:2024:FTF

[1698] Madina Otkel, Soundararajan Ganesan, Rakkiyappan Rajan, and Ardak Kashkynbayev. Finite-time/fixed-time synchronization of memristive shunting inhibitory cellular neural networks via sliding mode control. Mathematics and Computers in Simulation, 222:252–263, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003506.

Janani:2024:MDM

[1699] K. Janani, S. S. Mohanrasu, Ardak Kashkynbayev, and R. Rakkiyappan. Minkowski distance measure in fuzzy PROMETHEE for ensemble feature selection. Mathematics and Computers in Simulation, 222: 264–295, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003531.

Li:2024:LFC

[1700] Le Li, Lifen Yan, Chuangxia Huang, Jinde Cao, and Xiaodan Ding. Linear formation of Cucker–Smale model with distributed time delays. Mathematics and Computers in Simulation, 222:296–310, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003713.

Keerthana:2024:DDL

[1701] N. Keerthana, R. Saranya, and N. Annapoorani. Dynamics and diffusion limit of traveling waves in a twospecies chemotactic model with logarithmic sensitivity. Mathematics and Computers in Simulation, 222:311– 329, August 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423003701.

Chen:2024:PTS

[1702] Shuting Chen, Ying Wan, Jinde Cao, and Jürgen Kurths. Predefined-time synchronization for competitive neural networks with different time scales and external disturbances. Mathematics and Computers in Simulation, 222:330–349, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003877.

Ragupathi:2024:RLN

[1703] E. Ragupathi and D. Prakash. Role of linear and non-linear thermal radiation over the rotating porous disc with the occurrence of non-uniform heat source/sink: HAM analysis. Mathematics and Computers in Simulation, 222:350–378, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003762.

Zhao:2024:SAD

[1704] Xian Zhao, Chuangxia Huang, Bingwen Liu, and Jinde Cao. Stability analysis of delay patch-constructed Nicholson’s blowflies system. Mathematics and Computers in Simulation, 222:379–392, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004123.

Anonymous:2024:NIh

[1705] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 222:393, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001733.

Anonymous:2024:ICEh

[1706] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 222:394, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424001745.

Anonymous:2024:EBi

[1707] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 223:ii, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002064.

Chen:2024:HBF

[1708] Weixin Chen, Xinzhong Xu, and Qimin Zhang. Hopf bifurcation and fixed-time stability of a reaction– diffusion echinococcosis model with mixed delays. Mathematics and Computers in Simulation, 223:1–19, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001150.

Anonymous:2024:PS

[1709] Anonymous. Pages 1–706 (September 2024). Mathematics and Computers in Simulation, 223:1–706, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Alaraifi:2024:CIE

[1710] Surour Alaraifi, Kaouther Moussa, and Seddik Djouadi. Chemo and immunotherapy effects on stability regions of tumor models. Mathematics and Computers in Simulation, 223: 20–33, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001083.

Li:2024:RSM

[1711] Shunjie Li, Xuebing Zhang, and Qi An. A rumor spreading multidelay model with delay-dependent parameter. Mathematics and Computers in Simulation, 223:34–49, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001307.

Buhmann:2024:NMQ

[1712] Martin Buhmann, Janin Jäger, Joaquín Jódar, and Miguel L. Rodríguez. New methods for quasi-interpolation approximations: Resolution of odddegree singularities. Mathematics and Computers in Simulation, 223: 50–64, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001125.

Wolf:2024:CAM

[1713] Felix L. Wolf, Griselda Deelstra, and Lech A. Grzelak. Consistent asset modelling with random coefficients and switches between regimes. Mathematics and Computers in Simulation, 223:65–85, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001010.

Dong:2024:ESM

[1714] Yunda Dong. Extended splitting methods for systems of three-operator monotone inclusions with continuous operators. Mathematics and Computers in Simulation, 223:86– 107, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001046.

Atef:2024:FTG

[1715] Mohamed Atef and Sifeng Liu. Four types of grey β-covering models and their applications. Mathematics and Computers in Simulation, 223:108– 129, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001137.

Hussain:2024:HKB

[1716] Manzoor Hussain, Abdul Ghafoor, Arshad Hussain, Sirajul Haq, Ihteram Ali, and Shams Ul Arifeen. A hybrid kernel-based meshless method for numerical approximation of multidimensional Fisher’s equation. Mathematics and Computers in Simulation, 223: 130–157, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001265.

Arandiga:2024:NUW

[1717] F. Aràndiga, D. Barrera, S. Eddargani, M. J. Ibáñez, and J. B. Roldán. Non-uniform WENO-based quasi-interpolating splines from the Bernstein–Bézier representation and applications. Mathematics and Computers in Simulation, 223:158–170, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001332.

Zhang:2024:HOE

[1718] Haiqing Zhang and Hong lin Liao. High-order energy stable variablestep schemes for the time-fractional Cahn–Hilliard model. Mathematics and Computers in Simulation, 223: 171–182, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001319.

Aimi:2024:NSN

[1719] A. Aimi, M. A. Leoni, and S. Remogna. Numerical solution of nonlinear Fredholm–Hammerstein integral equations with logarithmic kernel by spline quasi-interpolating projectors. Mathematics and Computers in Simulation, 223:183–194, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001356.

Chowdhury:2024:CID

[1720] Sourav Chowdhury, Jayanta Kumar Ghosh, and Uttam Ghosh. Co-infection dynamics between HIV-HTLV-I disease with the effects of Cytotoxic Tlymphocytes, saturated incidence rate and study of optimal control. Mathematics and Computers in Simulation, 223:195–218, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001411.

Difonzo:2024:CAS

[1721] Fabio V. Difonzo and Sabrina F.

Pellegrino. Convergence analysis of a spectral numerical method for a peridynamic formulation of Richards’ equation. Mathematics and Computers in Simulation, 223:219–228, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001344.

Jha:2024:IAE

[1722] Anjali Jha, Soumitra Pal, and A. K. Misra. Impact of anthropogenic emissions of carbon dioxide and related temperature rise on wildlife population: a modeling study. Mathematics and Computers in Simulation, 223: 229–252, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001320.

He:2024:BCA

[1723] Ke He, Jianping Shi, and Hui Fang. Bifurcation and chaos analysis of a fractional-order delay financial risk system using dynamic system approach and persistent homology. Mathematics and Computers in Simulation, 223: 253–274, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001393.

Saillot:2024:BSC

[1724] Mathis Saillot, Dominique Michel, and Ahmed Zidna. B-spline curve approximation with transformer neural networks. Mathematics and Computers in Simulation, 223:275–287, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001368.

Acal:2024:DPA

[1725] C. Acal, A. M. Aguilera, F. J. Alonso, J. E. Ruiz-Castro, and J. B. Roldán. Different PCA approaches for vector functional time series with applications to resistive switching processes. Mathematics and Computers in Simulation, 223:288–298, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001435.

Oka:2024:LSM

[1726] Tomoyuki Oka. Level set method via positive parts for optimal design problems of two-material thermal conductors. Mathematics and Computers in Simulation, 223:299– 306, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001149.

Rufai:2024:EFO

[1727] Mufutau Ajani Rufai, Bruno Carpentieri, and Higinio Ramos. An efficient fifth-order block method for solving third-order BVPs. Mathematics and Computers in Simulation, 223: 307–321, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400140X.

Mostafazadeh:2024:CAJ

[1728] Mahdi Mostafazadeh and Sedaghat Shahmorad. Convergence analysis of Jacobi spectral tau-collocation method in solving a system of weakly singular Volterra integral equations. Mathematics and Computers in Simulation, 223:322–337, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001496.

Guo:2024:HVL

[1729] Yuling Guo, Xiaoyu Xu, Zicheng Wang, and Zhongqing Wang. An hpversion Legendre collocation method for the third-kind VIEs with nonlinear vanishing delay. Mathematics and Computers in Simulation, 223: 338–350, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001423.

Zhang:2024:UDS

[1730] Chuanlin Zhang, Guoju Ye, Wei Liu, and Xuelong Liu. Upwind difference scheme for fuzzy convection– diffusion equations. Mathematics and Computers in Simulation, 223: 351–367, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001484.

Cuomo:2024:RST

[1731] Salvatore Cuomo, Mariapia De Rosa, Francesco Piccialli, and Laura Pompameo. Railway safety through predictive vertical displacement analysis using the PINN-EKF synergy. Mathematics and Computers in Simulation, 223:368–379, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001526.

Viera-Diaz:2024:MCS

[1732] Rosa Iris Viera-Díaz, Mario González-García, Ricardo Álvarez-Salas, Homero Miranda, Yuniel León-Ruiz, and Víctor Cárdenas. Modeling and control of a single-phase grid-tied mediumfrequency isolated converter for active and reactive power management in photovoltaic applications. Mathematics and Computers in Simulation, 223:380–393, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400137X.

Berenguer:2024:SDI

[1733] M. I. Berenguer, D. Gámez, H. Kunze, D. La Torre, and M. Ruiz Galán. Solving direct and inverse problems for Fredholm-type integro-differential equations with application to pollution diffusion modeling. Mathematics and Computers in Simulation, 223: 394–404, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001472.

Maurya:2024:NAG

[1734] Rahul Kumar Maurya, Dongxia Li, Anant Pratap Singh, and Vineet Kumar Singh. Numerical algorithm for a general fractional diffusion equation. Mathematics and Computers in Simulation, 223:405–432, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001447.

Ordaz-Rivas:2024:IPS

[1735] Erick Ordaz-Rivas and Luis Torres-Treviño. Improving performance in swarm robots using multi-objective optimization. Mathematics and Computers in Simulation, 223:433–457, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001538.

Shang:2024:CDF

[1736] Zuchong Shang and Yuanhua Qiao. Complex dynamics of a four-species food web model with nonlinear top predator harvesting and fear effect. Mathematics and Computers in Simulation, 223:458–484, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001514.

Zerroudi:2024:ASF

[1737] Benaissa Zerroudi, Otheman Nouisser, and Domingo Barrera. Approximate solution of Fredholm integral equations of the second kind through multinode Shepard operators. Mathematics and Computers in Simulation, 223: 485–493, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001642.

Yang:2024:ANS

[1738] Wenbin Yang, Danqing Li, and Xin Chang. Analysis and numerical simulation of computer virus propagation model based on limited resources. Mathematics and Computers in Simulation, 223:494–508, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001617.

Cordero:2024:ITU

[1739] Alicia Cordero, Miguel A. Leonardo-Sepúlveda, Juan R. Torregrosa, and María P. Vassileva. Increasing in three units the order of convergence of iterative methods for solving nonlinear systems. Mathematics and Computers in Simulation, 223:509– 522, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001666.

Cacace:2024:ROC

[1740] Simone Cacace and Alessio Oliviero.

Reliable optimal controls for SEIR models in epidemiology. Mathematics and Computers in Simulation, 223: 523–542, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001605.

Setoudehtazangi:2024:TSC

[1741] F. Setoudehtazangi, T. Manouchehri, A. R. Nematollahi, and M. Caporin. Time series clustering based on latent volatility mixture modeling with applications in finance. Mathematics and Computers in Simulation, 223: 543–564, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001575.

El-Amrani:2024:MSP

[1742] Mofdi El-Amrani, Abdelouahed Ouardghi, and Mohammed Seaid. Modeling and simulation of pollution transport in the Mediterranean Sea using enriched finite element method. Mathematics and Computers in Simulation, 223: 565–587, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400154X.

Alba-Fernandez:2024:TPW

[1743] M. V. Alba-Fernández, M. D. Jiménez-Gamero, and F. Jiménez-Jiménez. Testing for proportions when data are classified into a large number of groups. Mathematics and Computers in Simulation, 223:588–600, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001459.

Yang:2024:LCS

[1744] Hanhua Yang, Mengqing Yan, Wenyong Duan, and Chong Chen. Low conservative stability criteria for discretetime Lur’e systems with sector and slope constrained nonlinearities. Mathematics and Computers in Simulation, 223:601–616, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001630.

Chen:2024:BAC

[1745] Wei Chen, Long Zhang, Ning Wang, and Zhidong Teng. Bifurcation analysis and chaos for a double-strains HIV coinfection model with intracellular delays, saturated incidence and Logistic growth. Mathematics and Computers in Simulation, 223: 617–641, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001502.

Fortes:2024:ANG

[1746] M. A. Fortes, M. Raydan, M. L.

Rodríguez, and A. M. Sajo-Castelli. An assessment of numerical and geometrical quality of bases on surface fitting on Powell–Sabin triangulations. Mathematics and Computers in Simulation, 223:642–653, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001654.

Stocco:2024:PTM

[1747] Davide Stocco, Francesco Biral, and Enrico Bertolazzi. A physical tire model for real-time simulations. Mathematics and Computers in Simulation, 223:654–676, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001460.

Hu:2024:AFE

[1748] Zesen Hu and Xiaolin Li. Analysis of a fast element-free Galerkin method for the multi-term time-fractional diffusion equation. Mathematics and Computers in Simulation, 223:677– 692, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001812.

Cao:2024:ICMb

[1749] Jinde Cao, Raja Ramachandran, Chee Peng Lim, and Fathalla A. Rihan. 2nd International Conference on Mathematical Modeling and Computational Methods in Science and Engineering (ICMMCMSE-2020). Mathematics and Computers in Simulation, 223:693–694, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004627.

urRahman:2024:CEG

[1750] Mujeeb ur Rahman, Tasawar Hayat, Sohail A. Khan, and A. Alsaedi. Corrigendum to “Entropy generation in Sutterby nanomaterials flow due to rotating disk with radiation and magnetic effects” [Math. Comput. Simul. 197 (2022) 151–165]. Mathematics and Computers in Simulation, 223: 695–696, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475422004876. See

[837].

Vujovic:2024:CIE

[1751] Vuk Vujović. Corrigendum to “Influence of environmental fluctuations on Hepatitis C transmission” [Math. Comput. Simulation 191 (2022) 201–218]. Mathematics and Computers in Simulation, 223:697–702, September 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002112. See

[672].

Gao:2024:CRA

[1752] Shan Gao. Corrigendum to “Reliability analysis and optimization for a redundant system with dependent failures and variable repair rates” [Math. Comput. Simulation 208 (2023) 637–659]. Mathematics and Computers in Simulation, 223:703, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003944. See

[1264].

Anonymous:2024:NIi

[1753] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 223:704, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400209X.

Anonymous:2024:ICEi

[1754] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 223:705, September 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475424002106.

Anonymous:2024:EBj

[1755] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 224 (Part A):ii, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002313.

Anonymous:2024:EBk

[1756] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 224 (Part B):ii, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002258.

Pierfederici:2024:PEEa

[1757] Serge Pierfederici. Preface ELECTRI-MACS 2022 — Energy. Mathematics and Computers in Simulation, 224 (Part A):1, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001587.

Pierfederici:2024:PEEb

[1758] Serge Pierfederici. Preface Electrimacs 2022 — Engineering. Mathematics and Computers in Simulation, 224 (Part B):1, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001599.

Celades:2024:TOS

[1759] Eloy Celades, Emilio Pérez, Néstor Aparicio, and Ignacio Peñarrocha-Alós. Tool for optimization of sale and storage of energy in wind farms. Mathematics and Computers in Simulation, 224 (Part B):2–18, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001167.

Pniak:2024:PSM

[1760] Lucas Pniak, Bertrand Revol, Loïc Quéval, Jean-Sylvio Ngoua Teu Magambo, and Olivier Béthoux. Presizing of a modular high power density DC/DC converter with GaN components. Mathematics and Computers in Simulation, 224 (Part A):2– 19, October 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423001337.

Becerra:2024:LBC

[1761] Gerardo Becerra, Fredy Ruiz, Diego Patino, Minh T. Pham, and Xuefang Lin-Shi. Learning-based current estimation for power converters operating in continuous and discontinuous conduction modes. Mathematics and Computers in Simulation, 224 (Part B):19–32, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001155.

Utrilla:2024:SMF

[1762] Candelaria Utrilla, Hossein Dehghani Tafti, Anusha Kumaresan, Jérôme Buire, Vincent Debusschere, Josep Pou, and Nouredine Hadjsaid. A simplified model of flexible power point tracking algorithms in double-stage photovoltaic systems. Mathematics and Computers in Simulation, 224 (Part A):20–33, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001568.

DAmore:2024:IMP

[1763] G. D’Amore, A. Cabrera-Tobar, G. Petrone, A. Massi Pavan, and G. Spagnuolo. Integrating model predictive control and deep learning for the management of an EV charging station. Mathematics and Computers in Simulation, 224 (Part B):33–48, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001714.

Igourzal:2024:ETD

[1764] Ayoub Igourzal, François Auger, Jean-Christophe Olivier, and Clément Retière. Electrical, thermal and degradation modelling of PEMFCs for naval applications. Mathematics and Computers in Simulation, 224 (Part A):34–49, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423001817.

Nazer:2024:VBP

[1765] Afshin Nazer, Patrizio Manganiello, and Olindo Isabella. A virtual bus parallel differential power processing configuration for photovoltaic applications. Mathematics and Computers in Simulation, 224 (Part B):49– 62, October 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002446.

Lappalainen:2024:PMS

[1766] Kari Lappalainen, Michel Piliougine, Seppo Valkealahti, and Giovanni Spagnuolo. Photovoltaic module series resistance identification at its maximum power production. Mathematics and Computers in Simulation, 224 (Part A):50–62, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002409.

LaTona:2024:DAF

[1767] G. La Tona, M. Luna, and M. C.

Di Piazza. Day-ahead forecasting of residential electric power consumption for energy management using Long Short-Term Memory encoder– decoder model. Mathematics and Computers in Simulation, 224 (Part B):63–75, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002720.

Moratalla:2024:FTS

[1768] Manuel Colmenero Moratalla, Ricardo Vidal-Albalate, Francisco R. Blanquez Delgado, and Ramon Blasco-Gimenez. Fault-tolerant strategies in MMCbased high power magnet supply for particle accelerator. Mathematics and Computers in Simulation, 224 (Part A):63–79, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002951.

ElHajj:2024:DIF

[1769] Anthony El Hajj, Eric Semail, Abdelmounaim Tounzi, Darius Vizireanu, and Jalal Cheaytani. Detection of incipient faults in nine-phase machines: Impact of the star winding configuration. Mathematics and Computers in Simulation, 224 (Part B):76–86, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423002847.

Bergeron:2024:OMD

[1770] Célien Bergeron and Stéphane Vivier. Optimization of material distribution applied to magnet-free rotors of synchronous motors. Mathematics and Computers in Simulation, 224 (Part A):80–95, August 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423003804.

Almeida:2024:CEA

[1771] José Almeida, Joao Soares, Fernando Lezama, Zita Vale, and Bruno François. Comparison of evolutionary algorithms for solving risk-based energy resource management considering conditional value-at-risk analysis. Mathematics and Computers in Simulation, 224 (Part B):87– 110, October 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423002902.

Sabrie:2024:TPB

[1772] Antoine Sabrié, Alexandre Battiston, Jean-Yves Gauthier, and Xuefang Lin-Shi. Three-phase bidirectional active split source inverter for automotive traction application. Mathematics and Computers in Simulation, 224 (Part A):96–118, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423003816.

Rigo-Mariani:2024:TSC

[1773] Rémy Rigo-Mariani and Vincent Debusschere. A two-stage coordination strategy for the control of distributed storage at the household level — Arbitrage between users preferences and distribution grid objectives. Mathematics and Computers in Simulation, 224 (Part B):111– 127, August 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003695.

Diab:2024:NME

[1774] Haidar Diab, Yacine Amara, and Georges Barakat. A 3D nonlinear magnetic equivalent circuit model for an axial field flux focusing magnetic gear: Comparison of fixedpoint and Newton–Raphson methods. Mathematics and Computers in Simulation, 224 (Part A):119– 137, October 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423004251.

Aouad:2024:SMO

[1775] Anthony Aouad, Khaled Almaksour, and Dhaker Abbes. Storage management optimization based on electrical consumption and production forecast in a photovoltaic system. Mathematics and Computers in Simulation, 224 (Part B):128– 147, October 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423004391.

Anonymous:2024:NIj

[1776] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 224 (Part A):138, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475424002349.

Anonymous:2024:ICEj

[1777] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 224 (Part A): 139, October 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424002350.

Hoang:2024:DAM

[1778] Quang-Manh Hoang, Bào-Huy Nguy˜ ên, Thanh Vo-Duy, João Pedro F. Trovão, and Minh C. Ta. Decoupled average model-based sliding mode current control of LC-filtered inverters in rotating frame. Mathematics and Computers in Simulation, 224 (Part B):148–164, October 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423005499.

Anonymous:2024:EBl

[1779] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 225:ii, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002714.

Gao:2024:NAF

[1780] Yuefeng Gao, Zhichao Tang, Yuanyuan Ke, and Predrag S. Stanimirović. New activation functions and Zhangians in zeroing neural network and applications to time-varying matrix pseudoinversion. Mathematics and Computers in Simulation, 225:1–12, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001794.

Anonymous:2024:PN

[1781] Anonymous. Pages 1–1278 (November 2024). Mathematics and Computers in Simulation, 225:1–1278, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liu:2024:COP

[1782] Hanjie Liu and Yuanguo Zhu. Carbon option pricing based on uncertain fractional differential equation: a binomial tree approach. Mathematics and Computers in Simulation, 225:13–28, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001800.

Fina:2024:QSV

[1783] Alessandro Fina, Alessandro Gnoatto, and Athena Picarelli. Quantization of stochastic volatility models: Numerical tests and an open source implementation. Mathematics and Computers in Simulation, 225:29– 51, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001563.

Giannelli:2024:GFR

[1784] Carlotta Giannelli, Sofia Imperatore, Lisa Maria Kreusser, Estefanía Loayza-Romero, Fatemeh Mohammadi, and Nelly Villamizar. A general formulation of reweighted least squares fitting. Mathematics and Computers in Simulation, 225:52– 65, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001551.

Urbano-Leon:2024:RMF

[1785] Cristhian Leonardo Urbano-Leon, Ana María Aguilera, and Manuel Escabias. Repeated measures in functional logistic regression. Mathematics and Computers in Simulation, 225: 66–77, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001757.

Kumar:2024:PFD

[1786] Vikas Kumar. Pattern formation and delay-induced instability in a Leslie– Gower type prey–predator system with Smith growth function. Mathematics and Computers in Simulation, 225: 78–97, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001770.

Boushabi:2024:NBS

[1787] M. Boushabi, S. Eddargani, M. J.

Ibáñez, and A. Lamnii. Normalized B-spline-like representation for low-degree Hermite osculatory interpolation problems. Mathematics and Computers in Simulation, 225: 98–110, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001848.

Elahi:2024:FTC

[1788] Arezou Elahi, Alireza Alfi, and Mohammed Chadli. Fixed-time consensus control for uncertain heterogeneous multi-agent systems with high-order dynamics and time-varying delay under generic topologies. Mathematics and Computers in Simulation, 225: 111–128, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001836.

Bossu:2024:CGB

[1789] Adrien Bossu, Benoit Durillon, Arnaud Davigny, Hervé Barry, Benoît Robyns, Fateh Belaïd, and Christophe Saudemont. Coalitional game-based gain generation and distribution for collective self-consumption in an energy community. Mathematics and Computers in Simulation, 225:129– 147, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001381.

Elgindy:2024:FGPb

[1790] Kareem T. Elgindy. Fourier– Gegenbauer pseudospectral method for solving periodic fractional optimal control problems. Mathematics and Computers in Simulation, 225: 148–164, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001769.

Mazzoccoli:2024:RHP

[1791] Alessandro Mazzoccoli, Jorge Andres Rivero, and Pierluigi Vellucci. Refining Heisenberg’s principle: a greedy approximation of step functions with triangular waveform dictionaries. Mathematics and Computers in Simulation, 225:165–176, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542400185X.

Singh:2024:NRS

[1792] Harvindra Singh, L. K. Balyan, A. K. Mittal, and P. Saini. A numerically robust and stable time–space pseudospectral approach for multidimensional generalized Burgers–Fisher equation. Mathematics and Computers in Simulation, 225:177–194, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001782.

Tan:2024:ETA

[1793] Yunshun Tan and Hui Yu. Eventtriggered adaptive tracking for constrained multi-agent systems with saturated inputs and actuator faults. Mathematics and Computers in Simulation, 225:195–215, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001885.

Lv:2024:UBO

[1794] Linjing Lv, Bo Zhang, and Hui Li. An uncertain bi-objective meanentropy model for portfolio selection with realistic factors. Mathematics and Computers in Simulation, 225: 216–231, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001861.

Huang:2024:SSD

[1795] Xin Huang, Siu-Long Lei, Dongfang Li, and Hai-Wei Sun. A stabilized SAV difference scheme and its accelerated solver for spatial fractional Cahn–Hilliard equations. Mathematics and Computers in Simulation, 225: 232–249, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001903.

Li:2024:NAN

[1796] Wenli Li, Xing Liu, and Yanhua Lang. Numerical analysis of a nonlinear agestructured HBV model with saturated incidence and spatial diffusion. Mathematics and Computers in Simulation, 225:250–266, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001952.

Jaglan:2024:SSP

[1797] Jyoti Jaglan, Ankit Singh, Vikas Maurya, Vivek S. Yadav, and Manoj K. Rajpoot. Strong stability preserving multiderivative time marching methods for stiff reaction–diffusion systems. Mathematics and Computers in Simulation, 225:267–282, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424001939.

Yadav:2024:STC

[1798] Reeta Yadav and Moitri Sen. Spatiotemporal complexity in a prey– predator system with Holling type-IV response and Leslie-type numerical response: Turing and steadystate bifurcations. Mathematics and Computers in Simulation, 225:283– 302, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001927.

Casimiro:2024:LCD

[1799] Joyce A. Casimiro and Jaume Llibre. Limit cycles of discontinuous piecewise differential Hamiltonian systems separated by a circle, or a parabola, or a hyperbola. Mathematics and Computers in Simulation, 225: 303–312, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001940.

Rafales:2024:JDP

[1800] Jonatan Ráfales and Carlos Vázquez.

Jump–diffusion productivity models in equilibrium problems with heterogeneous agents. Mathematics and Computers in Simulation, 225:313– 331, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001915.

Liu:2024:NDD

[1801] Guifen Liu, Yangrong Li, and Fengling Wang. Numerical dynamics for discrete nonlinear damping Korteweg– de Vries equations. Mathematics and Computers in Simulation, 225: 332–349, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001988.

Baldi:2024:LGM

[1802] Mauro Maria Baldi, Cristiana Mammana, and Elisabetta Michetti. The κ-logistic growth model. Qualitative and quantitative dynamics. Mathematics and Computers in Simulation, 225:350–369, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001897.

Hmarrass:2024:SBB

[1803] Hanane Hmarrass and Redouane Qesmi. Stability and backward bifurcation for an HIV model with macrophages and CD4+T cells with latent reservoirs. Mathematics and Computers in Simulation, 225:370– 395, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542400199X.

Schmid:2024:CCL

[1804] Harald Schmid. On the connection coefficients for linear differential systems with applications to the spheroidal and ellipsoidal wave equation. Mathematics and Computers in Simulation, 225: 396–411, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002015.

Park:2024:NIA

[1805] Jaesung Park and Gihun Son. Numerical investigation of acoustic droplet vaporization and tissue deformation. Mathematics and Computers in Simulation, 225:412–429, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002118.

Blanc-Blocquel:2024:ELE

[1806] Augusto Blanc-Blocquel, Luis Ortiz-Gracia, and Rodolfo Oviedo. Efficient likelihood estimation of Heston model for novel climate-related financial contracts valuation. Mathematics and Computers in Simulation, 225: 430–445, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001976.

Habibirad:2024:NMM

[1807] Ali Habibirad, Omid Baghani, Hadis Azin, and Esmail Hesameddini. A novel meshless method for time Caputospace Riesz fractional Schrödinger equation. Mathematics and Computers in Simulation, 225:446–460, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002003.

Jiang:2024:LTB

[1808] Dandan Jiang, Mejdi Azaiez, Alain Miranville, Chuanju Xu, and Hui Yao. On the long-time behavior of the continuous and discrete solutions of a nonlocal Cahn–Hilliard type inpainting model. Mathematics and Computers in Simulation, 225: 461–479, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001964.

Hu:2024:PEI

[1809] Po Hu, Hongbin Ma, Xiaoguang Yang, and Yifen Mu. Pareto efficiency of infinite-horizon cooperative stochastic differential games with Markov jumps and Poisson jumps. Mathematics and Computers in Simulation, 225: 480–494, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001629.

Xu:2024:DSF

[1810] Xin Xu, Baodan Tian, Xingzhi Chen, and Yanhong Qiu. Dynamics of a stochastic food chain chemostat model with Monod–Haldane functional response and Ornstein– Uhlenbeck process. Mathematics and Computers in Simulation, 225:495– 512, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001873.

Arandiga:2024:EWC

[1811] F. Aràndiga, D. Barrera, and S. Eddargani. ENO and WENO cubic quasiinterpolating splines in Bernstein– Bézier form. Mathematics and Computers in Simulation, 225:513–527, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542400212X.

Durante:2024:NSV

[1812] D. Durante, S. Marrone, D. Brömmel, R. Speck, and A. Colagrossi. Numerical simulation of 3D vorticity dynamics with the Diffused Vortex Hydrodynamics method. Mathematics and Computers in Simulation, 225:

528–544, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002143.

Chahal:2024:GWA

[1813] Parmeet Kaur Chahal, Kamlesh Kumar, and Bhavneet Singh Soodan. Grey wolf algorithm for cost optimization of cloud computing repairable system with N-policy, discouragement and two-level Bernoulli feedback. Mathematics and Computers in Simulation, 225:545–569, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002167.

Yin:2024:NHO

[1814] Fengli Yin, Zhuangzhi Xu, and Yayun Fu. Novel high-order explicit energy-preserving schemes for NLStype equations based on the Liegroup method. Mathematics and Computers in Simulation, 225:570– 585, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002027.

Cuevas:2024:IAM

[1815] Erik Cuevas, Oscar Barba-Toscano, Héctor Escobar, Daniel Zaldívar, and Alma Rodríguez-Vázquez. An initialization approach for metaheuristic algorithms by using Gibbs sampling. Mathematics and Computers in Simulation, 225:586–606, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424001824.

Sefti:2024:CBS

[1816] Rania Sefti, Driss Sbibih, and Rachid Jennane. A CNN-based spline active surface method with an afterbalancing step for 3D medical image segmentation. Mathematics and Computers in Simulation, 225:607– 618, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424002131.

Yuan:2024:ASS

[1817] Shilei Yuan, Yantao Wang, Xian Zhang, and Xin Wang. H_{∞} antisynchronization of switching inertial neural networks with leakage delays and mixed time-varying delays. Mathematics and Computers in Simulation, 225:619–647, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002155.

Kim:2024:SIF

[1818] Hyundong Kim, Jyoti, Soobin Kwak, Seokjun Ham, and Junseok Kim. In silico investigation of the formation of multiple intense zebra stripes using extending domain. Mathematics and Computers in Simulation, 225: 648–658, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002210.

Safdar:2024:EID

[1819] Muhammad Umair Safdar, Tariq Shah, and Asif Ali. Enhancing image data security with chain and non-chain Galois ring structures. Mathematics and Computers in Simulation, 225: 659–694, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002192.

Wang:2024:URO

[1820] Yan Wang, Hongjun Peng, and Xin Chen. Uncertain random optimal control model for deteriorating inventory with the finite horizon. Mathematics and Computers in Simulation, 225: 695–715, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002180.

Babaei:2024:SCT

[1821] Maryam Babaei, Alireza Afzal Aghaei, Zahra Kazemi, Mahdieh Jamshidi, Reza Ghaderi, and Kourosh Parand. Solving a class of Thomas–Fermi equations: a new solution concept based on physics-informed machine learning. Mathematics and Computers in Simulation, 225:716– 730, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002209.

Wei:2024:DBS

[1822] Su Wei, Daqing Jiang, and Yaxin Zhou. Dynamic behavior of a stochastic HIV model with latent infection and Ornstein–Uhlenbeck process. Mathematics and Computers in Simulation, 225:731–759, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002362.

Meena:2024:IRR

[1823] V. P. Meena, V. P. Singh, and Josep M. Guerrero. Investigation of reciprocal rank method for automatic generation control in two-area interconnected power system. Mathematics and Computers in Simulation, 225: 760–778, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002179.

Ferri-Garcia:2024:ERP

[1824] Ramón Ferri-García, Jorge L. Rueda-Sánchez, María del Mar Rueda, and Beatriz Cobo. Estimating response propensities in nonprobability surveys using machine learning weighted models. Mathematics and Computers in Simulation, 225:779– 793, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002374.

Toledo:2024:PSD

[1825] F. Javier Toledo, Vicente Galiano, Jose M. Blanes, Victoria Herranz, and Efstratios Batzelis. Photovoltaic singlediode model parametrization. An application to the calculus of the Euclidean distance to an I–V curve. Mathematics and Computers in Simulation, 225:794–819, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423000058.

Magdalena:2024:NAB

[1826] I. Magdalena, D. N. Haloho, and M. B. Adityawan. Numerical approaches for Boussinesq type equations with its application in Kampar River, Indonesia. Mathematics and Computers in Simulation, 225: 820–834, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423002070.

Lu:2024:MSP

[1827] Kang-Ya Lu and Cun-Qiang Miao.

Matrix splitting preconditioning based on sine transform for solving twodimensional space-fractional diffusion equations. Mathematics and Computers in Simulation, 225:835–856, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423003828.

Kumar:2024:CSF

[1828] Ajay Kumar, Sunil Kumar, Shaher Momani, and Samir Hadid. A chaos study of fractal–fractional predator–prey model of mathematical ecology. Mathematics and Computers in Simulation, 225:857–888, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300410X.

Ruby:2024:CAN

[1829] Ruby and Moumita Mandal. Convergence analysis and numerical implementation of projection methods for solving classical and fractional Volterra integro-differential equations. Mathematics and Computers in Simulation, 225:889–913, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004159.

Mederos:2024:ISA

[1830] Victoria Hernández Mederos, Eduardo Moreno Hernández, Jorge Estrada Sarlabous, Isidro A. Abelló Ugalde, and Domenico Lahaye. Isogeometric simulation of acoustic radiation. Mathematics and Computers in Simulation, 225:914–938, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004184.

Das:2024:QAT

[1831] Riya Das, Dhiraj Kumar Das, and Tapan Kumar Kar. Qualitative analysis of TB transmission dynamics considering both the age since latency and relapse. Mathematics and Computers in Simulation, 225: 939–967, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004214.

Mondal:2024:GPP

[1832] Bapin Mondal, Uttam Ghosh, Susmita Sarkar, and Pankaj Kumar Tiwari. A generalist predator–prey system with the effects of fear and refuge in deterministic and stochastic environments. Mathematics and Computers in Simulation, 225:968– 991, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004226.

Mokni:2024:RDD

[1833] Karima Mokni, Mohamed Ch-Chaoui, Bapin Mondal, and Uttam Ghosh. Rich dynamics of a discrete two dimensional predator–prey model using the NSFD scheme. Mathematics and Computers in Simulation, 225: 992–1018, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542300424X.

Wu:2024:MAA

[1834] Chia-Huang Wu, Dong-Yuh Yang, and Ting-En He. Matrix-augmentation approach for machine repair problem with generally distributed repair times during working breakdown periods. Mathematics and Computers in Simulation, 225:1019– 1038, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004263.

Wu:2024:AAL

[1835] Juntao Wu, Xiao Wang, and Yicheng Liu. Asymptotic analysis of the linear formation model with an undirected connected topology. Mathematics and Computers in Simulation, 225:

1039–1055, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004305.

Butt:2024:NSS

[1836] M. M. Butt and S. Roy. A numerical scheme to solve Fokker–Planck control collective-motion problem. Mathematics and Computers in Simulation, 225: 1056–1071, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004317.

Tatlicioglu:2024:SCS

[1837] Buğçe Eminağa Tatlıcıoğlu. A simple chaotic system using signum function. Mathematics and Computers in Simulation, 225:1072–1088, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475423004640.

Chukwu:2024:FMH

[1838] C. W. Chukwu, Fatmawati, M. I. Utoyo, A. Setiawan, and J. O. Akanni. Fractional model of HIV transmission on workplace productivity using real data from Indonesia. Mathematics and Computers in Simulation, 225: 1089–1103, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475423004755.

Huang:2024:BBO

[1839] Song Huang, Zhijun Liu, and Lianwen Wang. Backward bifurcation and optimal control problem for a tuberculosis model incorporating LTBI detectivity and exogenous reinfection. Mathematics and Computers in Simulation, 225: 1104–1123, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004792.

Houssein:2024:EBW

[1840] Essam H. Houssein, Mahmoud Khalaf Saeed, and Mustafa M. Al-Sayed. EWSO: Boosting White Shark Optimizer for solving engineering design and combinatorial problems. Mathematics and Computers in Simulation, 225:1124–1153, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475423004809.

Janani:2024:OBF

[1841] K. Janani, R. Baranitha, Chee Peng Lim, and R. Rakkiyappan. Observerbased fuzzy integral sliding mode control for bilateral teleoperation systems with time-varying delays. Mathematics and Computers in Simulation, 225: 1154–1169, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475423004822.

Yamni:2024:ECI

[1842] Mohamed Yamni, Achraf Daoui, and Ahmed A. Abd El-Latif. Efficient color image steganography based on new adapted chaotic dynamical system with discrete orthogonal moment transforms. Mathematics and Computers in Simulation, 225:1170– 1198, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424000351.

Zhang:2024:TFD

[1843] Hao Zhang, Mengmeng Liu, Tao Guo, and Da Xu. Three finite difference schemes for generalized nonlinear integro-differential equations with tempered singular kernel. Mathematics and Computers in Simulation, 225: 1199–1217, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000387.

Zhang:2024:DKA

[1844] Xuetian Zhang, Chunrui Zhang, and Yazhuo Zhang. Discrete kinetic analysis of a general reaction–diffusion model constructed by Euler discretization and coupled map lattices. Mathematics and Computers in Simulation, 225: 1218–1236, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001095.

Fusi:2024:SCS

[1845] Lorenzo Fusi and Antonio Giovinetto. A spectral collocation scheme for the two dimensional flow of a regularized viscoplastic fluid: Numerical results and comparison with analytical solution. Mathematics and Computers in Simulation, 225:1237– 1256, November 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424001101.

Garralon-Lopez:2024:NAP

[1846] Rubén Garralon-López, Francisco Rus, and Francisco R. Villatoro. Numerical adiabatic perturbation theory for the absolute |K|(p, p) equation. Mathematics and Computers in Simulation, 225: 1257–1276, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424001113.

Anonymous:2024:NIk

[1847] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 225:1277, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002866.

Anonymous:2024:ICEk

[1848] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 225:1278, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424002878.

Anonymous:2024:EBm

[1849] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 226:ii, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003306.

Qian:2024:EHH

[1850] Shiwen Qian, Jing Wang, and Longxing Qi. The effect of human heterogeneity on the transmission of viral respiratory diseases under air pollution. Mathematics and Computers in Simulation, 226:1–23, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002404.

Anonymous:2024:PD

[1851] Anonymous. Pages 1–664 (December 2024). Mathematics and Computers in Simulation, 226:1–664, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Lee:2024:ORS

[1852] Hyosun Lee, Arsen Abdulali, Haeyoung Park, and Sunmi Lee. Optimal region-specific social distancing strategies in a complex multi-patch model through reinforcement learning. Mathematics and Computers in Simulation, 226:24–41, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002386.

Wu:2024:CAC

[1853] Jingwen Wu and Zhijun Tan. A conservative Allen–Cahn model for a hydrodynamics coupled phase-field surfactant system. Mathematics and Computers in Simulation, 226:42– 65, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002465.

Mohana:2024:NSA

[1854] C. M. Mohana and B. Rushi Kumar. Numerical and semi-analytical approaches for heat transfer analysis of ternary hybrid nanofluid flow: a comparative study. Mathematics and Computers in Simulation, 226: 66–90, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002507.

Dey:2024:GDT

[1855] Subir Dey, Tapan Kumar Kar, and Toshikazu Kuniya. Global dynamics and threshold behavior of an SEIR epidemic model with nonlocal diffusion. Mathematics and Computers in Simulation, 226:91– 117, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002532.

Kulekci:2024:RPH

[1856] Bükre Yıldırım Külekci, Ralf Korn, and A. Sevtap Selcuk-Kestel. Ruin probability for heavy-tailed and dependent losses under reinsurance strategies. Mathematics and Computers in Simulation, 226:118–138, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002489.

Sun:2024:DSP

[1857] Yanru Sun, Hao Sun, Panfei Sun, and Xuanzhu Jin. A dynamic surge pricing model throughout product lifecycle. Mathematics and Computers in Simulation, 226:139–151, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424002477.

Raeisi:2024:RBP

[1858] Banafsheh Raeisi, Mojtaba Fardi, and Mohammadreza Ahmadi Darani. RBFbased partition of unity methods for two-dimensional time-dependent PDEs: Numerical and theoretical aspects. Mathematics and Computers in Simulation, 226:152–171, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424002520.

Shokrollahi:2024:PAO

[1859] F. Shokrollahi, D. Ahmadian, and L. V. Ballestra. Pricing Asian options under the mixed fractional Brownian motion with jumps. Mathematics and Computers in Simulation, 226: 172–183, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002398.

Ahmadinia:2024:NSM

[1860] M. Ahmadinia, M. Abbasi, and A. R. Emrani. A novel A-stable method of order of accuracy three, based on optimization for solving initial value problems. Mathematics and Computers in Simulation, 226:184– 203, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002519.

Rakkiyappan:2024:STS

[1861] R. Rakkiyappan, V. Sharmila, K. Janani, and Ardak Kashkynbayev. Stabilization of Takagi–Sugeno fuzzy Hidden Markov Jump Systems with memory sampled-data control. Mathematics and Computers in Simulation, 226: 204–217, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002660.

Bhargava:2024:TCD

[1862] Masoom Bhargava and Balram Dubey.

Trade-off and chaotic dynamics in a two-prey one-predator model with refuge, environmental noise and seasonal effects. Mathematics and Computers in Simulation, 226:218– 245, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002623.

Molina-Diaz:2024:DSM

[1863] Oscar Emilio Molina-Díaz, Gerard Olivar-Tost, and Deissy Milena Sotelo-Castelblanco. Dynamics of a stockless market through piecewise smooth models. Mathematics and Computers in Simulation, 226:246–269, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542400260X.

Zhang:2024:PIS

[1864] Zidie Zhang, Daiyong Wu, and Nishan Li. Predator invasion in a spatially heterogeneous predator-prey model with group defense and prey-taxis. Mathematics and Computers in Simulation, 226:270–282, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002659.

Rashi:2024:EFS

[1865] Rashi, Harendra Pal Singh, and Suruchi Singh. Effect of fear with saturated fear cost and harvesting on aquatic food chain model (plankton–fish model) in the presence of nanoparticles. Mathematics and Computers in Simulation, 226: 283–305, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002611.

Zhao:2024:REP

[1866] Meiling Zhao, Xiaochen Fei, Liqun Wang, and Zhanbin Yuan. Reconstruction of electromagnetic parameters of the scattering from overfilled cavities using a complex-valued convolutional neural network. Mathematics and Computers in Simulation, 226: 306–322, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002672.

Liu:2024:SDC

[1867] Xiaoyuan Liu and Min Cai. Spectral deferred correction method for fractional initial value problem with Caputo–Hadamard derivative. Mathematics and Computers in Simulation, 226:323–337, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002581.

Sanfelici:2024:FMV

[1868] Simona Sanfelici and Giacomo Toscano.

The Fourier–Malliavin Volatility (FMVol) MATLAB(R) library. Mathematics and Computers in Simulation, 226: 338–353, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002568.

Salihu:2024:ENL

[1869] Nasiru Salihu, Poom Kumam, Ibrahim Mohammed Sulaiman, Ibrahim Arzuka, and Wiyada Kumam. An efficient Newton-like conjugate gradient method with restart strategy and its application. Mathematics and Computers in Simulation, 226: 354–372, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002593.

Yue:2024:BAS

[1870] Hongwei Yue and Peiping Shen. Biaffine scaling iterative method for convex quadratic programming with bound constraints. Mathematics and Computers in Simulation, 226:373– 382, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002647.

Feugeas:2024:APK

[1871] Jean-Luc Feugeas, Julien Mathiaud, Luc Mieussens, and Thomas Vigier. An asymptotic preserving kinetic scheme for the M1 model of linear transport. Mathematics and Computers in Simulation, 226: 383–398, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002751.

Li:2024:GMC

[1872] Haiyang Li and Junjie Ma. On generalized multistep collocation methods for Volterra integro-differential equations. Mathematics and Computers in Simulation, 226:399–412, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002763.

Ran:2024:GSS

[1873] Jie Ran, Yonghui Zhou, and Hao Pu. Global stability and synchronization of stochastic discrete-time variable-order fractional-order delayed quaternionvalued neural networks. Mathematics and Computers in Simulation, 226: 413–437, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400274X.

Kaur:2024:NSV

[1874] Manpreet Kaur and Mamta Kapoor.

Numerical simulation of Volterra PIDE with singular kernel via modified cubic exponential and uniform algebraic trigonometric tension Bspline DQM. Mathematics and Computers in Simulation, 226:438– 451, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002829.

Kim:2024:DOL

[1875] In Ae Kim, Yong Jun Kim, Yong Son Choe, and Chol Su Han. Discreteordinate-lattice-Boltzmann Method for analyzing radiative heat transfer: Application to two-dimensional irregular enclosure. Mathematics and Computers in Simulation, 226:452– 463, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002490.

Hernandez-Castorena:2024:LSA

[1876] Gerardo Armando Hernandez-Castorena, Maria Aracelia Alcorta-Garcia, Jose Armando Saenz-Esqueda, and Gerardo Maximiliano Mendez. Lyapunov’s stability analysis for first degree polynomial systems, subject to risk-sensitive control. Mathematics and Computers in Simulation, 226: 464–473, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542400257X.

Li:2024:DAF

[1877] Zhenlei Li and Yue Zhang. Dynamic analysis of a fast slow modified Leslie–Gower predator–prey model with constant harvest and stochastic factor. Mathematics and Computers in Simulation, 226:474–499, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424002842.

Arora:2024:ISA

[1878] Gourav Arora, Rajesh Kumar, and Helmi Temimi. An implicit semianalytical technique: Development, analysis and applications. Mathematics and Computers in Simulation, 226: 500–510, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002775.

Oubarka:2024:RWB

[1879] Ismail Oubarka, Imad Kissami, Imad Elmahi, and Eric Deleersnijder. A robust and well-balanced finite volume solver for investigating the effects of tides on water renewal timescale in the Nador lagoon, Morocco. Mathematics and Computers in Simulation, 226:511–523, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002556.

Culda:2024:DDT

[1880] Loredana Camelia Culda, Eva Kaslik, Mihaela Neamţu, and Nicoleta Sîrghi. Dynamics of a discrete-time mixed oligopoly Cournot-type model with three time delays. Mathematics and Computers in Simulation, 226:524– 539, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002830.

Rebollo:2024:DDS

[1881] Tomás Chacón Rebollo and Daniel Franco Coronil. Data-driven stabilized finite element solution of advectiondominated flow problems. Mathematics and Computers in Simulation, 226: 540–559, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002787.

Yesudhas:2024:SSD

[1882] Anto Anbarasu Yesudhas, Subramanian Kuppusamy, Syeong Ryong Lee, Jae Hoon Jeong, and Young Hoon Joo. Switched sampled-data-based membership function-dependent H_{∞} control for PMSG-based WTS with actuator failures. Mathematics and Computers in Simulation, 226:560– 577, December 2024. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002805.

Hao:2024:DAO

[1883] Wenhui Hao, Juping Zhang, and Zhen Jin. Dynamic analysis and optimal control of HIV/AIDS model with ideological transfer. Mathematics and Computers in Simulation, 226: 578–605, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002635.

El-Mesady:2024:IPI

[1884] A. El-Mesady and Hegagi Mohamed Ali. The influence of prevention and isolation measures to control the infections of the fractional Chickenpox disease model. Mathematics and Computers in Simulation, 226: 606–630, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002854.

Zhou:2024:SOE

[1885] Yongtao Zhou and Wei Xu. Secondorder error analysis of a corrected average finite difference scheme for time-fractional Cable equations with nonsmooth solutions. Mathematics and Computers in Simulation, 226: 631–644, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400288X.

Liang:2024:URC

[1886] Yanming Liang, Yongfeng Guo, and Zifei Lin. Using reservoir computing to solve FPK equations for stochastic dynamical systems under Gaussian or Non-Gaussian excitation. Mathematics and Computers in Simulation, 226:645–662, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002817.

Anonymous:2024:NIl

[1887] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 226:663, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003331.

Anonymous:2024:ICEl

[1888] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 226:664, December 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424003343.

Anonymous:2025:EBa

[1889] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 227:ii, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003537.

Zhang:2025:EDD

[1890] Xuebing Zhang, Qi An, and Ali Moussaoui. Effect of density-dependent diffusion on a diffusive predator– prey model in spatially heterogeneous environment. Mathematics and Computers in Simulation, 227: 1–18, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424002799.

Anonymous:2025:PJa

[1891] Anonymous. Pages 1–570 (January 2025). Mathematics and Computers in Simulation, 227:1–570, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ramasamy:2025:SHB

[1892] Sivasamy Ramasamy, David Banjerdpongchai, and PooGyeon Park. Stability and Hopf-bifurcation analysis of diffusive Leslie–Gower prey– predator model with the Allee effect and carry-over effects. Mathematics and Computers in Simulation, 227:19–40, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002982.

Ha:2025:PPA

[1893] Mijin Ha, Donghyun Kim, Ji-Hun Yoon, and Sun-Yong Choi. Pricing for perpetual American strangle options under stochastic volatility with fast mean reversion. Mathematics and Computers in Simulation, 227:41–57, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424002908.

Zourmba:2025:FOM

[1894] Kotadai Zourmba, Joseph Yves Effa, Clovis Fischer, José David Rodríguez-Muñoz, Maria Fernanda Moreno-Lopez, Esteban Tlelo-Cuautle, and Jean De Dieu Nkapkop. Fractional order 1D memristive timedelay chaotic system with application to image encryption and FPGA implementation. Mathematics and Computers in Simulation, 227:58– 84, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002970.

Chen:2025:NMF

[1895] C. S. Chen and Andreas Karageorghis. Novel method of fundamental solutions formulation for polyharmonic BVPs. Mathematics and Computers in Simulation, 227:85– 102, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002969.

Gao:2025:DID

[1896] Yujie Gao, Malay Banerjee, and Ton Viet Ta. Dynamics of infectious diseases in predator–prey populations: a stochastic model, sustainability, and invariant measure. Mathematics and Computers in Simulation, 227:103–120, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002891.

Balam:2025:TFD

[1897] R. Itzá Balam, M. Uh Zapata, and U. Iturrarán-Viveros. Triangular finite differences using bivariate Lagrange polynomials with applications to elliptic equations. Mathematics and Computers in Simulation, 227: 121–148, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003008.

Song:2025:RAD

[1898] Xiaona Song, Danjing Zheng, Shuai Song, Vladimir Stojanovic, and Inés Tejado. Robust anti-disturbance interval type-2 fuzzy control for interconnected nonlinear PDE systems via conjunct observer. Mathematics and Computers in Simulation, 227: 149–167, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003021.

Dwivedi:2025:NFS

[1899] Himanshu Kumar Dwivedi and Rajeev. A novel fast second order approach with high-order compact difference scheme and its analysis for the tempered fractional Burgers equation. Mathematics and Computers in Simulation, 227:168– 188, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003057.

Taugynbayeva:2025:FCD

[1900] Galiya Taugynbayeva, Shapen Azhgaliyev, Aksaule Zhubanysheva, and Nurlan Temirgaliyev. Full C(N)Dstudy of computational capabilities of Lagrange polynomials. Mathematics and Computers in Simulation, 227: 189–208, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424002957.

Tajudeen:2025:NFC

[1901] M. Mubeen Tajudeen, M. Syed Ali, Ganesh Kumar Thakur, Bandana Priya, and R. Perumal. Non-fragile control of discrete-time conic-type nonlinear Markovian jump systems under deception attacks using event-triggered scheme and Its application. Mathematics and Computers in Simulation, 227:209–225, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003100.

Li:2025:WCS

[1902] Yan Li, Qiuhong Xu, and Wanrong Cao. Weak convergence of the split-step backward Euler method for stochastic delay integrodifferential equations. Mathematics and Computers in Simulation, 227: 226–240, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003070.

Singh:2025:HON

[1903] Ajeet Singh, Hanz Martin Cheng, Naresh Kumar, and Ram Jiwari. A high order numerical method for analysis and simulation of 2D semilinear Sobolev model on polygonal meshes. Mathematics and Computers in Simulation, 227:241– 262, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003124.

Xing:2025:DOB

[1904] Ying Xing, Xinyi He, and Xiaodi Li. Disturbance observer-based eventtriggered impulsive control for nonlinear systems with unknown external disturbances. Mathematics and Computers in Simulation, 227:263– 271, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003148.

Cobo:2025:NTH

[1905] Beatriz Cobo, Jorge Luis Rueda-Sánchez, Ramón Ferri-García, and María del Mar Rueda. A new technique for handling non-probability samples based on model-assisted kernel weighting. Mathematics and Computers in Simulation, 227:272– 281, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003094.

Farman:2025:FOF

[1906] Muhammad Farman, Khadija Jamil, Changjin Xu, Kottakkaran Sooppy Nisar, and Ayesha Amjad. Fractional order forestry resource conservation model featuring chaos control and simulations for toxin activity and human-caused fire through modified ABC operator. Mathematics and Computers in Simulation, 227: 282–302, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400301X.

Liang:2025:PEL

[1907] Yijie Liang, Qinghui Zhang, and Shaojie Zeng. A piecewise extreme learning machine for interface problems. Mathematics and Computers in Simulation, 227:303–321, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003112.

Cuevas:2025:BIC

[1908] Erik Cuevas, Mario Vásquez, Karla Avila, Alma Rodriguez, and Daniel Zaldivar. Balancing individual and collective strategies: a new approach in metaheuristic optimization. Mathematics and Computers in Simulation, 227:322–346, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003069.

Stein:2025:STL

[1909] Andreas Stein and Andrea Barth.

Stochastic transport with Lévy noise fully discrete numerical approximation. Mathematics and Computers in Simulation, 227:347–370, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002994.

Imai:2025:DRQ

[1910] Junichi Imai and Ken Seng Tan. Dimension reduction for Quasi-Monte Carlo methods via quadratic regression. Mathematics and Computers in Simulation, 227:371–390, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003185.

Niu:2025:MOS

[1911] Yanbiao Niu, Xuefeng Yan, Weiping Zeng, Yongzhen Wang, and Yanzhao Niu. Multi-objective sand cat swarm optimization based on adaptive clustering for solving multimodal multi-objective optimization problems. Mathematics and Computers in Simulation, 227:391– 404, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003252.

Huang:2025:CAN

[1912] Qianru Huang, Qinwei Fan, Zhiwei Xing, Xiaofei Yang, and Xingshi He. Convergence analysis of a novel high order networks model based on entropy error function. Mathematics and Computers in Simulation, 227: 405–419, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003161.

Fan:2025:DAT

[1913] Yongqing Fan, Yu Zhang, and Zhen Li. Distributed adaptive tracking consensus control for a class of heterogeneous nonlinear multi-agent systems. Mathematics and Computers in Simulation, 227:420–441, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003240.

Paul:2025:EFN

[1914] Biswajit Paul, Gopal Chandra Sikdar, and Uttam Ghosh. Effect of fear and non-linear predator harvesting on a predator–prey system in presence of environmental variability. Mathematics and Computers in Simulation, 227:442–460, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003239.

Kelly:2025:SCC

[1915] Cónall Kelly, Gabriel J. Lord, and Fandi Sun. Strong convergence of a class of adaptive numerical methods for SDEs with jumps. Mathematics and Computers in Simulation, 227:461–476, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003227.

Zhu:2025:CCN

[1916] Zirui Zhu and Xingbo Liu. Canard cycle and nonsmooth bifurcation in a piecewise-smooth continuous predator-prey model. Mathematics and Computers in Simulation, 227:477–499, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003197.

Xue:2025:LQO

[1917] Xiaomin Xue, Juanjuan Xu, and Huanshui Zhang. Linear quadratic optimal control of stochastic 2-D Roesser models. Mathematics and Computers in Simulation, 227:500– 510, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003410.

Jyoti:2025:ERT

[1918] Jyoti and Mandeep Singh. An efficient recursive technique with Padé approximation for a kind of Lane– Emden type equations emerging in various physical phenomena. Mathematics and Computers in Simulation, 227:511–526, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003355.

Mandal:2025:DCP

[1919] Gourav Mandal, Lakshmi Narayan Guin, Santabrata Chakravarty, and Renji Han. Dynamic complexities in a predator–prey model with prey refuge influenced by double Allee effects. Mathematics and Computers in Simulation, 227:527– 552, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003173.

Huang:2025:CDB

[1920] Zhipeng Huang, Balint Negyesi, and Cornelis W. Oosterlee. Convergence of the deep BSDE method for stochastic control problems formulated through the stochastic maximum principle. Mathematics and Computers in Simulation, 227:553– 568, January 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003045.

Anonymous:2025:NIa

[1921] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 227:569, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003574.

Anonymous:2025:ICEa

[1922] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 227:570, January 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424003586.

Tasse:2025:NFD

[1923] Arsène J. Ouemba Tassé, Vuyiswa B.

Kubalasa, Berge Tsanou, Jean M.-S, and Lubuma. Nonstandard finite difference schemes for some epidemic optimal control problems. Mathematics and Computers in Simulation, 228 (??):1–22, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003380.

Romanengo:2025:DHP

[1924] Chiara Romanengo, Bianca Falcidieno, and Silvia Biasotti. Discretisation of the Hough parameter space for fitting and recognising geometric primitives in 3D point clouds. Mathematics and Computers in Simulation, 228 (??):73–86, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003458.

Blanco:2025:MLI

[1925] Beatriz Blanco, Roberto Palma, Manuel Hurtado, Gema Jiménez, Carmen Griñán-Lisón, Juan Melchor, Juan Antonio Marchal, Hector Gomez, Guillermo Rus, and Juan Soler. Modeling low-intensity ultrasound mechanotherapy impact on growing cancer stem cells. Mathematics and Computers in Simulation, 228(??): 87–102, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003434.

Bracco:2025:LSR

[1926] Cesare Bracco, Carlotta Giannelli, Francesco Patrizi, and Alessandra Sestini. Local spline refinement driven by fault jump estimates for scattered data approximation. Mathematics and Computers in Simulation, 228(??): 103–123, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003422.

Vincent:2025:MCI

[1927] Julien Vincent, Alberto Tenore, Maria Rosaria Mattei, and Luigi Frunzo. Modelling the comparative influence of conjugation and transformation on plasmid spread in biofilms. Mathematics and Computers in Simulation, 228(??):156– 177, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003203.

Cristiani:2025:KDM

[1928] Emiliano Cristiani, Nadia Loy, Marta Menci, and Andrea Tosin. Kinetic description and macroscopic limit of swarming dynamics with continuous leader–follower transitions. Mathematics and Computers in Simulation, 228 (??):362–385, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003616.

Abbas:2025:PMV

[1929] Mudassar Abbas, Francesco Giannino, Annalisa Iuorio, Zubair Ahmad, and Francesco Calabró. PDE models for vegetation biomass and autotoxicity. Mathematics and Computers in Simulation, 228(??):386– 401, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424002544.

Barrena:2025:FMW

[1930] E. Barrena, S. Bermudo, A. G.

Hernández-Díaz, A. D. López-Sánchez, and J. A. Zamudio. Finding the minimum k-weighted dominating sets using heuristic algorithms. Mathematics and Computers in Simulation, 228 (??):485–497, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003653.

Su:2025:GTS

[1931] Yichen Su and Leevan Ling. Greedy trial subspace selection in meshfree time-stepping scheme with applications in coupled bulk-surface pattern formations. Mathematics and Computers in Simulation, 228(??): 498–513, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003732.

Bastl:2025:ARI

[1932] Bohumír Bastl and Kristýna Slabá. Adaptive refinement in incompressible fluid flow simulation based on THB-splines-powered isogeometric analysis. Mathematics and Computers in Simulation, 228(??):514– 533, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003719.

Naimi:2025:SLE

[1933] M. Naimi and M. Lamnii. Solving the Laplace equation on the disc using the UAT spline. Mathematics and Computers in Simulation, 228(??): 534–548, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003598.

Anonymous:2025:NIb

[1934] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 228(??):549, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004208.

Anonymous:2025:ICEb

[1935] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 228(??):550, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542400421X.

Anonymous:2025:EBb

[1936] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 228:ii, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004178.

Anonymous:2025:PF

[1937] Anonymous. Pages 1–550 (February 2025). Mathematics and Computers in Simulation, 228:1–550, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Zhao:2025:SCF

[1938] Ailiang Zhao, Junmin Li, and Aili Fan. Stabilization for a class of fractional-order nonlinear reaction– diffusion systems with time-varying delay: Event-triggered boundary control approach. Mathematics and Computers in Simulation, 228:23– 38, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003367.

Zhang:2025:UCF

[1939] Jin Zhang and Chunxiao Zhang. Uniform convergence of finite element method on Bakhvalov-type mesh for a 2-D singularly perturbed convection–diffusion problem with exponential layers. Mathematics and Computers in Simulation, 228:39– 51, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003446.

Campos:2025:CPP

[1940] Beatriz Campos, Jordi Canela, Alberto Rodríguez-Arenas, and Pura Vindel. Computing parameter planes of iterative root-finding methods with several free critical points. Mathematics and Computers in Simulation, 228:52–72, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400315X.

Cerqueti:2025:CMD

[1941] Roy Cerqueti, Alessandro Ramponi, and Sergio Scarlatti. A compartmental model for the dynamic simulation of pandemics with a multiphase vaccination and its application to Italian COVID-19 data. Mathematics and Computers in Simulation, 228:124–146, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003136.

Devroye:2025:AOA

[1942] Luc Devroye and Dimitrios Los. An asymptotically optimal algorithm for generating bin cardinalities. Mathematics and Computers in Simulation, 228:147–155, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400346X.

Zwerger:2025:ILM

[1943] Tanja Zwerger and Paolo Mercorelli.

Inner loop model predictive control and outer loop PI reference governor for PMSMs with input and state saturation for torque control. Mathematics and Computers in Simulation, 228:178–201, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003082.

Zhao:2025:QMV

[1944] Longbin Zhao, Pengde Wang, and Qiongqi Fan. A quadrature method for Volterra integral equations with highly oscillatory Bessel kernel. Mathematics and Computers in Simulation, 228:202–210, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003483.

DiCrescenzo:2025:CAE

[1945] A. Di Crescenzo, A. Gómez-Corral, and D. Taipe. A computational approach to extreme values and related hitting probabilities in level-dependent quasi-birth–death processes. Mathematics and Computers in Simulation, 228:211–224, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003215.

Hoang:2025:GDN

[1946] Manh Tuan Hoang and Hoai Thu Pham. Global dynamics and numerical simulation of a modified epidemiological model for viral marketing on social networks. Mathematics and Computers in Simulation, 228: 225–244, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003264.

Wang:2025:ESD

[1947] Yurun Wang and Fei Liu. An exponential spectral deferred correction method for multidimensional parabolic problems. Mathematics and Computers in Simulation, 228: 245–262, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003495.

Kotb:2025:ESS

[1948] M. S. Kotb and M. A. Al Omari.

Estimation of the stress–strength reliability for the exponential-Rayleigh distribution. Mathematics and Computers in Simulation, 228:263–273, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003604.

Roy:2025:BPP

[1949] Subarna Roy and Pankaj Kumar Tiwari. Bistability in a predator–prey model characterized by the Crowley– Martin functional response: Effects of fear, hunting cooperation, additional foods and nonlinear harvesting. Mathematics and Computers in Simulation, 228:274–297, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003471.

Zhang:2025:STG

[1950] Fan Zhang, Po-Wei Li, Yan Gu, and Chia-Ming Fan. A space-time generalized finite difference scheme for long wave propagation based on high-order Korteweg–de Vries type equations. Mathematics and Computers in Simulation, 228:298– 312, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003677.

Mandal:2025:BAE

[1951] Sayan Mandal, Sudip Samanta, Pankaj Kumar Tiwari, and Ranjit Kumar Upadhyay. Bifurcation analysis and exploration of noise-induced transitions of a food chain model with Allee effect. Mathematics and Computers in Simulation, 228:313– 338, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003707.

Jiao:2025:RDR

[1952] Xubin Jiao, Li Liu, and Xiao Yu. Rich dynamics of a reaction–diffusion Filippov Leslie–Gower predator–prey model with time delay and discontinuous harvesting. Mathematics and Computers in Simulation, 228:339– 361, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003768.

Wu:2025:GWP

[1953] Peng Wu. Global well-posedness and dynamics of spatial diffusion HIV model with CTLs response and chemotaxis. Mathematics and Computers in Simulation, 228:402– 417, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003744.

Wu:2025:TLE

[1954] Longyuan Wu, Xufeng Xiao, and Shuying Zhai. Two linear energy stable lumped mass finite element schemes for the viscous Cahn–Hilliard equation on curved surfaces in 3D. Mathematics and Computers in Simulation, 228:418–430, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003756.

Midhun:2025:SSP

[1955] T. A. Midhun and K. Murugesan.

A study of stochastically perturbed epidemic model of HPV infection and cervical cancer in Indian female population. Mathematics and Computers in Simulation, 228:431– 465, February 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542400363X.

Liu:2025:NSV

[1956] Lei Liu, Xuan-Xuan Zhou, Xi-Yang Xie, and Wen-Rong Sun. Numerical study of vector solitons with the oscillatory phase backgrounds in the integrable coupled nonlinear Schrödinger equations. Mathematics and Computers in Simulation, 228: 466–484, February 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003641.

Anonymous:2025:EBc

[1957] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 229:ii, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004610.

Beneit-Barajas:2025:ANU

[1958] Adrián Beneit-Barajas, Patricia Penades-Huesca, and Jaime Martínez-Turégano. Analysis of non-uniform grid-forming control techniques for the HVDC connection of renewable energy. Mathematics and Computers in Simulation, 229:1–14, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003690.

Anonymous:2025:PMa

[1959] Anonymous. Pages 1–898 (March 2025). Mathematics and Computers in Simulation, 229:1–898, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Navarro-Martinez:2025:CSW

[1960] Gala Navarro-Martínez, Jaime Martínez-Turégano, and Ramon Blasco-Gimenez. Control of a 15MW off-shore wind turbine for black-start operation. Mathematics and Computers in Simulation, 229:15–31, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003689.

Yao:2025:DLG

[1961] Yong Yao. Dynamics of a Leslie– Gower type predator–prey system with herd behavior and constant harvesting in prey. Mathematics and Computers in Simulation, 229:32– 49, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003811.

Janani:2025:EFS

[1962] K. Janani, S. S. Mohanrasu, Ardak Kashkynbayev, and R. Rakkiyappan. Ensemble feature selection via CoCoSo method extended to interval-valued intuitionistic fuzzy environment. Mathematics and Computers in Simulation, 229:50–77, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003781.

Liu:2025:SMS

[1963] Minggang Liu, Ning Xu, Ben Niu, and Naif D. Alotaibi. Sliding-mode surface-based fixed-time adaptive critic tracking control for zero-sum game of switched nonlinear systems. Mathematics and Computers in Simulation, 229:78–95, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400380X.

UlKarim:2025:EAG

[1964] Md Aktar Ul Karim, Ruqaiya Altaf Shaikh, and Amiya Ranjan Bhowmick. Efficient approximation of global population dynamic models through statistical inference using local data. Mathematics and Computers in Simulation, 229:96–128, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003793.

Li:2025:MPP

[1965] Jiang Li, Xianning Liu, and Yangjiang Wei. Modelling the prudent predation in predator–prey interactions. Mathematics and Computers in Simulation, 229:129–150, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003860.

Dao:2025:AOC

[1966] Phuong Nam Dao, Quang Phat Nguyen, and Manh Hung Vu. Adaptive optimal coordination control of perturbed Bilateral Teleoperators with variable time delays using Actor–Critic Reinforcement Learning algorithm. Mathematics and Computers in Simulation, 229:151– 175, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003628.

Rujivan:2025:ACC

[1967] Sanae Rujivan, Nopporn Thamrongrat, Parun Juntanon, and Boualem Djehiche. Analytical computation of conditional moments in the extended Cox– Ingersoll–Ross process with regime switching: Hybrid PDE system solutions with financial applications. Mathematics and Computers in Simulation, 229:176–202, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003872.

Raina:2025:AEA

[1968] Aayushman Raina, Srinivasan Natesan, and Şuayip Toprakseven. Anisotropic error analysis of weak Galerkin finite element method for singularly perturbed biharmonic problems. Mathematics and Computers in Simulation, 229:203–221, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003720.

Campagna:2025:HSM

[1969] Rosanna Campagna, Mariantonia Cotronei, and Domenico Fazzino. A Hermite spline model for data regression. Mathematics and Computers in Simulation, 229:222–234, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424003665.

Kumar:2025:DAH

[1970] Rakesh Kumar and Sudarshan Dhua.

Dynamic analysis of Hashimoto’s Thyroiditis bio-mathematical model using artificial neural network. Mathematics and Computers in Simulation, 229:235–245, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003902.

Settati:2025:SSE

[1971] A. Settati, T. Caraballo, A. Lahrouz, I. Bouzalmat, and A. Assadouq. Stochastic SIR epidemic model dynamics on scale-free networks. Mathematics and Computers in Simulation, 229:246–259, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003823.

Ardeuan:2025:DAT

[1972] Andreea-Maria Ardeuan, Mihaela Neamţu, and Adriana Loredana Tănasie. A dynamic analysis of a tourism-based socioecological system. Mathematics and Computers in Simulation, 229:260– 272, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003033.

Li:2025:TDN

[1973] Lanlan Li, Zhuojia Fu, Ming Qin, Shuainan Liu, Weihong Zeng, and Xiaoting Liu. Three-dimensional numerical wave tank containing submerged breakwaters based on the localized method of fundamental solutions. Mathematics and Computers in Simulation, 229:273– 287, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003951.

Zhang:2025:FPE

[1974] Qian Zhang, Zhongbo Hu, Nan Hong, and Qinghua Su. A fixed point evolution algorithm based on expanded Aitken rapid iteration method for global numeric optimization. Mathematics and Computers in Simulation, 229:288–303, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003379.

Chen:2025:CGS

[1975] Shi-Shun Chen and Xiao-Yang Li.

Comparison of global sensitivity analysis methods for a fire spread model with a segmented characteristic. Mathematics and Computers in Simulation, 229:304–318, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004014.

Wang:2025:ASD

[1976] Jingjing Wang, Yunfeng Jia, and Majun Shi. Analysis and simulation on dynamics of a pest-natural enemy model with additional food, pest-taxis and degeneracy. Mathematics and Computers in Simulation, 229:319–339, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003884.

Ducharne:2025:AED

[1977] B. Ducharne and G. Sebald. Analytical expressions of the dynamic magnetic power loss under alternating or rotating magnetic field. Mathematics and Computers in Simulation, 229:340–349, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003975.

Belenguer:2025:STE

[1978] E. Belenguer, J. Segarra-Tamarit, E. Pérez, and R. Vidal-Albalate. Short-term electricity price forecasting through demand and renewable generation prediction. Mathematics and Computers in Simulation, 229:350–361, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003938.

Wang:2025:GBH

[1979] Jiaxin Wang, Kun Wang, Zhen-Guo Yan, Xiaofeng He, and Tiegang Liu. GPGPU-based heterogeneous parallel implementation of direct discontinuous Galerkin methods. Mathematics and Computers in Simulation, 229:362–391, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003896.

Seck:2025:ASM

[1980] Rama Seck, Diene Ngom, Benjamin Ivorra, and Angel M. Ramos. An age-structured mathematical model for studying Malaria transmission dynamics: Applications to some areas of Senegal. Mathematics and Computers in Simulation, 229:392– 408, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003963.

Dujardin:2025:NSV

[1981] Guillaume Dujardin, Ingrid Lacroix-Violet, and Anthony Nahas. A numerical study of vortex nucleation in 2D rotating Bose–Einstein condensates. Mathematics and Computers in Simulation, 229:409– 434, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003847.

Mohapatra:2025:TFI

[1982] Dhabaleswar Mohapatra and S. Chakraverty.

Type-2 fuzzy initial value problems under granular differentiability. Mathematics and Computers in Simulation, 229:435–447, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424003914.

Chang:2025:MTS

[1983] Chih-Wen Chang, Sania Qureshi, Ioannis K. Argyros, Francisco I. Chicharro, and Amanullah Soomro. A modified two-step optimal iterative method for solving nonlinear models in one and higher dimensions. Mathematics and Computers in Simulation, 229:448–467, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400377X.

Wang:2025:SBO

[1984] Libin Wang and Lixia Liu. Some bivariate options pricing in a regimeswitching stochastic volatility jumpdiffusion model with stochastic intensity, stochastic interest and dependent jump. Mathematics and Computers in Simulation, 229:468– 490, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004026.

Trong:2025:TPT

[1985] Dang Duc Trong, Dinh Nguyen Duy Hai, Nguyen Dang Minh, and Nguyen Nhu Lan. A two-parameter Tikhonov regularization for a fractional sideways problem with two interior temperature measurements. Mathematics and Computers in Simulation, 229:491–511, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004038.

Sgheri:2025:DEP

[1986] Luca Sgheri, Cristina Sgattoni, and Chiara Zugarini. Determination of emissivity profiles using a Bayesian data-driven approach. Mathematics and Computers in Simulation, 229:512–524, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004051.

Cammarasana:2025:ACH

[1987] Simone Cammarasana and Giuseppe Patané. Analysis and comparison of high-performance computing solvers for minimisation problems in signal processing. Mathematics and Computers in Simulation, 229:525– 538, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003926.

Mondal:2025:CAP

[1988] Santana Mondal and Subhas Khajanchi. Can adaptive prey refuge facilitate species coexistence in Bazykin’s prey–predator model? Mathematics and Computers in Simulation, 229:539–552, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004105.

Kumari:2025:ENA

[1989] Parvin Kumari, Satpal Singh, and Devendra Kumar. An effective numerical approach for solving a system of singularly perturbed differential– difference equations in biology and physiology. Mathematics and Computers in Simulation, 229:553–573, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424004002.

Meng:2025:FPE

[1990] Qingyan Meng, Yejuan Wang, Peter E.

Kloeden, and Xiaoying Han. Fokker– Planck equation and Feynman–Kac formula for multidimensional stochastic dynamical systems with Lévy noises and time-dependent coefficients. Mathematics and Computers in Simulation, 229:574–593, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400404X.

Hoai:2025:NPG

[1991] Pham Thi Hoai. A new proximal gradient method for solving mixed variational inequality problems with a novel explicit stepsize and applications. Mathematics and Computers in Simulation, 229:594– 610, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003999.

Kaur:2025:TPT

[1992] Supinder Kaur, Kalpana Dahiya, and Anuj Sharma. Two-phase time minimization transportation problem with the restricted flow. Mathematics and Computers in Simulation, 229:611– 635, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424003859.

Abia:2025:CAA

[1993] Luis M. Abia, Óscar Angulo, Juan Carlos López-Marcos, and Miguel Ángel López-Marcos. A convergence analysis for the approximation to the solution of an age-structured population model with infinite lifespan. Mathematics and Computers in Simulation, 229:636–651, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424003987.

Gupta:2025:SNC

[1994] R. P. Gupta, Shristi Tiwari, and Arun Kumar. The study of non-constant steady states and pattern formation for an interacting population model in a spatial environment. Mathematics and Computers in Simulation, 229:652–672, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004129.

Kamrani:2025:APP

[1995] Minoo Kamrani and Erika Hausenblas. An adaptive positive preserving numerical scheme based on splitting method for the solution of the CIR model. Mathematics and Computers in Simulation, 229:673– 689, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004117.

Nouri:2025:PBS

[1996] Bahareh Nouri, Imre Juhász, and Jamshid Saeidian. A polynomial basis with a shape parameter for curve and surface modeling. Mathematics and Computers in Simulation, 229:690–705, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004270.

Jia:2025:CDS

[1997] Chunping Jia, Xia Wang, and Yuming Chen. Complex dynamics of an SIHR epidemic model with variable hospitalization rate depending on unoccupied hospital beds. Mathematics and Computers in Simulation, 229:706–724, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004130.

ElMokhtari:2025:GAA

[1998] F. El Mokhtari, M. Lamnii, and D. Barrera. A genetic algorithm approach based on spline quasiinterpolation for solving Fredholm integral equations. Mathematics and Computers in Simulation, 229:725– 742, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004312.

He:2025:LCC

[1999] Dongping He and Jaume Llibre. Limit cycles in a class of planar discontinuous piecewise quadratic differential systems with a non-regular line of discontinuity (I). Mathematics and Computers in Simulation, 229: 743–757, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004075.

Rajpal:2025:EUI

[2000] Akanksha Rajpal, Sumit Kaur Bhatia, Shashank Goel, Sanyam Tyagi, and Praveen Kumar. Epidemic and unemployment interplay through bi-level multi delayed mathematical model. Mathematics and Computers in Simulation, 229:758– 788, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004269.

Mokni:2025:NDD

[2001] Karima Mokni, Halima Ben Ali, Bapan Ghosh, and Mohamed Ch-Chaoui. Nonlinear dynamics of a Darwinian Ricker system with strong Allee effect and immigration. Mathematics and Computers in Simulation, 229:789–813, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004087.

Irum:2025:CIT

[2002] Saba Irum, Naif Almakayeel, and Wejdan Deebani. Computational insights into tumor invasion dynamics: a finite element approach. Mathematics and Computers in Simulation, 229:814–829, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004257.

Alidrissi:2025:CDD

[2003] Youssef Alidrissi, Houda El Ouadoud, Serge Pierfederici, Matthieu Urbain, Radouane Ouladsine, and Mohamed Bakhouya. A cooperative distributed droop gains adjustment in DC microgrids. Mathematics and Computers in Simulation, 229:830– 841, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542400394X.

Lee:2025:COC

[2004] Taeyong Lee, Hee-Dae Kwon, and Jeehyun Lee. Constrained optimal control problem of oncolytic viruses in cancer treatment. Mathematics and Computers in Simulation, 229: 842–854, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004099.

Settembre:2025:LCC

[2005] Gaetano Settembre, Nicolò Taggio, Nicoletta Del Buono, Flavia Esposito, Paola Di Lauro, and Antonello Aiello. A land cover change framework analyzing wildfire-affected areas in bitemporal PRISMA hyperspectral images. Mathematics and Computers in Simulation, 229:855– 866, March 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424004324.

Chen:2025:UAD

[2006] Bingzhen Chen and Wenjuan Zhai. Unified algorithms for distributed regularized linear regression model. Mathematics and Computers in Simulation, 229:867–884, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004063.

Peri:2025:MOO

[2007] Daniele Peri. Multi-objective optimization of the appendages of a sailing yacht using the Normal Boundary Intersection method. Mathematics and Computers in Simulation, 229:885–895, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004397.

Anonymous:2025:NIc

[2008] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 229:896, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004646.

Anonymous:2025:ICEc

[2009] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 229:897, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424004658.

Anonymous:2025:EBd

[2010] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 230:ii, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424005019.

Li:2025:DAD

[2011] Ruqi Li, Yurong Song, Min Li, Hongbo Qu, and Guo-Ping Jiang. Dynamic analysis and data-driven inference of a fractional-order SEIHDR epidemic model with variable parameters. Mathematics and Computers in Simulation, 230:1–19, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004415.

Anonymous:2025:PAa

[2012] Anonymous. Pages 1–560 (April 2025).

Mathematics and Computers in Simulation, 230:1–560, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Sarumi:2025:ESO

[2013] Ibrahim O. Sarumi, Khaled M. Furati, and Abdul Q. M. Khaliq. Efficient second-order accurate exponential time differencing for time-fractional advection–diffusion–reaction equations with variable coefficients. Mathematics and Computers in Simulation, 230:20–38, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004439.

Shweta:2025:LTS

[2014] Shweta, Gourav Arora, and Rajesh Kumar. Large time solution for collisional breakage model: Laplace transformation based accelerated homotopy perturbation method. Mathematics and Computers in Simulation, 230:39–52, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004427.

Sharma:2025:IAW

[2015] Deepika Sharma and Randheer Singh. Interaction of an acceleration wave with a characteristic shock in twophase real modified Chaplygin model containing a variable source term. Mathematics and Computers in Simulation, 230:53–67, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004221.

Wu:2025:IFG

[2016] Wen-Ze Wu, Jie Xu, Wanli Xie, and Tao Zhang. An innovative fractional grey system model and its application. Mathematics and Computers in Simulation, 230:68–79, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004440.

DiCola:2025:INN

[2017] Vincenzo Schiano Di Cola, Vittorio Bauduin, Marco Berardi, Filippo Notarnicola, and Salvatore Cuomo. Investigating neural networks with groundwater flow equation loss. Mathematics and Computers in Simulation, 230:80–93, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004373.

Liu:2025:FTS

[2018] Xianghu Liu, Yanfang Li, and Guangjun Xu. Finite-time synchronization analysis for the generalized Caputo fractional spatio-temporal neural networks. Mathematics and Computers in Simulation, 230:94– 110, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004476.

Bruni:2025:PGC

[2019] Vittoria Bruni, Giulia Dominijanni, Domenico Vitulano, and Giuliana Ramella. A perception-guided CNN for grape bunch detection. Mathematics and Computers in Simulation, 230:111–130, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004464.

Gao:2025:LTB

[2020] Jianping Gao and Wenyan Lian. Long time behavior of a Lotka– Volterra competition system with two dynamical resources and densitydependent motility. Mathematics and Computers in Simulation, 230: 131–148, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542400449X.

Chakib:2025:NRS

[2021] Abdelkrim Chakib, Ibrahim Khalil, and Azeddine Sadik. On numerical resolution of shape optimization bi-Laplacian eigenvalue problems. Mathematics and Computers in Simulation, 230:149–164, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004488.

Ruiz-Castro:2025:AMC

[2022] Juan Eloy Ruiz-Castro and Hugo Alaín Zapata-Ceballos. Algorithmic modelling of a complex redundant multistate system subject to multiple events, preventive maintenance, loss of units and a multiple vacation policy through a MMAP. Mathematics and Computers in Simulation, 230:165–192, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004452.

Bee:2025:PDM

[2023] Marco Bee. A parsimonious dynamic mixture for heavy-tailed distributions. Mathematics and Computers in Simulation, 230:193–206, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847542400452X.

Hao:2025:CBT

[2024] Lupeng Hao, Weihang Peng, Junhua Liu, Wei Zhang, Yuan Li, and Kaixuan Qin. Competition-based two-stage evolutionary algorithm for constrained multi-objective optimization. Mathematics and Computers in Simulation, 230:207–226, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004506.

Zhang:2025:NSE

[2025] Lihong Zhang, Keke Lu, and Bashir Ahmad. Numerical simulation and error analysis for a novel fractal– fractional reaction diffusion model with weighted reaction. Mathematics and Computers in Simulation, 230:227–240, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004543.

Zong:2025:MLL

[2026] Tiancheng Zong, Junhong Li, and Guoping Lu. Maximum likelihood LM identification based on particle filtering for scarce measurement-data MIMO Hammerstein Box-Jenkins systems. Mathematics and Computers in Simulation, 230:241–255, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424004531.

Nieto-Hernandez:2025:MCI

[2027] David Nieto-Hernández and César-Fernando Méndez-Barrios. Modeling and control of an interactive tilt-rotor MAV. A robust delaybased strategy. Mathematics and Computers in Simulation, 230:256– 277, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004518.

Aguinaga:2025:ASE

[2028] Ismael Chirino Aguinaga, Nicolas Patin, Patrice Gomez, Vincent Lanfranchi, and Jeanne-Marie Dalbavie. Algorithm for sequences exploration and optimization of a Multi-Active-Bridge. Mathematics and Computers in Simulation, 230:278– 288, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542400435X.

Li:2025:NAA

[2029] Mengna Li and Zhanwen Yang. Numerical analysis of an age-structured model for HIV viral dynamics with latently infected T cells based on collocation methods. Mathematics and Computers in Simulation, 230: 289–305, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424003835.

Fumagalli:2025:NVA

[2030] Alessio Fumagalli and Francesco Saverio Patacchini. Numerical validation of an adaptive model for the determination of nonlinear-flow regions in highly heterogeneous porous media. Mathematics and Computers in Simulation, 230:306–322, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004233.

Das:2025:RSM

[2031] Saduri Das, Prashant K. Srivastava, and Pankaj Biswas. The role of social media in a tuberculosis compartmental model: Exploring Hopf-bifurcation and nonlinear oscillations. Mathematics and Computers in Simulation, 230:323–349, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004567.

Printechapat:2025:CMP

[2032] Tanes Printechapat, Pakorn Aiewsakun, and Tipaluck Krityakierne. Caged Markov process — a continuoustime framework for modeling a constrained Markov process within a freely-evolving Markov process. Mathematics and Computers in Simulation, 230:350–369, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004555.

DellAccio:2025:IME

[2033] Francesco Dell’Accio, Allal Guessab, and Federico Nudo. Improved methods for the enrichment and analysis of the simplicial vector-valued linear finite elements. Mathematics and Computers in Simulation, 230:370–385, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000260.

Sushko:2025:RCD

[2034] Iryna Sushko and Fabio Tramontana. Regular and chaotic dynamics in a 2D discontinuous financial market model with heterogeneous traders. Mathematics and Computers in Simulation, 230:386–399, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542400034X.

Grammont:2025:IOB

[2035] Laurence Grammont and Paulo B. Vasconcelos. Influence of the order between discretization and regularization in solving ill-posed problems. Mathematics and Computers in Simulation, 230:400–412, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424000594.

Allouch:2025:CSQ

[2036] C. Allouch, D. Barrera, A. Saou, and M. Tahrichi. Cubic spline quasiinterpolation operator to numerically solve integro-differential equations with weakly singular kernels. Mathematics and Computers in Simulation, 230:413–422, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424000934.

Jelti:2025:RBS

[2037] S. Jelti, A. Serghini, and I. Elmahi.

Resolution of the bidimensional shallow water equations with horizontal temperature gradients by a well Balanced Roe scheme. Mathematics and Computers in Simulation, 230:423–437, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004245.

Boda:2025:NEN

[2038] Lívia Boda and István Faragó. New efficient numerical methods for some systems of linear ordinary differential equations. Mathematics and Computers in Simulation, 230:438– 455, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004282.

Jin:2025:FAF

[2039] Jie Jin, Xiaoyang Lei, Chaoyang Chen, and Zhijing Li. A fuzzy activation function based zeroing neural network for dynamic Arnold map image cryptography. Mathematics and Computers in Simulation, 230: 456–469, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424004294.

Saidi:2025:TMB

[2040] S. Saidi, A. Abbadi, and M. Tahrichi. Two methods based on spline quasiinterpolants to estimate volumes enclosed by parametric surfaces. Mathematics and Computers in Simulation, 230:470–483, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004300.

Massoun:2025:ASS

[2041] Y. Massoun, A. K. Alomari, and C. Cesarano. Analytic solution for SIR epidemic model with multi-parameter fractional derivative. Mathematics and Computers in Simulation, 230: 484–492, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424004336.

Gonzalez:2025:NSL

[2042] P. González, A. Kouibia, B. Mustafa, and M. Pasadas. Numerical solution of a linear Volterra integrodifferential problem. Mathematics and Computers in Simulation, 230: 493–500, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424004348.

intHout:2025:NNA

[2043] Karel J. in ’t Hout. A note on the numerical approximation of Greeks for American-style options. Mathematics and Computers in Simulation, 230:501–516, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004361.

Singh:2025:DMF

[2044] Sarbjit Singh, Kulwinder Singh Parmar, and Jatinder Kumar. Development of multi-forecasting model using Monte Carlo simulation coupled with wavelet denoising-ARIMA model. Mathematics and Computers in Simulation, 230:517–540, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475424004385.

Valentino:2025:SST

[2045] Carmine Valentino, Giovanni Pagano, Dajana Conte, Beatrice Paternoster, Francesco Colace, and Mario Casillo. Step-by-step time discrete Physics-Informed Neural Networks with application to a sustainability PDE model. Mathematics and Computers in Simulation, 230:541– 558, April 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004403.

Anonymous:2025:NId

[2046] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 230:559, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004968.

Anonymous:2025:ICEd

[2047] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 230:560, April 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S037847542400497X.

Anonymous:2025:EBe

[2048] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 231:ii, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000242.

Ren:2025:TDD

[2049] Huarong Ren and Rui Xu. Transmission dynamics of dengue with asymptomatic infection and a case study in Bangladesh. Mathematics and Computers in Simulation, 231: 1–18, May 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424004749.

Anonymous:2025:PMb

[2050] Anonymous. Pages 1–354 (May 2025). Mathematics and Computers in Simulation, 231:1–354, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

An:2025:OBS

[2051] Yaxin An and Yongchao Liu. Observerbased self-triggered adaptive decentralized control for uncertain interconnected nonlinear systems. Mathematics and Computers in Simulation, 231:19–31, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004671.

Shen:2025:SEF

[2052] Hao Shen, Zhen Mei, Lei Su, Jing Wang, and Kaibo Shi. H_{∞} state estimation for fuzzy affine systems with PDT switching-based DoS attacks. Mathematics and Computers in Simulation, 231:32–45, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004701.

Kudenatti:2025:LST

[2053] Ramesh B. Kudenatti, Bharathi M.

C., and Noor-E-Misbah. Linear stability on transpiration effect of selfsimilar boundary layer flow for non-Newtonian fluids over a moving wedge. Mathematics and Computers in Simulation, 231:46–61, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004579.

Selicato:2025:TDA

[2054] Laura Selicato, Alessandro Pagano, Flavia Esposito, and Matteo Icardi. Topological data analysis for resilience assessment of water distribution networks. Mathematics and Computers in Simulation, 231: 62–70, May 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424004725.

Filelis-Papadopoulos:2025:AME

[2055] Christos K. Filelis-Papadopoulos, John P. Morrison, and Philip O’Reilly. Adaptive multilayer extreme learning machines. Mathematics and Computers in Simulation, 231:71– 98, May 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424004750.

Wang:2025:SFW

[2056] Shuo Wang, Jie Ma, and Ning Du. A stabilizer-free weak Galerkin finite element method for an optimal control problem of a time fractional diffusion equation. Mathematics and Computers in Simulation, 231:99–118, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004683.

Remogna:2025:SEI

[2057] Sara Remogna, María José Ibáñez, Paola Lamberti, and Driss Sbibih. Second Edition of the International Conference on Mathematical and Computational Modelling, Approximation and Simulation: New trends, recent developments and applications in environment and natural resources. Mathematics and Computers in Simulation, 231:119, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004713.

Ramos:2025:MOM

[2058] Sérgio Ramos, Luís A. C. Roque, António Gomes, João Soares, José Calvo Rolle, and Zita Vale. Multiobjective model for residential energy management in context of individual self-consumption. Mathematics and Computers in Simulation, 231:120–127, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400466X.

Chou:2025:SFC

[2059] Lot-Kei Chou, Wei Qu, Yuan-Yuan Huang, and Siu-Long Lei. A study on fractional centered difference scheme for high-dimensional integral fractional Laplacian operator with ωcirculant preconditioner. Mathematics and Computers in Simulation, 231:128–143, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004737.

Zhang:2025:MSE

[2060] Chao Zhang, Yi Yang, and Wei Chen. Multi-strategy ensemble wind driven optimization algorithm for robot path planning. Mathematics and Computers in Simulation, 231: 144–159, May 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004798.

Devi:2025:HHB

[2061] Arpita Devi and Praveen Kumar Gupta. Hopf-Hopf bifurcation and hysteresis in a COVID-19 transmission model implementing vaccination induced recovery and a modified Holling type-III treatment response. Mathematics and Computers in Simulation, 231:160–184, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004816.

Wang:2025:SEP

[2062] Junjie Wang. Some energy-preserving schemes for fractional Hamiltonian system with fractional Laplacian. Mathematics and Computers in Simulation, 231:185–208, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004762.

Yao:2025:SPL

[2063] Zichen Yao, Zhanwen Yang, and Lixin Cheng. On the stability preserving of L1 scheme to nonlinear time-fractional Schrödinger delay equations. Mathematics and Computers in Simulation, 231:209–220, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004695.

Zhu:2025:EPC

[2064] Xinhe Zhu, Yuanyou Shi, and Yongmin Zhong. An EKF prediction of COVID-19 propagation under vaccinations and viral variants. Mathematics and Computers in Simulation, 231:221–238, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004841.

Berrone:2025:EPM

[2065] Stefano Berrone and Fabio Vicini.

Effective polygonal mesh generation and refinement for VEM. Mathematics and Computers in Simulation, 231:239–258, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004786.

Thakur:2025:MQM

[2066] Sonali Thakur and Anamika Jain.

Markovian queueing model under Bernoulli vacation and servers’ malfunctioning: Metaheuristic optimization technique. Mathematics and Computers in Simulation, 231:259– 275, May 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004890.

El-Khatib:2025:DOP

[2067] Youssef El-Khatib, Zororo S. Makumbe, and Josep Vives. Decomposition of the option pricing formula for infinite activity jump-diffusion stochastic volatility models. Mathematics and Computers in Simulation, 231: 276–293, May 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004828.

Akay:2025:SEM

[2068] Rustu Akay and Mustafa Yusuf Yildirim. SBA*: an efficient method for 3D path planning of unmanned vehicles. Mathematics and Computers in Simulation, 231: 294–317, May 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424004889.

Uzunca:2025:LIM

[2069] Murat Uzunca and Bülent Karasözen. Linearly implicit methods for the nonlinear Klein–Gordon equation. Mathematics and Computers in Simulation, 231:318–330, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004919.

Wang:2025:MMO

[2070] Juan Wang, Feiyan Zhao, Jichao Wang, and An Li. Modeling and multiobjective optimal state-dependent control of a continuous double-bioreactor in series fermentation. Mathematics and Computers in Simulation, 231:331–351, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424004804.

Anonymous:2025:NIe

[2071] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 231:352, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000278.

Anonymous:2025:ICEe

[2072] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 231:353, May 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847542500028X.

Anonymous:2025:EBf

[2073] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 232:ii, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000618.

Ma:2025:ESM

[2074] Suna Ma and Huiyuan Li. Efficient spectral methods for the fourth-order elliptic eigenvalue problems. Mathematics and Computers in Simulation, 232:1–16, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004774.

Anonymous:2025:PJb

[2075] Anonymous. Pages 1–506 (June 2025).

Mathematics and Computers in Simulation, 232:1–506, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Dolui:2025:TAN

[2076] Soumini Dolui, Bivas Bhaumik, and Soumen De. Thermal analysis of nanolayer interfaces and nanoparticle shape reactivity in EMHD micromagnetorotational ternary nanofluid flow within deceased bifurcated artery. Mathematics and Computers in Simulation, 232:17–49, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004853.

Sagar:2025:ECE

[2077] B. Sagar and S. Saha Ray. Exploring computationally efficient stable numerical techniques for fractional Keller– Segel system modeling chemotaxis. Mathematics and Computers in Simulation, 232:50–74, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400483X.

Pal:2025:MWE

[2078] Saheb Pal and Nikhil Pal. Modelling the wind effect in predator– prey interactions. Mathematics and Computers in Simulation, 232:75– 90, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004865.

Wang:2025:CMS

[2079] Wentao Wang, Jihui Wu, and Wei Chen. The characteristics method to study global exponential stability of delayed inertial neural networks. Mathematics and Computers in Simulation, 232:91–101, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004932.

Wu:2025:ARD

[2080] Shipei Wu and Shaojun Lan. Analysis of repairable discrete-time queueing systems with negative customers, disasters, balking customers and interruptible working vacations under Bernoulli schedule. Mathematics and Computers in Simulation, 232:102– 122, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004877.

Hao:2025:BBT

[2081] Yating Hao and Ben Gao. Bilinear Bäcklund transformation and soliton solutions to a (3 + 1)-dimensional coupled nonlinear Schrödinger equation with variable coefficients in optical fibers. Mathematics and Computers in Simulation, 232:123– 139, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542400507X.

Li:2025:AMR

[2082] Hesong Li, Zhaoting Li, Hongbo Zhang, and Yi Wang. An adaptive mesh refinement method considering control errors for pseudospectral discretization. Mathematics and Computers in Simulation, 232:140– 159, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000059.

Hajri:2025:DWH

[2083] Youssra Hajri and Saida Amine. Dynamics of a within-host HIV/SARS-CoV-2 co-infection model with two intracellular delays. Mathematics and Computers in Simulation, 232: 160–191, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424005111.

Carfora:2025:TPI

[2084] M. F. Carfora, F. Iovanna, and I. Torcicollo. Turing patterns in an intraguild predator–prey model. Mathematics and Computers in Simulation, 232:192–210, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424005068.

Liu:2025:SHO

[2085] Chein-Shan Liu, Yung-Wei Chen, Jian-Hung Shen, and Yen-Shen Chang. Solving higher-order nonlocal boundary value problems with high precision by the fixed quasi Newton methods. Mathematics and Computers in Simulation, 232:211– 226, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475424005081.

Zhao:2025:DCH

[2086] Zhihong Zhao and Yuwei Shen. Dynamic complexity of Holling–Tanner predator–prey system with predator cannibalism. Mathematics and Computers in Simulation, 232:227– 244, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424005093.

Kumar:2025:NFO

[2087] Pushpendra Kumar, Tae H. Lee, and Vedat Suat Erturk. A novel fractionalorder neutral-type two-delayed neural network: Stability, bifurcation, and numerical solution. Mathematics and Computers in Simulation, 232:245–260, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000011.

Alla:2025:SDR

[2088] Alessandro Alla, Marco Berardi, and Luca Saluzzi. State Dependent Riccati for dynamic boundary control to optimize irrigation in Richards’ equation framework. Mathematics and Computers in Simulation, 232: 261–275, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475424004920.

Zhang:2025:FPT

[2089] Hai Zhang, Panqing Gao, Renyu Ye, Ivanka Stamova, and Jinde Cao. Fixed/predefined time synchronization of fractional quaternion delayed neural networks with disturbances. Mathematics and Computers in Simulation, 232:276–294, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000102.

Wu:2025:RMS

[2090] Hanghang Wu and Hongqi Yang. Regularization methods for solving a time-fractional diffusion inverse source problem. Mathematics and Computers in Simulation, 232:295– 310, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000023.

Bharti:2025:MPC

[2091] Jyotsna Kumari Bharti, K. Murugesan, and P. Balasubramaniam. Matrix projective combination synchronization of time-delayed chaotic system and its application in image encryption. Mathematics and Computers in Simulation, 232:311–345, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000047.

Vidal:2025:WTI

[2092] Marc Vidal and Ana M. Aguilera. Wavelet thresholding on independent subspace factorizations of spatially indexed wide functional data for robust estimation of cortical activity. Mathematics and Computers in Simulation, 232:346–361, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000126.

S:2025:DMB

[2093] Kavitha S., Kendra N., Satheeshkumar J., Amudha T., and Balachandran Manavalan. Decision making based ensemble feature selection approach through a new score function in q-rung orthopair hesitant fuzzy environment. Mathematics and Computers in Simulation, 232:362–381, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475424004907.

Zhang:2025:IEA

[2094] Yong Zhang. Image encryption algorithm based on butterfly module and chaos. Mathematics and Computers in Simulation, 232:382– 407, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000114.

Ounamane:2025:FTE

[2095] Said Ounamane, Lakhlifa Sadek, Bouchra Abouzaid, and El Mostafa Sadek. Fractional truncated exponential method for linear fractional optimal control problems. Mathematics and Computers in Simulation, 232:408–426, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000096.

Hu:2025:GDL

[2096] Sengen Hu and Liangqiang Zhou.

Global dynamics in the lateral oscillation model of pedestrian walking on a vibrating surface. Mathematics and Computers in Simulation, 232:427–453, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542400510X.

Zhang:2025:LSI

[2097] Hongyang Zhang, Shuting Wang, Yuanlong Xie, Hu Li, and Shiqi Zheng. Leadership succession inspired adaptive operator selection mechanism for multi-objective optimization. Mathematics and Computers in Simulation, 232:454–474, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000072.

Shanbhag:2025:GEC

[2098] Soham Shanbhag and Dong Eui Chang. Globally exponentially convergent observer for systems evolving on matrix Lie groups. Mathematics and Computers in Simulation, 232:475– 482, June 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000138.

Han:2025:IST

[2099] Peng-Fei Han, Ru-Suo Ye, and Yi Zhang. Inverse scattering transform for the coupled Lakshmanan– Porsezian–Daniel equations with nonzero boundary conditions in optical fiber communications. Mathematics and Computers in Simulation, 232:483–503, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000084.

Anonymous:2025:NIf

[2100] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 232:504, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000643.

Anonymous:2025:ICEf

[2101] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 232:505, June 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475425000655.

Anonymous:2025:EBg

[2102] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 233:ii, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542500103X.

Chen:2025:USL

[2103] Xin Chen, Zeyu Zhang, and Peiqi Huang. Uncertain stochastic linear quadratic control subject to forward and backward multi-stage systems. Mathematics and Computers in Simulation, 233:1–20, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000151.

Anonymous:2025:PJc

[2104] Anonymous. Pages 1–568 (July 2025). Mathematics and Computers in Simulation, 233:1–568, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Cao:2025:QPC

[2105] Dingding Cao and Changpin Li. Quenching phenomenon in the Caputo– Hadamard time-fractional Kawarada problem: Analysis and computation. Mathematics and Computers in Simulation, 233:21–38, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542500014X.

Mokni:2025:MPB

[2106] Karima Mokni, Hajar Mouhsine, and Mohamed Ch-Chaoui. Multiparameter bifurcations in a discrete Ricker-type predator–prey model with prey immigration. Mathematics and Computers in Simulation, 233: 39–59, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000205.

Garcia-Chan:2025:NMT

[2107] Néstor García-Chan, Lino J. Alvarez-Vázquez, Aurea Martínez, and Miguel E. Vázquez-Méndez. A nonconservative macroscopic traffic flow model in a twodimensional urban-porous city. Mathematics and Computers in Simulation, 233:60–74, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000163.

Li:2025:IPE

[2108] Xiaoliang Li, Jing Yang, and Ally Quan Zhang. Influence of price elasticity of demand on monopoly games under different returns to scale. Mathematics and Computers in Simulation, 233:75–98, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000175.

Shao:2025:ESV

[2109] Yan Shao, Fuqiang Wu, and Qingyun Wang. Excitability and synchronization of vanadium dioxide memristorinspired neurons. Mathematics and Computers in Simulation, 233:99– 116, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000308.

Acotto:2025:MNE

[2110] Francesca Acotto, Fabiana Camattari, and Ezio Venturino. Mitigating negative effects of EBHSVinfected eastern cottontail invasion in Italy using Z-type control on a four-population system. Mathematics and Computers in Simulation, 233:117–136, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000321.

Bersani:2025:AEB

[2111] Alberto M. Bersani, Alessandro Borri, Giovanna Tomassetti, and Pierluigi Vellucci. Asymptotic expansions beyond the tQSSA for the double phosphorylation mechanism. Mathematics and Computers in Simulation, 233:137–164, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000199.

Ranjbar:2025:SSI

[2112] Hassan Ranjbar. Split-step θ integrator for generalized stochastic Volterra integro-differential equations. Mathematics and Computers in Simulation, 233:165–186, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000291.

Dinku:2025:SAO

[2113] Tarekegn Dinku, Boka Kumsa, Jyotirmoy Rana, and Aiyappan Srinivasan. Stability analysis and optimal control of tumour-immune interaction problem using fractional order derivative. Mathematics and Computers in Simulation, 233:187–207, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475424005123.

Deeb:2025:NIN

[2114] Ahmad Deeb and Denys Dutykh. Numerical integration of Navier–Stokes equations by time series expansion and stabilized FEM. Mathematics and Computers in Simulation, 233: 208–236, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542500031X.

Li:2025:UEB

[2115] Jiyong Li and Minghui Yang. Uniform error bounds of an exponential wave integrator for the Klein–Gordon– Schrödinger system in the nonrelativistic and massless limit regime. Mathematics and Computers in Simulation, 233:237–258, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000357.

Mandal:2025:SMM

[2116] Uttam Kumar Mandal, Biren Karmakar, Sukanya Dutta, and Amiya Das. Symmetry methods and multistructure solutions for a (3 + 1)- dimensional generalized nonlinear evolution equation. Mathematics and Computers in Simulation, 233:259– 275, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000187.

Qin:2025:MEO

[2117] Wenjie Qin, Yi Zhao, and Jin Yang.

Modeling the effects of obesity on a tumor-immune model with combined therapy. Mathematics and Computers in Simulation, 233:276– 295, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000400.

Farouki:2025:APP

[2118] Rida T. Farouki, Marjeta Knez, Vito Vitrih, and Emil Žagar. On L^{2} approximation by planar Pythagoreanhodograph curves. Mathematics and Computers in Simulation, 233:296– 310, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000394.

Zhang:2025:IFC

[2119] Hongkang Zhang and Shao-Lun Huang. Improved fuzzy C-means clustering algorithm based on fuzzy particle swarm optimization for solving data clustering problems. Mathematics and Computers in Simulation, 233: 311–329, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000485.

Zeng:2025:ESD

[2120] Xianfu Zeng and Haiyan Song. Effect of space discretization on the parareal algorithm for advection– diffusion equations. Mathematics and Computers in Simulation, 233: 330–340, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000436.

Das:2025:IIO

[2121] Dipam Das and Debasish Bhattacharjee. An investigation into the impact of odour: a dynamical study of two predators and one prey model, taking into account both integer order and fractional order derivatives. Mathematics and Computers in Simulation, 233:341–368, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000345.

Bohaienko:2025:VMC

[2122] Vsevolod Bohaienko, Fasma Diele, Fabio V. Difonzo, Carmela Marangi, Angela Martiradonna, and Antonello Provenzale. Vertical modeling of carbon sequestration in coastal wetlands using fractional-order derivatives and moisture dynamics. Mathematics and Computers in Simulation, 233:369–388, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000412.

Kuang:2025:SNS

[2123] Daipeng Kuang, Dongdong Gao, and Jianli Li. Stabilization of nonlinear stochastic systems via eventtriggered impulsive control. Mathematics and Computers in Simulation, 233:389–399, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000333.

Shi:2025:MPM

[2124] Longyu Shi, Yunyun Wang, Wenyue Li, and Zhimin Zhang. Multi-period mean–variance portfolio optimization with capital injections. Mathematics and Computers in Simulation, 233:400–412, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000424.

Ren:2025:PFC

[2125] Jianlong Ren, Qiming Liu, and Ping Li. Pattern formation of Cucker– Smale system with nonlinear velocity couplings. Mathematics and Computers in Simulation, 233:413– 432, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000370.

Guzman-Torres:2025:DTA

[2126] J. A. Guzmán-Torres, F. J. Domínguez-Mota, E. M. Alonso Guzmán, G. Tinoco-Guerrero, and J. G. Tinoco-Ruíz. A digital twin approach based method in civil engineering for classification of salt damage in building evaluation. Mathematics and Computers in Simulation, 233:433– 447, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000382.

Zhang:2025:EPE

[2127] Pingrui Zhang, Xiaoyun Jiang, and Junqing Jia. Energy-preserving exponential wave integrator method and the long-time dynamics for the twodimensional space fractional coupled Klein–Gordon–Dirac equation. Mathematics and Computers in Simulation, 233:448–474, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000448.

Guo:2025:DGZ

[2128] Pengfei Guo, Yunong Zhang, and Zheng an Yao. Discrete gradientzeroing neural network algorithm for solving future Sylvester equation aided with left–right four-step rule as well as robot arm inverse kinematics. Mathematics and Computers in Simulation, 233:475–501, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000461.

Kannan:2025:NSM

[2129] Sushmitha Kannan and Vallampati Ramachandra Prasad. Numerical study of mixed convection and thermal enhancement in Williamson ternary nanofluid flow over a nonisothermal wedge using the Keller box method. Mathematics and Computers in Simulation, 233:502– 529, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000527.

Kumar:2025:MMT

[2130] Harsh Kumar and Sourav Rakshit. Multi-material topology optimization using isogeometric method based reaction–diffusion level set techniques. Mathematics and Computers in Simulation, 233:530–552, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S037847542500045X.

Azioune:2025:CCM

[2131] Mourad Azioune and Mohammed-Salah Abdelouahab. Controlling chaos and mixed mode oscillations in a Bertrand duopoly game with homogeneous expectations and quadratic cost functions. Mathematics and Computers in Simulation, 233:553– 566, July 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000515.

Anonymous:2025:NIg

[2132] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 233:567, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001065.

Anonymous:2025:ICEg

[2133] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 233:568, July 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475425001077.

Bulai:2025:NMS

[2134] Iulia Martina Bulai, Maria Carmela De Bonis, and Concetta Laurita. A new MATLAB software for numerical computation of biological observables for metastatic tumor growth. Mathematics and Computers in Simulation, 234(??):31–49, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000503.

Monti:2025:PSO

[2135] Angela Monti, Fasma Diele, Deborah Lacitignola, and Carmela Marangi. Patterns in soil organic carbon dynamics: Integrating microbial activity, chemotaxis and data-driven approaches. Mathematics and Computers in Simulation, 234(??):86– 101, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000552.

Braun:2025:NST

[2136] Elishan Christian Braun, Daniela Mansutti, and Kumbakonam R. Rajagopal. Numerical solution of a twodimensional rock-glacier flow model via the pressure method. Mathematics and Computers in Simulation, 234 (??):102–112, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000369.

Groselj:2025:FCT

[2137] Jan Grošelj and Marjeta Knez. A family of C^{1} Clough–Tocher spline spaces on C^{0} piecewise quadratic domain partitions. Mathematics and Computers in Simulation, 234(??): 368–389, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000813.

Cavoretto:2025:NUK

[2138] R. Cavoretto, A. De Rossi, F. Dell’Accio, A. Sommariva, and M. Vianello. Nonsingularity of unsymmetric Kansa matrices: Random collocation by MultiQuadrics and Inverse Multi-Quadrics. Mathematics and Computers in Simulation, 234(??):390– 395, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000801.

Nikpour:2025:NIG

[2139] Mehdi Nikpour, Mohammad Taghi Kheirabadi, and Ali Nodehi. A novel image gravity transform based on least significant bit in image steganography. Mathematics and Computers in Simulation, 234(??):396– 418, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000849.

Molla:2025:TDD

[2140] Md. Mamun Molla and Md. Mahadul Islam. Three-dimensional D3Q27 multiple-relaxation-time lattice Boltzmann simulation of Herschel–Bulkley viscoelastic fluids in a cubic cavity with top lid driven diagonally. Mathematics and Computers in Simulation, 234 (??):419–437, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000898.

Du:2025:ERA

[2141] Feifei Du, Jun-Guo Lu, and Qing-Hao Zhang. Estimating the region of attraction on fractionalorder complex networks with timevarying delay. Mathematics and Computers in Simulation, 234(??):438– 458, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000746.

Mignot:2025:CPS

[2142] Sarah Mignot and Frank Westerhoff. Contagious popular stories, stock market participation, and boom– bust cycles. Mathematics and Computers in Simulation, 234(??):459– 471, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000886.

Moatimid:2025:IMS

[2143] Galal M. Moatimid and Yasmeen M.

Mohamed. An innovative methodology in scrutinizing the nonlinear instability of two immiscible MHD viscoelastic liquids. Mathematics and Computers in Simulation, 234(??): 472–496, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000916.

Gurbuz:2025:DBB

[2144] Burcu Gürbüz, Aytül Gökçe, Segun I.

Oke, Michael O. Adeniyi, and Mayowa M. Ojo. Dynamical behavior and bifurcation analysis for a theoretical model of dengue fever transmission with incubation period and delayed recovery. Mathematics and Computers in Simulation, 234(??): 497–513, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000783.

Anonymous:2025:NIh

[2145] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 234(??):514, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001284.

Anonymous:2025:ICEh

[2146] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 234(??):515, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475425001296.

Anonymous:2025:EBh

[2147] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 234:ii, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001259.

Bakhvalov:2025:SFV

[2148] P. A. Bakhvalov and M. D. Surnachev. On the stability of finite-volume schemes on non-uniform meshes. Mathematics and Computers in Simulation, 234:1–14, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000539.

Anonymous:2025:PAb

[2149] Anonymous. Pages 1–516 (August 2025). Mathematics and Computers in Simulation, 234:1–516, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Wei:2025:SWP

[2150] Zhijian Wei and Lihui Guo. The singular wave in a pressureless hydrodynamic model. Mathematics and Computers in Simulation, 234: 15–30, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000540.

Hao:2025:ASH

[2151] Yicheng Hao, Yantao Luo, Jianhua Huang, Long Zhang, and Zhidong Teng. Analysis of a stochastic HIV/AIDS model with commercial heterosexual activity and Ornstein– Uhlenbeck process. Mathematics and Computers in Simulation, 234:50– 72, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000564.

Li:2025:DAS

[2152] Chunru Li, Xuesong Yuan, and Yu Gong. Dynamic analysis of a Solow–Swan model with capitalinduced labor migration. Mathematics and Computers in Simulation, 234:73–85, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000576.

Attias:2025:AMF

[2153] Laurent Attias, Vincent Siess, and Stéphane Labbé. An agile modeling framework for population dynamics. Mathematics and Computers in Simulation, 234:113–134, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475425000497.

Verma:2025:PES

[2154] Lalchand Verma, Ramakanta Meher, and Darshak P. Pandya. Parameter estimation study of temporal fractional HIV/AIDS transmission model with fractal dimensions using real data in India. Mathematics and Computers in Simulation, 234:135– 150, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000473.

Pene:2025:AML

[2155] Serigne Daouda Pene, Antoine Picot, Fabrice Gamboa, Nicolas Savy, Christophe Turpin, and Amine Jaafar. Aging modeling and lifetime prediction of a proton exchange membrane fuel cell using an extended Kalman filter. Mathematics and Computers in Simulation, 234:151– 168, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000667.

Du:2025:SPF

[2156] Zhao sheng Du, Jun qing Li, Hao nan Song, Kai zhou Gao, Ying Xu, Jia ke Li, and Zhi xin Zheng. Solving the permutation flow shop scheduling problem with sequence-dependent setup time via iterative greedy algorithm and imitation learning. Mathematics and Computers in Simulation, 234:169–193, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000692.

Gong:2025:TTC

[2157] Junjie Gong, Jian Chen, Dengsheng Cai, Haibo Xie, Wei Wei, and Yu Long. Trajectory tracking considering model uncertainty with interconnection and damping assignment passivity-based control for electrohydraulic servo systems. Mathematics and Computers in Simulation, 234:194–218, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000710.

Liu:2025:MML

[2158] Yang Liu and Zhongfeng Qin. Modified maximum likelihood approach in uncertain regression analysis and application to factors analysis of urban air quality. Mathematics and Computers in Simulation, 234:219– 234, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000680.

Coclite:2025:SMD

[2159] A. Coclite, L. Lopez, and S. F. Pellegrino. A spectral method for dispersive solutions of the nonlocal Sine– Gordon equation. Mathematics and Computers in Simulation, 234:235– 247, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000722.

Wang:2025:SAO

[2160] Leimin Wang, Jian Shen, Xiaofang Hu, Guodong Zhang, and Genping Wu. Stability analysis and optimal control of SEAIQR infectious disease model with nonlinear treatment term based on BA scale-free network. Mathematics and Computers in Simulation, 234:248–261, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000758.

Wang:2025:WPD

[2161] Teng Wang and Jun Zhou. Wellposedness and decay of the energy of the viscoelastic porous elastic system with dual phase-lag model. Mathematics and Computers in Simulation, 234:262–285, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542500076X.

Yi:2025:HOD

[2162] Qian Yi, An Chen, and Hengfei Ding.

High order difference method for fractional convection equation. Mathematics and Computers in Simulation, 234:286–298, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000679.

Chu:2025:DAO

[2163] Huidi Chu, Xinmiao Rong, Liu Yang, and Meng Fan. Dynamical analysis and optimal control strategy of seasonal brucellosis. Mathematics and Computers in Simulation, 234:299– 324, August 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000771.

Han:2025:MIE

[2164] Chang Yong Han and Song-Hwa Kwon. Motion interpolation with Euler–Rodrigues frames on extremal Pythagorean-hodograph curves. Mathematics and Computers in Simulation, 234:325–341, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000734.

Li:2025:SBA

[2165] Shanwei Li and Yimamu Maimaiti.

Stability and bifurcation analysis of a time-order fractional model for water-plants: Implications for vegetation pattern formation. Mathematics and Computers in Simulation, 234:342–358, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000825.

Kim:2025:AAS

[2166] Philsu Kim, Sangbeom Park, Seonghak Kim, and Soyoon Bak. An asymptotic approximation of the solution for nearly tridiagonal quasi-Toeplitz linear systems. Mathematics and Computers in Simulation, 234:359–367, August 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000709.

Anonymous:2025:EBi

[2167] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 235:ii, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001594.

Manzoor:2025:RRB

[2168] Sajida Manzoor, Muhammad Rehan, and Ijaz Ahmed. Robust regulation of blood glucose for Type-1 diabetes using insulin infusion limiter-based feedback control via artificial Pancreas. Mathematics and Computers in Simulation, 235:1–15, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425000941.

Anonymous:2025:PS

[2169] Anonymous. Pages 1–238 (September 2025). Mathematics and Computers in Simulation, 235:1–238, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Yuan:2025:GDC

[2170] Xiaoyue Yuan, Xuebing Zhang, Wenjun Liu, Ali Moussaoui, and Pierre Auger. Global dynamics of a chemotaxis system with toxicity in invasive species. Mathematics and Computers in Simulation, 235:16– 36, September 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000837.

Kafash:2025:CMF

[2171] B. Kafash and Z. Nikooeinejad. A computational method for finding feedback Nash equilibrium solutions (FBNES) in Nonzero-Sum Differential Games (NZSDG) Based on the Variational Iteration Method (VIM). Mathematics and Computers in Simulation, 235: 37–59, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000904.

Hakim:2025:CGQ

[2172] Arief Hakim, A. N. M. Salman, and Khreshna Syuhada. Conditional generalized quantiles as systemic risk measures: Properties, estimation, and application. Mathematics and Computers in Simulation, 235:60–84, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000850.

Wu:2025:MBA

[2173] Xi-Hu Wu and Yi-Tian Gao. Multibreather asymptotics for a nonlinear Schrödinger equation arising in oceanography, plasma physics, Bose–Einstein condensation and fiber optics. Mathematics and Computers in Simulation, 235:85–94, September 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000928.

Kumar:2025:DSM

[2174] Vineesh Kumar, Arvind Patel, and Monu Kumar. Dynamics of solitons and modulation instability in a (2 + 1)-dimensional coupled nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 235: 95–113, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000965.

Ladubec:2025:VAM

[2175] Chris Ladubec and Robert Gracie. Vertically averaged multi-constituent flow simulations of geological CO _{2} sequestration — Stabilized finite element methods and quadratic elements. Mathematics and Computers in Simulation, 235:114–131, September 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425000977.

Barreto:2025:MAM

[2176] Estevão G. C. Barreto, Tânia S. Klein, and Sávio S. V. Vianna. Modelling of the additional motion resistance term in CFD simulations using porosity distributed resistance (PDR). Mathematics and Computers in Simulation, 235: 132–144, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000990.

Costabile:2025:LBA

[2177] Massimo Costabile, Ivar Massabó, Emilio Russo, Alessandro Staino, Rogemar Mamon, and Yixing Zhao. A lattice-based approach for life insurance pricing in a stochastic correlation framework. Mathematics and Computers in Simulation, 235:145– 159, September 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001120.

Kim:2025:SPS

[2178] Yejin Kim, Hyoseon Yang, and Jungho Yoon. Shape-preserving subdivision scheme with the third-order accuracy and C^{2} smoothness. Mathematics and Computers in Simulation, 235: 160–174, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001156.

Ustun:2025:DSB

[2179] Deniz Ustun and Serap Sahinkaya. Design of secure S-Boxes via novel 2D-Zettle chaotic map and ABC algorithm for robust image encryption. Mathematics and Computers in Simulation, 235:175–204, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542500093X.

ElYazidi:2025:SBM

[2180] Youness El Yazidi and Shengda Zeng. A splitting based method for the numerical identification of a nonlinear convection coefficient in elliptic equations. Mathematics and Computers in Simulation, 235:205– 218, September 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000862.

Lin:2025:GSP

[2181] Xinyi Lin, Jian Song, Na Zhao, and Shenquan Liu. Geometric singular perturbation analysis of a threetimescale coupled reduced Hodgkin– Huxley system. Mathematics and Computers in Simulation, 235:219– 236, September 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000035.

Anonymous:2025:NIi

[2182] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 235:237, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001624.

Anonymous:2025:ICEi

[2183] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 235:238, September 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475425001636.

Anonymous:2025:EBj

[2184] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 236:ii, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002046.

Yang:2025:OMR

[2185] Ping Yang, Zhao-Hong Wang, and Yao-Lin Jiang. H_{2} optimal model reduction of linear dynamical systems with quadratic output by the Riemannian BFGS method. Mathematics and Computers in Simulation, 236:1–11, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000953.

Anonymous:2025:PO

[2186] Anonymous. Pages 1–476 (October 2025). Mathematics and Computers in Simulation, 236:1–476, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Yi:2025:ETS

[2187] Chengbo Yi, Jialun Li, Jiayi Cai, Ze You, and Chenglin Jing. Eventtriggered synchronization for multiagent networked systems with timevarying coupling strength and nondifferentiable delays. Mathematics and Computers in Simulation, 236: 12–28, October 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001144.

Yang:2025:PBC

[2188] Zheng Yang, Rui Zhou, HuaiRong Qu, Liang Liu, and QingBin Wu. Prediction of breast cancer using metaheuristicdriven ensemble learning: a novel classification approach. Mathematics and Computers in Simulation, 236:29–51, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000989.

Akyildiz:2025:CDS

[2189] Fahir Talay Akyildiz, Kuppalapalle Vajravelu, Cemil Tunç, and John Abraham. Chebyshev delta shaped and Chebyshev pseudo-spectral methods for solutions of differential equations. Mathematics and Computers in Simulation, 236:52– 69, October 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542500117X.

Polwang:2025:NSW

[2190] Apipoom Polwang, Kanyuta Poochinapan, and Ben Wongsaijai. Numerical simulation of wave flow: Integrating the BBM-KdV equation using compact difference schemes. Mathematics and Computers in Simulation, 236:70–89, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425000874.

Djilali:2025:DSE

[2191] Salih Djilali, Yuming Chen, and Shaofen Zou. A diffusive SIS epidemic model in a heterogeneous environment: Random dispersion vs. nonlocal dispersion. Mathematics and Computers in Simulation, 236: 90–110, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001181.

Li:2025:ELT

[2192] Yi Li, Ning Li, Dandan Xue, and Yi Qin. Efficient and long time stable IMEX schemes for the unsteady dual-porosity-Stokes system. Mathematics and Computers in Simulation, 236:111–134, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001119.

Xie:2025:EVJ

[2193] Jiayi Xie, Zhimin Zhang, and Zhenyu Cui. Efficient valuation of joint life variable annuities with guaranteed minimum death benefits. Mathematics and Computers in Simulation, 236: 135–153, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542500120X.

Martinez-Aparicio:2025:SPS

[2194] Pedro J. Martínez-Aparicio, Pedro Ortiz, and Juan Carlos Trillo. On the stability of the PPH subdivision scheme on nonuniform grids. Mathematics and Computers in Simulation, 236:154–169, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001132.

Lorin:2025:DDF

[2195] Emmanuel Lorin and Howl Nhan. Data-driven fractional algebraic system solver. Mathematics and Computers in Simulation, 236:170– 182, October 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425000795.

Menon:2025:ICD

[2196] Sidharth Menon and Sangeeta Kumari. Impact of cross-diffusion and Allee effect on modified Leslie– Gower model. Mathematics and Computers in Simulation, 236:183– 199, October 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425001168.

Yoshioka:2025:GRD

[2197] Hidekazu Yoshioka. Generalized replicator dynamics based on mean-field pairwise comparison dynamic. Mathematics and Computers in Simulation, 236:200–220, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001405.

Mondal:2025:OPB

[2198] Rajan Mondal, Amalesh Kumar Manna, Md Akhtar, and Asoke Kumar Bhunia. Optimization of a price break policy and advertisement effort based non-instantaneous deteriorating inventory problem with partial backlogged via metaheuristic algorithms. Mathematics and Computers in Simulation, 236:221–247, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001375.

Liu:2025:SBC

[2199] Jiayi Liu, Ruihong Li, and Dongmei Huang. Stability, bifurcation and characteristics of chaos in a new commensurate and incommensurate fractionalorder ecological system. Mathematics and Computers in Simulation, 236: 248–269, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001430.

Lai:2025:CDE

[2200] Qiang Lai, Chongkun Zhu, Minghong Qin, and Zhiqiang Wan. Complex dynamics and encryption application of a 3D dual-memristor oscillatory hyperchaotic map. Mathematics and Computers in Simulation, 236:270–283, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001363.

Qin:2025:DAL

[2201] Xiaofei Qin, Hongjun Li, and Changnv Zeng. Dynamic analysis of layered elastic medium by TD-BEM using the iteration technique in the whole time domain. Mathematics and Computers in Simulation, 236: 284–304, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542500148X.

Nazer:2025:HPP

[2202] Afshin Nazer, Patrizio Manganiello, and Olindo Isabella. Hierarchical PV to PV series differential power processing for photovoltaic applications: Power flow and converter rating analysis. Mathematics and Computers in Simulation, 236:305– 319, October 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001399.

Chen:2025:NSD

[2203] Ruyun Chen, Yu Li, and Yongxiong Zhou. Numerical steepest descent method for computing oscillatory-type Bessel integral transforms. Mathematics and Computers in Simulation, 236:320–333, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001466.

Zhai:2025:CIO

[2204] Shidong Zhai, Zhenmei Zhang, and Jun Ma. Convergence of individual opinions during a bi-virus epidemic and its impact. Mathematics and Computers in Simulation, 236: 334–353, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542500134X.

Goswami:2025:DTE

[2205] Anindya Goswami and Kuldip Singh Patel. Domain truncation error analysis for a multidimensional system of PDEs of option prices. Mathematics and Computers in Simulation, 236:354–378, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001211.

Song:2025:GDO

[2206] Caihong Song and Zigen Song. Global dynamics and optimal control strategies for the brucellosis model with environmental saturation infection: an assessment of severely affected provinces in China. Mathematics and Computers in Simulation, 236: 379–410, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001387.

Marin-Sanchez:2025:VEC

[2207] Freddy H. Marín-Sánchez, Alejandro Pinilla Barrera, Cristhian Montoya Zambrano, and Santiago Medina Hurtado. Valuation of European call options for the Scott’s stochastic volatility model: an explicit finite difference scheme. Mathematics and Computers in Simulation, 236: 411–425, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001193.

Maes:2025:NAR

[2208] Frederick Maes and Karel Van Bockstal. Numerical algorithms for the reconstruction of space-dependent sources in thermoelasticity. Mathematics and Computers in Simulation, 236:426–454, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001351.

Ocampo:2025:DSM

[2209] Lanndon Ocampo, Dharyll Prince Abellana, Julius Selle, Rhoda Namoco, Fatima Maturan, John Kevin Padro, and Samantha Shane Evangelista. Dynamical system modelling for picture fuzzy multi-criteria public opinion on transportation infrastructure. Mathematics and Computers in Simulation, 236:455–474, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001545.

Anonymous:2025:NIj

[2210] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 236:475, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002071.

Anonymous:2025:ICEj

[2211] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 236:476, October 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475425002083.

Anonymous:2025:EBk

[2212] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 237:ii, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002290.

Zaabouli:2025:NTF

[2213] Z. Zaabouli, L. Afraites, and A. Laghrib.

A novel time-fractional decomposition model for image denoising integrating Caputo derivative. Mathematics and Computers in Simulation, 237:1–17, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001442.

Anonymous:2025:PN

[2214] Anonymous. Pages 1–392 (November 2025). Mathematics and Computers in Simulation, 237:1–392, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Hussain:2025:CMB

[2215] Manzoor Hussain and Abdul Ghafoor. On the computational modeling of the behavior of a three-dimensional Brusselator system using a localized meshless method. Mathematics and Computers in Simulation, 237: 18–41, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542500151X.

Singh:2025:PRM

[2216] Vijay Pratap Singh, Madhu Jain, and Richa Sharma. N-policy for redundant machining system with double retrial orbits using soft computing techniques. Mathematics and Computers in Simulation, 237:42– 69, November 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001557.

Bu:2025:LCA

[2217] Weiping Bu and Xin Zheng. Local convergence analysis of L1/finite element scheme for a constant delay reaction-subdiffusion equation with uniform time mesh. Mathematics and Computers in Simulation, 237: 70–85, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001454.

Sans:2025:HBG

[2218] Eysan Sans, Melisa Akdemir, Ayse Tiryakioglu, Ayse Peker-Dobie, and Cihangir Ozemir. Hopf bifurcation in a generalized Goodwin model with delay. Mathematics and Computers in Simulation, 237:86–106, November 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001429.

Liang:2025:EKO

[2219] Ziyuan Liang and Zhenlei Wang. An enhanced Kepler optimization algorithm with global attraction model and dynamic neighborhood search for global optimization and engineering problems. Mathematics and Computers in Simulation, 237:107– 144, November 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001338.

Zheng:2025:NBS

[2220] Jihui Zheng, Zhenlan Pan, and Jing An. A novel BDF-spectral method and its error analysis for Cahn–Hilliard equation in polar geometry. Mathematics and Computers in Simulation, 237:145–166, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001533.

Yu:2025:CLE

[2221] Weiyang Yu and Haitao Zheng.

Composite likelihood expectation– maximization algorithm for the firstorder multivariate integer-valued autoregressive model with multivariate mixture distributions. Mathematics and Computers in Simulation, 237: 167–187, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001478.

Zhang:2025:NMG

[2222] Linfeng Zhang, Guorong Wang, Lin Zhong, Qingping Li, Qiang Fu, Mengdi Wu, and Lei Chen. Nonlocal modeling for gas-fluid seepage coupled with heat transfer in gas hydrate by using peridynamic differential operator. Mathematics and Computers in Simulation, 237:188–212, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001508.

Hoque:2025:MID

[2223] Ashabul Hoque, Razia Sultana, Hamidul Islam, and Abdul Malek. A modeling investigation of the disease severity driven by COVID-19-induced diabetic patients. Mathematics and Computers in Simulation, 237: 213–230, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001417.

Huang:2025:EPI

[2224] Xujia Huang, Fajie Wang, Benrong Zhang, and Hanqing Liu. Enriched physics-informed neural networks for dynamic Poisson–Nernst– Planck systems. Mathematics and Computers in Simulation, 237:231– 246, November 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001752.

Fan:2025:HNI

[2225] Yamin Fan, Ximei Liu, and Meihang Li. Hierarchical Newton iterative identification methods for a class of input multi-piecewise Hammerstein models with autoregressive noise. Mathematics and Computers in Simulation, 237:247–262, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001491.

Aslam:2025:OEF

[2226] Ch Muhammad Afaq Aslam, Ikram Ullah, and Muhammad Ishaq. An optimal and efficient framework for the construction of nonlinear components based on a polynomial ring. Mathematics and Computers in Simulation, 237:263–280, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001521.

Zhu:2025:PFL

[2227] Yuqi Zhu, Shiting Zhang, Yang Hu, Xiaoqiang Yue, Shi Shu, Qiang He, and Decai Li. Phase field-lattice Boltzmann model for axisymmetric twophase ferrofluid flows. Mathematics and Computers in Simulation, 237: 281–315, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001685.

Ye:2025:IHA

[2228] Guodong Ye, Shoukang Liu, Xiuchun Xiao, and Xiaoling Hunag. Image hiding algorithm based on local binary pattern and compressive sensing. Mathematics and Computers in Simulation, 237:316–334, November 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425001879.

Gine:2025:LCH

[2229] Jaume Giné and Joan Torregrosa. Limit cycles of homogeneous polynomial Kukles differential systems. Mathematics and Computers in Simulation, 237:335–343, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001831.

Lopez:2025:ICL

[2230] Miguel López, Sergio Valero, and Carolina Senabre. Impact of COVID-19 lockdown in short-term load forecasting. Mathematics and Computers in Simulation, 237:344–354, November 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425001739.

Yuan:2025:VRV

[2231] Cong Yuan, Jean-Philippe Martin, Serge Pierfederici, Matheepot Phattanasak, Farid Meibody-Tabar, and Shengzhao Pang. Voltage reference varying-based adaptive IDA-PBC design and stability analysis for DC microgrids. Mathematics and Computers in Simulation, 237: 355–372, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001673.

Lekhel:2025:GRL

[2232] Cheikh Elekbir Sidi Lekhel, Rita Mbayed, Oleksandr Velihorskyi, Oleksandr Husev, and Eric Monmasson. Generic residential load profile generator based on weather data and occupancy. Mathematics and Computers in Simulation, 237:373– 389, November 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542500182X.

Anonymous:2025:NIk

[2233] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 237:390, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002320.

Anonymous:2025:ICEk

[2234] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 237:391, November 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475425002332.

Anonymous:2025:EBl

[2235] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 238:ii, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425003179.

Stollenwerk:2025:ERM

[2236] Nico Stollenwerk, Luís Mateus, Vanessa Steindorf, Bruno V. Guerrero, Rubén Blasco-Aguado, Aitor Cevidanes, Joseba Bidaurrazaga Van-Dierdonck, and Maíra Aguiar. Evaluating the risk of mosquito-borne diseases in non-endemic regions: a dynamic modeling approach. Mathematics and Computers in Simulation, 238:1– 24, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425001648.

Anonymous:2025:PD

[2237] Anonymous. Pages 1–638 (December 2025). Mathematics and Computers in Simulation, 238:1–638, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liu:2025:MCP

[2238] Xinwu Liu, Xiang Yu, and Jiangli Liang. Minimax concave penaltybased TGV model for Poissonian image restoration. Mathematics and Computers in Simulation, 238:25– 44, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425001843.

Vigo:2025:CAS

[2239] Daniel G. Alfaro Vigo, Daniele C. R.

Gomes, Bruno A. do Carmo, and Mauro A. Rincon. Convergence analysis of a Schrödinger problem with moving boundary. Mathematics and Computers in Simulation, 238:45– 64, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001806.

Fernandes:2025:NAK

[2240] A. Fernandes, J. A. Ferreira, and L. Pinto. Numerical analysis of a Keller–Segel-Flow model for tumor cell migration. Mathematics and Computers in Simulation, 238: 65–81, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001727.

Oluwasakin:2025:LTV

[2241] Ebenezer O. Oluwasakin, Abdul Q. M. Khaliq, and Khaled M. Furati. Learning time-varying parameters of stiff dynamical systems using physics-informed transfer neural network. Mathematics and Computers in Simulation, 238:82–102, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001764.

Kubra:2025:NCF

[2242] Khadija Tul Kubra, Rooh Ali, Samra Gulshan, and Hafiza Hafsa Muzaffar. A novel comparative fractionalorder modeling of omicron dynamics: Vaccination impact and control strategies in the USA. Mathematics and Computers in Simulation, 238: 103–135, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542500179X.

Kumawat:2025:NAQ

[2243] Sangeeta Kumawat, Sanjay Bhatter, Bhamini Bhatia, Sunil Dutt Purohit, Haci Mehmet Baskonus, and Daya Lal Suthar. Novel application of q-HAGTM to analyze Hilfer fractional differential equations in diabetic dynamics. Mathematics and Computers in Simulation, 238: 136–149, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001661.

Timofeyev:2025:AML

[2244] Ilya Timofeyev, Alexey Schwarzmann, and Dmitri Kuzmin. Application of machine learning and convex limiting to subgrid flux modeling in the shallow-water equations. Mathematics and Computers in Simulation, 238: 163–178, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001703.

Zheng:2025:PIA

[2245] Jinyang Zheng and Xiufang Feng. A parallel iterative algorithm for solving the two-dimensional Helmholtz equation with discontinuous coefficients using a high-order compact finite difference scheme. Mathematics and Computers in Simulation, 238: 179–188, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001697.

Miao:2025:TPN

[2246] Peng Miao, Shuai Li, and Chenghang Li. Tracking problem of nonparticle morphology based on fixedtime ZNN model. Mathematics and Computers in Simulation, 238:189– 200, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001776.

Jiang:2025:RAC

[2247] Defeng Jiang and Shaohui Yan. Reversible adaptive chaotic multi-image privacy protection scheme for efficient batch encryption of various images. Mathematics and Computers in Simulation, 238:201–221, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002009.

Hanini:2025:SIM

[2248] Amine Hanini and Abdelaziz Beljadid. A semi-implicit method for modeling salinity and temperature in variable-density flow. Mathematics and Computers in Simulation, 238: 222–239, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001715.

Manosa:2025:KHK

[2249] Víctor Mañosa and Chara Pantazi. Kahan–Hirota–Kimura maps preserving original cubic Hamiltonians. Mathematics and Computers in Simulation, 238:240–254, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001855.

Marwan:2025:PCH

[2250] Muhammad Marwan, Ning Wang, Dur e Zehra Baig, and Feng Li. Practical computation of higher codimension Bogdanov–Takens bifurcations in energy supply–demand system. Mathematics and Computers in Simulation, 238:255–268, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001880.

Serry:2025:SEB

[2251] Mohamed A. Serry, Sean D. Peterson, and Jun Liu. Static Euler– Bernoulli beams with contact forces: Existence, uniqueness, and numerical solutions. Mathematics and Computers in Simulation, 238:269– 279, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425002204.

Horng:2025:ASO

[2252] Shih-Cheng Horng and Shieh-Shing Lin. Adapted skill optimization algorithm to solve constrained discrete stochastic optimization problems. Mathematics and Computers in Simulation, 238:280–295, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002472.

Sun:2025:PTA

[2253] Tao Sun, Michael K. Ng, and Hai-Wei Sun. Pre-tensorial alternative direction implicit method for nonlocal diffusion equations. Mathematics and Computers in Simulation, 238: 306–321, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542500240X.

He:2025:NCV

[2254] Dongping He, Ziheng Duan, Huidong Xu, Xianglan Kong, Yuanming Liu, and Tao Wang. Nonlinear coupled vibration and control of roll system for twenty-high rolling mill. Mathematics and Computers in Simulation, 238:322–341, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002484.

Pei:2025:HEC

[2255] Yansong Pei and Li Li. The hydra effect in a cannibalistic reaction– diffusion predator–prey model: Influence of space-time white noise on population dynamics. Mathematics and Computers in Simulation, 238: 342–360, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542500237X.

Samy:2025:RMS

[2256] Stephen Arockia Samy, Jae Hoon Jeong, Xiaoshan Bai, and Young Hoon Joo. Reliable memory sampled-data control for leader-following consensus of nonlinear multi-agent systems and its application. Mathematics and Computers in Simulation, 238: 361–382, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001867.

Diab:2025:ZHB

[2257] Zouhair Diab, Juan L. G. Guirao, and Jaume Llibre. Zero-Hopf bifurcation of a 5D hyperchaotic quadratic polynomial differential systems. Mathematics and Computers in Simulation, 238:383–387, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002538.

Huang:2025:ADE

[2258] W. Huang, J. J. Yang, and P. H. Wen. An automatic differentiationenhanced meshfree finite block method for nonlinear problems. Mathematics and Computers in Simulation, 238: 388–402, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002630.

Zhou:2025:DCF

[2259] Junwei Zhou, Dianguang Ma, Feng Ma, Yiqun Sun, Dandan Shen, and Yu Duan. Does the constant friction coefficient fit open-channel friction calculation? Mathematics and Computers in Simulation, 238:403– 426, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002563.

Zhao:2025:SPO

[2260] Hongyong Zhao and Huixia Li. Spatiotemporal patterns and optimal control in a diffusive Alzheimer’s disease model with pharmacokinetics. Mathematics and Computers in Simulation, 238:427–456, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002228.

Kulwant:2025:OPP

[2261] Aditya Kulwant, Madhu Jain, Parmeet Kaur Chahal, Rakesh Kumar Meena, and Arvind Kumar Singh. Optimization and performance prediction of fault-tolerant repairable system in fuzzy environment with reboot, feedback, and retrial. Mathematics and Computers in Simulation, 238:457– 478, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002411.

Aghazadeh:2025:ACB

[2262] Arezu Aghazadeh and Mehrdad Lakestani. Application of cubic Bsplines for second order fractional Sturm–Liouville problems. Mathematics and Computers in Simulation, 238: 479–496, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425002587.

Wu:2025:TWD

[2263] Xin Wu, Zeyu Ding, Rong Yuan, and Fangyuan Chen. Traveling waves in delayed lattice systems: Effects of case report delays and pathogen detection. Mathematics and Computers in Simulation, 238: 516–535, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425002551.

Zhou:2025:GHW

[2264] Yue Zhou and Hang Xu. A generalized Haar wavelet-homotopy method for solving steady and unsteady differential and partial differential equations. Mathematics and Computers in Simulation, 238:536– 554, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425002642.

Roque:2025:RSO

[2265] E. Roque and S. M. Torba. Representation of solutions of the one-dimensional Dirac equation in terms of Neumann series of Bessel functions. Mathematics and Computers in Simulation, 238: 568–584, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002617.

Fathi:2025:MMH

[2266] Ahmed Fathi, Maria Rosaria Mattei, Fernando G. Fermoso, and Luigi Frunzo. A multiscale model of heavy metal biosorption in porous media by one-dimensional biofilms. Mathematics and Computers in Simulation, 238: 585–603, December 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001892.

Zhao:2025:BAE

[2267] Shijie Zhao, Fanshuai Meng, Liang Cai, and Ronghua Yang. Boomerang aerodynamic ellipse optimizer: a human game-inspired optimization technique for numerical optimization and multilevel thresholding image segmentation. Mathematics and Computers in Simulation, 238:604–636, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542500271X.

Padron:2025:ICE

[2268] Miguel A. Padrón, Agustín Trujillo-Pino, and Jose Pablo Suárez. IMACS calendar of events. Mathematics and Computers in Simulation, 238: 638, December 2025. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425003222.

Choi:2026:DLS

[2269] Jae-Hoon Choi, Byung-Chai Lee, and Gi-Dong Sim. Discrete least square method for nonmatching mesh problems. Mathematics and Computers in Simulation, 239:1– 25, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425002113.

Din:2026:NDB

[2270] Qamar Din and Iqra Fareed. Nonlinear dynamics and bifurcation analysis of a discrete plant-herbivore system: Empirical validation and ecological implications. Mathematics and Computers in Simulation, 239: 44–82, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425002186.

Wang:2026:TTS

[2271] Lihong Wang, Tingting Chen, and Jingsong He. Three types of solitons propagating in saturable nonlinear media. Mathematics and Computers in Simulation, 239:83– 95, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425001934.

Bi:2026:OEE

[2272] Hui Bi, Chunmei Hu, Feilong Zhao, and Fengyu Fu. Optimal error estimates of the ultra weak local discontinuous Galerkin method for nonlinear sixthorder boundary value problems. Mathematics and Computers in Simulation, 239:96–114, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001788.

Kumar:2026:SSM

[2273] Rakesh Kumar, Kapil Kant, and B. V. Rathish Kumar. Superconvergent spectral methods for system of modified Volterra Integral Equations. Mathematics and Computers in Simulation, 239:115–134, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001946.

Liu:2026:RTI

[2274] Qingqing Liu and Fangfang Dou.

Reconstructing the time-independent volatility in Black–Scholes equation. Mathematics and Computers in Simulation, 239:135–154, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001818.

Chun:2026:TMM

[2275] Sehun Chun and Jae-Hun Jung. Time map method for spatiotemporal simulation of cardiac and neural signal propagation. Mathematics and Computers in Simulation, 239:155– 171, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001740.

Li:2026:NFD

[2276] Mengmeng Li, Guang an Zou, and Bo Wang. A new fully-decoupled energy-stable BDF2-FEM scheme for the electro-hydrodynamic equations. Mathematics and Computers in Simulation, 239:172–191, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001922.

Jia:2026:SRC

[2277] Chengwang Jia, Hongwei Zhang, Guang hua Gao, and Tongke Wang. Singularity-removing Chebyshev collocation methods for nonlinear fractional differential equations with blow-up. Mathematics and Computers in Simulation, 239:192–210, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425001971.

Marzban:2026:ICO

[2278] Hamid Reza Marzban and Mahtab Manochehri Naeini. An innovative class of orthogonal functions based on the fractional Vieta–Fibonacci functions and its utilization in optimal control of piecewise constant order systems consisting of delay. Mathematics and Computers in Simulation, 239:223– 244, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002125.

Barker:2026:TWD

[2279] William Barker. Traveling waves for a diffusive SIR epidemic model with delay in the diffusion term. Mathematics and Computers in Simulation, 239:245–262, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542500165X.

Chen:2026:NMI

[2280] Shi-Shun Chen and Xiao-Yang Li. A new moment-independent uncertainty importance measure based on cumulative residual entropy for developing uncertainty reduction strategies. Mathematics and Computers in Simulation, 239:263–281, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002253.

Han:2026:FTI

[2281] Luxia Han, Runzi Luo, Panpan Zhang, and Xianzhou Liu. Fixed-time intermittent control of a class of chaotic systems via a single controller. Mathematics and Computers in Simulation, 239:282–297, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001983.

S:2026:PES

[2282] Rubasri S., Kathiresan S., Ardak Kashkynbayev, and Rakkiyappan R. Pinning exponential synchronization of coupled neural networks with delay under stochastic deception attacks using the Markov switched control. Mathematics and Computers in Simulation, 239:298–317, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425002137.

Zhao:2026:MSP

[2283] Yafei Zhao, Hui Wu, Michael Y. Li, and Jie Lou. Model selection and parameter identification analysis in epidemiological models. Mathematics and Computers in Simulation, 239: 318–339, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425001958.

Liu:2026:CIE

[2284] Xingbin Liu, Shuyi Zheng, and Jing Yang. Color image encryption scheme based on a novel 2D-CLCM chaotic system and RNA encoding. Mathematics and Computers in Simulation, 239:340–360, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425002381.

Cai:2026:NLS

[2285] Min Cai and Weiwei Tong. Nonuniform L1/spectral element algorithm for the time fractional diffusion equation. Mathematics and Computers in Simulation, 239:361–375, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002198.

Pradhan:2026:NFS

[2286] Dickcha Pradhan, Khalid K. Ali, Seydi Battal Gazi Karakoc, and Asit Saha. New families of soliton solutions and dynamics of nonlinear traveling waves for the Whitham– Broer–Kaup equation. Mathematics and Computers in Simulation, 239: 376–390, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425001995.

Ye:2026:USH

[2287] Changlun Ye, Hai Bi, and Xianbing Luo. An unconditionally stable hybrid discontinuous Galerkin scheme for a Cahn–Hilliard phase-field model of two-phase incompressible flow with variable densities. Mathematics and Computers in Simulation, 239:403– 419, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002216.

Lee:2026:NTD

[2288] K. C. Lee, I. Hashim, M. N. Mohd Aris, and N. Senu. A novel two-derivative multistep collocation method with fitting-techniques with application to Duffing problem. Mathematics and Computers in Simulation, 239:420– 441, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475425002174.

Dhayal:2026:TCR

[2289] Rajesh Dhayal and Avadhesh Kumar. Total controllability results for impulsive fractional integro-differential systems of order 1 < ρ ≤ 2. Mathematics and Computers in Simulation, 239:442–457, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847542500196X.

Yu:2026:ACF

[2290] Qiang Yu, Junfeng Xiao, and Kaiqi Zhao. On accurate closed-form Coiflet solutions of nonlinear Korteweg–de Vries–Burgers equation for shallow water wave by a time-integration wavelet homotopy method. Mathematics and Computers in Simulation, 239:458–482, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425002162.

Liu:2026:DAS

[2291] Maoxing Liu, Yu Peng, Xuejie Ren, and Juping Zhang. Dynamical analysis of simplicial epidemic model with demographic and aware effect. Mathematics and Computers in Simulation, 239:483–504, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425002435.

Peng:2026:UEE

[2292] Shanshan Peng and Yanping Chen.

Unconditional error estimate of a linearized variable-time-step BDF2 nonconforming virtual element method for nonlinear coupled predator–prey equations. Mathematics and Computers in Simulation, 239:505–524, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847542500223X.

Yang:2026:MIA

[2293] Yafei Yang, Xin Zhao, and Meng Fan. Modeling the impact of Allee effect, herd behavior and harvesting on kelp beds: a cross-diffusion-ODE dynamical approach. Mathematics and Computers in Simulation, 239:525–554, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002460.

Jiang:2026:NST

[2294] Zhicheng Jiang, Wenjing Li, and Bo Zeng. A novel structure timedelay multivariable discrete grey prediction model and its application in input-output analysis. Mathematics and Computers in Simulation, 239: 572–589, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847542500254X.

Xu:2026:UBS

[2295] Liguang Xu, Hongxiao Hu, and Danhua He. Uniform boundedness and stability of fractional state-dependent delayed systems and applications to complex neural networks. Mathematics and Computers in Simulation, 239:599–615, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002496.

Suksai:2026:AAS

[2296] Sayaphat Suksai, Suthep Suantai, Somkiat Phornphisutthimas, and Chunchom Salikupata. Analyzing atmospheric stability and pollution dispersion through impulsive models. Mathematics and Computers in Simulation, 239:616– 628, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002745.

Wang:2026:MDS

[2297] Chuan Wang, Zhongqing Wang, and Chao Zhang. A multi-domain spectral collocation method for PDEs in curved domains with holes. Mathematics and Computers in Simulation, 239:629–645, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002836.

Yang:2026:DST

[2298] Yadong Yang, Jianye Gong, and Qikun Shen. Dynamic self-triggered fixed-time consensus for stochastic multiagent systems with unknown function powers under hybrid cyber attacks. Mathematics and Computers in Simulation, 239:646– 660, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002903.

Wang:2026:MRH

[2299] Yun Wang, Yicong Bian, and Hu Shao.

A modified Riemannian hybrid conjugate gradient method for nonconvex optimization problems. Mathematics and Computers in Simulation, 239:679–695, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475425002927.

Hoegele:2026:CPR

[2300] Wolfgang Hoegele. Combinatorial potential of random equations with mixture models: Modeling and simulation. Mathematics and Computers in Simulation, 239:696– 715, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475425002988.

Iqbal:2026:NDG

[2301] Muhammad Iqbal, Muhammad Bilal Riaz, Muhammad Aziz ur Rehman, and Syeda Sarwat Kazmi. Nonlinear dynamics of the generalized Hirota– Satsuma–Ito equation: Analytical solutions and graphical insights. Mathematics and Computers in Simulation, 239:729–744, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425002794.

Saqib:2026:AHC

[2302] Sana Ullah Saqib, Yin-Tzer Shih, Muhammad Wajahat Anjum, and Muhammad Shoaib. Advanced heuristic computing with Gudermannian neural networks for mathematical modeling of divorced dynamics in social networks. Mathematics and Computers in Simulation, 239:745– 765, January 2026. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847542500312X.

Amor:2026:AMR

[2303] Hanen Amor and Fayssal Benkhaldoun. Adaptive mesh refinement for flow and transport problem in heterogeneous porous media. Mathematics and Computers in Simulation, 239:766–789, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475425002514.

Anonymous:2026:PJ

[2304] Anonymous. In progress (January 2026). Mathematics and Computers in Simulation, 239:??, January 2026. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).